A Complete Bibliography of Publications in Linear Algebra and its Applications: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

07 April 2025 Version 1.51

Title word cross-reference

(0, 1) [BBD^{+}08, BD07, KTR08, SS06b, Živ06, LB04, RS03a, Sbu05, WJL02]. (1, 2, 1) [Vid08]. (2, 2) [LA05]. (2r − 1) [ML02]. (A + tB)^{m} [JLMC05]. (A, B) [PP02, OdAJ02]. (C, A) [CP07b]. (m, k) [WGL09]. (n − 5) [BS01]. (p, k) [TUC04]. (±1) [CHKS05]. (q − 1)^{13} [VLAOVL08]. (R, S) [Tre04e]. (R_{1}, R_{2}, , R_{N}) [DV04]. (SL_{5} × GL_{4}, Λ^{2}(C^{5}) ⊗ C^{4}) [AFK02]. (w) [AGP08]. ∗ [MT04, ZX05a]. + [vdW01]. {−1, 0, 1} [AAGK06]. −2 [BS04a, Cve02, SS08a]. 0 [FJ03, MW03b]. {0, 1} [BX05]. 0, ±1 [Kar07]. 1 [CGH04, HC03b, JS07a, MW03b, Sta07, Ste03b]. 2 [AY03, Arp05, Aud09b, BFPW02, CN02b, CLY08, DoMP09, Eld02b, FVP06, FK07, KSH05, MQ03, Omi08a, RM01, SG08, SMS09, XL04, Zha04c, ZFW06, vdH08a]. 2 + 5 [GOTR07]. 2n [DHH^{+}07]. 2n − 2 [Bir04]. 2 × 2 [BV02, BM05, Cho04, CI08, Dom00, Dru02, ES02, KN04, KZ03, LS07a, LSD06, LW07, OS01, Wat06]. 2 × N [Aud08]. 2 × q [Val02b]. 3 [pBhSzZ09, CN09, LRWW07, NBdP08b, PBR06, SLP01, Siv09, SMS09, TBK09, Yag01]. 322 [WHAB08]. 3 × 3 [BdPT08, BC00, Cal08, Hil07a, Mat04, MS04d, NBdP05, RS05a]. 3 × n [SH08]. 4 [Aga02a, BDD^{+}01, OB01, Zha04c]. 4.5 [WHBM09b]. 4 × 4 [OPS07]. 5 [LT03b]. 5 × 5 [BAD09, DA10, Kob05, LM04b, BX04]. 6 [AGKM09, CHP03]. 600 [HD03]. 6 × 6 [BN08]. 8 [BDD^{+}01]. [0, 1] [DS02]. [0, ∞) [DS02]. [A^{T} XA, B^{T} XB] = [C, D] [pLLfY06]. [f]([x]) = [A][x] + [b] [MW03a]. ^{(13)} [WWJ09]. ^{∗} [Lav03b]. ^{⋆} [HS06a]. _{q} [ZZ05].

T (AB) =_{T} (BA) [CCH02b]. A [CC00, – DJ07, JT04a, JS07a, MW08a, NT06a, RA08b]. A(R, S) [CHK^{+}01]. A + B [BK09a]. A + iB [BK09a]. A − XY ∗ [Sv05]. A = p(X) [Dra07]. A^{k} = J − I [WL03]. A^{m} [WJL02]. A^{⋆} [NT06a]. A_{1}X_{1}B_{1} + A_{2}X_{2}B_{2} + ⋯ + A_{l}X_{l}B_{l} = C [hPyHZ07]. A_{i}XB_{j} = C_{ij}, (i, j) ⊂ Γ [vdW03]. A_{n} [Zhu08]. A_{T,S}^{(2)} [CC00]. AA^{#} ≥ 0 [JT04a]. AA^{⋆} − A^{⋆}A [NT06a]. A log A [Fur03]. α [Mat05c, Wan06, Tre09b]. αX^{2}AY^{2} + βXAY + A = 0 [Ili08]. A → |A| [CS03b]. AX − XF = BY [ZD05]. AX = B [MHZ05, yPyH03, ZHZ08]. AX = C [Xu08, DK08a]. AXA^{∗} = B [LTT09, Gro00b, ZyC03]. AXY B = 0 [Dob01]. B(H) [Ji03a, Ji03b, DHB07, Pet05, Wu09, YZ09]. B(X) [AH02]. C [DHK06, CC01, CN01]. C[x] [SS04a]. C[Y] [RSS00]. C^{∗} [AR06b, CXZL02, FYY09, HJM09, Li08d, O’u02, PR07, Sha09, Sud02, Sud04, XS08, Xu09]. C^{1} [HMR09]. C^{n} [HvP05, BP06b]. C^{r} [PPF02]. C_{0} [WGT09]. C_{00} [Kar09]. C_{4} [Nik08a, Nik09]. C_{5} [CS05]. C_{n}^{2} [HWC06]. C_{P} [CZ07]. H_{∞} [PGRO06]. M [JO05]. D [ÁFMNW03, ÁFMNW06, FGL^{+}00, LY05, PWZ05, AH09b, Guo07b]. D_{k} [JSU03]. D_{n} [Zhu08]. det(A(G)) = deg(G) [Doo02]. DH(2n − 1, 4) [De 08]. E [Udd09, Fie05b]. e^{A+B} [BS07]. e^{A} = e^{B} [Sch03]. e^{A}e^{B} [BS07]. E_{7} [KKK07]. F [Dun04, Yah06, EdRM04]. f(A) + f(B) [BU07]. f(A + B) [BU07]. F^{N} [Sal09]. F_{2} [PM08]. F_{4} [Dra08c]. FC_{2} [SSSG06]. G [FIL07, Lav03a, Nie04b, TH01, GWW06, WG02a, WG02b, WJL02, MMT04]. G(c) [NT01]. G_{2} [BS02b, BDE03, DM06]. G_{2}(2) [KKK07]. Γ [LL08a, BBMX07]. GF (2) [Ose08b]. GL(n, q) [VLAOVL08]. GL_{n}(F) [NY00]. H [AH08, And04c, BOP08, BCPT05, BPS04, BCGM08, CT04, GH03, Had04, HSL05, KY00, Kol09, LS04b, LH04, Xia03, XZ06, YLX09, yZXtL07]. H^{2} [Tre04f, Bör07]. H^{∞} [Dam02]. H^{n} [Dam04]. h_{1}(x_{1}) + ⋯ + h_{n}(x_{n}) = y [Cai03]. H_{2} [ABR02, PS03, GM09a]. h_{i} [Cai03]. H_{∞} [BBMX07, FS02]. HyperlatT = Hyperlatf(T) [Prǎ02]. I [FS06b]. I × 2 × 2 [tBSR04]. I × 3 × 3 [tBSR04]. J [AADL03, ADL04, AD08, BNdP^{+}05, Fre08, FKRS09, FKR09, LPR04, MH06, NBdP05, NBdP08a, Rod09, Str03b, HQ08, LLM03, LL08b, LL03, WL07]. J^{α} [DM02]. J_{ℓ} [ADL06]. J_{n} [WJL02]. J_{q} [FKM02]. j_{qq} [FKM02]. k [BB03a, BT00, BK03, BG09a, BT05b, BM04b, BG06, Che07a, De 09c, DWD07, GS09b, GL08, Guo05d, Guo06b, GW09c, Hua09, KN03a, KR03, LT07a, Lee00, LK03, LZS09, SC08, SBS02, Stu03, Tre08, WY08, WXH05, YW07, YL06, ZgC06, ZWY00]. K(T) [LT05]. K_{2,l} [SS07c]. K_{3} × C_{n} [HLW08]. K_{m} × P_{n} [LPW08]. K_{n} [Doo02]. L [BC04b, Ish04, Sat06]. L(H) [CS03b]. L^{2} [Di 03]. L_{p} [Doe07, Pro08]. L_{r}^{s} [Han03]. λ [AMS05, Oku06]. Λ_{n}^{k} [Wan03b]. LR [PVV08]. M [Ano03-44, AT09, BCEG09, CNX05b, Che04b, Che07b, Che07c, CH07, DPZ06, ES08, GN03, HNT03, HTK07, Her03, Hog02, JS01a, JS01b, JS07b, JS07c, Kal03b, KH09, Kol09, LM06a, LHSL07, LCW09, Nab07, Son00, WZZ00, Wan09c, Xia03, BBS06, Di 03, GW09b, Jai06, Meh00, Wu05a]. M_{2} [RSW02]. M_{n} [Leg04]. M_{n}(A) [CXZL02]. M_{N}(C) [GM07b]. M_{n}(Z_{+}) [BGLS04]. μ [NT09b]. Mx = (M − γI)v [SP04]. N [Ano02-27, ATU03, ATU04, ATU05b, AT09, pBhSzZ09, Hag07, Bai03, pBlWyXwA08, BAL09, BKK07, CCH02b, Cho04, GL08, Guo05d, Guo07b, GW09c, JLY09, KMS00, LZS09, LM01, Yam06, YSL08, PCVT04]. n + 1 [KN06b]. n − 1 [FL07a, CW03a]. n − 4 [YSL08]. N = 1 [Zhu08]. n = 5 [DHH^{+}07]. n ≡ −1 mod 4 [ÁFMNW03]. n ≡ 0mod4 [LY05]. n ≡ 0 (mod k) [ZWY00]. n × n [Bir04]. O(n^{2}) [GL03]. O_{n}(V) [Ish04]. ω [Ots01, Wan09a]. Ω(V) [Ell04]. P [AT09, BEH^{+}06, DH00, GST04, JTU03, MS04c, MN07, Rum03a, Sch04, SG06, TSOA07, DJ07, Dug08, EL01, FF04b, Hou03, Joc09b, Kov02, LV03, Li03, TU02, VLAOVL08, Wat04a, FJTU00]. P (P_{0}) [HMT^{+}08]. P + Q + R + S [CXW09]. p^{m} [Kov02]. ϕ [BD06b]. ϕ_{J} [DoMP09, MPP10]. π [XZ06]. P LU [GT02a]. ±1 [SXKM03]. p → q [Aud09b]. ψ_{S} [HMP09]. P SL(2, q) [LL06b]. P SU(3, q) [Liu03b]. P U(2, 1) [XJ08]. Q [BC03, LR03, MM05, Nak00, NY07b, SG06, AC08, BLP06, BR09, CN02a, CN06, CN07, FN09, IT07a, Joh05, Liu03b, Mas09, Nis02, Siv09, Ter04, Wan09a, YLZ09]. Q^{+}(9, q) [De 07b]. R [Tre04c, Tre09a, YLZ06, Far08, TR03, YLS05]. R^{n} [CKM05, BKK07, QT06]. R_{0} [MM05]. S [Bon03b, XS04, Roj07b]. S(a, b) [dLdAOdF07]. S^{2}NS [SSY05, SHS03a]. sin 2θ [Li00b, TL03]. Sl(2, H) [For04]. sl(3, C) [HNPT06]. SL_{n} [KN06b]. SL_{n}(F) [NY00]. SO(2l + 1)/U(l − m) × SO(2m + 1) [Arv08]. SO(V) [Ell04]. so_{n} [Dvo03]. st + 1 [Roj07b]. T [BO09, LT05, Omi09, WX06b, WX06c, BT05a]. T^{∗} [LWD09]. T^{∗}|T^{2}|T ≥ T^{∗}|T |^{2}T [JK06a]. T^{2} [FKP00]. θ [RBTR09]. T r[(a^{∗}a)^{n}] [MLR00]. U(V U)^{†}V [Čer09]. u_{q}^{+} [Gun03]. vecX [NT02]. V × V → V, dim(V) = 3 [EdRM04]. W [KMS00]. W (H_{4}) [KKABAF06]. W (T) [Del01]. X [vdW03]. X + A^{∗}F (X)A = Q [RR02]. X + A^{∗}X^{−1}A = P [Xu01]. X + A^{∗}X^{−2}A = I [Zha03b]. X + A^{∗}X^{−α}A = I [ESP05]. X + A^{∗}X^{−δ}A = Q [HES06]. X + A^{∗}X^{1}A = P [gSX03]. X + Q + A^{H}(X − C)^{(} − 1)A [gS03b]. X − A^{∗}X^{−2}A = I [CX05]. X − A^{∗}X^{−p}A = Q [LZ09b]. X − AXB = C [JW03, JW03]. X − ∑_{i=1}^{m} A_{i}^{∗}X^{δ_{i}}A_{i} = Q [DLT08]. X = Q + ∑_{i=1}^{m} M_{i}^{∗}X^{δ_{i}}M_{i}^{∗} [Lim09c]. X^{n} = f(X) [JKL09]. X^{s} + A^{∗}X^{−t}A = Q [CC09, DL08, YLF09].

X^{s} ± A^{T} X^{−t}A = I_{n} [LG03]. XA − AX = X^{p} [Bur05, Ikr07a]. XB = D [Xu08, DK08a]. XBX^{2}B^{3}X^{2}BX = A [Lim07b]. ξ [QH09]. ξ(G) ≤ 2 [HvdH07]. X ⟶ ∑_{i=1}^{n} A_{i}XB_{i} − X [Dug01]. X ± A ∗ X^{−n}A = Q [Iva05]. X ± A ∗ X^{−q}A = Q [Has05]. X ± A^{∗}X^{−1}A = I [FG08]. X ± A^{∗}X^{−1}A = Q [HI06, HIU04]. X ± A^{∗}X^{2}A = I [IHM01]. xy − yx∗ [CFL05a]. XY − Y X^{∗} [CL09d]. Z [Fan02, LS00b, Nab07, Smi01a, PCVT04]. Z_{+} [Val02a]. Z_{2} [DT09]. Z_{3} ∗ Z_{3} [LR00]. Z_{n} [SHZ05]. Z_{p} [ACM03]. ||A^{−1}||_{∞} [CH07, Wan09c]. ||f(A) + f(B)|| [Kos06]. ||f(A + B)|| [Kos06].

* [BN08, BC00, GR02]. *-cocharacter [BC00]. *-rings [GR02]. *-subalgebras [BN08].

-adic [EL01, YLZ09, FK07]. -algebra [CXZL02, Li08d]. -algebras [Nak00, NY07b, O’u02, Sud02, Sud04, Yah06]. -Aluthge [AMS05, Oku06]. -Analogs [Siv09]. -analogue [BLP06]. -approximate [RA08b]. -Basis [AAGK06]. -Bernstein [Wan09a]. -by- [AY03, CN02b, CN09, NBdP08b]. -by-2-by-2 [BBS06]. -central [FKRS09]. -circulant [RS03a, WJL02]. -circulants [Tre09b]. -cocycles [ZFW06]. -coloring [SG08]. -Commuting [LL08a]. -complementable [MS04c]. -conjecture [NT09b]. -connected [BDD^{+}01, LRWW07]. -continuity [HMR09]. -contraction [OS01]. -contractions [And04c, Kar09]. -contractive [Rod09, AD08]. -control [FS02]. -corona [Tre04f]. -cube [Bai03]. -cycle [OB01]. -cyclic [Li03]. -D [Yag01]. -determinant [Nis02]. -diagonal [ML02]. -dimensional [Aga02a, GW09b]. -distance [BR09]. -divergence [FS06b]. -doubles [MT04]. -doubly [Nie04b]. -edge [BM04b]. -elementary [FKM02]. -eliminated [BT00]. -extension [LWD09]. -extensions [Joh05]. -factorization [GT02a]. -family [PPF02]. -Fibonacci [LK03]. -fields [Far08]. -filiform [BS01]. -flows [AGKM09]. -fold [KN03a]. -free [Nik08a, Nik09, SS07c]. -functions [Sat06]. -generalized [BT05b, DWD07, LT07a, TR03]. -graded [Bon03b, DT09]. -graphs [RBTR09]. -harmonic [PBR06]. -Hermitian [BNdP^{+}05, Fre08, NBdP08a]. -Hessenberg [Meh00]. -hypertournament [KR03]. -hypertrees [Siv09]. -hyponormal [CLY08, Dug08]. -Indecomposable [YLS05]. -instability [BD06b]. -invariance [Lav03a]. -Invariant [CHKS05, FIL07, TH01]. -inverses [GWW06, WG02a, WG02b, JS07a]. -Inverting [IT07a]. -involutory [Tre08]. -iteration [BBMX07]. -left [ZX05a]. -Lie [pBlWyXwA08, BAL09, pBhSzZ09, QH09]. -linear [OdAJ02, YLZ06]. -local [XL04]. -Lucas [Lee00]. -marked [CP07b]. -matrices [BBD^{+}08, Bru02b, BD07, JS07b, KTR08, LB04, SHRC03, AH08, Ano03-44, AT09, BCEG09, BCPT05, Bör07, BEH^{+}06, BPS04, BCGM08, CNX05b, CT04, Che04b, Che07b, Che07c, CH07, DPZ06, ES08, Fan02, GH03, GN03, HNT03, Had04, HTK07, Her03, HSL05, HMT^{+}08, JS01a, JS01b, JO05, JS07c, JTU03, KH09, Kol09, LS00b, LH04, Nab07, Rum03a, Sch04, Smi01a, TSOA07, WZZ00, Wan09c, XZ06, XS04, yZXtL07]. -matrix [BFPW02, FJTU00, ATU03, ATU04, ATU05b, DH00, Hog02, KY00, LHSL07, LCW09, MN07, Son00, Xia03, YLX09]. -metric [Jai06]. -Metrics [FF04b]. -modules [AR06b, FYY09, Gun03, HJM09, PR07, Sha09, XS08, Xu09]. -nearly [YLS05]. -norm [LV03]. -norms [Pro08]. -numerical [CC01, CN01, CN02a, CN06, CN07, DHK06, NBdP08a]. -Observers [BO09]. -operators [JLY09, Yam06, Kal03b]. -optimal [ÁFMNW03, ÁFMNW06, LY05, MW08a, PWZ05, Udd09]. -orthostochastic [NBdP05]. -periodic [Ots01]. -Pfaffian [Mat05c]. -positive [Aud09b]. -Potapov [FKRS09, FKR09]. -potence [Hua09, SC08, YW07]. -potences [WY08]. -potent [BT05a, Hua09, SBS02, Stu03, ZgC06]. -potential [Mas09]. -power [ZgC06]. -power/ [ZgC06]. -primitive [BK03]. -primitivity [GS09b]. -properties [GST04]. -property [SG06]. -quasi-hyponormal [DJ07]. -quasihyponormal [TU02, TUC04]. -Racah [Ter04]. -radii [NT01]. -ranks [Hou03]. -reducible [Wu05a]. -reflectors [MMT04]. -regular [AH09b]. -rings [Ish04]. -scalar [Wan06]. -selfadjoint [BOP08]. -separation [vdH08a]. -Serre [AC08]. -shape [Omi09, WX06b, WX06c]. -simple [Fie05b]. -skew [Tre04c, Tre04e]. -solutions [PCVT04]. -stable [FGL^{+}00]. -statistic [Dun04]. -statistics [FKP00]. -step [LA05, RM01]. -subgroups [VLAOVL08]. -subspace [HQ08, LLM03, LL08b, LL03, WL07]. -symmetric [LPR04, MH06, Str03b, Tre04c, Tre04e, Tre09a]. -symmetrizable [JSU03]. -system [vdW01]. -tensors [SMS09]. -tetrahedron [FN09, IT07a]. -th [Cho04]. -Transitivity [LM01]. -Tuples [CCH02b, BG09a]. -type [Dam02, LM06a]. -unitary [AADL03, ADL04, ADL06, LS04b]. -variable [CLY08]. -vertex [HD03]. -widths [MQ03].

0 [Tům06, Böt06b]. 0-521-83828-2 [Tům06]. 0-691-11946-5 [Böt06b].

1 [GV09b, vGKN09, Lan05, gLgWmW08, Moh05]. 1-inverse [gLgWmW08]. 1-type [BM04d]. 10th [Uhl04]. 11th [dOQSZ06]. 12th [Ano04t, Ano07s, BGKW07]. 13th [EMS06, Ano05p, BBVR08]. 14th [Ano09p, IMS^{+}09]. 16th [AHST09, Ano09q]. 1999 [dSQS00].

2 [Tům06]. 2000 [Uhl00]. 2001 [Ano03-46, Sch02b]. 2002 [Uhl04]. 2004 [EMS06, dOQSZ06]. 2005 [Ano07s, BLL06, BGKW07, Ros06]. 2006 [Ano05-33, BBVR08]. 2007 [AHST09, Ano09p, Ano09q, IMS^{+}09]. 2D [RO02].

3-540-21100-4 [Kol05]. 3-by-3 [Lás07]. 3D [GHO02].

5 [Böt06b].

7-03-012437-5 [Moh05]. 70th [Olk02].

8th [Pue01].

90 [Zaï04]. 978 [Moh05]. 978-0-521-83828-3 [Tům06]. 978-0-521-88300-9 [Rod09]. 978-7-03-012437-1 [Moh05].

= [WJL02].

A. [BLN04, Bot06a]. abelian [BN08, DKX03, KO08, RG03, CKK08, FZ04]. ABLE [Wu06a]. Absolute [Ips00, MM06a, Tre03b, Tre03a, EMERP06, HDA09, Ips03, RST01, Rad03b]. absorption [KN01]. Abstract [Zem07, Per03, Prz02]. abstraction [Per03]. Accelerated [SK08a, XZ06]. Acceleration [Rih01, Cen06]. according [Wan09e]. accretive [Nie09]. Accuracy [CCTT00, BNT05, Koz09]. Accurate [DK06, Fer03, KD07, BPV00, Hig00, MM07, Sla03]. accurately [OS00]. achievable [– DJ07]. achieve [Das04a, YLT05a]. Action [AM05, BFZ05, Bra03, DW05a, LK00]. actions [DDI03, Pre05]. acyclic [BFS08, BOvdD04, CdFM08b, JTU03, KS08, Moh07, NS07a, CdFM09]. Adamjan [IO01]. Adamjan-Arov-Krein [IO01]. Adaptive [KHG04, CPPS00, FLMR00, LCF06, PEPS00, SK08a, Wd09]. adaptive-order [LCF06]. added [RvdGM05]. addendum [DL01]. adding [Guo06a]. addition [AGO09, LW03]. additional [OCK07]. Additive [BS00c, BS09a, gCZ03, CL05b, CL06, Cui07, CIDW06, Gon05, HC03a, HQ08, HJ08, JLL03, Kuz02a, Kuz02b, KO06, LH03, Leš02, Lim05a, Lim08a, PY09, QH09, RST01, Rad03b, Tan05b, YT04, Zha05a, AGK06, BPS04, CD00a, Cal05a, DJ01, DG02, DH04, HWW01, HZ04b, HZ05a, HBLZ06, LK09, Lim05b, Lim06, Zaï04, Zha04c, Zha05b, dSG07, AH09a]. Additivity [JL09, Ji09, LL05b, Lu02a, CL05b, YT04]. adic [EL01, YLZ09, FK07]. Adjacency [AHZ05, CLP03, CH06, CLOvdD01, Gri04, Jør05, PFL05, RS05b, Sav05, vDHK07]. adjacent [Har05]. adjoint [CWW09, Del01, HZ05a, Lan08, LPR04, Mol06b, RA08b, Rob05, Tan03b, YLZ06]. adjointable [Xu08, XS08, Xu09]. adjointness [MPT04]. adjoints [Del01, IS08]. Admissibility [HG09, Wu08c]. Admissible [LS02d, GM04, ZGL09]. admit [AT09]. admitting [CDJ08, Liu03b, LL06b]. Advanced [Kra05]. Advances [Hil07b]. advantage [CM09]. adventures [Dax03a]. affected [PA04]. Affine [Jel03, RG03, BM06a, Ber04, BNV05, DDI03, Gew03, GN09, OP05, Ple07, PPZ09, RVZ00, smbH02]. affine-symplectic [GN09]. age [PSvdW05]. agent [ZTK09]. aggregates [DHFS00]. Aggregation [BBFP04, LT04a, MM03, MP06, PW04, Pul07, SK08b]. aggregation/disaggregation [MM03]. Aided [CFM03]. Alan [Bar05]. Alexandrov [Kle07]. Algebra [Ano02-28, Ano03v, Ano03-31, Ano03-43, Ano03-46, CGN01, FMN04, HL08b, JS07b, SP09, Yan07, AAA^{+}02, AD02a, AGDR05, AFR06, ADC08, BES02, Ber04, BV07, BS03b, BBP03, But03, BCG07, CW00a, Cao01, CXZL02, CT03a, CW03b, Cap01a, Cap01b, CEK^{+}01, CKK08, CK04, CSK01, CGB04, CIDW06, CI09c, DDvG05, DV04, De 00, DT09, DHK06, Dra08c, Eic04, EvdD01, EvdD04, FN09, Gav00, Gav02, GP03, GQ09b, GC04, GL05b, HH00, Har07b, IS09, IT07a, Kye03, LK03, Li08d, LM06c, Lur05, LY09b, Mat04, Mei05, Mei06, Mik09, Mol05b, Okt04, OWY07, PG01, Pap00, Sem06a, Sem07, SSB09, SC07, Sob08, Str03a, Str06, TR01, TX09, Ter02, Uhl03, WSL02, WY04, WY06b, WY06a, WYZ07, WL09a, WZCL08, ZXZ08, FMPT04, Zav07]. Algebraic [BL09c, Bou09, DR08, FO01, Kra03a, LZ08b, MS04b, PJ07, PRV05, Sal09, SH00a, Zha09, BP05a, BPT07, BILM06, Bra07, BP09, BCS02, BP06b, But08, CNR00, CFL05b, CLY07, DSW07a, DRGR04, FKP03, FPP01, GB06, GKvS04, Guo02, Guo07a, HM02, HL09, KMNS02, Kir03b, Knu00, KTKR06, Leš04, MN08, Mol06a, NZ06, Nie01a, OCK07, PL08, RM06, Roj09a, Ser09, SGS08, SU09, gS02, TI08, Tre09d, Xu07, Yah06, ZF07, Zha07c, ZW00, dA07]. algebraically [EdRM04]. algebras [Aga02a, AJF09, ABW05, Ali09, AC03, AS09, Bae07, pBlWyXwA08, BAL09, pBhSzZ09, Bal07, BCES05, BBC00, BLS05, BBL^{+}05, BS00c, Ben09, BS01, BC03, BCSV06, BKU06, BD03b, BPP08, BM05, BP06a, BGO09, Brü03c, BDV08, CM00a, CGN01, CS03a, CFN04, CLP03, CL03b, CT09, CD05, Cos07, CHL01, CGH04, CL09d, DZ01, DLZ03, DP02a, Dol03, DM02, Dom03, Drä03, Dun07, EMERP06, Eld00, FYL07, FP02a, FMM05, FM07b, FLRZ02, FMZ06, GML09, GR03, GZ07, GJMR01, GST04, GTM09, Ha03, Han03, HPP00, HL07b, Her03, HC03a, HC03b, HZ04b, HQ08, HJ08, HQ09, HN09, HM06, IS08, JW05, JW06, JL09, Ji09, JGPI08, Kim05, Kle09, KMMZ04, LR05a, LZ08b, LN08, LS06b, LLM03, LL05b, LL08b, LL09b, Lin03, LL04, Lip08, LH05, LR04, LP00, Lu02a, Lu02b, LL03]. algebras [Lu09, MT04, MG09, Mol06b, Nak00, NY02, NY07b, Ohn08, Oik08, OZ03b, OZ03c, Oku05, O’u02, Pei07, QH09, RST01, Rad03b, RM01, Sch09b, SS09b, SS06d, SSW08, Spi05, Str03b, Sud02, Sud04, WY08, WL07, Wil09, XX03, XL04, Yah06, ZJC04, ZYP05, ZFLW06, ZY06, ZFW06, ZPY06, ZYP06, ZWC06, ZZ08a, Zho00, ZG09, ZX05a, ZX05b, Zhu07, ZX07, ZXZ09, Zit09, Zus05, dlMW06]. algèbres [BCSV06]. AlgLatT} [Del01]. AlgLatT}/ [Del01]. algorithm [Aga02a, Aga02b, ABK02, AADL03, ADL04, ADL06, AGI00, AY08, Are09, BBD05, Bau03, BEG^{+}07, Bra03, BZ03, BQW09, CW00b, CFL05b, DW08, Dax00, Dax03a, DM07, FPP02, FNV08, FB04, Gav03, GL03, GZ08, HT01, HOP06, HC08, HS06b, Kle09, KMS01, KLMS03, Kre06, LQ00, LT08, MS01a, MM07, MVV08, McG06, Mel01, MZ06, Mül03, PM00, PHO07, PS07, PVV08, Pop01, Pop08, Prz02, Sid01, Ste07c, Van04, VWJ08, Wan01, rWQ02, WX06a, Wil07, Wu06a, XQ08, Yag01, YH04b]. Algorithmic [BD08, BV02, DMM00a, Dub09]. Algorithms [Ano03-45, Ano06f, AT08, FLQ04, Pen06, jYxL09, Arn04, BKDL01, BKO01, BBP03, CFM03, Cen01, Cen06, CGXL08, CCSS03, EG03, FMM01, GG09, GS03a, HR04, HL09, JPB08, MLK^{+}04, PZ00, PT05, Pot03, Zar06, HT09a]. alignment [HS09]. All-derivable [Zhu07, ZXZ08, ZXZ09, Wu09]. all-pass [HOP06]. Allow [COvdD09, Hua09, JS00b, SG03]. allowing [GLS07a]. Almost [JS01a, AFL00, AHL^{+}09, DHFS00, E^{+}00, Gar03, GM08a, Kir03a, PT03, PT06b, Ste05a, Ste00b, Zho00]. also [Li01]. alternants [Ash05]. alternate [YT04, Zha04c, Zha05a, Zha05b]. Alternating [Zar06, BPS07, CS08a, CW00b, CL02a, EdRM04, FS06b, GQ09a, HS03, NB03, Wat05]. alternating-Hamiltonian [Wat05]. Alternating-projection [Zar06]. Alternative [TY07, BL04, CC03, Rob06, SMK09]. Aluthge [AY03, AMS05, INOY03, Oku03, Oku06, Wu02, Yam02]. always [RSR00]. among [CHRS03, HS07b]. Amsterdam [BBVR08]. analog [SvdS01]. analogs [Zus05, Siv09]. analogue [BLP06, GK09, dH07, Igl01]. analogues [BP06a, MMT04]. analyses [Gri00, Gri04]. Analysis [Ano03-45, CP03, GM07a, KH01a, THH04, ADC08, AS02, Baz03, BM06a, Bén02, Cap03, Car09, CNX05a, CX05, CC06, DNR04, DKS04, DW05b, GvdV04, HNX03, HT07a, HB00, Hen07, HM04, HS09, Iva05, KFGK04, LZ09b, LgS01, MP06, NT02, PPP06, PW04, Pfa08, PSvdW05, Pul07, Rod06, RSW02, SC06, Str02, gSX03, gS03b, Vir08, VWJ08, WZ07, XZ06, XWD06, Xu01, XDS08, YTI06, Yeu04, vdVN00]. analytic [BES02, BES09, BR01, Bol02, CL07b, FKLN07, MGWN06, Wal05a, Wil07]. analytical [JP04]. analytically [Cao05, FKS02]. analyzing [BBP03]. and/or [SS07c]. Anderson [BT07b]. ando [XXZ09, BBM02, FK06, Lim08b, MPS00, ST08, Yam06]. angle [AEK06, RW01]. angles [BT09b, CFRS02]. angular [ZQ07]. anisotropic [FH06]. annihilating [BGS09, Obl08]. annihilation [HP00]. anniversary [Olk02]. ANOVA [Jia04]. answer [HL07a]. Anti [Meh00, ADC08, DW09b, Fie01a, Gem03, Hol05, KN03a, yPyH03]. anti- [Meh00]. anti-bidiagonal [Hol05]. anti-invariant [KN03a]. anti-Monge [Fie01a]. anti-reflective [ADC08]. anti-reflexive [yPyH03]. anti-stable [Gem03]. Anti-triangular [Meh00, DW09b]. antiderivations [Ben05]. Antieigenvalue [Sed09]. antirings [DO09a, DO09b, Tan07, Tan08]. antisymmetric [Yin04, Yin05]. any [Roj06b, RR07a, Roj09b]. AOR [LS01]. appealing [fur08b]. Appl [Ano03-31, HL08b, JS07b, Yan07]. application [AH03, BdPK07, BI00, BP09, CFM03, Čer09, DMMV09, DPZ06, Dun04, Est05, FWG08, HFR04, HC08, Jia00a, JOR01, KMS05b, MB07, MGWN06, PA04, SvdD05, Ser09, Xue01, ZJL02, WLBZ06]. Applications [Ano03-45, Hol00, HHT06, LXW04, SR06, AKL07, BB04d, BHNR03, Bat06, BKPW08, Bon03a, Bul08, Cao09, Cap03, Che05, CNSM08, CXW09, CNF08, CBD^{+}04, DL04, DK08a, EMT05, Elf04, Fan02, FK07, Fur08a, Fur06b, GH00a, Gol05, Gra02, vGKN09, Had05, Han07, HT02, HP03b, Hun06, Hun08, JMT04, JS04, KV09, KL07b, LS01, MPP06, Mül03, Nah06, Nie08, Pfa08, RL03, RW02, Reu06, Rie00, Rod08, RR07b, SSW05, SC06, SU09, SS09d, Ste05b, TR01, TR02, Tak04, Tan05b, TBK09, TS09, Tsa00, VWD08, WvdWC09, WlW07, Xu03, YFK05, YTI06, YM07, Zen01, Zha05a, tBS06, Tům06, vdD09]. Applied [FMN04, Di 03, KS09, Li03]. Applying [BBP03, RT06, TVB08]. appreciation [BHN^{+}07, Ros06]. Approach [vdD09, AGDR05, AVV05, AP04, Arp03, Arp05, Bak04b, Baz03, BPT07, BSS09, BZ03, BP06b, Cap01a, CTW01, CC06, DMM00a, DS03a, Dub09, EG04, FL04, FLPW07, FN07, FP02b, FM04b, FVP06, FS02, GN05, GTT01, GL05b, HNV00, Haz05, Hei01a, LS09b, LCP02, NPP00, OT06, OT01, PRW01, QC04, RR05, Rob06, RZ05, Rod05, RW01, SH09, Uhl03, Val00, VS04b, VZ02, WL03, XM02, XDS09, YLZ09, ZM08, dST06]. approximants [CL01a]. Approximate [AHW05, Bur06, DV00, NS03, PW04, Cao07a, DTGV06, Gha01, IK06, Pop01, RA08b, SA02, Ste02b, Wd09]. approximately [OV06a]. Approximating [HW05a, HO01, BC04b, Dah05, SvdE03]. Approximation [FKP00, HLT04, LZ08a, PJ08, ABB00, BS09b, BNT05, Bör07, Cen06, CC00, CL01b, CFP03a, CHP07, DLMO07, DP00, Dru08c, EF07, ESC08, Fei04, GS04b, Gu01, IBI09, JLMP02, Koz09, Li06b, MV07, NB03, Ovt06a, Ovt06b, pPyHZ05, hPyHZ07, RW01, SC07, Tic05, Tre04d, Wer08, ZW03, ZHZ08]. approximations [BKK08, Gri00, GV06b, RW06, RS02, Ste01b, Tyr04, XHpS06, Ye08]. April [Ano05-33]. AR-representations [Kar04]. arbitrarily [DS08c]. arbitrary [AMZ09, BD05, BOvdD07, CV05, CVV09, DHH^{+}07, DJOvdD00, EGO05b, HR04, KOvdD07, MTvdD05, MS08, ML02, Pop01, Pot03, PG04, Son02, VW00, VR06, Wan04]. Arc [KO08]. Arc-transitive [KO08]. arcs [SW03]. argument [FS07]. arguments [Chi04]. arise [De 09a, MMX04, SG09]. arising [BLdP04, BLM02, BLM03, BILM06, CBIXG08, CLB09, CP09, De 08, FB05, FH04, Fuj05, GG09, HNX03, JLMP02, LCN04, MM06b, OZ03a, Wan07b]. arithmetic [BK00b, Bha06, BK08, CE09, FNT05, Jan01, KL07a, Kul01]. arithmetic-geometric [BK00b]. arithmetical [Hon06]. ARMA [KV09, KMS05a, KS06b]. ARMAX [KS05]. Arnoldi [Bag08, BLR00, LCF06, MZ06, Ste07c, VV06]. Arov [IO01, Rod09, FKM02]. Arov-normalized [FKM02]. arrangements [BHL^{+}04]. array [Jai06, Jai07]. arrays [CtBDC09, TBK09, tBSR04, tBS06]. articles [Sch00b]. artifacts [TKX08]. Artinian [LZ08b]. Arveson [BR02]. Ascent [SdSW07, BS09a]. Aschbacher [CS05]. Askey [Nom05b, Vid07]. Aspects [Huh05, FO01]. assessment [CCTT00]. assignability [SSSG08]. Assignment [AMZ04, AMM06, GW02, GT06b, LK09, RZ01, RZ05]. assisted [Plu01]. associated [Alt09, BFL05, BSW06, BSL04, BLLS06, CFN04, DMM00a, FMM05, Fur03, Fur06b, GLW09, HV03, HTV02, Hon06, IHM08, JTU03, KH01b, KH03, KM07, Lee00, pLLfY06, NT06b, O’u02, pPyHZ05, Pou01, Sal09, Sat05, Tre04d, VPV01, WL08, XDS08, XDS09, ZW00]. Associating [GPM01]. Association [GWL09a, GWL09b, Tag04]. associative [Are09, BLS05, CC04, CHP03, CHJP05, FM07b, LN08]. assumptions [ATU04, FJTU00, YW07, Zha04b]. Asymmetric [FPP01, Wat01a]. Asymptotic [AL09a, Bar08, Ji03a, MW03b, NCV05, RSW05, Tre06, You07, GZ01, HBHS05, HB00, HT07b, Lur05, PRP02, SV07, SC07, Tre03c, Tre04b]. asymptotically [FLPW07, Pit09, OP05]. Asymptotics [Mal09]. asynchronous [BKS01, Str02]. atlas [LST08]. atomic [Ha03]. Atoms [Ros09]. attached [Bap05, Roj09b, RM09a, Tag04]. attaining [EKS01]. Auburn [Uhl04]. augmented [Cao09, Bag08]. Auslander [Kra03b]. Author [Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano01o, Ano02p, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano03r, Ano03s, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano04r, Ano04s, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano05o, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano06a, Ano06b, Ano06c]. Author [Ano06d, Ano06e]. auto [Del01]. auto-adjoints [Del01]. automata [BBFP04, CCG01]. Automated [BHS06]. automatic [FLMR00]. automorphic [BN05]. automorphism [Abd05, Chl06, Eic04, FJ04, Kes00, TX09]. Automorphisms [ALMS06, Cao01, CT03a, CSWM01, HP02, Kim05, Leg04, Lip08, WYZ07, CW00a, CCH02a, CLY07, DKX03, Eld00, Ish06, Jel03, Kov02, WY04, WY06b, XCX01, XL04, ZYP05]. autonomous [McC05, Val00]. autonormal [Udd09]. autoregressive [KS01]. Aveiro [Ano05-33, Ste07b]. average [BS08, CGHG06, KS01]. Avram [BGM08]. AZB [Bou05].

BA [DK04]. back [And07a, Gol05]. Backward [AD03b, HP03a, Ste02b, Tis00, Wie01, BR02, BKO01, CCTT00]. Baker [BC04a]. Baksalary [BS05a, Cal05b]. balance [BC08a]. Balanced [BD06a, HOP06, NT07a, Sty06, BPS09, CLVV06, HT08b, Jia04, Roj02, RP03, RS05b, SH00b, SA02, Uhl03]. Balancing [BK06b, CD00b, SK08b]. ball [AD02b, ABK02, AD03b, AD03c, AD04, Bol02]. balls [MQ03]. Banach [ATV07, BES02, CIDW06, CI09c, DH03, Din03, Ed05, GGM^{+}09, Her03, HM04, HM06, Jam07, Lu09, NY02, Sch09b, SS09b, WZ07, ZZJ08]. Banachiewicz [BS02a, TT09]. band [BG02, BES03, EG04, EG08, FLM02, GR00, Lin08a, LRR04, McM09]. Banded [Kra01, CG00b, DH09, GGCB07, HV01, Kar05a, KT08, NB03, Sch09c]. banks [CMX05, Jia00b]. bar [FLROV05]. Barcelona [Pue01]. Barker [SMK09]. Barnett [VPV01]. Barnett-type [VPV01]. Bartholdi [MS07]. base [BDM03, CL08a, LY06]. based [BZ06, BEG^{+}07, Bol01, BDKW04, CGHG06, CHP06, DG07, DM07, FM04b, Gug08, GWL09a, GWL09b, HHM09, HV01, IWL07, IK06, Jan01, JL03, KM04, Kir04, KTKR06, LMV03, Lim09b, LLT05, LzZT07, MVV08, MPS00, MPS03, PW04, RW06, RO02, Smi08b, Sun02, Ter01a, Wri06, YKF05, ZM08]. Bases [CGB04, AVV05, AP04, BOP08, BKU06, BSS07, CNSM07, DJ06, HJM09, Kro01, LXW04, Ma09, NT09a, Pit03, PR04b, YSS07, Zob00, dlMW06]. Basic [ARR06, CNSM08, Fie03a, Dom03, Fie04, FM04a, Har00b]. Basis [AAGK06, BPR03, Bor08, Bou09, BP09, Gu01, NT06b, OF02, Pou01, Yan01b, ZJL02]. Bauer [Ano05q, Ano06h, wCL04, FGH^{+}06]. Baxter [FLROV05]. BBP [SMK09]. be [BW05, Mir03, Nar05, Ste00b]. Beam [EHR08]. beams [DN00]. Bebiano [FNT07, Fur07, MF09]. Bedlewo [EMS06]. Behavior [Böt06b, Bar08, BES09, Cap01c, De 00, EN09, HT07b, LMT09, NCV05, PW08b, RSW05, SV07, SP04, TE05]. behavioral [PRV05, RW01, Val00]. Behaviors [WY07, Fuh02, FRY07, Val02a, Yek04]. behaviour [Woo02]. Behrens [IV04]. Beidar [Che08]. Beijing [Moh05]. Belitskii [Kob05]. Bell [WW07]. bellman [XXZ09]. below [SS08a]. benchmarks [LRK06]. Berlin [Kol05]. Berman [SMK09]. Bernstein [MM07, Wan09a, Win00, Win03]. Bessis [Hil07b, LS09b]. best [Bak04a, BD09a, Cen06, LwW07, MV04]. Bethe [RM06, RR07a, Roj07c, Roj08b, Roj09b, RM09a]. between [AEK06, AAB^{+}07, AB05b, BBL03b, BH04, BB04c, BHLL04, BT09b, Bal07, BFH05, Bri05, BAD09, CZ05, CE09, Dax06, FNV08, GPMT06, GHI09, Gol06, GSR03, Joc09a, Kos06, LRŠ02, Lim05a, Lim09a, McE02, MW01b, Nab07, NT09a, Tan03b, Tan05b, Xu07, ZgC06, ZQ07]. betweenness [CG07b]. Bezout [BM04a, DMMV09, VPV01]. Bezoutian [KS06b, Yan01b]. Bezoutians [HR01, HR03a, OO05]. Bhatia [And07b]. bi [Pfa08, WSL02, Yin04, vdMSR01]. bi-antisymmetric [Yin04]. bi-infinite [Pfa08, vdMSR01]. bibliography [BHN^{+}07]. Bicircular [Jam07, SZ04, FIL07, Ili08]. bicommutant [Str03b]. Bicyclic [HLZ09, PBA09, Guo05c, Guo05d, Guo07b, HXL08, LZ07, ZWS09]. Biderivations [Ben09, ZFLW06]. Bidiagonal [DK08b, GT08, BBD05, Fer07b, GL03, Hol05, Ral03]. bidiagonalization [BQW09, MS03a]. bidisk [Tre04f]. Biehler [HDB00, Hol03, OTC03]. biextensions [Drä03]. bifurcation [GPS01]. big [BW05]. biggest [Sav04]. bijections [AT07]. Bijective [LM06c]. Bilateral [ACS06]. bilinear [AC08, Are09, BD03a, BS06a, BBK03, BM06b, CLL05, – Dok06, EdRM04, GQ09a, HS06b, HS08, IS08, KKP08, KLY09, Ose08b, Tan09a, ZS05]. bimodal [MSR06]. Bimonotone [QT06]. Binary [Cen01, FMZ06, dMFZ09, DV01, HLR08, Kun06, Lem04, Roj02, RP03, VHK01, Wu06c]. binomial [BB04d, Hou03, Sev08]. biography [Ano09r, BKR^{+}09, BHN^{+}07]. BIOME [AFCC04]. Biorthogonal [CMX05, Fie03b]. Bipartite [Jef05, BOvdD04, CCRS08, DGN02, FQ09, Lee00, Sat06, Ste00b, Ste02a, THZ09, ZLS09, ZZ08b, dW03]. Birkhoff [BdPK07, Car04, CC03, CdFM08b, CdFM09, Lim07a, MOSZ02]. birth [Blo08]. bisubmodular [Kar07]. bisymmetric [hPyHZ07, ZHZ08]. bitangential [BB02b]. bivariate [HW00, dBS08]. bivectors [LT09]. Björck [BEG^{+}07]. black [CEK^{+}01]. Blaschke [GW03b]. blind [CW00b, MLK^{+}04, AFCC04]. Block [ADT06, Bag08, Bor08, BD04, KY00, Smo08, Wir09, Aud06, Aud08, BW00, Bér05, BDG08, Bol04, Bom08, BW01, BES03, BD03d, Cal05b, CGD05, CL04a, CL02b, CL03b, CL06, CHP08, CI08, Dan02, DP00, Dru02, FT00, Fie05b, FKM02, GGCB07, GS09e, HC06, IO01, Kle04, Kol02, Kol03b, LMMT01, Las06, LW07, Lop07, MSW09, Mol08a, PP01a, Peñ03b, RS02, RS06b, SSSG08, SS08b, Tia00, Tre09b, Tre09c, Tri02, Wan06, Wim01, XZ06, YH04a, ZJL02, yZtLC06, dF00, vdMSR01, vdMSR03]. block-diagonal [Tri02]. Block-diagonalisation [Wir09]. block-matrices [Bol04]. block-structured [Peñ03b]. block-tridiagonal [KY00]. blocks [CS00b, CS03c, CS03d, FGM03, KN04, Kun06]. Blockwise [Xue01]. blowing [SV06b]. blowing-ups [SV06b]. BLUE [KP06]. blur [CCSS03]. board [Ano02u, Ano02v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano03-28, Ano03-29, Ano03-30, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano04-37, Ano04-38, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30, Ano02t]. Bode [Igl01]. bodies [Pue07]. Bohr [ABP09]. Boltzmann [TVB08]. Book [Böt06b, Kol05, Moh05, And07b, SS07c]. book-free [SS07c]. books [BHN^{+}07]. boolean [Liu03a, AFS09, GL09, HLR04, LYY07, Wat01b]. bordered [DP01, GWW06, PES07, WG02a]. bordering [Bal00]. Bose [WF04]. both [CH01, FKRS09, NT07b]. Böttcher [Lás07]. bound [AFS09, BGS00, BFH^{+}08, BV00, BH08b, CCRS08, Che04a, CH07, cD03, Das04a, DB05, Das07, Das09, Eis06, FLQ04, Guo05a, HOZ01, KP01, Kir03a, KN09, Li08c, LB04, LLT06, Luz07, MV04, RSR00, Roj07a, SHWR02, Van07, Wan09c, Wan03b, YLT05a, dSS08]. boundaries [LSH08]. Boundary [AD02b, BdPT08, Bol02, BS04c, CWW09, Cap01c, CSSZ01, CH01, CY00, CNP02, CN09, CCG01, ĆDR01, FJ03, GS04b, HdABC07, JY07a, JY07b, Kru04, Li06b, Mou06, RS05a, SS06e]. bounded [BF08, Buc06, Dra08a, Dra08b, Lás03, LT07b, LT09, Mol02, Nik08a, PR04a, SS08a, Wer08, Wie01, XH08]. Bounding [Gol06, Ste03a]. Bounds [EvdD08, Fan07a, GH06, Gil04, HT03, HL02, Kol09, KH01b, LY06, LLT05, Ma09, Mor08, Nik07a, Nik07b, PT06b, Shi07a, Thi04, Ves03, YSS07, BS00a, BZ00, BJ09a, BHS06, CS03b, CW07a, CW07b, CKW08, CM01, CL04b, CL01b, CG07b, Das04b, DSW07a, DMM00b, Fio05, Fre04, GB06, Gre09, GJW02, HL04d, HLS00, HW07b, Ind09, Ips00, Ips03, JN06, Jai07, Kol08, KAP^{+}03, LP01c, LHSL07, LCW09, LHF06, gLgWmW08, LzZT07, MVV08, MSF05, MK03, MK05b, MK05a, NX04, Pan02, PR01, PP01b, RA08a, RS02, RO02, RR04b, Roj05, Roj06b, RM06, RR07a, Roj08a, Roj09a, Rum01a, SM09a, SS07c, SW04, SS00, SS03c, Sou05, Ste02b, Tia02, Til03, Ton04, TS06, Wal05b, WV08, WLBZ06, Wu05a, Wu06b, Wu08b, XX09, Yam01, Zha04f, ZH02, dSR05]. Bowen [Che09, MSR06]. box [CEK^{+}01]. bracelets [SH08]. Branges [AD03b]. Brauer [Kol03a, SR06]. breakdowns [RS06b]. bridge [Qui07]. brief [BHN^{+}07]. Browder [Cao07a]. Brualdi [vdD09, BBCH08, HWKS03, HP07]. Bruhat [BD07, MH07]. Brune [ABDD01]. Brunovsky [BBdH09]. BTTB [FH06]. building [CM09]. bundle [Gra04a]. bundles [KKL08, KK01]. Burnside [LR04, Lon04, RRY04]. Burr [MPS06]. Bursts [Jai06, JN06]. Busemann [FF04b]. Bush [Par02b]. Byers [Meh08].

C [FJ02]. C-S [FJ02]. Cache [BZ06]. Caen [LP01c]. calculating [MZ06]. calculations [HO02, Nör03, Str01, Zar06]. calculus [BR00b, Bul08, Dru08a, KS09, Kra05]. Call [Ano02q, Ano02r, Ano03v, Ano03t, Ano03u, Ano04u, Ano04v, Ano04t, Ano05p, Ano05q, Ano05r, Ano06f, Ano06g]. Cambridge [Rod09, Tům06]. Campbell [BC04a]. Campopiano [Jai07]. Campopiano-type [Jai07]. Can [BG02, BD01b, BD06b, BW05, Nar05, Ste00b]. cancellation [DN05a, DN05b, OS09, OA09]. cancellative [RGG01]. cancer [CCW^{+}08]. Candecomp [BSD09, SS07f, Ste09]. Candecomp/Parafac [BSD09, SS07f, Ste09]. Canonical [BBS06, BKU06, BRS00, DNR04, GH09, HS06a, HS08, LR05b, PHO07, Rod07, Ser00, Ser08, TG02, BOP08, BPZ08, Dri04, Dru02, GS04a, GH00a, Hol00, KPK04, Kob05, MH07, OO00, PPF02, RS01, Ser04, SMS09, WW08b, ZM03, dMFZ09, vdW02, Rod08]. capability [KP09]. Caractérisation [BR00a]. Carathéodory [AD03c, AD04, ATV07, ATV09, BKR04, HdABC07, Las06, Woe02]. Carathéodory-Toeplitz [Woe02]. cardinal [RR05]. Carlson [CF02]. carrier [LCP02]. Carroll [BSD09]. Cartan [Cos07, XX03]. Cartesian [BK00a, CHL02, Gol09, KSH08]. case [AH08, Alf05, ABDD01, BBdH09, BF05, CMP03, CH00a, FKS09, FJ02, FJL08, HB00, HDB00, JDS03a, KN09, PM00, TR03, Val00, Yua00]. Cases [Nie04b, BdHZ08b, DTGV06, FKRS09, Ney06, OTZ09, RS01, Van04]. catacondensed [Rad03a]. categories [Kra03a]. category [Gun03]. caterpillars [Sta09a]. Cauchy [BKO01, GVD08, Lee09a, PZ00, VS04b, Wad07, YYZ05, zY08, YC00a]. Cauchy-like [PZ00, GVD08]. Cayley [AD03a, BR00b, HT08a, HT09b, Hig08, Ili09a, Kah06, Kon08, Koš03, RS05c]. Cellular [XX03, CCG01, PPZ02]. center [Yah06]. Centers [OS03, CT01]. Central [CK04, Che08, FKRS09, HKKZ03, HKKL05, WSL02, Zav08, ZHZ08]. centralisers [Haz05]. Centralizers [WL07]. centrally [CMX05]. Centro [HR01, HR03a, XHpS06, ZHZ03]. centro-skewsymmetric [HR01, HR03a]. Centro-symmetric [HR01, XHpS06, ZHZ03]. Centrogonal [KS00b]. centrohermitian [FI03]. Centroid [BAL09]. Centrosymmetric [DS03b, HR03a, FI03, GR07a, Han07, Hei01a, Mel00, pPyHZ05, RR04c]. Certain [Chl06, AKYD16, ALMS06, AHLR09, BBL03b, BFH04, BPR03, BDG08, Cao01, CSWM01, Del01, EN09, Fen03, FT07, eFLW09, FK07, Gra02, GT09, HNX03, KD07, Kol09, KBdP07, Lav03a, Lav03b, Lav05, Nie04b, Nis02, Ose08b, Pou01, PH09, Roj06a, Roj09a, ZPY06, ZG09]. certaines [Del01]. Cesaro [FKS02, Yan02a]. CF [Hag07]. Chain [LW03, CNX05a, CM00c, CM01, GL05b, Hun08, Hun09, PW04, Shi01, dW03, Ste07a]. Chains [Bar01, Ben04, BM04d, CZ08, CNF08, CDS08, DHFS00, DHZ03a, GS09d, Har06, Hun05, Hun06, KN00a, Kir04, KNX07, Kru04, MS04b, MSF05, NX04, RA08a, RT03, Ser04, VM04, Wil07, Xue01, Yan02a, Yan09a, Zha04e]. Challenges [Ano02s, Uhl00, Uhl04]. Chan [CJ04, CJW05, CJ05, JW08]. Chang [BSD09]. change [BdHZ05, BdHZ08b, BdHZ08a, DS08a, DS07, Sav05]. changes [Bo02]. chaotic [FKN02, Fur02, Lin07]. character [HB08, Pre05]. characterisation [KM04]. Characteristic [CDS01, DW05a, Hol08, KK01, LK00, Rad03a, Alf06a, pBlWyXwA08, BBS06, Bor06, DSW07b, Eld02b, EV05, FGWG07, Gao03, Gil04, Gil08, HT08b, KKL08, KT08, KZ03, OdAJ02, TMARAE^{+}07, Tup09, Wat04a, Wu08d]. characteristically [BC03]. characteristics [MM08, Rod06]. Characterization [BCEG09, BT07a, Čer09, CZ00, DK07, Tre04a, Tre08, YLT05a, ZGL09, ZX05a, ACK05, BL04, BvDK08, BLL03, BG09a, BD05, CW03a, CT04, CL00b, Das04a, Das07, Dri08, DRR02, DD05, DL05, GOTR07, IPP07, JS01b, KTKR06, KLY09, Luz07, MS05, NT09c, OS01, Omi08a, Omi08b, Omi09, Peñ03b, RBTR09, SSSG06, VR06, WX06c, WX07a, WHBM09a, ZM03, Zha01, Zip03]. Characterizations [AH09a, BH04, BT05b, BKS03, BG06, DW09a, GM04, Lu09, MK03, Peñ04, Sch08, ACV09, CT03c, HLZ08, TY07, Tar05]. characterizing [Bak04a, BLNP01, Smi01a]. characters [HLS00, Pou01]. Charles [Bar05]. Chebysheff [LdlP09]. Chebyshev [EFS01, EF07, Hei01a, Li01, LA05, Li08a, Pug04, rWQ02, WlW07, YLZ09, Zyg02].

Chebyshev-Hankel [Hei01a].

checkerboard [ACDR04]. Chen [Tům06]. Chevalley [KKK07]. chief [Hof08, Bru00, Bru02a, Bru03a, Bru04, Bru05, Bru07, Bru09a]. chirp [Yer04]. Choice [AG08]. Cholesky [BB05c, PVV08]. chordal [Qi09]. Chromatic [Big02, Nik07c]. Chvátal [JK06b]. circle [AADL03, CMV03, CMV05, FMM01, Par09a]. circles [Kol05, Var04]. circuit [RM09b]. circuits [LRWW07]. Circulant [Doo02, ALSS08, ACM03, BS08, CDS08, CR02b, DDvG05, KSU03, Kov02, Mul05, OT06, RS03a, RR04c, RS09, SS06b, Sbu05, Sbu08, SSMQ^{+}00, Shp06, SS05b, ZJL02, WJL02]. circulants [Tre09b, Tre09c]. circumdiameter [DOvdD05]. circumscribed [GW04b]. CIS [AGI00]. Clar [AA07]. class [Bag06, BM01a, BN05, BPR03, Bol02, BS04b, ÇS02a, CZ07, CGD05, CQ08, CS05, Dug05b, Dun07, EG03, EL01, FL07a, FP07, FH04, FJ02, FJ05, GKL07, Guo07a, Ha03, HT08b, HvdH07, HHT06, HL08a, JSU03, JY07a, Kat04, LMZ03, Li03, Nak03, OTC03, OS05, pPyHZ05, Pat04, RO02, SS07b, SC06, Shi06, Sol06, Sza07, YMP01]. Classes [BCGM08, FGL^{+}00, MR01, Aga02a, AT00, ARR06, BLNP01, BL05, BKS03, BEH^{+}06, De 08, – DZ04, Ell04, EK08a, FJ02, FJ05, GS09b, Hog03, Hon06, JS04, JS01b, JKRT02, MM03, Mil03, MND01b, Nab07, NY00, ND05, OdAJ02, Pou01, RS06a, Sch08, Ser09, SSG04, Tad06, VLA03, VLAOVL08, Wu05c, Yan06, Yan07]. classical [Bou06b, FKS09, GWL09a, GLW09, GW09b, HNT03, HPP00, Joh05, gLxW02, Rob05]. Classification [BS01, FS07, Živ06, Est05, Han03, NT08d, Val02b]. classified [Kir05]. Classifying [DMD03, BLS05, BBL^{+}05]. clean [YZ07]. Clifford [BKU06, Pap00]. clique [AJF09]. cliques [AGM09, Row07]. close [BMS07, EH03, WHAB08]. close-to-Schwarz [EH03]. closed [BPS07, Dau08, DNR04, EdRM04, GS09c, Hon02, Hon06, LR05c, Mir03, Val08, WZ07]. closed-shells [Dau08]. closer [CL03c]. closure [Che08, Dug04a, Dug05a]. closures [BB03b, Ser07]. Cluster [Ovt06a, Ovt06b, Ovt06c, DW05b]. clustered [AdLV06]. clustering [FT00, SK08a, VWD08]. Co [Row07, Wan09b, AGM09, Sav03, Sav04, Sav05, Sav06b]. co- [AGM09]. Co-cliques [Row07]. Co-maximal [Wan09b]. co-order [Sav03, Sav04, Sav05, Sav06b]. cocharacter [BC00]. cocharacters [Mat04]. Cochran [Leš04]. cocircuits [Lem04]. cocirculations [Kar05b]. cocycles [HN09, ZFW06]. Cocyclic [MW09]. code [HK02a, Wu06c]. codes [CG00a, CHP07, CHP08, FZ04, GL05a, GLS07b, HST08, JNL05, Jai06, JN06, MOS07, Par02b, PR04c, Rvz09, VR06]. codeterminants [Doo02]. codimension [GZ07]. coding [CKZ05, FP04, Jai07]. Coefficient [SSSG08, Brü09b, Car04, CC03, JL03, Tre09a]. Coefficients [AK09a, AH04, BG09b, BD01b, BD06b, Ful00, II09, Ili09b, LRR04, Moh07, Pit09, SI09, TMARAE^{+}07, ZLC09]. Cognitive [Per03]. cograph [CHY08]. Coherence [CL02b]. coherent [KMMZ04]. cohomology [GR03, Wym02, Zus05]. Coimbra [dOQSZ06, dSQS00]. Colin [Gol09]. collapsing [PL08]. Collatz [AGDR05, TS03]. collections [Yah08]. collocation [CP03, KH01a]. colored [GS05, GS09a, HL08a]. coloring [SG08]. Colorings [Dia02]. Column [BSKS05, BGJS06, BBD^{+}08, BDD^{+}01, Dax00, HV08, Joc09a, Joc09b, JS00b, MW03b, Ste09]. columns [CSZ01, Hei02b, Ned01a, SLP01]. columnwise [BB07a]. combination [HGC07, Rab04]. combinations [BB00, BBS02a, Bak04c, BB04b, BB04a, BBÖ04, BBG06, BB07b, BT05b, BT05a, KR06]. Combinatorial [CFN04, CT04, vDS05, vdD09, BSW06, CK02, dST06, dSQS00, HKSS03]. combinatorics [But03]. come [FMM01]. coming [NCV08]. Commensurability [JKRT02]. Comment [XLG^{+}13]. Commentary [Bar05]. comments [BS05a]. Common [LS07a, LŠ09a, Val08, Xu08, And04c, AP04, CL03c, DTGV06, EMT05, KM04, KN06a, KS06a, LS07c, MSS07, Sci09, Tsa01, vdW03]. communication [BSL04, CCM00, HLW06, YFK05]. communications [DPZ06, SSW05]. commutant [Bra07, KR05]. commutants [Bra07]. Commutative [GM09b, OZ03b, OZ03c, Bal00, BBC00, CW00a, Cao01, CT03a, CZ05, CHASG02, CKK08, CHP03, Don01, HL07a, HB08, KVV09, LM09b, NDS02, NY02, OWY07, PH04, Pro05, RGG01, WY06b, WY06a, WYZ07, Wan09b, WL09a, XCX01]. Commutativity [BB02a, CCH02a, Lin03, Šem08, BK06a, CI09c, LL08c, Pet02, PS03, TTM02]. Commutativity-preserving [Lin03]. Commutator [Hir09, BW05, BW08b, Hoo07, Lás07]. Commutators [Nie09, Bot06a, DRR02, FMS01, FMS05, Kit09, Zhe02, ZY02]. commute [BT05a, GH09]. Commuting [Abd08, AR06a, GS00, AGHM04, AMRR06, AEI00, BBÖ04, BB07b, Bou07, CM00a, Fre04, HO01, KP02, MA06, McC01, Nov09, Oml04, Šiv08, Ste05a, Tic05, Val08, LL08a]. compact [ES01, Had04, Haz05, LS02c, LS08b, Shi01, Yah08, ZZ05]. Compactification [She00]. compactifications [FF04b]. Companion [GW04a, ASK^{+}08, Baz03, Bos02, Bos07, Cal08, Fie03c, GW07, Kes00, Lin02a, TR03, Wim01, Win03]. Comparison [BKDL01, BS07, CM01, LT04a, LS01, MS00a, MN08, Pry09a, SSA05, EvdD04, FGR04, Joc09b, Lin02b, Mar02, Mol05a, SBM05, Stȩ00a]. Comparisons [LEL02, EFN^{+}03, JY07a, JY07b]. compatibility [Mol02]. compensation [BMS04, BMS08a]. complement [AAB^{+}07, BS04a, BZ03, DH02, IK09, JMT04, Kra05, Lu02c, ST08, Sta07, Wat00, yZtLC06]. complementable [MS04c]. Complementarity [GP00a, HT09a, Bal07, CSY01, GSR03, Kak08, MM05, MND01a, PK08, YLX09]. Complementary [DJ06, Fie04, MPS00, Bén02, FG03, Par09a, Seo00]. complementation [LM01, SG09]. complements [BBS04, BN04, BGKR05, CMS02, CRS07a, FN01, Fan02, Gol03, Hog08, LHZ04, LH04, LLHK08, MS04c, Neu00, RJ02, Row07, Zha00, yZXtL07]. Completable [AC03, DJKM00, JTeG09]. Complete [DJL01, ZOE03, BCSV06, But08, CS03a, Hig00, MS03a, OvdDV00, WM02, WM03]. Completely [BX04, BX05, HHY04, Xu04, Auj03, Bar01, BB03a, BAD09, Dam04, DA10, DJKM00, Gol03, LT03b, PR04a, RSW02, YH00a, YH00b]. completely-positive [RSW02]. complètement [BCSV06]. Completing [Spi00, Tia00]. Completion [Cra09, KC06, MHZU05, Zha06, AHW05, AY08, ATU03, ATU04, ATU05a, ATU05b, BEH^{+}06, CDS04, Dod08a, Dod08b, DJKM00, EG04, Fan07b, FKM02, Hog01, Hog02, Hog03, HJR05, Hog07, JS00a, JTU01, JT04b, JTeG09, Kro01, LM04a, MS00b, OB01, QC04, Smi08a]. Completions [DH00, JTU03, OCK09, Alf03, Alf05, ATU06, BW01, FJTU00, GR00, HHW00, HLT04, Hog00, OCK07, Qi09, RDxC03, Urb04, Woe04]. Complex [Kal06, ASK^{+}08, Aga02a, Bér05, BM04c, BV00, BPT09, BGGC06, CHL02, Chi04, Cof03, Dan06, Dra05, EH09, GŞV02, GG09, GH09, HT09b, HDB00, HS06a, HW07b, JLN08, JKLP03, LS07a, Lip01, MP05, MS05, RS03b, Rum03b, SSW05, SS05a, SLR06, Wat06, Yeu04, YLZ06, dSR05, Rod06]. Complexity [BS03a, BLG01, CCM00, FF02, Sat08, dW03]. component [Omi08b]. components [BSD09, LSF09, MD02, Sem07, Ste00b, Ste02a]. Componentwise [MS04a, CLL05]. composed [DKS04]. composite [BRR01, SG06, Yag01]. Composition [Eld02b, Abd05, BKLM06, DJ01, Eld00, FQ04, HN04, KYAA08, Mat01, PW04]. compositions [GV09a]. compounds [Pat08a]. compression [Aud06, Aud08, CKŻ06, GW04c, LMV03]. compressions [AM02, Bou03a, Lim07a]. Computation [BFH04, BMP00, CR02b, Fer07b, Ser04, AVV05, BR06, Bor08, BDKW04, DMMV09, Hig00, KPK04, KHG04, KM04, Mae08, Ois01, Pan00, SS03a, ZJL02, Par09b]. Computational [Cen06, Rum01a, BLG01, FG02, HHM09, Kol05, WSW06]. computationally [FK04]. Computations [BGM01, Moh05, BB03c, Kra03a, PZ00, WWQ04]. compute [DP04, vdW01]. Computer [Plu01, CFM03]. Computer-assisted [Plu01]. Computers [HM02, Vin00]. Computing [CGGGSR01, DTGV06, Eic04, FI03, RS06a, Roj09a, VLA04, Boj03, CD00b, Dri04, GT06a, KYAA08, Kri05, MM03, MVV08, MSW09, OS00, PS07, Sch09c, Wil07, Wu06a]. comtrans [IS08, IS09]. concatenated [CHP07, CHP08]. Concave [Kar05b, DH03]. concavity [Bou06a, GRB04, Wan03a]. concept [Bak04b, LMMT01]. conceptual [Uhl03]. concerning [AKYD16, Bot06a, COvdD09, Che08, Che04b, Che07c, CL00a, Dit00, Kar05a, Mah07, SN09]. Concrete [fur08b, SvW05]. concurrences [Bag06]. condensation [Abe08]. Condensed [Ste01a, AEI00, Ji05]. condiagonalizable [Ikr07b]. Condition [GW03a, MS03a, gS02, BS00d, CR08, Gha01, Gus09, HP03a, JO02, LA05, NX04, Nik08b, NCV05, SS07f, TY09, Tis00, Var03]. Conditional [XS04, Udd09]. Conditionally [Mol02, CM09]. conditioned [FH01, KNX07, Kol02, Ots01, PPZ02]. Conditioning [Kir06, Kir04, NCV08, ÖA00]. Conditions [Dru04a, EFN^{+}03, BES09, BS00b, CWW09, CCG01, CHP06, GS04b, JY07b, KV06, KS06a, LS07c, MPS07, Mir03, OP05, Peñ04, QD09, Ste09, SS06e, TW05, VV06, WB01, WB08, XHpS06, Yan03, ZHZ03, Zit09]. cone [Alf06b, BPS09, CC01, Dah03, Dat08, Kal03a, LN06, MM05, Nie00, NT09c, Pat04, ST03a, She03, Tam04, Tar05, YH00b, ZG09]. cone-preserving [Tam04]. Cones [And04a, BPS07, BSS07, Ceg01, CPR03, CHP06, CL00b, CL07b, EMT05, GK09, Hil07a, Hil08a, Hil03, KFGK04, LM06b, LR05c, Lim07a, SSA05, SHS03b, SV06b, Stu00, SG06, TS03, Val08, Vil08]. Conference [Ano03-46, Ano07s, Ano09p, BBVR08, BLL06, BGKW07, IMS^{+}09, Pue01, Uhl04, dOQSZ06, KMS05c, Pac02, Ano04t, Ano05p, HKSS03, LWZ05]. configuration [Don01]. configurations [FMM05]. Confluent [Joh05, YH04b, VS04b, YC00a]. conformal [Abd05, ABR02]. conformation [DW05b]. Congruence [BS06b, BFJ07, CS09a, Cof03, – DI02, FJ03, FJ05, FJL08, HS06a, LR05b]. Congruences [HS04, TW00, Wat06]. congruent [GHS09]. Congruential [FJ04]. conics [Nie04a]. coninvolutory [AMP07]. conjecture [Alt09, AKYD16, AGDR05, BSD09, BH08b, Che07c, Che07a, CL00a, DL09, DHH^{+}07, Dit00, GT02b, Hil07b, KG09, LS09b, LR02, Lás07, LY05, Lie09, LB01, LJ01, NT09b, PX03, SS04a, Wan09d, YYZ05, YC00b, Yon00]. conjectures [BHS06, Kle07, Zha04d]. Conjugacy [Ell04, For04, VLA03, Yan06, Yan07, EL01, NY00, VLAOVL08]. Conjugate [FI08a, ABR02, CI09a, FI07, GH09]. Conjugate-normal [FI08a, CI09a, FI07]. conjugation [LT03a, Tri02]. conjunctivity [Mil03]. connected [BZ02, BBC00, BM04b, BDD^{+}01, CZ05, LRWW07, LLY09, RR04a, SGS08, Ste00b, Ste02a, WD06, YC00b, vdH08b]. connection [CH01, McE02]. Connections [AB05b, Dod07, DS08c]. connectivity [BP05a, BL09c, DWD07, FKP03, GB06, KMNS02, Kir03b, LSS00, Mol06a, PL08, RM06, Roj09a, SGS08, YFL09, Zha07c, Zha09, dA07]. consecutive [De 00, PG01]. consecutively [RvdGM05]. consensus [ZTK09]. consideration [JN06]. considerations [CL01a]. Consimilarity [Lip01]. Consistency [WSL02, XDS08, XDS09]. Consistent [EN09]. Constacyclic [Rvz09]. constant [AFM03, CL00b, Fur06b, JL03, Pit09]. constant-coefficient [JL03]. constants [BD09a, Fen03, FT07, Ton00]. Constrained [Pop08, BL09a, BD03c, Čer09, DW00, GJW02, HHM09, JLMP02, Leš01, NPP00, Nie01a, OF02, pPyHZ05, ZyC03, Sch06]. constraint [Qi09, ST03a, ZHZ08]. constraints [CR07, GZZ00, HG09, LS02d, MOS07, Sch09a, Wu08c]. Constructably [Kir07]. constructed [KMS05b]. Constructing [Bai08, Bor06, BC07, BC08b, CFRS02, dSF09, McC05]. Construction [AFK02, KMZ00, LŠ07b, Sch00a, Yin04, Yin05, Ala05, CFKP08, Han02a, JLY09, OT06, yPyHZ03, Pit03, RS03b, SLP01, SSW05, Sot03, Wal09]. constructions [CT09, SH00a]. Constructive [Uhl01]. contain [GW03a]. containing [BPT09, GJR03, LN08, LP09]. Containment [Lin02a]. contains [Fer07a]. continuation [Reb07]. Continued [RL03, EK08b, RK00, Van04, Zyg02]. Continuity [gLgWmW08, LY09b, Bal07, HMR09, HM06, RA08b]. Continuous [GGL08, GL06, Car09, CFL05b, DSW07a, FLPW07, IV04, Igl01, MA03, MZ08, RY08, RO02, SV07, Xu06, Xu07]. continuous-discrete [RO02]. continuous-time [CFL05b, Igl01, MA03, SV07, Xu06, Xu07]. Contour [IBI09]. contoured [IWL07]. Contractibility [LS02c]. contracting [Wu08a]. contraction [Car04, CC03, LS08b, Lim09c, OS01, Pat08b]. contractions [And04c, Dru08b, GW03b, GW04a, GW09a, HNPT06, Kar09, LS02c, PP09, Val08, WGT09]. Contractive [Reu06, AD08, BB02b, QC04, Rod09, XXZ09]. contractive-valued [BB02b]. contraharmonic [Lim05c]. contravariant [Pat08b]. contribution [CI09b]. Contributions [NS08b, NS08c, Cal05b, GTT01]. Control [Myš06, AAA^{+}02, BBMX07, BF07b, BHRV02, BRS00, But06, DH01, Dam02, FN07, FB05, FH04, FS02, HT02, IK06, Jef05, LV06, MZK00, PGRO06, Sil00, Ste01a, SK08b, TM00, TVB08, Wan07b, Xu06, YWS00]. controlability [CDS01]. Controllability [BRR01, DS08c, AD02a, CCZ06, JTU01, PP02, PPZ05, VZ02]. Controllability-observability [BRR01]. controllable [CG00a, PPZ05]. controlled [GV06a, Ois01, Ots01]. controller [Sch02a, YWS00]. converge [AY03, DL92, DL01]. Convergence [BW03, CW00b, Fri06, Fur05a, LWZ09, MM03, Mat02, RK00, VWJ08, Wan01, WB08, Yan02a, AM04, BC04a, CKMS02, CHP06, KN03b, Leš01, LZ09c, LWWZ08, LMT09, MS00a, MP06, MH06, Ney06, Pry09b, Pul07, Til03, XZ06, ZBY09, dMR07b, Ney01b]. convergent [She00, WB01, Yag01, ZZY06]. convergents [EK08b]. converse [Alz01, BT07b, KL07a, Mal01a]. converting [CD06]. Convex [CL07b, Han02b, LRY05, ATV09, AS03, AB07, BKDL01, BF05, BB05c, BDD^{+}01, CL00b, Dax06, DB06, GK07, GB06, HT07a, JP04, MOSZ02, Nar05, NS08b, NS08c, OT09, Pue07, Sch02a, Tik06]. Convexity [Bae07, GJK04, MLR00, PR02c, Bek04, Bul08, GRB04]. Convexoid [Bar02]. convolutional [CHP07, CHP08, FZ04, FP04, GL05a, GLS07b, HST08, VZ02]. convolutions [Sev08]. cooperative [XDS08]. copies [Roj07c, Roj09b]. copositive [Bom08, HJR05, Hog07, JR05, Kap00, MND01b, YH00a]. copositivity [Bom08, BD08, Kap01, jYxL09]. Copson [CW07a, CW07b, CKW08]. corank [vdH08a]. core [CFJ01, HS07a, TBK09]. corners [KBdP07]. corona [Tre04f]. correcting [JN06, Par02b]. correction [AB05b, LSS00, PR04c, Pul07, RT01a]. correlated [AFL00, MW08b, Sen04]. Correlation [SS06a, BP03, GH00a, GP07, Gus09, KP08, KC03, KC06, Ruk01, SS07b, SvKHO09, ZW03]. correlations [Dru02, Kes00, Kes07, OO00, Wat01a]. correspondence [Bén02, CC06, GHI09, GvdV04, vdVN00]. corresponding [AMLM^{+}09, BS00d, DHZ03b, HN09, MM08, SS03a]. Corrigendum [DL01, JS07b]. Corrigendum/addendum [DL01]. Cosine [NCCY00, PT06a, PT05]. cosines [Pas01]. Cospectral [vDHK07, FJ05, GRTR08, Ste00b]. cospectral-congruence [FJ05]. Cospectrality [DJ01]. cost [BF07b, PGRO06, YWS00]. cost/reward [BF07b]. could [RS05c]. countable [Yan09a]. Counterexample [LR02, BHL09, Fur05b, KR08, Wan09d]. Counterexamples [EG02, Fur05b]. counting [CdFM09, Han07, KB02a]. Coupled [HZ06, KAP^{+}03, SS06e]. Coupling [Hun09, FNV08]. Courant [BNdP09]. course [Okt04]. covariance [CP06, CM09, HB00, MD02, Nah06, SvR02, VC04]. covariant [Pat08b]. cover [BFH04, BFH05]. Covering [Fer02, CHY05, Fer01a, KN06b, Sat08]. coverings [DKX03, KS06e]. covers [KO08]. Coxeter [DL03, JKRT02, KKABAF06, Lad08, LdlP09, MOSZ02, Sat05, dlP03]. cp [BB03a, SM09a, BR06, LT03b]. cp-rank [BB03a, SM09a, BR06]. Craig [CL09a, Li00a, Mat03]. Craig-Sakamoto [Li00a, Mat03]. Cramer [GAE02, HL04b, Ji05]. CRC [vdD09]. created [LM06a]. Creation [HP00]. criteria [AI00, BMS08a, DKX03, GH03, Kol03b, KFGK04, Peñ01, SBM05]. criterion [AH08, Bom08, Had04, KKP08, Sav06a, Sot06, Tsa01, WLL07, Zho03]. Critical [Cal07, DG03, AA09, Bai03, BM04b, Brü03c, CR02a, LPW08, Sem07, dST06, dSR06]. crossed [Ala00]. crosspolytopes [Pac02]. crystal [BM01b]. CSL [ZPY06]. cube [Bai03, BBL^{+}05, BPT09]. cubes [Ost03, Pac02]. cubic [KH01a]. Cullis [NY07a]. current [Zus05]. Curriculum [Ano06h, Ant07, FGH^{+}06]. curvature [And07a, CSSZ01, DRGR04, Dit00, FM04b, SS07b]. curve [BZ06, DMMV09, ZZ00, ZGL09]. Curves [OZ03a, AMLM^{+}09, Jøn09, MBLV01, MS05, Var06]. Customizable [Hao04]. cut [LLT04b, Mol08b, BHL09]. cuts [Chu07]. Cvetković [vdD09]. cycle [BS04a, BOvdD07, DJKM00, Fer07a, HWC06, OB01]. cycles [GS05, Gew03, HL08a, LW03, Nik08b, Ser04]. Cyclic [Pue07, CPR03, Kal06, Li03, LMT09, Ser09]. cyclicity [RSS00]. Cyclizable [SS04a]. cyclization [SS07a]. cylindrical [DV01].

D [FVP06, SLP01, Yag01]. D. [LS09b]. DAEs [BRM01]. Damir [Rod09]. damped [Bai08]. Darboux [BM04e]. Data [Bör07, BIL06, CH00a, CH01, FNT05, FNC06, GS04b, IL06, IBI09, Nie04a, PPP06, PA04, Woe02, Yin04, Yin05]. Data-sparse [Bör07]. David [Sun08]. Davidson [BSS09, FS08, HS03, Hoc09, Vos07]. Davis [CN02a, CN09]. Davis-Wielandt [CN02a]. death [vDP08]. deblurring [CCSS03, MLK^{+}04]. December [Sch02b]. decide [BP02]. deciding [JPB08]. deck [Sci09]. decomposability [LSF09, Liv04]. Decomposable [LZ00, dSR06, BF07a, Lim08a, YH00a, YLS05, dCdS05, dSdC05]. Decomposing [SSMQ^{+}00, CQ08, CM09]. Decomposition [BJT04, CL02c, KRS03, WY04, WY06b, Zhe02, BST09, BW00, BB05c, Bos07, Bra06, CHL02, CD00d, DoMP09, De 07a, DV01, DM07, FI08b, HMP09, HT07b, IPP09, JHL05, JO05, KSU03, LV06, LY09a, MPP10, Nom05a, NT07c, PRS02, PPZ02, SSSG08, SAAF08, SS09a, Sch09a, Sch00a, Sla03, Smi08b, Sou05, SS07f, TR02, Wat04a, Whi00, Xu03, sV06a]. Decompositions [CP08, AB03b, BK00a, DTGV06, DK08b, HPP05]. deconvolution [CW00b]. decoupling [CH00b, CFP09]. decreasing [Kuz02b, Lim06, RR04b]. decreasingly [Mou06]. dedicated [Ano07u, BKMZ07, FMN04]. Dedication [BN00, BLOT02, DMM^{+}05]. Defect [TŻ08, BK09b, Kra03b, McG06, SdSW07]. defective [Ala05]. deficient [DV07a, DV08, Yua00]. define [Lom09]. defined [DH09, Kes07, OCK09, WY07]. Definite [And07b, BGN08, CMP03, CC09, CR08, CX05, CM08b, DS08b, DL08, FF02, Gro00a, Gro00b, Gro01, GV06b, Has05, HES06, HP09, HTV02, HZY05, JS00a, KL07a, KC03, LL08a, LM06b, LG03, MVV08, O’L03, OB01, PVV08, Ser07, Smi08a, Var03, WB01, XS08, Zha03b, ZyC03, vdH03b]. definition [Tik06, YTI06]. Deflated [DMW08]. deflation [DM03]. deformation [Far08]. deformations [CFP03b, CHPP06, FP01, Mai01, PPT04, Pue05]. degenerate [EdRM04, FKRS09, Sen04]. Degeneration [PW08a]. degree [BFP07, Bir04, BL09c, BG06, Cal07, Cao09, CSWM01, Chl06, GW02, GM07a, II09, KPP07, LM09a, Mue09, MS02b, Nik06a, Nik08a, OS03, Ste03a, TFZ08, YL08]. degrees [CD00a, Nik06b, Pre05]. delay [CP05, FS02, JH09, MGWN06, OTC03]. delay-differential [JH09]. delays [Sun02, Yer04]. deleted [SS08a]. deletion [DS08a]. Delsarte [Luz07]. Denjoy [LN06]. dense [BGGC06]. Denseness [EKS01]. densities [BDKW04, HO02, Mat05a]. density [BM01b, Chi02, FKP03, JN06, MT06, PR02c]. dependence [LDS07, Wal05a]. dependent [CŠ08b, MŠ04e, OP05, RvdGM05]. depending [BFPW02]. Derivable [ZX07, AH09a, HQ08, Wu09, ZX05b, Zhu07, ZXZ08, ZXZ09]. derivation [CC03, Fer01b, HQ09, KN00b, TX09, WL09a, dB00]. Derivations [Dun07, OWY07, WY06a, Ali09, AH09a, Bar02, Ben05, Cal05a, Cal07, CGN01, CLY07, DG02, DG03, Dug00, Eld00, Har00b, JW05, JGPI08, Kah04, LL09b, Lu09, QH09, ZJC04, ZY06, ZPY06, ZWC06, Zho00, ZX05a, dSG07]. derivatives [AD04, BJ09a, DH03, Dun04, Sen07]. derived [Drä03, Kra03a]. Desarguesian [Kes00, Kes07]. descent [BS09a, Cao07b, SdSW07]. described [FH07a]. Description [Del01, DMM00a, SvW05]. Descriptions [Sha09, BFPW02, FP04, Hil08a]. descriptor [BS00b, Brü09b, CCZ06, FS02, HMH01, Vir08]. design [EHR08, GM09a, Kle04, LV06, PWZ05, Sch02a, XM02, CFM03]. designs [ÁFMNW03, ÁFMNW06, ASK^{+}08, Bag06, Cal05b, DGN02, Dru04a, Kob00, LY05, MCEC06, MW08a, MW08b, SSW05, Udd09, vDS05]. Detecting [HTV02]. Detection [SLP01, BD08]. Determinant [Bap06, DP02a, AG00, AFM03, Bri05, CDJ08, DJ06, GHS09, Kra05, Li05, Myu09, NY07a, Nis02, NT06a, OvdDV00, Par02a, Sav06a, SXKM03, TW03, Wen09, Wu08d, YTI06]. Determinantal [Peñ01, PH04, BS05b, Che03, KP01, KBdP07, LSH08, LSF09, Mat03, SS05a, SLR06, SD00]. Determinants [Ano03u, BFL05, CP05, Fan03a, Mol08a, AFR06, BE01, BdPK07, BJ09a, BKLM06, CR02b, DMMV09, EN09, FLPW03, GN03, HK05, Joh05, KH01b, MSS08, RS06a, Sal00, jYzWbZ01]. determination [CS03a, Sta09a]. determine [Aga02a, BD01b, BD06b]. determined [BGO09, Bru02b, CHRS03, DH02, OT07, Sak01, SHZ05, WX06a, WX06b, Wil09, ZZ09, vDH03a]. determining [MZ06, WO05, jYxL09]. deterministic [Gue08]. Deutsch [KN09]. development [Abe08, Bak04a]. developments [Cve02]. deviation [BJ07, Nik06a]. devoted [Ano04t, Ano05-33]. Dewilde [EG01]. Dewilde-van [EG01]. DGMRES [Sid01]. diagonal [ALdV05, AdLV06, Bor06, BC07, BC08b, BW01, BPP08, Bru03b, CMV03, DJ01, Fie01b, Hog00, Hog07, HL04e, JMT05, JLN08, Kah06, Kes07, KN04, LLHK08, MS06, ML02, Mir00, MG09, NBdP08b, NS02, PV06a, Phi04, PES07, PVV08, PP09, SSMQ^{+}00, She03, SMK09, SN09, Tri02, Wan06, dF00, dSZ00b]. diagonal-plus-semiseparable [PVV08]. diagonal-Schur [LLHK08]. diagonalisation [Wir09]. diagonalizability [SG03]. diagonalizable [Bot00, Eis06, Nak03, OV06a, WLBZ06]. diagonalization [ADT06]. Diagonalizing [HPP05]. diagonally [CNX05b, CH07, Dah00, Kol03b, LHZ04, LH04, LLHK08, PP01b, Wan09c, yZtLC06]. diagonals [CT01]. Diagrammatic [Kul01]. diagrams [GPS01]. dialgebras [FLROV05]. Diameter [lHH08, AT07, BFS08, CdFM08b, Guo06c, GS06, Guo07b, LL09a, LLT07, LzZT07, wThY09, YSL08, ZSL09, ZLS09, Zha07c, vD07, vDK07]. diameters [AMRR06]. dichotomy [GS03a]. dictionaries [AEB06]. Difference [Har00a, BFH05, BCS02, BAD09, Bur06, Den09, Di 03, DHZ03b, HZ03, JY07a, JY07b, KN06a, KTKR06, KRS04, KMS05b, Kra01, LŠ09a, MT02, RSW05, RS01, SW09, SS06e, Tre09a, WZCL08]. difference-algebraic [BCS02]. differences [Kit04, Li08d, Wen09]. different [EFN^{+}03, GPMT06, Jia00b, KA06, RS03a]. differentiability [Bae07, ZM08]. Differential [HT07a, Bar08, Cap03, DJMS09, FB05, FH04, GPMSZ06, GJK04, HL05b, JH09, KMZ00, KS09, Lee09b, Lom07b, MT02, McC05, NCV08, OCK07, Plu01, Prz02, TI08, Tre09a, Tre09d]. differential-difference [MT02]. differentiation [Lom07b]. differentiation-invariant [Lom07b]. diffusion [BKK08]. digital [SvdS01]. Digraph [Kir04, MS07, Sat08, Sav05, XL04]. Digraph-based [Kir04]. digraphs [AK09b, ATU04, DOvdD05, DW05a, DSW07b, EH09, FGM06, FM07c, GS05, GS09a, GS09b, Hor07, HL08a, JMN08, JTeG09, KSH08, Kov02, Ma09, MNOvdD08, MZ00, MS01b, MS02a, Rad09, SH00a, Sha00, SG08, SW03, TMARAE^{+}07, WD06, XX09]. dihedral [KK01]. dilation [Han02a, QC04]. dilations [Bou03a, Bou03b, Bou03c]. Dimension [CM08a, CFP04, FM07a, AMLM^{+}09, Arn04, AS09, CFN04, CHP03, Dom03, GH00b, GT09, Prǎ02, SM09b, dST05]. dimensional [Aga02a, AD03b, AGI00, pBlWyXwA08, BKS01, BR02, BLG01, Cap01a, CD05, Cos07, Dom03, Fao01a, FB05, Fri00, GGM^{+}09, GZHT05, GW09b, Har07b, Hil08a, Hol08, JGPI08, Kin05, Kye03, Lom09, MT04, Nak00, NY02, NY07b, O’u02, Sch02a, She03, VW00, Wie01, YY03, Zus05]. Dimensionality [DV04, Sav06a, tBSR04]. dimensionally [Zho00]. dimensions [BO05, Bou07, Buk07, D– D05, FLRZ02, LL02, smbH02]. Direct [EG04, Ben04, Bér05, HV01, HK07, HZ05b, KN00b]. Directed [KMMZ04, TK09, Woj03, BK03, EH09, Jør05, TP09a, Wu05b]. direction [NB03]. directional [LP01a]. directions [TW05, VHK01]. Dirichlet [FQ04]. disaggregation [HW00, MM03, MP06, Pul07]. disc [GW03a]. disconnected [Sud02]. discovery [HM02]. discrepancy [But06, Doe07]. Discrete [DGLS01, LCN03, AFL00, BKDL01, BLNP01, BF07b, Bıy03, BD03b, BKK08, BRS00, Brü09b, BDD^{+}01, Cen01, CY00, CP09, DHZ03b, FL04, FO01, FPP02, FN07, FS04b, GHO02, GZHT05, HT01, HO02, HOP06, HZ03, HZ06, IO01, Kar04, KB02a, KTKR06, KN01, LS09c, Mal09, Ots01, PGRO06, PT05, Pot03, RA08a, RR05, RO02, SC01, Shi06, Val02a, VEB00, Wer09, Xu06, Xu07, Yag01, YWS00, LGH01]. discrete-time [AFL00, BF07b, BRS00, Brü09b, FPP02, FN07, FS04b, HOP06, IO01, KTKR06, Ots01, Xu06, Xu07, YWS00]. discretization [HZ03]. discretizations [CLB09]. discretized [Cap01c, Cap03]. discriminant [Par02a]. discussions [Okt04]. disguised [Kre06]. disjoint [Bak04c]. dispersal [LS06c]. Dispersion [Jia04]. Displacement [YLZ09, BCES05, BM01a, DLZ03, YH04b]. Disproof [YC00b]. Dissection [BKK07]. Dissipative [BPT07, Che02b]. Distance [Bap05, AHW05, Alf00, AW02, Alf03, Alf05, Alf06a, Alf06b, BB07c, BKN05, Bap06, BLP06, BR09, Bén06, Dat08, Dax06, Fio01, Fio05, GH06, HV03, Hol02, Hol08, Ind09, JMT05, KS07a, LT07b, LT09, Mik09, Okt04, PP08, Pas01, PT06a, Peñ03b, Siv09, Tar05, Ter02, Xia06]. distance-education [Okt04]. distance-regular [Mik09, Pas01, PT06a, Ter02]. distance-regularity [Fio05]. distances [BT09b, CGPM01]. distinct [CO09]. distinguished [BDM03]. distributed [GN00, WA09]. Distribution [Bol04, Cap01a, GMR^{+}07, BM01b, CM01, DV07b, FKP00, GL05a, IWL07, Kir05, MPS06, Olk02, SCS09, Tre02, Tre03b, Tre03a, Tre03c, yZXtL07]. distributions [CM09, GHO02, Hun05, IWL07, MP05]. distributive [Tan01, Tan03a, Tan05a]. disturbance [CH00b, CFP09]. divergence [FS06b]. divergent [APESS06]. Divisibility [Hon08, BK07, KH08]. Division [WY05, HC06, HL07c, HL08b, HL08c, JGPI08, RRY04, WZCL08, Yah06]. divisor [Kat04]. divisors [DTGV06, HC03a, Hou03]. DNA [FNC06]. Dobrushin [KN09]. Dodgson [Abe08]. does [Guo05c, WHBM09b]. dog [Bar01]. dog-ears [Bar01]. Dolezal [Gra04a]. Domain [Whi00, Bıy03, DGLS01, Yag01]. domains [ALC06, BHL^{+}04, DH06, HAT03, Kat04, Rao09]. dominance [MG09]. dominant [CNX05b, CH07, Dah00, Hor02, Kol03b, LHZ04, LH04, LLHK08, PP01b, Wan09c, yZtLC06]. domination [FYL06, LLT05, SAH08]. d’opérateurs [Del01]. Dose [ZM08, CCW^{+}08]. Dose-volume-based [ZM08]. Double [CNSM07, OT09, CHN06, JDS03a, McM09, PJ07, ZTK09]. double-integrator [ZTK09]. doubles [MT04]. doubling [CFL05b, HL09]. Doubly [HP01a, ABH03, ATU05a, BD03c, BAD09, Dah03, Dah04, DJKM00, Fan10, HP04c, HL04e, JMT05, LT04b, LHZ04, MMX04, Mou06, Nie04b, PV06b, SH02, XLG^{+}13, Zha05d, Zha09]. down [CP06]. dqds [PM00]. Dragos [vdD09]. Drazin [Bu05, CGD05, CXW09, Cui07, CIDW06, CI08, CI09c, DW09a, Den09, DW09b, DD05, GKW00, Gon05, HWW01, KKP07, LW07, gLgWmW08, PP04a, PS07, PG04, RW02, Sid01, SD00, Wan02, Zha01, ZZJ08]. Drazin-inverse [Sid01]. Driessche [Ano04u]. driven [IL06]. Dual [LW00, Alf08, BS02b, BFP07, De 07a, De 08, De 09b, gLxW02, Lom09, SG06, TK09, Tup08b, Tup09, XDS09, dSS08]. Duality [CGQ04, PW08b, EGMG07, Hol00, IBI09, Woo02]. duals [CJPS02, LM06b]. due [DS08a, FFN^{+}07]. dumbbell [WHBM09a]. Dunkl [PR07]. duplicate [Qui07]. Dym [Rod09]. Dynamic [ALC06, ALCT05, DHZ03a, GPFC04, HAT03, KS06a, SdL03, Sun02, Wu06a]. Dynamical [AFR06, ACK05, Bad08, BS06a, BKS01, BK01, ÇS02a, Fre08, Gug08, PV06a, Pen06, RW06]. Dynamics [Bon09, Daw02, DW05b, Igl01, OCK07, ZTK09].

each [Ter01b]. Early [Hig08]. ears [Bar01]. Eaton [NT01]. EBL [TMARAE^{+}07]. echelon [CRU09]. ecology [Log05]. econometrics [Gus02]. Edge [AGO09, BM04b, DS08a, FKP03, HD03, dW03]. edge-isoperimetric [HD03]. edges [Guo06a, KSH07, LLT04b, PL08]. edited [Bar05]. editor [Hof08, Bru00, Bru02a, Bru03a, Bru04, Bru05, Bru07, Bru09a]. editor-in-chief [Hof08, Bru00, Bru02a, Bru03a, Bru04, Bru05, Bru07, Bru09a]. Editorial [Ano02t, Ano02u, Ano02v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano03-28, Ano03-29, Ano03-30, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano04-37, Ano04-38, Ano04-39, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30]. Editors [Ano03-40, Ano03-41, Ano06i, Ano06j, Ano06k, Ano06l, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano03-33, Ano03-37, Ano02w, Ano03-32, Ano03-35, Ano03-36, Ano03-38, Ano03-39, Ano03-34, Ano04-40, Ano04-41]. education [Okt04]. Effect [AB03a, Guo06a, Kim05, PL08]. effective [RP05]. effects [Mol03]. efficiency [ACDR04]. Efficient [CEK^{+}01, GP07, JPB08, LMV03, MCEC06, FK04, WO05]. Efficiently [Sch09c, BC04b]. eigen [CKMS02, CKM05]. eigen-problem [CKMS02]. eigen-structure [CKM05]. eigenbases [NT07b]. eigenbasis [Ter01b]. eigencomputations [CCTT00]. eigendata [Bai08]. eigendecomposition [CG00b]. eigendecompositions [AB05a]. Eigenfunction [Gri00]. eigenfunctions [Gri04]. Eigenmatrices [MA06]. eigenmatrix [Zem06]. eigenpairs [yPyHZ03, SvdE03]. eigenproblem [LgS01, pPyHZ05, PVV08, Tan03a]. eigenproblems [BK06b, FMM01, Tre04d, Vos07, XHpS06]. eigenprojections [GKW00]. eigensharp [GM08a, Wat00]. eigensolvers [Ney09, Ovt06c]. eigenspace [Gav02, PT03, Tam04]. eigenspaces [PY09, Ser07]. Eigenstructure [EGO05a]. eigensystem [Hol02]. eigensystems [CLB09, DMM00b]. eigentriplets [Wu06a]. Eigenvalue [AB07, CLL07, HSZ08, JY07a, JY07b, JKO04, MSF05, AH03, BL09a, BPV00, BS00a, BCRS08a, BCRS08b, Bén02, BLLS06, BO05, BV07, BNV05, BMS04, BMS08a, BSS09, BHS06, CKLX09, CCRS08, Che04a, CLOvdD01, CHMM06, CO09, CP09, ĆDR01, Cve02, Das04a, DB05, DW08, DHH^{+}06, DT03, DN00, DS03b, ELN04, Fan10, Fer07a, GGL08, GZZ00, GL06, GHM03, GV09b, GT02b, GT04, Guo04, Guo05c, Guo05b, Guo06b, HV03, HTV02, HT02, HP03a, Hol05, Hor02, HvP05, HP04c, JH09, JDS03a, JDS08, KN03b, Kol08, Lan08, LP01c, LHSL07, LCW09, LB00, LZS09, LM04b, LMT09, LQ00, MPS07, MVV08, Ose08a, PP08, Peñ09, PHZ06, PBA09, Phi04, PES07, RR04b, Roj05, Roj06b, RR07a, Roj07b, Roj07a, Sci09, SCS09, SHWR02, ST00, Sla03, Šmi04, Sol06, SBM05, SR06, Sot06, Sta07, Sta09b]. eigenvalue [Ste03a, Tis00, TMARAE^{+}07, Van07, WV08, XLG^{+}13, YFL09, Zem06, ZZY06, Zha04e, yZXtL07, ZHZ03, xZxGpL09, vdVN00, vdW01]. Eigenvalues [AG05a, BT09a, BH05, Buc06, Cio06, CS00b, CS03c, Ful00, Nik06a, Nik06b, Nik07d, Rum06, SS06e, AA09, AdLV06, AB03b, AB05a, AKYD16, BMNA08, BI03, BZ00, BS09b, Bol04, Bra03, BH08b, CSSZ01, CH03, CD00b, Che07a, CDS04, wCL04, CO09, CG07a, CFRS02, cD03, Das04b, Dru02, Dun04, Eis06, EH07, EvdD08, FL03, FWG08, Fan07a, FYL06, Fer03, Fer07b, Fur08a, Gol06, Gra05, Gra02, Gra04b, GG07, HT03, HL02, HLZ09, HS01, HW07b, IHM08, Jia05, JD02, JDS03a, JDS^{+}03b, Kel08, KH01a, KD07, Kru04, Li03, LL05a, LSC09b, LG06, Lin02b, Lip05, LHF06, LS09c, Mal01a, Mal09, MS03b, MM08, McM09, Mon08, MZ06, NN01, Nie08, Nik07a, Nik07b, Ois01, Ovt06b, Pan02, PBT04, Pry09a, RSR00, RP03, RR07a, Rum01a, ST03a, ySZyY06, Ste07a, Ste03b, Ste04, THZ09, Til03]. eigenvalues [Toy09, TZ01, Ves00, Wal05b, WLBZ06, Wu05a, WZCL08, WSL09, XD03, Yam01, ZL03, Zha04f, Zha05c, Zho08, ZWY00]. eigenvector [Bap03b, But08, Das09, Hor02, PR00, Ye08, ZH02]. Eigenvectors [Zha05c, BMP00, DP04, DV00, Dun04, Fer03, HLG04, SS09a]. elaborate [CM09]. Elasticities [Kat04]. elasticity [CGGGSR01]. electrical [RM09b]. electromagnetic [BGGC06]. electromagnetics [WSW06]. electron [BM01b]. electron-density [BM01b]. element [BZ06, BKK08, FGWG07, Fer01a, NT08c, ST07]. Elementary [Hou03, Tur00, Bak04a, DKX03, Dug01, Dug04a, Dug04b, Dug05a, Dug08, FKM02, Haz05, Joc09a, Joc09b, KO08, LJ04, LL05b, LL08b, MZ04, Nör03, Sed01, Sed07, Tre04b, fur08b]. elements [BPP08, CNR00, GKvS04, LP00, Mir00, Mol06b, MDK09, MD09, NY00, Ros05, Tri02, XJ08]. eliminated [BT00]. Elimination [GML09, ACHR01, ACDR04, CP07c, GT02a, Gem08, KMS00, Yeu04]. ellipsoid [BNT05]. ellipsoidal [ZGL09]. Elliptic [Cal08, HV03, Fie01b, GW03a, HTV02, KH01a, Li01, NN01, NCV08, Plu01, XJ08]. elliptical [Liu02a]. elliptically [IWL07]. Elsner [LR02, Ste04]. Elsner-like [Ste04]. embeddable [De 09a]. Embedded [Har06]. embedding [De 07a, De 08]. embeddings [De 09a, De 09b]. Embree [Böt06b]. Embry [JKLP03]. emission [AG05b]. emphasis [FMN04]. employed [Gri00, Gri04]. enclosure [NN01, Ois01, Plu01, Toy09]. endomorphism [Lip08]. endomorphisms [CHPP06, OZ03b, OZ03c]. endpoint [Ter02]. energies [Wan09e, WK09]. energy [AG08, AGO09, Bal04, BS08, DS08a, GZ06, Gut08, GKMZ09, Hae08, HGW02, HW07a, Ili09a, Ind09, KTR08, LZ07, LL09a, Nik08a, RT03, Rad09, Shp06, YL06]. engineering [HFR04]. Enhancements [Kom08]. Enriched [CRS03]. ensuring [Bén06]. entangled [Ter01a]. entire [Gil04, Gil08, GKL07]. entries [CL09b, CHRS03, Dug08, Hog07, Hor02, Kir06, LR00, MSS08, NBdP08b, Nou06, PR00, Phi04, SHS03a, TRH01, Zha07a, ZH02]. entropies [Fur05c]. entropy [BLdP05, Bha03b, FYK05, Fur06b, HOZ01, Okn01, YKF05]. entry [Das09]. entrywise [FNS07, Hia09, KNR05]. Enumeration [Gew03, MW03b]. Enumerator’ [Mül03]. envelope [BGM08]. EP [CT03c, DKP08, Leš00, MDK09]. Epidemic [PSvdW05]. equal [BSD09, FKP03, GKW00, KMNS02, MW03b, PBT04, Tre03b, Tre03a]. equalities [AR06b, BDK08, CHL02, CLL07, LP01b, TS01, TS09]. Equality [Dat08, dCdS05, dSdC05, CP06, DW00, Dia02, GA06a, KN09, MZ08, Nie04b, PR07]. equally [Hol02]. equals [BFS08, Sta07]. equation [Bak04a, BKR06, BT04b, Bur05, CC09, CL03a, CX05, CNP02, DH01, DSW07a, Dob01, Dra07, DL08, DLT08, FPP01, FB05, GRB04, GHS06, GZHT05, Gro00b, Guo02, HES06, Ikr07a, Ili08, JKL09, KS01, KS05, KTKR06, LS07a, LS07c, Li06b, LZ09b, pLLfY06, Lim07b, Lim09c, LTT09, MHZ05, NZ06, yPyH03, hPyHZ07, RR02, RR04a, Sch03, SA02, Sty06, gSX03, gS03b, TVB08, Wan04, Wer09, WJL02, Xu01, YLF09, Zha03b, ZyC03, ZD05]. Equations [DK08a, DOS01, SS07e, AHLR09, AYC06, BLM02, BLM03, BILM06, BMP00, BKO01, BCS02, Cap03, Cec00, CCG02, CFL05b, CP05, CM08b, DJMS09, DvdV00, ESP05, FYY09, FLROV05, FG08, Fre02, FH04, GRV06, GPMSZ06, GKL07, GX09b, Guo07a, Har00a, HIU04, Has05, HI06, HP07, Hua00a, HL09, IHM01, JH09, JR06, JY07a, JY07b, JW03, KH01a, KAP^{+}03, Kra01, LL06a, LL08a, Lim08b, Lim09b, LWZ09, LG03, MM06a, MT02, MM06b, McC05, Mei06, Mul05, Myš05, Myš06, Ned01a, Neu01, NCV08, OCK07, Plu01, Prz02, Reu06, RS01, SS07d, Ste05b, Sty02a, Sty02b, gS02, SS06e, Tab01, TI08, TS03, Tov00, Tre09a, VB09a, WSL02, Wan04, WL09b, WvdWC09, Xu07, Xu08, XSG08, vdW03]. Equiangular [BPT09, Ren07, Kal06, Str08, STDH07, Wal09]. equidistribution [TZ03]. Equilibrated [Fie01a]. equilibrium [Gra04b, Ost06]. equipped [Zav03]. Equitable [Ter03]. Equivalence [BZ05, Fen03, FT07, PP09, RA08b, Ton00, ADT06, BFZ04, BS00b, CM08a, Fer04, FP07, FJ02, GPMT06, GLS07b, Had05, HAT03, JS04, LR05b, Lom02, Mil03, Tad06, TP09b, Xu06, Zob00]. equivalent [ALCT05, MR02b]. erasures [BP05b, HP04a, Kal06, Mas09]. ergodic [GAM06]. ergodicity [AM09, AK09a]. Erratum [Ano03-31, EFN10, HL08b, Yan07]. Error [Baz03, LS09c, AS02, Gha01, HLW06, HP03a, Kos00, Ovt06a, Ovt06b, Par02b, PR04c, RS02, Rum01a, SS00, SS03c, Ste02b, Tis00, Yam01]. errors [CCTT00, PA04, Sen04, Udd09]. essential [BBS08, Cao07a]. essentially [NT09c]. Estimate [BS03b, Ho03, KN03b, LS09c, MH06, Mat05b]. Estimates [CNR00, HLS00, AC09, BS00d, HIU04, HI06, HB00, LD02, MT05, Ney01b, OTZ09, Ovt06a, Ovt06b, YLF09, Ovt06a, Ovt06b]. Estimating [CGPM01, dlPGR07]. Estimation [GP00b, MPS06, Mar02, AL09a, AG05b, CL04a, Chi02, JSWT05, Sch06, Vol04, WA09, WZCL08, Yer04, ZZ00]. estimations [TP09b]. estimator [PRP02]. estimators [Bak04a, GM04, HG09, HDA09, IPP07, LS02d, MD02, Wu08c, ZGL09]. Estrada [DL09, dlPGR07]. Étude [Fao01b]. euclidean [BM01b, AHW05, Alf00, AW02, Alf03, Alf05, Alf06a, Alf06b, Bae07, BB07c, Bal07, Bén06, Dat08, Dra06, Fie03b, GST04, GTM09, JMT05, KS07a, MG09, SdSW05, Str06, Tar05, ZG09, Zip03]. Euler [Lad08]. Evaluation [KM07, Nis02, AEEG08]. Even [Wu06c, BOvdD07, CW04, EV05, HP04b, Kun06, Liu03b, Mel01]. even-odd [Mel01]. even-weight [Kun06]. Eventually [NM04, HP01a, LM06a, ZM03]. Every [Rab04, Hei01b]. evolution [Daw02, Gue08]. evolutionary [Ost06]. Exact [CLB09, Dra07, RS06b, AHW05, CM09, Fer07b, GA06a, KMZ00, KN00b, NN01, OP05, VLAOVL08, VM04]. exactly [FGWG07, Fer07a, HLZ09, Peñ09]. Exactness [Hil07a, DS08b]. example [AAA^{+}02, BBM02]. examples [fur08b]. exceed [Guo05c, WHBM09b]. exceeding [Lim05a]. exceeds [Gut08]. exceptional [Cve02, CRS07a]. exchange [Arn04]. exchangeability [FNT05]. excluding [WX06a]. Existence [Brü09b, ÇS02a, LPR04, RS03b, Sot03, AMZ06, AC04b, CPR03, CSY01, DL08, KN06a, KS06a, LS07c, Pan00, STDH07, Tsa01, TW05, YC00b, vdW03]. expanded [BRR01]. expanding [DR05]. expansion [FN03, RL03, Ye08]. expansions [Gin09, LS02a]. expansive [RA08b]. expected [CZ08]. experimental [Abe08, Dru04a, MCEC06]. experiments [CSK01, Kle04, Mar02, Stȩ00a]. explain [BM01b]. Explicit [EF07, Ji05, OP00, dFP01, AT00, BR07, CT09, Hua00a, KS06c, Lin08a, RR07a, TR02, ZD05]. Exploiting [RM09b]. exploration [DKS04]. exponent [CW03a, DOvdD05, FL07a, KSH05, KN00a, Kir00, LYY07, MNOvdD08, MZ00, Sch02c]. Exponential [OS00, Bha03a, DP00, Dru08c, HL04d, RS00, So04, TR02, TR03, ZW00]. exponentially [Ves03]. exponentials [Bou07, CM00b, CL01b]. Exponents [GS05, HL08a, OSvdD02, SS03b, SG08, GS09a, KSH07, Liu03a, PP09, SW03]. expression [KS06c]. expressions [Ji05, Tia02]. extended [AKL07, Gus00, Had04, JR09, KS09, KN06b, Nar05, Pop08, TY07, TY09]. Extension [MRR07, Min01, Pop01, AFL00, Bat06, BdPK07, BS04b, DH09, LWD09, Nie07a, Woo02, Yam06, Zav08]. extensions [Ahm05, Del01, FH07b, Fur04, GR03, GQ09a, GS09c, HBHS05, JK03, JW08, Joh05, Mal01a, Mos09, PZ00, SMK09, Yua00]. exterior [De 09c]. external [CS02b]. extinction [HLR08]. Extraction [Cho04]. Extrapolated [HT09a]. extrapolation [BZ03, Had05, HT08a, HT09b, JMT04]. extrapolatory [HMS01]. extrema [EF07, Ney01a]. Extremal [Das07, FKS09, LM06b, PES07, Pro08, Sak04, Zha07a, Bar08, BPZ08, Che07a, CL08b, DK07, Gol06, GZ08, JSWT05, LYY07, Nik06b, Row07, Wir02, Wu05a, dSS08]. extremals [BSS07]. Extreme [Gri04, HH09, Pat04, BLS02, BM04b, GRB04, GL06, HI06, KP08, Mir00, Til03, Zha03a]. Extremes [SHRC03, BGJS06].

Face [CdFM09]. Faces

[Tar05, Alf06b, Pat04]. facets [HP04b].

Facial [Kye03]. factor [BCPT05, CL09d, CI09c, FS07, HB00, MS00a, RZ01, ZYP05, ZFW06, ZZ08a]. factorability [Smi01b]. Factorial [CK02, Kob00, dMFZ09]. factorials [FK07]. Factoring [CQ08]. factorised [Ves00]. Factorization [ADL04, Gin09, Jøn09, LX08, McC01, ADL06, AMZ09, AT08, Bal00, BB03c, Bot09, BJ07, CM00b, CRU09, CFM03, CHNP04, CD00d, DvdV00, ES02, FM00a, FS06b, GT02a, GVD08, Hig00, HV08, Hon02, Kar05a, KY00, MP00, OP00, Pat03, PPP06, SS00, SS03c, SvKHO09, Ste06a, VWD08, YH04a, vdMSR01]. Factorizations [CK03, DKP08, Adu02, ABH07, DJM06, Fie08, GT08, Hao04, KVV09, LPR04, MG03, NS03, OS09, Pan00, PR02b, Reb07, SH05, WW08a, Yan09b]. factors [Bol01, CD00d, FKM02, Ha02, Har07a, Li08c, MZ05, PR02a, PR02b, RZ05, SSMQ^{+}00, Sil00]. failures [TM00]. faithful [Dom03]. Families [Fon00, BS03b, EE04, FM07c, GPS01, GZ08, KMS05b, LL02, Pet08, Tre03a, YZ07, ZW00]. family [AC04a, CFKP08, CKZ05, GZ01, GZ03, Mon08, PPF02, Ter01a, WX06a]. Fan [Fan07a, Hua08b, LC09, NV03]. Fast [Bér05, Bra06, CG00b, DMMV09, EG03, GVD08, GS04b, HV01, Ois01, PT05, Pot03, vdMRS06, BKO01, CFM03, EGG07, FM04b, HC08, LQ00, LG04, MM07, XQ08, YH04b, ZM08]. fastest [Ney06]. fat [Ano03-31, CGG02]. Favard [DS02]. FD [NCV08]. feasibility [HC08, JK06b, QX08]. feasible [LCP02]. features [Val02a]. February [Sch02b]. Feedback [BZ02, CSZ01, Dod05, RZ04, SS07a, ALCT05, ALC06, BFZ05, BZ05, BK01, CDMGP06, Dam02, DS07, DS09, EG02, HAT03, HZM02, LV06, MZK00, RSS00, RRH02, RZ01, RZ05, Ste01a, Sun02]. Fejér [AD03c, AD04, ATV09]. Ferrers [Dah03]. few [Wu06a]. FFT [Car09]. FG [Prz02]. Fibonacci [CHKS05, LK03, TR03]. Fiedler [Ano07n, BD05, BHN^{+}07, Hen07, KS08, Mol05a, SG09, Yon00]. Fiedler- [KS08]. field [pBlWyXwA08, BGL09, BBS06, EdRM04, FJ03, Gre02, MS00b, MZ08, Par02b, SS06c, SvW05, VW00, WL08, YY03, Zav08]. fields [AG00, Abd08, AMZ09, BS02c, BD05, Bu05, CFKP08, CSWM01, CS02b, Far08, MST06, OCK09, SvW05, Tup08a, VLA04, Wat01a, Zha04c, Zha05a, Zha05b]. Fike [wCL04]. filiform [AC03, BS01, GJMR01]. filling [CZZ09]. filter [CMX05, GM09a, Jia00b, Sen04]. filtering [DL04]. find [KMS01, KLMS03]. Finding [Amb07, Gra04b, Ple07, GZ08, Pet08]. Finite [BS02b, BR02, Fri00, GW03b, LRR04, Liu03b, LL06b, NY07b, Smi01c, Wen09, AAA^{+}02, AG00, Abd08, AJF09, AD03b, Arn04, BC08a, BGL09, BM06a, BF07b, BKK08, Cap01a, CFN04, CZ08, CHY05, CFKP08, Chl06, CL01b, CL02c, CD05, CS02b, Cos07, CRS07b, DvdV00, Di 03, Dom03, EG01, FZ04, Fao01a, FN07, FLRZ02, Gao03, GX09a, GKvS04, GN09, GW09b, Har07a, Har07b, JGPI08, Kar04, Kat04, Kes00, Kes07, Kin05, KPP07, LL02, LZ09a, gLxW02, MA06, MSR06, MR09, MT04, MW09, MST06, O’u02, Prǎ02, SS07a, SS06c, Sch02a, ST07, Tag04, Tre03a, VLA04, WSL02, WL08, WGL09, Wie01, YY03, Zhu08, vB09b]. finite-dimensional [Cap01a, Cos07, Fao01a, Har07b, JGPI08, MT04, O’u02, Sch02a, Wie01]. finite-horizon [BF07b, FN07]. finite-rank [MA06]. Finite-reductive [BS02b]. finite-time [AAA^{+}02]. Finitely [FM07b, AH09b, RGG01]. Finiteness [Lee09b, JB08, LS08b, ZG09, Zit09]. First [DDM08, AH04, CM00c, CHP07, FS07, Hof08, Hun05, Hun08, JL03]. first-order [JL03]. Fischer [BNdP09]. Fisher [GII08, IV04, KMZ00, KN00b, KS01, KMS05a, KS05, KS06b, KS06c, VWJ08]. fit [BSD09, FLM02]. fitted [Nie01a]. Fitting [Nie04a]. Five [CMV03]. Five-diagonal [CMV03]. Fixed [SS09c, FG08, Guo06c, GS06, HV08, HL04e, JN06, LLT07, MW03a, NY00, Pet09, Rih01, Ros09, wThY09, ZWS09]. fixed-point [FG08]. Fixing [Lip05]. flag [AMLM^{+}09, Arv08, Pue05]. Flanders [GJ04b]. Flat [BdPT08, CN09, BS04c, RS05a]. flats [CS02b, GN09]. flaws [SLP01]. flow [CDS04, Gla02, MA03]. flowers [Log05]. flows [AGKM09, DS03b, Dri04, DG07, GPFC04, GR07b, GR07a]. fluence [ZM08]. fluid [BILM06]. fold [KN03a]. folding [smbH02]. follow [Udd09]. Forbidden [HvdH07, Nik07d, SH00a]. forcing [Gro08]. Forest [CA02, AAGK06]. form [AC08, ACK05, BS02a, BBdH09, BBS06, BPZ08, CRU09, Cao09, DJL01, FI07, FH00, GS04a, Har04, Hol00, KPK04, Kob05, MH07, MSW09, Mue09, PH00, PPF02, Ral03, Sav05, Tan09a, Tre09d, Tri02, Zaï04, dB00, vdW02]. formal [BT04b, MT02]. forme [BR00a]. forming [PSvdW05]. forms [AEI00, Are09, BdlP06, BBL^{+}05, BS06b, BM04c, BBK03, BRS00, CLOvdD01, CSWM01, Chl06, CFP09, D– D05, – DZ04, DV09, Dri04, EGG07, Eld02b, FYL07, FS07, GR00, GH09, GQ09a, GP00a, Har04, HPP00, HS06a, HS06b, HS08, KTKR06, KLY09, Lad08, LR05b, MT05, MW09, Meh00, MR02a, OS03, Obl08, Ose08b, PHO07, Pre05, Rod07, Rod08, Sat05, Ste01a, SMS09, TG02, TT09, Tov00, Tup08a, VC04, Var06, ZM03, dMFZ09]. formula [AKL07, BT04a, BC04a, BR07, CC03, EK08a, EK08b, GH00b, GS09c, Kle00, KS09, Koz09, Kra03b, LdlP09, Lim07a, Lin08a, Map07, RR07a, SdS09, So04, TP09a, Tka05]. formulae [BBT03, BB07a, DP02b, Fur05a, KMS01, KLMS03, KT08, NT06c, Ose08b, XH08]. Formulas [SH09, DS03a, DD02, TR02, VPV01, YH04b]. formulation [CD00c, Ren01]. Fornasini [BKS01]. Fornasini-Marchesini [BKS01]. Foundations [Vin00]. Four [Bul08, Str01, HMR09, HL05a, IO01, MZ04, Sta09a]. four-block [IO01]. four-point [HMR09]. Fourier [DV07a, DV08, McE02, NS03, SC06, Ste08, Tad06, Yag01]. Fourth [BHRV02, GN00, NT02]. fourth-order [GN00]. fraction [FP04, MR02c, RL03, Van04]. Fractional [Adu02, FS03, Bur06]. fractions [EK08b, RK00, Zyg02]. frame [CL01b, KL07b]. Frames [BP05b, CL01b, ACRS05, BPT09, CL02c, HJM09, HP04a, Kal06, LCN04, PW02, Ren07, Str08, STDH07, Wal09]. framework [CPPS00, HHM09, MA07]. Franks [Che09, MSR06]. Fréchet [Zob00]. Fredholm [GV09a]. Free [LR00, BM04d, Bri01, DW05a, FM07b, Lip08, NY00, Nik08a, Nik09, SS07c, VW00, YY03]. French [BCSV06, Del01, Han00]. frequency [Pfa08]. Friedland [Ano09r, BKR^{+}09, CW03b, Wan09d]. Friedrich [Ano05q, Ano06h, FGH^{+}06]. Friedrichs [DH09]. Frobenius [AD03a, Bom08, BW08b, CR08, CKK08, GN05, Gus09, Lin02b, MPT02, MW09, MNV08, Nou06, Nou08, PT04, Ros09, Rum03b, Ste05a, gS02, TRH01]. Fuhrmann [Ano07o, Ant07]. Full [CRU09, SD00, Abd08, Ali09, BPR03, De 09b, FMM01, Ste09, Tri02]. Full-rank [SD00]. fullerene [CFRS02]. fully [MND01b]. Function [MPS03, Bri05, Che09, Dun07, Eld02b, Fer05, Fur02, GKL07, GL05b, GM09b, HG09, IL06, KN06a, KS06a, LRY05, MW03a, MS07, Mol06a, MFPPSS08, OZ03b, OZ03c, PR02b, PT06b, Ste08, Tyr04, Wu05c, ZS05, Zho03]. function-related [Tyr04]. Functional [EH07, BT00, CMP03, CQ08, CH08, GN03, Mul05, YM07, Yan09b, ZW00]. functionals [BM04e, BM06b, DHZ03b, Fan03a, HZ03, Mat05b, Nar05, VM04]. Functions [GGCB07, Par09b, Rod09, VS05, AEEG08, AADL03, ADL04, ADL06, ATV07, Alt09, AFR06, AD08, AP03b, Auj03, AS03, AB07, Bae07, BT09a, BB02b, BES02, BES09, Bek04, BN05, BHdST06, BD01a, BDG08, BR01, Bol02, BKR03, BKR04, BGM08, CH00a, CCW^{+}08, CL07b, Di 03, Dru04b, ES02, EKS01, ES01, FN00, FKL03, FKRS09, Fur08c, GW09a, Gen00, GHI09, Gin09, GN03, Gre02, Gre09, Han02b, HT07a, Han09, HZ02, Hia09, Hon06, HP03b, Hor07, HN09, HdABC07, IHM08, IK09, KVV09, KH01b, KH03, KS06e, Las06, LH03, LPR04, LS02a, McC05, MGWN06, MS01b, MS02a, Mul05, Nar05, OS09, OA09, OP05, OT09, Pat04, PM05, Pro08, Sak01, Sat06, Sen07, Sev08, SV01, Smi01c, Sza07, Thi04, Tik06, Tre04f, Wal05a, Wil07, WY07, fur08b]. Fundamental [Ros05, KV09, Šem03]. Further [BBL03a, BBL03b, BBL04, BBLT08, YLT09, GTT01]. Furuichi [Fur05b]. Furuta [FKN02, FK06, IK09, Yan01a]. Furuta-type [IK09]. fuzzy [Cec00, LPG03].

G [BM04d, HvdH07]. gain [YWS00]. Gale [AW02]. Galerkin [NB03]. Galois [GQ09a, GQ09b, Par02b, SvW05]. Gamas [Ber09]. game [XDS08]. games [MM06b, MQ06, Ost06]. GAMM [FMN04]. gamma [Mat05a]. gap [Bel05, Gol06, Lur06b]. gaps [Dug04b, Ros05]. Gauss [Kur04, LS00b, ZZ00]. Gaussian [DP02b, IV04, IPP07, KMZ00, KN00b, KMS00, She01, Yeu04]. gcd [Hon02, Alt09, AKYD16, Hon08]. gcd-closed [Hon02]. GCU [SSSG06]. Gelfand [Koz09, Ovs03]. Gelfand-Zetlin [Ovs03]. gene [EMT05, O’L08]. General [DDHL08, FYL07, GPFC04, Alf05, BHdST06, BS00d, BZ03, BCGM08, FJ04, Gav03, GP03, GP00b, GLP00, HN04, Han02a, Har00a, Mat07, OA09, SSA05, TP09b, TZ03, WM05, WY06a, WL09b, WGL09, Yan01b, ZZ00, ZGL09, ZTK09]. Generalisation [MNV08, Ala00, Han00]. generalisations [HGC07]. generalised [Dra05]. Generalization [GJ04b, jYzWbZ01, CJW05, DM07, HZ05b, JF01, Kle00, Lás03, LB01, Sav06a, Ste07a, Ste01b, Zen08, dMR07a, Han00]. Generalizations [BBK07, Chu08, ES08, Gre02, HDB00, Kol03a, MF09, Ano03-44, Fri00, Hol00]. Generalized [BS02a, Cal05a, Cap03, Che09, CMS02, DG02, FFM03, FM00b, FM02, FV04, FNT07, GS09a, Gu01, HL07b, Her03, HQ09, Ili08, KKL08, Liu03a, MS04c, MM06b, Moh05, Mol05b, Mul05, Neu01, Ots01, PV06a, PP04b, SS06c, Sch09b, WWQ04, YKF05, YC00a, YM07, Yan09b, YC01, ZFLW06, ZX05b, dSG07, ALSS08, AADL03, ADL04, AVV05, AH03, BW08a, BL04, BBL04, BB04b, BB04c, BBLT08, BJT04, Bar02, BM04a, BT05b, BV07, Boj03, CL00a, CH06, CT01, CTW01, CIDW06, CI09c, Czo05, DMM00a, DD04, DG03, DL05, DWD07, Dun04, Els05, FKP00, FT05, Fer05, FN07, Fur05a, Fur06b, Fur07, GGL08, GPM01, GS08, GAE02, GT06c, HLW06, HMR09, Hoc09, HM04, HSL05, HM06, INOY03, IK09, Ji05, JDS03a, KKP07, KN03b, KS04, Las06, LT07a]. generalized [LY09a, LL09b, LHZ04, LwW07, Lur06a, LY09b, Mah07, MPS06, MA07, Pap00, Pep08, PR04b, PP09, Pue05, QH09, Rie00, Roj07c, Roj08b, Roj09b, RM09a, SSW05, SC06, SH00b, SS01, SH02, SS02, SHS03a, SSW08, Ste04, Ste06b, Ste07c, Sty02a, Sty02b, Sud02, TR03, Tam04, Tia02, TT09, TS09, Tre02, Tre04a, Tre04d, Uhl01, VS04a, Wal05a, WSL02, WX06a, WX07a, WZ07, Wer09, Wir02, Wir05, Wu05b, XWD06, XHpS06, YH04a, YSS07, ZW00, ZX05a, vDHK07]. Generalized-confluent [YC00a]. generated [Cap01a, FM07b, Gao03, GX09a, Gem03, GN09, KA06, LZ09a, RGG01, Ros05, dSR05]. generating [Uhl01]. Generators [AS09, BSS07, BF08, VW00, YY03]. Generic [CtBDC09, GS04a, GPS01, IS08, GR00, RP05, vdW02, BdHZ08b]. genetic [Jef05]. genus [Mol06a]. geodesics [Arv08]. Geometric [ALM04, CFM03, Con08, Sob08, dDG02, dSQS00, Bad08, BK00b, Bha06, BH06, BK08, CE09, DS03a, FFN^{+}07, GJK04, JHL05, JLY09, KL07a, KN03b, Lim05c, Lim08b, Mat05b, Ney01a, Ney01b, Ney06, RL03, Yam06, YH04a, vDS05]. Geometrical [CR08, CTW01]. Geometry [Fie05a, HW05b, HC06, BLdPS05, BIL06, BH06, De 07b, DHZ03a, Don01, GC04, lHH08, Knu00, NS08b, NS08c, PPZ09, PP02, Šem03, Tar05, dS05]. Georg [Ros06]. Gersgorin [Kol05, Smi08b, Zha07b, Hof00, JSWT05, Par09a, BBCH08, MG09, Var04, sV06a]. Gersgorin-style [JSWT05]. Gersgorin-type [Smi08b]. Gershgorin [HP07]. Ghost [KLY09]. Giambelli [WY05]. Gilbert [BGS00, Blo08]. girth [LLT06, SU09, ZWS09]. given [BD03d, Dru02, FYL06, FYZ07, HL02, HW07a, II09, JS00b, JDS^{+}03b, KSH07, KS06d, LŠ07b, LL09a, Nik09, SAH08, XHSZ09, jYxL09, YFL09, YL08, ZSL09, ZLS09, dSF09, vD07, vDK07]. gives [BSD09]. GKK [HS02]. GKK-matrices [HS02]. Global [CKMS02, PPF02, Sun02, MP06]. globally [GSR03]. GLT [SC06]. GMRES [BGGC06, CLR00, DMW08, LZ09c, RS06b, Sid01, ZOE03]. GMRES-type [CLR00, Sid01]. goes [GT09]. Goldberg [Ton00]. Golden [BS07, Seo00]. Golden-Thompson [Seo00]. Golub [O’L08]. Good [MN07]. Google [Kir06]. Google-type [Kir06]. Graciano [Ano03t, Ano05-31]. grade [GS09e, MR02c]. Graded [BD02a, HNPT06, BD03a, Bon03b, Dah03, DT09, Mat04, TL03]. Gradient [CDS04, FNV08, Ovt06c]. gradient-type [FNV08]. grading [Eld06b]. Gradings [DM06, BTT09, Eld09, HPP00, HPP00]. Graduate [Moh05]. Graffiti [Zha04d]. grafting [Guo06a, PL08]. Gram [Ceg01, Dax00, KS07b, Pat04, PW02]. Graph [Alf00, Ano05-33, BHL^{+}04, DH06, DS08a, Hog01, SHZ05, Ste07b, AAB^{+}07, Bal04, BJ08, BM06a, BFH^{+}08, BOvdD04, BH08b, CS03a, CCRS08, Car09, CR02a, CG07b, cD03, Das07, EFN09, EFN10, FH07a, Fer07a, FP07, Gol09, GHK01, Guo06a, GZ06, GKMZ09, HK02a, Hag02, HM02, Hog08, HLW08, JD02, JDS03a, JDS^{+}03b, JTU03, KKL08, KK01, KS06e, KO08, LP01c, gLxW02, Li05, LB00, LLT04a, Mik09, MW01b, Nik07a, Nik07b, OvdDV00, Pan02, Pas01, PT06a, RSR00, RR04b, Roj07c, Roj09b, Shi07b, SHWR02, Ter02, TK09, Toy09, Wan09b, YC00b, ZL02, Zha05d, dW03]. Graphs [AABS09, BCRS08a, BCRS08b, CO09, Cve02, GRTR08, JTU01, OvdDV00, Sci09, vD07, vdH03b, Abd08, ÁFMNW06, AGHM04, AR06a, AMRR06, AGKM06, AK07, AG08, AJF09, AGO09, AGM09, AGKM09, Alf01, BvDK08, Bap05, BP05a, BFH04, BFH05, BF07a, BFS08, BHL09, BK03, BMS06, BMS07, BS04a, BS08, Bo02, BP05b, BHS06, CR07, CT03b, Che07a, CH06, Chu07, Cio06, CFRS02, CRS07a, CRS07b, Das04a, DB05, DK07, Das07, Dau08, Doo02, DJKM00, EN09, FKP03, FL03, Fan03b, FWG08, FQ09, FYZ07, FGM08, FN03, GL08, GLL09, GOTR07, GM08a, Gol03, Gol06, Gri04, GM08b, Guo03a, Guo05c, Guo05a, Guo05d, Guo07b, Guo08, GW09c, GYY09, Gut08, HLZ08, Hae08, HT08b, Hig06, HvdH07, HGW02, HT05a, HXL08, HLZ09, HW07a, Ili09a, Ind09, Jør05, KYAA08, KSH07, KMNS02, Kir07, KMMZ04, Kon08, KS06e]. graphs [LK00, Lee00, LZ07, Li08b, LM09a, LSC09a, LSC09b, LL09a, LW00, Lin08b, LLT04b, LLT07, LL08d, LZS09, LLY09, LSS00, LLT05, LLT06, LzZT07, Mer05, MQ06, Moh07, Mol06a, Mol08b, NS07a, Nik06a, Nik06b, Nik06c, Nik08a, Nik08b, Nik09, ORvdD02, OdAJ02, Omi08a, Omi08b, PR00, PBT04, PBR06, PBA09, Qi09, RBTR09, Roj07a, Roj08a, Roj09a, RJ02, Row07, Sat06, SGS08, SU09, SHZ05, Shi07a, SS07c, Shi09, Shp06, SW04, SS08a, ST07, Sta07, Sta09b, Ste00b, Ste02a, Ste03b, SS05b, SdAdFDV07, SAH08, SI09, SG09, TFZ08, TP09a, Ter03, THZ09, Wal09, WX06a, WX07a, WHAB08, WHBM09b, WHBM09a, Wu05b, WSL09, XHSZ09, XL05, YFL09, YLT04, YLT05a, YSL08, ZSL09, ZWS09, ZLS09, ZL03, Zha05c, ZZ08b, ZZ09, Zho08, dA07, dLdAOdF07, vDH03a, vDHK07, vDK07, vdH08a, vdH08b, vB09b]. Grassmann [Cal05a, Cal07, Chi02, De 08, DT09, McG06]. Grassmannian [BGLS04, Fin00]. Grassmannians [CS02b, De 09a, GR02]. greater [FL03, PBT04]. greatest [DTGV06]. Greedy [Hig06]. Greenberg [Cio06]. grid [Kar05b]. grids [Pot03]. Gröbner [AP04, Bou09, Kro01, LXW04, OF02, ZJL02]. Group [Ano07a, Ano08a, BTT09, Cai03, AM05, Bai03, BFZ05, Big02, BJ09b, CM00b, CG00a, CNX05b, CKS08, CK08, CL02a, CR02a, – Dok06, DR08, Eic04, EL01, Ell04, FM07a, GHO02, Gao03, HH00, HWC06, HLW08, Ish04, Kat04, KKABAF06, KKK07, KS07a, KK01, LW01, LPW08, Liu03b, LR03, LL06b, Nie00, Nie01b, Sat08, Sch00b, SH08, Sud02, VW00, WD05, YY03, ZZ05]. group-monotone [BJ09b]. Groups [VS04a, ALMS06, BC08a, BGM04, Che09, CHY05, Chl06, Cig01, Cig07, CDJ08, Dam04, DDI03, FZ04, FJ04, GX09a, Gew03, GN09, GW09b, Haz05, JKRT02, Kra04, MSR06, MR09, MOSZ02, RG03, Sud02, TX09, WGL09, Yan06, Yan07]. Growth [KMS00, Bol01, GZ07, JPB08, KMS05b, KMS05c, LS06c, VJ06, ZZ00, ZGL09]. Grunenfelder [CL07a]. Grüss [PR04a, Ren01]. Guaranteed [YWS00, PGRO06]. Guo [RS09, BH08b]. Guorong [Moh05]. Guttman [TY05, TY07, TY09]. gyroscopic [LMZ03].

Hadamard [AG05a, Che03, Che04a,

Che04b, Che07c, CLL07, CS01, CFJ01, FJ07, Fan07a, FS03, Hua08b, KMS01, KLMS03, LHSL07, LCW09, LC09, MW09, PW02, QZ07, RT01b, Seo00, Son00, SDL06, Tan09b, Vis00, WZZ00, Xia03]. Hafnian [Sch09c]. Hägele [LS09b]. Haifa [Ano03-46, BLL06]. half [BD09a, De 07b, FF04a, FF04c, Tag04]. half-line [BD09a]. half-spin [De 07b]. halfspace [QX08]. halfspace-relaxation [QX08]. Hall [HL04c, Nör03]. Halmos [Fao01a]. Hamilton [AD03a, BR00b, FMM01, Koš03]. Hamiltonian [BK06b, Che02b, FMM01, FI05, Ikr01, MSW09, Shi04, Shi06, Wat05, Wat04b]. Hamming [BvDK08, Jai07, VR06]. hand [DMW08]. Hankel [ADT06, BQW09, CI09b, FT00, Gen00, Hei01a, Hei02a, HR03a, KG09, LQ00, MT05, PEPS00, Sal00, VPV01, Var03, XQ08]. hard [CGHG06, BP02]. hard-on-average [CGHG06]. hardback [Rod09]. hardness [BC04b]. Hardy [OT01, dMR07a]. Harmonic [Car09, ES01, Mol09, MZ06, PBR06, SvdE03, Smi01c]. Harnack [ES01]. Harry [Rod09]. Hartley [BD03b, DLZ03]. Hartley-type [BD03b, DLZ03]. Hartwig [BST09, VWJ08]. Haukkanen [AKYD16]. Hausdorff [BC04a, CW07a, CW07b, Kes01]. having [BKR04, DK07, DW05a, KK01, LŠ06a, LK00, Nou06, SS02, WX07a]. Heidelberg [Kol05]. height [Nom05c]. Heinig [Ros06]. Heinz [Aud07, Yos07]. Heisenberg [DDI03]. Helmholtz [BSS09]. Helmut [Sch02b]. help [Log05]. Henk [Ano09n]. hereditary [Drä03]. Herglotz [BHdST06]. Hermite [BFZ04, BFZ05, Bat06, Chi04, HDB00, Hol03, OTC03, RR05]. Hermite-Biehler [HDB00, OTC03]. Hermitian [HL09, dSQS00, HR04, Rod08, SHS03b, Str03a, Tan05b, Xu08, BGN08, BH04, BS00a, BNdP^{+}05, BM04c, Buc06, BFJ07, CMP03, CC09, CX05, Dan02, De 09b, DMM00b, DL08, EGO05b, FJ06, Fre08, Ful00, Gro01, GGCB07, HK02a, HTV02, HW05b, Jia09, JD02, JDS^{+}03b, LL05a, Lim05a, LTT09, Meh00, MD09, NBdP08a, Nak03, SCS09, Str06, Tic05, Toy09, Tre03b, Tre04c, TL03, TS06, WB01, WB08, Wer08, ZZY06, Zha03b, ZyC03, dF00]. hermitian-preserving [SHS03b]. Hersonsky [CW03b]. Hessenberg [BEG^{+}07, BM06b, EGG07, Els05, FV04, God09, Meh00, Zem06]. Hessian [CKK08]. Heydar [Ano02r, BORŠ04]. Hiai [FK06, ST08]. hidden [Chu08, MN07, VWD08]. Hierarchical [Peñ08, GX09b]. hierarchy [FO01, VR06]. high [CCSS03, FB04, LCN04, NCCY00, So04]. high-resolution [CCSS03, FB04, LCN04]. Higham [Par09b]. Higher [BJ09a, CNF08, CKŻ06, Buk07, CDMGP06, CSWM01, Chl06, DV04, KL07a, LPS05, Lom09, Mei05, OS03, Pet08, Sen07, SC01, TX09]. higher-dimensional [Lom09]. Higher-order [CNF08, DV04, Sen07, SC01]. Higher-rank [CKŻ06]. Highly [ALSS08, Sla03]. Hilbert [ABP09, ABK02, AR06b, ĆDR01, DK08a, Dra06, Dra08a, Dra08b, DD05, DD07, Dug05b, FYY09, FP02a, FQ04, Fur04, GMP08, HJM09, HK00, HK07, Hir09, HHT06, Kim05, Kit09, LS02c, Lim07a, Mol03, PR07, Sha09, Str06, XS08, Xu09, YLZ06, Zaï04]. Hilbertian [ACG08, Suc09]. historical [Abe08, Cve02]. hitting [CZ08]. HIV [HL05b]. Hochschild [GR03]. Hoffman [Bar05, HT05a, WD06]. Hoffman-type [HT05a]. Hölder [BLFS09, MQ03]. holes [CZZ09]. holomorphic [BP06b]. homeomorphic [FPP01]. Homogeneous [Arv08, DT03, GZZ00, Lim09b, Lip08, SU09]. Homomorphisms [DSW07b, Buk07, Cai03, DH06, KB02b]. homotopies [Wri06]. homotopy [Lom02]. Hong [Alt09]. honor [Ano02r, Ano02z, Ano03t, Ano04u, Ano04v, Ano05q, Ano05r]. Hopf [Adu02, AMZ09, CM00b, FM07b]. horizon [BF07b, FN07, Kar04]. Horn [CKS06, Ano05r, Ano07t, Ano07p, Ano07u, BKMZ07, Knu00]. Hotelling [Gus09]. Householder [Bag08, MMT04, SAAF08, Wri06, xZxGpL09]. Howard [O’L08]. HR [Ano04-43]. Hu [CL00a]. Hu-Hurley-Tam [CL00a]. hua [XXZ09, Mos09, Šem03]. hua-marcus-bellman-ando [XXZ09]. Hückel [Wan09e]. hull [BG06]. hulls [DS04a, DLS08, FGM03, MOSZ02]. Hurley [CL00a]. Hurwitz [AYC06, CGR09, Eld02b, Hol03, Peñ04, Yan03]. Hurwitz-Radon [Eld02b]. hv [BKDL01]. hv-convex [BKDL01]. Hwang [Nie07a]. Hybrid [HZM02, NB03]. hyperbolic [BQS00, Boj03, HV03, HTV02, JKRT02, Lie09, Mul05, ST00, Sla03, ST03b]. hypercube [dS05]. hypercyclicity [Cao07a]. hyperenergetic [SS05b]. hypergeneralized [BBL04, BBG06, BBLT08, Ste06b]. Hypergraph [BDKW04]. Hypergraph-based [BDKW04]. hypergraphs [Bru03b, KLM01, MS02b]. hyperinvariant [DM02]. hyperplane [BHL^{+}04]. Hyperplanes [De 09a, De 08, Nie01a, NS07b]. Hyperspheres [Nie01a]. hypertournament [KPT01, KR03]. hypertrees [BCDF08, Siv09]. hypoenergetic [LM09a]. hyponormal [Cao05, CLY08, DJ07, Dug08, MRR07]. hypotheses [CM09].

ideal [ALC06, pBhSzZ09, HAT03]. ideals [AMS05, Con08, Con09, Dah01, LZ09a, LL09b, Lin03, LL08c, RY08, Roc09, RGG01]. Idempotence [Zha04b, Kuz02a, SZZ07, XCX01]. Idempotence-preserving [Zha04b]. Idempotency [BB00, BBS02a, Bak04c, BB07b, BT05a, FJP07, WFJ08]. Idempotent [DO09a, BB00, BBS02a, Bak04c, BB04a, BBS04, BB07b, BT05a, CGQ04, CV07, Dup03, Foš05, HCLC09, Hua08a, KR06, LZ09a, Rao09, TS01]. idempotents [AH09a, DW09a, Den09, Dol03, DRR02, Rab04]. Identification [AFCC04, DHFS00, LD02, AD05, GX09b, JLMP02, TG02]. identify [FLPW07]. identité [Han00]. Identities [BD03a, Gil08, BD02a, BB04d, Bap03a, BE01, Bir04, BP09, DT09, – DS08, Hil08b, HT05a, HK05, JMT04, Mat04, PH04, TTM02, Wen09]. identity [BM04a, FM00a, Han00, Kes00, Wan04, WY05]. Ihara [Hor07, TP09a]. II [Auj03, BB05b, BCRS08b, Bou03c, CW07b, CS01, CS03c, Dod08b, FT07, FS06a, FF04a, GW04c, GT06b, Kes01, KLMS03, LZ09c, Ney01b, Nik07b, NS08c, OZ03c, Ovt06b, Sed01, SS07e, gSX03, Tka05, wW02, Yan01a, dCdS05, dSG07, vDS05]. III [AADL03, BBCH08, CH00a, HPP00, KN03b, gLxW02]. ILAS [Ano04t, Ano05p, Ano07s, Ano09p, BBVR08, BGKW07, IMS^{+}09, Pue01, Uhl04, dOQSZ06]. Ill [CLPT06, ÖA00, CRS03]. Ill-conditioning [ÖA00]. ill-posed [CRS03]. Ill-posedness [CLPT06]. Ilmonen [AKYD16]. ILU [Bol01, KY00]. image [CCSS03, FLM02, FM04b, FB04, GMR^{+}07, HHM09, HNV00, LXW04, LCN04, MLK^{+}04, NCCY00, NPP00, Pop08]. image-based [HHM09]. images [DV01, SvdS01, VHK01]. imaginary [AYC06, Ste07a]. imaging [TKX08]. Imbedding [QD09]. Immanant [CD06]. Immanantal [Mer05, dCdS05, dSdC05]. impact [LMT09]. Implementation [AEEG08, MLK^{+}04, LMV03, PM00, PA04]. implications [Dau08, Smi01b]. implicit [Boj03, DJMS09, Dek09, NB03]. implicitization [DMMV09]. implicitly [OCK09]. imply [TTM02]. implying [Fur07]. importance [Har00b]. impossible [PFL05]. Imprimitive [CL08b, Zha07a]. Improved [HIU04, NX04, Roj05, cD03, HDA09, LT08]. improvement [LW01, LHF06]. improvements [HNT03, PP01b]. Improving [HW07b, Wal05b]. impurities [BES01]. IMRT [Ano06f, CGXL08, CCW^{+}08, SK08a, ZM08]. Incidence [Wil09, AGKM06, GR03, Gri00, GKMZ09, IHM08, KH01b, KH03, LP00, Spi05]. incline [HL04b, HL04a, HLG04, HL04c, HLW05, KR04]. Inclines [KR04, HL04b, HL07a]. Inclusion [BI03, LHL08, Bap00, HSZ08, Kol08, Mel09]. inclusions [Bar08, Fer07b, O’u02, VEB00]. Incomplete [CS08a, ZGL09]. inconsistent [BS00b, CLR00, CS08a]. increasing [Bha03a, Wir05]. Indecomposable [GM08b, YLS05, CHK^{+}01, vGKN09, Ter01a]. indefinite [ARR06, ADL06, And04b, BLdPS05, BdPT08, BT04a, BKR03, BKR04, FGR07, HHZ09, Kin05, MR01, MRR07, Rod06, RS01, Sch09a, Ser08, SS00, TL03, TS06]. independence [VC04, XHSZ09]. independent [CL02c, Cig01, CM09, EE04, HHL^{+}09, OTZ09]. Index [Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano01d, Ano01p, Ano02c, Ano03h, CS01, Dvo03, TI08, Unk01, AFS09, AK09a, AK09b, Ano01a, Ano01b, Ano01c, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01n, Ano02p, Ano02a, Ano02b, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano03r, Ano03s, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano04r, Ano04s, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l]. index [Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano05o, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, AMM06, BMS06, BMS07, BFPW02, DJMS09, DL09, Ed05, GLS07a, GGM^{+}09, GLL09, GOTR07, GV09a, HL07a, Kes01, LWZ09, LDS08, LCP02, Omi08a, Omi08b, Rod05, SdS09, SM09b, WHBM09b, dlPGR07, vdMSR01, vdMSR03, vB09b, Ano01o, Ano03g]. index- [BFPW02]. Indices [HL04a, AMZ09, BFZ04, BFZ05, Bo02, BRVS08, DK07, KYAA08, OS09, dSZ00a, dSZ00b]. individual [TW00]. indscal [tBSR04, BSD09, Ste09]. induced [BdlP06, BD02b, CDL05, EN09, Nie00, ST03a, Tar05]. Inequalities [BN04, BNdP^{+}05, BLdP05, Boc03a, Che07b, CHL02, Dra08a, HZ02, Hir05, HK07, Kos06, PV06b, AB05b, Auj03, AS03, AB07, BDK08, BHNR03, BGJS06, BdL04, BK00b, Bha03b, BT07a, Bon03a, BD09a, Bou06b, BS07, CJPS02, Che03, CKZ05, Con09, CSK01, Dra06, Dra08b, DD07, EH07, FNS01, FST04, FNT07, FYK05, Fur06a, FKY09, FG03, Fur03, Fur05a, Fur05b, Fur05c, Fur06b, GJR03, GN03, Gut01, HK00, HK02b, Hir09, HZY05, Hua08b, JMP09, KN04, Kit04, Kit09, Kly00, LP01b, LM02, Li06a, Liu02b, LC09, Mah07, Mal01a, MPS00, MZ08, NBdP08a, Nie04b, Nie07b, OT01, Pat08b, Per07, PWZ05, QT06, Rao00, Sed04, Seo00, Tao06, Uch02, WlW07, XXZ09, zY08, YKF05, YFK05, YLT09, Zha04a, Zha00]. inequality [ABP09, Alz01, And04b, Aud06, Aud07, Aud08, Aud09a, Auj00, Bén06, BD01a, Bha06, BK08, Bou06a, BU07, BS07, BLFS09, CE09, CS09b, Dra05, Dru08b, EFN^{+}03, FP02a, Fuj05, FK06, FFN^{+}07, Fur04, Fur07, Fur09, HG09, HZ05b, JP04, Jia00a, KL07a, Kle07, Kom08, KS04, Lee09a, LP01c, LCP02, LS02d, MPT04, Mal01a, MPP06, MF09, Mos09, NY02, Pas01, PR04a, Rad09, Ren01, Sch02c, ST08, Son00, TP05, Wad07, wW02, Wu08c, WGT09, Xia03, Yam06, Yan01a, Yan02b, Yos07]. Inertia [FS04a, FS06a, SSG04, AS01, BHL09, CL00b, Dau08, Fer07b, GTM09, HLW01, JF01, MSS07, Sty02b, dF00]. inertial [IK06]. Inertially [KOvdD07, ML02, CV05, CVV09]. inertias [Dan02]. Inexact [BMGS06, BGN08, BX08, BW08a, DW08, RS06b]. inexpensive [Bra03]. infeasibility [CBIXG08, Peñ03b]. infeasible [AB05b]. infectivity [HL05b]. inference [Vin00]. Infinite [BES01, DOS01, Tab01, AMZ04, BK07, BGM01, DL92, DL01, FLMR00, FB05, Fri00, KP08, KMS05b, Pfa08, PSvdW05, Sak01, SS06c, She00, SS09d, Wer08, vdMSR01, vdMSR03]. infinitely [CH00a, Tre04f]. infinity [GT09, Kol09]. influence [BLR00, FLPW03, Liu02a]. influential [LD02]. information [Fuj05, GII08, JB00, KMZ00, KN00b, Kle00, KS01, KMS05a, KS05, KS06b, KS06c, KS09]. Ingram [Ano04v]. Inherited [ABH07]. inhomogeneous [THH04]. initial [BS00b]. injection [RZ01, RZ04, RZ05]. injective [LS06b]. Inner [DvdV00, DM03, GW09a, Gla04, ARR06, BLdPS05, Dra05, Eld02a, FGR07, GZZ00, GL09, MR01, MRR07, Pat08b, Rod06, Ser08, Zho00]. Inner-outer [DvdV00, GZZ00]. inner-product [GL09]. input [Zen01]. input-output [Zen01]. inseparable [Ahm05, Pit03]. instability [BD01b, BD06b, SvdD05]. integer [BW00, BP02, EK08a, FM07a, Gue08, JPB08, LX08, PW08a, PP09, RP03, WW08a, Zaï04]. integer-valued [Gue08]. integrable [AL09b]. Integral [PRS02, AABS09, BP05a, BdlP06, BV02, BH08a, Bru02b, Che09, CT09, Fan03b, Fan07b, GM08b, GS09c, Igl01, Kir07, Ose08a, PR01, PBR06, Pih04, Rao09, SdAdFDV07, Yan06, Yan07, Zha06, dSF09, dLdAOdF07, fur08b]. integrals [Igl01]. integration [VB09a]. integrator [ZTK09]. Integrators [BM07]. Intensity [Ano06f, CBIXG08, CGXL08, HC08, SK08b]. Intensity-Modulated [Ano06f, CBIXG08, CGXL08, HC08]. Inter [HMS01]. Inter- [HMS01]. interaction [GZHT05]. interactor [XM02]. interconnection [Bal07]. interconnections [GSR03]. interest [Kob00]. interior [CH01, GL06, OS05]. Interlace [AvdH04]. Interlacing [LM02, WSL09, Li06a]. internal [Lom07a, SvdE03]. International [AHST09, Ano03-46, Ano09q, EMS06]. Interpolating [Bha06, Bri05]. Interpolation [AT00, Joc09a, PS07, AD02b, ABR02, AD03c, AD04, ATV09, BB02b, BN05, CH00a, CH01, Con08, FKLN07, FKS09, GVV03, GJ04a, GM07a, HdABC07, JS01b, KR07, McE02, OF02, OT01, Pop02, RR04a, RR05, Wu05c]. interpolatory [CGR09]. interpretation [Pic07]. intersection [LSD06, Mil03, RDxC03]. Intersections [FLRZ02, LT03a]. Interval [CCG02, HCLC09, Myš05, AM04, CWW09, DS02, Di 03, GG04, Jan01, Kul01, MW03a, Myš06, Rih01]. Intervals [Gar03, HZ06, TH00, TK09]. interview [Blo08]. Intrinsic [Fie08]. Introduction [Sch00b]. Invariance [GT06c, BB05a, Las06, Lav03a]. invariances [Ots01]. Invariant [CHKS05, DM02, Kle04, RZ01, Wu08d, ABW05, AFK02, And06, AP04, Bad08, BR02, BC07, BC08b, Bra03, BCS02, CD00a, CG00a, CLS05, CL02a, CL02b, CM08a, CHPP06, CS03d, CH08, DHFS00, DDHL08, DH04, EMT05, FYL07, Fao01a, FP01, FIL07, FGR07, FH01, GV06a, HZ02, HZ04a, HN09, HZM02, Ips00, KS06a, KN03a, KL07b, KP02, LCF06, Li08c, Lom07b, MZ05, MS03c, MRR07, MBLV01, Nah06, Oku03, Oku06, Ovt06a, PR04c, Poo08, PPZ02, PPF02, Rvz09, RS02, RZ05, RT01b, Sil00, TH01, Tsa01, Val08, WLBZ06, ZWPR06, BCES05]. Invariants [BPP08, AMZ06, BZ02, BdHZ08b, BdHZ08a, Boc03a, Boc03b, BK01, CSZ01, CDMGP06, Dod05, DS07, Dom00, FM07b, For04, FMS01, FMS05, KZ03, LW00, Mer05, RRH02, RZ04, SH09]. Inverse [CSK01, GvdV04, Hog00, IWL07, JDS03a, KH09, LR05c, Tre04d, AD06, AT09, BX08, BBT03, BB07a, Bap00, BHL09, BES02, BMGS06, Bol01, BMS04, BMS08a, Bri01, BOvdD04, CKLX09, CGD05, Ceg01, CBIXG08, CC00, Che04b, Che07b, Che07c, CXW09, CL01b, Cui07, CIDW06, CI08, CI09c, DW09a, DW09b, DS03b, ELN04, ES08, Fan10, Fei04, Fie01b, FNS07, GKW00, Gon05, Gro00a, Gu01, HFR04, HWW01, Hog02, Hol05, HP04c, JS01a, JO05, JS07b, JS07c, KKP07, KV09, KN03b, KDC07, Kol09, Koš03, Kro01, LW01, LHSL07, LW07, LCW09, Lim05c, Lin08a, LwW07, gLgWmW08, LM04b, Mal01b, MPS07, MHZU05, Mor08, Neu00, Ney01a, Ney01b, Ney06, pPyHZ05, PY09, Pat01, Pat03, PP01b, PHZ06, PS07, Phi04, PES07, RW02, SLR06, Sid01, Šmi04, Son00, SBM05, SR06, Sot06]. inverse [SD00, TVB08, TMARAE^{+}07, WZZ00, Wd09, WD05, Xia03, XHpS06, XLG^{+}13, Zha01, ZHZ03, xZxGpL09]. Inverses [BR09, LM06a, Moh05, Neu00, Alt09, AH03, BS02a, BB04c, Bap03a, BCEG09, Bu05, CGP04, CNX05b, CTW01, Den09, Din03, DD05, Gha01, Gin09, GT06c, GWW06, HTK07, Hei02b, HSZ08, HM04, HM06, JS07a, Ji05, KS07b, Li08d, Mah07, Rie00, Rob02, SS06c, SS01, SH02, SS02, SHS03a, Tia02, TT09, TS09, Wan02, WWQ04, WZ07, WG02a, WG02b, XH08, Xu09, YC01, ZJL02, dFP01]. Inversion [ACM03, Bas02, Fie05b, Gha01, DvdV00, EG01, EG03, GKvS04, NRW02, OTZ09, PH09, YH04b]. Invert [Ney09]. inverted [Dam03]. invertibilities [HC03a]. Invertibility [LLM03, EGMG07, Gil01, PP04a, PP04b, Pfa08, PG04]. Invertible [DO09b, HL04b, CB06, CL07b, LRY05, MD09, Oku06, SS06c, Tan07]. inverting [Wil07, IT07a]. investigation [CC09]. involutary [TS01]. involution [DK08a, EMERP06, KDC07, Tup08a, YZ09]. Involutions [Spi05, ABR02, Bot06a, CM00a, CKS08, CK08, Ell04, Fur08a, Zhe02]. involutory [AG04b, Tre05, Tre08]. involving [BN04, CW05, DK06, DW09a, Fur05c, GN03, GT06c, HZ02, Hil07a, LP01b, Liu02b, MZ08, Nie04b, NT06c, WlW07, YLT09]. Irreducible [BBK03, DK04, Kra04, LHS02, SBS02, AFK02, Bre03, CJK05, FM07a, FLQ04, Har07b, LM06a, Lon04, Sav03, Ser09, Wu05a, WD06, Yah04, YSS07, vdW01]. irregular [Shi09]. ISBN [Böt06b, Kol05, Moh05, Rod09, Tům06]. ISBN-10 [Böt06b, Kol05, Moh05, Tům06]. ISBN-13 [Moh05, Rod09, Tům06]. iso [DG07]. iso-spectral [DG07]. Isolated [TU02, TUC04, WX07a]. isometric [De 09b, Ser08, Ste07c]. Isometries [CLS05, GS08, ACG08, Jel03, Lav03a, Lav05, Sha09, Suc09]. isometry [– Dok06, DR08]. isomorphic [BB04d, BB05b]. isomorphism [Aga02a]. isomorphisms [BBC00, HZ04b, HBLZ06, LM03, LT07c, LL03, Mol06b, PJ07, XCX01]. isoperimetric [HD03]. Isospectral [GR07b, GR07a, Lan05, DS03b, GPFC04, LST08]. Isotropic [GHO02, GWL09a, GWL09b, GLW09, GW09b]. Israel [Ano03-46]. Issue [Ano02z, Ano02-28, Ano02-27, Ano02-29, Ano03v, Ano03t, Ano03u, Ano03-44, Ano03-45, Ano04u, Ano04v, Ano04t, Ano05q, Ano06f, Ano06g, DOS01, HKSS03, Ano02r, Ano04-43, Ano07u, BFSV06, BKMZ07, BHRV02, BKPW08, CNTV08, FMPT04, LGH01, Ano05r, Ano05-33]. Issues [Ano05-34, LB01]. Iterated [LN06, AY03, BHNR03]. iterates [Mat02, MH06]. Iteration [Dax03b, LV03, BBMX07, BMGS06, CKMS02, Dek09, EGO05b, FG08, FS08, Had05, HLR08, KN03b, Ney01a, Ney01b, Ney06, NB03, Ovt06c, PRW01, PY09, Rih01, SSA05, Str02, WB08, YLX09]. Iterations [MST06, BPS04, FLMR00, GZZ00, HT05b, Pit09, Yua00]. Iterative [ESP05, SS03a, AH08, BW03, Cen01, CHP06, Dax00, FS08, GG09, Gug08, HNT03, Had04, Huh05, Li03, LWZ09, MS00a, MM03, MP06, Pul07, Sol06, Wan07a, XZ06, Yag01]. IV [BCES05, Li00b, Ney06]. IW [PP09]. IW-contractions [PP09].

J [Ano09o, BN00, Par09b]. Jacobi [Bap03a, Boj03, BSS09, BW06, DD04, FMM01, FS08, HS03, Hoc09, Kle09, LS00b, Mal09, Mat02, MH06, yPyHZ03, Vos07]. Jacobi-like [Boj03]. Jacobians [Har00b]. James [Tur05]. Jensen [Mal01a, MPP06, MPS00]. Jerzy [BS05a, Cal05b]. Jet [BKR04]. join [BF07a, IHM08, KYAA08, KH03, KH08]. joined [Roj08b]. Joint [BM04c, Bek04, BNT05, Blo08, BKPW08, CZ00, Fre04, GJK04, KS06d, Mae08, MOS07, Myu09, PJ08, PW08b]. jointly [GN00, Lim09b]. Jordan [Ali09, Bae07, Bal07, BR00a, BBC00, BdHZ05, BOP08, Ben05, BPZ08, DKL^{+}07, FGM03, FMZ06, FMS05, FJL08, GR00, GZ07, GST04, GTM09, HL07b, Hol00, HJ08, HQ09, HBLZ06, JL09, Ji09, KPK04, Kru04, LS06b, LJ04, LL05b, LL08b, LL04, LT07c, Lu02a, LL03, Lu03, Lu09, MS03b, MS03c, MG09, Mol06b, Obl08, PRS02, Pei07, PH00, Roc09, Sav05, Sch00a, Str03a, WY04, WFJ08, Wat04a, Wil07, Wu09, XCX01, ZM03, ZY06, ZG09, ZX05a, dB00]. July [dSQS00]. jump [BG02]. June [Uhl04]. justification [Vos07].

Kac [BB09, BR07, BT04b, Tre02]. Kaczmarz [Pop08]. Kalman [MSS07, Sen04]. Kantorovich [Dra08b, FST04, FG03, Fur06b, HZ05b, PWZ05, Yam06]. Kantorovich-type [PWZ05]. Karatsuba [Ose08b]. Karatsuba-like [Ose08b]. Karla [Ros06]. Kauffman [RT06]. KdV [GHS06]. Keller [Che02a, LDS08]. Kernel [Hei02a, ABK02, Bre03, ĆDR01, DMM00a, DD07, Dug00, Dug01, Dug04b, Fer01b, Li06b, Ose08a]. kernels [Ahm05, AGKM06, BHNR03, BCES05, BP06b]. Key [Woo02, BD01a]. Kharitonov [OO05]. Khatri [Liu02b, YLT09]. kind [PHZ06, dSZ00a]. kinds [ACHR01]. Kleene [SSB09]. Klyachko [Fri00, Har07a]. kneading [AFR06]. Kneser [dH07]. knot [LW00]. knots [NS03]. known [Yan09b]. Kolmogorov [FF02]. König [Shi01]. Kostant [SP09]. KP [FO01]. Krawtchouk [IT07b, Sev08]. Krein [And09, BNdP09, BT04a, IO01, NBdP05, Str03b]. Kronecker [AG05b, BD04, CD00a, DV07a, eFLW09, Han05, Huh06, JM09, OZ03a, OZ03b, OZ03c, PPF02, RO02, SMS09, SH09, Tad06, Tyr04, VC04, Vis00, Wat01b, Zav07, dSZ00a]. Kronecker-product [Tyr04]. Krull [CGGGSR01, Kat04]. Kruskal [SS07f]. Krylov [AEEG08, CRS03, CW02, DW08, Dek09, Gug08, GS09e, HLY06, JR09, RS02, SL06, SP04, Ste02b, Zem07]. Kubo [Dit00]. Kubo-Mori [Dit00]. Kuriyama [Fur05b]. Kutta [Dek09]. Ky [NV03]. Ky-Fan [NV03].

L [Ano05q, SH00b]. L-matrices [SH00b].

LAA [Ano09r, BKR^{+}09, BMS08b, CNTV08, Hof08]. Laffey [Ano09o]. Lagrange [VS04b, dBS08]. Laguerre [Nie07b, Chi04]. Lancaster [Ano02z, Ano04-44]. Lanczos [VV06, AS01, BF01, BQW09, DBK05, Dru08c, GZZ00, HS07a, MS03a, Ney09, PM08, Wat04b]. Laplace [AD03a]. Laplacian [Shi07a, AC05, BJ08, Bap05, BZ00, BHS06, BH08b, CCRS08, CA02, cD03, Das04a, Das04b, DB05, Das07, FL03, FYL06, Fie05a, Gol06, GM08b, Guo03a, Guo03b, Guo05a, Guo05b, Guo06a, Guo06c, Guo08, GW09c, GZ06, II09, Ili09b, Kir07, LP01c, LSC09a, LSC09b, Lin08b, LLT04a, LLT05, LzZT07, MW01b, Moh07, Mol05a, Mol08b, OdAJ02, OT07, Omi08a, Omi09, Pan02, PBT04, RSR00, Roj02, RP03, RR04b, RS05b, Roj07b, Roj07a, ySZyY06, SHWR02, Sta09a, SI09, SG09, TFZ08, THZ09, WX06b, WHBM09b, Wu05b, YLT05a, YLT05b, YL08, ZSL09, ZL02, ZL03, Zha04f, ZZ08b, ZLC09, Zho08, dLdAOdF07]. Laplacians [BKN05, BHL^{+}04, CRS07b, Li06a]. Large [Ano02-27, Ano02-29, CG07a, GZZ00, ÁFMNW03, AG05a, BFSV06, BES03, BDKW04, CS08a, GRV06, GA06b, Gug08, JR06, KN00a, Kir00, MNOvdD08, Mal09, OS05, Pen06, RM09b, SL06, SU09, SW03, Vos07, WSW06, Wu06a, Yan02a, Yan09a]. Large-Scale [Ano02-29, BFSV06, GRV06, GA06b, Gug08, OS05, RM09b]. largest [AEK06, AK09b, BMNA08, BHS06, CH03, CO09, Das04a, DB05, GH06, GT02b, GT04, Guo05c, Guo05b, HL02, LRWW07, LP01c, LG06, LB00, LYY07, Pac02, RR04b, Roj05, Roj06b, RR07a, Roj07a, SHWR02, Sta07, Sta09b, Ste03a, Van07, WV08, ZWY00]. Latent [Hen07]. Latimer [BV02]. Lattice [BM05, BP09, BLG01, HQ08, LLM03, LL08b, LL03, Nie00, SHS03b, Tan01, WL07]. Lattice-ordered [BM05]. Lattices [Gao03, GX09a, GLW09, GN09, WL08, ABW05, DDHL08, LM03, PJ07, Tan03a, Tan05a]. Laurent [AY08, AH04, BM04d, BES01, JR09]. law [BB05a, JF01, Wan02, WWJ09, Yan02a, Yan09a]. laws [WG02b]. Lawson [FFN^{+}07]. Lax [Lie09]. LCM [Hon02, Hon06, Hon08]. LCM-closed [Hon06]. leading [Bal03]. leaping [EK08b]. learning [AAA^{+}02, Per03]. Least [Eld02a, Eld06a, ZHZ08, AB05b, Bak04b, BCRS08a, BCRS08b, BD09b, CCRS08, Čer09, Cve02, DW00, EF07, FWG08, FL04, FLPW07, FB04, GA06b, GJW02, GWW06, JB00, KM04, LMV03, LD02, LCN04, LZS09, LLT06, MLK^{+}04, NPP00, Nie01a, OCK07, PBA09, PA04, VHK03, Wal05a, WL09b, WG02b, XQW08, YFL09, Yer04, ZZ00, ZM08, dSS08, Sch06]. least-norm [WL09b]. Least-squares [Eld02a, Eld06a, BD09b, EF07, JB00, Nie01a, ZZ00, ZM08, Sch06]. leaves [II09]. leaving [CL02a, CH08]. Lebesgue [CMP03]. Lee [JNL05, JN06]. Left [EGMG07, MV07, BBL03a, HS01, LSD06, ZX05a]. left-star [BBL03a]. Legacy [Ano03u]. Leibniz [LH05, Yan09b]. lemma [Lás03, MSS07]. Lemos [FNT07, Fur07, MF09]. Length [Ish06, GJMR01, GM09b, JN06]. Lengths [CD05, Jia00b, Kru04]. Leonard [Cur07, Har05, NT06a, NT06b, NT06c, NT07a, NT07b, NT08c, NT09a, Tan09a, Ter04, Vid07]. Leontief [Peñ08, SdL03]. lesion [PSvdW05]. lesion-forming [PSvdW05]. Leslie [CL05a]. less [CO09, Omi08a, Omi08b, Sta09b]. Level [MM08, ACM03, LG04, MN08, NCV05]. level/phase [LG04]. Leverrier [rWQ02]. Leverrier-Chebyshev [rWQ02]. Levinger [PT03, PT06b]. Levinson [Mel01]. Lewin [LJ01]. Li [Lim08b, Yam06, HWKS03]. Lie [FMM05, JW05, WL09a, ZWC06, BPP08, GJMR01, HN09, Mou06, Aga02a, AKL07, AD02a, AC03, BC08a, pBlWyXwA08, BAL09, pBhSzZ09, BLS05, BS01, BC03, BCSV06, BP06a, BDV08, CGN01, CS03a, Cao01, CCH02a, CT03a, CFN04, DDI03, DHK06, Dra08c, Eic04, Eld06b, Eld09, Fur05a, Han03, HPP00, Kle09, LL09b, Lin03, LL08c, LWD09, OWY07, QH09, RG03, RM01, SP09, TX09, WM02, WM03, WY06b, WY06a, WYZ07, WL09a, ZZ08a, Zho00, Zus05]. Lieb [Aud09a]. life [BS05a, BLN04]. lifting [DKX03]. like [BFH05, Cig07, FGM08, GVD08, Ose08b, SSW08, Ste04, TX09, Xu03, YH04b, YLZ09, ZZY06, vdW01, BLM03, Boj03, PZ00]. likelihood [HB00]. Lim [FFN^{+}07]. limit [CS04, GZ03, Guo03a, Guo08, Kar05a, Kir03b]. limited [VHK01]. limiting [Shi04]. limits [FKS02, Sak01, dBS08]. Lin [WF04]. line [ADL04, BD09a, BD03d, JLY01, KS06e, Sat06, Shi07b, dSF09]. Lineability [MFPPSS08]. linéaires [Fao01b]. Linear [APESS06, AT07, AL09b, AG04a, Ano02-28, Ano02-27, Ano03v, Ano03-31, Ano03-43, Ano03-46, Ano06f, AS02, Bal03, BL00, BS02c, BLS02, BGLS04, BGLS05, BGJS06, BBS08, Bon03b, Bru03b, Bu05, CZ05, CGXL08, CL02a, CCG01, CHP08, CHL01, DOS01, DKX03, FMPT04, FJ06, FMN04, Gav00, GM07b, Gut01, HL08b, JS01b, JS07b, KS00a, LP01a, LTT02, LRŠ02, LS04b, LPS05, Lim09a, MD02, Nie01b, NT07b, RP00, SP09, TR01, TS03, Tan03b, Wal05a, Wan03a, WFJ08, Yan07, Zha03a, Zha04c, Zha05b, ZYP06, ZgC06, AAA^{+}02, AKGC09, AB03a, AB03b, AGDR05, AFR06, AB05b, Amb07, AMZ06, AdO02, Bad08, BRM01, BGN08, BB00, BBS02a, Bak04a, Bak04c, BB04b, BB04a, BBÖ04, BBG06, BB07b, Bal07, Bar08, BPT07, BIL06, BT05b, BT05a, Bér05, Ber04, BD02b, BHRV02, BP02, Bol05, Bon09]. linear [BKO01, BK09b, BKK08, BK01, BRS00, BCS02, Brü09b, BM04e, BP06b, BBP03, BGGC06, But03, CD00a, ÇS02a, CLPT06, CCH02a, CP03, CHASG02, CDMGP06, CJPS02, CCG02, CS08a, Cen06, CBIXG08, CEK^{+}01, CSY01, Che02a, CHRS03, Che02b, CN02b, CL03b, CCZ06, CPPS00, CHP07, CM08a, CQ08, CHJP05, CGH04, DL04, Dam02, Din03, Dod07, DS08c, Doe07, Dra08a, Dra08b, DLM^{+}00, DH04, Elf04, EK08a, EGMG07, Fao01a, Far08, FLMR00, FL04, FP02b, FS04b, Fre08, FS02, GA06a, GRB04, GH00a, GPM01, GPMSZ06, GPFC04, GJ04a, Gem03, Gew03, GJR03, God09, GZZ00, GQ09b, GSR03, GST04, GP00b, GM04, GC04, GL05b, GLP00, Ha03, HHY04, HvS07, HR07, HG09, HC08, HDA09, HC03b, HZ04b, HZM02, Hua00a, HM04, HvP05, Huh06, IPP09, IT07a, JP04, Ji03a, Ji03b, JW06, JK06b, Kak08]. linear [KKP07, KRS09, KS06a, KN03a, KR06, Kye03, LT04a, LK03, LCF06, LV03, Leš00, LLM03, LZ09c, LT04b, LWWZ08, LWZ09, LN06, Lip08, Liu02a, Liu03b, LL06b, LwW07, LDS07, LDS08, Lom07a, Lom07b, Lom09, LCP02, LS02d, Luz07, MM05, MH07, Map07, MM07, MT02, MS06, McC05, MMX04, Mei06, MST06, MND01a, MZK00, Mue09, Myš05, Myš06, Nah06, Ned01a, NB03, NZ06, Nie07c, OCK07, OCK09, Okt04, OdAJ02, OS05, Ots01, PV06a, PGRO06, Pet05, PA04, PK08, PCVT04, Pro08, Pry09a, Psa00, PPF02, QT06, Rab04, RY08, RSW05, Reu06, RA08b, RW06, RO02, RW01, Sal09, SdSW05, SdSW07, SdS09, Sch09a, Ser00, Ser04, SC07, Sha00, SSY05, Shi04, Shi06, Sid01, Smi01b, Ste01a, Stȩ00a, Tab01, TY06, Ter01a, Ter01b, TP09b, TW00, Tov00, Tre09a, Tru07]. linear [Uhl03, Val00, VHK03, Var06, VR06, VEB00, Vol04, Wan04, WY06a, WYZ07, WZ07, Wan07a, WB08, WGL09, Wat01a, Wie01, Wu06c, Wu08c, XH08, YWS00, YLX09, YH00a, YH00b, YLZ06, ZWPR06, ZQ07, ZGL09, ZG09, HT09a, Zav07]. Linear/additive [Zha04c, Zha05b]. linearise [Daw02]. linearity [YW07, Zha04b]. linearization [ĆDR01, Dra07]. linearizations [BMX08, Lop07]. linearized [NN01]. Linearly [EE04, HHL^{+}09, CŠ08b, CL02c, MŠ04e]. lines [CS02b, PPZ09]. linesums [EJLS02]. linked [BHJ08]. Liouville [CWW09, Kos00, LS09c, SC01, SW09]. Lipschitz [Bal07, BOP08, CS03b]. Lipschitz-type [CS03b]. List [Ano02w, Ano03-32, Ano03-33, Ano03-34]. Lists [Ano03-35, Ano03-36, Ano03-37, Ano03-38, Ano03-39, Ano03-40, Ano03-41, Ano04-40, Ano04-41, Ano06i, Ano06j, Ano06k, Ano06l, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, JDS03a]. Littlewood [FL07b, Ful00, OT01]. LLL [LT08]. Lloyd [Böt06b]. LMI [FS02, OP05]. Local [AMZ07, AMZ09, Liu02a, Pul07, ZJC04, ZYP05, ZFW06, ZPY06, BBL^{+}05, Bör07, Bra07, CS09a, Chu07, GM07b, HS09, Ish06, KP06, Ma09, MP06, MZ00, MOS07, WY04, XL04, ZY02]. locality [LMT09]. Localization [LN04, GML09]. localized [BES01]. Locally [MŠ04e, AL09b, ATV09, Cap03, CŠ08b, DvdV00, Kah04, Pac02, vB09b]. location [KPT01, MPS06]. Loewy [Alt09]. Log [Fur09, AB07, Dug08, Per07, Wan03a]. log-concavity [Wan03a]. log-convex [AB07]. log-hyponormal [Dug08]. log-majorization [Per07]. logarithm [CL01a]. Logarithmic [BES02, Igl01, Fur05a, HL04d, Ose08a]. logic [Tab01]. lollipop [HLZ08]. long [OS00]. look [CL03c]. Looking [SdL03]. loop [AMZ04, Dod07, Zus05]. Lorentz [Ell04, Hil07a]. Lorentzian [Lim07a]. Lorenz [Lim07a]. loss [HG09]. lossless [PHO07]. Low [DD09, EOvdD03, Li06b, MR02c, Zus05, AMLM^{+}09, pBlWyXwA08, BM01a, Bou07, Bra06, CFP03a, D– D05, Hil08a, LCN04, MV07, ZW03]. Low-dimensional [Zus05, Hil08a]. low-rank [Bra06, MV07, ZW03]. low-resolution [LCN04]. Lower [BZ00, CW07a, CW07b, CKW08, JN06, LHSL07, LzZT07, MK05b, Ton04, BH08b, CCRS08, Che04a, CL01b, Das04b, DSW07a, Gre09, KP01, LCW09, MK03, MV04, Tia02, Van07, WV08, Wan03b]. Löwner [And04b, Gro01, KP09, Zha00, fur08b]. LQ [BF07b, FN07, Sch09a]. LS [Pug04]. LU [ABH07, ADT06, BJ07, DJM06, MG03, Pan00]. LU-equivalence [ADT06]. LU-factorizations [ABH07]. Lucas [Lee00]. Ludwig [Ano06h, FGH^{+}06]. lumpable [FLQ04]. Lyapunov [AS01, ACK05, CL03c, CL09c, Dam03, GP00a, GX09b, JR06, KN06a, KS06a, KMP00, LS07a, LS07c, LŠ09a, Log05, MSS07, McC05, Neu01, OP05, SS07d, SS07e, Sty02a, Sty02b]. Lyapunov-Schmidt [Neu01].

M [BM04d, Ges02, Hog00, Neu00]. M-matrices [Neu00, Ges02]. M-matrix [Hog00]. M/G/1 [BM04d]. M/G/1-type [BM04d]. MacDuffee [BV02]. machinery [LL06a]. MacMahon [FH07b, Min01]. Magic [LCTS09, GS09d]. magnetized [GG09]. Mahalanobis [GH06]. Mahler [Per07]. Main [Ano05-31, Bor06, HLZ09, dF00]. main-diagonal [Bor06]. majority [Ste03b]. Majorization [BD03c, AS03, BL00, BLL03, Bon03a, Bul08, Dah01, Dah08, Fur09, HR07, HP01b, KS07a, LP01a, LLY09, Nie01b, Per07, PMS05]. Majorization-constrained [BD03c]. majorized [Ful00]. makes [GV09b]. Malamud [FNS01]. Malcev [BP06a]. Manhattan [CDFM08a]. manifold [Arv08, Chi02, PPZ02]. manifolds [AMLM^{+}09, IK06, LST08, Pue05, Zus05]. manipulation [SLP01]. many [CH00a, DS08c, Tre04f]. map [Bal07, Bre03, CS03b, CKM05, LRŠ02, MPS07, Nie04b, Tam04, Tka05]. mapping [KN03a, MHZU05, Yag01]. Mappings [DKL^{+}07, AGK06, Bae07, BD03a, BS00c, CL02b, vGKN09, JLL03, Kuz02b, KO06, Lim05a, Lim05b, Lim06, Lim09b, Lu02b, Pet02, PS03, Pet05, RST01, Rad03b, Ser04, ZX05b, ZX07]. Maps [CH08, CL09d, Dol03, FJP07, LŠS07, Mol09, Šem06b, SZZ07, WY08, YZ09, AT07, AH02, AHZ05, AH09a, Aud09b, BBS08, BS09a, Bon09, BBK03, Bu05, CCH02a, CLL05, CFL05a, Che02a, CL02a, CLP03, CD06, DMD03, DP02a, DH04, DHB07, EdRM04, FJ06, Foš05, GM07b, Ha03, HHY04, HC03a, HC03b, HZ04b, HZ05a, HQ08, HJ08, HHZ09, HBLZ06, Ji03a, Ji03b, JW06, JL09, Ji09, KNR05, Kye03, LV03, LS06b, LTT02, LLM03, LJ04, LS04b, LL05b, LPS05, LP07, Lin03, LL08c, LL04, LN06, LDS07, LDS08, Lu02a, Lu03, MST06, Mol02, Myu09, Nie01b, Nie07c, Nie08, NT09a, Pei07, PR04a, Poo08, Reu06, RA08b, RSW02, Šem08, SC08, Ter01a, WFJ08, YLZ06, YW07, Zha04b, ZYP06, ZZ08a]. Marchesini [BKS01]. marcus [XXZ09]. margin [dDG02]. Marginal [Nah06]. Mark [Böt06b]. marked [CFP03b, CFP04, CP07b]. Markham [Yon00]. Markov [BLN04, Ben04, BM04d, BD09a, BDKW04, CNX05a, CZ08, CNF08, CM00c, CM01, CDS08, DMM00a, DHFS00, DHZ03a, GAM06, Gue08, GL05b, GS09d, Har06, Hun05, Hun06, Hun08, Hun09, KN00a, Kir04, KNX07, MS04b, MSF05, NX04, PW04, RA08a, THH04, VWD08, VM04, Xue01, Yan02a, Yan09a, ZZ00, Zha04e]. Markovian [HLR08]. masks [CGR09]. Master [FH07b, Min01]. matching [CNSM07, FYZ07, GYY09, HL02, MW01b, SS09c, WK09]. matchings [BH05, CT03b, CG07a, GLL09, GT02b, GT04, Han07, MN03]. mathematical [Abe08, CPPS00, Log05, Tab01].

Mathematics [Moh05, Per03, Kol05]. Mathias [Lim08b, Yam06]. Matrices [AHST09, And07b, Ano03-45, Ano09q, Böt06b, BD03d, CHRS03, CDJ08, EMS06, Fer05, GHS09, JMP09, LRW08, PS08, Par09b, Sev08, SS04b, SS01, SH02, Tak04, Var06, WW07, AMP07, AG00, ALSS08, ACM03, AM02, ACV09, AC05, ASX08, ALdV05, AdLV06, AFM03, AGKM06, AGO09, AHW05, AB03a, Ala05, AB05a, AH08, Alf00, AW02, Alf06a, Alf06b, ARR06, AEI00, AI00, Alt09, AKYD16, ALMS06, Amb07, And04c, AMLM^{+}09, Ano03-44, ACS04, ATU06, AT09, ABH07, AP03a, AM04, AG05a, ADT06, Aud06, Aud08, AH04, AHH^{+}09, ACK05, BD02a, BD03a, BB00, BS02a, BBL03b, BBT03, BH04, Bak04c, BB04a, BBÖ04, BBS04, BHLL04, BB07b, BHJ08, BB07c, BFL05, Bal00, BDK08, Bap00, BJT04, BKN05, BR09, Bar01, BB03a, BM01a, BP03, BW00, Bas02, BE01]. matrices [Bau03, BR00b, BLS02, BSKS05, BN08, BNdP^{+}05, BdPT08, BV02, BLS05, BBS06, Bel05, BEG^{+}07, BOP08, BC00, BCEG09, BFS07, Ben05, BD05, Ben08, BB05c, BBD^{+}08, BZ00, BX04, BSL04, BX05, BLLS06, BGM04, BO05, BMGS06, BM07, BD01a, BK07, BKS03, BGM01, BDG08, Bir04, BJ09b, Boc03a, BPT09, Bol04, Bol05, BKR03, BKR04, BS04b, BCPT05, Bom08, Bon09, BF08, Bör07, Bor06, BC07, BC08b, Bos02, Bos07, BW01, Bot00, BES01, BG02, BES03, BW05, Bou09, BEH^{+}06, BL02, Bra03, BOvdD04, BQW09, BPS04, BCGM08, Bru02b, BD03c, BD07, Bu05, Buc06, BM06b, BFJ07, BJ07, BKN01, BAD09, But06, CSZ01, CW03a, Cal08, CM00b, CMV03, CRU09, CW00a, Cao01, CT03a, CGM05, Cao07a, CS09a, Cap01b, CZ07, CGD05, CHNP04, CNX05b, COvdD09, CP07a, Ceg01]. matrices [CG00b, CLB09, CP06, CA02, CR08, CD00b, Che03, CT04, CL04a, Che04a, Che04b, CC04, CMX05, CHN06, Che07b, CW07a, CW07b, Che07c, CNSM08, CKW08, CZZ09, Che09, CHL02, CT03c, CLL07, CJVW07, CH07, CL08a, CL08b, CL09b, CS00a, CHSS03, CHR^{+}03, CB06, CLOvdD01, CT01, CFL02, CN01, CN02b, CN09, CHY05, CNTV08, CK03, CJK05, CL04b, CDS08, Cho04, Co08, Chu03, CBD^{+}04, CM09, CL00b, CL03c, CL09c, CFP03b, CS04, CH05, CS01, CP07c, CP08, CFJ01, CS00b, CDS01, CS03c, CS03d, CV07, CI08, Dah00, Dah03, Dah04, Dah05, Dah08, Dah09, Dan06, Dan02, DoMP09, DL92, DL01, DS04a, DLS08, DMM00a, DV07a, DV08, DMMV09, DK06, DD04, DM03, DP04, DJ01, DPZ06, DHK06, Dod05, Doe07, – DI02, – DZ04, – DS08, DO09a, DO09b, Dom00, DA10, DMM00b, DJ06]. matrices [DH09, DJL01, Dun04, DB06, DB08, EMT05, EG01, EG03, EGO05a, EGO05b, EG08, EFS01, ES08, EFN^{+}03, EH03, EvdD04, Els05, EH07, E^{+}00, FGM03, FN01, FFM03, Fal04, FL07a, FJ07, FH07a, Fan02, Fan03a, Fan04, FJ06, Fan07a, Fan10, FT05, FGR04, FT00, FG02, FI03, FI05, FI07, FI08a, FLM02, FP07, Fer07b, FS04a, FS06a, FM00a, FM00b, Fie01a, Fie01b, FM02, Fie03a, Fie03c, Fie04, FV04, FM04a, Fie05b, Fie05c, Fie06, Fie08, Fie09, FKS02, FS03, FGL^{+}00, Flo09, FNS07, Foš05, FGR07, FNV08, Fre04, Fri06, FF02, Ful00, FIS02, FJ04, FKY09, GS04a, GA06a, GH03, GL01, Gar03, GT02a, GT08, GW04a, GW07, Gav03, GP03, GR00, GJ04a, GVD08, Gem08, Gen00, GŞV02, GI02, Ges02, Gha01, GN00, Gil01]. matrices [Gla04, GKvS04, GN03, GKW00, GHM03, GH09, GT09, Gre09, Gri00, Gri04, Gro01, Gro06, GP07, GL09, GZ01, GZ03, GG07, GS00, GS09d, GGCB07, HNT03, Had04, HLW01, HLR04, HL04b, HL04a, HLG04, HL04c, HLW05, HL07a, Han07, Hao04, HPP05, HTK07, Hei01a, HR01, Hei02a, HR03a, HR04, Her03, Hia09, HP09, Hil07a, Hil08b, HH05, HSZ08, Hog03, HS07b, Hol08, HS02, Hol05, Hon02, Hon06, Hon08, HP04b, Hor02, HS06b, HS08, HL01, Hou03, HZY05, HLS00, Hua00b, HSL05, HW05b, HT05b, HC06, HBLZ06, HW07b, lHH08, Hua08b, HMT^{+}08, HST08, HP01a, HKKZ03, HP04c, HL04e, Ikr01, Ikr07b, IHM08, Ips00, JT04a, Jef05, JSU03, JS01a, JS01b, JDS03a, JO04, JMT05, JO05, JR05, JS07b, JS07c, JDS08, JTU03, Jør05, JPB08, JB08, Kap00, Kar05a]. matrices [KSU03, Kar07, KV06, KH09, Kes01, Kes07, KTR08, KS02, KR04, KSH05, KS08, KN04, KPT01, KNX07, KP08, KS01, Kle04, KS09, Kle07, Kob05, KD07, KR03, KPV04, KR06, Kol02, Kol03b, Kol09, KH01b, KH03, KH08, KP02, KMS00, KMS01, KLMS03, KMS05b, Kra01, KT08, KMS05c, KM07, KS06d, KLM01, KS07a, Kur04, KC03, KO06, KZ03, KLY09, LŠ07b, LS07a, LŠ08a, LŠ09a, LMMT01, LR02, Las06, LZ08a, Lás07, LL08a, LM06a, LX08, LN03, LM06b, LS00b, LP01b, LM02, LHS02, LTT02, LS02b, Li03, LS04b, LR05c, LL05a, LHS05, LSD06, LW07, LP07, Li08a, LR09, LB04, pLLfY06, Lim09a, LCN03, LT04b, Lin02a, Lip01, Liu02b, Liu03a, LHZ04, LH04, LY06, LHF06, LYY07, LSH08, LLHK08, LC09, Liv04, LT03b, LM04b, LM09b, LMT09, LQ00, LPG03, Lur05, Lur06b, LM09c, Mae08]. matrices [Mai01, Mal09, MB07, MS00a, MM03, MSS07, MP05, Mat04, MW09, MW03b, McM09, Meh00, MR01, MR02a, MMX04, MRR07, Mel09, MS04d, MPP10, MQ06, Mir00, Mir03, MSS08, MND01b, MND01a, Mol08a, Mol05a, Mol08b, Mol04, Mou06, Mul05, MR02c, Nab07, NDS02, Nah06, NY07a, NBdP05, NM04, Ned01b, ND05, NT02, NS08a, NRW02, NS03, NS02, Nie08, Nou06, O’L03, OV06a, OPS07, OZ03b, OZ03c, OS01, OSvdD02, OS00, OP05, Olk02, OOT06, OV06b, Ost03, OWY07, Ovs03, Pac02, PRW01, pPyHZ05, PW08a, PRS02, Pat04, PP01a, Pat01, PR01, PP01b, Peñ03a, Peñ09, PW02, PHZ06, PV06b, Pet02, PS07, Pet09, Pfa08, Phi04, PES07, PR02c, PR04b, Pop01, Pro05, PH09, Pry09b, PT03, PT06b, PH00, PG04, QD09, Rao09, RS03a, RS06a, Rob02, RS05a]. matrices [Rod08, RS03b, RR04c, RS09, Ruk01, Rum03a, Rum03b, Sak01, SS06a, SS07b, SS06b, Sat05, Sav03, Sav06b, Sbu05, Sbu08, Sch04, Sch02c, Sch09c, SXKM03, SSA05, Sem07, Šem03, Ser09, Ser00, Ser04, Ser08, SH00a, SH00b, SS02, SHS03a, SS05a, SSY05, SLR06, She00, She01, SZZ07, SS09d, Sil00, Siv09, Šmi04, Smi01a, Smi01b, Smo08, SHRC03, SKBS08, Sot03, Sou05, Spi00, SZ08, SK00, Ste06a, Ste05a, Ste07a, Str03a, SBS02, Stu03, SM09b, Tad06, TR01, Tak00, Tan01, Tan03a, Tan05a, Tan07, Tan08, TY06, TSOA07, Tao06, TRH01, Tar05, TH00, Tia00, TS01, TT09, Ton04, TE05, Tre02, Tre03b, Tre04a, Tre04b, Tre04d, Tre04e, Tre05, Tre06, Tre08, Tre09a, TL03, TS06, Tsa01, Tyr04, Uhl01, Urb04, Val02b, VPV01, Var03, VLA03, VLA04, VS04a]. matrices [VS04b, VS05, Ves00, Wal05b, WW08a, Wan06, WZZ00, WB01, wW02, WM05, WA09, Wan09c, Wat05, Wat06, WG02b, WD05, WLBZ06, Wer08, Wim01, Wir09, Woj03, Wu05a, Wu06a, WD06, WLL07, WZCL08, Wu08a, XX03, XZ06, XCX01, XHpS06, XS04, Xu04, XQ08, XS08, XXZ09, XLG^{+}13, Yah04, Yah06, Yam01, YYZ05, YC00a, YH04b, YH04a, YLZ09, Yan09b, Yin04, Yin05, YT04, YLS05, YSS07, Zal03, ZM03, Zen01, Zha07a, ZJL02, Zha03a, Zha05b, Zha05d, yZtLC06, yZXtL07, Zha07b, ZZJ08, Zha09, ZW00, ZHZ08, Zhe02, ZHZ03, ZXZ08, xZxGpL09, Živ06, dF00, dFP01, dSF09, dHM07, dlP03, vdH03b, vdMSR01, vdMRS06, Ano06g, FMN04, KS00b, Lin02b, Neu00, SH00b]. Matricial [CF02]. Matrix [Ano02s, Arp05, BHNR03, BdL04, BLL06, Bou06b, Buk07, CXZL02, Cap01b, CNX05a, Che05, Cig01, Cod00, CCM00, DS02, DZ01, Dod08a, Dod08b, FGT03, FP04, Hig08, HK00, Hog03, HS09, HP01b, HKSS03, JS04, KS09, Mei05, NT06b, OTZ09, Pap00, Pih04, Rod09, Ste05b, Tům06, Uhl04, VR06, XDS08, XDS09, Zyg02, dMR07b, vdD09, Abd08, AEEG08, AA09, AJF09, AFS09, ABW05, Ala05, Alf03, Alf05, Alf08, Ali09, ABB00, ABR02, AADL03, ADL04, Alz01, AMZ07, AY03, And04a, ATU03, ATU04, ATU05a, ATU05b, AHK^{+}05, AHLR09, AD08, Arp03, AG04b, AS09, AYC06, AM05, BMNA08, BZ06, BBS02a, Bak04b, BB04c, BB07a, BST09, Bap03b, Bap05, Bap06, BLP06, Bat06, BT00, Baz03, BLL03, BGLS05, BG09a, BI03, BBC00, BN05, BS03a, BS00c, BHdST06, BI00, Bén06]. matrix [BB03b, BT05a, BB09, BK00b, BH06, BK08, BLM02, BB03c, BLM03, BM04d, BR01, Bol02, Bor08, BR07, BD03b, Bot09, BU07, BLFS09, BL02, BFPW02, BM05, BT04b, BGO09, Bri05, Bri01, Bur05, BLO02, BCG07, BMX08, CC09, CCH02a, gCZ03, CW03b, CZ05, Cao09, Cap01c, CL01a, CGP04, CHNP04, CC03, CA02, CH00a, CC00, CEK^{+}01, CL05a, CX05, CK01, CK02, CNP02, CN06, Chi04, CL02b, CLP03, CL03b, CL05b, CL06, CDS04, wCL04, Cig07, Cof03, CK04, CSK01, CHL01, CM08b, Dah03, DLMO07, DH01, Das09, DSW07a, DH00, De 00, DDM08, DD09, DW09b, DS03a, DTGV06, DP00, DP01, Dob01, DP02a, Dol03, DJM06, Dra07, Dru02, Dru04b, Dru08c, DL08, DLT08, Dun07, DP02b, DD02, DV07b, ES02, EGG07, Eis06, ESP05, FJTU00, Fan07b, FP02a, Far04, FM07a]. matrix [FI03, FLROV05, Fen03, FT07, Fer05, Fer07a, FS06b, FG08, FVP06, FN00, FN03, FH04, FKL03, FKS09, FMS01, FJ03, FMS05, FJ05, Fur05b, GPM01, GPS01, Gav00, GJ04b, GZ07, Gil04, Gil08, Gin09, GR07b, GKL07, Gra05, Gra02, Gre02, GT06a, Gro00a, Gro00b, GTT01, GT06c, Gu01, GX09b, GZ08, GWW06, Gut01, GM09b, Ha03, HHW00, HLT04, HH00, Han02a, HT07a, HP00, HR07, HIU04, Has05, HES06, HMS01, HB00, Haz05, HG09, Hei01b, Hei02b, HWKS03, HT03, Hil08a, HV08, HLY06, Hog01, Hog02, Hol02, HP03b, HS04, HS06a, Hor04, HC04, HL04d, HdABC07, Hua00a, HS01, HL07c, HL08b, HL08c, Huh07, HKKL05, Ikr07a, IHM01, JH09, JMT04, JR06, JW06, Jia00a, JW03, Jia04, Jia05, JMN08, JK06b, JS01b, JLY01, JOR01, JD02, JO02, JDS^{+}03b]. matrix [JO05, JKL09, Kal03a, KPK04, KKP07, KKP08, KV09, KR05, KYAA08, KY00, KN00a, Kir00, Kir03a, Kir06, KMZ00, KN00b, Kle00, KS01, KMS05a, KS05, KS06b, KS06c, Kly00, KB02b, KAP^{+}03, Kro01, KRS03, KS04, Kye03, LR05b, Las06, LL06a, LK03, LN08, Lee09a, LPR04, LS06b, LK09, LRY05, LRŠ02, LTW02, Li05, LS06c, LHSL07, LCW09, LZ09b, pLLfY06, Lim05a, LT07b, Lim07b, Lim09c, LgS01, Lin08a, LG03, LT07c, LSF09, LTT09, LM06c, Log05, LR04, LP00, Mal01a, Mal01b, MS04a, MS01a, Mar02, MP00, MZ05, MT02, MVV08, Mat05a, MP05, Mat07, McD03, MM08, MNV08, Mei06, Mel00, MHZ05, MK03, MGWN06, Mol05b, Mon08, Mor08, MN07, MR02c, NP01, NBdP08a, NBdP08b, Nak03, NV03, NB03, NCV05, NCV08, NT09c, OS09, OA09, Ohn08, Ois01]. matrix [OP00, OT09, Ost06, PZ00, PFL05, PG01, PP08, Par02a, PB06, Par02b, PV06a, PEPS00, PPP06, yPyHZ03, yPyH03, hPyHZ07, PMS05, Prz02, PT04, Pug04, Qi09, QC04, Rab04, RRY04, RK00, RL03, RR02, RR04a, Rao00, Ren01, Reu06, RS00, Roc09, Rod07, Roj02, RP03, RS05b, SS06c, SvdD05, Sav05, SS04a, SSMQ^{+}00, SS03a, SS09b, Šem06b, Ser07, Ser00, SS03b, Shi04, Shi01, SHWR02, SV01, SvW05, Sob08, Son00, Son02, SvKHO09, SvR02, SZ04, Sv05, Stȩ00a, Str01, gSX03, gS03b, SDL06, Sza07, TŻ08, TR02, TR03, Tan03b, Tan05b, Tia02, TS09, Tik06, Ton00, Tov00, Toy09, Tri02, Uch02, VC04, VPV01, Van04, Van07, VWD08, Ver08, WV08, Wan02, WY04, Wan04, WY06b, WY08, WL09a, WL09b, WvdWC09, WG02a, Wil09, Wil07, Win00, Win03, Wri06]. matrix [WJL02, Wu05c, Wu05b, Wu06b, Wu08b, Xia03, XD03, XM02, Xu01, Xu03, YMP01, YTI06, Yan01b, Yan02b, YM07, jYxL09, YLX09, YLF09, Zav08, Zha04a, Zha06, Zha00, Zha03b, ZW03, ZyC03, Zha04c, Zha05a, ZgC06, ZH02, Zho03, ZD05, ZXZ09, dSZ00b, dlMW06, vDHK07, vdW02, vdW03, Cra09, Uhl00, Hog00]. matrix-functions [BN05]. Matrix-traces [CXZL02]. matrix-valued [BHdST06, Bol02, CH00a, FKL03, Las06, Wu05c]. matrix-variate [Mat05a, MP05, Mat07]. matrix-vector [FI03, Hor04, Mel00]. matroid [Fer01a, Fer04]. matroids [Lem04, LRWW07, Vil08]. Mautner [Sud02]. Max [But03, EvdD04, Ser07, Ser09, BV07, BSS07, BCG07, But08, Cec00, CCG02, CGB04, Dan06, De 00, EvdD01, EH07, EvdD08, GK07, Gav00, Gav02, Gav03, GP03, LT04a, Lur05, Lur06a, LY09b, Mol05b, Myš05, Myš06, NS07b, NS08b, NS08c, PG01, Pep08, Sem06a, Sem07, SSB09, vdW01]. Max-algebra [But03, EvdD04, BCG07, CGB04]. max-algebraic [But08]. max-min [Cec00, Gav02, Gav03]. Max-plus [Ser07, De 00, GK07, Gav00, LT04a, Mol05b, NS07b]. max-separable [CCG02, Myš05, Myš06]. Maximal [KKABAF06, ORvdD02, Rie00, BB03a, BN08, BC04b, CS05, Cve02, Das09, FGR07, GWL09a, GWL09b, Hae08, HGW02, LZ07, MRR07, PR00, SXKM03, gSX03, TFZ08, Tia00, Wan09b, Woj03, Xu01, ZH02]. Maximizing [Dru04b, ZSL09, ZWS09, dKNvD08, TFZ08, vD07]. Maximum [HL07a, BFH05, BKK08, CHMM06, DHH^{+}06, Dra08a, Fer07a, FGS05, GJMR01, HB00, JDS08, Kal03b, Lav03b, Lav05, LM09a, Nik09, SM09b, Wan03b, YL08, vdH08a, vdH08b]. maximum-likelihood [HB00]. Maxwell [VB09a]. May [Ros06, ES08]. Mayer [PLG02]. McClelland [Rad09]. McMillan [GM07a]. MDP [HST08]. Mead [HNX03]. Mean [BK07, BK00b, Bha06, BK08, CE09, CM00c, CCW^{+}08, FNT05, FFN^{+}07, HLW06, HZY05, Hun05, JHL05, JLY09, Lim05c, Lim08b, Mol09, MR02b, PW04, RL03, Yam06]. mean-tail-dose [CCW^{+}08]. means [AM09, AL09a, ALM04, Aud07, BH06, HP09, HK02b, KL07a, KMZ00, KS04, Lim05c, Sch00a]. measure [Mat05c, MFPPSS08, Per07]. measured [CM00c]. measures [DS02, DS03a, HH05, HH09, MBLV01, TZ01]. mechanical [PR02c]. mechanics [BdL04, Fur06a]. mechanistic [PSvdW05]. meet [IHM08, KH01b, KH03, KH08]. meets [SP09]. Mellin [HFR04]. Mellin-type [HFR04]. members [CHK^{+}01]. memoir [Ros06, Sch02b]. MEMS [LRK06]. Mercer [MPP06]. Merikoski [AKYD16]. Meromorphic [Pop02, BES02]. mesh [Wd09]. meshes [ST07]. messages [CKZ05]. method [Abe08, AEEG08, BF01, BS06a, BX08, BBMX07, Ben04, Boj03, BLR00, CS08a, Dah05, DH01, DH03, Dax03b, EG01, EGO05b, Elf04, EvdD01, FK04, FS08, GM09a, GT06a, GX09b, Gug08, Had04, HNX03, Hoc09, Hol00, JR09, KYAA08, KH01a, Kos00, Lan05, LCF06, LMV03, Li01, Li03, LA05, LMT09, Mat02, MH06, MSW09, MPS00, MPS03, NN01, QX08, RS06b, SK08a, Sch00a, SK08b, TI08, VHK03, Vos07, Yam01, ZZY06, fur08b, vdW01, Bag08]. Methods [Ano03-43, Chi02, FMN04, AAA^{+}02, BRM01, BW03, BGN08, BW08a, BKPW08, BLO02, CLR00, CRS03, CW02, CDS04, CHP06, Dek09, DTGV06, ESP05, FH06, GGL08, GS03a, GL06, Gra04b, HV01, HLY06, Hog01, Huh05, JH09, JMT04, JR06, KKP07, KHG04, LS00b, LS01, LA05, LWZ09, MS00a, MM03, MS04b, MP06, MN08, MR02b, OS05, Pet08, Plu01, Pul07, Rih01, RT01a, Rum01b, SL06, SP04, Sol06, TR03, Wan07a, WB08, XZ06, Zem07, ZBY09, dSQS00, vdVN00]. Metric [Fin00, PPZ09, BT04a, Bha03a, Dit00, HHT06, JNL05, Jai06, Lim07a]. Metrics [FF04b, And07a, Dit00, GMR^{+}07, HP09, ZQ07]. Micchelli [Bar05]. Mićić [FG03]. microarrays [FNC06]. middle [KS06e]. midpoint [LL08a]. mildly [BW03, WM05]. min [Cec00, Dan06, Gav02, Gav03, GP03, NS08b, NS08c, Sem06a, Sem07]. Minc [CW05]. Minimal [AD06, AMLM^{+}09, BB03b, Bru02b, CMV05, CŠ08b, CLY07, Dam03, FZ04, HMH01, Kun06, OS09, Ost03, PR02a, ÁFMNW06, Aga02b, AVV05, BFH04, BCRS08a, BCRS08b, BW01, Cal07, Dob01, GI02, GW02, Guo02, HW07a, KVV09, LPR04, LL09a, Lip05, MR09, OV06b, RT03, RP00, SH00a, Tia00, Val02b, Var03, VJ06, Wan09e, WK09, Wie01, YSL08, ZJL02, vDK07]. Minimal-volume [Ost03]. minimally [Ma09]. minimax [Xia06]. Minimization [Tre04e, CW00b, FS06b, Sav06a]. Minimizing [FWG08, FYL06]. Minimum [BJ07, DHH^{+}06, YL06, AHK^{+}05, BF07a, BFS08, BGL09, BFH^{+}08, Che04a, FKP03, FH07a, FP07, GWW06, HHL^{+}09, Hog08, JMN08, KSH07, KR08, LHSL07, PBA09, WG02b, XHSZ09, BGL09]. Miniversal [CFP03b]. Minkowski [Chu08, GK07]. minor [Bal07, GT06b, JLMC05]. minors [BF08, Chu03, EMBV03, GT06a, GT06b, HK03, HvdH07, JS01a, KMS01, KLMS03, KMS05b, KM07, Mal01a]. minus [BBL03a, BH04, BJ09b]. Miroslav [Ano07n, BHN^{+}07]. missing [FNC06]. Mixed [HZY05, Uch02, WWJ09, AD04, Fan03b, Jia04, Kle00, yPyHZ03, Rod07, ZL02, ZL03]. Mixed-type [WWJ09, yPyHZ03]. Mixing [Hun06, Hun08, Hun09]. mixture [Kle04]. MixturEs [AFCC04]. ML [Lin02b].

ML-matrices [Lin02b]. mode

[Ois01, Ste09]. Model

[AD05, GVV03, GA06b, Lan08, RW06, WSW06, AdO02, BS06a, CNX05a, FKLN07, Fre08, GH00a, GP00b, GM04, Gug08, HL05b, IPP09, Kir05, LCF06, LRK06, Mat07, Pen06, PSvdW05, SS09c, SdL03, Sty06, TVB08, TP09b, ZZ00, ZGL09]. model-order [LCF06]. Model-updating [Lan08]. modeling [BF01, PA04]. modelization [CHP08]. Modelling [smbH02]. Models [Ano06f, BDE03, Dra08c, BB04d, BDKW04, CGXL08, DKS04, FMZ06, FK04, GA06a, HDA09, Jia04, KMZ00, KN00b, LS06c, Liu02a, Lom02, LS02d, MD02, Nah06, Peñ08, Smi01b, SvR02, THH04, TBK09, TP09b, Vol04, Wu08c, Zen01, dMFZ09]. moderate [JN06]. moderate-density [JN06]. modern [Vin00]. modification [EG01]. modifications [BB04c, Bra06, HNT03]. Modified [LS00b, BBT03, Bo02, Dax00, Yos07]. Modifying [EvdD01]. Modular [Cur07, XCX01]. Modulated [Ano06f, CBIXG08, CGXL08, HC08, SK08b]. Module [VZ02, Pic07, Woo02]. module-behaviour [Woo02]. modules [AR06b, BS03b, CZ05, Dom03, FYY09, FP02a, vGKN09, Gun03, HJM09, Har07b, Ish06, OZ03a, OZ03b, OZ03c, Ovs03, PR07, Sha09, Ter02, XS08, Xu09]. moduli [Bad08, TW00]. modulo [CGP04, PW08a, VLAOVL08]. Modulus [HT09a]. Moerbeke [BT04b]. moment [CH00a, CH01, DS02, FKL03, HC04, JKLP03, MS01a, NT02, PH09, Pue07, Tka05]. moments [GN00, LRW08, Mei05]. Mond [MPS00, MPS03]. Mond-Pecarić [MPS00, MPS03]. Monge [Fie01a, Fie06]. monitoring [Bol01]. monoid [CGGGSR01]. Monoidal [Gun03]. monoids [RGG01]. monomial [EMBV03, GLS07b, SV06b, Vil08].

Monotone [BB05c, Gav02, SV01, Auj03, BJ09b, CKMS02, Dit00, Fur02, GHI09, Han09, HZ02, OT09, Sza07, fur08b]. Monotonicity [Fur08c, Hia09, Lav03b, BT07a, Dru08c, GRB04, GSR03, IK09, JS07a]. Moore [PP04a, BBT03, BB07a, Bap00, BOvdD04, Fie03b, FGM06, Gro00a, KDC07, KS07b, Li08d, MD09, Pat01, Pat03, XH08, Xu09]. Mori [Dit00]. morphisms [GV09a, YC01]. Morrison [Map07]. Moser [CL03a]. most [GOTR07, LM09a, MCEC06, PBT04, Ste03b, vdH03b, vdH08b]. motion [TKX08]. Moussa [Hil07b, LS09b]. Moving [HS07b, KS01]. MST [LRK06]. Muir [Ano03u, FL07b, Mar05]. Multi [Bos02, AMM06, Bos07, CWW09, CtBDC09, Kes01, KFGK04, MN08, WA09, ZTK09, vdMSR01, vdMSR03]. multi-agent [ZTK09]. Multi-companion [Bos02, Bos07]. multi-criteria [KFGK04]. multi-index [AMM06, Kes01, vdMSR01, vdMSR03]. multi-interval [CWW09]. multi-level [MN08]. multi-spectral [WA09]. multi-way [CtBDC09]. Multialternating [GZ07]. multicollinearity [ÖA00]. multicomponent [GG09]. multidimensional [AT08]. multifilar [EFN09, EFN10]. multifilter [Jia00b]. Multigrid [FH06]. multiindex [vdMRS06]. Multilevel [Tre05, Wd09, Cap01b, FT00, OOT06, Tre04b, Tre06, Tre09b]. Multilinear [GN03, BS06b, DV04, DV09, MT05]. multimode [TL02]. multiparameter [HP03a, Koš03]. Multipartite [FGM06]. Multiple [CH01, DP04, DJM06, Kel08, Yer04, AA09, DMW08, GV09b, JDS^{+}03b, KM04, KV06, KS06c, KS09, PP08, Rum01a, Toy09, Wan02, Wu05c, XD03]. multiples [Dru07]. multiplication [BZ06, CS01, McE02, Mel00, SC07]. Multiplicative [CFL02, CDL05, Har07a, Lu02b, Mol06b, dlMW06, Alt09, AH02, BPS04, DR05, DJ01, Fan03a, Mol02, O’u02, Poo08, YLX09]. Multiplicativity [KNR05]. multiplicities [JD02, JDS03a, JKO04]. multiplicity [BFH05, CHMM06, DHH^{+}06, Fer07a, JDS08, MZ06, Pic07, dDG02]. Multipliers [Joc09b, HN04]. multiscale [AAV06]. multisplitting [BRM01, BW03, LS01]. multisplittings [WM05]. multistage [Vol04]. multistate [LS06c]. multivariable [GPM01]. Multivariate [FNT05, KP09, RvdGM05, BL00, CMX05, CNF08, GA06a, Har00b, Liu02a, NT02, TP05, Vin00, Wu08c, YTI06, Ziz04]. Murdock [Tre02]. Mutually [PR04b, BGS09, Har05, Obl08].

N [Böt06b]. natural [De 07a]. near [CNR00, CL05a, CM09, Neu01, Tov00]. near-exact [CM09]. near-periodic [CL05a]. Nearest [Gra05, GV09b, Ala05]. Nearly [SH00b, DHFS00, HB00, HL01, Li01, LY06, MSF05, Rum01a, Xue01, YLS05]. nearness [HTV02, pLLfY06]. nécessairement [BCSV06]. necessarily [BCSV06]. Necessary [VV06, Yan03, TY09]. necessity [SH05]. negative [ATU05a, CRU09, FM07a, Gro01, GG07, HV08, LŠ06a, Liv04, Mik09, NS02, Nou06, Pat04, TRH01, wW02]. Negativity [BMS04, Lad08]. neighborhood [Li05]. Nelder [HNX03]. NEPS [Ste00b, Ste02a]. nest [CGH04, FLRZ02, HC03b, HZ04b, HJ08, HQ09, LL05b, LL04, QH09, ZJC04, ZYP05, ZFLW06, ZFW06, ZYP06, ZWC06, ZX05a, ZX07]. nested [Sta09b]. network [FLPW03]. networks [AD05, BBFP04, BSL04, CDFM08a, DN00, EMT05, EH09, RT06]. Neumann [CTW01, CL09d, Dru08b, GS04b, JY07b, Kom08, Mol06b, NY02, Oik08, PP01a, Pat01, Rad03b, ZJC04, ZYP05, ZFW06, ZYP06, ZZ08a]. Neumann-type [CTW01]. neural [EMT05]. neutral [FYL07]. Nevai [YMP01]. Nevanlinna [ADL04, ADL06, BHdST06, Bol02, CH00a, CH01, GM07a, HdABC07, Wu05c]. Neville [ACHR01, ACDR04, CP07c, GT02a, Gem08]. Newton [BX08, CKMS02, DH01, DH03, HNV00, HLR08, JMP09, Sch00a, PRW01]. Newton-type [BX08]. Neyman [IL06]. Nicholas [Par09b]. nilalgebras [CHP03, CHJP05, HL05a]. nilindependent [BPP08]. nilindex [HL05a]. nilpotence [GLS07a]. nilpotency [CGH04, LŠS07, LDS08]. Nilpotent [CHJP05, Urb04, Bal03, BC03, BOvdD07, CM00b, DHK06, DO09b, GW08, GR00, HLW05, JLL03, Kah04, KOvdD07, LPG03, Obl08, Ovs03, RM01, Str03b, SM09b, Tan08, Zho00, FJL08]. nilradicals [BPP08]. Ninth [Ano03-46]. no [Hor02, KOvdD07, Omi08b]. nodal [Bıy03, BHL^{+}04, DH06, DGLS01]. node [MQ06]. nodes [EFS01, EF07, Li08a]. noisy [Bol05]. Non [Ash05, BM04d, DGN02, EdRM04, HV08, LN08, Lem04, ABDD01, BGN08, Bör07, BC04b, Bri01, CCRS08, Don01, ES02, Eld09, EK08b, FKP00, FM07a, GHI09, Gro01, GG07, Har04, Jai07, JTU03, KS06a, Kob00, Lad08, Lim08a, Liv04, Ma09, McC05, McC01, MN08, Pat04, SCS09, Sid01, Ste06a, Val08, WB01, wW02, WB08, Wan09b, XM02, Zha05c]. non- [BC04b]. non-acyclic [JTU03]. Non-associative [LN08]. non-autonomous [McC05]. non-bipartite [CCRS08]. non-commutative [Don01, Wan09b]. non-commuting [McC01]. Non-degenerate [EdRM04]. non-existence [KS06a]. non-Hamming [Jai07]. non-Hermitian [BGN08, SCS09, WB01, WB08]. non-local [Bör07]. Non-negative [HV08, FM07a, Gro01, GG07, Liv04, Pat04, wW02]. non-negativity [Lad08]. non-null [FKP00]. non-powerful [Ma09]. non-product [Har04]. non-rational [ES02]. non-real [GG07]. Non-regular [DGN02, EK08b, GHI09, Zha05c]. non-semigroup [Eld09]. Non-separating [Lem04]. non-singular [Bri01]. Non-skip-free [BM04d]. non-square [XM02]. non-stationary [ABDD01]. non-symmetric [MN08, Sid01, Ste06a]. non-trivial [Val08]. non-uniform [Kob00]. Non-vanishing [Ash05]. non-zero [Lim08a]. nonabelian [MR09]. nonautonomous [Tre09a]. noncommutative [AM09, Jøn09, KVV09, Kro01, YFK05]. noncommuting [Dru08b]. nonconstant [CS03d]. nondegeneracy [Qi09]. nondegenerate [FKRS09]. nonequispaced [NS03]. nonexistence [Myu09]. nonexpansive [LV03]. nonextensive [Fur06a]. nonhomogeneous [Yan02a, Yan09a]. nonincreasing [HHY04, Lim05b]. Nonlinear [Ano06f, DHB07, HHZ09, Lim09b, Mei06, ZZ08a, BW03, BLM03, BSS09, CGXL08, CY00, CKMS02, CKM05, DLT08, ESP05, HES06, HHT06, JKL09, LV06, Lim09c, NN01, RR02, RR04a, Reu06, RW06, Sol06, WM05, ZF07]. nonmaximal [LHF06, Wal05b]. Nonnegative [Ano03-44, BJ09b, FS06b, FN00, Gro00b, JT04a, KSH05, KS07b, LŠ06a, PPP06, Pit09, SvKHO09, Val00, YTI06, AHH^{+}09, Bau03, BSKS05, BL09b, BNV05, Bon09, BMS04, BMS08a, BD09b, BL02, Bru02b, BAD09, But06, CHNP04, Ceg01, CC01, CL09b, Chu03, CFJ01, Das09, DJKM00, ELN04, ES08, EH07, FN01, Fan02, FM02, FNS07, GN03, Gro00a, Guo02, HNV00, Hua08a, HMT^{+}08, JS07a, JTeG09, LŠ07b, LŠ08a, LM06a, LM02, LTW02, LR09, LM04b, Lur05, MPS07, McD03, MM08, Nab07, NM04, Nou08, NT09c, OSvdD02, PG01, PT03, PT06b, RS03b, RS09, Sav03, Sch02c, SS03b, SdL03, Šmi04, Sot03, SBM05, SR06, Sot06, Sou05, TMARAE^{+}07, VWD08, WD06, ZM03, Zen01, Zha07a, ZyC03, ZH02]. Nonnegative-definite [Gro00b, ZyC03].

Nonnegativity [DHZ03b]. Nonnormal [Böt06b, TE05]. nonnormality [Huh05]. nonograms [BSW06]. nonpositive [ATU06, CRU09, HL01]. nonreflecting [Li06b]. nonsingular [BJ07, Fan04, FM00a, Fie05c, FJ03, GH03, LR05c, Peñ03a, SH00a, ZS05]. Nonsingularity [BB04a, Hon06, Kol03b, DB08]. Nonsingularity/singularity [Kol03b]. Nonsquare [PR02b, LPR04, PR02a]. Nonstationary [WM05, HT09a]. nonstrictly [Kol03b]. Nonsubstitution [HBHS05]. nonsymmetric [AC05, Fre02, Guo02, Guo07a, RT01a, Wu06a, XD03]. nontrivial [RSR00, Roj07a]. nonuniqueness [CD00d]. nonzero [SHS03a, Wil09]. Nörlund [CKW08]. Norm [BDK08, CL04b, HK02b, Kit04, MT05, AR06b, Aud06, Aud08, Auj03, BNT05, BV00, BW08b, Con09, Dra08a, EKS01, GWW06, HK00, HL04d, Joc09a, KN04, Kol09, LV03, LR09, Mah07, Mat05b, MZ08, NT01, Oku03, Oku06, Sed04, Sed07, gS02, WL09b, WG02b]. Normal [Con09, GR00, Hor02, Mue09, Sav06b, Tov00, AM02, Ala05, Cao09, CT03c, CT01, CN06, CHY05, CL04b, CI09b, CI09a, DS04a, – DZ04, Dri08, FI07, FI08a, FMZ06, FJ02, GW04c, GHM03, GH09, JK03, Mat05a, MR01, MK03, MD09, NBdP08b, NS08a, Nie07c, Nie08, PH00, Pug04, QD09, Tre09d, Woe04, Zal03, dB00]. Normalisers [Haz05]. Normalized [Vid07, BJ08, FKM02, Li06a]. normed [Con08, LS08b, LwW07, Reu06]. Norms [Fit01, And04a, And06, Aud09b, Bar08, Bel05, BK00a, Bou06a, CLS05, Fen03, FT07, FIL07, Gol05, Gre09, GZ08, HZ02, HZ04a, Joc09a, Joc09b, Lás07, Lav03b, Lav05, Li08c, Lur06b, Mae08, Mor08, OT01, PR01, Poo08, Pro08, RT01b, SSA05, TH01, Ton00, Wir02, dSS08]. Note [WX06b, WX07a, AM09, Ano04-39, AC04b, Aud09b, BBÖ04, BK03, BR06, BPR03, BC08b, CJ04, CW00a, CW02, CP07a, Cec00, CJVW07, CDS08, CI08, Dah00, DW09b, DRGR04, DA10, EFN09, EFN10, Els05, Fan10, FI08b, Fie03c, Fie05c, FG08, FYK05, GAE02, Guo02, GT04, Guo06b, Hof08, HT05a, Ikr07a, Ips03, Kir03b, KP08, KS04, LT07a, Leš01, Leš04, LL05a, LW07, LY09a, Lim06, LHF06, LM09b, Lur06b, MP06, Mou06, Mül03, NS08a, PG01, PX03, Pet05, RSS00, RBTR09, RVZ00, RP03, Sch06, SM09a, Smi08b, Son02, SC08, Ste06b, Str08, Tan09b, THZ09, Tik06, Tre03c, Tre04b, WHBM09a, Xia06, XL04, XQW08, XLG^{+}13, Zha05d]. Notes [BK00b, BKMR04a, BMS07, WHAB08]. notions [Log05, SdSW07]. November [Ros06]. Novikov [BDV08]. NP [BP02, But08]. NP-complete [But08]. NP-hard [BP02]. nucleus [GH00b]. Null [BP03, AAGK06, AH03, CT04, FKP00, KS02, NS07a]. null-spaces [KS02, NS07a]. nullities [Bap03a]. nullity [FQ09, FGS05, GYY09, HXL08, KR06, Li08b, SdSW07, XL05, vdH03b, vdH08b]. nullspace [Alf06a, Bor08]. Nullstellensatz [RP05]. Number [SHS03a, AA07, ÁFMNW03, AG08, BFH04, BFH05, BS03b, CR08, CS03d, FL03, FYL06, FYZ07, Fer01a, FJ02, Gha01, GYY09, Gut08, HW07a, HKKL05, HL05b, II09, KSH07, KN06b, LLT05, MZ05, MW01b, NX04, Nik07c, NCV05, RS03a, Ros09, SW03, Sil00, SAH08, VLAOVL08, XHSZ09, ZF07, dSZ00a]. numbers [AJF09, BS09b, CL08b, CL09b, CHY05, EL01, Fer02, Gol09, HP03a, Lee00, MS03a, MS00b, Ste07a, gS02, TS03, Yan02a, Yan09a, Zha07a]. numerators [Hag07]. Numerical [ABB00, AVV05, BLdP04, FQ04, FMN04, GW04b, GW04c, GW07, GW08, Guo04, HHW00, Mat01, MBLV01, Psa00, Sty02a, Toy09, Wu02, WGT09, Abd05, And06, BLdPS05, BdPT08, BBMX07, Boc03a, BAGS06, BS04c, BBP03, BLO02, BG06, Cal08, CL01a, CSSZ01, CC01, CL00a, CT01, CN01, CN02a, CN02b, CNP02, CN06, CN07, CKŻ06, CH00b, DS04a, DLS08, DKS04, DHK06, Dra08a, Dru08a, Ed05, FGM03, GGM^{+}09, GW03a, GW04a, GW09a, GS08, GJK04, HHZ09, HLS00, INOY03, JO02, Kar09, Kru04, KS06d, LMMT01, Leš02, LZ00, LTW02, LPS05, LP07, LR09, Lin02a, MS04b, MPT02, MK05b, NP01, NBdP05, NBdP08a, NT09c, RS05a, Ros05, Ros09, Sed01, Yam02]. numerically [PT05].

objects [ÁFMNW03]. oblique [AC09, Čer09, CMS02, LW01, WZ07]. oblivious [BZ06]. observability [BRR01]. observable [CG00a, PPZ02]. observables [Mol02]. observation [Vis00]. observations [BS05b, FI05, FI07, LD02]. Observer [Fuh08, FH01, Sun02]. observer-based [Sun02]. Observers [BO09, HvS07, Tru07]. obtain [VV06]. obtained [KMMZ04, Roj07c, fur08b]. Obtaining [CJPS02]. Obtuse [Ceg01]. occurring [DH01]. octahedron [KO08]. octonionic [KKK07]. octonions [BDE03]. odd [Kar07, Kov02, Mel01, Nik08b, ySZyY06]. off [BC07, BC08b, dSZ00b]. off-diagonal [BC07, BC08b, dSZ00b]. oil [HFR04]. Old [dA07, BMS08b]. Oliveira [Ano03t, Ano05-31]. Olkin [Ano04v, Mat03]. OLPS [Hag07]. OMC [Dru07]. omitting [Hog00]. Omladic [CL07a]. One [CL09a, Ral03, BB04c, BdHZ05, BS00c, BS03b, gCZ03, CL06, CDJ08, DS07, DJ06, EGO05a, EL01, FGWG07, Fer07a, Fur04, Fur07, GR03, GHS09, Hol08, JLL03, KN06a, Kuz02b, LŠ09a, LRŠ02, Lim05a, Lim05b, Lim06, Lim09a, LWZ09, Nom05c, Ose08a, Peñ09, RY08, Sav03, Sav04, Sav05, Sav06b, Sed07, Ste09, Str03b, Ter02, Ton04, YLZ06, xZxGpL09]. one-parameter [BS03b]. one-point [GR03]. One-sided [Ral03, RY08]. ones [BD03d, CL08b, Sbu05]. only [CDJ08, GHS09, LP09, Ose08a]. onto [Ha02, Ple07]. Open [HS02, AMZ04, CW05, CK08, HL07a, Peñ08]. open-loop [AMZ04]. opérateurs [Fao01b]. operations [FMZ06, Huh06, WSL09, dMFZ09]. Operator [CZ07, FNS01, Fur03, Fur07, Gus02, Gus09, LT05, Mos09, Abd05, AKGC09, AD06, AD03b, Auj00, BKR06, BDK08, Bha06, Bra07, CGM05, Cao07a, CZZ09, CN06, CL01b, CHL01, CH08, Di 03, DM02, Dra06, DK04, Dug01, Dug04a, Dug04b, Dug05a, Dug08, Dup03, EG03, EKS01, ES01, FYL07, FYY09, Far04, Fei04, Fer01b, FP02b, FST04, FYK05, Fur02, Fur05c, Fur06b, Fur08c, Fur09, GHS06, GHI09, Gus00, Han09, HZ02, Ho03, HC03a, HZ05a, Ili08, IK09, JL09, Kri05, Kru04, LMMT01, LSD06, Lin07, LwW07, Lu02a, Lu02b, MPT04, McC01, MŠ04e, MZ08, NN01, Oku06, Ose08a, RST01, RDxC03, Sch03, Sed04, Smi01c, SZ04, Uch02, Wu02, Xu08, XSG08, YKF05, Zar06, ZZJ08, ZX05b, Zhu07, ZX07, Ziz04, fur08b]. operator-theoretic [Zar06]. operator-valued [EKS01, ES01, McC01, Smi01c]. Operators [Böt06b, ABP09, AGP08, AHZ05, AH09b, ACS06, BL00, BS02c, BLS02, BG09a, BLdP04, BS09a, BD02b, Bon03b, BGS09, Bör07, Bou05, BK09b, Bur06, CD00a, Cal07, CMV05, Cao05, CPR03, CŠ08b, CDL05, CN07, CCH02b, Con09, DK08a, Dam03, DH03, Del01, Din03, DKL^{+}07, DH09, Dra06, Dra08a, Dra08b, DKP08, Dru08a, DD05, Dug05b, Dug06, DJ07, FJP07, Fao01a, FQ04, Fur04, GW08, GMP08, GV06b, HHY04, HN04, HK02b, Hir05, HK07, Hir09, HO01, HHZ09, HM04, HHT06, JK06a, Jia09, JLL03, Joc09a, Joc09b, JLY09, KP06, Kal03b, KL07a, Kit04, Kit09, KMP00, KS07b, LP01a, LŠS07, LL08b, LS08b, LP09, Lin07, LRR04, MB07, MA06, MS04c, Mat01, MPP06, MŠ04e, MPS03, MT06, Mol09, MZ08, NBdP05, NZ06, Nie09, OT01, PR01, PM05, PWZ05]. operators [Pro08, RY08, Sch08, Sed01, Sed07, Ser08, SC07, She03, SW09, Str06, TY08, TU02, TUC04, Tan03b, Tic05, TE05, Tup08b, Tup09, Tur00, Tur05, WZ07, WFJ08, Wie01, Wir09, XH08, XS08, Xu09, Yah08, Yam06, YLZ06, Zha03a]. optics [CPPS00, PEPS00]. Optimal [BS00a, HH05, HP04a, JSWT05, Mae08, Mas09, MQ06, Ose08b, RW01, SAAF08, TVB08, Tic05, Wu06b, Ye08, ZW03, ÁFMNW03, ÁFMNW06, BM01b, BF07b, CJVW07, DZ01, FN07, HT09b, HBHS05, HH09, Li01, LY05, MD02, MW08a, MW08b, hPyHZ07, PWZ05, Udd09, ZHZ08]. optimality [Bag06, Dru04a]. Optimally [Kol02]. Optimization [AHH^{+}09, HLW06, BC04a, CLPT06, CCW^{+}08, Dah01, DLZ03, GB06, KFGK04, Pue07, Sch02a, SK08a, ZM08]. Optimizing [CLOvdD01, BLO02, FNV08]. orbit [BB03b, CFP04, Fao01a, NBdP08b, PPT04]. orbital [dST05, dST06]. Orbits [Sem06a, Sem07, CM08a, Gao03, GX09a, Gew03, GN09, LT03a, LP09]. Order [Ano02-29, BFSV06, Foš05, SL06, TH00, BF01, BB05a, BB04d, Bar08, BGL09, BJ09a, BJ09b, BD07, CDMGP06, CLVV06, CNF08, CL09c, CK04, DJMS09, DV04, DDM08, Dun04, EGO05a, EGO05b, Fre08, FKN02, Fur02, FG03, Fur08c, Fur09, GPMSZ06, GN00, Gro06, HL07a, HZM02, HL04e, IK06, JY07a, JY07b, JL03, Kes07, KL07a, KS06a, KKK07, LCF06, LV06, Li06b, LRK06, Lin07, LT03b, LS09c, MM05, MPS06, Mei05, MPS03, Nik09, PM05, Pet08, RW06, Roj07b, Sav03, Sav04, Sav05, Sav06b, Sen07, ySZyY06, SC01, SS06e, Wan02, WWJ09, WG02b, WSW06, Yan01a, YSL08, Zho03, tBSR04]. order-index [HL07a]. order-one-quasiseparable [EGO05a]. Ordered [Hag07, BM05, CSWM01, DH03, HL00, Her03, Leg04, Mou06, SS06d]. Ordering [CH03, LG06, Wan09e, WK09, YLT05b, Zha07c, ZLC09, BZ06, Fer03, Gro01, Gro06, Ili09b, KP09, KS07a, Leg04, SGS08, Zha00]. orderings [BBL03a, GML09, Gut01, HMS01, Nie00]. orders [BBL03b, BH04, BHLL04, HP02, Woj03]. ordinary [CLL07]. orientation [GHO02]. oriented [GTT01, MQ06]. originating [CKZ05]. ORQI [Dax03b]. Ortega [Ben08]. orthant [BNV05, Lav03b]. orthant-monotonicity [Lav03b]. orthocentricity [Fie09]. Orthogonal [BN08, DD02, Hog08, Ver08, ASK^{+}08, BBS02b, BBK07, BT09a, BT09c, BBD^{+}08, BPR03, CMV03, CMV05, CS00a, CHR^{+}03, CDJ08, DP04, – DZ04, FMZ06, FK04, GHO02, HPP05, Ish04, JR09, Jia00b, KRS03, LV06, LSS00, MHZU05, MHZ05, OPS07, Ohn08, PB06, Ple07, Pou01, SSW05, Tre03c, TZ03, Uhl01, VV06, YMP01, dST05, Dax03b]. Orthogonality [BFS07, LS02b, BLR00, Dug00, Dug01, Dug04b, Rie00, Tur00, Tur05, dST06]. orthogonalization [Dax00, Eld06a, Pop01]. Orthogonally [vGKN09]. orthonormal [DK06, LZ00]. orthostochastic [Co08, NBdP05]. Orthosymmetric [SS08b]. oscillating [DK08b, LRR04]. oscillatory [FFM03, Fal04, FL07a, FM04a, GT08]. Ostrowski [Kol03a, Yua00]. Ostrowski-Brauer [Kol03a]. Ostrowski-Reich [Yua00]. other [BBS06, GN03, GS04b, Stu00, Ter01b]. Out-of-band [EG08]. Outer [Bap03a, Rob02, CC00, DvdV00, GZZ00]. outlinks [dKNvD08]. Output [RRH02, EG02, HZM02, LV06, RZ01, RZ04, RZ05, Ste01a, Sun02, Zen01]. over-relaxation [XZ06]. Overcomplete [AFCC04, AEB06]. Overlapping [BPS04]. Oxford [Böt06b].

P [GSR03]. P. [Ant07]. PA [Par09b].

packings [dW03]. Padé [BF01, CL01a, DP00, Fre08, Gu01]. Padé-type [Fre08]. Padé-via-Lanczos [BF01]. PageRank [KHG04, dKNvD08]. pages [Böt06b, Rod09, Tům06, Bar05]. pair [CSZ01, CL03c, DJ01, HTV02, MS06, Nom05a, NT06a, NT06b, NT06c, NT07b, NT07c, NT08c, NT08b, NT09a, PP02, PPZ05]. paires [Fao01b]. pairings [GHM03]. Pairs [BKR03, BGS09, LRWW07, Rod08, AC04a, AC08, AI00, BS02b, BR00b, BLS02, BZ05, BBL^{+}05, CNSM07, wCL04, DKL^{+}07, FS04a, FS06a, FN09, GS04a, God09, GS00, Har05, Hog03, IT07a, IT07b, KS06a, LR05b, LT07b, LT09, LgS01, Lon04, Nak03, NV03, Nom05b, Nom05c, NT07a, NT08a, NT08d, NT09b, Obl08, SS04a, SS03b, SvdE03, Ter04, Vid08, Vid07, Wu06b, Wu08b, Zha03a]. pairwise [EvdD04, FGR04]. Paley [Ren07]. paper [smbH02]. Papers [Ano02q, Ano03v, Ano03t, Ano03u, Ano04u, Ano04v, Ano04t, Ano05p, Ano05q, Ano05r, Ano06f, Ano06g, Bar05, Ano02r, Ano05-33, BHN^{+}07]. Pappus [Don01]. parabolic [Dvo03, RT01a, WY06a]. paracontractivity [VEB00]. Parafac [BSD09, SS07f, Ste09]. Parallel [BRM01, ACS06, BW03, BDKW04, DG07, LS01, Mol09]. parallelism [BM01b]. Parameter [BFPW02, FLPW03, BS03b, Dam02, GA06a, JLMP02, JH09, OP05, PEPS00, TG02]. parameter-dependent [OP05]. Parameterization [LT03a, KC03]. parameterized [BW08a, GRB04, ZBY09]. parameters [Bai08, MPS06]. Parametric [Fur04, CC06, DTGV06]. parametrization [FPP01, FH01, PPZ05]. Parametrizations [Jia00b]. parametrized [Neu01]. parity [Sbu08, Wu06c]. Parrott [BBM02]. part [Dra08a, GV09b, GT06a, GT06b, vGKN09, KS01, Lan05]. Parter [BGM08, KS08]. Parter-vertices [KS08]. Partial [ACG08, Bha03b, FH00, ATU06, BF01, BBL03a, BBL03b, BHLL04, BJ09b, BW01, Cap03, Fan07b, Gro01, Gro06, Gut01, HMS01, HMH01, JTU03, KC03, KC06, Lop07, Phi04, Sha09, Urb04, Woe02, Woj03, Zha06, vDS05]. partially [ASX08, AEI00, CDS01, HDA09, Kro01, Leg04, SS06d, Sil00, TP09a]. Particular [BB07a, XH08, BBdH09]. partition [BD08, Fon00, HBHS05, XZ06]. Partitioned [Dek09, Aud06, Aud08, BS02a, BBS04, BB07a, GP00b, Gro00a, MB07, TT09, XH08, Xu09]. Partitioning [dW03, ACHR01, ACDR04, ST07]. partitions [BD06a, BFP07]. partly [ZS05]. parts [LŠ06a, Ste07a]. Pascal [AG04b, AG05a, BT00, CK01, SZ08, VS04a, YM07, ZW00]. pass [HOP06]. passage [BDKW04, CM00c, Hun05, Hun08]. Passivity [FKLN07]. Passivity-preserving [FKLN07]. Path [DWD07, BFH04, BFH05, BS04a, BKK07, DH02, JS07b, JS07c, Omi08b, RM09a, YC00b]. path-recursive [YC00b]. path-simplex [BKK07]. path-subsimplices [BKK07]. pathlengths [Hig06]. pathogen [PSvdW05]. pathway [Mat05a]. Pattern [Ruk01, AT09, AHK^{+}05, COvdD09, CL08a, CJK05, EOvdD03, GL01, HLW01, HLR04, HL01, LHS02, LHS05, LY06, LSH08, Ned01a, SV07, SS05a, SLR06, SBS02, Stu03, Wil09, YSS07]. pattern-shaped [Ned01a]. Patterns [BAGS06, EJLS02, JS00b, AHL^{+}09, BOvdD07, CV05, CVV09, CHMM06, Dah08, DH00, DHH^{+}06, DHH^{+}07, – DJ07, DJOvdD00, GLS07a, Hog00, Hua08a, Hua09, KOvdD07, MTvdD05, MS08, ML02, SG03, SSG04]. Paul [Ano07o]. Pauli [HNPT06]. Pauline [Ano04u]. PBW [CD05, GML09]. PCG [VWJ08]. PDE [LV06]. PDEs [JL03, ZWPR06]. Peano [BZ06]. Pecarić [FG03, MPS00, MPS03]. peculiar [Val02a]. penalty [HDA09, IL06]. pencil [JH09, KKP07, KG09, XM02, vdW02]. pencils [BB03b, CN02b, Cof03, DDM08, Dod08a, Dod08b, GPM01, GPS01, Gil04, Gil08, KKP08, LK09, Meh00, MMX04, MNV08, Pap00, Psa00, Rod07, SS04a, rWQ02, XD03, Xu06, dSZ00a]. pendant [Guo05d, GW09c, LZS09, WXH05]. pendent [GL08, HW07a, Li08b, YL06]. Penrose [BBT03, BB07a, Bap00, BOvdD04, Fie03b, Gro00a, Gus09, KDC07, KS07b, Li08d, MD09, Pat01, Pat03, PP04a, XH08, Xu09]. Penrose-invertible [MD09]. pentagon [LM03]. per-antisymmetric [Yin05]. Pereyra [BEG^{+}07]. Pereyra-type [BEG^{+}07]. perfect [BSD09, BH05, CT03b, GLL09, GT02b, GT04, Han07, JNL05, MN03, WK09]. performance [ACHR01, RO02]. period [Gav00, Mol05b, YC00b]. Periodic [CJK05, Sat05, dlP03, AKGC09, BRVS08, CL05a, CN07, wCL04, CP05, Kre06, LgS01, Liv04, Ots01, Ser09]. periodically [AFL00]. periodicity [Lad08]. periods [AKGC09, HL04a]. peripheral [McD03, MM08]. permanent [AFM03, Bri05, LB04, Sch09c, Wan03b]. Permanental [Sou05, Pat08a]. Permanents [ABH03, CHSS03, HL04e, BN04, CW05, CR02b, GN03, Han00, RS03a, RS06a, Sbu05, Sbu08, SHRC03, WlW07]. permutability [AG04a, TY06]. permutable [AGK06]. Permutation [CB06, Cig07, Dah09, Tad06, ALSS08, CW04, HP04b, LTT02, Nah06]. Permutation-like [Cig07]. Permutationally [PR04c]. Permuted [But08]. Permutohedra [OV06b]. Perron [Bap03b, BN04, Bom08, CLL05, DW05b, EvdD08, FN01, GN05, Kir03a, Kol08, Lin02b, Lu02c, LN04, MPT02, MNV08, MV04, Neu00, Nou06, Nou08, PT03, PT04, Rum03b, Sav03, Sav04, Tam04, TRH01]. Persistently [HTK07]. Personal [BKMR04a, Ros06, Sch02b]. perspective [TG02]. Perturbation [AH03, CX05, wCL04, CMP02, DT03, Fre04, GKW00, GJW02, HJM09, HM04, Iva05, Jia05, KAP^{+}03, KL07b, LZ09b, LgS01, Ste06c, gSX03, gS03b, Ves00, WZ07, XWD06, Xu01, AB03a, AB03b, BBdH09, BS00a, BJ09a, BM04e, CMP03, CM01, DDM08, DD09, Din03, DS07, DMM00b, Eis06, Gon05, HIU04, HI06, HS09, Ips00, Ips03, LS00a, Li00b, LW01, Li08c, Lip05, gLgWmW08, Nak03, NS08a, RA08a, RR02, Rod06, RS09, SS00, SS03c, ST00, ST03b, Ste04, Sty02a, WLBZ06, Xue01, YLF09]. Perturbations [YMP01, AGP08, BdHZ05, BdHZ08b, BdHZ08a, BOP08, BM06b, EOvdD03, Fer02, HP07, Peñ03b, Rum06, Smo08, YWS00]. Perturbed [KFGK04, AH04, EN09, FKS02, FB05, HTK07, Hun05, Hun06, KH09, LL05a, LS09c, RSW05, Zem07, vdMSR03]. Perturbing [GG07]. Peter [Ano02z, Ano04-44]. pfaffian [Lop07, Mat05c]. Pfaffianization [GZHT05]. Pfaffians [KRS09]. pg [Brü03c]. pg-critical [Brü03c]. phase [KB02a, LG04]. phases [Had08]. Philadelphia [Par09b]. Photo [Ano04-42, Ano04-43, Ano04-44, Ano07n, Ano07o, Ano07p, Ano08a, Ano08r, Ano09n, Ano09o, Ano07a]. physical [Bai08]. physics [BLdP04, DLMO07]. PI [KYAA08]. Picard [HH00]. Pick [Las06, Bol02, BKR03, BKR04, CH00a, CH01, GM07a, HdABC07, Wu05c]. PID [HBLZ06]. piecewise [Bon09, ÇS02a, RW06]. piecewise-linear [RW06]. Pierce [HPP05]. piling [OT09]. pivot [ND05, Tsa00]. Pivoting [BKO01, Bol01, Dax00, Hig00, Yeu04]. placement [CR03, EG02]. places [PFL05]. plan [SK08a]. planar [Gri04, ST07]. plane [DMD03, FF04a, FF04c, MS05]. planes [Kes00, Kes07, LL06b, Tag04, Wat01a]. planning [CBIXG08, CCW^{+}08, SK08b, TVB08]. plant [PSvdW05]. Plemmons [BN00, SMK09]. PLS [PRP02]. plus [CG00b, De 00, GK07, Gav00, Hei01a, Hei02a, HR03a, LT04a, Mol05b, NS07b, PVV08, Ser07, SH05]. Poincaré [Dom00]. Point [CNP02, WA09, AFR06, BW08a, Cao07a, FG08, GVV03, GRB04, GR03, HMR09, Mat05c, NY00, OS05, Ple07, Rih01, TU02, XWD06, ZBY09, ZX05b]. points [AA09, AKGC09, Alf05, Ano03-31, BM04b, CGG02, De 07b, Guo03a, Guo08, HL03, HH09, Hol02, HQ08, Kir03b, MW03a, Neu01, Pug04, TUC04, Wie01, Wu09, Zhu07, ZXZ08, ZXZ09]. pointwise [dBS08]. Poisson [AL09a, Wer09]. polar [DoMP09, De 07a, De 08, De 09b, FI08b, GK09, HMP09, gLxW02, Li08c, LY09a, MPP10, Smi08b]. polarization [DV09]. Polaroid [Dug06, AGP08]. Pole [CR03, CHASG02, DN05a, EG02, GW02, Zho03]. Pole-shifting [CHASG02]. Poles [Wym02, SS09c]. polydisks [BR01]. polygon [Mir03]. Polygonal [RT03]. polygons [DHZ03a, GW04b]. polyhedra [Kar07]. Polyhedral [SV06b]. Polynomial [BB03c, BM06b, DLS08, JH09, Sci02, TTM02, Yan01b, Alf06a, AY08, AMZ07, AVV05, BD02a, BD03a, BJT04, Bas02, BM04a, BLM02, BMP00, BD01b, Bor06, BD06b, BL02, BP09, BG06, Cal07, CFL05a, CNP02, DS04a, DT03, DK06, DT09, DS08b, EF07, FGM03, FLM02, Fei04, FVP06, GGM^{+}09, Gem03, GŞV02, Gre02, HT08b, HT02, KV09, KT08, Kru04, OdAJ02, PP08, Per07, PS07, Pet08, Pot03, Pug04, Pul07, Rad03a, RZ05, SS08a, Ste06a, TR02, Tis00, TMARAE^{+}07, Wan07b, YH04b]. Polynomials [BHJ08, Wat04a, dSR05, Aga02b, AA09, AvdH04, AADL03, ADL04, Arp03, Arp05, Bat06, BGLS05, BG09b, Big02, BGM01, BC07, BC08b, BD09a, BFPW02, BMX08, CD00a, CMV03, CMV05, Chi04, CLY07, CK04, CDS01, CS03d, DD09, DK06, DW05a, DSW07b, DG03, DV09, DD02, DV07b, FN03, Gao03, GI02, GZ07, Gra02, HNX03, HL00, HT03, HLY06, Hol08, JR09, KPK04, KM04, KKL08, KK01, LK00, Lee09b, LdlP09, Lie09, Lin02a, LM09c, MP00, MS03c, McC01, Mon08, NP01, Nör03, PP08, PT04, Sak04, SZ08, Ter04, Tre03c, TZ03, VPV01, WW07, Wan09a, WO05, Win00, Win03, WD06, Wu08d, YMP01, Yan03, ZJL02, Zyg02]. polytope [BFP07, CdFM08b, CdFM09, CW04, DB08, GZ08, HP04b, MS02b]. polytopes [AMM06, BF05, DS04b, Hil03, LT03a, OP05, BM04b]. Poncelet [MBLV01, Mir03, MS05]. Pontrjagin [FYL07]. Popov [MSS07]. Population [Ost06, Daw02, Har06, Kir05, LS06c]. populations [DDvG05]. portion [RS05a]. portions [BdPT08, BS04c, CN09]. Portugal [dSQS00]. posed [CRS03]. posedness [CLPT06]. poset [Foš05, Wu06c]. posets [BFL05, Zav03]. position [Alf05, JDS^{+}03b]. positions [Hog00]. Positive [And07b, BSL04, BS04b, CHP06, CM08b, Dam04, FGR07, Ha02, Has05, HDA09, JS07b, JS07c, Peñ09, Qi09, Var03, XS08, Amb07, AC04b, Aud06, Aud09b, BGN08, Bal07, Bar01, BB03a, BX04, BX05, BF08, Bou05, BRS00, BCS02, BRVS08, Brü03c, Buc06, BAD09, CMP03, CC09, CLL05, CC03, CR08, CX05, CLL07, CL09b, CL07b, CS01, DH03, Dat08, DA10, DJKM00, DL08, FJ07, FJ06, FM00b, Fie06, FS03, FVP06, FF02, Gar03, GJ04a, Gen00, Gil01, Gla04, Gol03, Gro00b, GV06b, Ha03, HvS07, HES06, HTK07, HP09, HZY05, HHT06, JS00a, JO04, JO05, JT04b, KL07a, KNR05, Kit04, KLM01, KC03, KZ03, Kye03, LL08a, LM06b, Liu02b, LG03, LHF06, LT03b, MS06, MVV08, MPS03, MND01a, Mol09, NS08a, O’L03, OB01, Ose08a, PM00, PM05, PVV08, RSW02]. positive [SV07, She03, Smi08a, SDL06, Ter01a, Vir08, Wal05b, WB01, Xu04, Xu08, YH00b, Zha07a, Zha03b, ZyC03, ZF07, vdH03b]. positive-definite [Gro00b, ZyC03]. Positively [Ned01b]. Positivity [Ano03v, HK03, BR01, DJL01, Gla02, GG04, Hol03, Hor04, KV06, KH09, Peñ01, Peñ04, SS06a, FMPT04]. positron [AG05b]. possessing [BES02]. Possible [JLY01, JD02, MV04]. possibly [BS00b, FG02, Sen04]. Potapov [FKRS09, FKR09, Las06]. potence [Hua09, SC08, YW07]. potences [WY08]. potent [BT05a, Hua09, SBS02, Stu03, ZgC06]. potential [AH09b, GL01, Mas09]. Potentially [BOvdD07]. Potter [LM09b]. pour [Del01]. power [AG00, BM04d, CH00a, CH01, CHP03, CHJP05, De 09c, EvdD01, Gin09, HK02b, Hon08, LDS07, RM09b, Sak04, SDL06]. power-associative [CHP03, CHJP05]. power/ [ZgC06]. powerful [CJK05, LHS02, LHS05, Ma09]. powers [AG05a, BHLL04, BCG07, Che04b, Che07c, CLL07, De 00, FJ07, Fer04, FS03, HP01a, Kir00, PG01, Ser09, SS04b, TR03, Tan01, THZ09, Wer08, Zho08]. pp [Kol05, Par09b]. practical [AEB06]. pre [AR06b, Bén06, Min01, PR07]. pre-distance [Bén06]. pre-Hilbert [AR06b, PR07]. pre-semi-rings [Min01]. Preconditioned [Sol06, BGGC06, CJVW07, FS08, KN03b, Ney01a, Ney01b, Ney06, Ney09, Ovt06c]. preconditioner [CJ04, CJW05, CJ05, DV00, FLM02, JW08, NB03]. preconditioners [Cap01b, CEK^{+}01, Est05, JL03, KY00, NCCY00, OS05, OT06, Wd09]. Preconditioning [BL09a, LCN04, Tům06, CR03, CP03, Che05, DZ01, KH01a, PY09, vdMRS06]. predictable [KPP07]. Preface [AHST09, Ano02x, Ano03-42, Ano03-43, Ano03-46, Ano03-44, Ano03-45, Ano07s, Ano07q, Ano07r, Ano09p, Ano09q, AHR^{+}07, BKvdV06, BKM^{+}04, BPV06, BHKL05, BKMR04b, BBVR08, BLL06, BGKW07, BFW09, BKMR08, CNTV08, CHR07, CFKS08, DKNS05, EMS06, GLQS09, HHLP01, IMS^{+}09, KMOT05, LS04a, LWZ05, PSW00, PSW02, PSW04, PSW05, RRY01, VHM04, dOQSZ06]. prehomogeneous [AFK02]. preordering [Gro01]. prescribed [ASX08, AMZ06, AM05, Bor06, BC07, BC08b, CSZ01, CDS04, CS00b, CDS01, CS03c, CS03d, Dod05, Fon00, FIS02, Fur08a, Gra05, GV09b, Kro01, LM06b, O’L03, PP08, PS08, Sil00, Sot03, Yin04, Yin05]. prescription [CCW^{+}08]. presence [BP02, Ste06c]. presentations [BL02]. presented [Ano05-33]. preserve [AT07, BS02c, BLS02, DKL^{+}07, Fer05, GR07b, JLL03, Zha03a]. preserver [CL03b, GLP00, TW03]. Preservers [LR09, LT07b, LT09, Mol03, AG04a, BLL03, BSKS05, BGLS05, BGJS06, Bru03b, CZ05, CC01, CFL02, CDL05, CL05b, CL06, CHL01, CGH04, Cui07, DDHL08, Gut01, HR07, Kuz02a, Leš02, Lim08a, Lim09a, LT04b, Nie07c, RP00, SKBS08, Tan05b, TY06, YT04, Zha04c, Zha05a, Zha05b, ZgC06]. Preserving [BBG09, AH02, AHZ05, BL00, BGLS04, BG09a, BS00c, BOP08, BBS08, BS09a, Bu05, CCH02a, gCZ03, CPR03, CFL05a, CLP03, CFL05b, CD06, CL09d, DP02a, Dol03, DG07, DH04, DHB07, FJ06, FJP07, FKLN07, FL04, Foš05, FG03, Fur08c, Fur09, GR07a, GM07b, HC03a, HC03b, HZ04b, HZ05a, HJ08, HHZ09, HBLZ06, lHH08, Ji03a, Ji03b, JW06, LP01a, LTT02, LLM03, LŠS07, LP07, Lim05a, Lin03, LL08c, Mol02, Mol09, Nie01b, Pet02, PS03, Pet05, Poo08, RST01, Rad03b, RSW02, Šem08, SZZ07, SHS03b, SC08, Tam04, Tan03b, Tur05, Wan03a, WFJ08, WY08, Wan09a, XCX01, Yan01a, YZ09, YLZ06, YW07, Zha04b, ZYP06, ZZ08a, ZS05]. Press [Böt06b, Moh05, Rod09, Tům06, vdD09]. pretest [HDA09]. Primary [RGG01, HP03b]. prime [Kov02, LZ08b, LL09b, Lin03]. primeness [Str03a]. primer [Kri05, PT04]. Primitive [AK09b, CHY05, KSH07, MNOvdD08, AFS09, BK03, CW03a, CL08a, DOvdD05, GS09a, KSH05, LY06, Ma09, MZ00, Sch02c, SW03]. primitivity [BG09a, GS09b]. Princeton [Böt06b]. Principal [Dah01, GT06a, GT06b, JLMC05, ND05, Sav04, Tsa00, AEK06, ALC06, Bal07, BL05, Chu03, Das09, HAT03, HK03, JS01a, JKO04, LŠ08a, PR00, Sav03, Sav05, Sav06b, Vis00, ZH02, ZHZ08]. principle [BKK08, GAM06, GII08, HT08a, Lim09c]. principles [Kal03b]. priori [RS02]. prioritized [CCW^{+}08]. Probabilistic [Yeu04, McG06, Str02]. probabilities [Gra04b]. probability [BI00, HO02, HLR08, MM03]. probe [Kap01]. problem [AA07, AFL00, AGI00, AP04, ATU03, ATU04, ATU05a, ATU05b, AH03, BB02b, BS00a, BHL09, BGL09, BW00, BSW06, BM04b, BV07, BNV05, Bol02, BMS04, BMS08a, BD09b, BKK08, But08, CGHG06, CKLX09, CW03b, Cen06, Che08, CH01, CZZ09, CY00, CKMS02, CPPS00, CH00b, CI09b, CI09a, CF02, CFP09, DLMO07, DW08, DW00, DN00, DS09, ELN04, EG04, EF07, Fan10, FKL03, GJ04a, GM07a, GT06b, Guo04, HL07a, HD03, HC08, HTV02, HP03a, Hoc09, HJR05, Hog07, Hol05, HC04, HvP05, HP04c, IV04, IO01, JLMP02, Joc09a, JS00a, JT04b, JTeG09, JKLP03, KR07, Knu00, KR08, KMS05b, KMS05c, KC06, pLLfY06, Lim05c, LJ01, LDS07, LM04b, MS01a, MPS07, MS00b, MA07, MZ04, MHZU05, MND01a, Okt04, OB01, pPyHZ05, PES07, PK08, PR02c, Pug04, QX08]. problem [RZ05, RS05c, SS09c, SS02, ST00, ST03b, Šmi04, Smi08a, SBM05, SR06, Sot06, Ton00, TMARAE^{+}07, TL02, WF04, WX07b, Woe02, Woe04, XQW08, XLG^{+}13, YLX09, Zav08, ZZY06, dSQS00, HT09a, Zav07]. Problems [BLS05, AAA^{+}02, AHW05, BX08, BW08a, BL09a, Bal07, BPV00, BN05, BS03a, BBL^{+}05, BLM02, BS04b, BSS09, BEH^{+}06, BGGC06, CRS03, CWW09, Cap01c, COvdD09, CH00a, CSY01, CW05, CK08, CLOvdD01, CKŻ06, CL03b, CP09, ĆDR01, DT03, DD04, DS03b, Dod08a, Dod08b, DJKM00, FN07, Fie01b, FKM02, FKS09, GGL08, GW02, GZZ00, GL06, GSR03, GJW02, GLP00, HV03, HFR04, HTV02, HT02, HS07a, Hog01, Hog02, Hog03, HS02, Ish06, JH09, JY07a, JB00, JS01b, JDS03a, JTU01, Kak08, Kin05, KN03b, Kon08, Koš03, Kos00, KFGK04, Kro01, Lan08, LM04a, LCN04, Log05, LS09c, MM05, MV07, NN01, OF02, PHZ06, Phi04, PA04, Pue07, Rih01, RT01a, RT01b, RM09b, RS01, Sav06a, Ser00, SLR06, SC01, Sol06, Ste07b, TG02, TW03, Tis00, Tre04c]. problems [Tre04d, Tre04e, VHK03, Vol04, Woo02, Wu05c, Zav07, ZBY09, ZHZ03, xZxGpL09, Cra09]. procedure [JK06b]. Proceedings [Ano07s, Ano09p, BBVR08, BGKW07, IMS^{+}09, dOQSZ06, AHST09, Ano05-34, Ano09q, BLL06, EMS06]. process [GZZ00, IPP07, KP09, KS01, KMS05a, KS05, KS06b, KS06c, Mat05c, Pic07, Udd09, vDP08]. processes [AFL00, BLM03, GAM06, Gra04b, Har04, IV04, KS01, LG04, Ney09, PGRO06, Per03, Wat04b]. Processing [Ano02-28, BBP03, DV04, GGCB07, HLW06, LXW04]. Procrustes [Kin05, Tre04c, TL02]. Product [NBdP08b, ABB00, ARR06, And04a, AHH^{+}09, BT09c, BLdPS05, BGO09, BKN01, Che03, CL04a, Che04a, CD00d, CL09d, DLMO07, DKL^{+}07, Dra05, Eld02a, Fan07a, FGR07, Gin09, GT06c, GL09, GS09c, Har04, HZ05a, Hua08b, Huh06, JS07b, JS07c, JM09, KSH08, Kin05, LHSL07, LP07, LCW09, Lim05b, LR00, LC09, MMT04, Mat05b, Pit03, RO02, SdS09, Seo00, Ser08, Son00, Tyr04, VC04, Vis00, WZZ00, WZCL08, Xia03, YLT09]. Products [EL01, FIS02, Fur08a, NDS02, Oik08, Rao09, ZY02, Ala00, AG05b, Bon03a, Bot06a, Bou05, CC04, CLL07, CH08, DL92, DL01, DV07a, FJ06, FJP07, FI03, eFLW09, Fie08, Fri06, FK07, FL07b, FMS05, FKY09, GW03b, Gol09, Gu01, GZ03, Hir05, Hor04, Hou03, HZ04b, HJ08, HHZ09, IS09, Jia09, JPB08, LP01b, LŠS07, Li08d, Liu02b, MZ05, MS03b, MS03c, MR01, MRR07, OS00, Pat08b, PW02, QZ07, Rod06, RT01b, She00, Tad06, Tan09b, Uhl01, Wan02, WFJ08, Wat01b, WG02b, ZYP06, ZZ08a]. Professor [Ano04v]. profile [Had04, HR04]. programming [AHW05, BW00, Eld06a, JK06b, Kal03a, LV06, Luz07, OS05, Qi09, TW05]. programs [DS08b, Val09]. progression [HL05b, YH04a]. Projected [GG09, Baz03]. Projection [JR06, BS06a, Ben04, CS08a, Cen06, Elf04, Fin00, LW01, Ple07, QX08, Sun05, VHK01, WZ07, Zar06]. Projections [And09, ACRS05, AC09, BV00, BD09b, BDD^{+}01, Cof03, CMS02, DL05, Dug05b, FIL07, Ha02, Ili08, Jam07, KRS03, LT07a, Li08d, Oik08, Ost03, Sch09b, SZ04, SvdS01]. projective [Ber04, De 09a, Don01, Fri06, HHT06, Lim07a, Liu03b, LL06b, Wat01a]. projector [Čer09, MHZU05]. projectors [BB02a, BBS02b, BL04, BBL04, BB04b, BK06a, BBG06, BBK07, BBLT08, BT09a, BT09c, BT05b, DWD07, KRS04, Ste06b, dBS08]. Proof [BH08b, Che07c, Lás07, Yon00, ABP09, Alt09, Bén02, Ber09, CL09a, Che07a, Cio06, De 09a, DJ06, Fao01a, Ges02, HZ05b, Kra03b, Li00a, Li06a, LR04, PLG02, Sch02c, SMK09]. proofs [CK08, Rao00]. proper [Kha09, LV06]. Properties [BBS04, HWKS03, Sv05, Tre09b, Tre09c, Bae07, BBL03a, BBL04, BLS02, BEG^{+}07, BG09b, BCS02, COvdD09, CR08, CL05a, CJ05, CGPM01, DL04, DW00, FT00, FG02, eFLW09, FVP06, GQ09a, GSR03, GST04, Gri00, GT06c, GZ01, HMS01, Hon08, JLN08, KNR05, KMP00, Lee09b, LT05, LH04, gLgWmW08, LLHK08, MM05, MLR00, OOT06, Ruk01, SV07, Tre04a, Tre04b, Tre06, Tre08, Tsa00, Wan09a, Wu08a, ZM03, Zha03a, ZW00]. Property [AGP08, BBS02b, BB05a, BBK07, Bal07, BT00, Bén02, Bha03a, Bom08, BPZ08, Che04b, Chu08, GL05b, Han02a, JB08, KMS05a, KPP07, LS08b, LM01, Nie00, NT01, Nou06, SSSG06, SK00, SG06, vdVN00]. proportional [DK04]. proposed [Fur05b]. prostate [CCW^{+}08]. protocol [ZTK09]. Providência [FNT07, Fur07, MF09]. Proximity [CNX05b, KA06]. Pruess [Kos00]. PSD [LMT09]. Pseudo [Han07, O’u02, Din03, Fio01, GX09a, GWL09b, PT06a, Rie00]. Pseudo-centrosymmetric [Han07]. pseudo-cosine [PT06a]. pseudo-distance-regularity [Fio01]. pseudo-inverses [Din03]. Pseudo-multiplicative [O’u02]. pseudo-orthogonality [Rie00]. pseudo-symplectic [GX09a, GWL09b]. pseudoinverse [DP01]. pseudomatroid [SK00]. Pseudoprime [WW08a]. Pseudospectra [AA09, Böt06b, MGWN06, CCTT00, HT02, HP03a, TE05]. pseudospectrum [MS04a, Rum06]. psychometrics [Tak04]. Pták [Sch00b, dB00]. Publications [Ano07t, Ano05-31, BS05a]. pull [And07a]. pull-back [And07a]. Pupkov [Hof00]. Pure [Far04, AYC06, Tre09d]. purely [Ahm05]. Pushing [BGM08]. Pyo [Nie07a].

QBD [Gra04b]. QD [Van04]. Qiao [Moh05]. QR} [Wan01, EGO05b, GVD08, Hig00, HT05b, HT07b, Kar05a, Kre06, SS00, SS03c, ZZY06]. QR-like [ZZY06]. QSVD [CD00c]. Quadratic [Che02a, Kur04, LS02a, LDS08, MH06, BdlP06, BKN01, CKLX09, CLOvdD01, CP09, DMD03, DHZ03b, Eld02b, FT05, GAM06, GI02, Guo04, HV03, HTV02, HZ03, KN06a, KS06a, KTKR06, Kru04, Lan08, LMMT01, LM09c, MM05, Pre05, Sat05, VC04, VJ06, Zav08]. quadratically [ZZY06]. Quadrature [DS03a, DP02b, DD02, LRW08]. quadrics [CMP02]. quadruples [BZ05, CGPM01]. Qualitative [Pry09b, JLY01]. quantitative [Vis00]. Quantization [SSW08]. quantum [And07a, BLdP04, BLdP05, DLMO07, FH07b, Fuj05, GII08, HOZ01, Kri05, Par02b, PR02c, PR04c, Ter01a, ZZ05]. quartic [Ahm05]. Quasi [GJMR01, GHM03, HNV00, OPS07, Ohn08, Suc09, ALdV05, AdLV06, AD06, BJ07, CMP03, DJMS09, DJ07, GT08, HB08, LL08d, LM09b, Zob00, vDP08]. quasi-commutative [LM09b]. quasi-death [vDP08]. quasi-definite [CMP03]. quasi-diagonal [ALdV05, AdLV06]. quasi-equivalence [Zob00]. Quasi-filiform [GJMR01]. Quasi-isometries [Suc09]. quasi-LU [BJ07]. Quasi-Newton [HNV00]. Quasi-orthogonal [OPS07, Ohn08]. quasi-oscillatory [GT08]. Quasi-real [GHM03]. quasi-regular [DJMS09]. quasi-separable [AD06]. quasi-thin [HB08]. quasi-tree [LL08d]. Quasiconvex [Jan01]. quasihyponormal [TU02, TUC04]. quasiseparable [EGO05a, EGO05b, EGG07, EG08]. Quaternion [ASK^{+}08, CT09, Fan03a, Hua00b, HL07c, HL08b, HL08c, Lip01, Rod07, Str03a, WL09b, WvdWC09, WZCL08]. Quaternionic [– DS08, ARR06, AYC06, BM04c, HS01, PRV05, Rod08, Str06, Zha07b]. Quaternions [GTT01, FGT03, FP03, KKABAF06]. question [Ben08, CL07a, FL07b, FKN02, Fur05b]. questions [Gen00, Hua00b]. queueing [BLM02, Gra02]. queues [BILM06]. Quiescent [Had08]. quiver [Boc03b]. Quivers [Bad08, Hil03, LH03, Zhu08]. Quotient [Dax03b, GV09a, ALC06, FS08, LWZ09, MA03, Ney01a]. quotients [BZ02, KA06, Wal05a]. qustochastic [Co08].

R [Bot06a]. Racah [Ter04]. radial

[BDG08]. Radiation [Ano06f, CBIXG08, CGXL08, HC08, SK08b, TVB08]. Radical [MT04, BBL^{+}05]. radii [DL03, EFN^{+}03, GŞV02, Guo08, LEL02, MGWN06, NT01, Shi07b, YLT05b, ZSL09]. radionuclide [TKX08]. radiotherapy [EHR08]. radius [And06, AHH^{+}09, BBS08, BNT05, Blo08, BKPW08, BH08a, CT03b, CZ00, CL04a, CN07, Das07, Dra06, Dra08a, Dru08a, FYZ07, GL08, GS08, Guo03b, Guo05a, Guo05d, Guo06a, Guo06c, GS06, Guo07b, GW09c, Ho03, HHZ09, Ind09, Kar09, Koz09, LP07, LSC09a, LR09, Lin08b, LLT04a, LLT04b, LLT07, LL08d, LLT06, Lur06a, LY09b, Mae08, MK05b, MW01a, Mol05a, Mol08b, Nik06c, Nik07c, Nik09, ORvdD02, PJ08, Pep08, PW08b, Roj08a, SS06d, Shi04, Shi07a, SS07c, Shi09, SW04, SAH08, wThY09, WHAB08, Wir02, Wir05, WXH05, WGT09, XHSZ09, XX09, YLT04, YLT05a, YL08, YSL08, ZWS09, ZLS09, ZZ08b, Ziz04, vD07, vDK07]. Radjavi [Ano02r, BORŠ04, CL07a]. Rado [CKS06]. Radon [Eld02b, SvdS01, TZ01]. Rajendra [And07b]. Ralph [Meh08]. Randić [DK07, Rod05]. Random [BD09b, Chu07, FMN04, Kly00, Mat07, AG00, AEK06, Bol04, GN00, MW08b, MT05, Olk02, Pic07, She01, WA09]. randomized [DM07]. Range [Dug00, Dug01, Abd05, BS05b, BdPT08, BS04c, CSSZ01, CC01, CL00a, CN01, CN02a, CN02b, CNP02, CN06, DHK06, Dug04a, Dug04b, Dug05a, FQ04, GW03a, GW04b, GW04c, GJK04, INOY03, JS07a, JO02, Kru04, KS06d, LMMT01, Leš02, LTW02, LP07, MPT02, NP01, NBdP08a, Psa00, RS05a, SS07a, Sed01, SDL06, Wu02, Yam02]. Range-kernel [Dug01, Dug04b]. ranges [BLdP04, BLdPS05, BM04c, Cal08, CT01, CKŻ06, GW04a, GW07, GW08, HHW00, KBdP07, LZ00, LPS05, Lin02a, Mat01, MBLV01, NBdP05]. rangespace [Alf06a]. Rank [AGK06, AH02, AP03b, Bal00, CGH04, DV07a, DV08, HC03b, Jør05, Lim05b, Sed07, SMS09, TS01, YLZ06, AAB^{+}07, AFS09, AG04a, AHK^{+}05, AHL^{+}09, BB04c, BB03a, BFH04, BF07a, BFS08, BM01a, BGL09, BW00, BS02c, BLS02, BGJS06, BBG09, BL09b, BS00c, BR06, BFH^{+}08, BBC05, BW01, Bra06, Buc06, CRU09, gCZ03, CHNP04, CHY08, CKŻ06, CL05b, CL06, CFP03a, CCM00, DoMP09, DHH^{+}06, DV04, DD09, EOvdD03, FH07a, FP07, Fon00, Gem08, GQ09a, GP07, HHL^{+}09, HHY04, HR04, Hog08, HBLZ06, JMN08, JLL03, KN06a, KPV04, KR06, KR08, Kuz02b, LR05a, LŠ09a, Las06, LZ09a, Leg04, LRŠ02, Lim05a, Lim06, Lim09a, LYY07, LT03b, MV07, MA06, MG03, OS09, OTZ09, Pan00, Pet09, Pic07, RP00, Sav06a, SM09a, Shi04, Smi01c, SD00, Ste09, Sud04, Tan05b, TY06, TX09]. rank [Wat01b, YT04, Yua00, Zha03a, ZW03, ZyC03, Zha05a, Zha05b, Zho03, xZxGpL09, Zit09, tBSR04, tBS06]. Rank- [CGH04, HC03b, DV04]. rank-1 [BBG09, Tan05b]. rank-additivity [CL05b, YT04]. rank-constrained [ZyC03]. Rank-deficient [DV07a, DV08]. rank-nonincreasing [HHY04]. Rank-one [Lim05b, BS00c, gCZ03, CL06, xZxGpL09]. rank-one-modifications [BB04c]. Rank-permutable [AGK06]. Rank-preserving [AH02]. rank-revealing [Pan00]. rank-structured [Gem08]. Ranks [Ano03-31, CGG02, Dan02, LTT09, WL09b, dSZ00b, BSKS05, BGJS06, CtBDC09, FIS02, Hou03, Tia00, Tia02, Zha04c]. Rao [Liu02b, YLT09]. Rate [LG04, LZ09c, Til03]. Rate-tilting [LG04]. rating [Qui07]. ratio [FST04, Fur03, Luz07, Seo00]. Rational [AHK^{+}05, AHLR09, CHMM06, AI00, AGI00, DH01, Dra07, ES02, FH04, FKL03, JB08, KVV09, LRW08, LCF06, LPR04, LR00, OS09, OA09, OZ03b, OZ03c, PFL05, Sak01, WY07]. ratios [BF08, FK07, Wu05b]. Ray [SSY05, CJK05, Fan04, Hua09, LHS02, LR05c, LHS05, LSH08, MS08, SS05a, SLR06, SBS02, Stu03]. ray-nonsingular [Fan04, LR05c]. Rayleigh [Dax03b, FS08, KA06, MA03, Ney01a, Ovt06a, Ovt06b, Ste01b, Wu05b]. rays [BLNP01, LM06b, Pat04, BLG01]. Reachability [BRVS08, NT09c, CDMGP06]. reaction [BKK08]. Real [BL09b, Cof03, GŞV02, Gra02, Huh06, Wat06, Wri06, BSKS05, BCSV06, BKU06, BMS08a, CL07b, CL09c, – DI02, Dra08a, DH04, FGR07, GHM03, GG07, HPP00, HT05b, HvP05, Huh07, JGPI08, JR00, KSU03, LŠ06a, LR05b, Lás03, MPS07, MS00b, OB01, Pet09, Rod06, SBM05, Uhl01, Ver08, WvdWC09, Wer08, Yin04, Yin05, ZX05a]. Realizability [Sot06, Alf01, SBM05]. Realization [AAV06, AM05, BHdST06, CHMM06, FH00, HMH01, KOvdD07, LŠ06a, MA07, VWD08]. realizations [AD06, AMZ07, AHK^{+}05, AC04b, BGKR05, CF02, FN00, GLS07b, HOP06, KS06b]. reciprocal [Dah05, WO05]. reciprocation [LM09c]. Recognizing [Bol05]. recombination [Daw02]. reconstructing [BKDL01, BLG01]. Reconstruction [BDD^{+}01, KN01, RT06, SvdS01, Cen01, CCSS03, FLM02, FNC06, FB04, Hei02b, LCN04, Mas09, NCCY00, Pop08, SLP01, Sci02, SS08a, WX07b]. reconstructions [VHK01]. reconstructor [CPPS00]. recovering [Phi04]. Rectangular [EFS01, JO05, Bor08, CRU09, GT02a, GT08, MP05, NY07a, Pfa08, Pop01, YTI06]. recurrence [CHRS03, EGO05a, EK08a, EK08b, TR01, dSR05]. recurrent [BP02]. Recursion [KT08, LdlP09]. recursions [DBK05]. recursive [MSS08, YC00b]. Recursivity [dS05]. Redheffer [SN09]. reduce [FL04, FLPW07]. Reduced [IK06, LV06, BF01, BT04b, CHNP04, CFP09, Sav06a, VPV01]. reduced-order [BF01]. reducedness [KPP07]. Reducibility [AI00, GW04a, DHH^{+}07]. Reducible [LHS05, Stu03, AH08, CL09b, CN06, LY06, MSR06, Wu05a]. Reducing [TKX08]. Reduction [Ano02-29, BGM04, BS06a, BBD05, BFSV06, BP09, DV04, EGG07, FKLN07, Fre08, GVV03, GP07, GA06b, Gug08, HL07c, HL08b, HL08c, LRK06, Neu01, Pen06, PPF02, Ral03, RW06, RW01, SL06, Sav06a, SA02, Sty06, TI08, Tri02, WSW06]. reductions [LCF06, She03]. reductive [BS02b]. redundant [BDM03, CKS06]. réelles [BCSV06]. Rees [Vil08]. Referees [Ano02y, Ano03-47, Ano04-45, Ano05-32]. reference [Tak04]. refinable [Thi04]. refined [Wu06a]. refinement [ALdV05, AdLV06, Nom05a, Wad07, Wd09]. refinements [Car09]. reflections [Wri06]. reflective [ADC08]. reflectors [MMT04, SS08b]. reflexive [yPyH03, ZX05b]. Reflexivity [Bra07, BK09b, DR08, MŠ04e, You07]. reformulation [MR02b]. regime [Ser09]. Regina [Ano07s, BGKW07]. region [Li01]. regional [LSF09]. regions [BI03, HSZ08, Lin02a, LSH08, SS05a, SLR06]. registration [FM04b, HHM09]. Regression [SvR02, AS02, Liu02a, RvdGM05, Sav06a]. regressors [FK04]. Regular [HL03, KN00a, SKBS08, AJF09, AH09b, Bal00, BES09, BBG09, Boc03b, CL00b, CP07c, CP08, DJMS09, DGN02, DD09, DKX03, EK08b, E^{+}00, Gav03, GHI09, Hae08, HD03, Jør05, KPK04, Kir03a, KMMZ04, KD07, MSS07, Mik09, Ned01b, Pac02, Pas01, PT06a, Pat01, Peñ03a, Row07, Sch08, SdAdFDV07, Ter02, Wan04, Wu06b, Wu08b, XJ08, Zha05c, vdW02]. Regularity [CH06, CR07, Fio01, Fio05, GP03, PP01a]. Regularization [HFR04, HS06b]. regularized [FB04, LCN04, MLK^{+}04]. regularizing [Est05]. Reich [Yua00].

Related

[Rod09, AKYD16, And06, AD08, BB05a, BZ05, BDG08, BD03b, CK02, CLOvdD01, Dah09, De 07b, DG03, DM02, FGM03, Fur05b, GHS06, Gen00, GP00a, GS00, HP09, Hog03, Kol08, LA05, LB01, MHZU05, NBdP08a, OOT06, SdSW07, Šem03, SHZ05, Tyr04, WW07, Zav07, dMFZ09, vdVN00]. relating [Dat08, Tan09a]. relation [CHRS03, Fer04, KKP07, MW01b, dSR05]. relations [AAB^{+}07, BFZ04, EGO05a, GAM06, GPMT06, Nom05b, Oku05, SdSW05, SdSW07, SdS09, Stu00, TR01, Vid07, Zem06]. Relationship [GPMT06, Li03, CE09]. Relationships [BB04c, BHLL04, BBL03b, Nab07]. Relative [EMBV03, Li00b, Nak03, ST00, ST03b, TS06, WLBZ06, AFK02, AS02, BLdP05, DNR04, DMM00b, FYK05, Fur05c, Fur06b, Ips00, Ips03, JS00a, JDS^{+}03b, Kos00, LS00a, Mue09, YKF05]. Relatively [PD00, Sch08]. relaxation [BRM01, QX08, XZ06]. relaxations [DS08b, Hil07a, Jan01]. relevant [vdVN00]. remark [Alf06b, Fal04, PH00]. Remarks [Gro06, Lie09, Ste03b, SS05b, Nie08, SvdD05, Zob00]. reorthogonalization [MS03a]. Reparametrisation [Kob00]. repetitive [PGRO06]. replicas [BBFP04]. Report [Pue01, Uhl04, LWZ05]. Representability [JP04]. Representation [Ano03-43, CC00, Bor08, CM00a, CG00a, CI08, DD05, FGT03, Kul01, MM05, SD00, SH09, Xia06, Zhu08, fur08b]. Representations [CGD05, CXW09, FKRS09, LR05a, AMZ07, BS00b, Big02, BKU06, BD04, CMV05, CB06, CD05, Dam03, DSW07b, DW09a, DP04, DHK06, Han05, HL05a, Hog08, IS09, Kar04, KV09, LW07, LSS00, MV07, MR02c, PD00, PEPS00, VW00, YY03]. reproducing [ABK02, BP06b, ĆDR01, DD07]. reproductive [HL05b]. require [AHL^{+}09, HLW01]. Requirements [LRK06]. Research [Ste07b, Ano05-31, Jef05]. residual [TS06]. residue [ABR02, EK08a]. residues [BES02]. resistive [BCEG09]. résolubles [BCSV06]. resolution [CCSS03, FB04, LCN04, NCCY00]. resolvable [BCSV06, BCSV06]. resolvent [BS00d, DH03, EFN09, EFN10]. respect [ARR06, CLPT06, HG09, JNL05, LS02d, Mar02, MR02a, Mue09, NT07b, Ter01b, Wu08c, Yan01b, ZGL09]. respiratory [TKX08]. response [MMX04]. restarted [AEEG08, MZ06, Wu06a]. restoration [NPP00]. restorations [HNV00]. restraints [Ste05b]. restricted [Cal05a, Cal07, CM08a, LZ08a, OO00, YYZ05]. restricting [AP03b]. restriction [ZGL09]. restrictions [Amb07, BZ02, Bou03a, CS00a, MSF05]. result [Aga02b, Ala00, Bel05, Bot06a, Fao01a, FG03, GRB04, Kar05a, Lur06b, Mat03, Nie07a, Par09a, SMK09, Stȩ00a, Wu06c]. resultant [AVV05, Gem03, KMS05a, WO05, Win00, Win03]. results [BBLT08, BMS06, Bot09, Cap01a, CR07, CIDW06, Din03, DPZ06, FM04a, GH00a, Gon05, GP00a, Hag02, HWW01, LS01, LSC09b, MPT02, MN08, PW08b, RR04c, Šem03, Tao06, WM02, WX07b, dA07]. retrieval [Hen07, JB00]. retrieve [AEB06]. retrospective [TKX08]. Returnability [EH09]. revealing [MG03, OS09, Pan00]. Reverse [BS07, Fur05c, WG02b, BB05a, BLFS09, FFN^{+}07, Fur04, Fur06b, Lee09a, MF09, TY06, Wan02, WWJ09]. Reversed [Har04]. reverses [Dra05, FNT07]. reversible [DHFS00, MSF05, Zha04e]. reversing [PM05]. Review [And07b, Bar05, Böt06b, Kol05, Moh05, Par09b, Rod09, Tům06, vdD09, SS09d]. Revisitation [BT09c, BBT03]. revisited [AGKM06, BK08, Gro00b, Ikr01, MR02b, PM08, Smi08a, WG02a]. Revisiting [FF04a, FF04c, XXZ09]. reward [BF07b, THH04]. Riccati [BILM06, Bra03, CFL05b, DSW07a, DD07, FPP01, FB05, Fre02, GRV06, Guo02, Guo07a, HP07, HL09, KTKR06, KAP^{+}03, MM06b, gS02, Xu07]. Richard [vdD09, Ano08r, Kol05]. Richardson [Ful00, Had05, HR03b]. ridge [TY08]. Riemann [Zus05]. Riemannian [BH06, HP09]. Riesz [ACRS05, CL02c, Dug05b, HJM09]. Right [LZ09a, BBL03a, DMW08, HR07, HL05a, LSD06, MV07]. right-hand [DMW08]. right-star [BBL03a]. Rigid [GPMSZ06, BCSV06, CS03a]. rigides [BCSV06]. rigidity [Alf00, Alf01, Alf08, Cod00]. Ring [HZ04b, LM03, BBC00, BM06a, Boc03b, CT03a, FM00a, GKvS04, HC06, HL07c, HL08b, HL08c, KPP07, LN08, OWY07, Pat01, PP04b, PG04, WY06a, WYZ07, WL09a, Yah06]. Rings [KZ03, Abd08, AGHM04, AR06a, ALC06, AY08, AH09a, Bal00, BS02c, CW00a, Cao01, CZ05, CS09a, CHASG02, CC04, DK08a, GR02, Ish04, Ish06, KDC07, LJ04, LT07c, MW09, Min01, MDK09, MD09, NDS02, PP04a, Pro05, RRY04, RY08, Roc09, SSSG06, SS07a, SSSG08, Tag04, TTM02, Vil08, WY04, Wan04, WY06b, Wan09b, XCX01, YZ07, ZY02]. Riordan [CKS08, CK08, LM09c]. risk [Mar02]. Ritz [GV06b, KA06, Ovt06a, Ovt06b, PS08, SvdE03, Ste01b, VV06, Wu05b]. road [So04]. Robert [BN00]. Robertson [GII08]. Robertson-type [GII08]. Robust [CCZ06, DW05b, BM04a, BBMX07, Bol01, DS08b, GM09a, Ovt06a, Ovt06b, PD00, PA04]. Robustness [MW08a, MW08b, Ovt06c]. Roger [Ano05r, Ano07t, Ano07p, Ano07u, BKMZ07]. role [BF07b]. Root [BdlP06, KB02a, ABB00, BMP00, EvdD08, KKK07, Lu02c, MV04, Roj08b, Sav04, Kol08]. Root-induced [BdlP06]. rooted [FGM08, Roj06a, RR07b]. roots [AFM03, BPT09, Brü03c, Cho04, HNX03, Ikr01, JOR01, JO02, LN04, Sav03, TZ03]. Rosenberg [LEL02, Nou08]. Rosenbrock [Lom02]. Rost [Ros06]. rotation [LR03, Ste07a]. rotations [LR00]. rough [Ho03]. Rounding [SS00, SS03c, DS04b, Ois01]. Rounding-error [SS00]. Routh [Hol03, Peñ04]. Rovnyak [AD03b]. Row [BBD^{+}08, ACV09, BdHZ05, BdHZ08b, BdHZ08a, BDD^{+}01, CL05a, CH00b, DS07, HV08, Hol02, Joc09a, Joc09b, JS00b, KPP07, MW03b, O’L03, Sbu05, SH00b, VPV01, BGJS06]. row-by-row [CH00b]. row-stochastic [CL05a]. rows [Dod05, YH04a]. RSVD [CD00c]. rule [GAE02, HL04b, Sun08]. rules [FP02b, Ji05, KMZ00]. Runge [Dek09].

S [Ano08r, Kol05, FJ02]. Saad [Ste01b]. saddle [BW08a, XWD06, ZBY09]. Sakamoto [CL09a, Li00a, Mat03]. same [Che09, OvdDV00]. sampling [Ho03, Ziz04]. Samuelson [Nie07b, TP05]. sandpile [HWC06, HLW08, SH08]. sandwiches [MN07]. Sanzheng [Moh05]. Sarason [Ala00]. satisfied [BHJ08]. satisfy [BLS02, Zha03a]. satisfying [Bir04, CCH02b, CGPM01, Dug06, EK08a, JK06a, JMP09]. scalar [Dru07, Gu01, Kin05, KRS03, MMT04, Van04, Wan06]. scalars [BM04c]. Scale [Ano02-27, Ano02-29, BFSV06, GRV06, GA06b, Gug08, MPS06, OS05, RM09b, SL06]. scaled [Mat02, MH06, Win00, XQW08]. Scaling [O’L03, JK06b, JMT05, JLN08, Kal03a, SSB09]. scattered [GS04b]. scattering [BGGC06, CSK01]. Schatten [AMS05, BK00a, Joc09b]. scheme [Est05, HMR09, YLX09]. schemes [GWL09a, GWL09b, HNT03, HB08, Pry09a, Tag04]. Schmidt [Dax00, HK00, Neu01]. Schrödinger [AH09b]. Schur [JLN08, Poo08, ABK02, AADL03, ADL04, ADL06, BS02a, BBS04, BGKR05, Bon03a, BZ03, CMS02, DW08, Fan02, FGL^{+}00, HOP06, JMT04, LP07, LHZ04, LH04, LLHK08, MS04c, MVV08, Mat05c, McE02, MSW09, PHO07, Reb07, SG09, TT09, Zha00, yZtLC06, yZXtL07]. Schur-based [MVV08]. Schur-multiplicative [Poo08]. Schur-type [JLN08]. Schwarz [BPS04, EH03, Las06, Lee09a, MS04b, Wad07, zY08, YLX09]. Science [Moh05]. Scientific [Bar05]. Scott [Han00]. scrambling [AFS09, AK09a, AK09b]. seamlessly [Nie01a]. search [Pry09a, TW05]. seaweed [Dvo03]. secant [Ano03-31, CGG02, McG06]. Second [CLVV06, BLLS06, Dun04, GPMSZ06, GN00, GT02b, GT04, Guo05c, Guo05b, HZM02, JY07a, JY07b, Lim05b, LYY07, LS09c, MM05, SL06, ySZyY06, Sta07, Sta09b, SS06e]. Second-order [CLVV06, Dun04, GPMSZ06, GN00, HZM02, LS09c, MM05, SL06, SS06e]. Secret [CGHG06]. section [FMN04]. sections [ABDD01, LRR04]. sectorial [AT00, AP03a]. segmentation [GMR^{+}07]. Segments [NS08b]. Segre [Ano03-31, CGG02]. Seidel [BPT09, LS00b]. Selected [Bar05, Gri00, Wu06a]. selection [EHR08, IL06, MQ06, dHM07]. Self [Rum01b, CWW09, Del01, GHS09, HZ05a, Lan08, LPR04, MPT04, Mol06b, Pro08, RA08b, SG06, YLZ06]. self-adjoint [CWW09, Del01, HZ05a, Lan08, LPR04, Mol06b, RA08b, YLZ06]. self-adjointness [MPT04]. self-congruent [GHS09]. self-dual [SG06]. self-similar [Pro08]. Self-validating [Rum01b]. selfadjoint [BR00b, BOP08, GMP08]. semantic [Hen07]. semelparity [DDvG05]. Semi [vdMSR03, ACG08, AM04, BDKW04, CG00b, DW05a, DS08b, Dom00, HM06, LS06b, Min01, NCV08, Suc09, Wu06a, XS08, ZBY09, vdH03b]. semi-continuity [HM06]. semi-convergence [AM04, ZBY09]. semi-definite [DS08b, XS08, vdH03b]. semi-elliptic [NCV08]. semi-free [DW05a]. semi-Hilbertian [ACG08, Suc09]. Semi-infinite [vdMSR03]. semi-invariants [Dom00]. semi-Markov [BDKW04]. semi-refined [Wu06a]. semi-separable [CG00b]. semi-triple [LS06b]. semibounded [WW08b]. Semiconvergence [Wan07a]. Semidefinite [Hil08a, Kal03a, AHW05, Aud06, CLL07, Dat08, Eld06a, Elf04, FS03, FGR07, GSR03, HHL^{+}09, JMN08, LV06, LWWZ08, Liu02b, MRR07, Qi09, SDL06, TW05, Val09]. Semidefiniteness [JR00]. semidiscretized [Sty06]. semifree [LK00]. semigroup [Buk07, CBD^{+}04, Eld06b, Eld09, HL04c, KB02b]. Semigroups [Leš00, Okn01, ALMS06, Bel05, BO05, BAGS06, CFL02, CV07, Dri08, DLM^{+}00, DK04, Dup03, Lip08, Liv04, Lur06b, PP04b, Ros05, Ros09, Ves03, Yah04, Zal03]. semimodules [CGQ04]. semimonotone [Chu03, MND01b]. seminonnegative [NM04]. seminorms [GN05, Lav03a]. semiregular [Sat06]. semiring [HCLC09]. semirings [BGJS06, CK03, PH04, SS09c, SKBS08]. semiseparable [FG02, PVV08]. semisimple [AGHM04, Kle09, NY07b]. Semispaces [NS08c]. sense [Yan02a]. Sensitivity [Kir05, AB05a, CM00c, DBK05, Igl01, Sun05, XD03]. Seo [FG03]. separability [BIL06, PR02c, PR04b, Woe04]. separable [AD06, CCG02, CG00b, Hil08a, Myš05, Myš06]. separating [Lem04]. separation [CGQ04, vdH08a]. separatrices [Reb07]. sequence [BC00, BP02, BRZ04, De 00, Kir00, OT01, PG01, RR04b]. sequences [APESS06, BD06a, BB05c, BL09c, Cap01a, Cap01c, Cap03, CK02, CHKS05, EK08a, FKR09, KL07b, MS02b, NT06c, NCV05, NCV08, PT06a, SS09b, TR03]. sequential [Cen01]. Series [Moh05, AG00, APESS06, AH04, BZ02, BM04d, CTW01, DL04, DS08c, Dom00, FQ04, HW00, KMZ00, Kle00, KS01, KS06c, KS09, Kol05, Sak01, Ste08]. Serre [AC08, LB01]. sesquilinear [FS07, HS06b, HS08, MR02a, Tup08a]. Set [Sha03, AKGC09, Bap00, BD03b, DB06, EE04, FPP01, Har06, KM04, LRŠ02, LCP02, Mol02, Mou06, Nar05, Pat04, PP02, Ros09, She00, Shi04, wThY09, dW03]. set-chains [Har06]. Set-systems [Sha03]. Sets [And04c, DL92, KS06d, APESS06, AGM09, AM05, BKDL01, BLNP01, BIL06, BM04c, Boc03a, Bon09, BG02, BLG01, BDD^{+}01, CT03c, CL08a, CL02c, CS04, Dax06, De 07b, Flo09, GK07, Gro08, Hol08, Hon02, Hon06, JP04, Kol08, KMS05b, KN01, LRY05, LHL08, Mel09, MZ00, MT06, NS08c, SSG04, Shi01, Smi08b, Tre03a, dST05, dST06, sV06a, DL01]. setting [ADL06, ATV07, Har00a, KVV09, SCS09]. settings [Boc03b]. Several [FI05, Liu02b, BM04c, CS00b, CS03c, CS03d, Han02b, HZY05, McC01, Nab07, Pop02, Sak04, Wal05a]. Shanghai [Ano09p, IMS^{+}09]. Shannon [Fur04, YKF05]. Shape [Wan09a, BM01b, CPR03, NP01, Omi09, Vid08, WX06b, WX06c]. Shape-preserving [Wan09a, CPR03]. shaped [Ned01a]. shaping [Eld02a]. Shapley [XDS09]. sharing [CGHG06, CL03c, DKS04]. Sharp [Das04b, Ind09, NT08a, Pan02, SW04, XX09, CCRS08, DB05, Das09, Gro06, KN03b, SHWR02, Zha04f]. sharpness [BS00d, Kol08]. sheaf [Wym02]. shell [CN02a, CN09]. shells [Dau08]. Sherman [Map07]. Shift [SS07b, AD03b, BR02, BRVS08, CG00a, CN07, GX09b, KL07b]. shift-invariant [KL07b]. shift-similarity [BRVS08]. shift-splitting [GX09b]. shifted [HSZ08, Mat05c]. shifting [CHASG02]. shifts [CLY08, DMW08, ZZY06]. Shmuel [Ano09r, BKR^{+}09]. Short [Ano09r, BKR^{+}09, Ber09, De 09a, KN03b, Kra03b, Li06a, Str01]. Shorted [MB07, ACS04, ACS06]. Shorting [GMP08]. Should [GC04]. shrinkage [HDA09]. SIAM [Par09b]. sided [ABR02, HS03, Kol08, NP07, PP01b, RY08, Ral03]. sides [DMW08]. Siegel [EV05, FF04a, FF04c]. Sign [AT09, AHL^{+}09, BC08a, CP07c, GLS07a, Hua08a, Hua09, Kak08, SG03, AHK^{+}05, BOvdD07, COvdD09, CV05, CL08a, CHMM06, CP08, DHH^{+}06, EOvdD03, Fie05c, GL01, HLW01, HLR04, HK03, HL01, HKKZ03, HKKL05, KOvdD07, KD07, LRY05, LY06, MTvdD05, ML02, Peñ03a, Rod06, SH00a, Sha00, SH00b, SSG04, SS05a, SLR06, YSS07]. sign-central [HKKZ03, HKKL05]. sign-nonsingular [Fie05c]. Sign-solvable [Kak08]. Signal [Ano02-28, Baz03, BBP03, DV04, GGCB07, HLW06, LXW04, PA04]. signals [Yer04]. Signed [PW02, KS00a, KS02, Ma09, Sha03, SS04b, SS01, SH02, SS02, SHS03a]. Signless [CRS07b, CCRS08, Omi09, WHBM09b]. Silverman [FPP02]. similar [Hei01b, Hua08a, NM04, Pro08, XDS09, ZM03]. Similarity [BL05, DH04, HL07c, HL08b, Ji03b, JW06, Rod06, Stu00, Tup09, Wim01, AMZ06, BdHZ08b, BdHZ08a, BBC05, BRVS08, DJ01, – DZ04, – DS08, DHB07, Flo09, FMS01, FJ02, FMS05, FJ05, GW04a, GI02, Ji03a, Kob05, LP09, Mil03, Pet05, Pro05, Tov00, Tup08b, VV06, Ver08, HL08c]. Similarity-preserving [Ji03b, Ji03a]. Simple [CK08, GH03, Har00b, AR06a, BTT09, CL09a, Cio06, Dax03a, Fie05b, GS09c, LZ08b, Li00a, Per03, WLBZ06, XM02, Yam01]. simplest [LR04]. simplex [BKK07, HNX03, vdW01]. simplex-like [vdW01]. Simplicial [CPR03, BKK08, BD08]. Simplicity [TBK09, SK08b]. simplified [FS08]. simply [BES02, DDI03]. simulation [LG04, TKX08, VM04]. Simultaneous [CHPP06, Dob04, Flo09, AL09a, CJPS02, Cen01, CH00b, – DS08, Dub09, FNC06, GW02, MS06, Pet08, Yah08]. simultaneously [LPG03, OV06a]. Sinc [NB03, Ste08]. Sine [JL03, HV01]. Singapore [Bar05]. Single [LP00, CHN06, HW05a, KR05]. Singular [AC09, Bos07, Bou03a, Bou05, DS09, GRV06, Kit09, Mir00, sV06a, ASX08, AGO09, Aud07, Aud09a, BX08, BK09a, BR07, Bra06, Bri01, Bur06, But06, CW02, CD00d, CM08a, DDM08, DJM06, DM07, Eis06, FG02, FH00, Fer03, Fer07b, Gao03, GX09a, GWL09a, GWL09b, GW09b, HB00, Hoc09, IWL07, JSWT05, LX08, LS00a, LHL08, LWZ09, MS00a, MK05a, Mor08, NS07a, Neu01, Nie08, Ois01, OS00, SS03a, She01, Shi04, Sid01, ST03b, SvR02, Ste06c, Tao06, rWQ02, Wan07a, WGL09, WD05, WLBZ06, Xia06, ZBY09, vDS05]. Singularities [KVV09, BB03b]. Singularity [KS06a, CH05, Kol03b, VWJ08]. singularly [AH04]. Sir [Ano03u, Mar05]. size [HL02, Kir05, NBdP05, Nik09, OTZ09, Val02b]. size-classified [Kir05]. size-independent [OTZ09]. Skew [AMP07, AM05, BGN08, BLS05, FI05, Hil08b, HHZ09, HW05b, Ikr01, Kah06, KMS05c, LTT09, MHZ05, pPyHZ05, PT03, PT06b, Rod08, Tre04a, Tre04c, Tre04d, Tre04e, Tre05, WSL02, Zav08, MQ06]. Skew-coninvolutory [AMP07]. skew-Hamiltonian [FI05, Ikr01]. skew-hermitian [Rod08, BGN08, HW05b, LTT09]. skew-symmetric [BLS05, Hil08b, LR05b, MHZ05, PT03, PT06b]. skew-tableaux [AM05]. skewsymmetric [HR01, HR03a, Rod07]. skip [BM04d]. sliced [tBS06]. slices [SMS09, TBK09]. slope [SS07b]. slowly [LRR04]. SM [CNTV08]. Small [BL02, LL02, BO05, CNX05a, FL03, Fer02, Hig06, Kah06, Pet09, Ste06c, Živ06]. small-world [CNX05a]. smallest [AKYD16, MNOvdD08, MVV08, Mor08, SS03a, Zha04e]. Smith [Cao09]. Smooth [Reb07]. smoothness [Thi04]. Sobolev [DM02, KLY09, MS01a]. Society [Ano03-46]. Söderlind [PX03]. solid [HD03]. Solov’ev [Hof00]. Solution [BKR06, FLMR00, Sch09a, BRM01, Bal07, BPV00, BM04a, BILM06, BSS09, CX05, CH00b, DSW07a, Dra07, EF07, GHS06, Guo02, Guo04, HT09a, HV01, JKL09, Kar04, KS05, LS07a, LS07c, Map07, MS04b, MA07, MZ04, NN01, SS02, Sid01, Sty02a, gSX03, WL09b, Xu01, ZD05, vdW03]. Solutions [CKLX09, FYY09, YLF09, AT00, And04c, AHLR09, AYC06, Bar08, BT04b, Brü09b, CC09, ÇS02a, CJPS02, CSY01, CL03c, CF02, CP05, CM08b, Dob01, DL08, FPP01, GRB04, GM07a, Gro00b, Has05, HES06, HI06, Hua00a, HL09, Iva05, JW03, JM09, KS00a, KTKR06, LŠ09a, LM04a, LG03, LTT09, MT02, MSS07, MHZ05, OF02, yPyH03, hPyHZ07, PA04, PCVT04, RR02, RSW05, Sha03, SdL03, WSL02, Wer09, WJL02, Xu08, XSG08, Zha03b, ZyC03, ZF07, ZHZ08]. solvability [CLPT06, CHP03, CHJP05, Hua00a, Myš06, Wan07b, XHpS06, ZHZ03]. solvable [BPP08, Cai03, CS03a, Eic04, GSR03, Kak08, Sha00, SSY05]. solve [Log05]. solved [RS05c]. solver [Gem03]. solvers [Bér05]. solves [FS08]. Solving [AGI00, BLM02, BLM03, JB00, KRS09, LL06a, Lim09c, xZxGpL09, BKO01, DS08b, FK04, GX09b, HC08, JL03, Kro01, LWZ09, MS01a, MM07, MV07, Ned01a, SP04, VHK03, WL03, YLX09]. Some [AAB^{+}07, Auj03, BBM02, BS05a, BE01, BMS06, BMS07, BG09b, BKLM06, CC09, Che03, CJ05, CL03b, CT09, De 09c, DW00, DPZ06, Dra05, Dra06, FI07, Fie01b, GH00a, GS03a, GST04, GTM09, GLP00, Hag02, Han09, HWW01, HP04b, Hua08b, JMT04, Kon08, LSC09b, LH04, LLHK08, Mah07, MP05, Nie08, NT06c, PM05, PP01b, Phi04, RM01, RR04c, RS01, SvdD05, TR02, Vol04, WM02, WM03, WX07b, WHAB08, YZ07, Zen01, ZZ09, Zob00, AR06a, AR06b, BBG06, BLNP01, Bér05, Bot09, Bou06b, Cal05b, CLB09, CT01, CS05, CM08b, DS03b, – DS08, DR08, FGM08, FM07c, Fur08a, GR07a, GSR03, HGC07, Hog00, HQ08, HLS00, JNL05, Jam07, JW08, Joh05, Kar09, Kle07, KMS05b, LM04a, LSC09a, LB01, Mal09, MM03, ND05, Nou06, Pry09a, Rao00, RS03a, RS06a, RS05b, Roj06b]. some [RR07b, RT01b, Sch08, Sch09c, Sed04, SHZ05, Shp06, Smi01b, SvW05, SBM05, SZ04, TRH01, TS09, Tup09, Tyr04, Val02a, Wad07, Wu05c, Wu08a, WSL09, XSG08, Yan09b, Zha04a, ZZ08b, ZX05b, dFP01, dW03]. sont [BCSV06]. SOR [BD04, Yua00]. sort [Kle09]. sort-Jacobi [Kle09]. Soules [CHN06, CNSM07, CNSM08]. sources [AFCC04]. Sourour [Bot06a]. southeast [GV09b]. space [ABP09, AAGK06, ATV07, AFK02, AHZ05, And07a, BGKR05, Bas02, BB09, BFPW02, CCH02a, DK08a, De 08, De 09c, DKS04, DD05, Dug05b, FYL07, FQ04, Fur04, FS07, GPMT06, GLS07b, Gro01, GS09e, HO02, HK07, Hir09, dH07, HS09, JS07a, Joc09b, Kar04, Kim05, Kit09, KN00b, KS06b, LL06a, LS02c, gLxW02, LwW07, Mol03, NBdP05, PHO07, Str06, XM02, ZZJ08, Zip03]. spaced [Hol02]. Spaces [Pet09, ABK02, ABR02, AD03b, ARR06, ATV09, And04a, And09, ACG08, AH03, BB04d, BB05b, BDM03, BP03, BLdPS05, BNdP09, BM01b, BT04a, BR02, Bon03b, BK09b, Cal05a, Cal07, gCZ03, CP07a, CŠ08b, CT04, CL02b, CL05b, CL06, CS02b, ĆDR01, DH03, De 07a, De 09b, DS02, Din03, – Dok06, DM02, Dra05, Dra06, Dra08a, Dra08b, DD07, Ed05, Fao01a, Fie03b, FGR07, Fri06, GGM^{+}09, GMP08, GL09, GWL09a, GWL09b, GLW09, HHY04, HL03, HL00, HZ05a, HM04, IS08, Jam07, Joc09a, Kha09, KS02, Kin05, Kly00, KL07b, LRŠ02, Lim05a, Lim05b, Lim09a, Liu03b, Lom07b, MMT04, MŠ04e, MFPPSS08, NS07a, NP07, OT01, Pat08b, PPT04, Reu06, SdSW05, SdSW07, Šem06b, Ser08, SZZ07, SZ04, Str03b, Suc09, Tan03b, Tan05b, WZ07, WL08, Wie01, YT04, YLZ06, Zha04c, Zha05a]. spaces [Zha05b, ZgC06, Zob00]. span [QZ07, Tan09b]. Sparse [CVV09, CHR^{+}03, JM09, BK06b, Bör07, Bor08, CD00b, CR02b, GZZ00, Had04, RS03a, RS06a, SS06b, TBK09, Wd09, Wer08]. Sparsity [CS00a, BPZ08]. spatial [Gri00, Gri04]. spatially [CCSS03]. Specht [FST04, Fur03, Seo00]. Special [Ano02z, Ano02-28, Ano02-27, Ano02-29, Ano03v, Ano03t, Ano03u, Ano03-44, Ano03-45, Ano04u, Ano04v, Ano04t, Ano05q, Ano07u, BFSV06, BKMZ07, BKPW08, DOS01, Dod07, FMPT04, FMN04, HKSS03, LGH01, Ano02r, Ano09r, BKR^{+}09, BHRV02, CNTV08, DL04, DP01, DJL01, GHO02, GQ09a, HT08b, LP09, PHZ06, Tak04, Val02a, dSZ00a, Ano05r, Ano05-33, Ano05-34, Ano06f, Ano06g]. Specializations [FH07b]. specific [LRŠ02, Nie04a]. specified [BMP00]. Spectra [Ano05-33, Böt06b, LŠ08a, Roj08b, Roj09b, RM09a, TE05, AC05, AB03a, BCDF08, Cao07a, CJVW07, CH06, Cig01, CDFM08a, EH03, FGM08, FM07c, FN03, FIS02, Guo03a, Hag02, HT08b, JMP09, Kly00, LŠ06a, LZ08a, LSD06, LCTS09, Mou06, PV06b, RDxC03, RS05b, Roj06a, Roj06b, Roj07c, RR07b, SS08a, Ter03, Tre03b, WX06a, ZZ09, Ste07b]. Spectral [ALdV05, AdLV06, ACS04, ADC08, BvDK08, BM06a, Cap01c, CR07, CL05a, CG07b, DD04, FT00, FYZ07, eFLW09, Fio05, FJ03, GOTR07, HLZ08, JR05, KSU03, LP01b, Mel09, PFL05, SC01, SW09, ST07, Ste06a, Tre02, vdMSR01, AB03b, AT08, AHH^{+}09, Bae07, BP05a, BES09, BBS08, BS09b, BNT05, Blo08, BKPW08, BG02, BH08a, Cap03, COvdD09, CT03b, CZ00, CL04a, Czo05, Das07, DDM08, DL03, DG07, DL05, EFN^{+}03, Fan03b, FWG08, FP03, GL08, GS03a, Gra04b, Guo03b, Guo05a, Guo05d, Guo06a, Guo06c, GS06, Guo07b, Guo08, GW09c, Ho03, Ind09, Koz09, Lan05, LR02, LS02a, LEL02, LSC09a, LR09, Lin08b, LLT04a, LLT04b, LLT07, LL08d, LLT06, Lur06a, LY09b, Mae08, MW01a, Mol05a, Mol08b, Nik06c, Nik07c, Nik08b, Nik09, ORvdD02, Omi08a, Omi08b, Omi09, PJ08, PPP06, Pep08, PR02a, PR02b]. spectral [PW08b, RBTR09, Rod05, Roj08a, RS01, SS06d, Sen07, SC07, Shi07a, Shi07b, SS07c, Shi09, SW04, SAH08, Stu00, Sun05, Tam04, wThY09, Tre06, WX06c, WX07a, WX07b, WHAB08, WHBM09a, WA09, Wir02, Wir05, WXH05, Wu06b, Wu08b, XHSZ09, XX09, YLT04, YLT05a, YLT05b, YL08, YSL08, Zen01, ZSL09, ZWS09, ZLS09, ZZ08b, Ziz04, dSQS00, vD07, vDK07]. spectrale [BR00a]. Spectrally [CV05, DHH^{+}07, DJOvdD00, LN03, MTvdD05, MS08, BOvdD07]. Spectrum [GT09, Pet02, PS03, AABS09, AH09b, BJ08, CNR00, DP01, DH02, DBK05, EOvdD03, FGR04, FNS07, Fre04, GM07b, GHK01, JDS^{+}03b, KPT01, Kra04, LŠ07b, LLT05, LzZT07, McD03, MW01b, OT07, Roj02, RS03b, SHZ05, Son02, Sot03, Sta09a, TU02, TUC04, WX06b, Yin04, Yin05, ZL02, ZZJ08, ZG09, vDH03a]. speech [LMV03]. Speeding [VHK01]. sphere [MA03]. spherical [Li06b]. spin [AD02a, AD05, De 07b, Ha02, PR04b]. Spindelböck [BST09]. spline [KH01a, RR05]. Split [HR04, Mel01, Nom05a, NT06b, NT06c, NT07c, QX08, Sta09b]. splitting [BGN08, GX09b, Hei01a, WB08]. Splittings [Ben08, CHP06, EFN^{+}03, LWWZ08, Nou08, OS00, WB01, Whi00]. spread [FWG08, GHK01, MK03, PSvdW05]. Springer [Kol05]. Springer-Verlag [Kol05]. Square [HL01, ABB00, DGN02, DDM08, – DI02, HW05a, HLW06, HS04, HWC06, Ikr01, JOR01, JO02, Kes07, LCTS09, Nov09, Tak00, XM02, XQ08, dSZ00a, dSZ00b, tBS06]. square-zero [Nov09, Tak00]. squared [AS02]. squarefree [DTGV06]. squares [AB05b, AT08, BBL03b, BH04, Bak04b, BD09b, Čer09, DW00, EF07, Eld02a, Eld06a, FL04, FLPW07, Fit01, FB04, Gro06, GA06b, GJW02, GWW06, HW05a, Hil07a, JB00, LMV03, LD02, LCN04, MLK^{+}04, NPP00, Nie01a, OCK07, PJ08, PA04, VHK03, WG02b, XQW08, Yer04, ZZ00, ZM08, ZHZ08, Sch06]. SR [SAAF08]. Stability [AB03b, BKS01, Che02b, GV06a, LCP02, NS02, OP05, Sty02b, Vir08, VEB00, BOP08, BKO01, BS00d, BLO02, CJ05, CP09, CGPM01, CP07b, DB06, FL04, GL01, GŞV02, Had08, HK03, JH09, JLN08, KB02a, KP02, Log05, Lur05, MS06, MGWN06, PV06a, PK08, RA08a, RO02, SvdD05, SMK09, SN09, TSOA07, Wan06]. stabilizability [Sun02]. stabilization [FS04b, HZM02, KKP08, KB02a, WLL07]. Stabilizing [OTC03, GRV06, HL09]. Stable [GL00, Gol05, Sud04, BBD05, BKS03, CG00b, FL04, FLPW07, FGL^{+}00, Gem03, HOP06, Kir05, LMZ03, LN03, Peñ04, PT05, SS07a, VB09a, Ves03]. stacking [TKX08]. stage [BW03, Wan07a]. staged [HL05b]. stagnation [ZOE03]. staircase [Dah08, PHO07]. Standard [HLG04, Hil08b, CBD^{+}04, FH00, HC03a, JL09, Lu02a]. standardizable [GPMT06]. Star [CRS07a, RJ02, AGM09, BBL03a, BH04, BS04a, CT01, GL01, MTvdD05, Row07, Sta07, Sun08]. star-centers [CT01]. Starlike [OT07, GRTR08]. stars [JDS03a, SSB09]. State [Dam02, GLS07b, HO02, Lom07a, MZK00, And07a, BGKR05, Bas02, BFPW02, CG00a, DKS04, DS09, FS02, FRY07, KN00b, KS06b, PHO07, RZ01, RZ05, XM02]. state-delay [FS02]. State-feedback [Dam02]. state-sharing [DKS04]. state-space [Bas02, DKS04, PHO07, XM02]. states [Far04, MT06, Pit03, Ter01a]. static [ALCT05, EG02, HZM02]. Stationary [Hun05, ABDD01, AAV06, CM01, HW00, Kir06, KS09, MM03, Pic07, VM04]. statistic [Dun04]. Statistical [AAA^{+}02, AG05b, Rao00, AdO02, BdL04, Fur06a, Jia00a, MP05, Smi01b, Ste05b, dMR07b]. Statistics [AHST09, Ano09q, EMS06, BB04d, CM09, FKP00, Gus02, Har00b, MPS06, Sed09]. steering [Cen01]. Stein [And04c, GP00a, KS01, KS05, LEL02, Nou08, SS07d, SS07e, SG06]. Stein-type [SG06]. Steinberg [LH05]. step [CP06, LA05, Mel01, RM01]. step-down [CP06]. stepsize [HZ03]. Stickelberger [Zen08]. Stieltjes [AT00, HC04, Van04]. Stirling [CK01, CK02, SZ08]. stochastic [ABH03, AFL00, AH04, BBFP04, BLM03, BD03c, CL05a, CS04, Dah04, DH01, Fan10, FPP02, FKS02, FLQ04, FB05, FH04, GAM06, HP01a, HP04c, HL04e, JMT05, Kir00, LTT02, LT04b, MM03, MM06b, Mou06, Nie04b, PV06b, PA04, Sat08, VHK01, XLG^{+}13, Zha05d, Zha09]. stock [Gue08]. Stokes [Sty06]. stop [HP01a]. stories [Str01]. Strang [Blo08]. strategies [CP03, DLZ03, MQ06]. strategy [McC05]. stratification [Fin00]. street [CDFM08a]. Strengthened [TW05]. Strengthening [BGS00]. Strict [MG09, EFN^{+}03, GJR03, GSR03, LR05b]. strictly [CW00a, Cao01, CCH02a, CT03a, CH07, CP08, Fur02, OWY07, WY04, WL09a, Wan09c, Wir05]. Strong [GP03, LT04b, LL08c, MG03, RA08a, TY06, Yan09a, Bak04b, BLL03, FLQ04, Ma09, SH00a, SKBS08, Yan02a]. Stronger [Log05]. Stronger-than-Lyapunov [Log05]. Strongly [Hae08, Ovs03, Gav03, Jør05, KMMZ04, PJ07, Row07, WD06, YZ07]. structurable [Oku05]. Structural [ABW05, BCS02, CP07b, FG02, FVP06, HH00]. Structure [pBlWyXwA08, – Dok06, Ish04, LR03, MR09, Sud02, APESS06, AD02a, AMZ04, BCES05, Bar08, BS00b, BdHZ05, BOP08, Bol05, CG00a, CK08, CKM05, CFL05b, CV07, DRGR04, EFN09, EFN10, FPP02, FKR09, FJL08, Gav02, GR07b, Gun03, Hei02a, JDS08, MD02, NS07b, NT08b, OT09, PSvdW05, RM09b, SdSW05, Tar05, Wat05, YLZ09, Yek04, dSS08]. structure-preserving [BOP08, CFL05b]. Structured [Ano03-45, Ano06g, BBdH09, CFP03a, DOS01, FMN04, HL09, PRW01, PEPS00, Sch02a, VWD08, BE01, BS04b, BD04, BMX08, CNTV08, CSK01, DPZ06, EG01, EG03, FMM01, FB04, Gem08, LMV03, LK09, MLK^{+}04, MMX04, Mei06, Mel09, PW08a, Peñ03b, Rum06, SS09b, SC07, Yer04, vdW02]. structures [BAL09, BBC05, BP09, BDV08, CFN04, CT09, GR07a, Kye03, RG03, Sal09, VC04, YH04b]. Student [Okt04]. study [ACHR01, ACDR04, Fuh02, Gre02, KKP07, PRV05, TKX08]. Sturm [CWW09, Kos00, LS09c, SC01, SW09]. style [JSWT05]. sub [DMM00a]. sub-Markov [DMM00a]. subadditivity [BU07, HZ04a]. subalgebras [BN08, Bir04, Dvo03, GM09b, LZ08b, OPS07, Ohn08, WY06a, ZJC04, ZYP05, ZFW06, ZYP06]. subclass [Sch04]. subcollection [Wal05a]. subconstituent [Ter02]. Subconstituents [gLxW02]. subdefinite [MND01a]. subdeterminants [SXKM03]. Subdirect [HMT^{+}08]. subdivision [CGR09]. subdominant [Pry09a]. subeigenvectors [SSB09]. subfamilies [CL01b]. subfamily [Fon00]. subfield [Yah04, Yah06]. subgraph [BM04b]. subgraphs [BPS09, Nik07d, SS08a]. subgroups [CS05, Hoo07, KKABAF06, VLAOVL08]. subject [Amb07]. sublattices [QT06]. submatrices [Bap03a, BL05, BC04b, DV07a, DV08, HL04e, JK03, JKO04, JO05, LŠ08a, Mol05a, Mol08b, NS02, Sav03, Sav04, Sav06b, dSZ00b]. submatrix [Bor06, BC07, BC08b, GV09b, Sav05, Vis00, ZHZ08, dDG02]. submultiplicative [Kra04]. Submultiplicativity [HZ04a]. subnormal [CLY08]. subnorms [GL00]. Suborbits [GW09b, WGL09]. Suborthogonality [Fie09]. subpermanents [CE09]. subradius [Czo05]. Subrepresentations [Han05]. Subresultants [Kah04]. subrings [EMBV03]. Subset [dHM07]. subsets [AR06a, DR08, MFPPSS08]. subshifts [MSR06]. subsimplices [BKK07]. Subspace [Dug04b, RT01a, AEEG08, AP04, CRS03, CW02, CGP04, DNR04, Dek09, DDHL08, FS07, HLY06, HQ08, JR09, JS00a, LLM03, LM03, LL08b, LL03, MZ04, Ovt06c, PJ07, Ple07, RS02, SL06, SP04, SM09b, WL07, Wu08d, Xia06]. subspaces [AEK06, ABW05, BT09b, BDM03, BR02, Bra03, CFKP08, CFP04, CHPP06, CP07b, De 09c, DM02, DH04, FYL07, Fao01a, FP01, FGR07, FH01, Gao03, GX09a, GQ09a, GV06a, GWL09a, GWL09b, GLW09, GW09b, Ips00, Kha09, KN03a, KA06, KP02, LL02, LS00a, MRR07, Ovt06a, PPZ02, PP02, PPZ05, Rvz09, RS02, Ste02b, Tsa01, WGL09, Wat06, WLBZ06, Ye08, ZQ07]. substitution [TI08]. substochastic [LT04b]. subsystems [CJPS02]. Subtotally [Fie06]. subtractivity [Leg04]. subunitary [Li08c]. succession [FP02b]. such [Cai03, DK04]. Sudoku [Dah09]. sufficiency [SH05]. Sufficient [WB01, ACV09, IPP07, KV06, MPS07, TY09, VV06]. sum [AGKM09, AT08, BMNA08, Buc06, CD00a, CXW09, Den09, JO05, KRS04, KRS03, LR09, LP09, Mol09, PJ08, THZ09, WZCL08, YC01, Zho08, dSQS00]. summability [MR02b]. summation [Bur06, CL04b]. Sums [Bot00, JO04, Nov09, ACS06, Bon03a, BD03d, DG07, eFLW09, Fit01, Gil08, HW05a, Hil07a, Hir05, HK07, HV08, HMT^{+}08, JS00b, JLY01, JLMC05, Kit04, Kur04, MW03b, O’L03, Tak00, dSF09]. superalgebra [Tri02]. superalgebras [BTT09, WM02, WM03]. Superfast [PZ00, Gem03, VHK03]. superinvolution [BTT09]. superiority [LA05]. superlinear [Cap01b]. supermatrix [Dup03]. Supernormal [vGKN09]. Superoptimal [ESC08, CJVW07, VWJ08, vdMRS06]. superregular [HST08]. Superstochastic [GS09d]. supplementary [KMS05b, Zha03a]. support [Ost06]. supported [AH09b]. Surface [DS04b]. surfaces [Lin08b]. surjections [gCZ03, lHH08]. surjectivity [YW07, Zha04b]. survey [BORŠ04, COvdD09, CW05, Cve02, FH07a, FI08a, Fre02, GN00, JW08, KR04, LXW04, Olk02]. Survival [vDP08]. SVD [Boj03, BSS09, GL03, Gug08, Hig00, Sla03, Ste06c, Xu03, XQ08]. SVD-approach [BSS09]. SVD-like [Xu03]. SVDs [DK06]. switching [NT08c, RJ02, Ter03]. Sylow [VLAOVL08]. Sylvester [Hig08, BB09, BR07, Dam03, JF01, KMP00, SA02, WSL02, Gem03]. symbol [Abd05]. Symbolic [DKS04, Dru08a]. symbols [ESC08]. Symmetric [Boc03b, DV09, HLW01, MR02a, Mel00, Oku05, Pac02, SZ08, Tag04, TM00, ALSS08, AHZ05, ATU05b, AG05a, AM05, BMNA08, BJT04, BLS05, BCEG09, Ben08, BB05c, Bér05, BMGS06, Big02, BCPT05, Bou06a, CW03a, gCZ03, CS09a, CHNP04, CP07a, CG00b, CT04, CMX05, CL08a, CHMM06, CGR09, Dan06, Das09, DHH^{+}06, DM03, DP04, – DZ04, DH09, FH07a, Fan02, Fan10, GGL08, GG09, GZZ00, HR01, HV01, Hil08b, Hog02, Hol05, HZ05a, HBLZ06, HP04c, Jia00b, Jia05, Kle04, Kly00, KT08, KMS05c, KO06, KLY09, LŠ07b, LR05b, LL06a, LPR04, pLLfY06, Lim05b, Lim07b, LT04b, LHF06, LM04b, MVV08, MH06, MHZ05, MN08, MQ06, Mon08, Mou06, NB03, O’L03, pPyHZ05, PD00, Peñ09, PV06b, Phi04, PVV08, PT03, PT06b, Rod07, RR04c, RS09, SSG04, SG08, SZZ07, Sid01, SP04, Sla03]. symmetric [Smi01b, Smo08, Sot06, Ste06a, Str03b, Stu00, SG06, TBK09, Tre04c, Tre04e, Tre06, Tre09a, Van07, Ves00, WV08, WSL02, Wat06, XHpS06, XQ08, XLG^{+}13, Yam01, jYxL09, Yin05, YT04, ZH02, ZHZ03, tBSR04]. symmetric/skew [LR05b]. symmetric/skew-symmetric [LR05b]. symmetrically [Con08, Dah05]. symmetrically-normed [Con08]. symmetries [ABR02, Tre05, Tre08, Uhl01]. symmetrizable [JSU03, MS04d]. symmetrization [Dob04]. symmetrized [Bon03a]. Symmetry [Han02a, Val09, tBS06, BPR03, EJLS02, FJTU00, HK03, JR00, Pou01, Tre04a, Tre04d]. Symplectic [HZ03, Hoo07, CBD^{+}04, De 07a, De 09b, De 09c, DJ06, DHZ03b, EL01, FI05, GX09a, GWL09b, GN09, HZ06, Knu00, SAAF08, SS03c, Spi00, Wat04b, Yan06, Yan07]. synchronised [BSL04]. synthesis [BPT07]. System [KM04, AT00, AB05b, AMZ04, AMZ06, Bai08, BMP00, CHP08, CM08a, Gue08, HL03, Igl01, KKK07, LZ09c, Lom09, Qui07, RM09b, RW01, SV07, Shi04, Ste01a, Wan04, WL09b, WvdWC09, XM02, ZF07, vdW01]. systematic [FZ04]. Systems [Ano02-29, DOS01, TW00, AAA^{+}02, AL09b, AD02a, AAV06, AMZ06, AMZ07, Bad08, BW03, BS06a, BGN08, BRR01, Bal03, BZ02, BS00b, BPT07, BFSV06, Bér05, BKS01, BHRV02, Bou09, BK01, BRS00, BRVS08, Brü09b, BGGC06, CLR00, ÇS02a, CW02, CP03, CHASG02, CDMGP06, Cec00, CCG02, CS08a, Che02b, CCZ06, CHP07, CM08a, CQ08, DJMS09, Dam02, DMW08, DR05, DvdV00, DS08b, Dod07, DS08c, DS09, DK08b, DHZ03b, Elf04, Est05, EGMG07, Far04, FLMR00, FH00, FL04, FLPW07, FPP02, FN07, FS04b, Fie03b, FH06, FVP06, Fre08, FS02, GRV06, GPM01, GPMT06, GPMSZ06, GPFC04, Gem03, GG09, GJR03, GA06b, Gug08, HOP06, HvS07, HCLC09, Har06, HV01, HZ03, HZ06, HMH01, HZM02, Igl01, IO01, JP04, Jef05, JM09, KKP07, KRS09, KB02a]. systems [KS00a, KS06a, Koš03, KPP07, LMZ03, Lan05, LT04a, LCF06, LV06, LWWZ08, LWD09, Lom07a, LCP02, MH07, Map07, MM07, MS06, Mas09, Mel01, MN08, MGWN06, Mue09, Myš05, Myš06, Ned01a, NB03, OTC03, OS05, Ots01, PV06a, PHO07, Pen06, PRV05, PCVT04, PPF02, Rad03a, RSW05, RR05, RW06, RO02, RS01, SSSG08, SL06, Sch09a, Sha03, SS09c, Sha00, SSY05, Shi06, Sid01, Sun02, Tab01, TG02, TM00, Tan09a, Tre09a, Tre09d, Tru07, Val00, VZ02, Vir08, VWJ08, WSL02, WM05, Wan07b, Wan07a, WB08, WW08b, WSW06, XWD06, YWS00, ZWPR06, ZTK09, vdMSR03]. Szeged [KYAA08]. Szego [BEG^{+}07, BW06, BGM08, Sak01, Tre02].

T. [CJ04, CJW05, CJ05, JW08]. tableaux [AM05]. tail [CCW^{+}08]. Takagi [KR07]. Tam [CL00a]. Tame [Drä03, Zav03, BS03b]. tangent [CS02b, HS09]. Tangential [BN05, HOP06]. Taylor [Sak01]. TCP [BLLS06]. teach [GC04]. teaching [Uhl03]. technique [FM04b, HGC07, SSW05, Sta07]. Techniques [Tům06, ACHR01, BBP03, Che05, ES02, GLP00, Sed09]. Temporal [HW00]. Tensor [BCES05, LS07c, OOT06, And04a, Bon03a, CSK01, DLMO07, DM07, Hou03, OTZ09, WZCL08, Zus05]. tensor-based [DM07]. tensors [Ano03-31, BPR03, CGG02, Dia02, DRGR04, LZ00, Lim08a, Pat08b, Pou01, SMS09, dCdS05, dSdC05, dSR06]. term [Bal03, BS02c, BBG09, EGO05a, EK08a, dSR05]. terminal [BF07b, Sci02, Sta09a]. terms [IPP07, KC03, Ste03a]. Ternary [BP06a, JGPI08]. Terwilliger [Mik09]. test [BGM08, CM09, GA06a, Kap00]. testing [GA06a, JK06b, WX06a]. tests [CP06, IL06, Peñ04, RO02]. tetrahedron [FN09, Har07b, IT07a]. th [Che07a, Cho04, De 09c, Di 03, Guo06b, Roj07b, ZWY00, Hag07]. their [ABW05, Ano03-44, BBL03b, BHLL04, BSKS05, BLG01, CH03, CNSM08, CNF08, CFRS02, DJ06, FL04, FF04b, FNT07, Fur06b, GH00a, GPS01, Gin09, Gol03, GH09, GT09, KMS05b, KS06e, Lad08, LG06, Lom02, OT07, Reu06, Rod08, Ruk01, Sat05, Sav06b, Ser07, SU09, SHZ05, SS09d, SKBS08, Sta09a, Ste05b, WX06a, WlW07, Wil09, XHpS06, YFK05, YLT05b, ZZ09, dSS08, tBS06, vDH03a]. them [AEB06, CGPM01]. theme [BBCH08, HR03b, Hof00, Ste05a]. Theorem [Zaï04, BBM02, BR00b, BdPK07, BV02, BHdST06, Ber09, Bıy03, BT07b, CM00a, Cao05, CGM05, Cao07b, CL09a, Cio06, CS04, Cos07, Dan06, DS02, Dug06, ES01, FP03, FNS07, Fri00, GK07, GJ04b, Ges02, Gra04a, Har07a, HDB00, dH07, Kes01, KSH08, KPV04, Kol03a, Koš03, Leš04, LS00a, Li00a, LEL02, LR03, LLY09, LM09b, LR04, Lon04, Lur06a, Mat03, MOSZ02, MG09, NV03, OTC03, PLG02, Pep08, RRY04, SN09, SR06, Ste01b, Ste04, TY05, TY07, TY09, TRH01, Tre04f, TL03, TZ03, Yua00, Zen08, dMR07a, CLL05, CKS06, FH07b, HBHS05, Min01, OO05, Par09a]. theorems [AD03a, AW02, AAV06, AI00, BW03, BGM04, BGM08, wCL04, CGQ04, DGLS01, FS04a, FS06a, GP00a, GTM09, Li00b, LN06, MS00a, Nou08, SSA05, Šem03, Sty02b, Zha07b]. theoretic [Hog01, KM04, Pic07, VZ02, Zar06]. Theoretical [CL01a, CS03a, SH09]. théorie [Fao01b]. theorists [Kri05]. Theory [Ano02s, Ano03-43, BLL06, HKSS03, Moh05, Par09b, Uhl00, Uhl04, vdD09, AL09a, AT00, Bad08, BHNR03, BNdP09, Bou09, CD00a, Cal05a, Cal05b, Che02b, CHP07, DDM08, DT03, DG02, FH01, Fuh08, Fuj05, Gol05, GN05, GQ09b, GV06b, HLT04, HM02, HP00, HT02, HP03b, IBI09, JR05, KKP07, KN03b, Li00b, MM03, MMX04, MNV08, MZK00, Nak03, Ney01a, Ney01b, Ney06, PWZ05, PT04, RR02, RR04a, RT01b, Rum03b, SC01, Shi06, SW09, SP09, Sil00, ST00, ST03b, Sty02a, Ves00, WWQ04, WLBZ06, XDS08, Xue01, YFK05, Zen01, dSG07, vB09b, BKPW08, Hig08]. Therapy [Ano06f, CBIXG08, CGXL08, HC08, SK08b, TVB08]. there [Kah06]. thick [Wu06a]. thin [Bra06, HB08, Ter02]. Third [NT02, KS06a, LB00]. Thirring [Aud09a]. Thomas [Ano03u, Mar05, Ano09o]. Thompson [BS07, Seo00]. thoughts [BBM02]. Three [BB04d, BB05b, BS05b, BD01b, BD06b, EGO05a, Har05, vdH08b, Bag06, Bak04c, BB07b, Bar08, BLG01, CC04, CO09, EK08a, GZHT05, LDS08, NDS02, PBT04, Rab04, Sbu05, Tak00, TBK09, smbH02, dSR05, tBSR04, tBS06]. Three-connected [vdH08b]. three-dimensional [BLG01, GZHT05]. Three-term [EGO05a]. three-way [TBK09, tBSR04, tBS06]. threshold [CR02a]. Tight [De 07b, HKKZ03, RM06, BPT09, HKKL05, Pas01, Ren07, Roj09a, Str08, STDH07]. tiled [HP02]. tilting [LG04]. time [AAA^{+}02, AFL00, BF07b, BDKW04, BRS00, BCS02, Brü09b, CFL05b, CM08a, DL04, Dam02, FPP02, FN07, FS04b, HOP06, HW00, HZM02, IV04, Igl01, IO01, Kar04, KB02a, KS06a, KMZ00, Kle00, KS01, KS06c, KS09, KTKR06, LCF06, MA03, MGWN06, OCK09, OTC03, Ots01, Pfa08, PPF02, SV07, Sun02, Tru07, Val02a, Xu06, Xu07, YWS00]. time-delay [MGWN06]. time-invariant [BCS02, CM08a, HZM02, LCF06, PPF02]. time-varying [Dam02, Igl01, Sun02, Tru07]. times [CZ08, CM00c, Hun05, Hun06, Hun08, Hun09]. Titchmarsh [Shi06]. Toda [Gla02]. Toeplitz [BB05c, Bér05, BGM01, BB03c, BDG08, BES01, BG02, BES03, Cap01a, Cap01b, Cap03, CI09a, Dan02, Est05, FLM02, GKvS04, GGCB07, Hei01a, HR01, Hei01b, Hei02a, Hei02b, HR03a, HR04, HV01, HH05, Kar05a, Kle07, LZ09c, LCN03, Lin08a, MD02, MVV08, Mel01, NRW02, NCV05, OOT06, Sak01, Sch09c, Str01, Til03, Tre04b, Tre06, TZ01, VHK03, VWJ08, WD05, Wim01, Woe02, vdMSR01, vdMSR03, vdMRS06]. Toeplitz-plus-Hankel [Hei01a, Hei02a, HR03a]. tomographic [SvdS01]. tomography [AG05b, Cen01, HT01, Yag01, LGH01]. Tools [SCS09, PRV05]. topics [AD08, Rod09]. Töplitz [SS04a]. Topological [Cao07b, ATV09, GN05, GJK04]. toric [RVZ00]. Total [Bou03b, Bou03c, Gla02, Sch06, AB05b, FB04, GG04, Hol03, JB00, KV06, KS06e, LMV03, LD02, MLK^{+}04, NPP00, Nie01a, Peñ01, Peñ04, SS06a, XQW08, Yer04]. Totally [ATU06, DR05, KLM01, BF08, CRU09, CFJ01, Doe07, FN01, FJ07, FM00a, FM00b, FM02, Gar03, GJ04a, Gla04, GWL09b, GLW09, GW09b, HMT^{+}08, JO04, JT04b, JTeG09, LM02, Ost03, SS04b, Tre06]. tournament [E^{+}00, HWKS03, Kir03a]. tournaments [BBG09, Ren07]. Trace [ÁFMNW06, BES09, Fur06a, FKY09, HK05, Bek04, BT07a, CZ00, CS09b, – DS08, Fuj05, Fur05a, Han02b, Hil07b, KN04, Kom08, Lim09a, NT06c, RSW02, YFK05]. Trace-minimal [ÁFMNW06]. trace-preserving [RSW02]. traces [Bha03b, CXZL02, – DJ07, Dru04b, HW07b, MV04, Yah04]. trails [BPS07]. Train [Zit09, LR05a]. trajectories [AL09b, SV07, Val02a]. trajectory [RW06]. Transfer [Myu09, FN00]. transform [AD03a, AW02, HT09b, HV01, JL03, Kah06, NCCY00, SvdS01, Wu02, HT08a]. Transformation [ES02, Mai01, BM04e, DL04, INOY03, KKP08, Oku03, Oku06, PT03, Yam02, ZS05, dMR07b]. transformational [JS01b]. Transformations [BM01b, BBC05, MT06, Xu07, BGLS04, BQS00, BRZ04, CHJP05, DL03, EV05, FNT05, God09, GST04, IT07a, Lad08, Leš00, LK09, LRŠ02, Lon04, MMT04, NT07b, Nov09, SAAF08, Sal09, SG06, TBK09, Ter01b, Tov00, VV06, Wan03a, YH00a, YH00b, ZG09, smbH02, tBS06]. transformed [TP09b]. transforming [LS04b, LPS05]. transforms [AY03, AMS05, BD03b, LCN03, ND05, PEPS00, PT05, Pot03, Tsa00]. transient [PW08b]. Transition [KNX07, NT09a, CDS08, KN00a, Sav05]. transitive [Dah05, DDI03, DLM^{+}00, KO08, LL02, Sav05, Tan05a]. Transitivity [LM01]. Translations [DDI03]. transport [GG09, TVB08]. Transportation [Dah08]. transpose [HS04, Ver08]. transvections [EL01, ZY02]. transversal [CFKP08]. trapezoidal [Gav03]. treatment [CCW^{+}08, HC04]. treatments [MCEC06]. Tree [SS09a, AH09b, Bap05, Bap06, BLP06, BR09, BFH05, CHMM06, DHH^{+}06, GT02b, Guo03b, GT04, HLR08, JD02, JDS03a, LL08d, PL08, Roj02, RP03, Roj06b, RR07a, Roj07c, Roj09b, Toy09, WX06b, Wat00, Zha09, BLM03]. tree-like [BFH05, BLM03]. Trees [FGS05, ABH03, BHL09, Bıy03, BL09c, BH08a, CH03, Che07a, CLOvdD01, Das04b, FGWG07, FYL06, FGM08, GRTR08, Guo05b, Guo06c, Guo06b, GS06, HT08b, HL02, II09, Ili09b, LG06, MW01a, MN03, Mol05a, OT07, Omi09, RT06, Roj05, RS05b, Roj06a, Roj06b, RM06, RR07a, Roj07b, RR07b, Roj08b, RM09a, ySZyY06, SGS08, Ste03a, SG09, wThY09, WX06c, Wan09e, WK09, WXH05, YLT05b, YL06, YL08, Zha07c, ZLC09, ZWY00]. Trefethen [Böt06b]. tri [BW01]. tri-diagonal [BW01]. trial [KA06]. triality [Eld00, Oku05]. triangle [Dra05, FP02a, PJ07]. triangular [AH09a, BW00, BBC00, BS00c, Ben05, Ben09, BPP08, BM05, CW00a, Cao01, CCH02a, CT03a, CGM05, CC04, CL02b, CLP03, CL03b, CL06, CHL01, DW09b, DP00, Fer05, FMM05, Foš05, GR00, Hao04, HPP05, HC06, JW06, Kar05b, Kob05, LSD06, Lin08a, LT07c, Meh00, NDS02, OWY07, PP01a, Pei07, Pet02, TY06, Tia00, Urb04, WY04, WY06b, WY08, WL09a, XCX01, ZY06, ZXZ08]. triangularizability [KR05]. triangularization [– DS08, Dub09, Flo09, MR09, Yah08]. tribute [Sch00b]. tricyclic [GL08, GLL09, GW09c, LL09a]. Tridiagonal [Dah04, D– D05, – DZ04, FN09, IT07b, Nom05b, Nom05c, NT09b, Vid08, AC04a, AC08, BKR06, BD05, CN01, DM03, DP04, EGG07, EOvdD03, Fer07b, Gra02, GT09, Jia05, KH09, KY00, LST08, LZ09c, Mal01b, Mol08a, Nom05a, NT07b, NT07c, NT08a, NT08b, NT08d, PP01b, Ter01b, Wan01, dFP01]. tridiagonalization [HS07a]. tridiagonals [PD00]. trigonometric [Mol04, Sak04]. trigonometry [Gus00, Gus02, Gus09]. Trilinear [IS09]. Trimmed [BMX08]. tripartite [Jef05]. triple [DKL^{+}07, GT06c, JW05, LS06b, LWD09, Lu03, Mol06b, WFJ08, WL09a, ZWC06]. triples [BLS05, Cur07, GJ04b, GS00, LL08a, NT01, Oml04, Šiv08]. tripotent [BBS02a, BBÖ04]. trivial [Val08]. tropical [GK09]. Trotter [AKL07, Fur05a]. truncated [CHKS05, JKLP03]. truncation [CLVV06, GVV03, Sty06]. Tsallis [FYK05, Fur06b, YKF05]. Tsetlin [Bap03b]. Tucker [TBK09]. Tucker- [TBK09]. TUHF [JW05]. Tuples [CCH02b, BG09a, OSvdD02]. Turing [SvdD05]. Twisted [HN09, Pat08b, SH08]. twistor [Gus09]. Two [AW02, ABR02, BLO02, CP06, CT03c, De 08, DV01, FM04a, Fur06b, GGM^{+}09, HS03, HP03b, KP06, Nak00, NY02, NP07, PHZ06, She03, Ter01b, VW00, YY03, Zha04f, ACM03, ACHR01, AGI00, BMNA08, BW03, BB00, Bak04c, BB07b, BHJ08, BT09c, BBS06, BKS01, BW05, BLG01, BKN01, CD00a, CL04b, CK04, Dax06, Dra06, EGO05a, FL03, Fur08a, Fur05c, GA06a, GS05, GS09a, GS09b, GW04b, Gra05, HK02a, Har00b, HI06, HLZ09, Hua00b, HL08a, II09, JH09, Jia09, Kol08, Kye03, LŠ09a, Lin07, Lin02b, Lip05, LSF09, LC09, LMT09, Mel01, Mol04, NCV05, PFL05, PP04b, PEPS00, PP01b, PBT04, Ros05, Sak01, SSG04, SV01, Tan09a, Tup09, Wan07a, WG02b, Wu06b, Wu08b, Zha04d, vDS05, vdH03b]. two-colored [GS05, GS09a, HL08a]. two-cyclic [LMT09]. Two-dimensional [GGM^{+}09, Nak00, She03, VW00, YY03, AGI00, BKS01, BLG01, Kye03]. two-graph [HK02a]. two-level [NCV05]. two-parameter [JH09, PEPS00]. Two-sided [ABR02, HS03, NP07, Kol08, PP01b]. two-stage [BW03, Wan07a]. two-step [Mel01]. two-term [EGO05a]. Two-view [DV01]. type [AC08, And04b, BC08a, BX08, Bap03a, BEG^{+}07, BC03, BM04d, BD03b, BS07, CLR00, Cao07b, CS03b, CW07a, CW07b, CKW08, CGR09, Cos07, Dam02, DDI03, DLZ03, DMM00b, Dra08b, FI08b, FNS07, FNV08, Fre08, FST04, FG03, GII08, HFR04, Hoc09, HLY06, HT05a, HZ05b, IT07b, IK09, Jai07, JLN08, JPB08, KSH08, Kir06, LM06a, LS00b, LY09a, MS01a, MPT02, MSR06, MPP06, Mik09, MG09, Nie07b, Nou08, yPyHZ03, PWZ05, Seo00, Sid01, Smi08b, Str03b, Sud04, SG06, Van07, WWJ09, WGL09, WlW07, Zha07b, Zhu08, CTW01, VPV01]. types [BD09a, Nie04a, Zhu08]. Typical [tBSR04, CtBDC09, tBS06]. typically [BW05].

U [ZZ05]. UB [Mir03]. UB-matrices [Mir03]. Ullemar [Tka05]. ultimate [De 00]. ultrametric [DMM00a]. unbiased [Bak04a, PR04b]. unbounded [BLdP04, ESC08, Mal09]. uncertain [AAA^{+}02, BES03, CCZ06, Gha01, KFGK04, Sun02]. uncertainty [Dam02, GII08, PA04]. Unconditionally [VB09a]. uncoupled [DHFS00, MSF05, Xue01]. underdetermined [JM09]. Unextendible [Pit03]. Unicity [BW01]. Unicyclic [HGW02, HW07a, BFS08, BMS06, CT03b, Dau08, FWG08, Guo05d, GYY09, Li05, LSC09b, LLT07, LZS09, NS07a, Roj08a, SI09, XL05]. unification [Bot06a, Bot09]. unified [Bak04b, BMS08a, FLPW07, FN07, FM04b, HC04, PRW01, Uhl03, zY08]. unified-least-squares [Bak04b]. Uniform [Str03a, Cao07b, FNS07, Kob00]. Uniformly [MS04d, vB09b]. Unifying [BQS00, Ips03, OT06]. unilevel [Tre09c]. unimodular [AK07, AY08, Doe07, Fan07b, Ost03, Zha06]. Unimodularity [AA07]. union [BF05, CL02c]. unions [Kha09]. unique [AM09, AHL^{+}09, HLW01, HB00]. uniquely [Bru02b, GSR03, Wil09]. Uniqueness [JOR01, JO02, LM04a, AEB06, Alf03, Alf05, Brü09b, ÇS02a, SS07f, Ste09, TW05]. unistochastic [Co08]. unit [BB02b, Bol02, CMV03, CMV05, Dru08a, FLROV05, MQ03, ZX07]. unital [Kye03]. unitaries [Dru07, O’u02]. Unitarily [And06, – DJ07, CLS05, HZ02, HZ04a, Li08c, Oku03, Oku06, Poo08, RT01b]. Unitary [FJ05, GI02, JMN08, QC04, AADL03, ADL04, ADL06, BEG^{+}07, BBC05, BQS00, CMV05, DW08, FJ02, Gao03, Ili09a, LL08a, LS04b, LP09, LH05, NBdP08b, NS08a, Pet05, PR02b, Rod06, Ste05a, Tad06, TŻ08, – DS08]. unitoidness [Rob06]. Unitriangular [Pre05, VLA03, VLA04]. units [NT06b]. unity [AFM03, BPT09]. univariate [PRP02]. Universal [Til03]. universe [PPZ09]. University [Böt06b, Rod09, Tům06, Uhl04, dSQS00]. unknown [SvR02]. Unoriented [TFZ08]. unreplicated [MCEC06]. unsolvable [Cec00]. Unspecified [Hog07]. unusual [FI08b, LY09a]. unwrapping [KB02a]. update [CW05]. updating [Lan08]. Upper [GB06, Gre09, MK05a, SS07c, Tia02, BW00, BFH^{+}08, BHS06, CW00a, Cao01, CT03a, CGM05, CH07, CLP03, CHL01, cD03, Das04a, DB05, Das07, Das09, Fer05, Foš05, FF04a, FF04c, Guo05a, HOZ01, JW06, Kir03a, Kob05, LSD06, LB04, LT07c, LLT06, OWY07, Pan02, RSR00, RR04b, RR07a, Roj07a, Roj08a, Roj09a, SM09a, SHWR02, SW04, Tag04, Urb04, WY06b, WY08, WL09a, Wan09c, Wu08b, YLT05a, Zha04f, ZXZ08, dSS08]. upper-triangular [WY08]. ups [SV06b]. Urn [BB04d]. USAOR [Li03]. Use [CFP09, SvdE03]. useful [Gre02]. Using [But06, Kro01, AFS09, AHW05, ACHR01, ACDR04, AS02, BZ06, BMP00, CCW^{+}08, Dri04, Eld06a, EvdD08, HW07b, LV06, MPS06, Map07, MV04, OTC03, PJ08, SL06, SLP01, SK08a, Sch09a, SSA05, SvdS01, SH09, Tia02, Vin00, Yer04, smbH02]. Uzawa [BW08a, ZBY09].

V. [Gra04a]. vague [Ned01b, Ned01a]. validating [Rum01b]. valuations [FK07]. value [AC09, Aud07, BX08, Bos07, BG02, Bra06, Cap01c, CY00, CD00d, DM07, Fer05, Hoc09, JY07a, Kir03a, Kit09, MM06a, Mor08, PW04, SS03a, ST03b, Tao06, XDS09, sV06a]. Valued [Rod09, AT07, AD08, BB02b, BES02, BES09, BHdST06, Bol02, CH00a, EMERP06, EKS01, ES01, FKL03, Gue08, Las06, McC01, Smi01c, Wu05c]. Values [KMS05b, ASX08, AGO09, Aud09a, BK09a, BR07, Bou03a, Bou05, Bur06, But06, CH08, Fer03, Fer07b, FJ03, Gil04, Gil08, Gre02, HB08, JSWT05, KA06, KMS01, KLMS03, LHL08, MK05a, Mir00, Nie08, Ois01, OS00, PS08, RST01, Rad03b, Sbu05, She01, Ste06c, VV06, Xia06, vDS05]. Vandermonde [BEG^{+}07, DK06, EFS01, KS01, KPV04, Li08a, MM07, OP00, RS00, YC00a, jYzWbZ01, YH04b, YLZ09]. Vandermonde-like [YH04b, YLZ09]. vanish [Gra04b]. vanishing [Ash05, BG09b]. Varga [Kol05, Ano08r, BBCH08]. variable [Brü09b, CLY08, HZ03, SK08b]. variables [Han02b, Lom07a, MW08b, McC01, Pop02, RvdGM05, Sak04, SV01, Wal05a]. variance [AdO02, Bén06, MD02, PRP02]. Variances [Hun08, HB00]. variant [Kes01, MPP06]. variate [Mat05a, MP05, Mat07]. Variation [BD02b, FJL08, HR03b, FJ03, LR02, PFL05, Son02, Wu06b, Wu08b]. variational [CD00c, RT01b]. Variations [Ste05a, BP05a, BBCH08, Bul08, Fan03b, Hof00, dW03]. varieties [Ano03-31, CGG02, Cof03, GS00, Lip08, McG06, RVZ00, Šiv08, VJ06]. variety [GH00b, Oml04, CCTT00]. various [BEH^{+}06]. Varshamov [BGS00]. varying [CCSS03, Dam02, Igl01, OCK09, Sun02, Tru07]. Vector [Kha09, AT07, ABB00, ATV09, AFK02, BB04d, BB05b, BD06a, BDM03, BES09, Bon03b, BRZ04, De 09c, FI03, Gra04a, Hor04, dH07, HP01b, Kir05, Kir06, KN00b, Kle00, KMS05a, Mel00, Mol05a, Mue09, NP07, OCK09, Sal00, SC01, Van04, WL08, Wie01, Mül03]. vector-bundle [Gra04a]. vector-mixed [Kle00]. vector-valued [AT07, BES09]. vectorial [Fer04]. vectors [DK08b, Fer03, Fon00, Gue08, Kar07, KA06, Kun06, MM03, QZ07, SS03a, SG09, Tan09b]. Veen [EG01]. Verdière [Gol09]. Verification [CSY01]. Verifying [DS08b]. Verlag [Kol05]. Veronese [Cof03, GH00b]. Versal [FP01, PPT04, Pue05, CHPP06, Mai01]. version [BV02, Bha06, BT04b, CKS06, Gra04a, Leš04, Lur06a, Pep08, TP05, Vis00, zY08, Zha03a]. versions [Bou06b, Har07a]. versus [BL09b, JS07a]. vertex [HD03, KYAA08, KMNS02, Sav05, SS08a, Ste03a, WX07a]. vertex-deleted [SS08a]. vertex-transitive [Sav05]. Verticalité [Fao01b]. vertices [BHL09, GL08, Guo05d, Guo07b, GW09c, Gut08, HHL^{+}09, HW07a, II09, KS08, Kov02, Li08b, LZS09, Mol08b, Sci02, Sta09a, WXH05, YL06]. very [SS06b]. Veselov [CL03a]. via [AD03a, Alf00, AP04, BF01, Bas02, Ber04, BW06, BRVS08, CZ00, CK01, CN06, wCL04, CR02b, FST04, Fur03, Fur05a, Fur06b, Fur09, GVV03, GHS09, GB06, Gin09, HPP05, HP01b, Ili08, KRS09, KP09, LL06a, Lim09c, Ois01, OP05, PR02b, PWZ05, RS06a, RZ04, RZ05, SC07, Sou05, Ste08, Sun02, XM02, Yam06, Yan01a, xZxGpL09, dKNvD08]. vibrating [Bai08, Lan05]. view [AFR06, BMS08a, DV01, GVV03, HN04]. viewpoint [CHP07]. viewpoints [GJK04]. Villani [Hil07b, LS09b]. vines [KC03, KC06]. Virasoro [SSW08, TX09]. Virasoro-like [SSW08, TX09]. visualization [SSB09]. Vitae [Ant07, Ano06h, FGH^{+}06]. Vlastimil [Sch00b]. vol [Kol05, Moh05]. voltages [KK01]. Volume [AdO02, Ano03r, Ano04r, Ano09r, BI00, BKR^{+}09, Ost03, Sav06a, ZM08]. volumes [Ano02p, Ano02y, Ano03s, Ano03-47, Ano04s, Ano04-45, Ano05o, Ano05-32, Ano05-34, DS04b, EMBV03, Mat07]. Vorst [Ano09n]. vortices [CP09]. vs [HZ04a, MV07, Rod06].

Walks [Nik06c, SdAdFDV07, BPS07, Chu07, Kly00]. Wallach [SP09]. Wang [Moh05]. wave [BGGC06, GZHT05, Li06b]. waveform [BRM01]. Wavelet [CCSS03, LCN03]. wavelets [CL02c, Han02a]. way [AEB06, CtBDC09, TBK09, tBSR04, tBS06]. ways [Uhl01]. Weak [Ala00, AS03, JS07a, Per07, PMS05, FJTU00, XZ06]. weakly [HL09, Oku03, Oku06]. Wedderburn [TY05, TY07, TY09]. Wei [Moh05]. weighing [ÁFMNW03, ÁFMNW06, KMS00, KM07]. weighings [ÁFMNW03]. weight [GL05a, JN06, Kun06, wThY09, VR06]. Weighted [ACRS05, MS02a, AM09, Alf01, CN07, CLY08, DB05, Das07, DW00, FGM08, FFN^{+}07, GJW02, HK02b, HL04d, KL07a, MV07, MR02b, PR01, RW02, Roj07a, RR07b, Roj08b, Sat08, wThY09, ZZ00]. weighting [WM05]. weights [Bap06, BV00, CN07, DP02b]. well [KNX07, Yan09b]. well-conditioned [KNX07]. well-known [Yan09b]. Wenzel [Lás07]. Weyl [BR00b, Cao05, CGM05, Cao07b, DMM00b, Dug06, Shi06]. Weyl-type [DMM00b]. where [Mou06]. whether [WX06a]. Which [CLY08, vDH03a, AT07, Bak04c, BB07b, BLS02, BCRS08a, BCRS08b, Das04a, DL92, DL01, Eld06b, ES08, KS02, KN00a, PBA09, Val02a, YLT05a, ZM03, Zha03a]. Whittle [KS09, Kle00]. whose [AGM09, BFS08, BMS07, BS04a, CHK^{+}01, Fer07a, Guo05c, Gut08, HP01a, JD02, JDS03a, LŠ09a, SS08a, Sta07, Sta09b, Toy09, WHAB08, WHBM09b, Wat00, vdH03b, vdH08b]. wider [LMZ03]. width [BCPT05]. widths [MQ03]. Wielandt [Sch02b, CN02a, CN09, FNS01, KSH08, NV03, Sch02c, TS03, zY08]. Wiener [Adu02, AMZ09, CM00b]. wild [BLS05, BBL^{+}05]. Williams [PR07]. Wilson [Nom05b, Vid07]. Windsor [AHST09, Ano09q]. wireless [DPZ06, SSW05]. Wirtinger [BD09a]. Wishart [IWL07]. Wishartness [VC04]. WITHDRAWN [HL08c]. within [FJ05]. without [Fer03, JR00, Yeu04, YW07, ZZY06, Zha04b, fur08b]. Witt [Ahm05]. Wolff [LN06]. word [LL06a, Lim07b]. words [– DJ07]. work [BLN04]. works [ST07]. Workshop [AHST09, Ano05-33, Ano09q, EMS06, Ste07b, dSQS00, FMN04].

World [Bar05, CNX05a, Hig06]. woven [CHSS03]. Wronskians [Wal05a].

x [Kol05, BLNP01, BLG01]. X-rays [BLG01, BLNP01]. xi [Rod09]. XII [MPS06]. xliv [Bar05]. xvii [Böt06b]. xx [Par09b]. xxiv [Tům06].

Yacubovich [MSS07]. Yanagi [Fur05b]. Yang [FLROV05]. years [BMS08b]. yields [SS06a, SS07b]. Yimin [Moh05]. York [Kol05]. Youla [FI07]. Young [CS09b, HK00].

Z [Rod09]. ZAB [Bou05]. Zenger [KN09]. Zero [BGO09, DG07, DN05b, DV07b, Gro08, HZ05a, LK09, OA09, AGKM09, BS02c, BBL^{+}05, BM04c, BP02, BD03d, CL08a, DKL^{+}07, – DJ07, Fie01b, GGM^{+}09, GKW00, Hor02, HC03a, HZ04b, HJ08, Kah06, Lim08a, Lim09a, Nov09, Okn01, SV07, Tak00, Ton04, Tre03c, Wil09, ZYP06, ZX05b, dF00]. zeroone [Ton04]. Zero-preserving [DG07]. Zero-product [HZ05a]. zero-products [HJ08]. zero-sum [AGKM09]. zerosymmetric [CL08a]. zero-term [BS02c]. zeros [Baz03, BGLS05, BG09b, BD03d, CMV03, CFL05a, CL08b, FLPW03, GS04a, HS07b, HKKL05, LM06b, Lin02a, Pet08, WO05, Wym02, ZWPR06, dSR05]. Zeta [KS06e, MS01b, AFR06, BFL05, Che09, Hor07, MS02a, MS07, Ste08]. Zetlin [Ovs03]. zrank [YYZ05].

References

Abeledo:2007:UCN

[AA07] Hernán Abeledo and Gary W. Atkinson. Unimodularity of the Clar number problem. Linear Algebra and its Applications, 420(2–3):441– 448, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ahmad:2009:PCP

[AA09] Sk. Safique Ahmad and Rafikul Alam. Pseudospectra, critical points and multiple eigenvalues of matrix polynomials. Linear Algebra and its Applications, 430(4): 1171–1195, February 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Abdallah:2002:SLM

[AAA^{+}02] C. T. Abdallah, F. Amato, M. Ariola, P. Dorato, and V. Koltchinskii. Statistical learning methods in linear algebra and control problems: the example of finitetime control of uncertain linear systems. Linear Algebra and its Applications, 351– 352:11–26, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akbari:2007:SRB

[AAB^{+}07] Saieed Akbari, Alireza Alipour, Javad Ebrahimi Boroojeni, Ebrahim Ghorbani, and Mirhamed Mirjalalieh Shirazi. Some relations between rank of a graph and its complement. Linear Algebra and its Applications, 422(1):341– 347, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-

3795 (print), 1873-1856 (electronic).

Ahmadi:2009:GIS

[AABS09] Omran Ahmadi, Noga Alon, Ian F. Blake, and Igor E. Shparlinski. Graphs with integral spectrum. Linear Algebra and its Applications, 430 (1):547–552, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2003:SAG

[AADL03] Daniel Alpay, Tomas Azizov, Aad Dijksma, and Heinz Langer. The Schur algorithm for generalized Schur functions III: J-unitary matrix polynomials on the circle. Linear Algebra and its Applications, 369(C):113– 144, August 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
31/abstract.html.

Akbari:2006:BNS

[AAGK06] Saieed Akbari, Alireza Alipour, Ebrahim Ghorbani, and Gholamreza B. Khosrovshahi. {−1, 0, 1}-basis for the null space of a forest. Linear Algebra and its Applications, 414(2–3):506–511, April 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2006:RTS

[AAV06] Daniel Alpay, Haim Attia, and Dan Volok. Realization theorems for stationary multiscale systems. Linear Algebra and its Applications, 412(2–3):326–347, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alam:2003:ELP

[AB03a] R. Alam and S. Bora. Effect of linear perturbation on spectra of matrices. Linear Algebra and its Applications, 368(C):329–342, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/217/25/
                                        46/abstract.html.

Alam:2003:SES

[AB03b] R. Alam and S. Bora. Stability of eigenvalues and spectral decompositions under linear perturbation. Linear Algebra and its Applications, 364(C):189–211, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/213/25/
                                        29/abstract.html.

Alam:2005:SEE

[AB05a] R. Alam and S. Bora. On sensitivity of eigenvalues and eigendecompositions of matrices. Linear Algebra and its Applications, 396(1):273– 301, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Amaral:2005:CBT

[AB05b] Paula Amaral and Pedro Barahona. Connections between the total least squares and the correction of an infeasible system of linear inequalities. Linear Algebra and its Applications, 395(1): 191–210, January 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Aujla:2007:EIC

[AB07] Jaspal Singh Aujla and Jean-Christophe Bourin. Eigenvalue inequalities for convex and log-convex functions. Linear Algebra and its Applications, 424(1):25– 35, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Allen:2000:NAP

[ABB00] E. J. Allen, J. Baglama, and S. K. Boyd. Numerical approximation of the product of the square root of a matrix with a vector. Linear Algebra and its Applications, 310 (1–3):167–181, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/127/25/
                                        39/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/127/25/39/article.
                                        pdf.

Abdollahi:2005:NRC

[Abd05] A. Abdollahi. The numerical range of a composition operator with conformal automorphism symbol. Linear Algebra and its Applications, 408 (1):177–188, October 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Abdollahi:2008:CGF

[Abd08] Alireza Abdollahi. Commuting graphs of full matrix rings over finite fields. Linear Algebra and its Applications, 428(11–12):2947– 2954, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2001:BSN

[ABDD01] Daniel Alpay, Vladimir Bolotnikov, Patrick Dewilde, and Aad Dijksma. Brune sections in the non-stationary case. Linear Algebra and its Applications, 343–344(1): 389–418, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/

gej-ng/10/30/19/174/27/
45/abstract.html.

Abeles:2008:DCH

[Abe08] Francine F. Abeles. Dodgson condensation: The historical and mathematical development of an experimental method. Linear Algebra and its Applications, 429 (2–3):429–438, July 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ahmadi:2003:PDS

[ABH03] Mohammad H. Ahmadi, Jae-Hyun Baek, and Suk-Geun Hwang. Permanents of doubly stochastic trees. Linear Algebra and its Applications, 370(1):15–24, September 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arav:2007:ILF

[ABH07] Marina Arav, Jean Bevis, and Frank J. Hall. Inherited LU-factorizations of matrices. Linear Algebra and its Applications, 427(1):26– 41, November 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2002:SAR

[ABK02] Daniel Alpay, Vladimir Bolotnikov, and H. Turgay Kaptanoğlu. The Schur algorithm and reproducing kernel Hilbert spaces in the ball. Linear Algebra and its Applications, 342(1–3): 163–186, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/173/27/
38/abstract.html.

Abramovich:2009:NPI

[ABP09] Shoshana Abramovich, Josipa Barić, and Josip Pečarić. A new proof of an inequality of Bohr for Hilbert space operators. Linear Algebra and its Applications, 430(4): 1432–1435, February 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2002:TSR

[ABR02] D. Alpay, V. Bolotnikov, and L. Rodman. Two-sided residue interpolation in matrix H_{2} spaces with symmetries: conformal conjugate involutions. Linear Algebra and its Applications, 351–352:27– 68, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akkurt:2005:SMA

[ABW05] Mustafa Akkurt, George Phillip Barker, and Marcel Wild. Structural matrix algebras and their lattices of invariant subspaces. Linear Algebra and its Applications, 394 (1):25–38, January 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

AncocheaBermudez:2003:CFL

[AC03] José María Ancochea Bermúdez and Rutwig Campoamor. Completable filiform Lie algebras. Linear Algebra and its Applications, 367(C):185– 191, July 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
38/abstract.html.

Alnajjar:2004:FTP

[AC04a] Hasan Alnajjar and Brian Curtin. A family of tridiagonal pairs. Linear Algebra and its Applications, 390(1):369– 384, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Astolfi:2004:NEP

[AC04b] A. Astolfi and P. Colaneri. A note on the existence of positive realizations. Linear Algebra and its Applications, 390 (1):329–343, October 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Agaev:2005:SNL

[AC05] Rafig Agaev and Pavel Chebotarev. On the spectra of nonsymmetric Laplacian matrices. Linear Algebra and its Applications, 399(1):157– 168, April 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alnajjar:2008:BFT

[AC08] Hasan Alnajjar and Brian Curtin. A bilinear form for tridiagonal pairs of q-Serre type. Linear Algebra and its Applications, 428(11–12): 2688–2698, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Antezana:2009:SVE

[AC09] Jorge Antezana and Gustavo Corach. Singular value estimates of oblique projections. Linear Algebra and its Applications, 430(1):386– 395, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alonso:2004:NES

[ACDR04] Pedro Alonso, Raquel Cortina, Irene Díaz, and José Ranilla. Neville elimination: a study of the efficiency using checkerboard partitioning. Linear Algebra and its Applications, 393(1):3–14, December 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arias:2008:PIS

[ACG08] M. Laura Arias, Gustavo Corach, and M. Celeste Gonzalez. Partial isometries in semi-Hilbertian spaces. Linear Algebra and its Applications, 428(7):1460–1475, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alonso:2001:SPN

[ACHR01] Pedro Alonso, Raquel Cortina, Vicente Hernández, and José Ranilla. A study of the performance of Neville elimination using two kinds of partitioning techniques. Linear Algebra and its Applications, 332–334(1):111–117, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/31/article.
pdf.

Ayala:2005:DCL

[ACK05] Victor Ayala, Fritz Colonius, and Wolfgang Kliemann. Dynamical characterization of the Lyapunov form of matrices. Linear Algebra and its Applications, 402(1):272– 290, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Accettella:2003:ITL

[ACM03] Carlo J. Accettella, Gianna M. Del Corso, and Giovanni Manzini. Inversion of two level circulant matrices over Z_{p}. Linear Algebra and its Applications, 366(C):5–23, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/215/25/
                                        27/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Antezana:2005:WPR

[ACRS05] Jorge Antezana, Gustavo Corach, Mariano Ruiz, and Demetrio Stojanoff. Weighted projections and Riesz frames. Linear Algebra and its Applications, 402(1):367–389, June 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Antezana:2004:SSM

[ACS04] Jorge Antezana, Gustavo Corach, and Demetrio Stojanoff. Spectral shorted matrices. Linear Algebra and its Applications, 381(1):197– 217, April 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Antezana:2006:BSO

[ACS06] Jorge Antezana, Gustavo Corach, and Demetrio Stojanoff. Bilateral shorted operators and parallel sums. Linear Algebra and its Applications, 414(2–3):570–588, April 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Adler:2009:NCR

[ACV09] Ilan Adler, Richard W. Cottle, and Sushil Verma. New characterizations of row sufficient matrices. Linear Algebra and its Applications, 430(11–12):2950– 2960, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Albertini:2002:LAS

[AD02a] Francesca Albertini and Domenico D’Alessandro. The Lie algebra structure and controllability of spin systems. Linear Algebra and its Applications, 350(1–3):213– 235, July 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2002:BIB

[AD02b] Daniel Alpay and Chen Dubi. Boundary interpolation in the ball. Linear Algebra and its Applications, 340(1– 3):33–54, January 1, 2002.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
29/abstract.html.

Adkins:2003:CHF

[AD03a] William A. Adkins and Mark G. Davidson. The Cayley–Hamilton and Frobenius theorems via the Laplace transform. Linear Algebra and its Applications, 371(1): 147–152, September 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2003:BSO

[AD03b] D. Alpay and C. Dubi. Backward shift operator and finite dimensional de Branges Rovnyak spaces in the ball. Linear Algebra and its Applications, 371(1):277–285, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2003:CFI

[AD03c] D. Alpay and C. Dubi. Carathéodory–Fejér interpolation in the ball. Linear Algebra and its Applications, 359(1–3):1–19, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
26/abstract.html.

Alpay:2004:CFI

[AD04] D. Alpay and C. Dubi. Carathéodory Fejér interpolation in the ball with mixed derivatives. Linear Algebra and its Applications, 382(1): 117–133, May 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Albertini:2005:MIS

[AD05] Francesca Albertini and Domenico D’Alessandro. Model identification for spin networks. Linear Algebra and its Applications, 394(1):237– 256, January 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alijagic:2006:MQS

[AD06] E. Alijagic and P. Dewilde. Minimal quasi-separable realizations for the inverse of a quasi-separable operator. Linear Algebra and its Applications, 414(2–3):445– 463, April 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arov:2008:CMV

[AD08] Damir Z. Arov and H. (Harry) Dym. J-contractive matrix valued functions and related topics, volume 116 of Encyclopedia of mathematics and its applications. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-88300-8 (hardcover). xi + 575 pp. LCCN QA331.7 .A76 2008.

Arico:2008:SAA

[ADC08] A. Aricò, M. Donatelli, and Serra Capizzano. Spectral analysis of the anti-reflective algebra. Linear Algebra and its Applications, 428(2–3): 657–675, January 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2004:FUM

[ADL04] D. Alpay, A. Dijksma, and H. Langer. Factorization of J-unitary matrix polynomials on the line and a Schur algorithm for generalized Nevanlinna functions. Linear Algebra and its Applications, 387(1):313– 342, August 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2006:UFS

[ADL06] D. Alpay, A. Dijksma, and H. Langer. J_{ℓ}-unitary factorization and the Schur algorithm for Nevanlinna functions in an indefinite setting. Linear Algebra and its Applications, 419(2–3):675–709, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ahues:2006:SRC

[AdLV06] Mario Ahues, Filomena Dias d’Almeida, Alain Largillier, and Paulo B. Vasconcelos. Spectral refinement for clustered eigenvalues of quasidiagonal matrices. Linear Algebra and its Applications, 413(2–3):394–402, March 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Araujo:2002:VVL

[AdO02] I. C. Araújo and M. P. de Oliveira. Volume and variance in the linear statistical model. Linear Algebra and its Applications, 357(1–3): 303–306, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
48/abstract.html.

Atti:2006:BDL

[ADT06] Nadia Ben Atti and Gema M. Diaz-Toca. Block diagonalization and LU-equivalence of Hankel matrices. Linear Algebra and its Applications, 412(2–3):247–269, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Adukov:2002:FWH

[Adu02] Victor M. Adukov. Fractional and Wiener–Hopf factorizations. Linear Algebra and its Applications, 340(1– 3):199–213, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
41/abstract.html.

Aharon:2006:UOD

[AEB06] Michal Aharon, Michael Elad, and Alfred M. Bruckstein. On the uniqueness of overcomplete dictionaries, and a practical way to retrieve them. Linear Algebra and its Applications, 416(1): 48–67, July 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Afanasjew:2008:IRK

[AEEG08] Martin Afanasjew, Michael Eiermann, Oliver G. Ernst, and Stefan Güttel. Implementation of a restarted Krylov subspace method for the evaluation of matrix functions. Linear Algebra and its Applications, 429(10):2293– 2314, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpin:2000:CFP

[AEI00] Yu. A. Alpin, L. Elsner, and Kh. D. Ikramov. On condensed forms for partially commuting matrices. Linear Algebra and its Applications, 306(1–3):165–182, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/121/25/38/article.
pdf.

Absil:2006:LPA

[AEK06] P.-A. Absil, A. Edelman, and P. Koev. On the largest principal angle between random subspaces. Linear Algebra and its Applications, 414(1): 288–294, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Albera:2004:BIO

[AFCC04] Laurent Albera, Anne Ferréol, Pierre Comon, and Pascal Chevalier. Blind Identification of Overcomplete MixturEs of sources (BIOME). Linear Algebra and its Applications, 391(1):3–30, November 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Amano:2002:CIR

[AFK02] Katsutoshi Amano, Masaki Fujigami, and Takeyoshi Kogiso. Construction of irreducible relative invariant of the prehomogeneous vector space (SL_{5} × GL_{4}, Λ^{2}(C^{5}) ⊗ C^{4}). Linear Algebra and its Applications, 355(1–3): 215–222, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/201/27/
                                        40/abstract.html.

Alpay:2000:EPD

[AFL00] Daniel Alpay, Boris Freydin, and Philippe Loubaton. An extension problem for discrete-time almost periodically correlated stochastic processes. Linear Algebra and its Applications, 308(1– 3):163–181, March 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/125/25/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/125/25/34/article.
                                        pdf.

Akbari:2003:MCD

[AFM03] S. Akbari, H.-R. Fanaï, and K. Mahmoudian. On the matrices with constant determinant and permanent over roots of unity. Linear Algebra and its Applications, 375(1): 245–249, December 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Abrego:2003:OWD

[ÁFMNW03] Bernardo M. Ábrego, Silvia Fernández-Merchant, Michael G.

Neubauer, and William Watkins. D-optimal weighing designs for n ≡ −1 mod 4 objects and a large number of weighings. Linear Algebra and its Applications, 374(1): 175–218, November 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Abrego:2006:TMG

[ÁFMNW06] Bernardo M. Ábrego, Silvia Fernández-Merchant, Michael G. Neubauer, and William Watkins. Trace-minimal graphs and D-optimal weighing designs. Linear Algebra and its Applications, 412(2–3):161–221, January 15, 2006. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alves:2006:DZF

[AFR06] João F. Alves, J. L. Fachada, and J. Sousa Ramos. Dynamical zeta functions and kneading determinants: A linear algebra point of view. Linear Algebra and its Applications, 418(2–3):913–924, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akelbek:2009:BSI

[AFS09] Mahmud Akelbek, Sandra Fital, and Jian Shen. A bound on the scrambling index of a primitive matrix using Boolean rank. Linear Algebra and its Applications, 431(10):1923–1931, October 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Abdel-Ghaffar:2000:DRP

[AG00] Khaled A. S. Abdel-Ghaffar. The determinant of random power series matrices over finite fields. Linear Algebra and its Applications, 315(1– 3):139–144, August 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/132/25/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/132/25/32/article.
                                        pdf.

Alieva:2004:LPR

[AG04a] Anna A. Alieva and Alexander E. Guterman. Linear preservers of rank permutability. Linear Algebra and its Applications, 384(1):97– 108, June 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ashrafi:2004:IPM

[AG04b] Ashkan Ashrafi and Peter M. Gibson. An involutory Pascal matrix. Linear Algebra and its Applications, 387(1):277– 286, August 1, 2004. CO-DEN LAAPAW. ISSN 0024-

Ashrafi:2005:EHP

[AG05a] Ashkan Ashrafi and Peter M. Gibson. Eigenvalues of Hadamard powers of large symmetric Pascal matrices. Linear Algebra and its Applications, 405(1):60– 66, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Aston:2005:SEK

[AG05b] John A. D. Aston and Roger N. Gunn. Statistical estimation with Kronecker products in positron emission tomography. Linear Algebra and its Applications, 398(1): 25–36, March 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akbari:2008:CNE

[AG08] Saieed Akbari and Ebrahim Ghorbani. Choice number and energy of graphs. Linear Algebra and its Applications, 429(11–12):2687–2690, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Agaoka:2002:ADI

[Aga02a]  Yoshio Agaoka.  An al-

gorithm to determine the isomorphism classes of 4dimensional complex Lie algebras. Linear Algebra and its Applications, 345(1– 3):85–118, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/175/27/
32/abstract.html.

Agashe:2002:ARM

[Aga02b] S. D. Agashe. An algorithm for a result on minimal polynomials. Linear Algebra and its Applications, 357(1–3): 291–293, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
45/abstract.html.

Alves:2005:LAA

[AGDR05] J. F. Alves, M. M. Graça, M. E. Sousa Dias, and J. Sousa Ramos. A linear algebra approach to the conjecture of Collatz. Linear Algebra and its Applications, 394 (1):277–289, January 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akbari:2004:CGS

[AGHM04] S. Akbari, M. Ghandehari, M. Hadian, and A. Mohammadian. On commuting graphs of semisimple rings. Linear Algebra and its Applications, 390(1):345– 355, October 1, 2004. CO-

DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpin:2000:STD

[AGI00] Yurii A. Al’pin, Alan George, and Khakim D. Ikramov. Solving the two-dimensional CIS problem by a rational algorithm. Linear Algebra and its Applications, 312(1– 3):115–123, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/129/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/129/25/31/article.
pdf.

Alieva:2006:RPA

[AGK06] A. A. Alieva, A. E. Guterman, and B. Kuzma. Rankpermutable additive mappings. Linear Algebra and its Applications, 414(2–3):607– 616, April 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akbari:2006:KIM

[AGKM06] Saieed Akbari, Narges Ghareghani, Gholamreza B. Khosrovshahi, and Hamidreza Maimani. The kernels of the incidence matrices of graphs revisited. Linear Algebra and its Applications, 414(2–3):617– 625, April 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akbari:2009:ZSF

[AGKM09] S. Akbari, N. Ghareghani, G. B. Khosrovshahi, and A. Mahmoody. On zerosum 6-flows of graphs. Linear Algebra and its Applications, 430(11–12):3047– 3052, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akbari:2009:GWS

[AGM09] S. Akbari, E. Ghorbani, and A. Mahmoodi. On graphs whose star sets are (co- )cliques. Linear Algebra and its Applications, 430(1):504– 510, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akbari:2009:EAS

[AGO09] Saieed Akbari, Ebrahim Ghorbani, and Mohammad Reza Oboudi. Edge addition, singular values, and energy of graphs and matrices. Linear Algebra and its Applications, 430(8–9):2192– 2199, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Aiena:2008:PPP

[AGP08] Pietro Aiena, Jesús R. Guillen, and Pedro Peña.

Property (w) for perturbations of polaroid operators. Linear Algebra and its Applications, 428(8–9):1791– 1802, April 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

An:2002:RPM

[AH02] Guimei An and Jinchuan Hou. Rank-preserving multiplicative maps on B(X). Linear Algebra and its Applications, 342(1–3):59–78, February 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/173/27/
31/abstract.html.

Avrachenkov:2003:PNS

[AH03] Konstantin E. Avrachenkov and Moshe Haviv. Perturbation of null spaces with application to the eigenvalue problem and generalized inverses. Linear Algebra and its Applications, 369 (C):1–25, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
26/abstract.html.

Avrachenkov:2004:FLS

[AH04] Konstantin E. Avrachenkov and Moshe Haviv. The first Laurent series coefficients for singularly perturbed stochastic matrices. Linear Algebra and its Applications, 386(1): 243–259, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alanelli:2008:NIC

[AH08] M. Alanelli and A. Hadjidimos. A new iterative criterion for H-matrices: The reducible case. Linear Algebra and its Applications, 428(11– 12):2761–2777, June 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

An:2009:CDT

[AH09a] Runling An and Jinchuan Hou. Characterizations of derivations on triangular rings: Additive maps derivable at idempotents. Linear Algebra and its Applications, 431(5–7):1070–1080, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ando:2009:SSO

[AH09b] Kazunori Ando and Yusuke Higuchi. On the spectrum of Schrödinger operators with a finitely supported potential on the d-regular tree. Linear Algebra and its Applications, 431(10):1940–1951, October 15, 2009. CO-DEN LAAPAW. ISSN 0024-

Axtell:2009:OSR

[AHH^{+}09] Jonathan Axtell, Lixing Han, Daniel Hershkowitz, Michael Neumann, and Nung-Sing Sze. Optimization of the spectral radius of a product for nonnegative matrices. Linear Algebra and its Applications, 430(5–6):1442– 1451, March 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arav:2005:RRM

[AHK^{+}05] Marina Arav, Frank J. Hall, Selcuk Koyuncu, Zhongshan Li, and Bhaskara Rao. Rational realizations of the minimum rank of a sign pattern matrix. Linear Algebra and its Applications, 409(1):111– 125, November 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arav:2009:SPR

[AHL^{+}09] Marina Arav, Frank Hall, Zhongshan Li, Assefa Merid, and Yubin Gao. Sign patterns that require almost unique rank. Linear Algebra and its Applications, 430(1): 7–16, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arav:2009:RSC

[AHLR09] Marina Arav, Frank Hall, Zhongshan Li, and Bhaskara Rao. Rational solutions of certain matrix equations. Linear Algebra and its Applications, 430(2–3):660–663, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ahmad:2005:WKP

[Ahm05] Hamza Ahmad. The Witt kernels of purely inseparable quartic extensions. Linear Algebra and its Applications, 395(1):265–273, January 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Antoulas:2007:P

[AHR^{+}07] A. C. Antoulas, U. Helmke, J. Rosenthal, V. Vinnikov, and E. Zerz. Preface. Linear Algebra and its Applications, 425(2–3):217–219, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ahmed:2009:PPI

[AHST09] S. Ejaz Ahmed, Jeffrey J. Hunter, George P. H. Styan, and Götz Trenkler. Preface to the Proceedings of the 16th International Workshop on Matrices and Statistics, Windsor 2007. Linear Algebra and its Applications, 430 (10):2563–2565, May 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Al-Homidan:2005:AEC

[AHW05] Suliman Al-Homidan and Henry Wolkowicz. Approximate and exact completion problems for Euclidean distance matrices using semidefinite programming. Linear Algebra and its Applications, 406(1):109–141, September 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

An:2005:APM

[AHZ05] Runling An, Jinchuan Hou, and Liankuo Zhao. Adjacency preserving maps on the space of symmetric operators. Linear Algebra and its Applications, 405(1):311– 324, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpin:2000:RTP

[AI00] Yurii A. Al’pin and Khakim D. Ikramov. Reducibility theorems for pairs of matrices as rational criteria. Linear Algebra and its Applications, 313 (1–3):155–161, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/130/25/

                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/130/25/34/article.
                                        pdf.

Akbari:2009:CNR

[AJF09] S. Akbari, M. Jamaali, and S. A. Seyed Fakhari. The clique numbers of regular graphs of matrix algebras are finite. Linear Algebra and its Applications, 431(10): 1715–1718, October 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akbari:2007:UG

[AK07] S. Akbari and S. J. Kirkland. On unimodular graphs. Linear Algebra and its Applications, 421(1):3–15, February 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akelbek:2009:CES

[AK09a] Mahmud Akelbek and Steve Kirkland. Coefficients of ergodicity and the scrambling index. Linear Algebra and its Applications, 430(4): 1111–1130, February 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akelbek:2009:PDL

[AK09b] Mahmud Akelbek and Steve Kirkland. Primitive digraphs with the largest scrambling index. Linear Algebra and its Applications, 430(4):1099– 1110, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akbar:2009:SPP

[AKGC09] K. Ali Akbar, V. Kannan, Sharan Gopal, and P. Chiranjeevi. The set of periods of periodic points of a linear operator. Linear Algebra and its Applications, 431(1– 2):241–246, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ahn:2007:ELT

[AKL07] Eunkyung Ahn, Sejung Kim, and Yongdo Lim. An extended Lie–Trotter formula and its applications. Linear Algebra and its Applications, 427(2–3):190–196, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Altinisik:2016:CIH

[AKYD16] Ercan Altinisik, Ali Keskin, Mehmet Yildiz, and Murat Demirbüken. On a conjecture of Ilmonen, Haukkanen and Merikoski concerning the smallest eigenvalues of certain GCD related matrices. Linear Algebra and its Applications, 493 (??):1–13, March 15, 2016. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0024379515006977.

See [IHM08].

Ahmed:2009:ATS

[AL09a] S. Ejaz Ahmed and Shuangzhe Liu. Asymptotic theory of simultaneous estimation of Poisson means. Linear Algebra and its Applications, 430 (10):2734–2748, May 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akram:2009:LSL

[AL09b] M. Saeed Akram and Vakhtang Lomadze. Linear systems with locally integrable trajectories. Linear Algebra and its Applications, 430(8–9):2277– 2289, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alaimia:2000:WCP

[Ala00] M. R. Alaimia. Weak crossed products and a generalisation of a result of Sarason. Linear Algebra and its Applications, 318(1–3): 181–193, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/137/24/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/137/24/36/article.
                                        pdf.

Alam:2005:CND

[Ala05] Rafikul Alam. On the construction of nearest defective matrices to a normal matrix. Linear Algebra and its Applications, 395(1):367– 370, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alonso:2006:DFP

[ALC06] J. A. Hermida Alonso and M. M. López-Cabeceira. Dynamic feedback over principal ideal domains and quotient rings. Linear Algebra and its Applications, 413(1):235– 244, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alonso:2005:WDS

[ALCT05] J. A. Hermida Alonso, M. M. López-Cabeceira, and M. T. Trobajo. When are dynamic and static feedback equivalent? Linear Algebra and its Applications, 405(1):74– 82, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ahues:2005:SRQ

[ALdV05] Mario Ahues, Alain Largillier, Filomena Dias d’Almeida, and Paulo B. Vasconcelos. Spectral refinement on quasidiagonal matrices. Linear Algebra and its Applications, 401(1):109–117, May 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alfakih:2000:GRE

[Alf00] Abdo Y. Alfakih. Graph rigidity via Euclidean distance matrices. Linear Algebra and its Applications, 310 (1–3):149–165, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/127/25/38/article.
pdf.

Alfakih:2001:RRW

[Alf01] Abdo Y. Alfakih. On rigidity and realizability of weighted graphs. Linear Algebra and its Applications, 325(1– 3):57–70, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/147/24/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/147/24/27/article.
pdf.

Alfakih:2003:UED

[Alf03] Abdo Y. Alfakih. On the uniqueness of Euclidean distance matrix completions. Linear Algebra and its Applications, 370(1):1–14, September 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alfakih:2005:UED

[Alf05] Abdo Y. Alfakih. On the uniqueness of Euclidean distance matrix completions: the case of points in general position. Linear Algebra and its Applications, 397(1):265– 277, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alfakih:2006:NRC

[Alf06a] A. Y. Alfakih. On the nullspace, the rangespace and the characteristic polynomial of Euclidean distance matrices. Linear Algebra and its Applications, 416(2–3):348– 354, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alfakih:2006:RFC

[Alf06b] A. Y. Alfakih. A remark on the faces of the cone of Euclidean distance matrices. Linear Algebra and its Applications, 414(1):266– 270, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alfakih:2008:DRM

[Alf08] A. Y. Alfakih. On the dual rigidity matrix. Linear Algebra and its Applications, 428(4):962–972, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alizadeh:2009:JDF

[Ali09] R. Alizadeh. Jordan derivations of full matrix algebras. Linear Algebra and its Applications, 430(1):574– 578, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ando:2004:GM

[ALM04] T. Ando, Chi-Kwong Li, and Roy Mathias. Geometric means. Linear Algebra and its Applications, 385(1):305– 334, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alwill:2006:ACG

[ALMS06] Michael Alwill, Chi-Kwong Li, Christina Maher, and Nung-Sing Sze. Automorphisms of certain groups and semigroups of matrices. Linear Algebra and its Applications, 412(2–3):490– 525, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Abreu:2008:HSG

[ALSS08] M. Abreu, D. Labbate, R. Salvi, and N. Zagaglia Salvi. Highly symmetric generalized circulant permutation matrices. Linear Algebra and its Applications, 429(1): 367–375, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Altinisik:2009:IGM

[Alt09] Ercan Altınışık. On inverses of GCD matrices associated with multiplicative functions and a proof of the Hong–Loewy conjecture. Linear Algebra and its Applications, 430(4):1313–1327, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alzer:2001:CMI

[Alz01] Horst Alzer. A converse of a matrix inequality. Linear Algebra and its Applications, 323(1–3):195–199, January 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Adam:2002:CNM

[AM02] Maria Adam and John Maroulas. On compressions of normal matrices. Linear Algebra and its Applications, 341(1–3):403–418, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/

                                        gej-ng/10/30/19/171/27/
                                        60/abstract.html.

Arndt:2004:SCI

[AM04] Hans-Robert Arndt and Günter Mayer. On the semiconvergence of interval matrices. Linear Algebra and its Applications, 393(1):15– 37, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Azenhas:2005:ASG

[AM05] Olga Azenhas and Ricardo Mamede. Action of the symmetric group on sets of skewtableaux with prescribed matrix realization. Linear Algebra and its Applications, 401 (1):221–275, May 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Accardi:2009:NNU

[AM09] Luigi Accardi and Farrukh Mukhamedov. A note on noncommutative unique ergodicity and weighted means. Linear Algebra and its Applications, 430(2–3):782–790, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ambrozie:2007:FPM

[Amb07] C.-G. Ambrozie. Finding positive matrices subject to linear restrictions. Linear Algebra and its Applications, 426(2–3):716–728, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Andruchow:2009:MMC

[AMLM^{+}09] Esteban Andruchow, Luis E. Mata-Lorenzo, Alberto Mendoza, Lázaro Recht, and Alejandro Varela. Minimal matrices and the corresponding minimal curves on flag manifolds in low dimension. Linear Algebra and its Applications, 430(8–9):1906– 1928, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Appa:2006:MIA

[AMM06] G. Appa, D. Magos, and I. Mourtos. On multi-index assignment polytopes. Linear Algebra and its Applications, 416(2–3):224–241, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Abara:2007:SCM

[AMP07] Ma. Nerissa M. Abara, Dennis I. Merino, and Agnes T. Paras. Skew-coninvolutory matrices. Linear Algebra and its Applications, 426(2–3): 540–557, October 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akbari:2006:DCG

[AMRR06] S. Akbari, A. Mohammadian, H. Radjavi, and P. Raja. On the diameters of commuting graphs. Linear Algebra and its Applications, 418(1):161– 176, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Antezana:2005:ATS

[AMS05] Jorge Antezana, Pedro Massey, and Demetrio Stojanoff. λ-Aluthge transforms and Schatten ideals. Linear Algebra and its Applications, 405 (1):177–199, August 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Amparan:2004:AIS

[AMZ04] A. Amparan, S. Marcaida, and I. Zaballa. Assignment of infinite structure to an openloop system. Linear Algebra and its Applications, 379(1): 249–266, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Amparan:2006:ELS

[AMZ06] A. Amparan, S. Marcaida, and I. Zaballa. On the existence of linear systems with prescribed invariants for system similarity. Linear Algebra and its Applications, 413(2–3):510–533, March 1, 2006. CODEN

LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Amparan:2007:LRL

[AMZ07] A. Amparan, S. Marcaida, and I. Zaballa. Local realizations and local polynomial matrix representations of systems. Linear Algebra and its Applications, 425(2–3):757– 775, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Amparan:2009:LWH

[AMZ09] A. Amparan, S. Marcaida, and I. Zaballa. Local Wiener– Hopf factorization and indices over arbitrary fields. Linear Algebra and its Applications, 430(5–6):1700– 1722, March 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ando:2004:CNT

[And04a] T. Ando. Cones and norms in the tensor product of matrix spaces. Linear Algebra and its Applications, 379(1): 3–41, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ando:2004:LII

[And04b] T. Ando. Löwner inequality of indefinite type. Linear Algebra and its Applications, 385(1):73–80, July 1, 2004.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ando:2004:SMC

[And04c] T. Ando. Sets of matrices with common Stein solutions and H-contractions. Linear Algebra and its Applications, 383(1):49–64, May 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ando:2006:UIN

[And06] T. Ando. Unitarily invariant norms related to the numerical radius. Linear Algebra and its Applications, 417(1): 3–7, August 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Andai:2007:CQS

[And07a] Attila Andai. On the curvature of the quantum state space with pull-back metrics. Linear Algebra and its Applications, 423(2–3):287– 304, June 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ando:2007:RBR

[And07b] T. Ando. Review of the book of Rajendra Bhatia “Positive Definite Matrices”. Linear Algebra and its Applications, 426(2–3):855–858, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-

Ando:2009:PKS

[And09]  Tsuyoshi Ando.  Projec-

tions in Krein spaces. Linear Algebra and its Applications, 431(12):2346–2358, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2000:Ia

[Ano00a] Anonymous. Index. Linear Algebra and its Applications, 304(1–3):201, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/118/25/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/118/25/39/article.
pdf.

Anonymous:2000:Ib

[Ano00b] Anonymous. Index. Linear Algebra and its Applications, 305(1–3):203, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/120/25/
40/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/120/25/40/article.
pdf.

Anonymous:2000:Ic

[Ano00c] Anonymous. Index. Linear Algebra and its Applications, 306(1–3):211, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
42/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/121/25/42/article.
pdf.

Anonymous:2000:Id

[Ano00d] Anonymous. Index. Linear Algebra and its Applications, 307(1–3):195, March 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/39/article.
pdf.

Anonymous:2000:Ie

[Ano00e] Anonymous. Index. Linear Algebra and its Applications, 308(1–3):217, March 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/125/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/125/25/37/article.
pdf.

Anonymous:2000:If

[Ano00f] Anonymous. Index. Linear Algebra and its Applications, 309(1–3):363, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
43/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/43/article.
pdf.

Anonymous:2000:Ig

[Ano00g] Anonymous. Index. Linear Algebra and its Applications, 310(1–3):195, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
41/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/127/25/41/article.
pdf.

Anonymous:2000:Ih

[Ano00h] Anonymous. Index. Linear Algebra and its Applications, 311(1–3):201, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/128/25/
40/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/128/25/40/article.
pdf.

Anonymous:2000:Ii

[Ano00i] Anonymous. Index. Linear Algebra and its Applications, 312(1–3):203, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/129/25/
                                        39/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/129/25/39/article.
                                        pdf.

Anonymous:2000:Ij

[Ano00j] Anonymous. Index. Linear Algebra and its Applications, 313(1–3):207, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/130/25/
                                        39/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/130/25/39/article.
                                        pdf.

Anonymous:2000:Ik

[Ano00k] Anonymous. Index. Linear Algebra and its Applications, 314(1–3):205, July 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/131/25/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/131/25/32/article.
                                        pdf.

Anonymous:2000:Il

[Ano00l] Anonymous. Index. Linear Algebra and its Applications, 315(1–3):209, August 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/132/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/132/25/38/article.
pdf.

Anonymous:2000:Im

[Ano00m] Anonymous. Index. Linear Algebra and its Applications, 316(1–3):287, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
41/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/41/article.
pdf.

Anonymous:2000:In

[Ano00n] Anonymous. Index. Linear Algebra and its Applications, 317(1–3):241, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/136/25/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/136/25/39/article.
pdf.

Anonymous:2000:Io

[Ano00o] Anonymous. Index. Linear Algebra and its Applications, 318(1–3):227, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/137/24/
                                        39/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/137/24/39/article.
                                        pdf.

Anonymous:2000:Ip

[Ano00p] Anonymous. Index. Linear Algebra and its Applications, 319(1–3):211, November 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/138/24/
                                        37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/138/24/37/article.
                                        pdf.

Anonymous:2000:Iq

[Ano00q] Anonymous. Index. Linear Algebra and its Applications, 320(1–3):207, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/139/24/
                                        40/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/139/24/40/article.
                                        pdf.

Anonymous:2000:Ir

[Ano00r] Anonymous. Index. Linear Algebra and its Applications, 320(1–3):221–231, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
42/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/42/article.
pdf.

Anonymous:2000:Is

[Ano00s] Anonymous. Index. Linear Algebra and its Applications, 321(1–3):411–412, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
51/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/51/article.
pdf.

Anonymous:2001:AIa

[Ano01a]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 324(1–3): 235, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/144/24/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/144/24/37/article.
pdf.

Anonymous:2001:AIb

[Ano01b] Anonymous. Author index. Linear Algebra and its Applications, 325(1–3):221, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/147/24/
                                        39/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/147/24/39/article.
                                        pdf.

Anonymous:2001:AIc

[Ano01c] Anonymous. Author index. Linear Algebra and its Applications, 326(1–3):225, March 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/148/24/
                                        39/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/148/24/39/article.
                                        pdf.

Anonymous:2001:AId

[Ano01d]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 327(1–3): 225, April 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej%2Dng/10/
                                        30/19/152/25/42/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej%2Dng/10/30/19/152/
                                        25/42/article.pdf.

Anonymous:2001:AIe

[Ano01e]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 328(1–3): 223, May 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/153/25/34/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/153/
25/34/article.pdf.

Anonymous:2001:AIf

[Ano01f]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 329(1–3): 189, May 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/154/25/36/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/154/
25/36/article.pdf.

Anonymous:2001:AIg

[Ano01g] Anonymous. Author index. Linear Algebra and its Applications, 330(1–3):219, June 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/155/25/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/155/25/39/article.
pdf.

Anonymous:2001:AIh

[Ano01h] Anonymous. Author index. Linear Algebra and its Applications, 331(1–3):211, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
42/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/42/article.
pdf.

Anonymous:2001:AIi

[Ano01i]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 332–334(1): 611–612, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
58/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/58/article.
pdf.

Anonymous:2001:AIj

[Ano01j]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 335(1–3): 205, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/165/27/
40/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/165/27/40/article.
pdf.

Anonymous:2001:AIk

[Ano01k]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 336(1–3): 265, October 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
46/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/166/27/46/article.
pdf.

Anonymous:2001:AIl

[Ano01l]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 337(1–3): 269, November 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/167/27/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/167/27/39/article.
pdf.

Anonymous:2001:AIm

[Ano01m]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 338(1–3): 291, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
49/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/168/27/49/article.
                                        pdf.

Anonymous:2001:AIn

[Ano01n]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 339(1–3): 221, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/169/27/
                                        40/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/169/27/40/article.
                                        pdf.

Anonymous:2001:AIo

[Ano01o]  Anonymous.  Author In-

dex. Linear Algebra and its Applications, 343–344(1): 477–478, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/174/27/
                                        49/abstract.html.

Anonymous:2001:I

[Ano01p] Anonymous. Index. Linear Algebra and its Applications, 322(1–3):219, January 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/141/24/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/141/24/36/article.
                                        pdf.

Anonymous:2002:AIa

[Ano02a]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 340(1–3): 255, January 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
46/abstract.html.

Anonymous:2002:AIb

[Ano02b]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 341(1–3): 419–420, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
61/abstract.html.

Anonymous:2002:AIc

[Ano02c]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 342(1–3): 219, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/173/27/
41/abstract.html.

Anonymous:2002:AId

[Ano02d] Anonymous. Author index. Linear Algebra and its Applications, 345(1–3):269, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/175/27/
                                        42/abstract.html.

Anonymous:2002:AIe

[Ano02e] Anonymous. Author index. Linear Algebra and its Applications, 346(1–3):293, May 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/189/27/
                                        41/abstract.html.

Anonymous:2002:AIf

[Ano02f] Anonymous. Author index. Linear Algebra and its Applications, 347(1–3):293, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/194/27/
                                        45/abstract.html.

Anonymous:2002:AIg

[Ano02g]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 348(1–3): 289–290, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/195/27/
                                        50/abstract.html.

Anonymous:2002:AIh

[Ano02h]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 349(1–3): 283, July 1, 2002. CO-

Anonymous:2002:AIi

[Ano02i]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 350(1–3): 293, July 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2002:AIj

[Ano02j] Anonymous. Author index. Linear Algebra and its Applications, 351–352:825–826, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2002:AIk

[Ano02k]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 353(1–3): 317, September 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2002:AIl

[Ano02l]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 354(1–3): 291, October 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
50/abstract.html.

Anonymous:2002:AIm

[Ano02m]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 355(1–3): 315, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/201/27/
46/abstract.html.

Anonymous:2002:AIn

[Ano02n]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 356(1–3): 275, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/202/25/
44/abstract.html.

Anonymous:2002:AIo

[Ano02o]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 357(1–3): 309–310, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
50/abstract.html.

Anonymous:2002:AIV

[Ano02p] Anonymous. Author index to volumes 301–320. Linear Algebra and its Applications, 340(1–3):263–273, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
49/abstract.html.

Anonymous:2002:CP

[Ano02q] Anonymous. Call for papers. Linear Algebra and its Applications, 354(1–3): 289–290, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
49/abstract.html.

Anonymous:2002:CPS

[Ano02r] Anonymous. Call for papers: special issue in honor of Heydar Radjavi. Linear Algebra and its Applications, 357 (1–3):307–308, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
49/abstract.html.

Anonymous:2002:CMT

[Ano02s] Anonymous. Challenges in matrix theory 2002. Linear Algebra and its Applications, 345(1–3):261, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/175/27/
40/abstract.html.

Anonymous:2002:EBa

[Ano02t] Anonymous. Editorial Board. Linear Algebra and its Applications, 355(1–3):ii–iii, November 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/201/27/
                                        27/abstract.html.

Anonymous:2002:EBb

[Ano02u] Anonymous. Editorial board. Linear Algebra and its Applications, 356(1–3):ii–iii, November 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/202/25/
                                        25/abstract.html.

Anonymous:2002:EBc

[Ano02v] Anonymous. Editorial board. Linear Algebra and its Applications, 357(1–3):ii–iii, December 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/203/25/
                                        25/abstract.html.

Anonymous:2002:LE

[Ano02w] Anonymous. List of Editors. Linear Algebra and its Applications, 354(1–3): ii–iii, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
27/abstract.html.

Anonymous:2002:P

[Ano02x] Anonymous. Preface. Linear Algebra and its Applications, 356(1–3):1–2, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/202/25/
26/abstract.html.

Anonymous:2002:RV

[Ano02y] Anonymous. Referees for volumes 321–340. Linear Algebra and its Applications, 340(1–3):259–262, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
48/abstract.html.

Anonymous:2002:SIH

[Ano02z]  Anonymous.  Special is-

sue in honor of Peter Lancaster. Linear Algebra and its Applications, 353(1–3): 311, September 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2002:SILb

[Ano02-27] Anonymous. Special issue on large scale linear and N …. Linear Algebra and its Applications, 353(1–3):315–316, September 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2002:SILa

[Ano02-28] Anonymous. Special issue on linear algebra in signal processing. Linear Algebra and its Applications, 353(1–3): 313–314, September 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2002:SIO

[Ano02-29] Anonymous. Special issue on order reduction of large-scale systems. Linear Algebra and its Applications, 353(1–3): 309–310, September 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2003:AIa

[Ano03a]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 358(1–3): 425, January 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/207/25/
                                        46/abstract.html.

Anonymous:2003:AIb

[Ano03b]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 359(1–3): 293, January 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
48/abstract.html.

Anonymous:2003:AIc

[Ano03c] Anonymous. Author index. Linear Algebra and its Applications, 361(C):279, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
44/abstract.html.

Anonymous:2003:AId

[Ano03d]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 362(C):301– 302, March 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/211/25/
50/abstract.html.

Anonymous:2003:AIe

[Ano03e]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 363(C):333– 334, April 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
47/abstract.html.

Anonymous:2003:AIf

[Ano03f]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 364(C):325– 326, May 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
46/abstract.html.

Anonymous:2003:AIg

[Ano03g] Anonymous. Author Index. Linear Algebra and its Applications, 365(C):467, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
54/abstract.html.

Anonymous:2003:AIh

[Ano03h]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 366(C):483– 484, June 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
52/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Anonymous:2003:AIi

[Ano03i]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 367(C):349– 350, July 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
52/abstract.html.

Anonymous:2003:AIj

[Ano03j]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 368(C):387– 388, July 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
51/abstract.html.

Anonymous:2003:AIk

[Ano03k]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 369(C):351– 352, August 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
47/abstract.html.

Anonymous:2003:AIl

[Ano03l]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 370(1):379–380, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2003:AIm

[Ano03m]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 371(1):377–378, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2003:AIn

[Ano03n]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 372(1):345, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2003:AIo

[Ano03o]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 373(1):331– 332, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2003:AIp

[Ano03p]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 374(1):327–328, November 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2003:AIq

[Ano03q]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 375(1):319– 320, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2003:AIVa

[Ano03r] Anonymous. Author index to volume 360. Linear Algebra and its Applications, 360 (C):277, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
46/abstract.html.

Anonymous:2003:AIVb

[Ano03s] Anonymous. Author index to volumes 341–360. Linear Algebra and its Applications, 360(C):289–302, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
49/abstract.html.

Anonymous:2003:CPSb

[Ano03t] Anonymous. Call for papers: Special issue in honor of Graciano de Oliveira. Linear Algebra and its Applications, 360(C):273, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
44/abstract.html.

Anonymous:2003:CPSc

[Ano03u] Anonymous. Call for papers: Special issue on determinants and the legacy of Sir Thomas Muir. Linear Algebra and its Applications, 360(C):275– 276, February 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/209/25/
                                        45/abstract.html.

Anonymous:2003:CPSa

[Ano03v] Anonymous. Call for papers: Special issue on positivity in linear algebra. Linear Algebra and its Applications, 359 (1–3):291–292, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/208/25/
                                        47/abstract.html.

Anonymous:2003:EBa

[Ano03w] Anonymous. Editorial board. Linear Algebra and its Applications, 358(1–3):ii–iii, January 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/207/25/
                                        25/abstract.html.

Anonymous:2003:EBb

[Ano03x] Anonymous. Editorial board. Linear Algebra and its Applications, 362(C):ii–iii, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/211/25/
                                        25/abstract.html.

Anonymous:2003:EBc

[Ano03y] Anonymous. Editorial board. Linear Algebra and its Applications, 366(C):ii–iii, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
25/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Anonymous:2003:EBd

[Ano03z] Anonymous. Editorial board. Linear Algebra and its Applications, 367(C):ii–iii, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
25/abstract.html.

Anonymous:2003:EBe

[Ano03-27] Anonymous. Editorial board. Linear Algebra and its Applications, 369(C):ii–iii, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
25/abstract.html.

Anonymous:2003:EBf

[Ano03-28] Anonymous. Editorial board. Linear Algebra and its Applications, 370(1):ii–iii, September 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2003:EBg

[Ano03-29] Anonymous. Editorial board. Linear Algebra and its Applications, 373(1):ii–iii, November 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2003:EBh

[Ano03-30] Anonymous. Editorial board. Linear Algebra and its Applications, 374(1):ii–iii, November 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2003:ERT

[Ano03-31] Anonymous. Erratum to: “Ranks of tensors, secant varieties of Segre varieties and fat points” [Linear Algebra Appl. 355 (2002) 263– 285]. Linear Algebra and its Applications, 367(C):347– 348, July 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
51/abstract.html;   http:
//www.elsevier.nl/PII/S002437950200352X.

See [CGG02].

Anonymous:2003:LEa

[Ano03-32] Anonymous. List of Editors. Linear Algebra and its Applications, 359(1–3): ii–iii, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

gej-ng/10/30/19/208/25/
25/abstract.html.

Anonymous:2003:LEd

[Ano03-33] Anonymous. List of editors. Linear Algebra and its Applications, 363(C):ii–iii, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
25/abstract.html.

Anonymous:2003:LEj

[Ano03-34] Anonymous. List of Editors. Linear Algebra and its Applications, 375(1):ii–iii, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2003:LEb

[Ano03-35] Anonymous. Lists of Editors. Linear Algebra and its Applications, 360(C):ii– iii, February 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
25/abstract.html.

Anonymous:2003:LEc

[Ano03-36] Anonymous. Lists of Editors. Linear Algebra and its Applications, 361(C):ii–iii, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

gej-ng/10/30/19/210/25/
25/abstract.html.

Anonymous:2003:LEe

[Ano03-37] Anonymous. Lists of editors. Linear Algebra and its Applications, 364(C):ii–iii, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
45/abstract.html.

Anonymous:2003:LEf

[Ano03-38] Anonymous. Lists of Editors. Linear Algebra and its Applications, 365(C):ii–iii, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
53/abstract.html.

Anonymous:2003:LEg

[Ano03-39] Anonymous. Lists of Editors. Linear Algebra and its Applications, 368(C):ii–iii, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
25/abstract.html.

Anonymous:2003:LEh

[Ano03-40] Anonymous. Lists of editors. Linear Algebra and its Applications, 371(1):ii–iii, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2003:LEi

[Ano03-41] Anonymous. Lists of editors. Linear Algebra and its Applications, 372(1):ii– iii, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2003:P

[Ano03-42] Anonymous. Preface. Linear Algebra and its Applications, 358(1–3):1–3, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/207/25/
26/abstract.html.

Anonymous:2003:PLA

[Ano03-43] Anonymous. Preface: Linear algebra methods in representation theory. Linear Algebra and its Applications, 365(C):1–2, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
51/abstract.html.

Anonymous:2003:PSIa

[Ano03-44] Anonymous. Preface: Special issue on nonnegative matrices, M-matrices and their generalizations. Linear Algebra and its Applications, 363(C):1, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
26/abstract.html.

Anonymous:2003:PSIb

[Ano03-45] Anonymous. Preface: Special issue on “Structured Matrices: Analysis, Algorithms and Applications”. Linear Algebra and its Applications, 366(C):1–3, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
26/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Anonymous:2003:PNC

[Ano03-46] Anonymous. Preface: The Ninth Conference of the International Linear Algebra Society, Haifa, Israel 2001. Linear Algebra and its Applications, 361(C):1, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
26/abstract.html.

Anonymous:2003:RV

[Ano03-47] Anonymous. Referees for volumes 341–360. Linear Algebra and its Applications, 360(C):285–288, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
48/abstract.html.

Anonymous:2004:AIa

[Ano04a]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 376(1):305, January 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIb

[Ano04b]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 377(1):297, January 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIc

[Ano04c]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 378(1):293, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AId

[Ano04d]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 379(1):537– 538, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIe

[Ano04e]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 381(1):281, April 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIf

[Ano04f]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 382(1):279, May 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIg

[Ano04g]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 383(1):247, May 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIh

[Ano04h]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 384(1):243, June 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIi

[Ano04i]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 385(1):475– 476, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIj

[Ano04j]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 386(1):407– 408, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIk

[Ano04k]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 387(1):369, August 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIl

[Ano04l]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 388(1):399–400, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIm

[Ano04m]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 389(1):381–382, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIn

[Ano04n]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 390(1):385– 386, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIo

[Ano04o]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 391(1):287, November 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIp

[Ano04p]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 392(1):323, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIq

[Ano04q]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 393(1):431– 432, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIVa

[Ano04r] Anonymous. Author index to volume 380. Linear Algebra and its Applications, 380(1): 271, March 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:AIVb

[Ano04s] Anonymous. Author index to volumes 361–380. Linear Algebra and its Applications, 380(1):281–297, March 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:CPSc

[Ano04t] Anonymous. Call for papers: Special issue devoted to the 12th ILAS Conference. Linear Algebra and its Applications, 392(1):321, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:CPSa

[Ano04u] Anonymous. Call for papers: Special issue in honor of Pauline van den Driessche. Linear Algebra and its Applications, 378(1):289, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:CPSb

[Ano04v] Anonymous. Call for papers: Special issue in honor of Professor Ingram Olkin. Linear Algebra and its Applications, 378(1):291–292, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBa

[Ano04w] Anonymous. Editorial board. Linear Algebra and its Applications, 377(1):ii–iii, January 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBb

[Ano04x] Anonymous. Editorial board. Linear Algebra and its Applications, 379(1):ii–iii, March 1, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBc

[Ano04y] Anonymous. Editorial board. Linear Algebra and its Applications, 380(1):ii–iii, March 15, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBd

[Ano04z] Anonymous. Editorial board. Linear Algebra and its Applications, 381(1):ii–iii, April 1, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBe

[Ano04-27] Anonymous. Editorial board. Linear Algebra and its Applications, 382(1):ii–iii, May 1, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBf

[Ano04-28] Anonymous. Editorial board. Linear Algebra and its Applications, 383(1):ii–iii, May 15, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBg

[Ano04-29] Anonymous. Editorial board. Linear Algebra and its Applications, 384(1):ii–iii, June 1, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBh

[Ano04-30] Anonymous. Editorial board. Linear Algebra and its Applications, 385(1):ii–iii, July 1, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBi

[Ano04-31] Anonymous. Editorial board. Linear Algebra and its Applications, 386(1):ii–iii, July 15, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBj

[Ano04-32] Anonymous. Editorial board. Linear Algebra and its Applications, 387(1):ii–iii, August 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBk

[Ano04-33] Anonymous. Editorial board. Linear Algebra and its Applications, 388(1):ii–iii, September 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBl

[Ano04-34] Anonymous. Editorial board. Linear Algebra and its Applications, 389(1):ii–iii, September 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBm

[Ano04-35] Anonymous. Editorial board. Linear Algebra and its Applications, 390(1):ii–iii, October 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBn

[Ano04-36] Anonymous. Editorial board. Linear Algebra and its Applications, 391(1):ii–iii, November 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBo

[Ano04-37] Anonymous. Editorial board. Linear Algebra and its Applications, 392(1):ii–iii, November 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EBp

[Ano04-38] Anonymous. Editorial board. Linear Algebra and its Applications, 393(1):ii–iii, December 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:EN

[Ano04-39] Anonymous. Editorial note. Linear Algebra and its Applications, 386(1):xii–xiii, July 15, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:LEa

[Ano04-40] Anonymous. Lists of Editors. Linear Algebra and its Applications, 376(1):ii– iii, January 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:LEb

[Ano04-41] Anonymous. Lists of Editors. Linear Algebra and its Applications, 378(1):ii–iii, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:P

[Ano04-42] Anonymous. Photo. Linear Algebra and its Applications, 379(1):xiv, March 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:PHI

[Ano04-43]  Anonymous.  Photo for

HR issue. Linear Algebra and its Applications, 383(1): xii, May 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:PPL

[Ano04-44] Anonymous. Photo of Peter Lancaster. Linear Algebra and its Applications, 385 (1):xii, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2004:RV

[Ano04-45] Anonymous. Referees for volumes 361–380. Linear Algebra and its Applications, 380 (1):275–279, March 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIa

[Ano05a]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 394(1):353– 354, January 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIb

[Ano05b]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 395(1):371– 372, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIc

[Ano05c]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 396(1):391– 392, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AId

[Ano05d]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 397(1):389– 390, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIe

[Ano05e]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 398(1):245, March 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIf

[Ano05f]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 399(1):299– 300, April 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIg

[Ano05g]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 400(1):351– 352, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIh

[Ano05h]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 401(1):483– 484, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIi

[Ano05i]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 402(1):403– 404, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIj

[Ano05j]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 403(1):419– 420, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIk

[Ano05k]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 404(1):403– 404, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIl

[Ano05l]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 405(1):327– 328, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIm

[Ano05m]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 410(1):288–289, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIn

[Ano05n]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 411(1):390, December 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:AIV

[Ano05o] Anonymous. Author index to volumes 301–400. Linear Algebra and its Applications, 400(1):367–430, May 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:CPI

[Ano05p] Anonymous. Call for papers: 13th ILAS Conference. Linear Algebra and its Applications, 400(1):349, May 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:CPSa

[Ano05q] Anonymous. Call for papers: Special issue in honor of Friedrich L. Bauer. Linear Algebra and its Applications, 394(1):351, January 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:CPSb

[Ano05r] Anonymous. Call for papers: Special Issue in honor of Roger Horn. Linear Algebra and its Applications, 404 (1):401, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:EBa

[Ano05s] Anonymous. Editorial board. Linear Algebra and its Applications, 394(1):ii–iii, January 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:EBb

[Ano05t] Anonymous. Editorial board. Linear Algebra and its Applications, 395(1):ii–iii, January 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:EBc

[Ano05u] Anonymous. Editorial board. Linear Algebra and its Applications, 396(1):ii–iii, February 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:EBd

[Ano05v] Anonymous. Editorial board. Linear Algebra and its Applications, 397(1):ii–iii, March 1, 2005. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:EBe

[Ano05w] Anonymous. Editorial board. Linear Algebra and its Applications, 398(1):ii–iii, March 15, 2005. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:EBf

[Ano05x] Anonymous. Editorial board. Linear Algebra and its Applications, 399(1):ii–iii, April 1, 2005. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:EBg

[Ano05y] Anonymous. Editorial board. Linear Algebra and its Applications, 400(1):ii–iii, May 1, 2005. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:EBh

[Ano05z] Anonymous. Editorial board. Linear Algebra and its Applications, 401(1):ii–iii, May 15, 2005. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:EBi

[Ano05-27] Anonymous. Editorial board. Linear Algebra and its Applications, 402(1):ii–iii, June 1, 2005. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:EBj

[Ano05-28] Anonymous. Editorial board. Linear Algebra and its Applications, 403(1):ii–iii, July 1, 2005. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:EBk

[Ano05-29] Anonymous. Editorial board. Linear Algebra and its Applications, 404(1):ii–iii, July 15, 2005. CODEN LAA-PAW. ISSN 0024-3795

Anonymous:2005:EBl

[Ano05-30] Anonymous. Editorial board. Linear Algebra and its Applications, 405(1):ii–iii, August 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:MRP

[Ano05-31] Anonymous. Main research publications of Graciano de Oliveira. Linear Algebra and its Applications, 401(1): 5–9, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:RV

[Ano05-32] Anonymous. Referees for volumes 381–400. Linear Algebra and its Applications, 400 (1):359–363, May 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:SID

[Ano05-33]  Anonymous.  Special Is-

sue devoted to papers presented at the Aveiro Workshop on Graph Spectra, 10– 12 April 2006. Linear Algebra and its Applications, 405 (1):325, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2005:SIP

[Ano05-34] Anonymous. Special Issues and Proceedings in volumes 301–400. Linear Algebra and its Applications, 400(1):365– 366, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2006:AIa

[Ano06a] Anonymous. Author index. Linear Algebra and its Applications, 412(2–3):549–550, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2006:AIb

[Ano06b]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 413(2–3):619– 620, March 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2006:AIc

[Ano06c]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 414(2–3):670– 671, April 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2006:AId

[Ano06d]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 415(2–3):576– 577, June 1, 2006. CO-

Anonymous:2006:AIe

[Ano06e]  Anonymous.  Author in-

dex. Linear Algebra and its Applications, 416(2–3):1102– 1104, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2006:CPSa

[Ano06f] Anonymous. Call for papers: Special Issue on Linear and Nonlinear Models and Algorithms in Intensity-Modulated Radiation Therapy (IMRT). Linear Algebra and its Applications, 412(1): 75–76, January 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2006:CPSb

[Ano06g] Anonymous. Call for papers: Special Issue on Structured Matrices. Linear Algebra and its Applications, 414 (1):406, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2006:CVF

[Ano06h]  Anonymous.  Curriculum

vitae of Friedrich Ludwig Bauer. Linear Algebra and its Applications, 417(2–3):ii– iii, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2006:LEa

[Ano06i] Anonymous. Lists of editors. Linear Algebra and its Applications, 418(1):ii– iii, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2006:LEb

[Ano06j] Anonymous. Lists of editors. Linear Algebra and its Applications, 418(2–3):ii– iii, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2006:LEc

[Ano06k] Anonymous. Lists of editors. Linear Algebra and its Applications, 419(1):ii–iii, November 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2006:LEd

[Ano06l] Anonymous. Lists of editors. Linear Algebra and its Applications, 419(2–3):ii– iii, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:GP

[Ano07a] Anonymous. Group photo. Linear Algebra and its Applications, 423(1):ii–iii, May 1, 2007. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:LEa

[Ano07b] Anonymous. Lists of editors. Linear Algebra and its Applications, 420(1):ii– iii, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:LEb

[Ano07c] Anonymous. Lists of editors. Linear Algebra and its Applications, 420(2–3):ii– iii, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:LEc

[Ano07d] Anonymous. Lists of editors. Linear Algebra and its Applications, 421(2–3):ii– iii, March 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:LEd

[Ano07e] Anonymous. Lists of editors. Linear Algebra and its Applications, 422(1):ii– iii, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:LEe

[Ano07f] Anonymous. Lists of editors. Linear Algebra and its Applications, 422(2–3):ii– iii, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:LEf

[Ano07g] Anonymous. Lists of editors. Linear Algebra and its Applications, 423(2–3):ii– iii, June 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:LEg

[Ano07h] Anonymous. Lists of editors. Linear Algebra and its Applications, 424(2–3):ii– iii, July 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:LEh

[Ano07i] Anonymous. Lists of editors. Linear Algebra and its Applications, 425(1):ii– iii, August 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:LEi

[Ano07j] Anonymous. Lists of editors. Linear Algebra and its Applications, 426(1):ii– iii, October 1, 2007. CO-DEN LAAPAW. ISSN 0024-

Anonymous:2007:LEj

[Ano07k] Anonymous. Lists of editors. Linear Algebra and its Applications, 426(2–3):ii– iii, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:LEk

[Ano07l] Anonymous. Lists of editors. Linear Algebra and its Applications, 427(1):ii–iii, November 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:LEl

[Ano07m] Anonymous. Lists of editors. Linear Algebra and its Applications, 427(2–3):ii– iii, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:PMF

[Ano07n]  Anonymous.  Photo of

Miroslav Fiedler. Linear Algebra and its Applications, 421(2–3):xii, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:PPF

[Ano07o] Anonymous. Photo of Paul Fuhrmann. Linear Algebra and its Applications, 425(2– 3):ii–iii, September 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:PRHb

[Ano07p] Anonymous. Photo of Roger Horn. Linear Algebra and its Applications, 424(1):ii– iii, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:Pa

[Ano07q] Anonymous. Preface. Linear Algebra and its Applications, 423(1):xi, May 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:Pb

[Ano07r] Anonymous. Preface. Linear Algebra and its Applications, 425(2–3):x, September 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:PIC

[Ano07s] Anonymous. Preface to the 12th ILAS Conference Proceedings, Regina 2005. Linear Algebra and its Applications, 421(1):ii–iii, February 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:PRHa

[Ano07t] Anonymous. Publications of Roger Horn. Linear Algebra and its Applications, 424 (1):3–7, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2007:SID

[Ano07u] Anonymous. Special issue dedicated to Roger Horn. Linear Algebra and its Applications, 424(1):xii, July 1, 2007. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:GP

[Ano08a] Anonymous. Group photo. Linear Algebra and its Applications, 429(11–12):x, December 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEa

[Ano08b] Anonymous. Lists of editors. Linear Algebra and its Applications, 428(1):ii– iii, January 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEb

[Ano08c] Anonymous. Lists of editors. Linear Algebra and its Applications, 428(2–3):ii– iii, January 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEc

[Ano08d] Anonymous. Lists of editors. Linear Algebra and its Applications, 428(4):ii–iii, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEd

[Ano08e] Anonymous. Lists of editors. Linear Algebra and its Applications, 428(5–6):ii– iii, March 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEe

[Ano08f] Anonymous. Lists of editors. Linear Algebra and its Applications, 428(7):ii– iii, April 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEf

[Ano08g] Anonymous. Lists of editors. Linear Algebra and its Applications, 428(8–9):ii– iii, April 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEg

[Ano08h] Anonymous. Lists of editors. Linear Algebra and its Applications, 428(10):ii– iii, May 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEh

[Ano08i] Anonymous. Lists of editors. Linear Algebra and its Applications, 428(11–12): ii–iii, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEi

[Ano08j] Anonymous. Lists of editors. Linear Algebra and its Applications, 429(1):ii– iii, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEj

[Ano08k] Anonymous. Lists of editors. Linear Algebra and its Applications, 429(2–3):ii– iii, July 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEk

[Ano08l] Anonymous. Lists of editors. Linear Algebra and its Applications, 429(4):ii– iii, August 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEl

[Ano08m] Anonymous. Lists of editors. Linear Algebra and its Applications, 429(5–6):ii– iii, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEm

[Ano08n] Anonymous. Lists of editors. Linear Algebra and its Applications, 429(7):ii– iii, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEn

[Ano08o] Anonymous. Lists of editors. Linear Algebra and its Applications, 429(8–9):ii– iii, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEo

[Ano08p] Anonymous. Lists of editors. Linear Algebra and its Applications, 429(10):ii– iii, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2008:LEp

[Ano08q] Anonymous. Lists of editors. Linear Algebra and its Applications, 429(11–12): ii–iii, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-

Anonymous:2008:PRV

[Ano08r]  Anonymous.  Photo of

Richard S. Varga. Linear Algebra and its Applications, 429(10):x, November 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:LEa

[Ano09a] Anonymous. Lists of editors. Linear Algebra and its Applications, 430(1):ii– iii, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:LEb

[Ano09b] Anonymous. Lists of editors. Linear Algebra and its Applications, 430(2–3):ii– iii, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:LEc

[Ano09c] Anonymous. Lists of editors. Linear Algebra and its Applications, 430(4):ii–iii, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:LEd

[Ano09d] Anonymous. Lists of editors. Linear Algebra and its Applications, 430(7):ii– iii, April 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:LEe

[Ano09e] Anonymous. Lists of editors. Linear Algebra and its Applications, 430(8–9):ii– iii, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:LEf

[Ano09f] Anonymous. Lists of editors. Linear Algebra and its Applications, 430(11–12): ii–iii, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:LEg

[Ano09g] Anonymous. Lists of editors. Linear Algebra and its Applications, 431(1–2):ii– iii, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:LEh

[Ano09h] Anonymous. Lists of editors. Linear Algebra and its Applications, 431(3–4): ix, July 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:LEi

[Ano09i] Anonymous. Lists of editors. Linear Algebra and its Applications, 431(5–7):ii– iii, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:LEj

[Ano09j] Anonymous. Lists of editors. Linear Algebra and its Applications, 431(8):ii–iii, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:LEk

[Ano09k] Anonymous. Lists of editors. Linear Algebra and its Applications, 431(9):ii– iii, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:LEl

[Ano09l] Anonymous. Lists of editors. Linear Algebra and its Applications, 431(10):ii– iii, October 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:LEm

[Ano09m] Anonymous. Lists of editors. Linear Algebra and its Applications, 431(11):ii– iii, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:PHV

[Ano09n] Anonymous. Photo of Henk van der Vorst. Linear Algebra and its Applications, 431(3– 4):ii–iii, July 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:PTJ

[Ano09o]  Anonymous.  Photo of

Thomas J. Laffey. Linear Algebra and its Applications, 430(7):ix, April 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:PIC

[Ano09p] Anonymous. Preface to the 14th ILAS Conference Proceedings Shanghai 2007. Linear Algebra and its Applications, 430(5–6):ii–iii, March 1, 2009. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:PPI

[Ano09q] Anonymous. Preface to the Proceedings of the 16th International Workshop on Matrices and Statistics, Windsor 2007. Linear Algebra and its Applications, 430(10): ii–iii, May 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Anonymous:2009:SBS

[Ano09r] Anonymous. Short biography of Shmuel Friedland for his special LAA volume. Linear Algebra and its Applications, 431(12):ii– iii, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Antoulas:2007:CVP

[Ant07] A. C. Antoulas. Curriculum vitae P. A. Fuhrmann. Linear Algebra and its Applications, 425(2–3):220–225, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arlinski:2003:SM

[AP03a] Yu. M. Arlinski and A. B. Popov. On sectorial matrices. Linear Algebra and its Applications, 370(1):133– 146, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Atzmon:2003:RRF

[AP03b] Aharon Atzmon and Allan Pinkus. Rank restricting functions. Linear Algebra and its Applications, 372(1):305– 323, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arapura:2004:CIS

[AP04] Donu Arapura and Chris Peterson. The common invariant subspace problem: an approach via Gröbner bases. Linear Algebra and its Applications, 384(1):1– 7, June 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Aizpuru:2006:LSS

[APESS06] A. Aizpuru, C. Pérez-Eslava, and J. B. Seoane-Sepúlveda. Linear structure of sets of divergent sequences and series. Linear Algebra and its Applications, 418(2–3):595–598, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Akbari:2006:CGS

[AR06a] S. Akbari and P. Raja. Commuting graphs of some subsets in simple rings. Linear Algebra and its Applications, 416(2–3):1038–1047, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arambasic:2006:SNE

[AR06b] Ljiljana Arambašić and Rajna Rajić. On some norm equalities in pre-Hilbert C^{∗}- modules. Linear Algebra and its Applications, 414(1):19– 28, April 1, 2006. CO-

Arenas:2009:AAB

[Are09] Manuel Arenas. An algorithm for associative bilinear forms. Linear Algebra and its Applications, 430(1):286– 295, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arndt:2004:FDE

[Arn04] Holger Arndt. On finite dimension exchange algorithms. Linear Algebra and its Applications, 380(1):73– 93, March 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arponen:2003:MAP

[Arp03]  Teijo Arponen.  A ma-

trix approach to polynomials. Linear Algebra and its Applications, 359(1–3): 181–196, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
40/abstract.html.

Arponen:2005:MAP

[Arp05] Teijo Arponen. Matrix approach to polynomials 2. Linear Algebra and its Applications, 394(1):257–276, January 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2006:BCM

[ARR06] D. Alpay, A. C. M. Ran, and L. Rodman. Basic classes of matrices with respect to quaternionic indefinite inner product spaces. Linear Algebra and its Applications, 416 (2–3):242–269, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Arvanitoyeorgos:2008:HGF

[Arv08] Andreas Arvanitoyeorgos. Homogeneous geodesics in the flag manifold SO(2l + 1)/U(l − m) × SO(2m + 1). Linear Algebra and its Applications, 428(4):1117–1126, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Antoulas:2001:LLI

[AS01] A. C. Antoulas and D. C. Sorensen. Lyapunov, Lanczos, and inertia. Linear Algebra and its Applications, 326(1–3):137–150, March 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/148/24/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/148/24/32/article.
pdf.

Arnold:2002:LRA

[AS02] Bernhard F. Arnold and Peter Stahlecker. Linear regression analysis using the relative squared error. Linear Algebra and its Applications, 354(1–3):3–20, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
29/abstract.html.

Aujla:2003:WMI

[AS03] Jaspal Singh Aujla and Fernando C. Silva. Weak majorization inequalities and convex functions. Linear Algebra and its Applications, 369(C):217–233, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
38/abstract.html.

Aslaksen:2009:GMA

[AS09] Helmer Aslaksen and Arne B. Sletsjøe. Generators of matrix algebras in dimension 2 and 3. Linear Algebra and its Applications, 430(1): 1–6, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ash:2005:NVA

[Ash05] Avner Ash. Non-vanishing of alternants. Linear Algebra and its Applications, 411(1):

348–355, December 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Adams:2008:QOD

[ASK^{+}08] Sarah Spence Adams, Jennifer Seberry, Nathaniel Karst, Jonathan Pollack, and Tadeusz A. Wysocki. Quaternion orthogonal designs from complex companion designs. Linear Algebra and its Applications, 428(4):1056–1071, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Aguiar:2008:SVP

[ASX08] Pedro M. Q. Aguiar, Marko Stošić, and João Xavier. On singular values of partially prescribed matrices. Linear Algebra and its Applications, 429(8–9):2136–2145, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2000:ITS

[AT00] D. Alpay and E. Tsekanovskii. Interpolation theory in sectorial Stieltjes classes and explicit system solutions. Linear Algebra and its Applications, 314(1–3):91–136, July 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/131/25/

28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/131/25/28/article.
pdf.

Aizpuru:2007:LBW

[AT07] A. Aizpuru and M. Tamayo. Linear bijections which preserve the diameter of vectorvalued maps. Linear Algebra and its Applications, 424 (2–3):371–377, July 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Avelli:2008:AMS

[AT08] D. Napp Avelli and H. L. Trentelman. Algorithms for multidimensional spectral factorization and sum of squares. Linear Algebra and its Applications, 429(5–6):1114–1134, September 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Araujo:2009:SPM

[AT09] C. Mendes Araújo and Juan R. Torregrosa. Sign pattern matrices that admit M-, N-, P - or inverse Mmatrices. Linear Algebra and its Applications, 431(5– 7):724–731, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Araujo:2003:MCP

[ATU03] C. Mendes Araújo, Juan R. Torregrosa, and Ana M. Urbano. N-matrix completion problem. Linear Algebra and its Applications, 372(1):111– 125, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Araujo:2004:MCP

[ATU04] C. Mendes Araújo, Juan R. Torregrosa, and Ana M. Urbano. The N-matrix completion problem under digraphs assumptions. Linear Algebra and its Applications, 380 (1):213–225, March 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Araujo:2005:DNM

[ATU05a] C. Mendes Araújo, Juan R. Torregrosa, and Ana M. Urbano. The doubly negative matrix completion problem. Linear Algebra and its Applications, 401(1):295– 306, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Araujo:2005:SMC

[ATU05b] C. Mendes Araújo, Juan R. Torregrosa, and Ana M. Urbano. The symmetric Nmatrix completion problem. Linear Algebra and its Applications, 406(1):235–252, September 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Araujo:2006:TNC

[ATU06] C. Mendes Araújo, Juan R. Torregrosa, and Ana M. Urbano. Totally nonpositive completions on partial matrices. Linear Algebra and its Applications, 413(2–3):403– 424, March 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2007:CFB

[ATV07] Daniel Alpay, Olga Timoshenko, and Dan Volok. Carathéodory functions in the Banach space setting. Linear Algebra and its Applications, 425(2–3):700–713, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alpay:2009:CFI

[ATV09] Daniel Alpay, Olga Timoshenko, and Dan Volok. Carathéodory–Fejér interpolation in locally convex topological vector spaces. Linear Algebra and its Applications, 431(8):1257–1266, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Audenaert:2006:NCI

[Aud06] Koenraad M. R. Audenaert. A norm compression inequality for block partitioned positive semidefinite matrices. Linear Algebra and its Applications, 413(1):155–176, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Audenaert:2007:SVI

[Aud07] Koenraad M. R. Audenaert. A singular value inequality for Heinz means. Linear Algebra and its Applications, 422(1):279–283, April 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Audenaert:2008:NCI

[Aud08] Koenraad M. R. Audenaert. On a norm compression inequality for 2 × N partitioned block matrices. Linear Algebra and its Applications, 428(4):781–795, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Audenaert:2009:LTI

[Aud09a] Koenraad M. R. Audenaert. A Lieb–Thirring inequality for singular values. Linear Algebra and its Applications, 430(11–12):3053– 3057, June 1, 2009. CO-

Audenaert:2009:NNP

[Aud09b] Koenraad M. R. Audenaert. A note on the p → q norms of 2-positive maps. Linear Algebra and its Applications, 430(4):1436–1440, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Aujla:2000:OI

[Auj00] Jaspal Singh Aujla. On an operator inequality. Linear Algebra and its Applications, 310(1–3):43–47, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/127/25/30/article.
pdf.

Aujla:2003:SNI

[Auj03] Jaspal Singh Aujla. Some norm inequalities for completely monotone functions — II. Linear Algebra and its Applications, 359(1–3): 59–65, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
30/abstract.html.

Aigner:2004:IP

[AvdH04] Martin Aigner and Hein van der Holst. Interlace polynomials. Linear Algebra and its Applications, 377(1):11– 30, January 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Antoniou:2005:NCM

[AVV05] E. N. Antoniou, A. I. G. Vardulakis, and S. Vologiannidis. Numerical computation of minimal polynomial bases: A generalized resultant approach. Linear Algebra and its Applications, 405(1):264– 278, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Alfakih:2002:TTE

[AW02] Abdo Y. Alfakih and Henry Wolkowicz. Two theorems on Euclidean distance matrices and Gale transform. Linear Algebra and its Applications, 340(1–3):149–154, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/172/27/
                                        37/abstract.html.

Ando:2003:IAT

[AY03] T. Ando and T. Yamazaki. The iterated Aluthge transforms of a 2-by-2 matrix converge. Linear Algebra and its Applications, 375(1):299– 309, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Amidou:2008:AUC

[AY08] Morou Amidou and Ihsen Yengui. An algorithm for unimodular completion over Laurent polynomial rings. Linear Algebra and its Applications, 429(7):1687–1698, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Au-Yeung:2006:PIQ

[AYC06] Yik-Hoi Au-Yeung and Che-Man Cheng. On the pure imaginary quaternionic solutions of the Hurwitz matrix equations. Linear Algebra and its Applications, 419(2–3):630–642, December 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bader:2008:QGI

[Bad08] Markus Bader. Quivers, geometric invariant theory, and moduli of linear dynamical systems. Linear Algebra and its Applications, 428(11–12): 2424–2454, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Burer:2009:DBD

[BAD09] Samuel Burer, Kurt M. Anstreicher, and Mirjam Dür. The difference between 5 × 5 doubly nonnegative and completely positive matrices. Linear Algebra and its Applications, 431(9): 1539–1552, October 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See note [DA10].

Baes:2007:CDP

[Bae07]  Michel Baes.  Convexity

and differentiability properties of spectral functions and spectral mappings on Euclidean Jordan algebras. Linear Algebra and its Applications, 422(2–3):664–700, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bagchi:2006:OCD

[Bag06] Sunanda Bagchi. On the optimality of a class of designs with three concurrences. Linear Algebra and its Applications, 417(1):8– 30, August 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baglama:2008:ABH

[Bag08]  James Baglama.  Aug-

mented Block Householder Arnoldi Method. Linear Algebra and its Applications, 429(10):2315–2334, November 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bras-Amoros:2006:PNS

[BAGS06] Maria Bras-Amorós and Pedro A. García-Sánchez. Patterns on numerical semigroups. Linear Algebra and its Applications, 414(2–3): 652–669, April 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bai:2003:CGC

[Bai03] Hua Bai. On the critical group of the n-cube. Linear Algebra and its Applications, 369(C):251–261, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
40/abstract.html.

Bai:2008:CPP

[Bai08] Zheng-Jian Bai. Constructing the physical parameters of a damped vibrating system from eigendata. Linear Algebra and its Applications, 428(2–3):625–656, January 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2004:EDE

[Bak04a] Jerzy K. Baksalary. An elementary development of the equation characterizing best linear unbiased estimators. Linear Algebra and its Applications, 388(1):3–6, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2004:NAC

[Bak04b] Jerzy K. Baksalary. A new approach to the concept of a strong unified-least-squares matrix. Linear Algebra and its Applications, 388(1):7–15, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2004:ILC

[Bak04c] Oskar Maria Baksalary. Idempotency of linear combinations of three idempotent matrices, two of which are disjoint. Linear Algebra and its Applications, 388(1):67– 78, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ballico:2000:RFB

[Bal00] E. Ballico. Rank factorization and bordering of regular matrices over commutative rings. Linear Algebra and its Applications, 305(1–3): 187–190, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/120/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/120/25/38/article.
pdf.

Balser:2003:LSN

[Bal03] Werner Balser. Linear systems with nilpotent leading term. Linear Algebra and its Applications, 369(C): 145–152, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
32/abstract.html.

Balakrishnan:2004:EG

[Bal04] R. Balakrishnan. The energy of a graph. Linear Algebra and its Applications, 387 (1):287–295, August 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Balaji:2007:IBL

[Bal07] R. Balaji. On an interconnection between the Lipschitz continuity of the solution map and the positive principal minor property in linear complementarity problems over Euclidean Jordan algebras. Linear Algebra and its Applications, 426(1):83– 95, October 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bai:2009:CSL

[BAL09] Ruipu Bai, Hongwei An, and Zhenheng Li. Centroid structures of n-Lie algebras. Linear Algebra and its Applications, 430(1):229– 240, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:2000:MPI

[Bap00] R. B. Bapat. Moore–Penrose inverse of set inclusion matrices. Linear Algebra and its Applications, 318(1–3): 35–44, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/137/24/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/137/24/26/article.
                                        pdf.

Bapat:2003:OIJ

[Bap03a] R. B. Bapat. Outer inverses: Jacobi type identities and nullities of submatrices. Linear Algebra and its Applications, 361 (C):107–120, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/210/25/
                                        34/abstract.html.

Bapat:2003:PET

[Bap03b] R. B. Bapat. Perron eigenvector of the Tsetlin matrix. Linear Algebra and its Applications, 363(C):3–16, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
27/abstract.html.

Bapat:2005:DMLb

[Bap05]  R. B. Bapat.  Distance

matrix and Laplacian of a tree with attached graphs. Linear Algebra and its Applications, 411(1):295–308, December 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:2006:DDM

[Bap06] R. B. Bapat. Determinant of the distance matrix of a tree with matrix weights. Linear Algebra and its Applications, 416(1):2–7, July 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barioli:2001:CDE

[Bar01] Francesco Barioli. Chains of dog-ears for completely positive matrices. Linear Algebra and its Applications, 330 (1–3):49–66, June 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/155/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/155/25/30/article.
pdf.

Barraa:2002:CGD

[Bar02] M. Barraa. Convexoid and generalized derivations. Linear Algebra and its Applications, 350(1–3):289–292, July 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barnes:2005:RSP

[Bar05] Earl R. Barnes. Review of Selected Papers of Alan Hoffman with Commentary, edited by Charles Micchelli, World Scientific, Singapore, xliv + 446 pages. Linear Algebra and its Applications, 394(1):347–350, January 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barabanov:2008:ABE

[Bar08] N. E. Barabanov. Asymptotic behavior of extremal solutions and structure of extremal norms of linear differential inclusions of order three. Linear Algebra and its Applications, 428(10):2357– 2367, May 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Basilio:2002:IPM

[Bas02]  J. C. Basilio.  Inversion

of polynomial matrices via state-space. Linear Algebra and its Applications, 357 (1–3):259–271, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
43/abstract.html.

Batahan:2006:NEH

[Bat06] Raed S. Batahan. A new extension of Hermite matrix polynomials and its applications. Linear Algebra and its Applications, 419(1):82– 92, November 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bauch:2003:AAN

[Bau03] Manfred J. Bauch. Another algorithm for nonnegative matrices. Linear Algebra and its Applications, 365(C):3–12, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
40/abstract.html.

Bazan:2003:EAS

[Baz03]  F. S. V. Bazán.  Error

analysis of signal zeros: a projected companion matrix approach. Linear Algebra and its Applications, 369(C):

153–167, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
33/abstract.html.

Baksalary:2000:ILC

[BB00] Jerzy K. Baksalary and Oskar Maria Baksalary. Idempotency of linear combinations of two idempotent matrices. Linear Algebra and its Applications, 321(1–3): 3–7, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/25/article.
pdf.

Baksalary:2002:CP

[BB02a] Jerzy K. Baksalary and Oskar Maria Baksalary. Commutativity of projectors. Linear Algebra and its Applications, 341(1–3):129–142, January 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
37/abstract.html.

Ball:2002:BIP

[BB02b] Joseph A. Ball and Vladimir Bolotnikov. On a bitangential interpolation problem for contractive-valued functions on the unit ball. Linear Algebra and its Applications, 353(1–3):107–147, September 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barioli:2003:MCR

[BB03a] F. Barioli and A. Berman. The maximal cp-rank of rank k completely positive matrices. Linear Algebra and its Applications, 363 (C):17–33, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
28/abstract.html.

Bender:2003:MSO

[BB03b] Jens Bender and Klaus Bongartz. Minimal singularities in orbit closures of matrix pencils. Linear Algebra and its Applications, 365 (C):13–24, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
48/abstract.html.

Bini:2003:PFT

[BB03c] Dario A. Bini and Albrecht Böttcher. Polynomial factorization through Toeplitz matrix computations. Linear Algebra and its Applications, 366(C):25–37, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/215/25/
                                        28/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Baksalary:2004:NLCa

[BB04a] Jerzy K. Baksalary and Oskar Maria Baksalary. Nonsingularity of linear combinations of idempotent matrices. Linear Algebra and its Applications, 388(1):25– 29, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2004:LCG

[BB04b] Jerzy K. Baksalary and Oskar Maria Baksalary. On linear combinations of generalized projectors. Linear Algebra and its Applications, 388 (1):17–24, September 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2004:RBG

[BB04c] Jerzy K. Baksalary and Oskar Maria Baksalary. Relationships between generalized inverses of a matrix and generalized inverses of its rank-one-modifications. Linear Algebra and its Applications, 388(1):31–44, September 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Balasubramanian:2004:TIV

[BB04d] K. Balasubramanian and M. I. Beg. Three isomorphic vector spaces—applications to binomial identities, Urn models and order statistics. Linear Algebra and its Applications, 388(1):79– 89, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2005:IPR

[BB05a] Jerzy K. Baksalary and Oskar Maria Baksalary. An invariance property related to the reverse order law. Linear Algebra and its Applications, 410(1):64–69, November 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Balasubramanian:2005:TIV

[BB05b] K. Balasubramanian and M. I. Beg. Three isomorphic vector spaces—II. Linear Algebra and its Applications, 410(1):254–266, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berenhaut:2005:MCS

[BB05c] Kenneth S. Berenhaut and Dipankar Bandyopadhyay. Monotone convex sequences and Cholesky decomposition of symmetric Toeplitz matrices. Linear Algebra and its Applications, 403(1):75– 85, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2007:PFM

[BB07a] Jerzy K. Baksalary and Oskar Maria Baksalary. Particular formulae for the Moore–Penrose inverse of a columnwise partitioned matrix. Linear Algebra and its Applications, 421(1):16– 23, February 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2007:ILC

[BB07b] Oskar Maria Baksalary and Julio Benítez. Idempotency of linear combinations of three idempotent matrices, two of which are commuting. Linear Algebra and its Applications, 424(1):320– 337, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Balaji:2007:EDM

[BB07c] R. Balaji and R. B. Bapat. On Euclidean distance matrices. Linear Algebra and its Applications, 424(1):108– 117, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bevilacqua:2009:SKM

[BB09] Roberto Bevilacqua and Enrico Bozzo. The Sylvester– Kac matrix space. Linear Algebra and its Applications, 430(11–12):3131– 3138, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beidar:2000:JIT

[BBC00] K. I. Beidar, M. Brešar, and M. A. Chebotar. Jordan isomorphisms of triangular matrix algebras over a connected commutative ring. Linear Algebra and its Applications, 312(1–3):197–201, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/129/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/129/25/38/article.
pdf.

Bevilacqua:2005:TRS

[BBC05] Roberto Bevilacqua, Enrico Bozzo, and Gianna M. Del Corso. Transformations to rank structures by unitary similarity. Linear Algebra and its Applications, 402(1): 126–134, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Boros:2008:GVI

[BBCH08] Endre Boros, Richard A. Brualdi, Yves Crama, and A. J. Hoffman. Geršgorin variations III: On a theme of Brualdi and Varga. Linear Algebra and its Applications, 428(1):14–19, January 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barlow:2005:NSB

[BBD05] Jesse L. Barlow, Nela Bosner, and Zlatko Drmač. A new stable bidiagonal reduction algorithm. Linear Algebra and its Applications, 397(1): 35–84, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berliner:2008:RCO

[BBD^{+}08] Adam Berliner, Richard A. Brualdi, Louis Deaett, Kathleen P. Kiernan, and Michael Schroeder. Row and column orthogonal (0, 1)-matrices. Linear Algebra and its Applications, 429(11–12):2732– 2745, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baragana:2009:SPB

[BBdH09] Itziar Baragaña, M. Asunción Beitia, and Inmaculada de Hoyos. Structured perturbation of the Brunovsky form: A particular case. Linear Algebra and its Applications, 430(5–6):1613–1625, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benoit:2004:ASA

[BBFP04] Anne Benoit, L. Brenner, P. Fernandes, and B. Plateau. Aggregation of stochastic automata networks with replicas. Linear Algebra and its Applications, 386(1):111– 136, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2006:SLC

[BBG06] Jerzy K. Baksalary, Oskar Maria Baksalary, and Jürgen Groß. On some linear combinations of hypergeneralized projectors. Linear Algebra and its Applications, 413(2–3):264–273, March 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:2009:PRT

[BBG09] LeRoy B. Beasley, David E. Brown, and Alexander E. Guterman. Preserving regular tournaments and term rank-1. Linear Algebra and its Applications, 431(5–7): 926–936, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brenner:2003:IMB

[BBK03] Sheila Brenner, M. C. R. Butler, and Alastair D. King. Irreducible maps and bilinear forms. Linear Algebra and its Applications, 365 (C):99–105, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
28/abstract.html.

Baksalary:2007:GPO

[BBK07] Jerzy K. Baksalary, Oskar Maria Baksalary, and Paulina Kik. Generalizations of a property of orthogonal projectors. Linear Algebra and its Applications, 420(1): 1–8, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2003:FPS

[BBL03a] Jerzy K. Baksalary, Oskar Maria Baksalary, and Xiaoji Liu. Further properties of the star, left-star, rightstar, and minus partial orderings. Linear Algebra and its Applications, 375(1):83– 94, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-

Baksalary:2003:FRB

[BBL03b] Jerzy K. Baksalary, Oskar Maria Baksalary, and Xiaoji Liu. Further relationships between certain partial orders of matrices and their squares. Linear Algebra and its Applications, 375(1):171– 180, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2004:FPG

[BBL04] Jerzy K. Baksalary, Oskar Maria Baksalary, and Xiaoji Liu. Further properties of generalized and hypergeneralized projectors. Linear Algebra and its Applications, 389(1):295–303, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Belitskii:2005:PCP

[BBL^{+}05] Genrich Belitskii, Vitalij M. Bondarenko, Ruvim Lipyanski, Vladimir V. Plachotnik, and Vladimir V. Sergeichuk. The problems of classifying pairs of forms and local algebras with zero cube radical are wild. Linear Algebra and its Applications, 402(1): 135–142, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2008:FRG

[BBLT08] Jerzy K. Baksalary, Oskar Maria Baksalary, Xiaoji Liu, and Götz Trenkler. Further results on generalized and hypergeneralized projectors. Linear Algebra and its Applications, 429(5–6):1038– 1050, September 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bagchi:2002:STA

[BBM02] Bhaskar Bagchi, Tirthankar Bhattacharyya, and Gadadhar Misra. Some thoughts on Ando’s theorem and Parrott’s example. Linear Algebra and its Applications, 341 (1–3):357–367, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/171/27/
                                        56/abstract.html.

Benner:2007:RNM

[BBMX07] Peter Benner, Ralph Byers, Volker Mehrmann, and Hongguo Xu. A robust numerical method for the γ-iteration in H_{∞} control. Linear Algebra and its Applications, 425(2–3):548–570, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2004:NLCb

[BBÖ04] Jerzy K. Baksalary, Oskar Maria Baksalary, and Halim Özdemir. A note on linear combinations of commuting tripotent matrices. Linear Algebra and its Applications, 388(1):45–51, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bunch:2003:ANL

[BBP03] J. R. Bunch, R. C. Le Borne, and I. K. Proudler. Applying numerical linear algebra techniques to analyzing algorithms in signal processing. Linear Algebra and its Applications, 361 (C):133–146, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
36/abstract.html.

Baksalary:2002:ILC

[BBS02a] Jerzy K. Baksalary, Oskar Maria Baksalary, and George P. H. Styan. Idempotency of linear combinations of an idempotent matrix and a tripotent matrix. Linear Algebra and its Applications, 354(1–3): 21–34, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
30/abstract.html.

Baksalary:2002:POP

[BBS02b] Jerzy K. Baksalary, Oskar Maria Baksalary, and Tomasz Szulc. A property of orthogonal projectors. Linear Algebra and its Applications, 354(1–3): 35–39, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
31/abstract.html.

Baksalary:2004:PSC

[BBS04] Jerzy K. Baksalary, Oskar Maria Baksalary, and Tomasz Szulc. Properties of Schur complements in partitioned idempotent matrices. Linear Algebra and its Applications, 379(1):303– 318, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Belitskii:2006:CFM

[BBS06] Genrich Belitskii, Maxim Bershadsky, and Vladimir V. Sergeichuk. Canonical form of m-by-2-by-2 matrices over a field of characteristic other than two. Linear Algebra and its Applications, 418(1):15– 19, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bendaoud:2008:LMP

[BBS08] M. Bendaoud, A. Bourhim, and M. Sarih. Linear maps preserving the essential spectral radius. Linear Algebra and its Applications, 428(4): 1041–1045, February 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2003:RFM

[BBT03] Jerzy K. Baksalary, Oskar Maria Baksalary, and Götz Trenkler. A revisitation of formulae for the Moore– Penrose inverse of modified matrices. Linear Algebra and its Applications, 372(1):207– 224, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bart:2008:PIC

[BBVR08] Harm Bart, Jan Brandts, Paul Van Dooren, and André Ran. Preface to the 13th ILAS Conference Proceedings, Amsterdam 2006. Linear Algebra and its Applications, 429(5–6):949–950, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benanti:2000:CSM

[BC00] F. Benanti and M. G. Campanella. On the *-cocharacter sequence of 3 × 3 matrices. Linear Algebra and its Applications, 312(1–3): 101–114, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/129/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/129/25/30/article.
pdf.

Bermudez:2003:CNL

[BC03] José María Ancochea Bermúdez and Rutwig Campoamor. On characteristically nilpotent Lie algebras of type Q. Linear Algebra and its Applications, 367(C):193–212, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
39/abstract.html.

Blanes:2004:COB

[BC04a] Sergio Blanes and Fernando Casas. On the convergence and optimization of the Baker–Campbell–Hausdorff formula. Linear Algebra and its Applications, 378(1):135– 158, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brieden:2004:HEA

[BC04b] Andreas Brieden and Shawn Cokus. On the hardness of efficiently approximating maximal non-L submatrices. Linear Algebra and its Applications, 377(1):195– 205, January 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Borobia:2007:CMP

[BC07] Alberto Borobia and Roberto Canogar. Constructing matrices with prescribed offdiagonal submatrix and invariant polynomials. Linear Algebra and its Applications, 424(2–3):615–633, July 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See note [BC08b].

Bagno:2008:SBF

[BC08a] Eli Bagno and Yona Cherniavsky. Sign balance for finite groups of Lie type. Linear Algebra and its Applications, 429(1):224–233, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Borobia:2008:NCM

[BC08b] Alberto Borobia and Roberto Canogar. A note on “constructing matrices with prescribed off-diagonal submatrix and invariant polynomials”. Linear Algebra and its Applications, 429(7):1684– 1686, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See [BC07].

Barriere:2008:SH

[BCDF08] L. Barrière, F. Comellas, C. Dalfó, and M. A. Fiol. On the spectra of hypertrees. Linear Algebra and its Applications, 428(7):1499– 1510, April 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bendito:2009:CSM

[BCEG09] E. Bendito, A. Carmona, A. M. Encinas, and J. M. Gesto. Characterization of symmetric M-matrices as resistive inverses. Linear Algebra and its Applications, 430(4):1336–1349, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Banks:2005:TAD

[BCES05] T. Banks, T. Constantinescu, and Nermine El-Sissi. Tensor algebras and displacement structure. IV. Invariant kernels. Linear Algebra and its Applications, 405(1): 83–103, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Butkovic:2007:MPM

[BCG07] P. Butkovič and R. A. Cuninghame-Green. On matrix powers in max-algebra. Linear Algebra and its Applications, 421(2–3):370–381, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bru:2008:CGM

[BCGM08] R. Bru, C. Corral, I. Giménez, and J. Mas. Classes of general H-matrices. Linear Algebra and its Applications, 429(10):2358–2366, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Boman:2005:FWS

[BCPT05] Erik G. Boman, Doron Chen, Ojas Parekh, and Sivan Toledo. On factor width and symmetric H-matrices. Linear Algebra and its Applications, 405(1):239–248, August 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bell:2008:GWLa

[BCRS08a] Francis K. Bell, Dragoš Cvetković, Peter Rowlinson, and Slobodan K. Simić. Graphs for which the least eigenvalue is minimal, I. Linear Algebra and its Applications, 429(1):234–241, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bell:2008:GWLb

[BCRS08b] Francis K. Bell, Dragoš Cvetković, Peter Rowlinson, and Slobodan K. Simić. Graphs for which the least eigenvalue is minimal, II. Linear Algebra and its Applications, 429(8–9):2168–2179, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bru:2002:SPP

[BCS02] Rafael Bru, Carmen Coll, and Elena Sánchez. Structural properties of positive linear time-invariant difference-algebraic equations. Linear Algebra and its Applications, 349(1–3): 1–10, July 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bermudez:2006:ALR

[BCSV06] J. M. Ancochea Bermúdez, R. Campoamor-Stursberg, and L. García Vergnolle. Les algèbres de Lie résolubles rigides réelles ne sont pas nécessairement complètement résolubles. (french) [Resolvable real rigid Lie algebras are not necessarily complete resolvable]. Linear Algebra and its Applications, 418(2–3):657–664, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2001:KIF

[BD01a] Rajendra Bhatia and Chandler Davis. A key inequality for functions of matrices. Linear Algebra and its Applications, 323(1–3):1– 5, January 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Borobia:2001:TCP

[BD01b] Alberto Borobia and Sebastián Dormido. Three coefficients of a polynomial can determine its instability. Linear Algebra and its Applications, 338(1–3): 67–76, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/32/article.
pdf.

Bahturin:2002:GPI

[BD02a] Yuri Bahturin and Vesselin Drensky. Graded polynomial identities of matrices. Linear Algebra and its Applications, 357(1–3): 15–34, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
27/abstract.html.

Bhatia:2002:VIL

[BD02b] Rajendra Bhatia and J. A. Dias da Silva. Variation of induced linear operators. Linear Algebra and its Applications, 341(1–3): 391–402, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
59/abstract.html.

Bahturin:2003:IBM

[BD03a] Yu. A. Bahturin and V. Drensky. Identities of bilinear mappings and graded polynomial identities of matrices. Linear Algebra and its Applications, 369(C):95– 112, August 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
30/abstract.html.

Bortoletti:2003:SMA

[BD03b] Alessandro Bortoletti and Carmine Di Fiore. On a set of matrix algebras related to discrete Hartleytype transforms. Linear Algebra and its Applications, 366(C):65–85, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
30/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Brualdi:2003:MCD

[BD03c] Richard A. Brualdi and Geir Dahl. Majorizationconstrained doubly stochastic matrices. Linear Algebra and its Applications, 361 (C):75–97, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
32/abstract.html.

Brualdi:2003:MZO

[BD03d] Richard A. Brualdi and Geir Dahl. Matrices of zeros and ones with given line sums and a zero block. Linear Algebra and its Applications, 371(1): 191–207, September 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Buchholz:2004:BSK

[BD04] Peter Buchholz and Tuğrul Dayar. Block SOR for Kronecker structured representations. Linear Algebra and its Applications, 386(1):83– 109, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bento:2005:FCT

[BD05] Américo Bento and António Leal Duarte. On Fiedler’s characterization of tridiagonal matrices over arbitrary fields. Linear Algebra and its Applications, 401(1):467–481, May 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barany:2006:BPV

[BD06a] Imre Bárány and Benjamin Doerr. Balanced partitions of vector sequences. Linear Algebra and its Applications, 414(2–3):464–469, April 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Borobia:2006:TCP

[BD06b] Alberto Borobia and Sebastián Dormido. Three coefficients of a polynomial can determine its ϕ-instability. Linear Algebra and its Applications, 416(2–3):857–867, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:2007:MBO

[BD07] Richard A. Brualdi and Louis Deaett. More on the Bruhat order for (0, 1)-matrices. Linear Algebra and its Applications, 421(2–3):219–232, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bundfuss:2008:ACD

[BD08] Stefan Bundfuss and Mirjam Dür. Algorithmic copositivity detection by simplicial partition. Linear Algebra and its Applications, 428 (7):1511–1523, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bottcher:2009:BCI

[BD09a] Albrecht Böttcher and Peter Dörfler. On the best constants in inequalities of the Markov and Wirtinger types for polynomials on the half-line. Linear Algebra and its Applications, 430(4): 1057–1069, February 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Boutsidis:2009:RPN

[BD09b] Christos Boutsidis and Petros Drineas. Random projections for the nonnegative least-squares problem. Linear Algebra and its Applications, 431(5–7):760–771, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brunetti:2001:RCC

[BDD^{+}01] S. Brunetti, A. DelLungo, F. DelRistoro, A. Kuba, and M. Nivat. Reconstruction of 4- and 8-connected convex discrete sets from row and column projections. Linear Algebra and its Applications, 339(1–3):37–57, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/169/27/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/169/27/30/article.
pdf.

Benito:2003:MOG

[BDE03] Pilar Benito, Cristina Draper, and Alberto Elduque. Models of the octonions and G_{2}. Linear Algebra and its Applications, 371(1):333–359, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bini:2008:CBT

[BDG08] Dario A. Bini, Alessandra De Rossi, and Bruno Gabutti. On certain (block) Toeplitz matrices related to radial functions. Linear Algebra and its Applications, 428(2–3): 508–519, January 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beitia:2005:CJS

[BdHZ05] M. Asunción Beitia, Inmaculada de Hoyos, and Ion Zaballa. The change of the Jordan structure under one row perturbations. Linear Algebra and its Applications, 401 (1):119–134, May 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beitia:2008:CSIb

[BdHZ08a] M. Asunción Beitia, Inmaculada de Hoyos, and Ion Zaballa. The change of similarity invariants under row perturbations. Linear Algebra and its Applications, 429(5–6):1302–1333, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beitia:2008:CSIa

[BdHZ08b] M. Asunción Beitia, Inmaculada de Hoyos, and Ion Zaballa. The change of similarity invariants under row perturbations: Generic cases. Linear Algebra and its Applications, 429(2–3):482– 496, July 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bani-Domi:2008:NEI

[BDK08] Wathiq Bani-Domi and Fuad Kittaneh. Norm equalities and inequalities for operator matrices. Linear Algebra and its Applications, 429 (1):57–67, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bradley:2004:HBP

[BDKW04] Jeremy T. Bradley, Nicholas J. Dingle, William J. Knottenbelt, and Helen J. Wilson. Hypergraph-based parallel computation of passage time densities in large semi-Markov models. Linear Algebra and its Applications, 386 (1):311–334, July 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:2004:MIS

[BdL04] N. Bebiano, J. da Providência, Jr., and R. Lemos. Matrix inequalities in statistical mechanics. Linear Algebra and its Applications, 376(1):265– 273, January 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barot:2006:RII

[BdlP06] M. Barot and J. A. de la Peña. Root-induced integral quadratic forms. Linear Algebra and its Applications, 412(2–3):291–302, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barioli:2003:VSD

[BDM03] Francesco Barioli, Clorinda De Vivo, and Claudia Metelli. On vector spaces with distinguished subspaces and redundant base. Linear Algebra and its Applications, 374(1):107–126, November 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:2007:EBT

[BdPK07] Natália Bebiano, João da Providência, and Alexander Kovačec. An extension of Birkhoff’s theorem with an application to determinants. Linear Algebra and its Applications, 427(1): 130–137, November 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:2008:FPB

[BdPT08] N. Bebiano, J. da Providência, and R. Teixeira. Flat portions on the boundary of the indefinite numerical range of 3×3 matrices. Linear Algebra and its Applications, 428(11– 12):2863–2879, June 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Burde:2008:NAN

[BDV08] Dietrich Burde, Karel Dekimpe, and Kim Vercammen. Novikov algebras and Novikov structures on Lie algebras. Linear Algebra and its Applications, 429(1):31–41, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Basor:2001:SID

[BE01] Estelle L. Basor and Torsten Ehrhardt. Some identities for determinants of structured matrices. Linear Algebra and its Applications, 343– 344(1):5–19, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
28/abstract.html.

Bella:2007:BPT

[BEG^{+}07] Tom Bella, Yuli Eidelman, Israel Gohberg, Israel Koltracht, and Vadim Olshevsky. A Björck–Pereyratype algorithm for Szegő– Vandermonde matrices based on properties of unitary Hessenberg matrices. Linear Algebra and its Applications, 420(2–3):634–647, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bowers:2006:CPV

[BEH^{+}06] John Bowers, Job Evers, Leslie Hogben, Steve Shaner, Karyn Snider, and Amy Wangsness. On completion problems for various classes of P -matrices. Linear Algebra and its Applications, 413(2– 3):342–354, March 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bekjan:2004:JCT

[Bek04] Turdebek N. Bekjan. On joint convexity of trace functions. Linear Algebra and its Applications, 390(1):321– 327, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bell:2005:GRN

[Bel05] Jason P. Bell. A gap result for the norms of semigroups of matrices. Linear Algebra and its Applications, 402(1): 101–110, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benasseni:2002:CPE

[Bén02] J. Bénasséni. A complementary proof of an eigenvalue property in correspondence analysis. Linear Algebra and its Applications, 354(1– 3):49–51, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
33/abstract.html.

Benzi:2004:DPM

[Ben04] Michele Benzi. A direct projection method for Markov chains. Linear Algebra and its Applications, 386(1):27– 49, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benkovic:2005:JDA

[Ben05] Dominik Benkovič. Jordan derivations and antiderivations on triangular matrices. Linear Algebra and its Applications, 397(1):235– 244, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benasseni:2006:VIE

[Bén06] Jacques Bénasséni. A variance inequality ensuring that a pre-distance matrix is Euclidean. Linear Algebra and its Applications, 416(2–3): 365–372, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benzi:2008:SSM

[Ben08] Michele Benzi. Splittings of symmetric matrices and a question of Ortega. Linear Algebra and its Applications, 429(10):2340–2343, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benkovic:2009:BTA

[Ben09] Dominik Benkovič. Biderivations of triangular algebras. Linear Algebra and its Applications, 431(9):1587–1602, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bertram:2004:LAA

[Ber04] Wolfgang Bertram. From linear algebra via affine algebra to projective algebra. Linear Algebra and its Applications, 378(1):109–134, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bereux:2005:FDS

[Bér05] Natacha Béreux. Fast direct solvers for some complex symmetric block Toeplitz linear systems. Linear Algebra and its Applications, 404(1): 193–222, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berget:2009:SPG

[Ber09] Andrew Berget. A short proof of Gamas’s theorem. Linear Algebra and its Applications, 430(2–3):791–794, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bottcher:2001:ITL

[BES01] A. Böttcher, M. Embree, and V. I. Sokolov. Infinite Toeplitz and Laurent matrices with localized impurities. Linear Algebra and its Applications, 343–344(1): 101–118, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
31/abstract.html.

Bart:2002:LRA

[BES02] Harm Bart, Torsten Ehrhardt, and Bernd Silbermann. Logarithmic residues of analytic Banach algebra valued functions possessing a simply meromorphic inverse. Linear Algebra and its Applications, 341(1–3):327–344, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
53/abstract.html.

Bottcher:2003:LTB

[BES03] A. Böttcher, M. Embree, and V. I. Sokolov. On large Toeplitz band matrices with an uncertain block. Linear Algebra and its Applications, 366(C):87–97, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
31/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Bart:2009:TCR

[BES09] H. Bart, T. Ehrhardt, and B. Silbermann. Trace conditions for regular spectral behavior of vector-valued analytic functions. Linear Algebra and its Applications, 430(8–9):1945–1965, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bai:2001:PPL

[BF01] Zhaojun Bai and Roland W. Freund. A partial Padé-via-Lanczos method for reducedorder modeling. Linear Algebra and its Applications, 332–334(1):139–164, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/34/article.
pdf.

Barany:2005:CWU

[BF05] Imre Bárány and Komei Fukuda. A case when the union of polytopes is convex. Linear Algebra and its Applications, 397(1):381– 388, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barioli:2007:MRJ

[BF07a] Francesco Barioli and Shaun Fallat. On the minimum rank of the join of graphs and decomposable graphs. Linear Algebra and its Applications, 421(2–3):252–263, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bilardi:2007:RTC

[BF07b] Gianfranco Bilardi and Augusto Ferrante. The role of terminal cost/reward in finite-horizon discrete-time LQ optimal control. Linear Algebra and its Applications, 425(2–3):323–344, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Boocher:2008:GBR

[BF08] Adam Boocher and Bradley Froehle. On generators of bounded ratios of minors for totally positive matrices. Linear Algebra and its Applications, 428(7):1664– 1684, April 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barioli:2004:CMR

[BFH04] Francesco Barioli, Shaun Fallat, and Leslie Hogben. Computation of minimal rank and path cover number for certain graphs. Linear Algebra and its Applications, 392(1):289– 303, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barioli:2005:DBM

[BFH05] Francesco Barioli, Shaun Fallat, and Leslie Hogben. On the difference between the maximum multiplicity and path cover number for treelike graphs. Linear Algebra and its Applications, 409 (1):13–31, November 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:2008:UBM

[BFH^{+}08] Avi Berman, Shmuel Friedland, Leslie Hogben, Uriel G. Rothblum, and Bryan Shader. An upper bound for the minimum rank of a graph. Linear Algebra and its Applications, 429(7):1629–1638, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bueno:2007:CHM

[BFJ07] M. I. Bueno, S. Furtado, and C. R. Johnson. Congruence of Hermitian matrices by Hermitian matrices. Linear Algebra and its Applications, 425 (1):63–76, August 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ballantine:2005:DAZ

[BFL05] Cristina M. Ballantine, Sharon M. Frechette, and John B. Little. Determinants associated to zeta matrices of posets.

Linear Algebra and its Applications, 411(1):364–370, December 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhattacharya:2007:PDD

[BFP07] Amitava Bhattacharya, Shmuel Friedland, and Uri N. Peled. The polytope of dual degree partitions. Linear Algebra and its Applications, 426(2–3):458–461, October 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bracke:2002:PDS

[BFPW02] Martin Bracke, Sven Feldmann, and Dieter Prätzel-Wolters. Parameter depending state space descriptions of index-2-matrix polynomials. Linear Algebra and its Applications, 347 (1–3):59–80, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
30/abstract.html.

Benitez:2007:OM

[BFS07] Carlos Benítez, Manuel Fernández, and María L. Soriano. Orthogonality of matrices. Linear Algebra and its Applications, 422(1):155– 163, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-

Barioli:2008:AUG

[BFS08] Francesco Barioli, Shaun M. Fallat, and Ronald L. Smith. On acyclic and unicyclic graphs whose minimum rank equals the diameter. Linear Algebra and its Applications, 429(7):1568–1578, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benner:2006:SIO

[BFSV06] Peter Benner, Roland W. Freund, Danny C. Sorensen, and Andras Varga. Special issue on “Order reduction of large-scale systems”. Linear Algebra and its Applications, 415(2–3):231– 234, June 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brandts:2009:P

[BFW09] Jan Brandts, Bernd Fischer, and Andy Wathen. Preface. Linear Algebra and its Applications, 431(3–4): 299, July 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baragana:2004:HIE

[BFZ04] I. Baragaña, V. Fernández, and I. Zaballa. Hermite indices and equivalence relations. Linear Algebra and its Applications, 379(1):201– 211, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baragana:2005:HIA

[BFZ05] I. Baragaña, V. Fernández, and I. Zaballa. Hermite indices and the action of the feedback group. Linear Algebra and its Applications, 401 (1):401–427, May 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bottcher:2002:CSV

[BG02] A. Böttcher and S. M. Grudsky. Can spectral value sets of Toeplitz band matrices jump? Linear Algebra and its Applications, 351–352:99– 116, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Burke:2006:CPN

[BG06] James V. Burke and Anne Greenbaum. Characterizations of the polynomial numerical hull of degree k. Linear Algebra and its Applications, 419(1):37–47, November 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:2009:COP

[BG09a] LeRoy B. Beasley and Alexander E. Guterman. The characterization of operators preserving primitivity for matrix k-tuples. Linear Algebra and its Applications, 430 (7):1762–1777, April 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bialas:2009:SPZ

[BG09b] Stanisław Białas and Michał Góra. Some properties of zeros of polynomials with vanishing coefficients. Linear Algebra and its Applications, 430(8–9):1976–1991, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bunse-Gerstner:2006:PGC

[BGGC06] Angelika Bunse-Gerstner and Ignacio Gutiérrez-Cañas. A preconditioned GMRES for complex dense linear systems from electromagnetic wave scattering problems. Linear Algebra and its Applications, 416(1):135–147, July 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:2006:LPE

[BGJS06] LeRoy B. Beasley, Alexander E. Guterman, Young-Bae Jun, and Seok-Zun Song. Linear preservers of extremes of rank inequalities over semirings: Row and column ranks. Linear Algebra and its Applications, 413(2–3):495– 509, March 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bart:2005:SCS

[BGKR05] H. Bart, I. Gohberg, M. A. Kaashoek, and A. C. M. Ran. Schur complements and state space realizations. Linear Algebra and its Applications, 399(1):203–224, April 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2007:PIC

[BGKW07] Rajendra Bhatia, Robert Guralnick, Steve Kirkland, and Henry Wolkowicz. Preface to the 12th ILAS Conference Proceedings, Regina 2005. Linear Algebra and its Applications, 421(1):1–2, February 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barrett:2009:MRP

[BGL09] Wayne Barrett, Jason Grout, and Raphael Loewy. The minimum rank problem over the finite field of order 2: Minimum rank 3. Linear Algebra and its Applications, 430(4):890–923, February 1, 2009. CO-

Beasley:2004:LTP

[BGLS04] LeRoy B. Beasley, Alexander E. Guterman, Sang-Gu Lee, and Seok-Zun Song. Linear transformations preserving the Grassmannian over M_{\}(Z_{+}). Linear Algebra and its Applications, 393(1):39– 46, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:2005:LPZ

[BGLS05] LeRoy B. Beasley, Alexander E. Guterman, Sang-Gu Lee, and Seok-Zun Song. Linear preservers of zeros of matrix polynomials. Linear Algebra and its Applications, 401(1):325–340, May 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bini:2001:CIT

[BGM01] D. A. Bini, L. Gemignani, and B. Meini. Computations with infinite Toeplitz matrices and polynomials. Linear Algebra and its Applications, 343–344(1):21–61, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
29/abstract.html.

Bernik:2004:RTG

[BGM04] Janez Bernik, Robert Guralnick, and Mitja Mastnak. Reduction theorems for groups of matrices. Linear Algebra and its Applications, 383(1): 119–126, May 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bottcher:2008:PET

[BGM08] Albrecht Böttcher, Sergei M. Grudsky, and Egor A. Maksimenko. Pushing the envelope of the test functions in the Szegő and Avram–Parter theorems. Linear Algebra and its Applications, 429(1): 346–366, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bai:2008:IHS

[BGN08] Zhong-Zhi Bai, Gene H. Golub, and Michael K. Ng. On inexact Hermitian and skew-Hermitian splitting methods for non-Hermitian positive definite linear systems. Linear Algebra and its Applications, 428(2–3):413– 440, January 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bresar:2009:ZPD

[BGO09] Matej Brešar, Mateja Grašič, and Juana Sánchez Ortega. Zero product determined matrix algebras. Linear Algebra and its Applications, 430(5– 6):1486–1498, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barg:2000:SGV

[BGS00] Alexander Barg, Sugi Guritman, and Juriaan Simonis. Strengthening the Gilbert– Varshamov bound. Linear Algebra and its Applications, 307(1–3):119–129, March 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/33/article.
pdf.

Bondarenko:2009:PMA

[BGS09] Vitalij M. Bondarenko, Tatiana G. Gerasimova, and Vladimir V. Sergeichuk. Pairs of mutually annihilating operators. Linear Algebra and its Applications, 430(1): 86–105, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2004:CMS

[BH04] Jerzy K. Baksalary and Jan Hauke. Characterizations of minus and star orders between the squares of Hermitian matrices. Linear Algebra and its Applications, 388 (1):53–59, September 1, 2004.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brouwer:2005:EPM

[BH05] Andries E. Brouwer and Willem H. Haemers. Eigenvalues and perfect matchings. Linear Algebra and its Applications, 395(1):155– 162, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2006:RGM

[BH06] Rajendra Bhatia and John Holbrook. Riemannian geometry and matrix geometric means. Linear Algebra and its Applications, 413(2– 3):594–618, March 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brouwer:2008:ITS

[BH08a] A. E. Brouwer and W. H. Haemers. The integral trees with spectral radius 3. Linear Algebra and its Applications, 429(11–12):2710–2718, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brouwer:2008:LBL

[BH08b] Andries E. Brouwer and Willem H. Haemers. A lower bound for the Laplacian eigenvalues of a graph — proof of a conjecture by Guo. Linear Algebra and its Applications, 429(8–9):2131– 2135, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2003:EMI

[Bha03a] Rajendra Bhatia. On the exponential metric increasing property. Linear Algebra and its Applications, 375(1):211– 220, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2003:PTE

[Bha03b] Rajendra Bhatia. Partial traces and entropy inequalities. Linear Algebra and its Applications, 370(1):125– 132, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2006:IAG

[Bha06] Rajendra Bhatia. Interpolating the arithmetic–geometric mean inequality and its operator version. Linear Algebra and its Applications, 413(2– 3):355–363, March 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Belyi:2006:GRT

[BHdST06] Sergey Belyi, Seppo Hassi, Henk de Snoo, and Eduard Tsekanovskiĭ. A general realization theorem for matrixvalued Herglotz–Nevanlinna functions. Linear Algebra and its Applications, 419(2–3): 331–358, December 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2008:PST

[BHJ08] Oskar Maria Baksalary, Jan Hauke, and Charles R. Johnson. Polynomials satisfied by two linked matrices. Linear Algebra and its Applications, 429(10):2335–2339, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barrett:2005:P

[BHKL05] Wayne Barrett, Samad Hedayat, Christian Krattenthaler, and Raphael Loewy. Preface. Linear Algebra and its Applications, 411(1):1– 2, December 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Biyikoglu:2004:GLN

[BHL^{+}04] Türker Bıyıkoğlu, Wim Hordijk, Josef Leydold, Tomaž Pisanski, and Peter F. Stadler. Graph Laplacians, nodal domains, and hyperplane arrangements. Linear Algebra and its Applications, 390(1):155–174, October 1, 2004. CODEN LAAPAW. ISSN 0024-

Barrett:2009:IIP

[BHL09] Wayne Barrett, H. Tracy Hall, and Raphael Loewy. The inverse inertia problem for graphs: Cut vertices, trees, and a counterexample. Linear Algebra and its Applications, 431(8):1147–1191, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2004:RBP

[BHLL04] Jerzy K. Baksalary, Jan Hauke, Xiaoji Liu, and Sanyang Liu. Relationships between partial orders of matrices and their powers. Linear Algebra and its Applications, 379(1):277– 287, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Butkovic:2007:BBA

[BHN^{+}07] Peter Butkovič, Leslie Hogben, Reinhard Nabben, Zdenek Štrakoš, and Miroslav Tůma. A brief biography and appreciation of Miroslav Fiedler with a bibliography of his books and papers. Linear Algebra and its Applications, 421(2–3):173– 181, March 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Banks:2003:MIA

[BHNR03] William Banks, Asma Harcharras, Stefan Neuwirth, and Eric Ricard. Matrix inequalities with applications to the theory of iterated kernels. Linear Algebra and its Applications, 362 (C):275–286, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/211/25/
                                        47/abstract.html.

Blondel:2002:FSI

[BHRV02] V. D. Blondel, D. Hinrichsen, J. Rosenthal, and P. Van Dooren. Fourth special issue on linear systems and control. Linear Algebra and its Applications, 351–352:1– 9, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brankov:2006:ACU

[BHS06] V. Brankov, P. Hansen, and D. Stevanović. Automated conjectures on upper bounds for the largest Laplacian eigenvalue of graphs. Linear Algebra and its Applications, 414(2–3):407–424, April 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ben-Israel:2000:AMV

[BI00] Adi Ben-Israel. An application of the matrix volume in probability. Linear Algebra and its Applications, 321(1– 3):9–25, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/26/article.
pdf.

Beattie:2003:IRM

[BI03] Christopher Beattie and Ilse C. F. Ipsen. Inclusion regions for matrix eigenvalues. Linear Algebra and its Applications, 358(1–3): 281–291, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/207/25/
40/abstract.html.

Biggs:2002:CPR

[Big02] Norman Biggs. Chromatic polynomials and representations of the symmetric group. Linear Algebra and its Applications, 356(1–3): 3–26, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/202/25/
27/abstract.html.

Ben-Israel:2006:GLS

[BIL06] Adi Ben-Israel and Yuri Levin. The geometry of linear separability in data sets. Linear Algebra and its Applications, 416(1):75– 87, July 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bini:2006:SAR

[BILM06] Dario A. Bini, Bruno Iannazzo, Guy Latouche, and Beatrice Meini. On the solution of algebraic Riccati equations arising in fluid queues. Linear Algebra and its Applications, 413(2–3): 474–494, March 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Birmajer:2004:SMS

[Bir04] Daniel Birmajer. On subalgebras of n × n matrices not satisfying identities of degree 2n − 2. Linear Algebra and its Applications, 387(1):133– 141, August 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Biyikoglu:2003:DND

[Bıy03] Türker Bıyıkoğlu. A discrete nodal domain theorem for trees. Linear Algebra and its Applications, 360(C): 197–205, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/

Bueno:2007:MDQ

[BJ07] M. I. Bueno and C. R. Johnson. Minimum deviation, quasi-LU factorization of nonsingular matrices. Linear Algebra and its Applications, 427(1):99–118, November 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Banerjee:2008:SNG

[BJ08] Anirban Banerjee and Jürgen Jost. On the spectrum of the normalized graph Laplacian. Linear Algebra and its Applications, 428(11–12):3015– 3022, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2009:HOD

[BJ09a] Rajendra Bhatia and Tanvi Jain. Higher order derivatives and perturbation bounds for determinants. Linear Algebra and its Applications, 431(11):2102–2108, November 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Blackwood:2009:NGM

[BJ09b] Brian Blackwood and S. K. Jain. Nonnegative groupmonotone matrices and the minus partial order. Linear Algebra and its Applications, 430(1):121–132, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:2004:DGP

[BJT04] R. B. Bapat, S. K. Jain, and John Tynan. Decomposition of generalized polynomial symmetric matrices. Linear Algebra and its Applications, 376(1):69– 83, January 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2000:CDS

[BK00a] Rajendra Bhatia and Fuad Kittaneh. Cartesian decompositions and Schatten norms. Linear Algebra and its Applications, 318(1–3): 109–116, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/137/24/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/137/24/31/article.
pdf.

Bhatia:2000:NMA

[BK00b] Rajendra Bhatia and Fuad Kittaneh. Notes on matrix arithmetic-geometric mean inequalities. Linear Algebra and its Applications, 308(1– 3):203–211, March 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/125/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/125/25/36/article.
pdf.

Brewer:2001:FIL

[BK01] James Brewer and Lee Klingler. On feedback invariants for linear dynamical systems. Linear Algebra and its Applications, 325(1–3): 209–220, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/147/24/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/147/24/38/article.
pdf.

Beasley:2003:NPD

[BK03] LeRoy B. Beasley and Steve Kirkland. A note on kprimitive directed graphs. Linear Algebra and its Applications, 373(1):67–74, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2006:CP

[BK06a] Oskar Maria Baksalary and Paulina Kik. On commutativity of projectors. Linear Algebra and its Applications, 417(1):31–41, August 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benner:2006:BSH

[BK06b] Peter Benner and Daniel Kressner. Balancing sparse Hamiltonian eigenproblems. Linear Algebra and its Applications, 415(1):3–19, May 1, 2006. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2007:MMI

[BK07] Rajendra Bhatia and Hideki Kosaki. Mean matrices and infinite divisibility. Linear Algebra and its Applications, 424(1):36–54, July 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2008:MAG

[BK08] Rajendra Bhatia and Fuad Kittaneh. The matrix arithmetic–geometric mean inequality revisited. Linear Algebra and its Applications, 428(8–9):2177–2191, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2009:SV

[BK09a] Rajendra Bhatia and Fuad Kittaneh. The singular values of A + B and A + iB. Linear Algebra and its Applications, 431(9):1502–1508, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bracic:2009:RDS

[BK09b] Janko Bračič and Bojan Kuzma. Reflexivity defect of spaces of linear operators. Linear Algebra and its Applications, 430(1):344– 359, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Balogh:2001:CAR

[BKDL01] Emese Balogh, Attila Kuba, Csaba Dévényi, and Alberto Del Lungo. Comparison of algorithms for reconstructing hv-convex discrete sets. Linear Algebra and its Applications, 339(1– 3):23–35, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/169/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/169/27/29/article.
pdf.

Brandts:2007:DPS

[BKK07] Jan Brandts, Sergey Korotov, and Michal Křížek. Dissection of the path-simplex in R^{n} into n path-subsimplices.

Linear Algebra and its Applications, 421(2–3):382–393, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brandts:2008:DMP

[BKK08] Jan H. Brandts, Sergey Korotov, and Michal Křížek. The discrete maximum principle for linear simplicial finite element approximations of a reaction–diffusion problem. Linear Algebra and its Applications, 429(10):2344– 2357, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brunat:2006:SCD

[BKLM06] J. M. Brunat, C. Krattenthaler, A. Lascoux, and A. Montes. Some composition determinants. Linear Algebra and its Applications, 416(2–3):355–364, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:2004:P

[BKM^{+}04] R. B. Bapat, R. Kaashoek, R. Mathias, T. Y. Tam, and F. Uhlig. Preface. Linear Algebra and its Applications, 379(1):1–2, March 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bart:2004:PN

[BKMR04a] Harm Bart, Israel Koltracht, Alexander Markus, and Leiba Rodman. Personal notes. Linear Algebra and its Applications, 385(1):7– 21, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bart:2004:P

[BKMR04b] Harm Bart, Israel Koltracht, Alexander Markus, and Leiba Rodman. Preface. Linear Algebra and its Applications, 385(1):1–6, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:2008:P

[BKMR08] Richard Brualdi, Reza Khosrovshahi, Bojan Mohar, and Peter Rowlinson. Preface. Linear Algebra and its Applications, 429(11–12):2661– 2662, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2007:SID

[BKMZ07] Rajendra Bhatia, Fuad Kittaneh, Roy Mathias, and Xingzhi Zhan. Special issue dedicated to Roger Horn. Linear Algebra and its Applications, 424(1):1– 2, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bunger:2001:PTQ

[BKN01] Florian Bünger, Frieder Knüppel, and Klaus Nielsen. The product of two quadratic matrices. Linear Algebra and its Applications, 331 (1–3):31–41, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/28/article.
pdf.

Bapat:2005:DMLa

[BKN05] R. Bapat, S. J. Kirkland, and M. Neumann. On distance matrices and Laplacians. Linear Algebra and its Applications, 401(1):193– 209, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Boros:2001:PBS

[BKO01] Tibor Boros, Thomas Kailath, and Vadim Olshevsky. Pivoting and backward stability of fast algorithms for solving Cauchy linear equations. Linear Algebra and its Applications, 343–344 (1):63–99, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
30/abstract.html.

Blondel:2008:SIJ

[BKPW08] Vincent D. Blondel, Michael Karow, Vladimir Protassov, and Fabian R. Wirth. Special issue on the joint spectral radius: Theory, methods and applications. Linear Algebra and its Applications, 428(10): 2259–2260, May 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bolotnikov:2003:PFI

[BKR03] V. Bolotnikov, A. Kheifets, and L. Rodman. Pairs of functions with indefinite Pick matrices. Linear Algebra and its Applications, 367 (C):271–290, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
45/abstract.html.

Bolotnikov:2004:JFH

[BKR04] Vladimir Bolotnikov, Alexander Kheifets, and Leiba Rodman. Jet functions having indefinite Carathéodory–Pick matrices. Linear Algebra and its Applications, 385(1):215– 286, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Balasubramanian:2006:STO

[BKR06] R. Balasubramanian, S. H. Kulkarni, and R. Radha. Solution of a tridiagonal operator equation. Linear Algebra and its Applications, 414(1): 389–405, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:2009:SBS

[BKR^{+}09] Abraham Berman, Christian Krattenthaler, Siegfried M. Rump, Ilya Spitkovsky, and Fuzhen Zhang. Short biography of Shmuel Friedland for his special LAA volume. Linear Algebra and its Applications, 431(12):2277–2284, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhaya:2001:SAT

[BKS01] A. Bhaya, E. Kaszkurewicz, and Y. Su. Stability of asynchronous two-dimensional Fornasini-Marchesini dynamical systems. Linear Algebra and its Applications, 332–334 (1):257–263, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
40/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/40/article.
pdf.

Bhaya:2003:CCS

[BKS03] A. Bhaya, E. Kaszkurewicz, and R. Santos. Characterizations of classes of stable matrices. Linear Algebra and its Applications, 374(1):159– 174, November 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bilge:2006:CBR

[BKU06] A. H. Bilge, Ş. Koçak, and S. Uḡuz. Canonical bases for real representations of Clifford algebras. Linear Algebra and its Applications, 419(2–3):417–439, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bai:2006:P

[BKvdV06] Zhaojun Bai, Andrew Knyazev, and Henk A. van der Vorst. Preface. Linear Algebra and its Applications, 415(1): 1–2, May 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:2000:LOP

[BL00] LeRoy B. Beasley and Sang-Gu Lee. Linear operators preserving multivariate majorization. Linear Algebra and its Applications, 304(1– 3):141–159, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/118/25/
33/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/118/25/33/article.
                                        pdf.

Boyle:2002:SPM

[BL02] Mike Boyle and Douglas Lind. Small polynomial matrix presentations of nonnegative matrices. Linear Algebra and its Applications, 355(1–3):49–70, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/201/27/
                                        31/abstract.html.

Baksalary:2004:ACG

[BL04] Jerzy K. Baksalary and Xiaoji Liu. An alternative characterization of generalized projectors. Linear Algebra and its Applications, 388 (1):61–65, September 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:2005:SCP

[BL05] Abraham Berman and Thomas J. Laffey. Similarity classes and principal submatrices. Linear Algebra and its Applications, 401(1):341–351, May 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baker:2009:PCE

[BL09a] C. G. Baker and R. B. Lehoucq. Preconditioning constrained eigenvalue problems. Linear Algebra and its Applications, 431(3–4):396– 408, July 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:2009:RRV

[BL09b] LeRoy B. Beasley and Thomas J. Laffey. Real rank versus nonnegative rank. Linear Algebra and its Applications, 431(12):2330–2335, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Biyikoglu:2009:ACD

[BL09c] Türker Bıyıkoḡlu and Josef Leydold. Algebraic connectivity and degree sequences of trees. Linear Algebra and its Applications, 430(2–3):811– 817, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:2004:NRU

[BLdP04] N. Bebiano, R. Lemos, and J. da Providência. Numerical ranges of unbounded operators arising in quantum physics. Linear Algebra and its Applications, 381(1):259– 279, April 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:2005:IQR

[BLdP05] N. Bebiano, R. Lemos, and J. da Providência. Inequalities for quantum relative entropy. Linear Algebra and its Applications, 401(1):159– 172, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:2005:GNR

[BLdPS05] N. Bebiano, R. Lemos, J. da Providência, and G. Soares. On the geometry of numerical ranges in spaces with an indefinite inner product. Linear Algebra and its Applications, 399(1): 17–34, April 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourin:2009:MRH

[BLFS09] Jean-Christophe Bourin, Eun-Young Lee, Masatoshi Fujii, and Yuki Seo. A matrix reverse Hölder inequality. Linear Algebra and its Applications, 431(11):2154– 2159, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brunetti:2001:CCR

[BLG01] S. Brunetti, A. Del Lungo, and Y. Gerard. On the computational complexity of reconstructing threedimensional lattice sets from their two-dimensional Xrays. Linear Algebra and its Applications, 339(1–3): 59–73, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/169/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/169/27/31/article.
pdf.

Beasley:2003:CSP

[BLL03] LeRoy B. Beasley, Sang-Gu Lee, and You-Ho Lee. A characterization of strong preservers of matrix majorization. Linear Algebra and its Applications, 367 (C):341–346, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
50/abstract.html.

Berman:2006:PHM

[BLL06] Abraham Berman, Leonid Lerer, and Raphael Loewy. Preface to the 2005 Haifa Matrix Theory Conference Proceedings. Linear Algebra and its Applications, 416 (1):1, July 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:2006:SEM

[BLLS06] Abraham Berman, Thomas Laffey, Arie Leizarowitz, and Robert Shorten. On the second eigenvalue of matrices associated with TCP. Linear Algebra and its Applications, 416(1):175–183, July 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bini:2002:SMP

[BLM02] Dario A. Bini, Guy Latouche, and Beatrice Meini. Solving matrix polynomial equations arising in queueing problems. Linear Algebra and its Applications, 340(1–3): 225–244, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
43/abstract.html.

Bini:2003:SNM

[BLM03] Dario A. Bini, Guy Latouche, and Beatrice Meini. Solving nonlinear matrix equations arising in Tree-Like stochastic processes. Linear Algebra and its Applications, 366(C):39–64, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
29/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Basharin:2004:LWM

[BLN04] Gely P. Basharin, Amy N. Langville, and Valeriy A. Naumov. The life and work of A. A. Markov. Linear Algebra and its Applications, 386 (1):3–26, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barcucci:2001:XRC

[BLNP01] Elena Barcucci, Alberto Del Lungo, Maurice Nivat, and Renzo Pinzani. X-rays characterizing some classes of discrete sets. Linear Algebra and its Applications, 339(1– 3):3–21, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/169/27/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/169/27/28/article.
pdf.

Burke:2002:TNM

[BLO02] James V. Burke, Adrian S. Lewis, and Michael L. Overton. Two numerical methods for optimizing matrix stability. Linear Algebra and its Applications, 351–352:117– 145, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Blondel:2008:BJS

[Blo08] Vincent D. Blondel. The birth of the joint spectral radius: An interview with Gilbert Strang. Linear Algebra and its Applications, 428 (10):2261–2264, May 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2002:D

[BLOT02] R. Bhatia, C.-K. Li, K. Okubo, and M. J. Tsatsomeros. Dedication. Linear Algebra and its Applications, 341(1– 3):1–3, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/171/27/
                                        27/abstract.html.

Bapat:2006:QAD

[BLP06] R. B. Bapat, A. K. Lal, and Sukanta Pati. A qanalogue of the distance matrix of a tree. Linear Algebra and its Applications, 416 (2–3):799–814, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Braconnier:2000:IOA

[BLR00] T. Braconnier, P. Langlois, and J. C. Rioual. The influence of orthogonality on the Arnoldi method. Linear Algebra and its Applications, 309(1–3):307–323, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
40/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/40/article.
pdf.

Beasley:2002:LOPb

[BLS02] LeRoy B. Beasley, Sang-Gu Lee, and Seok-Zun Song. Linear operators that preserve pairs of matrices which satisfy extreme rank properties. Linear Algebra and its Applications, 350(1–3):263– 272, July 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Belitskii:2005:PCA

[BLS05] Genrich Belitskii, Ruvim Lipyanski, and Vladimir V. Sergeichuk. Problems of classifying associative or Lie algebras and triples of symmetric or skew-symmetric matrices are wild. Linear Algebra and its Applications, 407(1): 249–262, September 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barnabei:2001:CML

[BM01a] M. Barnabei and L. B. Montefusco. On a class of matrices with low displacement rank. Linear Algebra and its Applications, 325(1–3):

161–176, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/147/24/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/147/24/35/article.
pdf.

Behrends:2001:TOP

[BM01b] Ehrhard Behrends and Fritz Mädler. Transformations into optimal parallelism in euclidean spaces (or: how to explain the shape of the electron-density distribution inside a crystal). Linear Algebra and its Applications, 337 (1–3):139–156, November 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/167/27/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/167/27/34/article.
pdf.

Basilio:2004:RSG

[BM04a] J. C. Basilio and M. V. Moreira. A robust solution of the generalized polynomial Bezout identity. Linear Algebra and its Applications, 385(1): 287–303, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Biha:2004:ECS

[BM04b] M. Didi Biha and A. R. Mahjoub. The k-edge connected subgraph problem I: Polytopes and critical extreme points. Linear Algebra and its Applications, 381(1): 117–139, April 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Binding:2004:JZS

[BM04c] P. Binding and A. Markus. Joint zero sets and ranges of several Hermitian forms over complex and quaternionic scalars. Linear Algebra and its Applications, 385 (1):63–72, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bini:2004:NSF

[BM04d] Dario A. Bini and Beatrice Meini. Non-skip-free M/G/1-type Markov chains and Laurent matrix power series. Linear Algebra and its Applications, 386(1):187– 206, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bueno:2004:DTP

[BM04e] M. I. Bueno and F. Marcellán. Darboux transformation and perturbation of linear functionals. Linear Algebra and its Applications, 384 (1):215–242, June 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bradley:2005:LOT

[BM05] Walter Bradley and Jingjing Ma. Lattice-ordered 2 × 2 triangular matrix algebras. Linear Algebra and its Applications, 404(1):262– 274, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bell:2006:SAA

[BM06a] Jason Bell and Marvin Minei. Spectral analysis of the affine graph over the finite ring. Linear Algebra and its Applications, 414(1):244–265, April 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bueno:2006:PPB

[BM06b] M. I. Bueno and F. Marcellán. Polynomial perturbations of bilinear functionals and Hessenberg matrices. Linear Algebra and its Applications, 414(1):64– 83, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhat:2007:IM

[BM07] B. V. Rajarama Bhat and Mithun Mukherjee. Integrators of matrices. Linear Algebra and its Applications, 426 (1):71–82, October 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berns-Muller:2006:III

[BMGS06] Jörg Berns-Müller, Ivan G. Graham, and Alastair Spence. Inexact inverse iteration for symmetric matrices. Linear Algebra and its Applications, 416(2–3):389–413, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

B:2008:STL

[BMNA08] Javad Ebrahimi B, Bojan Mohar, Vladimir Nikiforov, and Azhvan Sheikh Ahmady. On the sum of two largest eigenvalues of a symmetric matrix. Linear Algebra and its Applications, 429(11–12):2781–2787, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bondyfalat:2000:CSR

[BMP00] Didier Bondyfalat, Bernard Mourrain, and Victor Y. Pan. Computation of a specified root of a polynomial system of equations using eigenvectors. Linear Algebra and its Applications, 319(1–3): 193–209, November 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/138/24/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/138/24/36/article.
                                        pdf.

Borobia:2004:NCN

[BMS04] Alberto Borobia, Julio Moro, and Ricardo Soto. Negativity compensation in the nonnegative inverse eigenvalue problem. Linear Algebra and its Applications, 393(1):73– 89, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Belardo:2006:SRI

[BMS06] Francesco Belardo, Enzo Maria Li Marzi, and Slobodan K. Simić. Some results on the index of unicyclic graphs. Linear Algebra and its Applications, 416(2– 3):1048–1059, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Belardo:2007:SNG

[BMS07] Francesco Belardo, Enzo Maria Li Marzi, and Slobodan K. Simić. Some notes on graphs whose index is close to 2. Linear Algebra and its Applications, 423(1): 81–89, May 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Borobia:2008:UVC

[BMS08a] Alberto Borobia, Julio Moro, and Ricardo L. Soto. A unified view on compensation criteria in the real nonnegative inverse eigenvalue problem. Linear Algebra and its Applications, 428(11–12): 2574–2584, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:2008:LYO

[BMS08b] Richard A. Brualdi, Volker Mehrmann, and Hans Schneider. LAA is 40 years old. Linear Algebra and its Applications, 428(1):1– 3, January 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Byers:2008:TLS

[BMX08] Ralph Byers, Volker Mehrmann, and Hongguo Xu. Trimmed linearizations for structured matrix polynomials. Linear Algebra and its Applications, 429(10):2373–2400, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benzi:2000:DRJ

[BN00] Michele Benzi and Jim Nagz. Dedication to Robert J. Plemmons. Linear Algebra and its Applications, 316(1– 3):1–12, September 1, 2000.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/24/article.
pdf.

Bapat:2004:IPI

[BN04] Ravindra Bapat and Michael Neumann. Inequalities for permanents involving Perron complements. Linear Algebra and its Applications, 385(1): 95–104, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bekker:2005:TIP

[BN05] M. B. Bekker and A. A. Nudel’man. Tangential interpolation problems for a class of automorphic matrixfunctions. Linear Algebra and its Applications, 397(1):171– 184, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beauchamp:2008:OMA

[BN08] Kyle Beauchamp and Remus Nicoara. Orthogonal maximal abelian *-subalgebras of the 6 × 6 matrices. Linear Algebra and its Applications, 428(8–9):1833–1853, April 15, 2008. CODEN LAAPAW. ISSN 0024-

Bebiano:2005:IHM

[BNdP^{+}05] N. Bebiano, H. Nakazato, J. da Providência, R. Lemos, and G. Soares. Inequalities for J-Hermitian matrices. Linear Algebra and its Applications, 407(1):125–139, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:2009:CFT

[BNdP09] N. Bebiano, H. Nakazato, and J. da Providência. On the Courant–Fischer theory for Krein spaces. Linear Algebra and its Applications, 430 (7):1867–1876, April 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Blondel:2005:AEN

[BNT05] Vincent D. Blondel, Yurii Nesterov, and Jacques Theys. On the accuracy of the ellipsoid norm approximation of the joint spectral radius. Linear Algebra and its Applications, 394(1):91– 107, January 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Blondel:2005:AEP

[BNV05] Vincent D. Blondel, Laure Ninove, and Paul Van Dooren. An affine eigenvalue problem on the nonnegative orthant. Linear Algebra and its Applications, 404(1): 69–84, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bo:2002:CIM

[Bo02] Zhou Bo. The changes in indices of modified graphs. Linear Algebra and its Applications, 356(1–3):95–101, November 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/202/25/
33/abstract.html.

Bernik:2005:SME

[BO05] Janez Bernik and Jan Okniński. On semigroups of matrices with eigenvalue 1 in small dimensions. Linear Algebra and its Applications, 405(1): 67–73, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Blumthaler:2009:O

[BO09] Ingrid Blumthaler and Ulrich Oberst. T -observers. Linear Algebra and its Applications, 430(8–9):2416–2447, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bochi:2003:INI

[Boc03a] Jairo Bochi. Inequalities for numerical invariants of sets of matrices. Linear Algebra and its Applications, 368 (C):71–81, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
29/abstract.html.

Bocklandt:2003:SQS

[Boc03b] Raf Bocklandt. Symmetric quiver settings with a regular ring of invariants. Linear Algebra and its Applications, 365(C):25–43, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
36/abstract.html.

Bojanczyk:2003:IJL

[Boj03] Adam W. Bojanczyk. An implicit Jacobi-like method for computing generalized hyperbolic SVD. Linear Algebra and its Applications, 358(1– 3):293–307, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/207/25/
41/abstract.html.

Bollhofer:2001:RIP

[Bol01] Matthias Bollhöfer. A robust ILU with pivoting based on monitoring the growth of the inverse factors. Linear Algebra and its Applications, 338 (1–3):201–218, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
41/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/41/article.
pdf.

Bolotnikov:2002:BNP

[Bol02]  Vladimir Bolotnikov.  A

boundary Nevanlinna–Pick problem for a class of analytic matrix-valued functions in the unit ball. Linear Algebra and its Applications, 346 (1–3):239–260, May 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/189/27/
38/abstract.html.

Bolla:2004:DER

[Bol04] Marianna Bolla. Distribution of the eigenvalues of random block-matrices. Linear Algebra and its Applications, 377(1):219–240, January 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bolla:2005:RLS

[Bol05] Marianna Bolla. Recognizing linear structure in noisy matrices. Linear Algebra and its Applications, 402(1):228– 244, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bomze:2008:PFP

[Bom08] Immanuel M. Bomze. Perron– Frobenius property of copositive matrices, and a block copositivity criterion. Linear Algebra and its Applications, 429(1):68–71, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bondar:2003:SMI

[Bon03a] James V. Bondar. Schur majorization inequalities for symmetrized sums with applications to tensor products. Linear Algebra and its Applications, 360(C):1– 13, February 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
26/abstract.html.

Bondarenko:2003:LOG

[Bon03b] Vitalij M. Bondarenko. Linear operators on S-graded vector spaces. Linear Algebra and its Applications, 365(C):45–90, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/214/25/
                                        50/abstract.html.

Bondarenko:2009:DPL

[Bon09] Ievgen Bondarenko. Dynamics of piecewise linear maps and sets of nonnegative matrices. Linear Algebra and its Applications, 431(5–7):495– 510, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bella:2008:LSC

[BOP08] T. Bella, V. Olshevsky, and U. Prasad. Lipschitz stability of canonical Jordan bases of H-selfadjoint matrices under structure-preserving perturbations. Linear Algebra and its Applications, 428(8– 9):2130–2176, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Borobia:2006:CMP

[Bor06] Alberto Borobia. Constructing matrices with prescribed main-diagonal submatrix and characteristic polynomial. Linear Algebra and its Applications, 418(2–3):886–890, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Borm:2007:DSA

[Bör07] Steffen Börm. Data-sparse approximation of non-local operators by H^{∈}-matrices.

Linear Algebra and its Applications, 422(2–3):380–403, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Borne:2008:BCR

[Bor08] Sabine Le Borne. Block computation and representation of a sparse nullspace basis of a rectangular matrix. Linear Algebra and its Applications, 428(11–12):2455– 2467, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhatia:2004:SHR

[BORŠ04] Rajendra Bhatia, Matjaz Omladič, Peter Rosenthal, and Peter Šemrl. A survey of Heydar Radjavi. Linear Algebra and its Applications, 383 (1):1–15, May 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Boshnakov:2002:MCM

[Bos02] Georgi N. Boshnakov. Multicompanion matrices. Linear Algebra and its Applications, 354(1–3):53–83, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
34/abstract.html.

Boshnakov:2007:SVD

[Bos07] Georgi N. Boshnakov. Singular value decomposition of multi-companion matrices. Linear Algebra and its Applications, 424(2–3):393– 404, July 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Botha:2000:SDM

[Bot00] J. D. Botha. Sums of diagonalizable matrices. Linear Algebra and its Applications, 315(1–3):1–23, August 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/132/25/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/132/25/25/article.
                                        pdf.

Botha:2006:URR

[Bot06a] J. D. Botha. On a unification result by A. R. Sourour concerning commutators and products of involutions. Linear Algebra and its Applications, 416(2–3):872– 879, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bottcher:2006:SPB

[Böt06b] Albrecht Böttcher. Book review: Spectra and Pseudospectra: The Behavior of Nonnormal Matrices and Operators by Lloyd N. Trefethen and Mark Embree, Princeton University Press, Princeton and Oxford, 2005, xvii + 606 pages, ISBN-10 0-691-11946-5. Linear Algebra and its Applications, 416(2–3):1098– 1101, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Botha:2009:USM

[Bot09] J. D. Botha. A unification of some matrix factorization results. Linear Algebra and its Applications, 431(10): 1719–1725, October 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourin:2003:SVC

[Bou03a] Jean-Christophe Bourin. Singular values of compressions, restrictions and dilations. Linear Algebra and its Applications, 360(C): 259–272, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
43/abstract.html.

Bourin:2003:TD

[Bou03b] Jean-Christophe Bourin. Total dilations. Linear Algebra and its Applications, 368 (C):159–169, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/217/25/
                                        35/abstract.html.

Bourin:2003:TDI

[Bou03c] Jean-Christophe Bourin. Total dilations II. Linear Algebra and its Applications, 374(1):19–29, November 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourin:2005:SVP

[Bou05] Jean-Christophe Bourin. Singular values of products of positive operators: AZB and ZAB. Linear Algebra and its Applications, 407(1):64– 70, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourin:2006:CIS

[Bou06a] Jean-Christophe Bourin. A concavity inequality for symmetric norms. Linear Algebra and its Applications, 413(1): 212–217, February 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourin:2006:MVS

[Bou06b] Jean-Christophe Bourin. Matrix versions of some classical inequalities. Linear Algebra and its Applications, 416 (2–3):890–907, July 15, 2006.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourgeois:2007:CEL

[Bou07] Gerald Bourgeois. On commuting exponentials in low dimensions. Linear Algebra and its Applications, 423 (2–3):277–286, June 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourgeois:2009:ASM

[Bou09]  Gerald Bourgeois.  Alge-

braic systems of matrices and Gröbner basis theory. Linear Algebra and its Applications, 430(8–9):2157–2169, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Britz:2004:MPI

[BOvdD04] T. Britz, D. D. Olesky, and P. van den Driessche. The Moore–Penrose inverse of matrices with an acyclic bipartite graph. Linear Algebra and its Applications, 390(1): 47–60, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bingham:2007:PNS

[BOvdD07] B. D. Bingham, D. D. Olesky, and P. van den Driessche. Potentially nilpotent and spectrally arbitrary even cycle sign patterns. Linear Algebra and its Applications, 421(1): 24–44, February 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Blondel:2002:PZI

[BP02] Vincent D. Blondel and Natacha Portier. The presence of a zero in an integer linear recurrent sequence is NP-hard to decide. Linear Algebra and its Applications, 351–352:91– 98, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barrett:2003:NSC

[BP03] Wayne Barrett and Stephen Pierce. Null spaces of correlation matrices. Linear Algebra and its Applications, 368 (C):129–157, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
34/abstract.html.

Barik:2005:ACS

[BP05a] Sasmita Barik and Sukanta Pati. On algebraic connectivity and spectral integral variations of graphs. Linear Algebra and its Applications, 397 (1):209–222, March 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bodmann:2005:FGE

[BP05b] Bernhard G. Bodmann and Vern I. Paulsen. Frames, graphs and erasures. Linear Algebra and its Applications, 404(1):118–146, July 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bremner:2006:TAL

[BP06a] Murray R. Bremner and Luiz A. Peresi. Ternary analogues of Lie and Malcev algebras. Linear Algebra and its Applications, 414 (1):1–18, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Buescu:2006:LAA

[BP06b] Jorge Buescu and A. C. Paixão. A linear algebraic approach to holomorphic reproducing kernels in C^{n}. Linear Algebra and its Applications, 412(2–3):270–290, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bremner:2009:ALB

[BP09] Murray R. Bremner and Luiz A. Peresi. An application of lattice basis reduction to polynomial identities for algebraic structures. Linear Algebra and its Applications, 430(2–3):642–659, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Boyko:2008:ISL

[BPP08] Vyacheslav Boyko, Jiri Patera, and Roman Popovych. Invariants of solvable lie algebras with triangular nilradicals and diagonal nilindependent elements. Linear Algebra and its Applications, 428(4):834–854, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bessenrodt:2003:NOB

[BPR03] C. Bessenrodt, M. R. Pournaki, and A. Reifegerste. A note on the orthogonal basis of a certain full symmetry class of tensors. Linear Algebra and its Applications, 370(1):369–374, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bru:2004:OAM

[BPS04] Rafael Bru, Francisco Pedroche, and Daniel B. Szyld. Overlapping additive and multiplicative Schwarz iterations for H-matrices. Linear Algebra and its Applications, 393(1):91–105, December 1, 2004. CODEN LAAPAW. ISSN 0024-

Bhattacharya:2007:CCA

[BPS07] Amitava Bhattacharya, Uri N. Peled, and Murali K. Srinivasan. Cones of closed alternating walks and trails. Linear Algebra and its Applications, 423(2–3):351–365, June 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhattacharya:2009:CBS

[BPS09] Amitava Bhattacharya, Uri N. Peled, and Murali K. Srinivasan. The cone of balanced subgraphs. Linear Algebra and its Applications, 431(1– 2):266–273, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Belur:2007:DSS

[BPT07] Madhu N. Belur, Harish K. Pillai, and H. L. Trentelman. Dissipative systems synthesis: A linear algebraic approach. Linear Algebra and its Applications, 425(2–3): 739–756, September 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bodmann:2009:ETF

[BPT09] Bernhard G. Bodmann, Vern I. Paulsen, and Mark Tomforde. Equiangular tight frames from complex Seidel matrices containing cube roots of unity. Linear Algebra and its Applications, 430 (1):396–417, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barlow:2000:ASE

[BPV00] J. L. Barlow, Beresford N. Parlett, and Krešimir Veselić. The accurate solution of eigenvalue problems. Linear Algebra and its Applications, 309(1–3):1–2, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/25/article.
pdf.

Barlow:2006:P

[BPV06] Jesse L. Barlow, Beresford N. Parlett, and Kresimir Veselić. Preface. Linear Algebra and its Applications, 417(2–3): 381, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:2008:ESP

[BPZ08] Richard A. Brualdi, Pei Pei, and Xingzhi Zhan. An extremal sparsity property of the Jordan canonical form. Linear Algebra and its Applications, 429(10):2367–2372, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bojanczyk:2000:UUH

[BQS00] Adam Bojanczyk, Sanzheng Qiao, and Allan O. Steinhardt. Unifying unitary and hyperbolic transformations. Linear Algebra and its Applications, 316(1–3): 183–197, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/36/article.
pdf.

Browne:2009:LBA

[BQW09] Kevin Browne, Sanzheng Qiao, and Yimin Wei. A Lanczos bidiagonalization algorithm for Hankel matrices. Linear Algebra and its Applications, 430(5–6):1531– 1543, March 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baribeau:2000:CSF

[BR00a] Line Baribeau and Sylvain Roy. Caractérisation spectrale de la forme de Jordan. Linear Algebra and its Applications, 320(1–3): 183–191, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/37/article.
pdf.

Baur:2000:WCC

[BR00b] Franziska Baur and Werner J. Ricker. The Weyl calculus and a Cayley–Hamilton theorem for pairs of selfadjoint matrices. Linear Algebra and its Applications, 319(1–3): 103–116, November 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/138/24/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/138/24/30/article.
pdf.

Bolotnikov:2001:PAM

[BR01] V. Bolotnikov and L. Rodman. On positivity of analytic matrix functions in polydisks. Linear Algebra and its Applications, 328(1– 3):69–94, May 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/153/25/27/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/153/
25/27/article.pdf.

Bolotnikov:2002:FDB

[BR02] V. Bolotnikov and L. Rodman. Finite dimensional backward shift invariant subspaces of Arveson spaces. Linear Algebra and its Applications, 349(1–3):265–282, July 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:2006:NCC

[BR06] Abraham Berman and Uriel G. Rothblum. A note on the computation of the CPrank. Linear Algebra and its Applications, 419(1):1– 7, November 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Boros:2007:EFS

[BR07] Tibor Boros and Pál Rózsa. An explicit formula for singular values of the Sylvester– Kac matrix. Linear Algebra and its Applications, 421(2– 3):407–416, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bapat:2009:IDM

[BR09] R. B. Bapat and Pritha Rekhi. Inverses of q-distance matrices of a tree. Linear Algebra and its Applications, 431(10):1932–1939, October 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brandts:2003:RAE

[Bra03] Jan Brandts. The Riccati algorithm for eigenvalues and invariant subspaces of matrices with inexpensive action. Linear Algebra and its Applications, 358(1–3): 335–365, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/207/25/
                                        43/abstract.html.

Brand:2006:FLR

[Bra06] Matthew Brand. Fast lowrank modifications of the thin singular value decomposition. Linear Algebra and its Applications, 415(1):20– 30, May 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bracic:2007:RCL

[Bra07] Janko Bračič. Reflexivity of the commutant and local commutants of an algebraic operator. Linear Algebra and its Applications, 420(1):20– 28, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brenner:2003:KIM

[Bre03] Sheila Brenner. On the kernel of an irreducible map. Linear Algebra and its Applications, 365(C):91–97, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
26/abstract.html.

Britz:2001:INS

[Bri01] Thomas Britz. The inverse of a non-singular free matrix. Linear Algebra and its Applications, 338(1–3): 245–249, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
44/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/44/article.
pdf.

Briand:2005:MFI

[Bri05] Emmanuel Briand. On a matrix function interpolating between determinant and permanent. Linear Algebra and its Applications, 411(1): 343–347, December 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bahi:2001:PSL

[BRM01] J. M. Bahi, K. Rhofir, and J.-C. Miellou. Parallel solution of linear DAEs by multisplitting waveform relaxation methods. Linear Algebra and its Applications, 332–334(1):

181–196, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/36/article.
pdf.

Bakule:2001:COE

[BRR01] Lubomír Bakule, José Rodellar, and Josep M. Rossell. Controllability-observability of expanded composite systems. Linear Algebra and its Applications, 332–334(1): 381–400, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
44/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/44/article.
pdf.

Bru:2000:CFP

[BRS00] Rafael Bru, Sergio Romero, and Elena Sánchez. Canonical forms for positive discretetime linear control systems. Linear Algebra and its Applications, 310(1–3):49–71, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/

10/30/19/127/25/31/article.
pdf.

Brualdi:2000:EC

[Bru00] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 320(1–3):209–216, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
41/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/41/article.
pdf.

Brualdi:2002:EC

[Bru02a] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 340(1–3):257–258, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
47/abstract.html.

Brualdi:2002:MNI

[Bru02b] Richard A. Brualdi. Minimal nonnegative integral matrices and uniquely determined (0,1)-matrices. Linear Algebra and its Applications, 341 (1–3):351–356, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
55/abstract.html.

Brualdi:2003:EC

[Bru03a] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 360(C):279–283, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/209/25/
                                        47/abstract.html.

Brualdi:2003:LPD

[Bru03b] Richard A. Brualdi. Linear preservers and diagonal hypergraphs. Linear Algebra and its Applications, 373 (1):51–65, November 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brustle:2003:PRP

[Brü03c] Thomas Brüstle. On positive roots of pg-critical algebras. Linear Algebra and its Applications, 365(C):107– 114, May 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/214/25/
                                        31/abstract.html.

Brualdi:2004:EC

[Bru04] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 380 (1):273–274, March 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:2005:EC

[Bru05] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 400(1):353–357, May 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:2007:EC

[Bru07] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 420(2–3):727–730, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brualdi:2009:EC

[Bru09a] Richard A. Brualdi. From the Editor-in-Chief. Linear Algebra and its Applications, 430 (1):587–593, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brull:2009:EUS

[Brü09b]  Tobias Brüll.  Existence

and uniqueness of solutions of linear variable coefficient discrete-time descriptor systems. Linear Algebra and its Applications, 431(1–2):247– 265, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bru:2008:RIP

[BRVS08] Rafael Bru, Sergio Romero-Vivó, and Elena Sánchez.

Reachability indices of periodic positive systems via positive shift-similarity. Linear Algebra and its Applications, 429(5–6):1288–1301, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brezinski:2004:NVS

[BRZ04] C. Brezinski and M. Redivo-Zaglia. New vector sequence transformations. Linear Algebra and its Applications, 389(1):189–213, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Barlow:2000:OPB

[BS00a] Jesse L. Barlow and Ivan Slapničar. Optimal perturbation bounds for the Hermitian eigenvalue problem. Linear Algebra and its Applications, 309(1–3):19–43, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/27/article.
pdf.

Baser:2000:ESD

[BS00b] U. Başer and J. M. Schumacher. The equivalence structure of descriptor representations of systems with possibly inconsistent initial conditions. Linear Algebra and its Applications, 318(1– 3):53–77, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/137/24/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/137/24/28/article.
pdf.

Bell:2000:ARO

[BS00c] Jeremy Bell and A. R. Sourour. Additive rank-one preserving mappings on triangular matrix algebras. Linear Algebra and its Applications, 312(1–3):13–33, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/129/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/129/25/26/article.
pdf.

Borovykh:2000:SSE

[BS00d] N. Borovykh and M. N. Spijker. The sharpness of stability estimates corresponding to a general resolvent condition. Linear Algebra and its Applications, 311(1– 3):161–175, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/128/25/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/128/25/36/article.
                                        pdf.

Bermudez:2001:CFL

[BS01] José María Ancochea Bermúdez and Otto Rutwig Campoamor Stursberg. Classification of (n − 5)-filiform Lie algebras. Linear Algebra and its Applications, 336(1–3): 167–180, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/166/27/
                                        37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/166/27/37/article.
                                        pdf.

Baksalary:2002:GIP

[BS02a] Jerzy K. Baksalary and George P. H. Styan. Generalized inverses of partitioned matrices in Banachiewicz– Schur form. Linear Algebra and its Applications, 354(1– 3):41–47, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/200/27/
                                        32/abstract.html.

Barchini:2002:FRD

[BS02b] L. Barchini and Mark R. Sepanski. Finite-reductive dual pairs in G_{2}. Linear Algebra and its Applications, 340(1– 3):123–136, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
35/abstract.html.

Beasley:2002:LOPa

[BS02c] Leroy B. Beasley and Seok-Zun Song. Linear operators that preserve zeroterm rank over fields and rings. Linear Algebra and its Applications, 341(1–3): 143–149, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
38/abstract.html.

Belitskii:2003:CMP

[BS03a] Genrich R. Belitskii and Vladimir V. Sergeichuk. Complexity of matrix problems. Linear Algebra and its Applications, 361(C):203– 222, March 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
40/abstract.html.

Brustle:2003:ENO

[BS03b] Thomas Brüstle and Vladimir V. Sergeichuk. Estimate of the number of one-parameter families of modules over a tame algebra. Linear Algebra and its Applications, 365 (C):115–133, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/214/25/
                                        27/abstract.html.

Bell:2004:GWS

[BS04a] Francis K. Bell and Slobodan K. Simiç. On graphs whose star complement for −2 is a path or a cycle. Linear Algebra and its Applications, 377(1):249– 265, January 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bolotnikov:2004:PEP

[BS04b] Vladimir Bolotnikov and Paul A. Smith. Positive extension problems for a class of structured matrices. Linear Algebra and its Applications, 381(1):165– 195, April 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brown:2004:FPB

[BS04c] Ethan S. Brown and Ilya M. Spitkovsky. On flat portions on the boundary of the numerical range. Linear Algebra and its Applications, 390(1): 75–109, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-

Baksalary:2005:SCL

[BS05a] Oskar Maria Baksalary and George P. H. Styan. Some comments on the life and publications of Jerzy K. Baksalary (1944–2005). Linear Algebra and its Applications, 410(1):3–53, November 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bebiano:2005:TOD

[BS05b] N. Bebiano and G. Soares. Three observations on the determinantal range. Linear Algebra and its Applications, 401(1):211–220, May 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bai:2006:PMM

[BS06a] Zhaojun Bai and Daniel Skoogh. A projection method for model reduction of bilinear dynamical systems. Linear Algebra and its Applications, 415(2–3):406–425, June 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Belitskii:2006:CMF

[BS06b] Genrich R. Belitskii and Vladimir V. Sergeichuk. Congruence of multilinear forms.

Linear Algebra and its Applications, 418(2–3):751–762, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourin:2007:RIG

[BS07] Jean-Christophe Bourin and Yuki Seo. Reverse inequality to Golden–Thompson type inequalities: Comparison of e^{A+B} and e^{A}e^{B}. Linear Algebra and its Applications, 426(2–3):312–316, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Blackburn:2008:AEC

[BS08] Simon R. Blackburn and Igor E. Shparlinski. On the average energy of circulant graphs. Linear Algebra and its Applications, 428(8– 9):1956–1963, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bendaoud:2009:AMP

[BS09a] M. Bendaoud and M. Sarih. Additive maps preserving the ascent and descent of operators. Linear Algebra and its Applications, 431(10):1740– 1744, October 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bhouri:2009:ANS

[BS09b] Imen Bhouri and Haı¨ kel Skhiri. On the approximation numbers and spectral eigenvalues. Linear Algebra and its Applications, 430(2–3):847– 854, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berge:2009:CCC

[BSD09] Jos M. F. Ten Berge, Alwin Stegeman, and Mohammed Bennani Dosse. The Carroll and Chang conjecture of equal Indscal components when Candecomp/ Parafac gives perfect fit. Linear Algebra and its Applications, 430(2–3):818–829, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Beasley:2005:CRT

[BSKS05] LeRoy B. Beasley, Seok-Zun Song, Kyung-Tae Kang, and B. K. Sarma. Column ranks and their preservers over nonnegative real matrices. Linear Algebra and its Applications, 399(1):3–16, April 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:2004:PMA

[BSL04] Abraham Berman, Robert Shorten, and Douglas Leith. Positive matrices associated with synchronised communication networks. Linear Algebra and its Applications, 393 (1):47–54, December 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Butkovic:2007:GEB

[BSS07] Peter Butkovič, Hans Schneider, and Sergeĭ Sergeev. Generators, extremals and bases of max cones. Linear Algebra and its Applications, 421(2– 3):394–406, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Botchev:2009:SAJ

[BSS09] M. A. Botchev, G. L. G. Sleijpen, and A. Sopaheluwakan. An SVD-approach to Jacobi– Davidson solution of nonlinear Helmholtz eigenvalue problems. Linear Algebra and its Applications, 431(3– 4):427–440, July 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2009:MDH

[BST09] Oskar Maria Baksalary, George P. H. Styan, and Götz Trenkler. On a matrix decomposition of Hartwig and Spindelböck. Linear Algebra and its Applications, 430(10): 2798–2812, May 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benton:2006:CPA

[BSW06] Jessica Benton, Rion Snow, and Nolan Wallach. A combinatorial problem associated with nonograms. Linear Algebra and its Applications, 412(1):30–38, January 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bayat:2000:PEF

[BT00] M. Bayat and H. Teimoori. Pascal k-eliminated functional matrix and its property. Linear Algebra and its Applications, 308(1–3): 65–75, March 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/125/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/125/25/27/article.
pdf.

Belyi:2004:KFI

[BT04a] Sergey Belyi and Eduard Tsekanovskii. On Krein’s formula in indefinite metric spaces. Linear Algebra and its Applications, 389(1): 305–322, September 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bradshow:2004:FSM

[BT04b] David Bradshow and Alexander Tovbis. On formal solutions to matrix version of a reduced Kac–van Moerbeke equation. Linear Algebra and its Applications, 384(1):207– 214, June 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benitez:2005:ILC

[BT05a] J. Benítez and N. Thome. Idempotency of linear combinations of an idempotent matrix and a t-potent matrix that commute. Linear Algebra and its Applications, 403(1):414–418, July 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Benitez:2005:CLC

[BT05b] Julio Benítez and Néstor Thome. Characterizations and linear combinations of kgeneralized projectors. Linear Algebra and its Applications, 410(1):150–159, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bikchentaev:2007:CTM

[BT07a] Airat M. Bikchentaev and Oleg E. Tikhonov. Characterization of the trace by monotonicity inequalities. Linear Algebra and its Applications, 422(1):274– 278, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-

Blanco:2007:CAT

[BT07b] A. Blanco and A. Turnšek. On the converse of Anderson’s theorem. Linear Algebra and its Applications, 424 (2–3):384–389, July 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2009:EFO

[BT09a] Oskar Maria Baksalary and Götz Trenkler. Eigenvalues of functions of orthogonal projectors. Linear Algebra and its Applications, 431(11):2172–2186, November 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2009:ADB

[BT09b] Oskar Maria Baksalary and Götz Trenkler. On angles and distances between subspaces. Linear Algebra and its Applications, 431(11):2243–2260, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Baksalary:2009:RPT

[BT09c] Oskar Maria Baksalary and Götz Trenkler. Revisitation of the product of two orthogonal projectors. Linear Algebra and its Applications, 430 (10):2813–2833, May 1, 2009.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bahturin:2009:GGS

[BTT09] Yu Bahturin, M. Tvalavadze, and T. Tvalavadze. Group gradings on superinvolution simple superalgebras. Linear Algebra and its Applications, 431(5–7):1054–1069, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bu:2005:LMP

[Bu05]  Changjiang Bu.  Linear

maps preserving Drazin inverses of matrices over fields. Linear Algebra and its Applications, 396(1):159–173, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bourin:2007:MSI

[BU07] Jean-Christophe Bourin and Mitsuru Uchiyama. A matrix subadditivity inequality for f(A + B) and f(A) + f(B). Linear Algebra and its Applications, 423(2–3):512– 518, June 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Buch:2006:EHM

[Buc06] Anders Skovsted Buch. Eigenvalues of Hermitian matrices with positive sum of bounded rank. Linear Algebra and its Applications, 418(2–3):480– 488, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bukovsek:2007:MSH

[Buk07] Damjana Kokol Bukovšek. Matrix semigroup homomorphisms into higher dimensions. Linear Algebra and its Applications, 420(1):34– 50, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Buliga:2008:FAM

[Bul08] Marius Buliga. Four applications of majorization to convexity in the calculus of variations. Linear Algebra and its Applications, 429(7): 1528–1545, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Burde:2005:MEX

[Bur05] Dietrich Burde. On the matrix equation XA − AX = X^{p}. Linear Algebra and its Applications, 404(1):147– 165, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Burman:2006:ASV

[Bur06] Prabir Burman. Approximate singular values of the fractional difference and summation operators. Linear Algebra and its Applications, 416(2–3):677–687, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Butkovic:2003:MAL

[But03] Peter Butkovič. Max-algebra: the linear algebra of combinatorics? Linear Algebra and its Applications, 367 (C):313–335, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
48/abstract.html.

Butler:2006:UDC

[But06] Steve Butler. Using discrepancy to control singular values for nonnegative matrices. Linear Algebra and its Applications, 419(2–3):486–493, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Butkovic:2008:PMA

[But08] P. Butkovič. Permuted maxalgebraic eigenvector problem is NP-complete. Linear Algebra and its Applications, 428(8–9):1874–1882, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bobrovnikova:2000:NBP

[BV00] Elena Y. Bobrovnikova and Stephen A. Vavasis. A norm bound for projections with complex weights. Linear Algebra and its Applications, 307(1–3):69–75, March 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/29/article.
pdf.

Behn:2002:AVT

[BV02] A. Behn and A. B. Van der Merwe. An algorithmic version of the theorem by Latimer and MacDuffee for 2×2 integral matrices. Linear Algebra and its Applications, 346(1–3):1–14, May 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/189/27/
27/abstract.html.

Binding:2007:GEP

[BV07] P. A. Binding and H. Volkmer. A generalized eigenvalue problem in the max algebra. Linear Algebra and its Applications, 422(2–3):360– 371, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bang:2008:SCH

[BvDK08] Sejeong Bang, Edwin R. van Dam, and Jack H. Koolen. Spectral characterization of the Hamming graphs. Linear Algebra and its Applications, 429(11–12):2678–2686, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bart:2000:IPP

[BW00] H. Bart and A. P. M. Wagelmans. An integer programming problem and rank decomposition of block upper triangular matrices. Linear Algebra and its Applications, 305(1–3):107–129, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/120/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/120/25/33/article.
pdf.

Bostian:2001:UMR

[BW01] A. A. Bostian and H. J. Woerdeman. Unicity of minimal rank completions for tri-diagonal partial block matrices. Linear Algebra and its Applications, 325(1– 3):23–55, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/147/24/

26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/147/24/26/article.
pdf.

Bai:2003:CTP

[BW03] Zhong-Zhi Bai and Chuan-Long Wang. Convergence theorems for parallel multisplitting two-stage iterative methods for mildly nonlinear systems. Linear Algebra and its Applications, 362 (C):237–250, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/211/25/
44/abstract.html.

Bottcher:2005:HBC

[BW05] Albrecht Böttcher and David Wenzel. How big can the commutator of two matrices be and how big is it typically? Linear Algebra and its Applications, 403(1):216– 228, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bottcher:2006:SJ

[BW06] Albrecht Böttcher and Harold Widom. Szegő via Jacobi. Linear Algebra and its Applications, 419(2–3):656–667, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bai:2008:PIU

[BW08a] Zhong-Zhi Bai and Zeng-Qi Wang. On parameterized inexact Uzawa methods for generalized saddle point problems. Linear Algebra and its Applications, 428(11–12): 2900–2932, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bottcher:2008:FNC

[BW08b] Albrecht Böttcher and David Wenzel. The Frobenius norm and the commutator. Linear Algebra and its Applications, 429(8–9):1864–1885, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:2004:CPM

[BX04] Abraham Berman and Changqing Xu. 5 × 5 completely positive matrices. Linear Algebra and its Applications, 393(1):55– 71, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:2005:CPM

[BX05] Abraham Berman and Changqing Xu. {0, 1} completely positive matrices. Linear Algebra and its Applications, 399(1): 35–51, April 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bai:2008:INT

[BX08] Zheng-Jian Bai and Shufang Xu. An inexact Newton-type method for inverse singular value problems. Linear Algebra and its Applications, 429 (2–3):527–547, July 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Berman:2000:LBE

[BZ00] Abraham Berman and Xiao-Dong Zhang. Lower bounds for the eigenvalues of Laplacian matrices. Linear Algebra and its Applications, 316(1– 3):13–20, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/25/article.
pdf.

Baragana:2002:FIR

[BZ02] I. Baragaña and I. Zaballa. Feedback invariants of restrictions and quotients: series connected systems. Linear Algebra and its Applications, 351–352:69–89, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Brezinski:2003:SCA

[BZ03] C. Brezinski and M. Redivo Zaglia. A Schur complement approach to a general extrapolation algorithm. Linear Algebra and its Applications, 368(C):279–301, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
43/abstract.html.

Beitia:2005:EQF

[BZ05] M. Asunción Beitia and Ion Zaballa. Equivalence of quadruples and feedback equivalence of related pairs. Linear Algebra and its Applications, 401(1):353–369, May 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bader:2006:COM

[BZ06] Michael Bader and Christoph Zenger. Cache oblivious matrix multiplication using an element ordering based on a Peano curve. Linear Algebra and its Applications, 417(2–3):301–313, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chebotarev:2002:FMA

[CA02] Pavel Chebotarev and Rafig Agaev. Forest matrices around the Laplacian matrix. Linear Algebra and its Applications, 356(1–3): 253–274, November 15, 2002.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/202/25/
43/abstract.html.

Cain:2003:GHS

[Cai03] Bryan E. Cain. Group homomorphisms h_{i} such that h_{1}(x_{1}) + ⋯ + h_{n}(x_{n}) = y is solvable. Linear Algebra and its Applications, 360(C): 191–195, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
36/abstract.html.

Caldeira:2005:GDR

[Cal05a] Cristina Caldeira. Generalized derivations restricted to Grassmann spaces and additive theory. Linear Algebra and its Applications, 401(1): 11–27, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Calinski:2005:SJB

[Cal05b] Tadeusz Caliński. On some of Jerzy Baksalary’s contributions to the theory of block designs. Linear Algebra and its Applications, 410(1):54– 63, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Caldeira:2007:COD

[Cal07]  Cristina Caldeira.  Criti-

cal operators for the degree of the minimal polynomial of derivations restricted to Grassmann spaces. Linear Algebra and its Applications, 424(2–3):492–509, July 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Calbeck:2008:ENR

[Cal08] William Calbeck. Elliptic numerical ranges of 3 × 3 companion matrices. Linear Algebra and its Applications, 428(11–12):2715– 2722, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cao:2001:ALA

[Cao01]  You’an Cao.  Automor-

phisms of the Lie algebra of strictly upper triangular matrices over certain commutative rings. Linear Algebra and its Applications, 329(1–3): 175–187, May 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej%2Dng/10/
                                        30/19/154/25/35/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej%2Dng/10/30/19/154/
                                        25/35/article.pdf.

Cao:2005:WTA

[Cao05] Xiaohong Cao. Weyl’s theorem for analytically hyponormal operators. Linear Algebra and its Applications, 405 (1):229–238, August 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cao:2007:BEA

[Cao07a]  Xiaohong Cao.  Browder

essential approximate point spectra and hypercyclicity for operator matrices. Linear Algebra and its Applications, 426(2–3):317–324, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cao:2007:TUD

[Cao07b] Xiaohong Cao. Topological uniform descent and Weyl type theorem. Linear Algebra and its Applications, 420 (1):175–182, January 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cao:2009:SNF

[Cao09] Wei Cao. Smith normal form of augmented degree matrix and its applications. Linear Algebra and its Applications, 431(10):1778–1784, October 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Capizzano:2001:DRA

[Cap01a] S. Serra Capizzano. Distribution results on the algebra generated by Toeplitz sequences: a finite-dimensional approach. Linear Algebra and its Applications, 328(1– 3):121–130, May 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/153/25/29/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/153/
25/29/article.pdf.

Capizzano:2001:MAP

[Cap01b] Stefano Serra Capizzano. Matrix algebra preconditioners for multilevel Toeplitz matrices are not superlinear. Linear Algebra and its Applications, 343–344(1): 303–319, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
40/abstract.html.

Capizzano:2001:SBM

[Cap01c] Stefano Serra Capizzano. Spectral behavior of matrix sequences and discretized boundary value problems. Linear Algebra and its Applications, 337(1–3):37–78, November 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/167/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/167/27/29/article.
pdf.

Capizzano:2003:GLT

[Cap03] S. Serra Capizzano. Generalized locally Toeplitz sequences: spectral analysis and applications to discretized partial differential equations. Linear Algebra and its Applications, 366 (C):371–402, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
46/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Carroll:2004:BCC

[Car04] Joseph E. Carroll. Birkhoff’s contraction coefficient. Linear Algebra and its Applications, 389(1):227–234, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Carlson:2009:HAG

[Car09] Robert Carlson. Harmonic analysis for graph refinements and the continuous graph FFT. Linear Algebra and its Applications, 430(11– 12):2859–2876, June 1, 2009.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cherniavsky:2006:PRI

[CB06] Yona Cherniavsky and Eli Bagno. Permutation representations on invertible matrices. Linear Algebra and its Applications, 419(2–3):494– 518, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chu:2004:SSS

[CBD^{+}04] M. Chu, N. Del Buono, F. Diele, T. Politi, and S. Ragni. On the semigroup of standard symplectic matrices and its applications. Linear Algebra and its Applications, 389(1):215–225, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Censor:2008:LIA

[CBIXG08] Yair Censor, Adi Ben-Israel, Ying Xiao, and James M. Galvin. On linear infeasibility arising in intensitymodulated radiation therapy inverse planning. Linear Algebra and its Applications, 428(5–6):1406–1420, March 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2000:RAO

[CC00] Yong-Lin Chen and Xin Chen. Representation and approximation of the outer inverse A_{T,S}^{(2)} of a matrix A. Linear Algebra and its Applications, 308(1–3): 85–107, March 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/125/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/125/25/29/article.
pdf.

Chan:2001:CNN

[CC01] Jor-Ting Chan and Kong Chan. The cone of nonnegative c-numerical range and its preservers. Linear Algebra and its Applications, 323(1– 3):75–85, January 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cavazos-Cadena:2003:ADB

[CC03] Rolando Cavazos-Cadena. An alternative derivation of Birkhoff’s formula for the contraction coefficient of a positive matrix. Linear Algebra and its Applications, 375(1):291–297, December 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2004:PTT

[CC04] Huanyin Chen and Miaosen Chen. On products of three triangular matrices over associative rings. Linear Algebra and its Applications, 387 (1):297–311, August 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cuadras:2006:PAC

[CC06] Carles M. Cuadras and Daniel Cuadras. A parametric approach to correspondence analysis. Linear Algebra and its Applications, 417 (1):64–74, August 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cai:2009:SIH

[CC09] Jing Cai and Guoliang Chen. Some investigation on Hermitian positive definite solutions of the matrix equation X^{s} + A^{∗}X^{−t}A = Q. Linear Algebra and its Applications, 430(8–9):2448– 2456, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chin:2001:LCA

[CCG01] William Chin, Barbara Cortzen, and Jerry Goldman. Linear cellular automata with boundary conditions. Linear Algebra and its Applications, 322(1–3):193–206, January 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/141/24/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/141/24/34/article.
pdf.

Cechlarova:2002:ISM

[CCG02] Katarína Cechlárová and R. A. Cuninghame-Green. Interval systems of maxseparable linear equations. Linear Algebra and its Applications, 340(1–3):215–224, January 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
42/abstract.html.

Cao:2002:CPL

[CCH02a] You’an Cao, Zhonghu Chen, and Chong Huang. Commutativity preserving linear maps and Lie automorphisms of strictly triangular matrix space. Linear Algebra and its Applications, 350(1–3): 41–66, July 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cho:2002:TOS

[CCH02b] Muneo Chō, Raúl E. Curto, and Tadasi Huruya. ntuples of operators satisfying _{T} (AB) =_{T} (BA). Linear Algebra and its Applications, 341(1–3):291–298, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/171/27/
                                        49/abstract.html.

Codenotti:2000:MRC

[CCM00] Bruno Codenotti, Gianna Del Corso, and Giovanni Manzini. Matrix rank and communication complexity. Linear Algebra and its Applications, 304(1–3):193–200, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/118/25/
                                        38/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/118/25/38/article.
                                        pdf.

Cardoso:2008:SLB

[CCRS08] Domingos M. Cardoso, Dragoš Cvetković, Peter Rowlinson, and Slobodan K. Simić. A sharp lower bound for the least eigenvalue of the signless Laplacian of a nonbipartite graph. Linear Algebra and its Applications, 429(11–12):2770–2780, December 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chan:2003:WDA

[CCSS03] Raymond H. Chan, Tony F.

Chan, Lixin Shen, and Zuowei Shen. Wavelet deblurring algorithms for spatially varying blur from highresolution image reconstruction. Linear Algebra and its Applications, 366(C):139– 155, June 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
34/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Chaitin-Chatelin:2000:AAE

[CCTT00] F. Chaitin-Chatelin, V. Toumazou, and E. Traviesas. Accuracy assessment for eigencomputations: Variety of backward errors and pseudospectra. Linear Algebra and its Applications, 309(1– 3):73–83, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/30/article.
pdf.

Clark:2008:ITP

[CCW^{+}08] V. H. Clark, Y. Chen, J. Wilkens, J. R. Alaly, K. Zakaryan, and J. O. Deasy. IMRT treatment planning for prostate cancer using prioritized prescription optimization and mean-tail-dose functions. Linear Algebra and its Applications, 428(5– 6):1345–1364, March 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chou:2006:RCL

[CCZ06] Jyh-Horng Chou, Shinn-Horng Chen, and Qing-Ling Zhang. Robust controllability for linear uncertain descriptor systems. Linear Algebra and its Applications, 414(2– 3):632–651, April 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Caldeira:2000:IPD

[CD00a] Cristina Caldeira and J. A. Dias da Silva. The invariant polynomials degrees of the Kronecker sum of two linear operators and additive theory. Linear Algebra and its Applications, 315(1– 3):125–138, August 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/132/25/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/132/25/31/article.
                                        pdf.

Chen:2000:BSM

[CD00b] Tzu-Yi Chen and James W. Demmel. Balancing sparse matrices for computing eigenvalues. Linear Algebra and its Applications, 309(1–3): 261–287, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/38/article.
pdf.

Chu:2000:VFQ

[CD00c] Delin Chu and Bart De Moor. On a variational formulation of the QSVD and the RSVD. Linear Algebra and its Applications, 311 (1–3):61–78, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/128/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/128/25/30/article.
pdf.

Chu:2000:NFF

[CD00d] Delin Chu and Bart De Moor. On the nonuniqueness of the factorization factors in the product singular value decomposition. Linear Algebra and its Applications, 314 (1–3):191–203, July 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/131/25/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/131/25/31/article.
                                        pdf.

Das:2003:IUB

[cD03] Kinkar ch. Das. An improved upper bound for Laplacian graph eigenvalues. Linear Algebra and its Applications, 368(C):269–278, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/217/25/
                                        42/abstract.html.

Constantine:2005:LFD

[CD05] D. Constantine and M. Darnall. Lengths of finite dimensional representations of PBW algebras. Linear Algebra and its Applications, 395(1):175–181, January 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Coelho:2006:IPI

[CD06] M. Purificação Coelho and M. Antónia Duffner. Immanant preserving and immanant converting maps. Linear Algebra and its Applications, 418(1):177–187, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Comellas:2008:SMS

[CDFM08a] F. Comellas, C. Dalfó, M. A. Fiol, and M. Mitjana. The spectra of Manhattan street networks. Linear Algebra and its Applications, 429(7): 1823–1839, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Costa:2008:DAB

[CdFM08b] Liliana Costa, C. M. da Fonseca, and Enide Andrade Martins. The diameter of the acyclic Birkhoff polytope. Linear Algebra and its Applications, 428(7):1524– 1537, April 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Costa:2009:FCA

[CdFM09] Liliana Costa, C. M. da Fonseca, and Enide Andrade Martins. Face counting on an Acyclic Birkhoff polytope. Linear Algebra and its Applications, 430(4):1216–1235, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Coakley:2008:MOG

[CDJ08] Edward S. Coakley, Froilán M. Dopico, and Charles R. Johnson. Matrices with orthogonal groups admitting only determinant one. Linear Algebra and its Applications, 428(4):796–813, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheung:2005:MPI

[CDL05] Wai-Shun Cheung, M. Antonia Duffner, and Chi-Kwong Li. Multiplicative preservers and induced operators. Linear Algebra and its Applications, 401(1):307– 323, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Carriegos:2006:HOL

[CDMGP06] Miguel V. Carriegos, Héctor Diez-Machío, and M. Isabel García-Planas. On higher order linear systems: Reachability and feedback invariants. Linear Algebra and its Applications, 413(2–3):285– 296, March 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Curgus:2001:LBE

[ĆDR01] Branko Ćurgus, Aad Dijksma, and Tom Read. The linearization of boundary eigenvalue problems and reproducing kernel Hilbert spaces. Linear Algebra and its Applications, 329(1–3): 97–136, May 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/

30/19/154/25/31/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/154/
25/31/article.pdf.

Cravo:2001:CPC

[CDS01] G. Cravo, J. A. Dias da Silva, and Fernando C. Silva. Characteristic polynomials and controlability of partially prescribed matrices. Linear Algebra and its Applications, 335(1–3):157–166, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/165/27/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/165/27/36/article.
pdf.

Chu:2004:GFM

[CDS04] Moody T. Chu, Fasma Diele, and Ivonne Sgura. Gradient flow methods for matrix completion with prescribed eigenvalues. Linear Algebra and its Applications, 379(1):85– 112, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chou:2008:NCT

[CDS08] Wun-Seng Chou, Bau-Sen Du, and Peter J.-S. Shiue. A note on circulant transition matrices in Markov chains. Linear Algebra and its Applications, 429(7):1699–1704, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheon:2009:RBS

[CE09] Gi-Sang Cheon and Andrew W. Eckford. A relationship between subpermanents and the arithmetic–geometric mean inequality. Linear Algebra and its Applications, 430 (1):114–120, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cechlarova:2000:NUS

[Cec00]  Katarína Cechlárová.  A

note on unsolvable systems of max-min (fuzzy) equations. Linear Algebra and its Applications, 310(1–3): 123–128, May 1, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/127/25/
                                        35/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/127/25/35/article.
                                        pdf.

Cegielski:2001:OCG

[Ceg01]  Andrzej Cegielski.  Ob-

tuse cones and Gram matrices with nonnegative inverse. Linear Algebra and its Applications, 335(1–3): 167–181, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/165/27/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/165/27/37/article.
pdf.

Chen:2001:EMP

[CEK^{+}01] Li Chen, Wayne Eberly, Erich Kaltofen, B. David Saunders, William J. Turner, and Gilles Villard. Efficient matrix preconditioners for black box linear algebra. Linear Algebra and its Applications, 343–344(1): 119–146, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
32/abstract.html.

Censor:2001:BSD

[Cen01] Yair Censor. Binary steering in discrete tomography reconstruction with sequential and simultaneous iterative algorithms. Linear Algebra and its Applications, 339(1–3): 111–124, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/169/27/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/169/27/34/article.
pdf.

Censor:2006:CAP

[Cen06]  Yair Censor.  Computa-

tional acceleration of projection algorithms for the linear best approximation problem. Linear Algebra and its Applications, 416(1):111– 123, July 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cerny:2009:COP

[Čer09] Aleš Černý. Characterization of the oblique projector U(V U)^{†}V with application to constrained least squares. Linear Algebra and its Applications, 431(9):1564–1570, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Compta:2002:MRS

[CF02] Albert Compta and Josep Ferrer. Matricial realizations of the solutions of the Carlson problem. Linear Algebra and its Applications, 353(1–3): 197–206, September 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Crans:2001:HCT

[CFJ01] Alissa S. Crans, Shaun M. Fallat, and Charles R. Johnson. The Hadamard core of the totally nonnegative matrices. Linear Algebra and its Applications, 328(1–3):

203–222, May 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/153/25/33/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/153/
25/33/article.pdf.

Chistov:2008:CFT

[CFKP08] Alexander Chistov, Hervé Fournier, Pascal Koiran, and Sylvain Perifel. On the construction of a family of transversal subspaces over finite fields. Linear Algebra and its Applications, 429(2– 3):589–600, July 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetkovic:2008:P

[CFKS08] Ljiljana Cvetković, Andreas Frommer, Lilia Kolotilina, and Daniel B. Szyld. Preface. Linear Algebra and its Applications, 429(10):2291– 2292, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheung:2002:MPS

[CFL02] Wai-Shun Cheung, Shaun Fallat, and Chi-Kwong Li. Multiplicative preservers on semigroups of matrices. Linear Algebra and its Applications, 355(1–3):173–186, November 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/201/27/
                                        38/abstract.html.

Chebotar:2005:MPZ

[CFL05a] Mikhail A. Chebotar, Yuen Fong, and Pjek-Hwee Lee. On maps preserving zeros of the polynomial xy − yx∗. Linear Algebra and its Applications, 408(1):230–243, October 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chu:2005:SPD

[CFL05b] E. K.-W. Chu, H.-Y. Fan, and W.-W. Lin. A structurepreserving doubling algorithm for continuous-time algebraic Riccati equations. Linear Algebra and its Applications, 396(1):55–80, February 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Casciola:2003:AFF

[CFM03] G. Casciola, F. Fabbri, and L. B. Montefusco. An application of fast factorization algorithms in Computer Aided Geometric Design. Linear Algebra and its Applications, 366(C):121–138, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/

gej-ng/10/30/19/215/25/
33/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Carriazo:2004:CSA

[CFN04] A. Carriazo, L. M. Fernández, and J. Núñez. Combinatorial structures associated with Lie algebras of finite dimension. Linear Algebra and its Applications, 389(1):43– 61, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chu:2003:SLR

[CFP03a] Moody T. Chu, Robert E. Funderlic, and Robert J. Plemmons. Structured low rank approximation. Linear Algebra and its Applications, 366(C):157–172, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
35/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Compta:2003:MDM

[CFP03b] Albert Compta, Josep Ferrer, and Ferran Puerta. Miniversal deformations of marked matrices. Linear Algebra and its Applications, 361 (C):181–201, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/210/25/
                                        39/abstract.html.

Compta:2004:DOM

[CFP04] Albert Compta, Josep Ferrer, and Marta Peña. Dimension of the orbit of marked subspaces. Linear Algebra and its Applications, 379(1):239– 248, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Compta:2009:URF

[CFP09] Albert Compta, Josep Ferrer, and Marta Peña. Use of reduced forms in the disturbance decoupling problem. Linear Algebra and its Applications, 430(5–6):1574– 1589, March 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetkovic:2002:CFG

[CFRS02] Dragoš Cvetković, Patrick Fowler, Peter Rowlinson, and Dragan Stevanović. Constructing fullerene graphs from their eigenvalues and angles. Linear Algebra and its Applications, 356(1–3): 37–56, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/

Campanella:2000:SSR

[CG00a] Matteo Campanella and Giovanni Garbo. The structure of the state representation of shift invariant controllable and observable group codes. Linear Algebra and its Applications, 319(1–3): 163–178, November 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/138/24/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/138/24/34/article.
pdf.

Chandrasekaran:2000:FSE

[CG00b] S. Chandrasekaran and M. Gu. Fast and stable eigendecomposition of symmetric banded plus semi-separable matrices. Linear Algebra and its Applications, 313(1– 3):107–114, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/130/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/130/25/30/article.
pdf.

Cioaba:2007:LME

[CG07a] Sebastian M. Cioabă and David A. Gregory. Large matchings from eigenvalues. Linear Algebra and its Applications, 422(1):308– 317, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Comellas:2007:SBB

[CG07b] F. Comellas and S. Gago. Spectral bounds for the betweenness of a graph. Linear Algebra and its Applications, 423(1):74–80, May 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cuninghame-Green:2004:BMA

[CGB04] R. A. Cuninghame-Green and P. Butkovič. Bases in max-algebra. Linear Algebra and its Applications, 389(1): 107–120, September 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Castro-Gonzalez:2005:RDI

[CGD05] N. Castro-González and E. Dopazo. Representations of the Drazin inverse for a class of block matrices. Linear Algebra and its Applications, 400(1):253– 269, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Catalisano:2002:RTS

[CGG02] M. V. Catalisano, A. V. Geramita, and A. Gimigliano.

Ranks of tensors, secant varieties of Segre varieties and fat points. Linear Algebra and its Applications, 355(1–3): 263–285, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/201/27/
43/abstract.html. See er-

ratum [Ano03-31].

Chapman:2001:CEK

[CGGGSR01] S. T. Chapman, J. I. García-García, P. A. García-Sánchez, and J. C. Rosales. Computing the elasticity of a Krull monoid. Linear Algebra and its Applications, 336 (1–3):191–200, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/166/27/39/article.
pdf.

Cui:2004:RNL

[CGH04] Jianlian Cui, Maozheng Guo, and Jinchuan Hou. Rank-1 nilpotency linear preservers on nest algebras. Linear Algebra and its Applications, 377(1):267–290, January 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Caballero-Gil:2006:SSB

[CGHG06] P. Caballero-Gil and C. Hernández-Goya. Secret sharing based on a hard-on-average problem. Linear Algebra and its Applications, 414(2–3):626– 631, April 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cao:2005:WTU

[CGM05] Xiaohong Cao, Maozheng Guo, and Bin Meng. Weyl’s theorem for upper triangular operator matrices. Linear Algebra and its Applications, 402(1):61–73, June 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Camacho:2001:ADL

[CGN01] L. M. Camacho, J. R. Gómez, and R. M. Navarro. Algebra of derivations of Lie algebras. Linear Algebra and its Applications, 332–334(1): 371–380, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/157/25/
                                        43/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/157/25/43/article.
                                        pdf.

Carriegos:2004:MIM

[CGP04] Miguel V. Carriegos and M. Isabel García-Planas. On matrix inverses modulo a subspace. Linear Algebra and its Applications, 379(1):229– 237, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Clotet:2001:EDQ

[CGPM01] Josep Clotet, M. Isabel García-Planas, and M. Dolors Magret. Estimating distances from quadruples satisfying stability properties to quadruples not satisfying them. Linear Algebra and its Applications, 332–334 (1):541–567, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
54/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/54/article.
pdf.

Cohen:2004:DST

[CGQ04] Guy Cohen, Stéphane Gaubert, and Jean-Pierre Quadrat. Duality and separation theorems in idempotent semimodules. Linear Algebra and its Applications, 379(1):395– 422, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Conti:2009:SSM

[CGR09] Costanza Conti, Luca Gemignani, and Lucia Romani. From symmetric subdivision masks of Hurwitz type to interpolatory subdivision masks. Linear Algebra and its Applications, 431(10):1971–1987, October 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Censor:2008:LNM

[CGXL08] Yair Censor, James M. Galvin, Ying Xiao, and Mark Langer. Linear and nonlinear models and algorithms in intensity-modulated radiation therapy (IMRT). Linear Algebra and its Applications, 428(5–6):1203–1206, March 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2000:NPI

[CH00a] Gong-Ning Chen and Yong-Jian Hu. The Nevanlinna– Pick interpolation problems and power moment problems for matrix-valued functions III: The infinitely many data case. Linear Algebra and its Applications, 306(1– 3):59–86, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/121/25/31/article.
pdf.

Chu:2000:NSS

[CH00b] Delin Chu and Y. S. Hung. A numerical solution for the simultaneous disturbance decoupling and row-by-row decoupling problem. Linear Algebra and its Applications, 320(1–3):37–49, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/27/article.
pdf.

Chen:2001:MNP

[CH01] Gong-Ning Chen and Yong-Jian Hu. Multiple Nevanlinna– Pick interpolation with both interior and boundary data and its connection with the power moment problem. Linear Algebra and its Applications, 323(1–3):167–194, January 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:2003:OTT

[CH03] An Chang and Qiongxiang Huang. Ordering trees by their largest eigenvalues. Linear Algebra and its Applications, 370(1):175–184, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Coppersmith:2005:SM

[CH05] Don Coppersmith and Alan J. Hoffman. On the singularity of matrices. Linear Algebra and its Applications, 411(1): 277–280, December 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chesnokov:2006:RGA

[CH06] Andrey A. Chesnokov and Willem H. Haemers. Regularity and the generalized adjacency spectra of graphs. Linear Algebra and its Applications, 416(2–3):1033– 1037, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:2007:UBS

[CH07] Guang-Hui Cheng and Ting-Zhu Huang. An upper bound for ||A^{−1}||_{∞} of strictly diagonally dominant M-matrices. Linear Algebra and its Applications, 426(2–3):667–673, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cui:2008:MLF

[CH08] Jianlian Cui and Jinchuan Hou. Maps leaving functional values of operator products invariant. Linear Algebra and its Applications, 428(7): 1649–1663, April 1, 2008. CODEN LAAPAW. ISSN

0024-3795 (print), 1873-1856 (electronic).

Carriegos:2002:PSL

[CHASG02] Miguel Carriegos, José A. Hermida-Alonso, and Tomás Sánchez-Giralda. Poleshifting for linear systems over commutative rings. Linear Algebra and its Applications, 346(1–3):97–107, May 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/189/27/
32/abstract.html.

Cheng:2002:QLK

[Che02a] Charles Ching-An Cheng. Quadratic linear Keller maps. Linear Algebra and its Applications, 348(1–3):203–207, June 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
42/abstract.html.

Chern:2002:STL

[Che02b]  Shuh-Jye Chern.  Stabil-

ity theory for linear dissipative Hamiltonian systems. Linear Algebra and its Applications, 357(1–3): 143–162, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
35/abstract.html.

Chen:2003:SDI

[Che03] Shencan Chen. Some determinantal inequalities for Hadamard product of matrices. Linear Algebra and its Applications, 368(C):99– 106, July 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/217/25/
                                        31/abstract.html.

Chen:2004:LBM

[Che04a] Shencan Chen. A lower bound for the minimum eigenvalue of the Hadamard product of matrices. Linear Algebra and its Applications, 378(1):159–166, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2004:PCH

[Che04b] Shencan Chen. A property concerning the Hadamard powers of inverse M-matrices. Linear Algebra and its Applications, 381(1):53–60, April 1, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2005:MPT

[Che05] Ke Chen. Matrix preconditioning techniques and applications, volume 19 of Cambridge monographs on applied and computational mathematics. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-83828-

2. xxiii + 568 + 8 pp. LCCN QA188 .C455 2005.

URL http://www.loc.gov/
catdir/enhancements/fy0632/
2006271080-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0632/2006271080-t.html;
 http://www.loc.gov/catdir/
enhancements/fy0733/2006271080-
b.html.

Chen:2007:PCE

[Che07a] Jiansheng Chen. The proof on the conjecture of extremal graphs for the kth eigenvalues of trees. Linear Algebra and its Applications, 426(1): 12–21, October 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2007:IMI

[Che07b] Shencan Chen. Inequalities for M-matrices and inverse M-matrices. Linear Algebra and its Applications, 426(2–3):610–618, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2007:PCC

[Che07c] Shencan Chen. Proof of a conjecture concerning the Hadamard powers of inverse M-matrices. Linear Algebra and its Applications, 422(2– 3):477–481, April 15, 2007.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chebotar:2008:PBC

[Che08] Mikhail Chebotar. On a problem by Beidar concerning the central closure. Linear Algebra and its Applications, 429(4):835–840, August 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2009:GBF

[Che09] Sheng Chen. Generalized Bowen–Franks groups of integral matrices with the same zeta function. Linear Algebra and its Applications, 431(8):1397–1406, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chikuse:2002:MDE

[Chi02]  Yasuko Chikuse.  Meth-

ods of density estimation on the Grassmann manifold. Linear Algebra and its Applications, 354(1–3): 85–102, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/200/27/
                                        35/abstract.html.

Chikuse:2004:HLP

[Chi04] Yasuko Chikuse. Hermite and Laguerre polynomials with complex matrix arguments. Linear Algebra and its Applications, 388(1):91– 105, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Correa:2005:NLT

[CHJP05] Ivan Correa, Irvin Roy Hentzel, Pedro Pablo Julca, and Luiz Antonio Peresi. Nilpotent linear transformations and the solvability of power-associative nilalgebras. Linear Algebra and its Applications, 396(1):35– 53, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cho:2001:AWM

[CHK^{+}01] Han Hyuk Cho, Chan Yong Hong, Suh-Ryung Kim, Chang Hoon Park, and Yunsun Nam. On A(R, S) all of whose members are indecomposable. Linear Algebra and its Applications, 332–334 (1):119–129, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/32/article.
pdf.

Choi:2005:IST

[CHKS05] Gyoung-Sik Choi, Suk-Geun Hwang, Ik-Pyo Kim, and Bryan L. Shader. (±1)- invariant sequences and truncated Fibonacci sequences. Linear Algebra and its Applications, 395(1):303–312, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cui:2001:LPU

[CHL01] Jianlian Cui, Jinchuan Hou, and Bingren Li. Linear preservers on upper triangular operator matrix algebras. Linear Algebra and its Applications, 336(1–3): 29–50, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/166/27/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/166/27/30/article.
                                        pdf.

Cheng:2002:IEC

[CHL02] Che-Man Cheng, Roger A. Horn, and Chi-Kwong Li. Inequalities and equalities for the Cartesian decomposition of complex matrices. Linear Algebra and its Applications, 341(1–3): 219–237, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
44/abstract.html.

Chlebowicz:2006:CFG

[Chl06] Agnieszka Chlebowicz. Certain finite groups as automorphism groups of forms of higher degree. Linear Algebra and its Applications, 419(2–3):326–330, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chowdhury:2006:RRM

[CHMM06] Atoshi Chowdhury, Leslie Hogben, Jude Melancon, and Rana Mikkelson. Rational realization of maximum eigenvalue multiplicity of symmetric tree sign patterns. Linear Algebra and its Applications, 418(2–3):380–393, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2006:SDS

[CHN06] Mei Q. Chen, Lixing Han, and Michael Neumann. On single and double Soules matrices. Linear Algebra and its Applications, 416(1):88– 110, July 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Catral:2004:RRN

[CHNP04] M. Catral, Lixing Han, Michael Neumann, and R. J. Plemmons. On reduced rank nonnegative matrix factorization for symmetric nonnegative matrices. Linear Algebra and its Applications, 393(1): 107–126, December 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Choudhry:2004:ERM

[Cho04] Ajai Choudhry. Extraction of n-th roots of 2 × 2 matrices. Linear Algebra and its Applications, 387(1):183– 192, August 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Correa:2003:SCP

[CHP03] Ivan Correa, Irvin Roy Hentzel, and Luiz Antonio Peresi. On the solvability of the commutative powerassociative nilalgebras of dimension 6. Linear Algebra and its Applications, 369 (C):185–192, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
35/abstract.html.

Climent:2006:PCC

[CHP06] Joan-Josep Climent, Victoria Herranz, and Carmen Perea. Positive cones and convergence conditions for iterative methods based on splittings. Linear Algebra and its Applications, 413(2–3):319– 326, March 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Climent:2007:FAC

[CHP07] Joan-Josep Climent, Victoria Herranz, and Carmen Perea. A first approximation of concatenated convolutional codes from linear systems theory viewpoint. Linear Algebra and its Applications, 425(2–3):673–699, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Climent:2008:LSM

[CHP08] Joan-Josep Climent, Victoria Herranz, and Carmen Perea. Linear system modelization of concatenated block and convolutional codes. Linear Algebra and its Applications, 429(5–6):1191–1212, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Compta:2006:SVD

[CHPP06] Albert Compta, Uwe Helmke, Marta Peña, and Xavier Puerta. Simultaneous versal deformations of endomorphisms and invariant subspaces. Linear Algebra and its Applications, 413(2–3): 303–318, March 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheon:2003:SOM

[CHR^{+}03] Gi-Sang Cheon, Suk-Geun Hwang, Seog-Hoon Rim, Bryan L. Shader, and Seok-Zun Song. Sparse orthogonal matrices. Linear Algebra and its Applications, 373(1):211– 222, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetkovic:2007:P

[CHR07] D. Cvetković, W. Haemers, and P. Rowlinson. Preface. Linear Algebra and its Applications, 423(1):1– 2, May 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheon:2003:MDL

[CHRS03] Gi-Sang Cheon, Suk-Geun Hwang, Seog-Hoon Rim, and Seok-Zun Song. Matrices determined by a linear recurrence relation among entries. Linear Algebra and its Applications, 373(1):89– 99, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheon:2003:PWM

[CHSS03] Gi-Sang Cheon, Suk-Geun Hwang, Bryan L. Shader, and Seok-Zun Song. Permanents of woven matrices. Linear Algebra and its Applications, 364(C):223–233, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
36/abstract.html.

Chu:2003:SMN

[Chu03] Teresa H. Chu. On semimonotone matrices with nonnegative principal minors. Linear Algebra and its Applications, 367(C):147– 154, July 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
35/abstract.html.

Chung:2007:RWL

[Chu07] Fan Chung. Random walks and local cuts in graphs. Linear Algebra and its Applications, 423(1):22–32, May 1, 2007. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chu:2008:GHM

[Chu08]  Teresa H. Chu.  Gener-

alizations of the hidden Minkowski property. Linear Algebra and its Applications, 429(8–9):2076–2088, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chillag:2005:PNM

[CHY05] D. Chillag, R. Holzman, and I. Yona. Primitive normal matrices and covering numbers of finite groups. Linear Algebra and its Applications, 403(1):165–177, July 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:2008:RC

[CHY08] Gerard J. Chang, Liang-Hao Huang, and Hong-Gwa Yeh. On the rank of a cograph. Linear Algebra and its Applications, 429(2–3):601– 605, July 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetkovic-Ilic:2008:NRD

[CI08] Dragana S. Cvetković-Ilić. A note on the representation for the Drazin inverse of 2 × 2 block matrices. Linear Algebra and its Applications, 429(1):242–248, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chugunov:2009:CNT

[CI09a] V. N. Chugunov and Kh. D. Ikramov. The conjugatenormal Toeplitz problem. Linear Algebra and its Applications, 430(8–9):2467– 2473, April 15, 2009. CO-

Chugunov:2009:CNH

[CI09b] V. N. Chugunov and Kh. D. Ikramov. A contribution to the normal Hankel problem. Linear Algebra and its Applications, 430(8–9):2094– 2101, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetkovic-Ilic:2009:GDI

[CI09c] Dragana S. Cvetković-Ilić. The generalized Drazin inverse with commutativity up to a factor in a Banach algebra. Linear Algebra and its Applications, 431(5–7):783– 791, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetkovic-Ilic:2006:ARG

[CIDW06] Dragana S. Cvetković-Ilić, Dragan S. Djordjević, and Yimin Wei. Additive results for the generalized Drazin inverse in a Banach algebra. Linear Algebra and its Applications, 418(1):53– 61, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cigler:2001:MGI

[Cig01] Grega Cigler. Matrix groups with independent spectra.

Linear Algebra and its Applications, 327(1–3):27–40, April 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/27/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/152/
25/27/article.pdf.

Cigler:2007:PLM

[Cig07] Grega Cigler. Permutationlike matrix groups. Linear Algebra and its Applications, 422(2–3):486–505, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cioaba:2006:EGS

[Cio06] Sebastian M. Cioabă. Eigenvalues of graphs and a simple proof of a theorem of Greenberg. Linear Algebra and its Applications, 416(2– 3):776–782, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cai:2004:NCP

[CJ04] Ming-Chao Cai and Xiao-Qing Jin. A note on T. Chan’s preconditioner. Linear Algebra and its Applications, 376(1):283–290, January 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:2005:SSP

[CJ05] Che-Man Cheng and Xiao-Qing Jin. Some stability properties of T. Chan’s preconditioner. Linear Algebra and its Applications, 395(1): 361–365, January 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cho:2005:PIP

[CJK05] Han Hyuk Cho, Jong Sam Jeon, and Hwa Kyung Kim. Periodic, irreducible, powerful ray pattern matrices. Linear Algebra and its Applications, 404(1):283–296, July 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Castillo:2002:OSS

[CJPS02] E. Castillo, F. Jubete, R. E. Pruneda, and C. Solares. Obtaining simultaneous solutions of linear subsystems of inequalities and duals. Linear Algebra and its Applications, 346(1–3):131–154, May 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/189/27/
34/abstract.html.

Cheng:2007:NSO

[CJVW07] Che-Man Cheng, Xiao-Qing Jin, Seak-Weng Vong, and Wei Wang. A note on spectra of optimal and superoptimal preconditioned matrices. Linear Algebra and its Applications, 422(2–3):482– 485, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cai:2005:GCP

[CJW05] Ming-Chao Cai, Xiao-Qing Jin, and Yi-Min Wei. A generalization of T. Chan’s preconditioner. Linear Algebra and its Applications, 407(1): 11–18, September 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheon:2001:SMP

[CK01] Gi-Sang Cheon and Jin-Soo Kim. Stirling matrix via Pascal matrix. Linear Algebra and its Applications, 329(1– 3):49–59, May 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/154/25/27/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/154/
25/27/article.pdf.

Cheon:2002:FSM

[CK02] Gi-Sang Cheon and Jin-Soo Kim. Factorial Stirling matrix and related combinatorial sequences. Linear Algebra and its Applications, 357 (1–3):247–258, December 15, 2002. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
42/abstract.html.

Cho:2003:FMS

[CK03] Han Hyuk Cho and Suh-Ryung Kim. Factorizations of matrices over semirings. Linear Algebra and its Applications, 373(1):289– 296, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Colombo:2004:CPM

[CK04] Jones Colombo and Plamen Koshlukov. Central polynomials in the matrix algebra of order two. Linear Algebra and its Applications, 377(1): 53–67, January 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheon:2008:SPO

[CK08] Gi-Sang Cheon and Hana Kim. Simple proofs of open problems about the structure of involutions in the Riordan group. Linear Algebra and its Applications, 428(4):930– 940, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Choi:2008:AHA

[CKK08] Yuncherl Choi, Hyuk Kim, and Kyunghee Kim. Abelian Hessian algebra and commutative Frobenius algebra. Linear Algebra and its Applications, 428(8–9):2236– 2258, April 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cai:2009:SQI

[CKLX09] Yun-Feng Cai, Yuen-Cheng Kuo, Wen-Wei Lin, and Shu-Fang Xu. Solutions to a quadratic inverse eigenvalue problem. Linear Algebra and its Applications, 430(5– 6):1590–1606, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Choi:2005:ESN

[CKM05] Y. S. Choi, I. Koltracht, and P. J. McKenna. On eigenstructure of a nonlinear map in R^{n}. Linear Algebra and its Applications, 399(1):141– 155, April 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Choi:2002:GMC

[CKMS02] Y. S. Choi, I. Koltracht, P. J. McKenna, and N. Savytska. Global monotone convergence of Newton iteration for a nonlinear eigenproblem. Linear Algebra and its Applications, 357(1–3): 217–228, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
40/abstract.html.

Casazza:2006:RVR

[CKS06] Peter G. Casazza, Gitta Kutyniok, and Darrin Speegle. A redundant version of the Rado–Horn Theorem. Linear Algebra and its Applications, 418(1):1–10, October 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheon:2008:RGI

[CKS08] Gi-Sang Cheon, Hana Kim, and Louis W. Shapiro. Riordan group involutions. Linear Algebra and its Applications, 428(4):941–952, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2008:LBC

[CKW08] Chang-Pao Chen, Meng-Kuang Kuo, and Kuo-Zhong Wang. Lower bounds of Copson type for Nörlund matrices. Linear Algebra and its Applications, 428(8–9):1939– 1948, April 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cohen:2005:FIO

[CKZ05] Joel E. Cohen, Johannes H. B. Kemperman, and Gheorghe H. Zbăganu. A family of inequalities originating from coding of messages. Linear Algebra and its Applications, 395(1):1–82, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Choi:2006:HRN

[CKŻ06] Man-Duen Choi, David W. Kribs, and Karol Życzkowski. Higher-rank numerical ranges and compression problems. Linear Algebra and its Applications, 418(2–3):828–839, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:2000:HHT

[CL00a] Che-Man Cheng and Chi-Kwong Li. On the Hu-Hurley-Tam conjecture concerning the generalized numerical range. Linear Algebra and its Applications, 305(1– 3):87–97, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/120/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/120/25/31/article.
pdf.

Cohen:2000:CCC

[CL00b] Nir Cohen and Izchak Lewkowicz. A characterization of convex cones of matrices with constant regular inertia. Linear Algebra and its Applications, 318(1– 3):23–33, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/137/24/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/137/24/25/article.
pdf.

Cardoso:2001:TNC

[CL01a] J. R. Cardoso and F. Silva Leite. Theoretical and numerical considerations about Padé approximants for the matrix logarithm. Linear Algebra and its Applications, 330(1–3):31–42, June 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/155/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/155/25/28/article.
pdf.

Christensen:2001:FEL

[CL01b] Ole Christensen and Alexander M. Lindner. Frames of exponentials: lower frame bounds for finite subfamilies and approximation of the inverse frame operator. Linear Algebra and its Applications, 323(1–3):117–130, January 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chiang:2002:LML

[CL02a] Hanley Chiang and Chi-Kwong Li. Linear maps leaving the alternating group invariant. Linear Algebra and its Applications, 340(1– 3):69–80, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/172/27/
                                        31/abstract.html.

Chooi:2002:CIM

[CL02b] W. L. Chooi and M. H. Lim. Coherence invariant mappings on block triangular matrix spaces. Linear Algebra and its Applications, 346 (1–3):199–238, May 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/189/27/
                                        37/abstract.html.

Christensen:2002:DRF

[CL02c] Ole Christensen and Alexander M. Lindner. Decomposition of Riesz frames and wavelets into a finite union of linearly independent sets. Linear Algebra and its Applications, 355(1–3): 147–159, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

Cardoso:2003:MVE

[CL03a] J. R. Cardoso and F. Silva Leite. The Moser–Veselov equation. Linear Algebra and its Applications, 360(C): 237–248, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
41/abstract.html.

Chooi:2003:SLP

[CL03b] Wai-Leong Chooi and Ming-Huat Lim. Some linear preserver problems on block triangular matrix algebras. Linear Algebra and its Applications, 370(1):25– 39, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cohen:2003:PMS

[CL03c] Nir Cohen and Izchak Lewkowicz. A pair of matrices sharing common Lyapunov solutions: A closer look. Linear Algebra and its Applications, 360(C):83– 104, February 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
31/abstract.html.

Chen:2004:ESR

[CL04a] Mei-Qin Chen and Xiezhang Li. An estimation of the spectral radius of a product of block matrices. Linear Algebra and its Applications, 379 (1):267–275, March 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Choi:2004:NBS

[CL04b] Man-Duen Choi and Chi-Kwong Li. Norm bounds for summation of two normal matrices. Linear Algebra and its Applications, 379(1):137– 157, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2005:SPN

[CL05a] Mei-Qin Chen and Xiezhang Li. Spectral properties of a near-periodic row-stochastic Leslie matrix. Linear Algebra and its Applications, 409(1): 166–186, November 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chooi:2005:APR

[CL05b] Wai-Leong Chooi and Ming-Huat Lim. Additive preservers of rank-additivity on matrix spaces. Linear Algebra and its Applications, 402 (1):291–302, June 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chooi:2006:ARO

[CL06] W. L. Chooi and M. H. Lim. Additive rank-one preservers on block triangular matrix spaces. Linear Algebra and its Applications, 416 (2–3):588–607, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chebotar:2007:QGO

[CL07a] Mikhail Chebotar and Victor Lomonosov. On a question by Grunenfelder, Omladič and Radjavi. Linear Algebra and its Applications, 426(2–3): 368–370, October 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cohen:2007:CIC

[CL07b] Nir Cohen and Izchak Lewkowicz. Convex invertible cones and positive real analytic functions. Linear Algebra and its Applications, 425(2–3):797–813, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:2008:BSP

[CL08a] Bo Cheng and Bolian Liu. The base sets of primitive zero-symmetric sign pattern matrices. Linear Algebra and its Applications, 428(4):715– 731, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:2008:IMZ

[CL08b] Bo Cheng and Bolian Liu. Imprimitive matrices of zeros and ones with extremal numbers of ones. Linear Algebra and its Applications, 429(1): 293–301, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Carrieu:2009:OMS

[CL09a] Hervé Carrieu and Patrice Lassère. One more simple proof of the Craig–Sakamoto theorem. Linear Algebra and its Applications, 431(9): 1616–1619, October 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:2009:NPE

[CL09b] Bo Cheng and Bolian Liu. On the numbers of positive entries of reducible nonnegative matrices. Linear Algebra and its Applications, 430(1):308– 317, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cohen:2009:LOR

[CL09c] Nir Cohen and Izchak Lewkowicz. The Lyapunov order for real matrices. Linear Algebra and its Applications, 430(7):1849–1866, April 1, 2009. CODEN

Cui:2009:MPP

[CL09d] Jianlian Cui and Chi-Kwong Li. Maps preserving product XY − Y X^{∗} on factor von Neumann algebras. Linear Algebra and its Applications, 431(5–7):833–842, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chang:2009:EES

[CLB09] Hung-Wen Chang, Sen-Eon Liu, and Robert Burridge. Exact eigensystems for some matrices arising from discretizations. Linear Algebra and its Applications, 430(4): 999–1006, February 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Carroll:2005:PTP

[CLL05] Joseph E. Carroll, Timothy Lauck, and Roland H. Lamberson. A Perron Theorem for positive componentwise bilinear maps. Linear Algebra and its Applications, 396(1): 259–272, February 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:2007:EEO

[CLL07] Che-Man Cheng, Ieng-Chi Law, and Sok-I Leong. Eigenvalue equalities for ordinary and Hadamard products of powers of positive semidefinite matrices. Linear Algebra and its Applications, 422(2– 3):771–787, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheung:2001:OQF

[CLOvdD01] Wai-Shun Cheung, Chi-Kwong Li, D. D. Olesky, and P. van den Driessche. Optimizing quadratic forms of adjacency matrices of trees and related eigenvalue problems. Linear Algebra and its Applications, 325(1–3): 191–207, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/147/24/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/147/24/37/article.
pdf.

Chooi:2003:APM

[CLP03] W. L. Chooi, M. H. Lim, and PeterŠemrl. Adjacency preserving maps on upper triangular matrix algebras. Linear Algebra and its Applications, 367(C):105–130, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
33/abstract.html.

Canovas:2006:IPR

[CLPT06] M. J. Cánovas, M. A. López, J. Parra, and F. J. Toledo. Ill-posedness with respect to the solvability in linear optimization. Linear Algebra and its Applications, 416(2– 3):520–540, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Calvetti:2000:GTM

[CLR00] D. Calvetti, B. Lewis, and L. Reichel. GMRES-type methods for inconsistent systems. Linear Algebra and its Applications, 316(1–3): 157–169, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/34/article.
pdf.

Chan:2005:IUI

[CLS05] Jor-Ting Chan, Chi-Kwong Li, and Nung-Sing Sze. Isometries for unitarily invariant norms. Linear Algebra and its Applications, 399 (1):53–70, April 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chahlaoui:2006:SOB

[CLVV06] Y. Chahlaoui, D. Lemonnier, A. Vandendorpe, and P. Van Dooren. Secondorder balanced truncation. Linear Algebra and its Applications, 415(2–3):373–384, June 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chuang:2007:MPA

[CLY07] Chen-Lian Chuang, Tsiu-Kwen Lee, and Chi-Tsuen Yeh. Minimal polynomials of algebraic derivations and automorphisms. Linear Algebra and its Applications, 423 (2–3):339–350, June 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Curto:2008:WHV

[CLY08] Raúl E. Curto, Sang Hoon Lee, and Jasang Yoon. Which 2-hyponormal 2variable weighted shifts are subnormal? Linear Algebra and its Applications, 429(8–9):2227–2238, October 16, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cabrera:2000:RTA

[CM00a] M. Cabrera and Amir A. Mohammed. A representation theorem for algebras with commuting involutions. Linear Algebra and its Applications, 306(1–3): 25–31, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/121/25/27/article.
pdf.

Camara:2000:WHF

[CM00b] M. C. Câmara and M. T. Malheiro. Wiener–Hopf factorization for a group of exponentials of nilpotent matrices. Linear Algebra and its Applications, 320(1–3): 79–96, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/30/article.
pdf.

Cho:2000:MCS

[CM00c] Grace E. Cho and Carl D. Meyer. Markov chain sensitivity measured by mean first passage times. Linear Algebra and its Applications, 316(1– 3):21–28, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/26/article.
pdf.

Cho:2001:CPB

[CM01] Grace E. Cho and Carl D. Meyer. Comparison of perturbation bounds for the stationary distribution of a Markov chain. Linear Algebra and its Applications, 335 (1–3):137–150, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/165/27/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/165/27/34/article.
                                        pdf.

Clotet:2008:DOL

[CM08a] Josep Clotet and M. Dolors Magret. Dimension of orbits of linear time-invariant singular systems under restricted system equivalence. Linear Algebra and its Applications, 429(5–6):1102–1113, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetkovic:2008:PDS

[CM08b] Aleksandar S. Cvetković and Gradimir V. Milovanović. Positive definite solutions of some matrix equations. Linear Algebra and its Applications, 429(10):2401–2414, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Coelho:2009:ADE

[CM09] Carlos A. Coelho and Filipe J. Marques. The advantage of decomposing elaborate hypotheses on covariance matrices into conditionally independent hypotheses in building near-exact distributions for the test statistics. Linear Algebra and its Applications, 430(10):2592– 2606, May 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Clotet:2002:PQ

[CMP02] Josep Clotet, M. Dolors Magret, and Xavier Puerta. Perturbation of quadrics. Linear Algebra and its Applications, 355(1–3):223–239, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/201/27/
41/abstract.html.

Cachafeiro:2003:LPQ

[CMP03] A. Cachafeiro, F. Marcellán, and C. Pérez. Lebesgue perturbation of a quasi-definite Hermitian functional. The positive definite case. Linear Algebra and its Applications, 369(C):235–250, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
39/abstract.html.

Corach:2002:GSC

[CMS02] G. Corach, A. Maestripieri, and D. Stojanoff. Generalized Schur complements and oblique projections. Linear Algebra and its Applications, 341(1–3):259–272, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/171/27/
                                        47/abstract.html.

Cantero:2003:FDM

[CMV03] M. J. Cantero, L. Moral, and L. Velázquez. Five-diagonal matrices and zeros of orthogonal polynomials on the unit circle. Linear Algebra and its Applications, 362 (C):29–56, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/211/25/
                                        28/abstract.html.

Cantero:2005:MRU

[CMV05] M. J. Cantero, L. Moral, and L. Velázquez. Minimal representations of unitary operators and orthogonal polynomials on the unit circle. Linear Algebra and its Applications, 408(1):40– 65, October 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2005:BMF

[CMX05] Qiuhui Chen, Charles A. Micchelli, and Yuesheng Xu. Biorthogonal multivariate filter banks from centrally symmetric matrices. Linear Algebra and its Applications, 402 (1):111–125, June 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chien:2001:NRT

[CN01] Mao-Ting Chien and Hiroshi Nakazato. The c-numerical range of tridiagonal matrices. Linear Algebra and its Applications, 335(1–3): 55–61, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/165/27/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/165/27/28/article.
pdf.

Chien:2002:DWS

[CN02a] Mao-Ting Chien and Hiroshi Nakazato. Davis-Wielandt shell and q-numerical range. Linear Algebra and its Applications, 340(1–3):15–31, January 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
28/abstract.html.

Chien:2002:NRL

[CN02b] Mao-Ting Chien and Hiroshi Nakazato. The numerical range of linear pencils of 2by-2 matrices. Linear Algebra and its Applications, 341(1– 3):69–100, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/171/27/
                                        33/abstract.html.

Chien:2006:NRR

[CN06] Mao-Ting Chien and Hiroshi Nakazato. The q-numerical range of a reducible matrix via a normal operator. Linear Algebra and its Applications, 419(2–3):440–465, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chien:2007:NRW

[CN07] Mao-Ting Chien and Hiroshi Nakazato. The qnumerical radius of weighted shift operators with periodic weights. Linear Algebra and its Applications, 422(1):198– 218, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chien:2009:FPB

[CN09] Mao-Ting Chien and Hiroshi Nakazato. Flat portions on the boundary of the Davis– Wielandt shell of 3-by-3 matrices. Linear Algebra and its Applications, 430(1):204– 214, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ching:2008:HOM

[CNF08] Wai-Ki Ching, Michael K. Ng, and Eric S. Fung. Higherorder multivariate Markov chains and their applications. Linear Algebra and its Applications, 428(2–3):492–507, January 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chien:2002:PEB

[CNP02] Mao-Ting Chien, Hiroshi Nakazato, and Panayiotis Psarrakos. Point equation of the boundary of the numerical range of a matrix polynomial. Linear Algebra and its Applications, 347(1– 3):205–217, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
39/abstract.html.

Chen:2000:ESN

[CNR00] Yin Chen, Abdelkrim Nokrane, and Thomas Ransford. Estimates for the spectrum near algebraic elements. Linear Algebra and its Applications, 308(1–3):153–161, March 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/125/25/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/125/25/33/article.
                                        pdf.

Chen:2007:DSP

[CNSM07] Mei Q. Chen, Michael Neumann, and Naomi Shaked-Monderer. Double Soules pairs and matching Soules bases. Linear Algebra and its Applications, 421(2–3):315– 327, March 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2008:BSM

[CNSM08] Mei Q. Chen, Michael Neumann, and Naomi Shaked-Monderer. Basic Soules matrices and their applications. Linear Algebra and its Applications, 429(5–6):951–971, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ching:2008:PLS

[CNTV08] Wai-Ki Ching, Michael K. Ng, Eugene Tyrtyshnikov, and Marc Van Barel. Preface for LAA special issue on structured matrices (SM). Linear Algebra and its Applications, 428(2–3):401–402, January 15, 2008. CO-

Catral:2005:MAM

[CNX05a] Minerva Catral, Michael Neumann, and Jianhong Xu. Matrix analysis of a Markov chain small-world model. Linear Algebra and its Applications, 409(1):126–146, November 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Catral:2005:PGI

[CNX05b] Minerva Catral, Michael Neumann, and Jianhong Xu. Proximity in group inverses of M-matrices and inverses of diagonally dominant Mmatrices. Linear Algebra and its Applications, 409(1):32– 50, November 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chterental:2008:OUQ

[Co08] Oleg Chterental and Dragomir Ž. D. oković. On orthostochastic, unistochastic and qustochastic matrices. Linear Algebra and its Applications, 428(4):1178–1201, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chuang:2009:GTD

[CO09] H. Chuang and G. R. Omidi. Graphs with three distinct eigenvalues and largest eigenvalue less than 8. Linear Algebra and its Applications, 430(8–9):2053–2062, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Codenotti:2000:MR

[Cod00] Bruno Codenotti. Matrix rigidity. Linear Algebra and its Applications, 304(1–3): 181–192, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/118/25/
                                        37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/118/25/37/article.
                                        pdf.

Coffman:2003:RCC

[Cof03] Adam Coffman. Real congruence of complex matrix pencils and complex projections of real Veronese varieties. Linear Algebra and its Applications, 370(1):41– 83, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Conde:2008:GIS

[Con08] Cristian Conde. Geometric interpolation in symmetricallynormed ideals. Linear Algebra and its Applications, 429 (4):819–834, August 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Conde:2009:NOI

[Con09] Cristian Conde. Normal operators and inequalities in norm ideals. Linear Algebra and its Applications, 431(10): 1774–1777, October 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Costara:2007:CTT

[Cos07] Constantin Costara. A Cartan type theorem for finitedimensional algebras. Linear Algebra and its Applications, 426(2–3):299–304, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Catral:2009:APC

[COvdD09] M. Catral, D. D. Olesky, and P. van den Driessche. Allow problems concerning spectral properties of sign pattern matrices: A survey. Linear Algebra and its Applications, 430(11–12):3080– 3094, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Capizzano:2003:APS

[CP03] Stefano Serra Capizzano and Cristina Tablino Possio. Analysis of preconditioning strategies for collocation linear systems. Linear Algebra and its Applications, 369 (C):41–75, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
28/abstract.html.

Crabb:2005:DPS

[CP05] M. C. Crabb and A. J. B. Potter. Determinants and periodic solutions of delay equations. Linear Algebra and its Applications, 411(1):356– 363, December 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chaudhuri:2006:TST

[CP06] Sanjay Chaudhuri and Michael D. Perlman. Two step-down tests for equality of covariance matrices. Linear Algebra and its Applications, 417(1): 42–63, August 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Causin:2007:NSS

[CP07a] Andrea Causin and Gian Pietro Pirola. A note on spaces of symmetric matrices. Linear Algebra and its Applications, 426(2–3):533–539, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Compta:2007:SSM

[CP07b] Albert Compta and Marta Peña. Structural stability of (C, A)-marked subspaces. Linear Algebra and its Applications, 421(1):45– 52, February 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cortes:2007:SRM

[CP07c] V. Cortes and J. M. Peña. Sign regular matrices and Neville elimination. Linear Algebra and its Applications, 421(1):53–62, February 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cortes:2008:DSS

[CP08] V. Cortés and J. M. Peña. Decompositions of strictly sign regular matrices. Linear Algebra and its Applications, 429(5–6):1071–1081, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chugunova:2009:QEP

[CP09] Marina Chugunova and Dmitry Pelinovsky. On quadratic eigenvalue problems arising in stability of discrete vortices. Linear Algebra and its Applications, 431(5– 7):962–973, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chu:2000:MFL

[CPPS00] Moody T. Chu, Victor P. Pauca, Robert J. Plemmons, and Xiaobai Sun. A mathematical framework for the linear reconstructor problem in adaptive optics. Linear Algebra and its Applications, 316(1–3):113–135, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/32/article.
pdf.

Chalmers:2003:SCE

[CPR03] B. L. Chalmers, M. P. Prophet, and J. M. Ribando. Simplicial cones and the existence of shape-preserving cyclic operators. Linear Algebra and its Applications, 375(1):157–170, December 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cluzeau:2008:FDC

[CQ08] Thomas Cluzeau and Alban Quadrat. Factoring and decomposing a class of linear functional systems. Linear Algebra and its Applications, 428(1):324–381, January 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Christianson:2002:CGT

[CR02a] Hans Christianson and Victor Reiner. The critical group of a threshold graph. Linear Algebra and its Applications, 349(1–3):233–244, July 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Codenotti:2002:CSC

[CR02b] B. Codenotti and G. Resta. Computation of sparse circulant permanents via determinants. Linear Algebra and its Applications, 355(1– 3):15–34, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/201/27/
29/abstract.html.

Calvetti:2003:PPP

[CR03] D. Calvetti and L. Reichel. Pole placement preconditioning. Linear Algebra and its Applications, 366 (C):99–120, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/215/25/
                                        32/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Cardoso:2007:SRG

[CR07] Domingos M. Cardoso and Paula Rama. Spectral results on graphs with regularity constraints. Linear Algebra and its Applications, 423 (1):90–98, May 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chehab:2008:GPF

[CR08] Jean-Paul Chehab and Marcos Raydan. Geometrical properties of the Frobenius condition number for positive definite matrices. Linear Algebra and its Applications, 429(8–9):2089–2097, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cravo:2009:MCP

[Cra09] Glória Cravo. Matrix Completion Problems. Linear Algebra and its Applications, 430(8–9):2511–2540, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Calvetti:2003:EKS

[CRS03] D. Calvetti, L. Reichel, and A. Shuibi. Enriched Krylov subspace methods for illposed problems. Linear Algebra and its Applications, 362 (C):257–273, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/211/25/
46/abstract.html.

Cvetkovic:2007:SCE

[CRS07a] D. Cvetković, P. Rowlinson, and S. K. Simić. Star complements and exceptional graphs. Linear Algebra and its Applications, 423(1):146– 154, May 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetkovic:2007:SLF

[CRS07b] Dragoš Cvetković, Peter Rowlinson, and Slobodan K. Simić. Signless Laplacians of finite graphs. Linear Algebra and its Applications, 423(1): 155–171, May 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Canto:2009:FRF

[CRU09] Rafael Cantó, Beatriz Ricarte, and Ana M. Urbano. Full rank factorization in echelon form of totally nonpositive (negative) rectangular matrices. Linear Algebra and its Applications, 431(11):2213–2227, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheon:2000:SOM

[CS00a] Gi-Sang Cheon and Bryan L. Shader. Sparsity of orthogonal matrices with restrictions. Linear Algebra and its Applications, 306(1–3): 33–44, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/121/25/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/121/25/29/article.
                                        pdf.

Cravo:2000:EMS

[CS00b] Glória Cravo and Fernando C. Silva. Eigenvalues of matrices with several prescribed blocks. Linear Algebra and its Applications, 311 (1–3):13–24, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/128/25/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/128/25/26/article.
                                        pdf.

Corach:2001:IHM

[CS01] Gustavo Corach and Demetrio Stojanoff. Index of Hadamard multiplication by positive matrices II. Linear Algebra and its Applications, 332–334 (1):503–517, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
51/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/51/article.
pdf.

Camlibel:2002:EUS

[ÇS02a] M. K. Çamlibel and J. M. Schumacher. Existence and uniqueness of solutions for a class of piecewise linear dynamical systems. Linear Algebra and its Applications, 351–352:147–184, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cossidente:2002:TSE

[CS02b] Antonio Cossidente and Alessandro Siciliano. On tangent spaces and external flats to Grassmannians of lines over finite fields. Linear Algebra and its Applications, 347 (1–3):81–89, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
31/abstract.html.

Campoamor-Stursberg:2003:GTD

[CS03a] R. Campoamor-Stursberg. A graph theoretical determination of solvable complete rigid Lie algebras. Linear Algebra and its Applications, 372(1):53–66, October 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cavaretta:2003:LTB

[CS03b] Alfred S. Cavaretta and Laura Smithies. Lipschitztype bounds for the map A → |A| on L(H). Linear Algebra and its Applications, 360(C): 231–235, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/209/25/
                                        40/abstract.html.

Cravo:2003:EMS

[CS03c] Glória Cravo and Fernando C. Silva. Eigenvalues of matrices with several prescribed blocks, II. Linear Algebra and its Applications, 364(C):81–89, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/213/25/
                                        34/abstract.html.

Cravo:2003:NNI

[CS03d] Glória Cravo and Fernando C. Silva. The number of nonconstant invariant polynomials of matrices with several prescribed blocks. Linear Algebra and its Applications, 368(C):107– 116, July 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
32/abstract.html.

Condon:2004:LTS

[CS04] Anne Condon and Michael Saks. A limit theorem for sets of stochastic matrices. Linear Algebra and its Applications, 381(1):61–76, April 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cossidente:2005:SMS

[CS05] A. Cossidente and A. Siciliano. On some maximal subgroups in Aschbacher’s class C_{▽}. Linear Algebra and its Applications, 403(1):285– 290, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cegielski:2008:IAP

[CS08a] Andrzej Cegielski and Agnieszka Suchocka. Incomplete alternating projection method for large inconsistent linear systems. Linear Algebra and its Applications, 428(5–6):1313–1324, March 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chebotar:2008:MLL

[CŠ08b] Mikhail Chebotar and Peter Šemrl. Minimal locally linearly dependent spaces of operators. Linear Algebra and its Applications, 429(4):887– 900, August 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cao:2009:CSM

[CS09a] Yonglin Cao and Fernando Szechtman. Congruence of symmetric matrices over local rings. Linear Algebra and its Applications, 431(9): 1687–1690, October 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cho:2009:YIT

[CS09b] Kazuki Cho and Takashi Sano. Young’s inequality and trace. Linear Algebra and its Applications, 431(8):1218– 1222, September 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Constantinescu:2001:ISE

[CSK01] Tiberiu Constantinescu, Ali H. Sayed, and Thomas Kailath. Inverse scattering experiments, structured matrix inequalities, and tensor algebra. Linear Algebra and its Applications, 343–344(1): 147–169, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
33/abstract.html.

Caston:2001:EBC

[CSSZ01] Lauren Caston, Milena Savova, Ilya Spitkovsky, and Nahum Zobin. On eigenvalues and boundary curvature of the numerical range. Linear Algebra and its Applications, 322(1–3):129–140, January 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/141/24/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/141/24/31/article.
pdf.

Chlebowicz:2001:ACF

[CSWM01] A. Chlebowicz, A. Sładek, and M. Wołowiec-Musiał. Automorphisms of certain forms of higher degree over ordered fields. Linear Algebra and its Applications, 331 (1–3):145–153, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/36/article.
pdf.

Chen:2001:VES

[CSY01] Xiaojun Chen, Yukihiro Shogenji, and Maretsugu Yamasaki. Verification for existence of solutions of linear complementarity problems. Linear Algebra and its Applications, 324(1–3): 15–26, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/144/24/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/144/24/26/article.
                                        pdf.

Cabral:2001:FIP

[CSZ01] Isabel Cabral, Fernando C. Silva, and Ion Zaballa. Feedback invariants of a pair of matrices with prescribed columns. Linear Algebra and its Applications, 332–334(1): 447–458, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/157/25/
                                        47/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/157/25/47/article.
                                        pdf.

Cheung:2001:SCS

[CT01] Gilbert Cheung and Nam-Kiu Tsing. On star-centers of some generalized numerical ranges and diagonals of normal matrices. Linear Algebra and its Applications, 337(1–3):109–119, November 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/167/27/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/167/27/32/article.
pdf.

Cao:2003:ALA

[CT03a] You’an Cao and Zuowen Tan. Automorphisms of the Lie algebra of strictly upper triangular matrices over a commutative ring. Linear Algebra and its Applications, 360(C): 105–122, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
32/abstract.html.

Chang:2003:SRU

[CT03b] An Chang and Feng Tian. On the spectral radius of unicyclic graphs with perfect matchings. Linear Algebra and its Applications, 370(1): 237–250, September 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:2003:TSN

[CT03c] Shizhen Cheng and Yongge Tian. Two sets of new characterizations for normal and EP matrices. Linear Algebra and its Applications, 375(1):181– 195, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2004:CCN

[CT04] Doron Chen and Sivan Toledo. Combinatorial characterization of the null spaces of symmetric Hmatrices. Linear Algebra and its Applications, 392(1):71– 90, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ciavarella:2009:SEC

[CT09] Miriam Ciavarella and Lea Terracini. Some explicit constructions of integral structures in quaternion algebras. Linear Algebra and its Applications, 431(5–7):1013– 1026, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Comon:2009:GTR

[CtBDC09] P. Comon, J. M. F. ten Berge, L. De Lathauwer, and J. Castaing. Generic and typical ranks of multi-way arrays. Linear Algebra and its Applications, 430(11–12):2997– 3007, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Climent:2001:GAG

[CTW01] Joan-Josep Climent, Néstor Thome, and Yimin Wei. A geometrical approach on generalized inverses by Neumann-type series. Linear Algebra and its Applications, 332–334(1):533–540, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
53/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/53/article.
pdf.

Cui:2007:ADI

[Cui07] Jianlian Cui. Additive Drazin inverse preservers. Linear Algebra and its Applications, 426(2–3):448–453, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Curtin:2007:MLT

[Cur07]  Brian Curtin.  Modular

Leonard triples. Linear Algebra and its Applications, 424 (2–3):510–539, July 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cavers:2005:SIA

[CV05] Michael S. Cavers and Kevin N. Vander Meulen. Spectrally and inertially arbitrary sign patterns. Linear Algebra and its Applications, 394(1):53–72, January 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetko-Vah:2007:SSI

[CV07] Karin Cvetko-Vah. On the structure of semigroups of idempotent matrices. Linear Algebra and its Applications, 426(1):204–213, October 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cvetkovic:2002:GLE

[Cve02] Dragoš Cvetković. Graphs with least eigenvalue −2: a historical survey and recent developments in maximal exceptional graphs. Linear Algebra and its Applications, 356(1–3):189–210, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/202/25/
40/abstract.html.

Cavers:2009:SIA

[CVV09] Michael S. Cavers, Kevin N. Vander Meulen, and Loretta Vanderspek. Sparse inertially arbitrary patterns. Linear Algebra and its Applications, 431(11):2024–2034, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-

Cao:2000:NAA

[CW00a] You’an Cao and Jingtong Wang. A note on algebra automorphisms of strictly upper triangular matrices over commutative rings. Linear Algebra and its Applications, 311 (1–3):187–193, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/128/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/128/25/38/article.
pdf.

Chan:2000:CAM

[CW00b] Tony F. Chan and C. K. Wong. Convergence of the alternating minimization algorithm for blind deconvolution. Linear Algebra and its Applications, 316(1–3): 259–285, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
40/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/40/article.
pdf.

Cao:2002:NKS

[CW02] Zhi-Hao Cao and Mei Wang. A note on Krylov subspace methods for singular systems. Linear Algebra and its Applications, 350(1–3):285– 288, July 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cai:2003:CSP

[CW03a] Jun-Liang Cai and Bo-Ying Wang. The characterization of symmetric primitive matrices with exponent n − 1. Linear Algebra and its Applications, 364 (C):135–145, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/213/25/
                                        37/abstract.html.

Cao:2003:FHP

[CW03b] Wensheng Cao and Xiantao Wang. Friedland–Hersonsky problem for matrix algebra. Linear Algebra and its Applications, 372(1):67– 78, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cunningham:2004:EPP

[CW04] William H. Cunningham and Yaoguang Wang. On the even permutation polytope. Linear Algebra and its Applications, 389(1):269–281, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheon:2005:UMS

[CW05] Gi-Sang Cheon and Ian M. Wanless. An update on Minc’s survey of open problems involving permanents. Linear Algebra and its Applications, 403(1):314–342, July 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2007:LBCa

[CW07a] Chang-Pao Chen and Kuo-Zhong Wang. Lower bounds of Copson type for Hausdorff matrices. Linear Algebra and its Applications, 420(1):208– 217, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2007:LBCb

[CW07b] Chang-Pao Chen and Kuo-Zhong Wang. Lower bounds of Copson type for Hausdorff matrices II. Linear Algebra and its Applications, 422(2– 3):563–573, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cao:2009:BCS

[CWW09] Xifang Cao, Zhong Wang, and Hongyou Wu. On the boundary conditions in selfadjoint multi-interval Sturm– Liouville problems. Linear Algebra and its Applications, 430(11–12):2877– 2889, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cheng:2005:PAH

[CX05] Mingsong Cheng and Shufang Xu. Perturbation analysis of the Hermitian positive definite solution of the matrix equation X −A^{∗}X^{−2}A = I. Linear Algebra and its Applications, 394(1):39– 51, January 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2009:RDI

[CXW09] Jianbiao Chen, Zhaoliang Xu, and Yimin Wei. Representations for the Drazin inverse of the sum P + Q + R + S and its applications. Linear Algebra and its Applications, 430(1):438– 454, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cao:2002:MTA

[CXZL02] Huai-Xin Cao, Zong-Ben Xu, Jian-Hua Zhang, and Wei-Hua Li. Matrix-traces on C^{∗}- algebra M_{n}(A). Linear Algebra and its Applications, 345(1–3):255–260, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/

Cheng:2000:DNB

[CY00] Sui Sun Cheng and Hung-Ta Yen. On a discrete nonlinear boundary value problem. Linear Algebra and its Applications, 313(1– 3):193–201, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/130/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/130/25/37/article.
pdf.

Chen:2000:CJS

[CZ00] Quande Chen and Xinlong Zhou. Characterization of joint spectral radius via trace. Linear Algebra and its Applications, 315(1–3): 175–188, August 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/132/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/132/25/35/article.
pdf.

Cao:2005:LPB

[CZ05] Chong-Guang Cao and Xian Zhang. Linear preservers between matrix modules over connected commutative rings. Linear Algebra and its Applications, 397(1):355– 366, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cassier:2007:OMC

[CZ07] G. Cassier and E. H. Zerouali. Operator matrices in class C_{P}. Linear Algebra and its Applications, 420(2–3): 361–376, January 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2008:EHT

[CZ08] Haiyan Chen and Fuji Zhang. The expected hitting times for finite Markov chains. Linear Algebra and its Applications, 428(11–12):2730– 2749, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Czornik:2005:GSS

[Czo05] Adam Czornik. On the generalized spectral subradius. Linear Algebra and its Applications, 407(1):242–248, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chen:2009:FHP

[CZZ09] Xiaoling Chen, Shifang Zhang, and Huaijie Zhong. On the filling in holes problem for operator matrices.

Linear Algebra and its Applications, 430(1):558–563, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Davidson:2005:TFL

[D– D05] Kenneth R. Davidson and Dragomir Ž. – Doković. Tridiagonal forms in low dimensions. Linear Algebra and its Applications, 407(1):169–188, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deAbreu:2007:ONR

[dA07] Nair Maria Maia de Abreu. Old and new results on algebraic connectivity of graphs. Linear Algebra and its Applications, 423(1):53– 73, May 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dong:2010:NCP

[DA10] Hongbo Dong and Kurt Anstreicher. A note on “5 × 5 completely positive matrices”. Linear Algebra and its Applications, 433(5): 1001–1004, October 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See [BAD09].

Dahl:2000:NDD

[Dah00] Geir Dahl. A note on diagonally dominant matrices. Linear Algebra and its Applications, 317(1–3):217– 224, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/136/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/136/25/36/article.
pdf.

Dahl:2001:PMI

[Dah01] Geir Dahl. Principal majorization ideals and optimization. Linear Algebra and its Applications, 331(1– 3):113–130, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/34/article.
pdf.

Dahl:2003:DGM

[Dah03]  Geir Dahl.  The doubly

graded matrix cone and Ferrers matrices. Linear Algebra and its Applications, 368 (C):171–190, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
36/abstract.html.

Dahl:2004:TDS

[Dah04] Geir Dahl. Tridiagonal doubly stochastic matrices. Linear Algebra and its Applications, 390(1):197–208, October 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dahl:2005:MAS

[Dah05] Geir Dahl. A method for approximating symmetrically reciprocal matrices by transitive matrices. Linear Algebra and its Applications, 403(1): 207–215, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dahl:2008:TMS

[Dah08]  Geir Dahl.  Transporta-

tion matrices with staircase patterns and majorization. Linear Algebra and its Applications, 429(7):1840–1850, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dahl:2009:PMR

[Dah09] Geir Dahl. Permutation matrices related to Sudoku. Linear Algebra and its Applications, 430(8–9):2457–2463, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Damm:2002:SFT

[Dam02]  Tobias Damm.  State-

feedback H^{∞}-type control of linear systems with timevarying parameter uncertainty. Linear Algebra and its Applications, 351–352:185– 210, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Damm:2003:MRI

[Dam03]  Tobias Damm.  Minimal

representations of inverted Sylvester and Lyapunov operators. Linear Algebra and its Applications, 363 (C):35–41, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
29/abstract.html.

Damm:2004:PGC

[Dam04]  Tobias Damm.  Positive

groups on H^{\} are completely positive. Linear Algebra and its Applications, 393(1):127– 137, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dancis:2002:RIH

[Dan02] Jerome Dancis. Ranks and inertias of Hermitian block Toeplitz matrices. Linear Algebra and its Applications, 353(1–3):21–32, September 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Danciger:2006:MMT

[Dan06] Jeffrey Danciger. A min–max theorem for complex symmetric matrices. Linear Algebra and its Applications, 412 (1):22–29, January 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Das:2004:CGW

[Das04a] Kinkar Ch. Das. A characterization on graphs which achieve the upper bound for the largest Laplacian eigenvalue of graphs. Linear Algebra and its Applications, 376 (1):173–186, January 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Das:2004:SLB

[Das04b] Kinkar Ch. Das. Sharp lower bounds on the Laplacian eigenvalues of trees. Linear Algebra and its Applications, 384(1):155–169, June 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Das:2007:EGC

[Das07] Kinkar Ch. Das. Extremal graph characterization from the upper bound of the Laplacian spectral radius of weighted graphs. Linear Algebra and its Applications, 427(1):55–69, November 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Das:2009:SUB

[Das09] Kinkar Ch. Das. A sharp upper bound on the maximal entry in the principal eigenvector of symmetric nonnegative matrix. Linear Algebra and its Applications, 431(8):1340–1350, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dattorro:2008:ERE

[Dat08] Jon Dattorro. Equality relating Euclidean distance cone to positive semidefinite cone. Linear Algebra and its Applications, 428(11–12):2597– 2600, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Daugherty:2008:IUG

[Dau08] Sean Daugherty. The inertia of unicyclic graphs and the implications for closedshells. Linear Algebra and its Applications, 429(4):849– 858, August 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dawson:2002:EPU

[Daw02] Kevin J. Dawson. The evolution of a population under recombination: how to linearise the dynamics. Linear Algebra and its Applications, 348 (1–3):115–137, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
35/abstract.html.

Dax:2000:MGS

[Dax00]  Achiya Dax.  A modi-

fied Gram–Schmidt algorithm with iterative orthogonalization and column pivoting. Linear Algebra and its Applications, 310(1–3): 25–42, May 1, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/127/25/29/article.
pdf.

Dax:2003:ASA

[Dax03a] Achiya Dax. The adventures of a simple algorithm. Linear Algebra and its Applications, 361(C):41–61, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
30/abstract.html.

Dax:2003:ORQ

[Dax03b] Achiya Dax. The Orthogonal Rayleigh Quotient Iteration (ORQI) method. Linear Algebra and its Applications, 358(1–3):23–43, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/207/25/
                                        28/abstract.html.

Dax:2006:DBT

[Dax06] Achiya Dax. The distance between two convex sets. Linear Algebra and its Applications, 416(1):184–213, July 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deBoor:2000:PDJ

[dB00] Carl de Boor. On Pták’s derivation of the Jordan normal form. Linear Algebra and its Applications, 310 (1–3):9–10, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/127/25/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/127/25/27/article.
                                        pdf.

Das:2005:SUB

[DB05] Kinkar Ch. Das and R. B. Bapat. A sharp upper bound on the largest Laplacian eigenvalue of weighted graphs. Linear Algebra and its Applications, 409(1):153– 165, November 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dzhafarov:2006:SCS

[DB06] Vakif Dzhafarov and Taner Büyükköroğlu. On the stability of a convex set of matrices. Linear Algebra and its Applications, 414(2–3):547– 559, April 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dzhafarov:2008:NPM

[DB08] Vakif Dzhafarov and Taner Büyükköroḡlu. On nonsingularity of a polytope of matrices. Linear Algebra and its Applications, 429(5–6):1174– 1183, September 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Druskin:2005:SLR

[DBK05] Vladimir Druskin, Liliana Borcea, and Leonid Knizhnerman. On the sensitivity of Lanczos recursions to the spectrum. Linear Algebra and its Applications, 396(1): 103–125, February 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deBoor:2008:PLB

[dBS08] C. de Boor and B. Shekhtman. On the pointwise limits of bivariate Lagrange projectors. Linear Algebra and its Applications, 429(1):311– 325, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

daCruz:2005:EID

[dCdS05] Henrique F. da Cruz and J. A. Dias da Silva. Equality of immanantal decomposable tensors, II. Linear Algebra and its Applications, 395 (1):95–119, January 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duran:2002:OMP

[DD02] Antonio J. Durán and Emilio Defez. Orthogonal matrix polynomials and quadrature formulas. Linear Algebra and its Applications, 345(1– 3):71–84, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/175/27/
                                        31/abstract.html.

Derevyagin:2004:SPG

[DD04] Maxim Derevyagin and Vladimir Derkach. Spectral problems for generalized Jacobi matrices. Linear Algebra and its Applications, 382(1):1– 24, May 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Du:2005:RCD

[DD05] Hong-Ke Du and Chun-Yuan Deng. The representation and characterization of Drazin inverses of operators on a Hilbert space. Linear Algebra and its Applications, 407(1):117–124, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dubi:2007:RIR

[DD07] Chen Dubi and Harry Dym. Riccati inequalities and reproducing kernel Hilbert spaces. Linear Algebra and its Applications, 420(2–3): 458–482, January 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeTeran:2009:LRP

[DD09] Fernando De Terán and Froilán M. Dopico. Low rank perturbation of regular matrix polynomials. Linear Algebra and its Applications, 430(1):579–586, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deDurana:2002:GMM

[dDG02] José María González de Durana and Juan-Miguel Gracia. Geometric multiplicity margin for a submatrix. Linear Algebra and its Applications, 349(1–3): 77–104, July 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dolinar:2008:GPI

[DDHL08] Gregor Dolinar, Shuanping Du, Jinchuan Hou, and Peter LegiŠa. General preservers of invariant subspace lattices. Linear Algebra and its Applications, 429(1):100– 109, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeCat:2003:TST

[DDI03] Tine De Cat, Karel Dekimpe, and Paul Igodt. Translations in simply transitive affine actions of Heisenberg type Lie groups. Linear Algebra and its Applications, 359(1–3): 101–111, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/208/25/
                                        34/abstract.html.

DeTeran:2008:FOS

[DDM08] Fernando De Terán, Froilán M. Dopico, and Julio Moro. First order spectral perturbation theory of square singular matrix pencils. Linear Algebra and its Applications, 429(2– 3):548–576, July 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Davydova:2005:CPA

[DDvG05] N. V. Davydova, O. Diekmann, and S. A. van Gils. On circulant populations. I. The algebra of semelparity. Linear Algebra and its Applications, 398(1):185– 243, March 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeSchutter:2000:UBS

[De 00] Bart De Schutter. On the ultimate behavior of the sequence of consecutive powers of a matrix in the maxplus algebra. Linear Algebra and its Applications, 307(1– 3):103–117, March 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/32/article.
pdf.

DeBruyn:2007:DNE

[De 07a] Bart De Bruyn. A decomposition of the natural embedding spaces for the symplectic dual polar spaces. Linear Algebra and its Applications, 426(2–3):462–477, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeBruyn:2007:TSP

[De 07b] Bart De Bruyn. Tight sets of points in the half-spin geometry related to Q^{+}(9, q). Linear Algebra and its Applications, 424(2–3):480–491, July 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeBruyn:2008:TNC

[De 08] Bart De Bruyn. Two new classes of hyperplanes of the dual polar space DH(2n − 1, 4) not arising from the Grassmann embedding. Linear Algebra and its Applications, 429(8–9):2030–2045, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeBruyn:2009:HEG

[De 09a] Bart De Bruyn. Hyperplanes of embeddable Grassmannians arise from projective embeddings: A short proof. Linear Algebra and its Applications, 430(1):418– 422, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeBruyn:2009:IFE

[De 09b] Bart De Bruyn. On isometric full embeddings of symplectic dual polar spaces into Hermitian dual polar spaces. Linear Algebra and its Applications, 430(8–9): 2541–2552, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeBruyn:2009:SSE

[De 09c] Bart De Bruyn. Some subspaces of the kth exterior power of a symplectic vector space. Linear Algebra and its Applications, 430(11–12): 3095–3104, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dekker:2009:PKS

[Dek09]  K. Dekker.  Partitioned

Krylov subspace iteration in implicit Runge–Kutta methods. Linear Algebra and its Applications, 431(3–4):488– 494, July 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Delai:2001:DAS

[Del01] M’hammed Benlarbi Delai. Description de AlgLatT / W (T) pour certaines extensions d’opérateurs autoadjoints. (French) [description of AlgLatT /W (T) for certain extensions of selfadjoint operators]. Linear Algebra and its Applications, 322(1–3):43–49, January 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/141/24/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/141/24/26/article.
pdf.

Deng:2009:DIS

[Den09]  Chun Yuan Deng.  The

Drazin inverses of sum and difference of idempotents. Linear Algebra and its Applications, 430(4):1282–1291, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

daFonseca:2000:IHB

[dF00] C. M. da Fonseca. The inertia of Hermitian block matrices with zero main diagonal. Linear Algebra and its Applications, 311(1– 3):153–160, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/128/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/128/25/35/article.
pdf.

daFonseca:2001:EIS

[dFP01] C. M. da Fonseca and J. Petronilho. Explicit inverses of some tridiagonal matrices. Linear Algebra and its Applications, 325 (1–3):7–21, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/147/24/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/147/24/25/article.
pdf.

DiasdaSilva:2002:GDA

[DG02] J. A. Dias da Silva and Hemar Godinho. Generalized derivations and additive theory. Linear Algebra and its Applications, 342(1– 3):1–15, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/173/27/
27/abstract.html.

DiasdaSilva:2003:CPR

[DG03] J. A. Dias da Silva and Hemar Godinho. Critical polynomials related to generalized derivations. Linear Algebra and its Applications, 370(1):315–339, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Driessel:2007:ZPI

[DG07] Kenneth R. Driessel and Alf Gerisch. Zero-preserving isospectral flows based on parallel sums. Linear Algebra and its Applications, 421(1): 69–84, February 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Davies:2001:DND

[DGLS01] E. Brian Davies, Graham M. L. Gladwell, Josef Leydold, and Peter F. Stadler. Discrete nodal domain theorems. Linear Algebra and its Applications, 336(1–3): 51–60, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/166/27/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/166/27/31/article.
                                        pdf.

DeSimone:2002:NRS

[DGN02] Caterina De Simone, Grigor Gasparyan, and Paolo Nobili. Non-regular square bipartite designs. Linear Algebra and its Applications, 348 (1–3):63–85, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/195/27/
                                        32/abstract.html.

DeAlba:2000:CMP

[DH00] Luz Maria DeAlba and Leslie Hogben. Completions of P - matrix patterns. Linear Algebra and its Applications, 319(1–3):83–102, November 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/138/24/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/138/24/29/article.
pdf.

Damm:2001:NMR

[DH01] T. Damm and D. Hinrichsen. Newton’s method for a rational matrix equation occurring in stochastic control. Linear Algebra and its Applications, 332–334(1): 81–109, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/30/article.
pdf.

Doob:2002:CPD

[DH02] Michael Doob and Willem H. Haemers. The complement of the path is determined by its spectrum. Linear Algebra and its Applications, 356(1– 3):57–65, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/202/25/
                                        30/abstract.html.

Damm:2003:NMC

[DH03] T. Damm and D. Hinrichsen. Newton’s method for concave operators with resolvent positive derivatives in ordered Banach spaces. Linear Algebra and its Applications, 363(C):43–64, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/212/25/
                                        30/abstract.html.

Du:2004:SIR

[DH04] Shuanping Du and Jinchuan Hou. Similarity invariant real linear subspaces and similarity preserving additive maps. Linear Algebra and its Applications, 377(1):141– 153, January 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Daneshgar:2006:GHN

[DH06] Amir Daneshgar and Hossein Hajiabolhassan. Graph homomorphisms and nodal domains. Linear Algebra and its Applications, 418(1):44– 52, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2007:VSA

[dH07] Xiang dong Hou. On a vector space analogue of Kneser’s theorem. Linear Algebra and its Applications, 426(1):214– 227, October 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dosly:2009:FEO

[DH09] Ondřej Došlý and Petr Hasil. Friedrichs extension of operators defined by symmetric banded matrices. Linear Algebra and its Applications, 430(8–9):1966–1975, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Du:2007:NMP

[DHB07] Shuanping Du, Jinchuan Hou, and Zhaofang Bai. Nonlinear maps preserving similarity on B(H). Linear Algebra and its Applications, 422(2–3):506–516, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Deuflhard:2000:IAI

[DHFS00] P. Deuflhard, W. Huisinga, A. Fischer, and Ch. Schütte. Identification of almost invariant aggregates in reversible nearly uncoupled Markov chains. Linear Algebra and its Applications, 315(1–3):39–59, August 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/132/25/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/132/25/27/article.
                                        pdf.

DeAlba:2006:MRM

[DHH^{+}06] Luz Maria DeAlba, Timothy L. Hardy, Irvin Roy Hentzel, Leslie Hogben, and Amy Wangsness. Minimum rank and maximum eigenvalue multiplicity of symmetric tree sign patterns. Linear Algebra and its Applications, 418(2–3):394–415, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DeAlba:2007:SAP

[DHH^{+}07] Luz M. DeAlba, Irvin R. Hentzel, Leslie Hogben, Judith McDonald, Rana Mikkelson, Olga Pryporova, Bryan Shader, and Kevin N. Vander Meulen. Spectrally arbitrary patterns: Reducibility and the 2n conjecture for n = 5. Linear Algebra and its Applications, 423 (2–3):262–276, June 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dirr:2006:LAR

[DHK06] G. Dirr, U. Helmke, and M. Kleinsteuber. Lie algebra representations, nilpotent matrices, and the Cnumerical range. Linear Algebra and its Applications, 413(2–3):534–566, March 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deHoog:2007:SSM

[dHM07] F. R. de Hoog and R. M. M. Mattheij. Subset selection for matrices. Linear Algebra and its Applications, 422(2– 3):349–359, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ding:2003:MCD

[DHZ03a] Jiu Ding, L. Richard Hitt, and Xin-Min Zhang. Markov chains and dynamic geometry of polygons. Linear Algebra and its Applications, 367 (C):255–270, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
44/abstract.html.

Dosly:2003:NDQ

[DHZ03b] Ondřej Došlý, Roman Hilscher, and Vera Zeidan. Nonnegativity of discrete quadratic functionals corresponding to symplectic difference systems. Linear Algebra and its Applications, 375(1):21– 44, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DiBenedetto:2003:DOA

[Di 03] F. Di Benedetto. The mth difference operator applied to L^{2} functions on a finite interval. Linear Algebra and its Applications, 366 (C):173–198, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
36/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

DiasdaSilva:2002:CET

[Dia02] J. A. Dias da Silva. Colorings and equality of tensors. Linear Algebra and its Applications, 342(1–3): 79–91, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/173/27/
32/abstract.html.

Ding:2003:NPR

[Din03] Jiu Ding. New perturbation results on pseudo-inverses of linear operators in Banach spaces. Linear Algebra and its Applications, 362 (C):229–235, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

Dittmann:2000:CMM

[Dit00] Jochen Dittmann. On the curvature of monotone metrics and a conjecture concerning the Kubo-Mori metric. Linear Algebra and its Applications, 315(1–3): 83–112, August 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/132/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/132/25/29/article.
pdf.

DiasdaSilva:2001:CSP

[DJ01] J. A. Dias da Silva and Charles R. Johnson. Cospectrality and similarity for a pair of matrices under multiplicative and additive composition with diagonal matrices. Linear Algebra and its Applications, 326(1– 3):15–25, March 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/148/24/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/148/24/25/article.
pdf.

Dopico:2006:CBS

[DJ06] Froilán M. Dopico and Charles R. Johnson. Complementary bases in symplectic matrices and a proof that their determinant is one. Linear Algebra and its Applications, 419(2–3):772–778, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duggal:2007:QHO

[DJ07] B. P. Duggal and In Ho Jeon. On p-quasi-hyponormal operators. Linear Algebra and its Applications, 422(1):331– 340, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drew:2000:CCG

[DJKM00] John H. Drew, Charles R. Johnson, Steven J. Kilner, and Angela M. McKay. The cycle completable graphs for the completely positive and doubly nonnegative completion problems. Linear Algebra and its Applications, 313 (1–3):141–154, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/130/25/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/130/25/33/article.
                                        pdf.

Drew:2001:CPM

[DJL01] John H. Drew, Charles R. Johnson, and Fumei Lam. Complete positivity of matrices of special form. Linear Algebra and its Applications, 327(1–3):121–130, April 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/36/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/152/
25/36/article.pdf.

Dopico:2006:MLF

[DJM06] Froilán M. Dopico, Charles R. Johnson, and Juan M. Molera. Multiple LU factorizations of a singular matrix. Linear Algebra and its Applications, 419(1):24– 36, November 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DAlfonso:2009:IOQ

[DJMS09] Lisi D’Alfonso, Gabriela Jeronimo, Gustavo Massaccesi, and Pablo Solernó. On the index and the order of quasi-regular implicit systems of differential equations. Linear Algebra and its Applications, 430(8–9):2102– 2122, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drew:2000:SAP

[DJOvdD00] J. H. Drew, C. R. Johnson, D. D. Olesky, and P. van den Driessche. Spectrally arbitrary patterns. Linear Algebra and its Applications, 308(1–3):121–137, March 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/125/25/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/125/25/31/article.
                                        pdf.

Drnovsek:2004:IOS

[DK04] R. Drnovšek and T. Košir. Irreducible operator semigroups such that AB and BA are proportional. Linear Algebra and its Applications, 383(1):113–118, May 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Demmel:2006:ASP

[DK06] James Demmel and Plamen Koev. Accurate SVDs of polynomial Vandermonde matrices involving orthonormal polynomials. Linear Algebra and its Applications, 417(2–3):382–396, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Das:2007:CGH

[DK07] Kinkar Ch. Das and Jin Ho Kwak. Characterization of graphs having extremal Randić indices. Linear Algebra and its Applications, 420 (1):124–134, January 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dajic:2008:ERR

[DK08a] Alegra Dajić and J. J. Koliha. Equations ax = c and xb = d in rings and rings with involution with applications to Hilbert space operators. Linear Algebra and its Applications, 429(7):1779–1809, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dopico:2008:BDO

[DK08b] Froilán M. Dopico and Plamen Koev. Bidiagonal decompositions of oscillating systems of vectors. Linear Algebra and its Applications, 428(11–12):2536– 2548, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dobovisek:2007:MPP

[DKL^{+}07] M. Dobovišek, B. Kuzma, G. Lešnjak, C. K. Li, and T. Petek. Mappings that preserve pairs of operators with zero triple Jordan product.

Linear Algebra and its Applications, 426(2–3):255–279, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dellnitz:2005:P

[DKNS05] Michael Dellnitz, Steve Kirkland, Michael Neumann, and Christof Schuette. Preface. Linear Algebra and its Applications, 398(1):1– 2, March 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deKerchove:2008:MPO

[dKNvD08] Cristobald de Kerchove, Laure Ninove, and Paul van Dooren. Maximizing PageRank via outlinks. Linear Algebra and its Applications, 429(5–6):1254–1276, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drivaliaris:2008:FEO

[DKP08] Dimosthenis Drivaliaris, Sotirios Karanasios, and Dimitrios Pappas. Factorizations of EP operators. Linear Algebra and its Applications, 429(7): 1555–1567, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Derisavi:2004:SSS

[DKS04] Salem Derisavi, Peter Kemper, and William H. Sanders. Symbolic state-space exploration and numerical analysis of state-sharing composed models. Linear Algebra and its Applications, 386(1):137– 166, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Du:2003:LCL

[DKX03] Shao-Fei Du, Jin Ho Kwak, and Ming-Yao Xu. Linear criteria for lifting automorphisms of elementary abelian regular coverings. Linear Algebra and its Applications, 373(1):101–119, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Daubechies:1992:SMA

[DL92] Ingrid Daubechies and Jeffrey C. Lagarias. Sets of matrices all infinite products of which converge. Linear Algebra and its Applications, 161(1):227–263, January 15, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See corrigendum [DL01].

Daubechies:2001:CAS

[DL01] Ingrid Daubechies and Jeffrey C. Lagarias. Corrigendum/addendum to: Sets of matrices all infinite products of which converge. Linear Algebra and its Applications, 327(1–3):69–83, April 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/31/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/152/
25/31/article.pdf.

Dlab:2003:SRC

[DL03] Vlastimil Dlab and Piroska Lakatos. On spectral radii of Coxeter transformations. Linear Algebra and its Applications, 365 (C):143–153, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
49/abstract.html.

Dagum:2004:LTP

[DL04] Estela Bee Dagum and Alessandra Luati. A linear transformation and its properties with special applications in time series filtering. Linear Algebra and its Applications, 388(1):107– 117, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Du:2005:SCG

[DL05] Hong-Ke Du and Yuan Li.

The spectral characterization of generalized projections. Linear Algebra and its Applications, 400(1):313– 318, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duan:2008:EHP

[DL08] Xuefeng Duan and Anping Liao. On the existence of Hermitian positive definite solutions of the matrix equation X^{s} + A^{∗}X^{−t}A = Q. Linear Algebra and its Applications, 429(4):673– 687, August 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Das:2009:EIC

[DL09] Kinkar Ch. Das and Sang-Gu Lee. On the Estrada index conjecture. Linear Algebra and its Applications, 431(8):1351–1359, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deLima:2007:LIG

[dLdAOdF07] Leonardo Silva de Lima, Nair Maria Maia de Abreu, Carla Silva Oliveira, and Maria Aguieiras Alvarez de Freitas. Laplacian integral graphs in S(a, b). Linear Algebra and its Applications, 423(1):136–145, May 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drnovsek:2000:TLS

[DLM^{+}00] R. Drnovšek, L. Livshits, G. MacDonald, B. Mathes, H. Radjavi, and P. Šemrl. On transitive linear semigroups. Linear Algebra and its Applications, 305(1–3): 67–86, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/120/25/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/120/25/30/article.
                                        pdf.

Dahl:2007:TPM

[DLMO07] Geir Dahl, Jon Magne Leinaas, Jan Myrheim, and Eirik Ovrum. A tensor product matrix approximation problem in quantum physics. Linear Algebra and its Applications, 420(2–3):711–725, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

delaMora:2006:MBM

[dlMW06] Carlos de la Mora and Piotr J. Wojciechowski. Multiplicative bases in matrix algebras. Linear Algebra and its Applications, 419(2–3):287– 298, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-

delaPena:2003:PCM

[dlP03] José A. de la Peña. Periodic Coxeter matrices. Linear Algebra and its Applications, 365(C):135–142, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
33/abstract.html.

delaPena:2007:EEI

[dlPGR07] José Antonio de la Peña, Ivan Gutman, and Juan Rada. Estimating the Estrada index. Linear Algebra and its Applications, 427(1):70– 76, November 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Davis:2008:PNH

[DLS08] Chandler Davis, Chi-Kwong Li, and Abbas Salemi. Polynomial numerical hulls of matrices. Linear Algebra and its Applications, 428(1):137– 153, January 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duan:2008:NME

[DLT08] Xuefeng Duan, Anping Liao, and Bin Tang. On the nonlinear matrix equation X − ∑_{i=1}^{m} A_{i}^{∗}X^{δ_{i}}A_{i} = Q. Linear Algebra and its Applications, 429(1):110–121, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DiFiore:2003:HTA

[DLZ03] Carmine Di Fiore, Filomena Lepore, and Paolo Zellini. Hartley-type algebras in displacement and optimization strategies. Linear Algebra and its Applications, 366 (C):215–232, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
38/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Domanov:2002:IHS

[DM02] I. Yu. Domanov and M. M. Malamud. Invariant and hyperinvariant subspaces of an operator J^{α} and related operator algebras in Sobolev spaces. Linear Algebra and its Applications, 348(1–3): 209–230, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
43/abstract.html.

Dhillon:2003:IDS

[DM03] I. S. Dhillon and A. N. Malyshev. Inner deflation for symmetric tridiagonal matrices. Linear Algebra and its Applications, 358(1–3): 139–144, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/207/25/
33/abstract.html.

Draper:2006:G

[DM06] Cristina Draper and Cándido Martín. Gradings on }_{∈}. Linear Algebra and its Applications, 418(1):85–111, October 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drineas:2007:RAT

[DM07] Petros Drineas and Michael W. Mahoney. A randomized algorithm for a tensor-based generalization of the singular value decomposition. Linear Algebra and its Applications, 420(2–3):553–571, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Diaz:2003:CQM

[DMD03] R. Durán Díaz, J. Muñoz Masqué, and A. Peinado Domínguez. Classifying quadratic maps from plane to plane. Linear Algebra and its Applications, 364(C): 1–12, May 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/213/25/
                                        39/abstract.html.

de:2009:BOC

[dMFZ09] Vera Jesus de, João Tiago Mexia, Miguel Fonseca, and Roman Zmyślony. Binary operations and canonical forms for factorial and related models. Linear Algebra and its Applications, 430(10):2781– 2797, May 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dellacherie:2000:DSM

[DMM00a] Claude Dellacherie, Servet Martínez, and Jaime San Martín. Description of the sub-Markov kernel associated to generalized ultrametric matrices. An algorithmic approach. Linear Algebra and its Applications, 318(1– 3):1–21, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/137/24/
                                        24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/137/24/24/article.
                                        pdf.

Dopico:2000:WTR

[DMM00b] Froilán M. Dopico, Julio Moro, and Juan M. Molera. Weyl-type relative perturbation bounds for eigensystems of Hermitian matrices. Linear Algebra and its Applications, 309(1–3):3–18, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/26/article.
pdf.

DiasdaSilva:2005:D

[DMM^{+}05] José Dias da Silva, Eduardo Marques de Sá, Russell Merris, Fernando Conceição Silva, and João Filipe Queiró. Dedication. Linear Algebra and its Applications, 401(1): 1–4, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Delvaux:2009:FCD

[DMMV09] Steven Delvaux, Ana Marco, José-Javier Martínez, and Marc Van Barel. Fast computation of determinants of Bézout matrices and application to curve implicitization. Linear Algebra and its Applications, 430(1):27– 33, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deMalafosse:2007:GHT

[dMR07a] Bruno de Malafosse and Vladimir Rakočević. A generalization of a Hardy theorem. Linear Algebra and its Applications, 421(2–3):306– 314, March 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deMalafosse:2007:MTS

[dMR07b] Bruno de Malafosse and Vladimir Rakocević. Matrix transformation and statistical convergence. Linear Algebra and its Applications, 420(2–3):377–387, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Darnell:2008:DGS

[DMW08] Dean Darnell, Ronald B. Morgan, and Walter Wilcox. Deflated GMRES for systems with multiple shifts and multiple right-hand sides. Linear Algebra and its Applications, 429(10):2415–2434, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dekoninck:2000:EPN

[DN00] Bertrand Dekoninck and Serge Nicaise. The eigenvalue problem for networks of beams. Linear Algebra and its Applications, 314(1– 3):165–189, July 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/

gej-ng/10/30/19/131/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/131/25/30/article.
pdf.

Dym:2005:PC

[DN05a] Harry Dym and Shahar Nevo. Pole cancellation. Linear Algebra and its Applications, 404(1):27–57, July 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dym:2005:ZC

[DN05b] Harry Dym and Shahar Nevo. Zero cancellation. Linear Algebra and its Applications, 404(1):1–26, July 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dauxois:2004:CAR

[DNR04] J. Dauxois, G. M. Nkiet, and Y. Romain. Canonical analysis relative to a closed subspace. Linear Algebra and its Applications, 388(1):119– 145, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dolzan:2009:IMA

[DO09a] David Dolžan and Polona Oblak. Idempotent matrices over antirings. Linear Algebra and its Applications, 431(5– 7):823–832, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dolzan:2009:INM

[DO09b] David Dolžan and Polona Oblak. Invertible and nilpotent matrices over antirings. Linear Algebra and its Applications, 430(1):271– 278, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dobovisek:2001:MSM

[Dob01] Mirko Dobovišek. On minimal solutions of the matrix equation AXY B = 0. Linear Algebra and its Applications, 325(1–3):81–99, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/147/24/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/147/24/30/article.
pdf.

Dobovisek:2004:SS

[Dob04] Mirko Dobovišek. Simultaneous symmetrization. Linear Algebra and its Applications, 383(1):107–112, May 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dodig:2005:FIM

[Dod05] Marija Dodig. Feedback invariants of matrices with prescribed rows. Linear Algebra and its Applications, 405 (1):121–154, August 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dodig:2007:SLC

[Dod07] Marija Dodig. Special loop connections of linear systems. Linear Algebra and its Applications, 424(2–3):555– 569, July 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dodig:2008:MPCa

[Dod08a] Marija Dodig. Matrix pencils completion problems. Linear Algebra and its Applications, 428(1):259–304, January 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dodig:2008:MPCb

[Dod08b] Marija Dodig. Matrix pencils completion problems II. Linear Algebra and its Applications, 429(2–3):633–648, July 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Doerr:2007:LDT

[Doe07] Benjamin Doerr. L_{p} linear discrepancy of totally unimodular matrices. Linear Algebra and its Applications, 420(2–3):663–666, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dolinar:2003:MMA

[Dol03] Gregor Dolinar. Maps on matrix algebras preserving idempotents. Linear Algebra and its Applications, 371(1): 287–300, September 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Domokos:2000:PSS

[Dom00] Mátyás Domokos. Poincaré series of semi-invariants of 2 × 2 matrices. Linear Algebra and its Applications, 310 (1–3):183–194, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
40/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/127/25/40/article.
pdf.

Domokos:2003:DFM

[Dom03] M. Domokos. On the dimension of faithful modules over finite dimensional basic algebras. Linear Algebra and its Applications, 365 (C):155–157, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

Dato-on:2009:PDM

[DoMP09] Jolle E. Dato-on, Dennis I. Merino, and Agnes T. Paras. The ϕ_{J} polar decomposition of matrices with rank 2. Linear Algebra and its Applications, 430(2–3):756–761, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See related work [MPP10].

Donati:2001:PCN

[Don01] Giorgio Donati. On Pappus’ configuration in noncommutative projective geometry. Linear Algebra and its Applications, 331(1– 3):203–209, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
41/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/41/article.
pdf.

Doob:2002:CGC

[Doo02] Michael Doob. Circulant graphs with det(A(G)) = deg(G): codeterminants with K_{n}. Linear Algebra and its Applications, 340(1–3): 87–96, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/172/27/
                                        33/abstract.html.

deOliveira:2006:PIC

[dOQSZ06] Graciano de Oliveira, João Filipe Queiró, Bryan Shader, and Ion Zaballa. Preface to the 11th ILAS Conference Proceedings, Coimbra 2004. Linear Algebra and its Applications, 413(2–3):245– 246, March 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dewilde:2001:SIS

[DOS01] Patrick Dewilde, Vadim Olshevsky, and Ali H. Sayed. Special issue on structured and infinite systems of linear equations. Linear Algebra and its Applications, 343– 344(1):1–4, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/174/27/
                                        27/abstract.html.

Dame:2005:ECP

[DOvdD05] L. F. Dame, D. D. Olesky, and P. van den Driessche. The exponent and circumdiameter of primitive digraphs. Linear Algebra and its Applications, 396(1):243– 258, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dieci:2000:PAE

[DP00] Luca Dieci and Alessandra Papini. Padé approximation for the exponential of a block triangular matrix. Linear Algebra and its Applications, 308(1–3): 183–202, March 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/125/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/125/25/35/article.
pdf.

Ding:2001:SPS

[DP01] Jiu Ding and Wallace C. Pye. On the spectrum and pseudoinverse of a special bordered matrix. Linear Algebra and its Applications, 331 (1–3):11–20, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/26/article.
pdf.

Dolinar:2002:DPM

[DP02a] Gregor Dolinar and PeterŠemrl. Determinant preserving maps on matrix algebras. Linear Algebra and its Applications, 348(1–3): 189–192, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/195/27/
                                        40/abstract.html.

Duran:2002:GQF

[DP02b] Antonio J. Durán and Beatriz Polo. Gaussian quadrature formulae for matrix weights. Linear Algebra and its Applications, 355(1–3): 119–146, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/201/27/
                                        35/abstract.html.

Dhillon:2004:MRC

[DP04] Inderjit S. Dhillon and Beresford N. Parlett. Multiple representations to compute orthogonal eigenvectors of symmetric tridiagonal matrices. Linear Algebra and its Applications, 387(1):1– 28, August 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ding:2006:SRS

[DPZ06] Jiu Ding, Wallace Pye, and Lian Zhao. Some results on structured M-matrices with an application to wireless communications. Linear Algebra and its Applications, 416(2–3):608–614, July 15, 2006. CODEN LAAPAW. ISSN 0024-

Denardo:2005:TEM

[DR05] Eric V. Denardo and Uriel G. Rothblum. Totally expanding multiplicative systems. Linear Algebra and its Applications, 406(1):142–158, September 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dutta:2008:ARS

[DR08] S. Dutta and T. S. S. R. K. Rao. Algebraic reflexivity of some subsets of the isometry group. Linear Algebra and its Applications, 429(7): 1522–1527, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Draxler:2003:TBD

[Drä03] Peter Dräxler. Tame biextensions of derived tame hereditary algebras. Linear Algebra and its Applications, 365 (C):159–167, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
38/abstract.html.

Dragomir:2005:SRG

[Dra05] Sever S. Dragomir. Some reverses of the generalised triangle inequality in complex inner product spaces. Linear Algebra and its Applications, 402(1):245–254, June 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dragomir:2006:SIE

[Dra06] Sever S. Dragomir. Some inequalities for the Euclidean operator radius of two operators in Hilbert spaces. Linear Algebra and its Applications, 419(1):256–264, November 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drazin:2007:ERS

[Dra07] Michael P. Drazin. Exact rational solution of the matrix equation A = p(X) by linearization. Linear Algebra and its Applications, 426(2–3):502–515, October 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dragomir:2008:INR

[Dra08a] S. S. Dragomir. Inequalities for the numerical radius, the norm and the maximum of the real part of bounded linear operators in Hilbert spaces. Linear Algebra and its Applications, 428(11–12): 2980–2994, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dragomir:2008:NIK

[Dra08b] S. S. Dragomir. New inequalities of the Kantorovich type for bounded linear operators in Hilbert spaces. Linear Algebra and its Applications, 428(11–12):2750– 2760, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Draper:2008:MLA

[Dra08c] Cristina Draper. Models of the Lie algebra F_{4}. Linear Algebra and its Applications, 428(11–12):2813– 2839, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Diaz-Ramos:2004:NSA

[DRGR04] J. Carlos Díaz-Ramos and Eduardo García-Río. A note on the structure of algebraic curvature tensors. Linear Algebra and its Applications, 382(1):271–277, May 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Driessel:2004:CCF

[Dri04] Kenneth R. Driessel. On computing canonical forms using flows. Linear Algebra and its Applications, 379(1): 353–379, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drissi:2008:CNS

[Dri08]  Driss Drissi.  A charac-

terization of normal semigroups. Linear Algebra and its Applications, 428(8–9): 1883–1885, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drnovsek:2002:CCI

[DRR02] Roman Drnovšek, Heydar Radjavi, and Peter Rosenthal. A characterization of commutators of idempotents. Linear Algebra and its Applications, 347(1–3):91–99, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
32/abstract.html.

Drury:2002:CCB

[Dru02] S. W. Drury. The canonical correlations of a 2 × 2 block matrix with given eigenvalues. Linear Algebra and its Applications, 354(1–3): 103–117, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
36/abstract.html.

Druilhet:2004:COE

[Dru04a] Pierre Druilhet. Conditions for optimality in experimental designs. Linear Algebra and its Applications, 388(1):

147–157, September 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drury:2004:MTM

[Dru04b] S. W. Drury. Maximizing traces of matrix functions. Linear Algebra and its Applications, 387(1):221– 234, August 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drury:2007:OSM

[Dru07] S. W. Drury. OMC for scalar multiples of unitaries. Linear Algebra and its Applications, 422(1):318–325, April 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drury:2008:SCO

[Dru08a]  S. W. Drury.  Symbolic

calculus of operators with unit numerical radius. Linear Algebra and its Applications, 428(8–9):2061–2069, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drury:2008:NIN

[Dru08b] S. W. Drury. Von Neumann’s inequality for noncommuting contractions. Linear Algebra and its Applications, 428 (1):305–315, January 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Druskin:2008:MLA

[Dru08c] Vladimir Druskin. On monotonicity of the Lanczos approximation to the matrix exponential. Linear Algebra and its Applications, 429(7): 1679–1683, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dette:2002:MMM

[DS02] Holger Dette and William J. Studden. Matrix measures, moment spaces and Favard’s theorem for the interval [0, 1] and [0, ∞). Linear Algebra and its Applications, 345(1– 3):169–193, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/175/27/
35/abstract.html.

Dette:2003:QFM

[DS03a] Holger Dette and William J. Studden. Quadrature formulas for matrix measures: a geometric approach. Linear Algebra and its Applications, 364(C):33–64, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
26/abstract.html.

Diele:2003:CIF

[DS03b] F. Diele and I. Sgura. Centrosymmetric isospectral flows and some inverse eigenvalue problems. Linear Algebra and its Applications, 366 (C):199–214, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
37/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Davis:2004:PNH

[DS04a] Ch. Davis and A. Salemi. On polynomial numerical hulls of normal matrices. Linear Algebra and its Applications, 383(1):151–161, May 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drton:2004:SVR

[DS04b] Mathias Drton and Udo Schwingenschlögl. Surface volumes of rounding polytopes. Linear Algebra and its Applications, 378(1):71– 91, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

daSilva:2005:RGH

[dS05] Ilda P. F. da Silva. Recursivity and geometry of the hypercube. Linear Algebra and its Applications, 397(1):223– 233, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dodig:2007:CFI

[DS07] Marija Dodig and Marko Stošić. The change of feedback invariants under one row perturbation. Linear Algebra and its Applications, 422(2–3):582–603, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Day:2008:GEC

[DS08a] Jane Day and Wasin So. Graph energy change due to edge deletion. Linear Algebra and its Applications, 428(8– 9):2070–2078, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dietz:2008:VER

[DS08b] S. G. Dietz and C. W. Scherer. Verifying exactness of relaxations for robust semi-definite programs by solving polynomial systems. Linear Algebra and its Applications, 429(7):1758– 1778, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dodig:2008:CSC

[DS08c] Marija Dodig and Fernando C. Silva. Controllability of series connections of arbitrarily many linear systems. Linear Algebra and its Applications, 429(1):122– 141, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dodig:2009:SSS

[DS09] Marija Dodig and Marko Stošić. Singular systems, state feedback problem. Linear Algebra and its Applications, 431(8):1267–1292, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

daSilva:2005:EID

[dSdC05] J. A. Dias da Silva and Henrique F. da Cruz. Equality of immanantal decomposable tensors. Linear Algebra and its Applications, 401(1): 29–46, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

daSilva:2009:CIM

[dSF09] J. A. Dias da Silva and Amélia Fonseca. Constructing integral matrices with given line sums. Linear Algebra and its Applications, 431(9):1553–1563, October 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

daSilva:2007:GDA

[dSG07] J. A. Dias da Silva and Hemar Godinho. Generalized derivations and additive theory II. Linear Algebra and its Applications, 420(1):117– 123, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deSa:2000:WGC

[dSQS00] E. Marques de Sá, J. F. Queiró, and A. P. Santana. Workshop on “Geometric and combinatorial methods in the Hermitian sum spectral problem” University of Coimbra, Portugal, 15–16 July 1999. Linear Algebra and its Applications, 319(1– 3):1–2, November 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/138/24/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/138/24/24/article.
pdf.

daSilva:2005:PGT

[dSR05] A. P. da Silva and A. Sri Ranga. Polynomials generated by a three term recurrence relation: bounds for complex zeros. Linear Algebra and its Applications, 397 (1):299–324, March 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

daSilva:2006:DCT

[dSR06] J. A. Dias da Silva and Fátima Rodrigues. Decomposable critical tensors. Linear Algebra and its Applications, 414(1):172–198, April 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deSa:2008:ESL

[dSS08] E. Marques de Sá and Virgínia Santos. On the extremal structure of least upper bound norms and their dual. Linear Algebra and its Applications, 428(8–9):1928– 1938, April 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

daSilva:2005:ODO

[dST05] J. A. Dias da Silva and Maria M. Torres. On the orthogonal dimension of orbital sets. Linear Algebra and its Applications, 401(1): 77–107, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

daSilva:2006:CAO

[dST06] J. A. Dias da Silva and Maria M. Torres. A combinatorial approach to the orthogonality on critical orbital sets. Linear Algebra and its Applications, 414(2– 3):474–491, April 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Davies:2007:NLS

[DSW07a] Richard Davies, Peng Shi, and Ron Wiltshire. New lower solution bounds of the continuous algebraic Riccati matrix equation. Linear Algebra and its Applications, 427(2–3):242–255, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Deng:2007:HRC

[DSW07b] Aiping Deng, Iwao Sato, and Yaokun Wu. Homomorphisms, representations and characteristic polynomials of digraphs. Linear Algebra and its Applications, 423(2– 3):386–407, June 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

deSa:2000:NKI

[dSZ00a] E. Marques de Sá and Yu-Lin Zhang. The number of Kronecker indices of square pencils of a special kind. Linear Algebra and its Applications, 305(1–3): 15–21, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/120/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/120/25/26/article.
pdf.

deSa:2000:RSD

[dSZ00b] E. Marques de Sá and Yu-Lin Zhang. Ranks of submatrices and the off-diagonal indices of a square matrix. Linear Algebra and its Applications, 305(1–3):1–14, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/120/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/120/25/25/article.
pdf.

Dedieu:2003:PTH

[DT03] Jean-Pierre Dedieu and Françoise Tisseur. Perturbation theory for homogeneous polynomial eigenvalue problems. Linear Algebra and its Applications, 358(1– 3):71–94, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/207/25/
30/abstract.html.

DiVincenzo:2009:GPI

[DT09] Onofrio M. Di Vincenzo and Viviane Ribeiro Tomaz da Silva. On Z_{∈}-graded polynomial identities of the Grassmann algebra. Linear Algebra and its Applications, 431(1– 2):56–72, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Diaz-Toca:2006:CGC

[DTGV06] Gema M. Diaz-Toca and Laureano Gonzalez-Vega. Computing greatest common divisors and squarefree decompositions through matrix methods: The parametric and approximate cases. Linear Algebra and its Applications, 412(2–3):222–246, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dubi:2009:AAS

[Dub09] C. Dubi. An algorithmic approach to simultaneous triangularization. Linear Algebra and its Applications, 430(11– 12):2975–2981, June 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duggal:2000:RKO

[Dug00] B. P. Duggal. Range kernel orthogonality of derivations. Linear Algebra and its Applications, 304(1–3): 103–108, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/118/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/118/25/29/article.
pdf.

Duggal:2001:RKO

[Dug01] B. P. Duggal. Range-kernel orthogonality of the elementary operator X ⟶ ∑_{i=1}^{n} A_{i}XB_{i} − X. Linear Algebra and its Applications, 337(1–3):79–86, November 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/167/27/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/167/27/30/article.
pdf.

Duggal:2004:CRE

[Dug04a] B. P. Duggal. The closure of the range of an elementary operator. Linear Algebra and its Applications, 392(1):305– 319, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duggal:2004:SGR

[Dug04b] B. P. Duggal. Subspace gaps and range-kernel orthogonality of an elementary operator. Linear Algebra and its Applications, 383(1):93– 106, May 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duggal:2005:RCE

[Dug05a] B. P. Duggal. On the range closure of an elementary operator. Linear Algebra and its Applications, 402(1):199– 206, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duggal:2005:RPC

[Dug05b] B. P. Duggal. Riesz projections for a class of Hilbert space operators. Linear Algebra and its Applications, 407(1):140–148, September 15, 2005. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duggal:2006:POS

[Dug06] B. P. Duggal. Polaroid operators satisfying Weyl’s theorem. Linear Algebra and its Applications, 414(1):271– 277, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duggal:2008:EOL

[Dug08]  B. P. Duggal.  An ele-

mentary operator with loghyponormal, p-hyponormal entries. Linear Algebra and its Applications, 428(4): 1109–1116, February 1, 2008. CODEN LAAPAW. ISSN

Dunajeva:2004:SOD

[Dun04] Olga Dunajeva. The secondorder derivatives of matrices of eigenvalues and eigenvectors with an application to generalized F -statistic. Linear Algebra and its Applications, 388(1):159–171, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duncan:2007:DCM

[Dun07] Benton L. Duncan. Derivations for a class of matrix function algebras. Linear Algebra and its Applications, 422(1):67–76, April 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Duplij:2003:SIO

[Dup03] Steven Duplij. On supermatrix idempotent operator semigroups. Linear Algebra and its Applications, 360 (C):59–81, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
30/abstract.html.

Drmac:2000:AEP

[DV00] Zlatko Drmač and Krešimir Veselić. Approximate eigenvectors as preconditioner.

Linear Algebra and its Applications, 309(1–3):191–215, April 15, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/36/article.
pdf.

DiGesu:2001:TVC

[DV01] Vito Di Gesù and Cesare Valenti. Two-view “cylindrical decomposition” of binary images. Linear Algebra and its Applications, 339(1–3): 205–219, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/169/27/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/169/27/39/article.
pdf.

DeLathauwer:2004:DRH

[DV04] Lieven De Lathauwer and Joos Vandewalle. Dimensionality reduction in higherorder signal processing and rank-(R_{1}, R_{2}, …, R_{N}) reduction in multilinear algebra. Linear Algebra and its Applications, 391(1):31– 55, November 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Delvaux:2007:RDS

[DV07a] Steven Delvaux and Marc Van Barel. Rank-deficient submatrices of Kronecker products of Fourier matrices. Linear Algebra and its Applications, 426(2–3):349– 367, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dym:2007:ZDM

[DV07b] Harry Dym and Dan Volok. Zero distribution of matrix polynomials. Linear Algebra and its Applications, 425(2–3):714–738, September 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Delvaux:2008:RDS

[DV08] Steven Delvaux and Marc Van Barel. Rank-deficient submatrices of Fourier matrices. Linear Algebra and its Applications, 429(7):1587– 1605, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Drapal:2009:SMF

[DV09] Aleš Drápal and Petr Vojtěchovský. Symmetric multilinear forms and polarization of polynomials. Linear Algebra and its Applications, 431(5–7):998– 1012, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dewilde:2000:IOF

[DvdV00] P. Dewilde and A.-J. van der Veen. Inner-outer factorization and the inversion of locally finite systems of equations. Linear Algebra and its Applications, 313 (1–3):53–100, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/130/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/130/25/28/article.
pdf.

Dvorsky:2003:IPS

[Dvo03] Alexander Dvorsky. Index of parabolic and seaweed subalgebras of ∫≀_{\}. Linear Algebra and its Applications, 374(1): 127–142, November 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DePierro:2000:SNP

[DW00] Alvaro R. De Pierro and Musheng Wei. Some new properties of the equality constrained and weighted least squares problem. Linear Algebra and its Applications, 320(1–3):145–165, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

gej-ng/10/30/19/139/24/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/34/article.
pdf.

deWerra:2003:PES

[dW03] D. de Werra. Partitioning the edge set of a bipartite graph into chain packings: complexity of some variations. Linear Algebra and its Applications, 368 (C):315–327, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
45/abstract.html.

Deng:2005:CPD

[DW05a] Aiping Deng and Yaokun Wu. Characteristic polynomials of digraphs having a semi-free action. Linear Algebra and its Applications, 408 (1):189–206, October 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Deuflhard:2005:RPC

[DW05b] Peter Deuflhard and Marcus Weber. Robust Perron cluster analysis in conformation dynamics. Linear Algebra and its Applications, 398 (1):161–184, March 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

David:2008:IKS

[DW08] Roden J. A. David and David S. Watkins. An inexact Krylov–Schur algorithm for the unitary eigenvalue problem. Linear Algebra and its Applications, 429(5–6):1213– 1228, September 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Deng:2009:CRD

[DW09a] Chunyuan Deng and Yimin Wei. Characterizations and representations of the Drazin inverse involving idempotents. Linear Algebra and its Applications, 431(9):1526– 1538, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Deng:2009:NDI

[DW09b] Chunyuan Deng and Yimin Wei. A note on the Drazin inverse of an anti-triangular matrix. Linear Algebra and its Applications, 431(10): 1910–1922, October 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Du:2007:PCG

[DWD07] Hong-Ke Du, Wen-Feng Wang, and Ying-Tao Duan. Path connectivity of kgeneralized projectors. Linear Algebra and its Applications, 422(2–3):712–720, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

DiFiore:2001:MAO

[DZ01] Carmine Di Fiore and Paolo Zellini. Matrix algebras in optimal preconditioning. Linear Algebra and its Applications, 335(1–3):1–54, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/165/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/165/27/27/article.
pdf.

Eschenbach:2000:ART

[E^{+}00] Carolyn Eschenbach et al. On almost regular tournament matrices. Linear Algebra and its Applications, 306(1–3): 103–121, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/121/25/33/article.
pdf.

Ed-dari:2005:NIB

[Ed05] Elmouloudi Ed-dari. On the numerical index of Banach spaces. Linear Algebra and its Applications, 403(1): 86–96, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Encinas:2004:NDB

[EdRM04] L. Hernández Encinas, A. Martín del Rey, and J. Muñoz Masqué. Non-degenerate bilinear alternating maps f: V × V → V, dim(V) = 3, over an algebraically closed field. Linear Algebra and its Applications, 387(1):69– 82, August 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eicker:2004:LIS

[EE04] Friedhelm Eicker and Guenter Ewald. Linearly independent set families. Linear Algebra and its Applications, 388(1): 173–191, September 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eisinberg:2007:ESP

[EF07] Alfredo Eisinberg and Giuseppe Fedele. Explicit solution of the polynomial least-squares approximation problem on Chebyshev extrema nodes. Linear Algebra and its Applications, 422(2–3):553–562, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feng:2009:SPS

[eFLW09] Jun e Feng, James Lam, and Yimin Wei. Spectral properties of sums of certain Kronecker products. Linear Algebra and its Applications, 431(9):1691–1701, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:2003:CSI

[EFN^{+}03] Ludwig Elsner, Andreas Frommer, Reinhard Nabben, Hans Schneider, and Daniel B. Szyld. Conditions for strict inequality in comparisons of spectral radii of splittings of different matrices. Linear Algebra and its Applications, 363(C):65–80, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
31/abstract.html.

Ejov:2009:NGR

[EFN09] Vladimir Ejov, Shmuel Friedland, and Giang T. Nguyen. A note on the graph’s resolvent and the multifilar structure. Linear Algebra and its Applications, 431(8):1367– 1379, September 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See erratum [EFN10].

Ejov:2010:ENG

[EFN10] Vladimir Ejov, Shmuel Friedland, and Giang T. Nguyen. Erratum to “A note on the graph’s resolvent and the multifilar structure”. Linear Algebra and its Applications, 432(9):2456, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See [EFN09].

Eisinberg:2001:RVM

[EFS01] A. Eisinberg, G. Franzé, and N. Salerno. Rectangular Vandermonde matrices on Chebyshev nodes. Linear Algebra and its Applications, 338(1–3):27–36, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/29/article.
pdf.

Eidelman:2001:MDV

[EG01] Y. Eidelman and I. Gohberg. A modification of the Dewilde-van der Veen method for inversion of finite structured matrices. Linear Algebra and its Applications, 343–344(1):419–450, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
46/abstract.html.

Eremenko:2002:CPP

[EG02] A. Eremenko and A. Gabrielov. Counterexamples to pole placement by static output feedback. Linear Algebra and its Applications, 351–352: 211–218, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eidelman:2003:FIA

[EG03] Y. Eidelman and I. Gohberg. Fast inversion algorithms for a class of structured operator matrices. Linear Algebra and its Applications, 371(1): 153–190, September 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eidelman:2004:DAB

[EG04] Y. Eidelman and I. Gohberg. Direct approach to the band completion problem. Linear Algebra and its Applications, 385(1):149–185, July 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eidelman:2008:BQM

[EG08] Y. Eidelman and I. Gohberg. Out-of-band quasiseparable matrices. Linear Algebra and its Applications, 429(1):266–289, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eidelman:2007:FRQ

[EGG07] Yuli Eidelman, Israel Gohberg, and Luca Gemignani. On the fast reduction of a quasiseparable matrix to Hessenberg and tridiagonal forms. Linear Algebra and its Applications, 420(1):86– 101, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Estrada:2007:LID

[EGMG07] Moisés Bonilla Estrada, Maricela Figueroa Garcia, Michel Malabre, and Juan Carlos Martínez García. Left invertibility and duality for linear systems. Linear Algebra and its Applications, 425(2–3):345–373, September 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eidelman:2005:EOO

[EGO05a] Y. Eidelman, I. Gohberg, and Vadim Olshevsky. Eigenstructure of order-one-quasiseparable matrices. three-term and two-term recurrence relations. Linear Algebra and its Applications, 405(1):1– 40, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eidelman:2005:QIM

[EGO05b] Yuli Eidelman, Israel Gohberg, and Vadim Olshevsky.

The QR iteration method for Hermitian quasiseparable matrices of an arbitrary order. Linear Algebra and its Applications, 404(1):305– 324, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:2003:SCS

[EH03] Ludwig Elsner and Daniel Hershkowitz. On the spectra of close-to-Schwarz matrices. Linear Algebra and its Applications, 363(C):81–88, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
32/abstract.html.

Elsner:2007:FIM

[EH07] Ludwig Elsner and Daniel Hershkowitz. Functional inequalities for max eigenvalues of nonnegative matrices. Linear Algebra and its Applications, 426(2–3):290–298, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Estrada:2009:RCD

[EH09] Ernesto Estrada and Naomichi Hatano. Returnability in complex directed networks (digraphs). Linear Algebra and its Applications, 430(8– 9):1886–1896, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ehrgott:2008:BSR

[EHR08] M. Ehrgott, A. Holder, and J. Reese. Beam selection in radiotherapy design. Linear Algebra and its Applications, 428(5–6):1272–1312, March 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eick:2004:CAG

[Eic04] Bettina Eick. Computing the automorphism group of a solvable Lie algebra. Linear Algebra and its Applications, 382(1):195–209, May 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eisenstat:2006:PBE

[Eis06] Stanley C. Eisenstat. A perturbation bound for the eigenvalues of a singular diagonalizable matrix. Linear Algebra and its Applications, 416(2–3):742–744, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Eischen:2002:PLS

[EJLS02] Ellen E. Eischen, Charles R. Johnson, Karen Lange, and David P. Stanford. Patterns, linesums, and symmetry. Linear Algebra and its Applications, 357(1–3):

273–289, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
44/abstract.html.

Elsner:2008:RCI

[EK08a] Carsten Elsner and Takao Komatsu. On the residue classes of integer sequences satisfying a linear three term recurrence formula. Linear Algebra and its Applications, 429(4):933–947, August 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:2008:RFL

[EK08b] Carsten Elsner and Takao Komatsu. A recurrence formula for leaping convergents of non-regular continued fractions. Linear Algebra and its Applications, 428(4):824– 833, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Enflo:2001:DNA

[EKS01] Per Enflo, Janice Kover, and Laura Smithies. Denseness for norm attaining operatorvalued functions. Linear Algebra and its Applications, 338(1–3):139–144, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/37/article.
pdf.

Ellers:2001:PTO

[EL01] Erich W. Ellers and Huberta Lausch. Products of transvections in one conjugacy class of the symplectic group over the p-adic numbers. Linear Algebra and its Applications, 327(1–3):151– 180, April 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/38/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/152/
25/38/article.pdf.

Elduque:2000:TAD

[Eld00] Alberto Elduque. On triality and automorphisms and derivations of composition algebras. Linear Algebra and its Applications, 314 (1–3):49–74, July 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/131/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/131/25/26/article.
pdf.

Eldar:2002:LSI

[Eld02a] Yonina C. Eldar. Leastsquares inner product shaping. Linear Algebra and its Applications, 348(1–3): 153–174, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
38/abstract.html.

Elduque:2002:CQF

[Eld02b] Alberto Elduque. Composition of quadratic forms and the Hurwitz-Radon function in characteristic 2. Linear Algebra and its Applications, 348(1–3):87–103, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
33/abstract.html.

Eldar:2006:LSO

[Eld06a] Yonina C. Eldar. Leastsquares orthogonalization using semidefinite programming. Linear Algebra and its Applications, 412(2–3):453– 470, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elduque:2006:LGW

[Eld06b] Alberto Elduque. A Lie grading which is not a semigroup grading. Linear Algebra and its Applications, 418(1):312– 314, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elduque:2009:MNS

[Eld09]  Alberto Elduque.  More

non-semigroup Lie gradings. Linear Algebra and its Applications, 431(9):1603–1606, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elfving:2004:PMS

[Elf04] Tommy Elfving. A projection method for semidefinite linear systems and its applications. Linear Algebra and its Applications, 391(1):57– 73, November 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ellers:2004:CCI

[Ell04] Erich W. Ellers. Conjugacy classes of involutions in the Lorentz group Ω(V) and in SO(V). Linear Algebra and its Applications, 383(1):77– 83, May 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Egleston:2004:NIE

[ELN04] Patricia D. Egleston, Terry D. Lenker, and Sivaram K. Narayan. The nonnegative inverse eigenvalue problem. Linear Algebra and its Applications, 379(1):475– 490, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:2005:NGH

[Els05] L. Elsner. A note on generalized Hessenberg matrices. Linear Algebra and its Applications, 409(1):147–152, November 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Escobar:2003:RVM

[EMBV03] Cesar A. Escobar, José Martínez-Bernal, and Rafael H. Villarreal. Relative volumes and minors in monomial subrings. Linear Algebra and its Applications, 374(1):275– 290, November 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

El-Mallah:2006:AVA

[EMERP06] Mohamed Lamei El-Mallah, Hader Elgendy, Abdellatif Rochdi, and Ángel Rodríguez Palacios. On absolute valued algebras with involution. Linear Algebra and its Applications, 414(1):295– 303, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:2006:PPI

[EMS06] Ludwig Elsner, Augustyn Markiewicz, and Tomasz Szulc. Preface to the Proceedings of the 13th International Workshop on Matrices and Statistics, Bedlewo 2004. Linear Algebra and its Applications, 417(1):ii– iii, August 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Edwards:2005:MCI

[EMT05] Roderick Edwards, Judith J. McDonald, and Michael J. Tsatsomeros. On matrices with common invariant cones with applications in neural and gene networks. Linear Algebra and its Applications, 398(1):37– 67, March 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ejov:2009:CBC

[EN09] Vladimir Ejov and Giang T. Nguyen. Consistent behavior of certain perturbed determinants induced by graphs. Linear Algebra and its Applications, 431(5–7):543–552, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:2003:LRP

[EOvdD03] L. Elsner, D. D. Olesky, and P. van den Driessche. Low rank perturbations and the spectrum of a tridiagonal sign pattern. Linear Algebra and its Applications, 374(1):219– 230, November 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Enflo:2001:HTH

[ES01] Per Enflo and Laura Smithies. Harnack’s theorem for harmonic compact operatorvalued functions. Linear Algebra and its Applications, 336(1–3):21–27, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/166/27/29/article.
pdf.

Ehrhardt:2002:TTT

[ES02] Torsten Ehrhardt and Frank-Olme Speck. Transformation techniques towards the factorization of non-rational 2 × 2 matrix functions. Linear Algebra and its Applications, 353(1–3):53–90, September 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elhashash:2008:GMW

[ES08] Abed Elhashash and Daniel B. Szyld. Generalizations of Mmatrices which may not have a nonnegative inverse. Linear Algebra and its Applications, 429(10):2435–2450, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Estatico:2008:SAU

[ESC08] Claudio Estatico and Stefano Serra-Capizzano. Superoptimal approximation for unbounded symbols. Linear Algebra and its Applications, 428(2–3):564–585, January 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

El-Sayed:2005:IMN

[ESP05] Salah M. El-Sayed and Milko G. Petkov. Iterative methods for nonlinear matrix equations X + A^{∗}X^{−α}A = I. Linear Algebra and its Applications, 403 (1):45–52, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Estatico:2005:CSR

[Est05] Claudio Estatico. A classification scheme for regularizing preconditioners, with application to Toeplitz systems. Linear Algebra and its Applications, 397(1):107– 131, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ellers:2005:STE

[EV05] Erich W. Ellers and Oliver Villa. Siegel transformations for even characteristic. Linear Algebra and its Applications, 395(1):163–174, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:2001:MPM

[EvdD01] Ludwig Elsner and P. van den Driessche. Modifying the power method in max algebra. Linear Algebra and its Applications, 332–334 (1):3–13, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/26/article.
pdf.

Elsner:2004:MAP

[EvdD04] L. Elsner and P. van den Driessche. Max-algebra and pairwise comparison matrices. Linear Algebra and its Applications, 385(1):47– 62, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Elsner:2008:BPR

[EvdD08] Ludwig Elsner and P. van den Driessche. Bounds for the Perron root using max eigenvalues. Linear Algebra and its Applications, 428(8–9):2000– 2005, April 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fallat:2004:ROM

[Fal04] Shaun M. Fallat. A remark on oscillatory matrices. Linear Algebra and its Applications, 393(1):139–147, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fan:2002:SCA

[Fan02] Yizheng Fan. Schur complements and its applications to symmetric nonnegative and Z-matrices. Linear Algebra and its Applications, 353(1–3):289–307, September 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fan:2003:DMF

[Fan03a] Jiangnan Fan. Determinants and multiplicative functionals on quaternion matrices. Linear Algebra and its Applications, 369(C):193– 201, August 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
36/abstract.html.

Fan:2003:SIV

[Fan03b] Yi-Zheng Fan. On spectral integral variations of mixed graphs. Linear Algebra and its Applications, 374(1):307– 316, November 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fanai:2004:RNM

[Fan04] Hamid-Reza Fanaï. On raynonsingular matrices. Linear Algebra and its Applications, 376(1):125–134, January 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fang:2007:BEH

[Fan07a] Maozhong Fang. Bounds on eigenvalues of the Hadamard product and the Fan product of matrices. Linear Algebra and its Applications, 425(1): 7–15, August 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fang:2007:CPI

[Fan07b] Maozhong Fang. On the completion of a partial integral matrix to a unimodular matrix. Linear Algebra and its Applications, 422(1): 291–294, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fang:2010:NIE

[Fan10] Maozhong Fang. A note on the inverse eigenvalue problem for symmetric doubly stochastic matrices. Linear Algebra and its Applications, 432(11):2925–2927, June 1, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See original paper [HP04c] and comments [Fan10, XLG^{+}13].

Faouzi:2001:OIS

[Fao01a] Abdelkhalek Faouzi. On the orbit of invariant subspaces of linear operators in finite-dimensional spaces (new proof of a Halmos’s result). Linear Algebra and its Applications, 329(1–3):171– 174, May 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/154/25/34/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/154/
25/34/article.pdf.

Faouzi:2001:VTO

[Fao01b] Abdelkhalek Faouzi. Verticalité en théorie des opérateurs linéaires (étude des paires). Linear Algebra and its Applications, 326(1–3):111–120, March 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/148/24/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/148/24/30/article.
pdf.

Farenick:2004:PMS

[Far04] Douglas R. Farenick. Pure matrix states on operator systems. Linear Algebra and its Applications, 393(1):149– 173, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Farid:2008:DLF

[Far08] F. O. Farid. On the deformation of linear r-fields. Linear Algebra and its Applications, 429(1):249–265, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fu:2004:RST

[FB04] Haoying Fu and Jesse Barlow. A regularized structured total least squares algorithm for high-resolution image reconstruction. Linear Algebra and its Applications, 391 (1):75–98, November 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fragoso:2005:IDP

[FB05] Marcelo D. Fragoso and Jack Baczynski. On an infinite dimensional perturbed Riccati differential equation arising in stochastic control. Linear Algebra and its Applications, 406(1):165–176, September 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feintuch:2004:PAI

[Fei04]  Avraham Feintuch.  On

polynomial approximation of the inverse of an operator. Linear Algebra and its Applications, 389(1):323–328, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feng:2003:ECC

[Fen03] Bao Qi Feng. Equivalence constants for certain matrix norms. Linear Algebra and its Applications, 374(1):247– 253, November 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fernandes:2001:CNM

[Fer01a] Rosário Fernandes. On the covering number of a matroid element. Linear Algebra and its Applications, 337(1– 3):21–35, November 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/167/27/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/

Fernandes:2001:KDO

[Fer01b] Rosário Fernandes. On the kernel of the derivation operator. Linear Algebra and its Applications, 337(1–3): 121–138, November 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/167/27/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/167/27/33/article.
pdf.

Fernandes:2002:CNU

[Fer02] Rosário Fernandes. Covering numbers under small perturbations. Linear Algebra and its Applications, 350(1– 3):105–124, July 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fernando:2003:AOE

[Fer03]  K. V. Fernando.  Accu-

rate ordering of eigenvectors and singular vectors without eigenvalues and singular values. Linear Algebra and its Applications, 374(1):1–17, November 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fernandes:2004:PVM

[Fer04] Rosário Fernandes. On the powers of a vectorial matroid — an equivalence relation. Linear Algebra and its Applications, 381(1):77– 95, April 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fernandes:2005:MPV

[Fer05] Rosário Fernandes. Matrices that preserve the value of the generalized matrix function of the upper triangular matrices. Linear Algebra and its Applications, 401(1): 47–65, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fernandes:2007:MME

[Fer07a] Rosário Fernandes. On the maximum multiplicity of an eigenvalue in a matrix whose graph contains exactly one cycle. Linear Algebra and its Applications, 422(1):1– 16, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fernando:2007:CEI

[Fer07b] K. V. Fernando. Computation of exact inertia and inclusions of eigenvalues (singular values) of tridiagonal (bidiagonal) matrices. Linear Algebra and its Applications, 422(1):77–99, April 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fujii:2002:KCP

[FF02] Jun Ichi Fujii and Masatoshi Fujii. Kolmogorov’s complexity for positive definite matrices. Linear Algebra and its Applications, 341(1–3): 171–180, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
40/abstract.html.

Freitas:2004:RSU

[FF04a] Pedro J. Freitas and Shmuel Friedland. Revisiting the Siegel upper half plane II. Linear Algebra and its Applications, 376(1):45–67, January 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:2004:MTB

[FF04b] Shmuel Friedland and Pedro J. Freitas. p-metrics on and their Busemann compactifications. Linear Algebra and its Applications, 376(1): 1–18, January 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:2004:RSU

[FF04c] Shmuel Friedland and Pedro J. Freitas. Revisiting the Siegel upper half plane I. Linear Algebra and its Applications, 376(1):19– 44, January 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fallat:2003:GOM

[FFM03] Shaun M. Fallat, Miroslav Fiedler, and Thomas L. Markham. Generalized oscillatory matrices. Linear Algebra and its Applications, 359(1–3):79–90, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
32/abstract.html.

Fujii:2007:RIW

[FFN^{+}07] Jun Ichi Fujii, Masatoshi Fujii, Masahiro Nakamura, Josip Pečarić, and Yuki Seo. A reverse inequality for the weighted geometric mean due to Lawson–Lim. Linear Algebra and its Applications, 427(2–3):272–284, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fasino:2002:SCP

[FG02] D. Fasino and L. Gemignani. Structural and computational properties of possibly singular semiseparable matrices. Linear Algebra and its Applications, 340(1–3): 183–198, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
40/abstract.html.

Furuta:2003:CRK

[FG03] Takayuki Furuta and Mariko Giga. A complementary result of Kantorovich type order preserving inequalities by Mićić–Pečarić–Seo. Linear Algebra and its Applications, 369(C):27–40, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
27/abstract.html.

Fital:2008:NFP

[FG08] Sandra Fital and Chun-Hua Guo. A note on the fixedpoint iteration for the matrix equations X ± A^{∗}X^{−1}A = I. Linear Algebra and its Applications, 429(8–9):2098– 2112, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fassbender:2006:CVF

[FGH^{+}06] Heike Fassbender, Michael Griebel, Olga Holtz, G. W. (Pete) Stewart, and Christoph Zenger. Curriculum vitae of Friedrich Ludwig Bauer. Linear Algebra and its Applications, 417(2–3):299–300, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special Issue in honor of Friedrich Ludwig Bauer.

Fleming:2000:CSS

[FGL^{+}00] R. Fleming, G. Grossman, T. Lenker, S. Narayan, and S.-C. Ong. Classes of Schur D-stable matrices. Linear Algebra and its Applications, 306(1–3):15–24, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/121/25/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/121/25/26/article.
                                        pdf.

Faber:2003:PNH

[FGM03] Vance Faber, Anne Greenbaum, and Donald E. Marshall. The polynomial numerical hulls of Jordan blocks and related matrices. Linear Algebra and its Applications, 374(1):231–246, November 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiol:2006:MMD

[FGM06] M. A. Fiol, J. Gimbert, and M. Miller. Multipartite Moore digraphs. Linear Algebra and its Applications, 419(1):234–250, November 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fernandes:2008:SSG

[FGM08] Rosário Fernandes, Helena Gomes, and Enide Andrade Martins. On the spectra of some graphs like weighted rooted trees. Linear Algebra and its Applications, 428(11– 12):2654–2674, June 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Farkas:2004:SPC

[FGR04] András Farkas, András György, and Pál Rózsa. On the spectrum of pairwise comparison matrices. Linear Algebra and its Applications, 385(1): 443–462, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fourie:2007:PRM

[FGR07] J. H. Fourie, G. J. Groenewald, and A. C. M. Ran. Positive real matrices in indefinite inner product spaces and invariant maximal semidefinite subspaces. Linear Algebra and its Applications, 424(2–3): 346–370, July 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiorini:2005:TMN

[FGS05] Stanley Fiorini, Ivan Gutman, and Irene Sciriha. Trees with maximum nullity. Linear Algebra and its Applications, 397(1):245– 251, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Farebrother:2003:MRQ

[FGT03] Richard William Farebrother, Jürgen Groß, and Sven-Oliver Troschke. Matrix representation of quaternions. Linear Algebra and its Applications, 362(C):251–255, March 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/211/25/
                                        45/abstract.html.

Fan:2007:TEO

[FGWG07] Yi-Zheng Fan, Shi-Cai Gong, Yi Wang, and Yu-Bin Gao. On trees with exactly one characteristic element. Linear Algebra and its Applications, 421(2–3):233–242, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feldmann:2000:PRS

[FH00] Sven Feldmann and Georg Heinig. Partial realization for singular systems in standard form. Linear Algebra and its Applications, 318(1–3): 127–144, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/137/24/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/137/24/33/article.
pdf.

Fuhrmann:2001:PCI

[FH01] P. A. Fuhrmann and U. Helmke. On the parametrization of conditioned invariant subspaces and observer theory. Linear Algebra and its Applications, 332–334(1): 265–353, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
41/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/41/article.
pdf.

Freiling:2004:CRM

[FH04] G. Freiling and A. Hochhaus. On a class of rational matrix differential equations arising in stochastic control. Linear Algebra and its Applications, 379(1):43–68, March 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fischer:2006:MMA

[FH06] Rainer Fischer and Thomas Huckle. Multigrid methods for anisotropic BTTB systems. Linear Algebra and its Applications, 417(2–3):314– 334, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fallat:2007:MRS

[FH07a] Shaun M. Fallat and Leslie Hogben. The minimum rank of symmetric matrices described by a graph: A survey. Linear Algebra and its Applications, 426(2–3):558–582, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Foata:2007:SEQ

[FH07b] Dominique Foata and Guo-Niu Han. Specializations and extensions of the quantum MacMahon Master Theorem. Linear Algebra and its Applications, 423(2–3):445– 455, June 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fassbender:2003:CMV

[FI03] Heike Fassbender and Khakim D. Ikramov. Computing matrixvector products with centrosymmetric and centrohermitian matrices. Linear Algebra and its Applications, 364 (C):235–241, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
44/abstract.html.

Fassbender:2005:SOS

[FI05] Heike Faßbender and Kh. D. Ikramov. Several observations on symplectic, Hamiltonian, and skew-Hamiltonian matrices. Linear Algebra and its Applications, 400(1): 15–29, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fassbender:2007:SOY

[FI07] H. Faßbender and Kh. D. Ikramov. Some observations on the Youla form and conjugate-normal matrices. Linear Algebra and its Applications, 422(1):29– 38, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fassbender:2008:CNM

[FI08a] H. Faßbender and Kh. D. Ikramov. Conjugate-normal matrices: A survey. Linear Algebra and its Applications, 429(7):1425–1441, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fassbender:2008:NUT

[FI08b] H. Faßbender and Kh. D. Ikramov. A note on an unusual type of polar decomposition. Linear Algebra and its Applications, 429(1): 42–49, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2001:EAM

[Fie01a] Miroslav Fiedler. Equilibrated anti-Monge matrices. Linear Algebra and its Applications, 335(1–3):151– 156, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/165/27/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/165/27/35/article.
pdf.

Fiedler:2001:SIP

[Fie01b] Miroslav Fiedler. Some inverse problems for elliptic matrices with zero diagonal. Linear Algebra and its Applications, 332–334(1): 197–204, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/37/article.
pdf.

Fiedler:2003:BM

[Fie03a] Miroslav Fiedler. Basic matrices. Linear Algebra and its Applications, 373(1):143– 151, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2003:MPB

[Fie03b] Miroslav Fiedler. Moore– Penrose biorthogonal systems in Euclidean spaces. Linear Algebra and its Applications, 362(C):137–143, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/211/25/
36/abstract.html.

Fiedler:2003:NCM

[Fie03c] Miroslav Fiedler. A note on companion matrices. Linear Algebra and its Applications, 372(1):325–331, October 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2004:CBM

[Fie04]  Miroslav Fiedler.  Com-

plementary basic matrices. Linear Algebra and its Applications, 384(1):199–206, June 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2005:GL

[Fie05a] Miroslav Fiedler. Geometry of the Laplacian. Linear Algebra and its Applications, 403(1):409–413, July 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2005:ISB

[Fie05b]  Miroslav Fiedler.  Inver-

sion of e-simple block matrices. Linear Algebra and its Applications, 400(1):231– 241, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2005:NSN

[Fie05c] Miroslav Fiedler. A note on sign-nonsingular matrices. Linear Algebra and its Applications, 408(1):14– 18, October 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2006:SPM

[Fie06] Miroslav Fiedler. Subtotally positive and Monge matrices. Linear Algebra and its Applications, 413(1):177– 188, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2008:IPF

[Fie08] Miroslav Fiedler. Intrinsic products and factorizations of matrices. Linear Algebra and its Applications, 428(1): 5–13, January 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2009:SOM

[Fie09] Miroslav Fiedler. Suborthogonality and orthocentricity of matrices. Linear Algebra and its Applications, 430(1):296– 307, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fosner:2007:INB

[FIL07] Maja Fosner, Dijana Ilisević, and Chi-Kwong Li. G-invariant norms and bicircular projections. Linear Algebra and its Applications, 420(2–3):596–608, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Finzel:2000:MPS

[Fin00] M. Finzel. Metric projection and stratification of the Grassmannian. Linear Algebra and its Applications, 308(1–3):109–119, March 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/125/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/125/25/30/article.
                                        pdf.

Fiol:2001:PDR

[Fio01] M. A. Fiol. On pseudodistance-regularity. Linear Algebra and its Applications, 323(1–3):145–165, January 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiol:2005:SBD

[Fio05] M. A. Fiol. Spectral bounds and distance-regularity. Linear Algebra and its Applications, 397(1):17–33, March 1, 2005. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furtado:2002:PMP

[FIS02] Susana Furtado, Laura Iglésias, and Fernando C. Silva. Products of matrices with prescribed spectra and ranks. Linear Algebra and its Applications, 340(1–3):137–147, January 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/172/27/
                                        36/abstract.html.

Fitzgerald:2001:NSS

[Fit01] Robert W. Fitzgerald. Norms of sums of squares. Linear Algebra and its Applications, 325(1–3):1–6, March 1, 2001.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/147/24/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/147/24/24/article.
pdf.

Furtado:2002:NUS

[FJ02] Susana Furtado and Charles R. Johnson. On the number of unitary similarity classes in a C-S equivalence class: the normal case. Linear Algebra and its Applications, 348(1– 3):193–202, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
41/abstract.html.

Furtado:2003:SVU

[FJ03] Susana Furtado and Charles R. Johnson. Spectral variation under congruence for a nonsingular matrix with 0 on the boundary of its field of values. Linear Algebra and its Applications, 359(1– 3):67–78, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
31/abstract.html.

Furtado:2004:CAG

[FJ04] Susana Furtado and Charles R. Johnson. Congruential automorphism groups of general matrices. Linear Algebra and its Applications, 376(1):291– 298, January 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furtado:2005:USC

[FJ05] Susana Furtado and Charles R. Johnson. Unitary similarity classes within the cospectralcongruence class of a matrix. Linear Algebra and its Applications, 394(1):291– 307, January 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fang:2006:LMP

[FJ06] Li Fang and Guoxing Ji. Linear maps preserving products of positive or Hermitian matrices. Linear Algebra and its Applications, 419(2–3):601– 611, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fallat:2007:HPT

[FJ07] Shaun M. Fallat and Charles R. Johnson. Hadamard powers and totally positive matrices. Linear Algebra and its Applications, 423(2–3):420– 427, June 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furtado:2008:VJS

[FJL08] Susana Furtado, Charles Johnson, and Jenna Le. Variation in Jordan structure under congruence: The Nilpotent case. Linear Algebra and its Applications, 429(8–9): 1970–1982, October 16, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fang:2007:MPI

[FJP07] Li Fang, Guoxing Ji, and Yongfeng Pang. Maps preserving the idempotency of products of operators. Linear Algebra and its Applications, 426(1):40–52, October 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fallat:2000:MCU

[FJTU00] Shaun M. Fallat, Charles R. Johnson, Juan R. Torregrosa, and Ana M. Urbano. P -matrix completions under weak symmetry assumptions. Linear Algebra and its Applications, 312 (1–3):73–91, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/129/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/129/25/28/article.
pdf.

Foschi:2004:CEM

[FK04] Paolo Foschi and Erricos J. Kontoghiorghes. A computationally efficient method for solving SUR models with orthogonal regressors. Linear Algebra and its Applications, 388(1):193–200, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fujii:2006:AHI

[FK06] Masatoshi Fujii and Eizaburo Kamei. Ando–Hiai inequality and Furuta inequality. Linear Algebra and its Applications, 416(2–3):541– 545, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:2007:AVC

[FK07] Shmuel Friedland and Christian Krattenthaler. 2-Adic valuations of certain ratios of products of factorials and applications. Linear Algebra and its Applications, 426 (1):159–189, October 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fritzsche:2003:MPR

[FKL03] Bernd Fritzsche, Bernd Kirstein, and Andreas Lasarow. On a moment problem for rational matrix-valued functions. Linear Algebra and its Applications, 372(1):1– 31, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fanizza:2007:PPM

[FKLN07] Giovanna Fanizza, Johan Karlsson, Anders Lindquist, and Ryozo Nagamune. Passivitypreserving model reduction by analytic interpolation. Linear Algebra and its Applications, 425(2–3):608–633, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fritzsche:2002:BCP

[FKM02] Bernd Fritzsche, Bernd Kirstein, and Mirko Mosch. On block completion problems for Arov-normalized j_{qq} J_{q}-elementary factors. Linear Algebra and its Applications, 346(1–3):273–291, May 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/189/27/
40/abstract.html.

Fujii:2002:QFC

[FKN02] Masatoshi Fujii, Eizaburo Kamei, and Ritsuo Nakamoto. On a question of Furuta on chaotic order. Linear Algebra and its Applications, 341 (1–3):119–127, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/171/27/
                                        36/abstract.html.

Fang:2000:ANN

[FKP00] Kai-Tai Fang, Tõnu Kollo, and Anne-Mai Parring. Approximation of the non-null distribution of generalized T^{2}-statistics. Linear Algebra and its Applications, 321(1– 3):27–46, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/140/24/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/140/24/27/article.
                                        pdf.

Fallat:2003:GAC

[FKP03] Shaun M. Fallat, Steve Kirkland, and Sukanta Pati. On graphs with algebraic connectivity equal to minimum edge density. Linear Algebra and its Applications, 373(1):31– 50, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fritzsche:2009:SPS

[FKR09] Bernd Fritzsche, Bernd Kirstein, and Uwe Raabe. On the structure of J-Potapov sequences. Linear Algebra and its Applications, 430(8– 9):2324–2361, April 15, 2009. CODEN LAAPAW. ISSN

Fritzsche:2009:RCP

[FKRS09] B. Fritzsche, B. Kirstein, U. Raabe, and K. Sieber. Representations of J-central J-Potapov functions in both nondegenerate and degenerate cases. Linear Algebra and its Applications, 431(5–7): 1027–1053, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Filar:2002:CLA

[FKS02] Jerzy Filar, Henry A. Krieger, and Zamir Syed. Cesaro limits of analytically perturbed stochastic matrices. Linear Algebra and its Applications, 353(1–3):227–243, September 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fritzsche:2009:ECI

[FKS09] B. Fritzsche, B. Kirstein, and L. A. Sakhnovich. Extremal classical interpolation problems (matrix case). Linear Algebra and its Applications, 430(2–3):762–781, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuichi:2009:TIP

[FKY09] Shigeru Furuichi, Ken Kuriyama, and Kenjiro Yanagi. Trace inequalities for products of matrices. Linear Algebra and its Applications, 430(8–9):2271– 2276, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fan:2003:GSN

[FL03] Yizheng Fan and Jiongsheng Li. On graphs with small number of Laplacian eigenvalues greater than two. Linear Algebra and its Applications, 360(C):207–213, February 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
38/abstract.html.

Feldmann:2004:LSA

[FL04] Sven Feldmann and Patrick Lang. A least squares approach to reduce stable discrete linear systems preserving their stability. Linear Algebra and its Applications, 381(1):141–163, April 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fallat:2007:COM

[FL07a] Shaun Fallat and Xiao Ping Liu. A class of oscillatory matrices with exponent n − 1. Linear Algebra and its Applications, 424(2–3):466– 479, July 15, 2007. CO-

Fu:2007:QML

[FL07b] Amy M. Fu and Alain Lascoux. A question of Muir and Littlewood’s products. Linear Algebra and its Applications, 420(2–3):508–516, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Favati:2002:PFP

[FLM02] P. Favati, G. Lotti, and O. Menchi. A polynomial fit preconditioner for band Toeplitz matrices in image reconstruction. Linear Algebra and its Applications, 346 (1–3):177–197, May 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/189/27/
36/abstract.html.

Favati:2000:SIL

[FLMR00] Paola Favati, Grazia Lotti, Ornella Menchi, and Francesco Romani. Solution of infinite linear systems by automatic adaptive iterations. Linear Algebra and its Applications, 318(1–3):209–225, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/137/24/

                                        38/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/137/24/38/article.
                                        pdf.

Florentino:2009:SST

[Flo09] Carlos A. A. Florentino. Simultaneous similarity and triangularization of sets of 2 by 2 matrices. Linear Algebra and its Applications, 431(9): 1652–1674, October 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feldmann:2003:PIZ

[FLPW03] Sven Feldmann, Patrick Lang, and Dieter Prätzel-Wolters. Parameter influence on the zeros of network determinants. Linear Algebra and its Applications, 371(1): 125–146, September 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feldmann:2007:ULS

[FLPW07] Sven Feldmann, Patrick Lang, and Dieter Prätzel-Wolters. A unified least squares approach to identify and to reduce continuous asymptotically stable systems. Linear Algebra and its Applications, 426(2–3):674– 689, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fourneau:2004:AIL

[FLQ04] J. M. Fourneau, M. Lecoz, and F. Quessette. Algorithms for an irreducible and lumpable strong stochastic bound. Linear Algebra and its Applications, 386(1):167– 185, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Felipe:2005:YBE

[FLROV05] Raúl Felipe, Nancy López-Reyes, Fausto Ongay, and Raúl Velásquez. Yang– Baxter equations on matrix dialgebras with a bar unit. Linear Algebra and its Applications, 403(1):31– 44, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fillmore:2002:INA

[FLRZ02] P. A. Fillmore, W. E. Longstaff, G. W. Mac-Donald H. Radjavi, and Y. Zhong. Intersections of nest algebras in finite dimensions. Linear Algebra and its Applications, 350(1–3):185– 197, July 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2000:FTN

[FM00a] Miroslav Fiedler and Thomas L. Markham. A factorization of totally nonsingular matrices over a ring with identity. Linear Algebra and its Applications, 304(1–3): 161–171, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/118/25/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/118/25/34/article.
pdf.

Fiedler:2000:GTP

[FM00b] Miroslav Fiedler and Thomas L. Markham. Generalized totally positive matrices. Linear Algebra and its Applications, 306(1–3):87–102, February 15, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/121/25/32/article.
pdf.

Fiedler:2002:GTN

[FM02] Miroslav Fiedler and Thomas L. Markham. Generalized totally nonnegative matrices. Linear Algebra and its Applications, 345(1–3):9–28, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/175/27/
28/abstract.html.

Fiedler:2004:TRB

[FM04a] Miroslav Fiedler and Thomas L. Markham. Two results on basic oscillatory matrices. Linear Algebra and its Applications, 389(1):175–181, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fischer:2004:UAF

[FM04b] Bernd Fischer and Jan Modersitzki. A unified approach to fast image registration and a new curvature based registration technique. Linear Algebra and its Applications, 380(1):107– 124, March 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Farhane:2007:DGN

[FM07a] Lotfi Farhane and Gérard Michon. Dimension group of a non-negative integer irreducible matrix. Linear Algebra and its Applications, 420(2–3):530–539, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ferreira:2007:FGI

[FM07b] Vitor O. Ferreira and Lucia S. I. Murakami. Finitely generated invariants of Hopf algebras on free associative algebras. Linear Algebra and its Applications, 420(1):70– 78, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiol:2007:SSF

[FM07c] M. A. Fiol and M. Mitjana. The spectra of some families of digraphs. Linear Algebra and its Applications, 423(1): 109–118, May 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fassbender:2001:HJC

[FMM01] Heike Faßbender, D. Steven Mackey, and Niloufer Mackey. Hamilton and Jacobi come full circle: Jacobi algorithms for structured Hamiltonian eigenproblems. Linear Algebra and its Applications, 332–334(1):37–80, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/29/article.
pdf.

Fernandez:2005:LAA

[FMM05] Luis M. Fernández and Laura Martín-Martínez. Lie algebras associated with triangular configurations. Linear Algebra and its Applications, 407(1):43–63, September 15, 2005. CODEN

Frommer:2004:SSD

[FMN04] Andreas Frommer, Volker Mehrmann, and Reinhard Nabben. Special section dedicated to the GAMM workshop Applied and Numerical Linear Algebra with special emphasis on Numerical Methods for Structured and Random Matrices. Linear Algebra and its Applications, 380(1):1–2, March 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fallat:2004:SIP

[FMPT04] Shaun Fallat, Judith McDonald, Juan Pena, and Michael Tsatsomeros. Special issue on Positivity in Linear Algebra. Linear Algebra and its Applications, 393(1):1–2, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furtado:2001:SIM

[FMS01] Susana Furtado, Enide Andrade Martins, and Fernando C. Silva. On similarity invariants of matrix commutators. Linear Algebra and its Applications, 335(1–3): 81–93, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/165/27/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/165/27/30/article.
                                        pdf.

Furtado:2005:SIM

[FMS05] Susana Furtado, Enide Andrade Martins, and Fernando C. Silva. On similarity invariants of matrix commutators and Jordan products. Linear Algebra and its Applications, 401(1):453– 466, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fonseca:2006:BOJ

[FMZ06] Miguel Fonseca, João Tiago Mexia, and Roman Zmyślony. Binary operations on Jordan algebras and orthogonal normal models. Linear Algebra and its Applications, 417(1): 75–86, August 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Forster:2000:NRM

[FN00] K.-H. Förster and B. Nagy. Nonnegative realizations of matrix transfer functions. Linear Algebra and its Applications, 311(1–3):107–129, May 15, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/

gej-ng/10/30/19/128/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/128/25/33/article.
pdf.

Fallat:2001:PCT

[FN01] Shaun M. Fallat and Michael Neumann. On Perron complements of totally nonnegative matrices. Linear Algebra and its Applications, 327(1– 3):85–94, April 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/32/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/152/
25/32/article.pdf.

Forster:2003:SEG

[FN03] K.-H. Förster and B. Nagy. On spectra of expansion graphs and matrix polynomials. Linear Algebra and its Applications, 363 (C):89–101, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
33/abstract.html.

Ferrante:2007:UAF

[FN07] Augusto Ferrante and Lorenzo Ntogramatzidis. A unified approach to finite-horizon generalized LQ optimal control problems for discretetime systems. Linear Algebra and its Applications, 425(2–3):242–260, September 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Funk-Neubauer:2009:TPT

[FN09] Darren Funk-Neubauer. Tridiagonal pairs and the qtetrahedron algebra. Linear Algebra and its Applications, 431(5–7):903–925, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:2006:SRM

[FNC06] Shmuel Friedland, Amir Niknejad, and Laura Chihara. A simultaneous reconstruction of missing data in DNA microarrays. Linear Algebra and its Applications, 416(1):8–28, July 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fujii:2001:OIM

[FNS01] Jun Ichi Fujii, Ritsuo Nakamoto, and Yuki Seo. Operator inequalities of Malamud and Wielandt. Linear Algebra and its Applications, 326(1– 3):101–109, March 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/148/24/
                                        29/abstract.html;   http:

//www.elsevier.nl/gej-ng/
10/30/19/148/24/29/article.
pdf.

Forster:2007:UTI

[FNS07] K.-H. Förster, B. Nagy, and M. Szilvási. A uniform type inverse spectrum theorem for entrywise nonnegative matrices. Linear Algebra and its Applications, 421(2–3):338– 344, March 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Frauendorf:2005:MDA

[FNT05] Eckehard Frauendorf, Heinz Neudecker, and Götz Trenkler. Multivariate data, the arithmetic mean and exchangeability of transformations. Linear Algebra and its Applications, 410(1):87– 95, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fujii:2007:GBL

[FNT07] Masatoshi Fujii, Ritsuo Nakamoto, and Masaru Tominaga. Generalized Bebiano– Lemos–Providência inequalities and their reverses. Linear Algebra and its Applications, 426(1):33–39, October 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fraikin:2008:GTA

[FNV08] Catherine Fraikin, Yurii Nesterov, and Paul Van Dooren. A gradient-type algorithm optimizing the coupling between matrices. Linear Algebra and its Applications, 429(5–6):1229–1242, September 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Felipe:2001:AAD

[FO01] R. Felipe and F. Ongay. Algebraic aspects of the discrete KP hierarchy. Linear Algebra and its Applications, 338(1– 3):1–17, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/27/article.
pdf.

Fonseca:2000:FVP

[Fon00] Amélia Fonseca. Families of vectors with prescribed rank partition and a prescribed subfamily. Linear Algebra and its Applications, 305(1–3):131–149, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/120/25/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/

Foreman:2004:CI

[For04] B. Foreman. Conjugacy invariants of Sl(2, H). Linear Algebra and its Applications, 381(1):25–35, April 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fosner:2005:OPM

[Foš05] Ajda Fošner. Order preserving maps on the poset of upper triangular idempotent matrices. Linear Algebra and its Applications, 403(1): 248–262, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ferrer:2001:VDI

[FP01] Josep Ferrer and Ferran Puerta. Versal deformations of invariant subspaces. Linear Algebra and its Applications, 332–334(1):569– 582, August 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
55/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/55/article.
pdf.

Farenick:2002:TIH

[FP02a] Douglas R. Farenick and Panayiotis J. Psarrakos. A triangle inequality in Hilbert modules over matrix algebras. Linear Algebra and its Applications, 341(1–3): 57–67, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
32/abstract.html.

Ferrari:2002:LOA

[FP02b] Luca Ferrari and Renzo Pinzani. A linear operator approach to succession rules. Linear Algebra and its Applications, 348(1–3): 231–246, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
44/abstract.html.

Farenick:2003:STQ

[FP03] Douglas R. Farenick and Barbara A. F. Pidkowich. The spectral theorem in quaternions. Linear Algebra and its Applications, 371(1):75–102, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fornasini:2004:MFD

[FP04] Ettore Fornasini and Raquel Pinto. Matrix fraction descriptions in convolutional coding. Linear Algebra and its Applications, 392(1):119– 158, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fernandes:2007:MRM

[FP07] Rosàrio Fernandes and Cecília Perdigão. The minimum rank of matrices and the equivalence class graph. Linear Algebra and its Applications, 427(2–3):161–170, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ferrante:2001:AAR

[FPP01] Augusto Ferrante, Michele Pavon, and Stefano Pinzoni. Asymmetric algebraic Riccati equation: A homeomorphic parametrization of the set of solutions. Linear Algebra and its Applications, 329(1–3): 137–156, May 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/154/25/32/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/154/
25/32/article.pdf.

Ferrante:2002:SAS

[FPP02] Augusto Ferrante, Giorgio Picci, and Stefano Pinzoni. Silverman algorithm and the structure of discrete-time stochastic systems. Linear Algebra and its Applications, 351–352:219–242, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Finet:2004:NRC

[FQ04] Catherine Finet and Hervé Queffélec. Numerical range of composition operators on a Hilbert space of Dirichlet series. Linear Algebra and its Applications, 377(1): 1–10, January 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fan:2009:NBG

[FQ09] Yi-Zheng Fan and Ke-Shi Qian. On the nullity of bipartite graphs. Linear Algebra and its Applications, 430(11– 12):2943–2949, June 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Freiling:2002:SNR

[Fre02] Gerhard Freiling. A survey of nonsymmetric Riccati equations. Linear Algebra and its Applications, 351–352:243– 270, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Freedman:2004:PBJ

[Fre04] D. Z. Freedman. Perturbation bounds for the joint spectrum of commuting matrices. Linear Algebra and its Applications, 387(1):29– 40, August 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Freund:2008:PTM

[Fre08] Roland W. Freund. On Padé-type model order reduction of J-Hermitian linear dynamical systems. Linear Algebra and its Applications, 429(10):2451–2464, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Friedland:2000:FID

[Fri00] Shmuel Friedland. Finite and infinite dimensional generalizations of Klyachko’s theorem. Linear Algebra and its Applications, 319(1–3): 3–22, November 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/138/24/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/138/24/25/article.
pdf.

Friedland:2006:CPM

[Fri06] Shmuel Friedland. Convergence of products of matrices in projective spaces. Linear Algebra and its Applications, 413(2–3):247–263, March 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fuhrmann:2007:SB

[FRY07] P. A. Fuhrmann, P. Rapisarda, and Y. Yamamoto. On the state of behaviors. Linear Algebra and its Applications, 424(2–3):570–614, July 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fridman:2002:CLS

[FS02] E. Fridman and U. Shaked. H_{∞}-control of linear statedelay descriptor systems: an LMI approach. Linear Algebra and its Applications, 351–352:271–302, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fischer:2003:FHP

[FS03] P. Fischer and J. D. Stegeman. Fractional Hadamard powers of positive semidefinite matrices. Linear Algebra and its Applications, 371(1): 53–74, September 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ferreira:2004:ITP

[FS04a] Cristina Ferreira and Fernando C. Silva. Inertia theorems for pairs of matrices. Linear Algebra and its Applications, 381(1):37– 52, April 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ferreira:2004:SLD

[FS04b] Cristina Ferreira and Fernando C. Silva. On the stabilization of linear discretetime systems. Linear Algebra and its Applications, 390(1): 7–18, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ferreira:2006:ITP

[FS06a] Cristina Ferreira and Fernando C. Silva. Inertia theorems for pairs of matrices, II. Linear Algebra and its Applications, 413(2–3):425– 439, March 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Finesso:2006:NMF

[FS06b] Lorenzo Finesso and Peter Spreij. Nonnegative matrix factorization and Idivergence alternating minimization. Linear Algebra and its Applications, 416(2– 3):270–287, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Futorny:2007:CSF

[FS07] Vyacheslav Futorny and Vladimir V. Sergeichuk.

Classification of sesquilinear forms with the first argument on a subspace or a factor space. Linear Algebra and its Applications, 424(1): 282–303, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Freitag:2008:RQI

[FS08] M. A. Freitag and A. Spence. Rayleigh quotient iteration and simplified Jacobi– Davidson method with preconditioned iterative solves. Linear Algebra and its Applications, 428(8–9):2049– 2060, April 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fujii:2004:KTO

[FST04] Jun Ichi Fujii, Yuki Seo, and Masaru Tominaga. Kantorovich type operator inequalities via the Specht ratio. Linear Algebra and its Applications, 377(1):69– 81, January 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fasino:2000:SCP

[FT00] Dario Fasino and Paolo Tilli. Spectral clustering properties of block multilevel Hankel matrices. Linear Algebra and its Applications, 306(1–3): 155–163, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/121/25/37/article.
pdf.

Farebrother:2005:GQM

[FT05] Richard W. Farebrother and Götz Trenkler. On generalized quadratic matrices. Linear Algebra and its Applications, 410(1):244–253, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feng:2007:ECC

[FT07] Bao Qi Feng and Andrew Tonge. Equivalence constants for certain matrix norms II. Linear Algebra and its Applications, 420(2–3):388–399, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fuhrmann:2002:SB

[Fuh02] P. A. Fuhrmann. A study of behaviors. Linear Algebra and its Applications, 351–352:303–380, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fuhrmann:2008:OT

[Fuh08] P. A. Fuhrmann. Observer theory. Linear Algebra and its Applications, 428(1):44– 136, January 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fujii:2005:TIA

[Fuj05] Jun Ichi Fujii. A trace inequality arising from quantum information theory. Linear Algebra and its Applications, 400(1):141–146, May 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fulton:2000:EMH

[Ful00] William Fulton. Eigenvalues of majorized Hermitian matrices and Littlewood– Richardson coefficients. Linear Algebra and its Applications, 319(1–3):23–36, November 1, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/138/24/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/138/24/26/article.
pdf.

Furuta:2002:OMF

[Fur02] Takayuki Furuta. An operator monotone function and strictly chaotic order. Linear Algebra and its Applications, 341(1–3):101–109, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
34/abstract.html.

Furuta:2003:OIA

[Fur03]  Takayuki Furuta.  Oper-

ator inequalities associated with A log A via Specht ratio. Linear Algebra and its Applications, 375(1):251–273, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuta:2004:PES

[Fur04]  Takayuki Furuta.  Para-

metric extensions of Shannon inequality and its reverse one in Hilbert space operators. Linear Algebra and its Applications, 381(1):219– 235, April 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuta:2005:CLT

[Fur05a] Takayuki Furuta. Convergence of logarithmic trace inequalities via generalized Lie–Trotter formulae. Linear Algebra and its Applications, 396(1):353–372, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuta:2005:CQP

[Fur05b] Takayuki Furuta. A counterexample to the question proposed by Yanagi– Furuichi–Kuriyama on matrix inequalities and related counterexamples. Linear Algebra and its Applications, 397(1):345–353, March 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuta:2005:RII

[Fur05c] Takayuki Furuta. Reverse inequalities involving two relative operator entropies and two relative entropies. Linear Algebra and its Applications, 403(1):24–30, July 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuichi:2006:TIN

[Fur06a] Shigeru Furuichi. Trace inequalities in nonextensive statistical mechanics. Linear Algebra and its Applications, 418(2–3):821–827, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuta:2006:TRI

[Fur06b] Takayuki Furuta. Two reverse inequalities associated with Tsallis relative operator entropy via generalized Kantorovich constant and their applications. Linear Algebra and its Applications, 412(2–3):526–537, January 15, 2006. CO-

Furuta:2007:OII

[Fur07]  Takayuki Furuta.  Op-

erator inequality implying generalized Bebiano–Lemos– Providência one. Linear Algebra and its Applications, 426(2–3):342–348, October 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furtado:2008:PTI

[Fur08a] Susana Furtado. Products of two involutions with prescribed eigenvalues and some applications. Linear Algebra and its Applications, 429(7): 1663–1678, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

furuta:2008:CEO

[fur08b] Takayuki furuta. Concrete examples of operator monotone functions obtained by an elementary method without appealing to Löwner integral representation. Linear Algebra and its Applications, 429(5–6):972–980, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuta:2008:MOP

[Fur08c] Takayuki Furuta. Monotonicity of order preserving operator functions. Linear Algebra and its Applications, 428(4): 1072–1082, February 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuta:2009:LMO

[Fur09] Takayuki Furuta. Log majorization via an order preserving operator inequality. Linear Algebra and its Applications, 431(1–2):132– 138, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fiedler:2004:GHM

[FV04] Miroslav Fiedler and Zdeněk Vavřín. Generalized Hessenberg matrices. Linear Algebra and its Applications, 380(1): 95–105, March 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fornasini:2006:PMA

[FVP06] Ettore Fornasini, Maria Elena Valcher, and Raquel Pinto. A polynomial matrix approach to the structural properties of 2D positive systems. Linear Algebra and its Applications, 413(2–3):458– 473, March 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fan:2008:MLE

[FWG08] Yi-Zheng Fan, Yi Wang, and Yu-Bin Gao. Minimizing the least eigenvalues of unicyclic graphs with application to spectral spread. Linear Algebra and its Applications, 429 (2–3):577–588, July 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Furuichi:2005:NOI

[FYK05] Shigeru Furuichi, Kenjiro Yanagi, and Ken Kuriyama. A note on operator inequalities of Tsallis relative operator entropy. Linear Algebra and its Applications, 407(1): 19–31, September 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feng:2006:MLE

[FYL06] Lihua Feng, Guihai Yu, and Qiao Li. Minimizing the Laplacian eigenvalues for trees with given domination number. Linear Algebra and its Applications, 419(2–3): 648–655, December 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fang:2007:GFO

[FYL07] Xiaochun Fang, Haitao Yang, and Shudong Liu. General forms of operator algebras on Pontrjagin space with neutral invariant subspaces. Linear Algebra and its Applications, 425(1):184– 209, August 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fang:2009:SOE

[FYY09] Xiaochun Fang, Jing Yu, and Hongliang Yao. Solutions to operator equations on Hilbert C^{∗}-modules. Linear Algebra and its Applications, 431(11):2142–2153, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Feng:2007:SRG

[FYZ07] Lihua Feng, Guihai Yu, and Xiao-Dong Zhang. Spectral radius of graphs with given matching number. Linear Algebra and its Applications, 422(1):133–138, April 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Fagnani:2004:MSC

[FZ04] Fabio Fagnani and Sandro Zampieri. Minimal and systematic convolutional codes over finite Abelian groups. Linear Algebra and its Applications, 378(1):31–59, February 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gamage:2006:ETT

[GA06a] Jinadasa Gamage and Malwane M. A. Ananda. An exact test for testing the equality of parameter matrices in two multivariate linear models. Linear Algebra and its Applications, 418(2–3):882– 885, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gugercin:2006:MRL

[GA06b] Serkan Gugercin and Athanasios C. Antoulas. Model reduction of large-scale systems by least squares. Linear Algebra and its Applications, 415 (2–3):290–321, June 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gong:2002:NGC

[GAE02] Zhiming Gong, M. Aldeen, and L. Elsner. A note on a generalized Cramer’s rule. Linear Algebra and its Applications, 340(1–3): 253–254, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
45/abstract.html.

Ganikhodjaev:2006:EPM

[GAM06] Nasir Ganikhodjaev, Hasan Akin, and Farrukh Mukhamedov. On the ergodic principle for Markov and quadratic stochastic processes and its relations. Linear Algebra and its Applications, 416(2–3): 730–741, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gao:2003:LGO

[Gao03] You Gao. Lattices generated by orbits of subspaces under finite singular unitary group and its characteristic polynomials. Linear Algebra and its Applications, 368 (C):243–268, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
41/abstract.html.

Garloff:2003:IAT

[Gar03] Jürgen Garloff. Intervals of almost totally positive matrices. Linear Algebra and its Applications, 363 (C):103–108, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
34/abstract.html.

Gavalec:2000:LMP

[Gav00] Martin Gavalec. Linear matrix period in max-plus algebra. Linear Algebra and its Applications, 307(1–3): 167–182, March 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/37/article.
pdf.

Gavalec:2002:MES

[Gav02] Martin Gavalec. Monotone eigenspace structure in maxmin algebra. Linear Algebra and its Applications, 345(1– 3):149–167, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/175/27/
34/abstract.html.

Gavalec:2003:GTA

[Gav03] Martin Gavalec. The general trapezoidal algorithm for strongly regular maxmin matrices. Linear Algebra and its Applications, 369 (C):319–338, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
45/abstract.html.

Ghosh:2006:UBA

[GB06] Arpita Ghosh and Stephen Boyd. Upper bounds on algebraic connectivity via convex optimization. Linear Algebra and its Applications, 418(2–3):693–707, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gueudet-Chartier:2004:SWT

[GC04] Ghislaine Gueudet-Chartier. Should we teach linear algebra through geometry? Linear Algebra and its Applications, 379(1):491–501, March 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cao:2003:ARO

[gCZ03] Chong guang Cao and Xian Zhang. Additive rankone preserving surjections on symmetric matrix spaces. Linear Algebra and its Applications, 362(C):145–151, March 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/211/25/
                                        37/abstract.html.

Gemignani:2003:SSS

[Gem03] L. Gemignani. A superfast solver for Sylvester’s resultant linear systems generated by a stable and an anti-stable polynomial. Linear Algebra and its Applications, 366 (C):233–255, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/

gej-ng/10/30/19/215/25/
39/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Gemignani:2008:NER

[Gem08] Luca Gemignani. Neville elimination for rank-structured matrices. Linear Algebra and its Applications, 428(4):978– 991, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Genin:2000:HMP

[Gen00] Yves V. Genin. Hankel matrices, positive functions and related questions. Linear Algebra and its Applications, 314 (1–3):137–164, July 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/131/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/131/25/29/article.
pdf.

Geskus:2002:NPT

[Ges02] Ronald B. Geskus. A new proof of a theorem on Mmatrices. Linear Algebra and its Applications, 348(1– 3):139–144, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/

                                        gej-ng/10/30/19/195/27/
                                        36/abstract.html.

Gewurz:2003:EOC

[Gew03] Daniele A. Gewurz. Enumeration of orbits on cycles for linear and affine groups. Linear Algebra and its Applications, 368(C):117–127, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/217/25/
                                        33/abstract.html.

Gladwell:2004:TPI

[GG04] G. M. L. Gladwell and K. Ghanbari. The total positivity interval. Linear Algebra and its Applications, 393(1):197–202, December 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2007:PNR

[GG07] Siwen Guo and Wuwen Guo. Perturbing non-real eigenvalues of non-negative real matrices. Linear Algebra and its Applications, 426(1):199– 203, October 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Giovangigli:2009:PIA

[GG09] Vincent Giovangigli and Benjamin Graille. Projected iterative algorithms for complex symmetric systems arising in magnetized multicomponent transport. Linear Algebra and its Applications, 430(4): 1404–1422, February 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gutierrez-Gutierrez:2007:FBH

[GGCB07] Jesús Gutiérrez-Gutiérrez, Pedro M. Crespo, and Albrecht Böttcher. Functions of banded Hermitian block Toeplitz matrices in signal processing. Linear Algebra and its Applications, 422(2– 3):788–807, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gao:2008:CMS

[GGL08] Xing-Bao Gao, Gene H. Golub, and Li-Zhi Liao. Continuous methods for symmetric generalized eigenvalue problems. Linear Algebra and its Applications, 428(2–3): 676–696, January 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garcia:2009:TDB

[GGM^{+}09] Domingo García, Bogdan C. Grecu, Manuel Maestre, Miguel Martín, and Javier Merí. Two-dimensional Banach spaces with polynomial numerical index zero. Linear Algebra and its Applications, 430(8–9):2488–2500, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gao:2000:SRC

[GH00a] Dao-De Gao and Rong-Bing Huang. Some results on canonical correlation and their applications to a linear model. Linear Algebra and its Applications, 321(1– 3):47–59, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/140/24/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/140/24/28/article.
                                        pdf.

Gmainer:2000:DFN

[GH00b] Johannes Gmainer and Hans Havlicek. A dimension formula for the nucleus of a Veronese variety. Linear Algebra and its Applications, 305(1–3):191–201, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/120/25/
                                        39/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/120/25/39/article.
                                        pdf.

Gan:2003:SCN

[GH03] Tai-Bin Gan and Ting-Zhu Huang. Simple criteria for nonsingular H-matrices. Linear Algebra and its Applications, 374(1):317–326, November 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gath:2006:BLM

[GH06] Eugene G. Gath and Kevin Hayes. Bounds for the largest Mahalanobis distance. Linear Algebra and its Applications, 419(1):93–106, November 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Goodson:2009:CFN

[GH09] Geoffrey R. Goodson and Roger A. Horn. Canonical forms for normal matrices that commute with their complex conjugate. Linear Algebra and its Applications, 430(4):1025–1038, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ghaoui:2001:IEC

[Gha01] Laurent El Ghaoui. Inversion error, condition number, and approximate inverses of uncertain matrices. Linear Algebra and its Applications, 343–344(1): 171–193, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/

                                        gej-ng/10/30/19/174/27/
                                        34/abstract.html.

Gibilisco:2009:CBR

[GHI09] P. Gibilisco, F. Hansen, and T. Isola. On a correspondence between regular and nonregular operator monotone functions. Linear Algebra and its Applications, 430(8– 9):2225–2232, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gregory:2001:SSG

[GHK01] David A. Gregory, Daniel Hershkowitz, and Stephen J. Kirkland. The spread of the spectrum of a graph. Linear Algebra and its Applications, 332–334(1):23–35, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/157/25/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/157/25/28/article.
                                        pdf.

Goodson:2003:QRN

[GHM03] Geoffrey R. Goodson, Roger A. Horn, and Dennis I. Merino. Quasi-real normal matrices and eigenvalue pairings. Linear Algebra and its Applications, 369(C):279– 294, August 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
42/abstract.html.

Gaffke:2002:IDO

[GHO02] Norbert Gaffke, Berthold Heiligers, and Robert Offinger. Isotropic discrete orientation distributions on the 3D special orthogonal group. Linear Algebra and its Applications, 354(1–3): 119–139, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
37/abstract.html.

Garimella:2006:SOE

[GHS06] Ramesh V. Garimella, Volodymyr Hrynkiv, and A. R. Sourour. A solution of an operator equation related to the KdV equation. Linear Algebra and its Applications, 418(2–3): 788–792, October 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gerasimova:2009:MSC

[GHS09] Tatyana G. Gerasimova, Roger A. Horn, and Vladimir V. Sergeichuk. Matrices that are self-congruent only via matrices of determinant one. Linear Algebra and its Applications, 431(9):1620–1632, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

George:2002:USM

[GI02] Alan George and Khakim D. Ikramov. Unitary similarity of matrices with quadratic minimal polynomials. Linear Algebra and its Applications, 349(1–3):11–16, July 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gibilisco:2008:RTU

[GII08] Paolo Gibilisco, Daniele Imparato, and Tommaso Isola. A Robertson-type uncertainty principle and quantum Fisher information. Linear Algebra and its Applications, 428(7):1706–1724, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gil:2001:IPI

[Gil01] M. I. Gil’. On invertibility and positive invertibility of matrices. Linear Algebra and its Applications, 327(1–3): 95–104, April 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej%2Dng/10/
                                        30/19/152/25/33/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej%2Dng/10/30/19/152/
                                        25/33/article.pdf.

Gil:2004:BCV

[Gil04] M. I. Gil’. Bounds for characteristic values of entire matrix pencils. Linear Algebra and its Applications, 390 (1):311–320, October 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gil:2008:ISC

[Gil08] M. I. Gil’. Identities for sums of characteristic values of entire matrix pencils. Linear Algebra and its Applications, 428(4):814– 823, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gingold:2009:FMF

[Gin09] H. Gingold. Factorization of matrix functions and their inverses via power product expansions. Linear Algebra and its Applications, 430(11–12): 2835–2858, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gekhtman:2004:LIP

[GJ04a] Michael Gekhtman and Charles R. Johnson. The linear interpolation problem for totally positive matrices. Linear Algebra and its Applications, 393(1):175–178, December 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gelonch:2004:GFT

[GJ04b] J. Gelonch and C. R. Johnson. Generalization of Flanders’ theorem to matrix triples. Linear Algebra and its Applications, 380(1):151– 171, March 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gutkin:2004:CJN

[GJK04] Eugene Gutkin, Edmond A. Jonckheere, and Michael Karow. Convexity of the joint numerical range: topological and differential geometric viewpoints. Linear Algebra and its Applications, 376 (1):143–171, January 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gomez:2001:QFL

[GJMR01] J. R. Gómez, A. Jiménez-Merchán, and J. Reyes. Quasi-filiform lie algebras of maximum length. Linear Algebra and its Applications, 335(1–3):119–135, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/165/27/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/165/27/33/article.
pdf.

Goberna:2003:LSC

[GJR03] Miguel A. Goberna, Valentín Jornet, and Margarita M. L. Rodríguez. On linear systems containing strict inequalities. Linear Algebra and its Applications, 360(C): 151–171, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
34/abstract.html.

Gulliksson:2002:PBC

[GJW02] M. Gulliksson, Xiao-Qing Jin, and Yi-Min Wei. Perturbation bounds for constrained and weighted least squares problems. Linear Algebra and its Applications, 349(1–3):221–232, July 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gaubert:2007:MTM

[GK07] Stéphane Gaubert and Ricardo D. Katz. The Minkowski theorem for maxplus convex sets. Linear Algebra and its Applications, 421(2–3):356–369, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gaubert:2009:TAP

[GK09] Stéphane Gaubert and Ricardo D. Katz. The tropical analogue of polar cones. Linear Algebra and its Applications, 431(5–7):608–625, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gohberg:2007:CEM

[GKL07] I. Gohberg, M. A. Kaashoek, and L. Lerer. On a class of entire matrix function equations. Linear Algebra and its Applications, 425(2–3):434– 442, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gutman:2009:IEG

[GKMZ09] Ivan Gutman, Dariush Kiani, Maryam Mirzakhah, and Bo Zhou. On incidence energy of a graph. Linear Algebra and its Applications, 431(8):1223–1233, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gohberg:2004:IFT

[GKvS04] I. Gohberg, M. A. Kaashoek, and F. van Schagen. On inversion of finite Toeplitz matrices with elements in an algebraic ring. Linear Algebra and its Applications, 385(1): 381–389, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gonzalez:2000:PDI

[GKW00] N. Castro González, J. J. Koliha, and Yimin Wei. Perturbation of the Drazin inverse for matrices with equal eigenprojections at zero. Linear Algebra and its Applications, 312(1–3): 181–189, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/129/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/129/25/36/article.
pdf.

Goldberg:2000:SS

[GL00] Moshe Goldberg and W. A. J. Luxemburg. Stable subnorms. Linear Algebra and its Applications, 307(1–3): 89–101, March 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/31/article.
pdf.

Gao:2001:PSS

[GL01] Yubin Gao and Jiongsheng Li. On the potential stability of star sign pattern matrices. Linear Algebra and its Applications, 327(1–3): 61–68, April 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej%2Dng/10/
                                        30/19/152/25/30/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej%2Dng/10/30/19/152/
                                        25/30/article.pdf.

Grosser:2003:ABS

[GL03] Benedikt Großer and Bruno Lang. An O(n^{2}) algorithm for the bidiagonal SVD. Linear Algebra and its Applications, 358(1–3): 45–70, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/207/25/
                                        29/abstract.html.

Gluesing-Luerssen:2005:WDC

[GL05a] Heide Gluesing-Luerssen. On the weight distribution of convolutional codes. Linear Algebra and its Applications, 408(1):298–326, October 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gurvits:2005:MPF

[GL05b] Leonid Gurvits and James Ledoux. Markov property for a function of a Markov chain: A linear algebra approach. Linear Algebra and its Applications, 404(1):85– 117, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-

Golub:2006:CME

[GL06] Gene H. Golub and Li-Zhi Liao. Continuous methods for extreme and interior eigenvalue problems. Linear Algebra and its Applications, 415(1):31–51, May 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Geng:2008:SRT

[GL08] Xianya Geng and Shuchao Li. The spectral radius of tricyclic graphs with n vertices and k pendent vertices. Linear Algebra and its Applications, 428(11–12):2639– 2653, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gudder:2009:BIP

[GL09] Stan Gudder and Frédéric Latrémolière. Boolean innerproduct spaces and Boolean matrices. Linear Algebra and its Applications, 431(1– 2):274–296, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gladwell:2002:TPT

[Gla02] G. M. L. Gladwell. Total positivity and Toda flow. Linear Algebra and its Applications, 350(1–3):279–284, July 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gladwell:2004:ITP

[Gla04] G. M. L. Gladwell. Inner totally positive matrices. Linear Algebra and its Applications, 393(1):179–195, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2008:CPI

[gLgWmW08] Xin guo Liu, Wei guo Wang, and Yi min Wei. Continuity properties of the 1-inverse and perturbation bounds for the Drazin inverse. Linear Algebra and its Applications, 429(5–6):1026–1037, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Geng:2009:ITG

[GLL09] Xianya Geng, Shuchao Li, and Xuechao Li. On the index of tricyclic graphs with perfect matchings. Linear Algebra and its Applications, 431(12):2304–2316, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guterman:2000:SGT

[GLP00] Alexander Guterman, Chi-Kwong Li, and PeterŠemrl. Some general techniques on linear preserver problems. Linear Algebra and its Applications, 315(1–3):61–81, August 15, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/132/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/132/25/28/article.
pdf.

Gow:2009:P

[GLQS09] Roderick Gow, Raphael Loewy, João Filipe Queiró, and Vladimir Sergeichuk. Preface. Linear Algebra and its Applications, 430(7): 1725–1729, April 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gao:2007:SPA

[GLS07a] Yubin Gao, Zhongshan Li, and Yanling Shao. Sign patterns allowing nilpotence of index 3. Linear Algebra and its Applications, 424(1): 55–70, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gluesing-Luerssen:2007:SSR

[GLS07b] Heide Gluesing-Luerssen and Gert Schneider. State space realizations and monomial equivalence for convolutional codes. Linear Algebra and its Applications, 425(2–3):518– 533, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2009:LAT

[GLW09] Jun Guo, Zengti Li, and Kaishun Wang. Lattices associated with totally isotropic subspaces in classical spaces. Linear Algebra and its Applications, 431(5–7):1088–1095, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2002:SDP

[gLxW02] Feng gao Li and Yang xian Wang. Subconstituents of dual polar graph in finite classical space III. Linear Algebra and its Applications, 349(1–3):105–123, July 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gross:2004:CAL

[GM04] Jürgen Groß and Augustyn Markiewicz. Characterizations of admissible linear estimators in the linear model. Linear Algebra and its Applications, 388(1):239–248, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gombani:2007:NPI

[GM07a] A. Gombani and György Michaletzky. On the Nevanlinna–Pick interpolation problem: Analysis of the McMillan degree of the solutions. Linear Algebra and its Applications, 425(2–3):486– 517, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gonzalez:2007:LMP

[GM07b] Manuel González and Mostafa Mbekhta. Linear maps on M_{N}(C) preserving the local spectrum. Linear Algebra and its Applications, 427(2–3):176–182, December 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ghorbani:2008:EAE

[GM08a] E. Ghorbani and H. R. Maimani. On eigensharp and almost eigensharp graphs. Linear Algebra and its Applications, 429(11–12):2746– 2753, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grone:2008:ILI

[GM08b] Robert Grone and Russell Merris. Indecomposable Laplacian integral graphs. Linear Algebra and its Applications, 428(7):1565–1570, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Geromel:2009:NMR

[GM09a] José C. Geromel and Rafael C. D. Martins. A new method to H_{∈} robust filter design. Linear Algebra and its Applications, 430(1):145– 154, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guterman:2009:CMS

[GM09b] A. E. Guterman and O. V. Markova. Commutative matrix subalgebras and length function. Linear Algebra and its Applications, 430(7): 1790–1805, April 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garcia:2009:EOL

[GML09] Juan Ignacio Garcia Garcia, Jesus Garcia Miranda, and F. J. Lobillo. Elimination orderings and localization in PBW algebras. Linear Algebra and its Applications, 430(8–9):2133–2148, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Giribet:2008:SSO

[GMP08] J. I. Giribet, A. Maestripieri, and F. Martínez Pería.

Shorting selfadjoint operators in Hilbert spaces. Linear Algebra and its Applications, 428(8–9):1899–1911, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Georgiou:2007:DMI

[GMR^{+}07] Tryphon Georgiou, Oleg Michailovich, Yogesh Rathi, James Malcolm, and Allen Tannenbaum. Distribution metrics and image segmentation. Linear Algebra and its Applications, 425(2–3):663– 672, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ghazal:2000:SOF

[GN00] Ghazal A. Ghazal and Heinz Neudecker. On secondorder and fourth-order moments of jointly distributed random matrices: a survey. Linear Algebra and its Applications, 321(1–3): 61–93, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/29/article.
pdf.

Goldberger:2003:MFI

[GN03] Assaf Goldberger and Michael Neumann. Multilinear functional inequalities involving permanents, determinants, and other multilinear functions of nonnegative matrices and M-matrices. Linear Algebra and its Applications, 369(C):295–310, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/218/25/
                                        43/abstract.html.

Goldberger:2005:PFT

[GN05] Assaf Goldberger and Michael Neumann. Perron–Frobenius theory of seminorms: a topological approach. Linear Algebra and its Applications, 399(1):245–284, April 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2009:LGO

[GN09] Jun Guo and Jizhu Nan. Lattices generated by orbits of flats under finite affinesymplectic groups. Linear Algebra and its Applications, 431(5–7):536–542, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Godjali:2009:HPL

[God09]  Ali Godjali.  Hessenberg

pairs of linear transformations. Linear Algebra and its Applications, 431(9):1579– 1586, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Goldberg:2003:CPG

[Gol03] Felix Goldberg. On completely positive graphs and their complements. Linear Algebra and its Applications, 371(1):45–51, September 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Goldberg:2005:SNT

[Gol05]  Moshe Goldberg.  Stable

norms — From theory to applications and back. Linear Algebra and its Applications, 404(1):223–250, July 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Goldberg:2006:BGB

[Gol06]  Felix Goldberg.  Bound-

ing the gap between extremal Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 416(1): 68–74, July 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Goldberg:2009:CVN

[Gol09] Felix Goldberg. On the Colin de Verdière numbers of Cartesian graph products.

Linear Algebra and its Applications, 431(12):2285–2290, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gonzalez:2005:APR

[Gon05] N. Castro González. Additive perturbation results for the Drazin inverse. Linear Algebra and its Applications, 397 (1):279–297, March 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ghareghani:2007:SCG

[GOTR07] N. Ghareghani, G. R. Omidi, and B. Tayfeh-Rezaie. Spectral characterization of graphs with index at most \sqrt{2 + \sqrt{5}}. Linear Algebra and its Applications, 420(2–3):483–489, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gowda:2000:CFT

[GP00a] M. Seetharama Gowda and T. Parthasarathy. Complementarity forms of theorems of Lyapunov and Stein, and related results. Linear Algebra and its Applications, 320 (1–3):131–144, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/

33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/33/article.
pdf.

Gross:2000:EUG

[GP00b] Jürgen Groß and Simo Puntanen. Estimation under a general partitioned linear model. Linear Algebra and its Applications, 321(1–3): 131–144, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/33/article.
pdf.

Gavalec:2003:SRM

[GP03] Martin Gavalec and Ján Plávka. Strong regularity of matrices in general max– min algebra. Linear Algebra and its Applications, 371(1): 241–254, September 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grubisic:2007:ERR

[GP07] Igor Grubišić and Raoul Pietersz. Efficient rank reduction of correlation matrices. Linear Algebra and its Applications, 422(2–3):629– 653, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garvey:2004:GIF

[GPFC04] Seamus D. Garvey, Uwe Prells, Michael I. Friswell, and Zheng Chen. General isospectral flows for linear dynamic systems. Linear Algebra and its Applications, 385(1):335–368, July 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garcia-Planas:2001:AMP

[GPM01] M. Isabel García-Planas and M. Dolors Magret. Associating matrix pencils to generalized linear multivariable systems. Linear Algebra and its Applications, 332–334(1): 235–256, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/157/25/
                                        39/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/157/25/39/article.
                                        pdf.

Garcia-Planas:2006:RSS

[GPMSZ06] M. Isabel García-Planas, M. Dolors Magret, Vladimir V. Sergeichuk, and Nadya A. Zharko. Rigid systems of second-order linear differential equations. Linear Algebra and its Applications, 414(2– 3):517–532, April 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garcia-Planas:2006:RBD

[GPMT06] M. I. García-Planas, M. D. Magret, and S. Tarragona. Relationship between different equivalence relations in the space of standardizable systems. Linear Algebra and its Applications, 413(2– 3):274–284, March 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Garcia-Planas:2001:GFM

[GPS01] M. Isabel García-Planas and Vladimir V. Sergeichuk. Generic families of matrix pencils and their bifurcation diagrams. Linear Algebra and its Applications, 332–334 (1):165–179, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/35/article.
pdf.

Gow:2009:GES

[GQ09a] Rod Gow and Rachel Quinlan. Galois extensions and subspaces of alternating bilinear forms with special rank properties. Linear Algebra and its Applications, 430(8– 9):2212–2224, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gow:2009:GTL

[GQ09b] Rod Gow and Rachel Quinlan. Galois theory and linear algebra. Linear Algebra and its Applications, 430 (7):1778–1789, April 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gekhtman:2000:NFG

[GR00] Michael I. Gekhtman and Leiba Rodman. Normal forms of generic triangular band matrices and Jordan forms of nilpotent completions. Linear Algebra and its Applications, 308(1– 3):1–29, March 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/125/25/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/125/25/25/article.
                                        pdf.

Golasinski:2002:GR

[GR02] Marek Golasiński and Francisco Gómez Ruiz. On Grassmannians over *-rings. Linear Algebra and its Applications, 355(1–3):161–172, November 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/201/27/
                                        37/abstract.html.

Gatica:2003:HCI

[GR03] María Andrea Gatica and María Julia Redondo. Hochschild cohomology of incidence algebras as one-point extensions. Linear Algebra and its Applications, 365(C):169– 181, May 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
39/abstract.html.

Gladwell:2007:IFPb

[GR07a] G. M. L. Gladwell and Oscar Rojo. Isospectral flows preserving some centrosymmetric structures. Linear Algebra and its Applications, 422(2– 3):839–853, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gladwell:2007:IFPa

[GR07b] G. M. L. Gladwell and Oscar Rojo. Isospectral flows that preserve matrix structure. Linear Algebra and its Applications, 421(1):85– 96, February 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grassmann:2002:REC

[Gra02] Winfried K. Grassmann. Real eigenvalues of certain tridiagonal matrix polynomials, with queueing applications. Linear Algebra and its Applications, 342(1–3): 93–106, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/173/27/
                                        33/abstract.html.

Grasse:2004:VBV

[Gra04a] Kevin A. Grasse. A vectorbundle version of a theorem of V. Doležal. Linear Algebra and its Applications, 392(1): 45–59, November 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grassmann:2004:FEP

[Gra04b] Winfried K. Grassmann. Finding equilibrium probabilities of QBD processes by spectral methods when eigenvalues vanish. Linear Algebra and its Applications, 386(1): 207–223, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gracia:2005:NMT

[Gra05] Juan-Miguel Gracia. Nearest matrix with two prescribed eigenvalues. Linear Algebra and its Applications, 401(1): 277–294, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ganesan:2004:EPR

[GRB04] Ashwin Ganesan, Sheila R.

Ross, and B. Ross Barmish. An extreme point result for convexity, concavity and monotonicity of parameterized linear equation solutions. Linear Algebra and its Applications, 390(1):61– 73, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Greenbaum:2002:GFV

[Gre02]  Anne Greenbaum.  Gen-

eralizations of the field of values useful in the study of polynomial functions of a matrix. Linear Algebra and its Applications, 347(1– 3):233–249, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
41/abstract.html.

Greenbaum:2009:ULB

[Gre09]  Anne Greenbaum.  Up-

per and lower bounds on norms of functions of matrices. Linear Algebra and its Applications, 430(1):52– 65, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Griffith:2000:EPA

[Gri00] Daniel A. Griffith. Eigenfunction properties and approximations of selected incidence matrices employed in spatial analyses. Linear Algebra and its Applications, 321(1–3): 95–112, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/30/article.
pdf.

Griffith:2004:EEA

[Gri04] Daniel A. Griffith. Extreme eigenfunctions of adjacency matrices for planar graphs employed in spatial analyses. Linear Algebra and its Applications, 388(1):201–219, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gross:2000:MPI

[Gro00a] Jürgen Groß. The Moore– Penrose inverse of a partitioned nonnegative definite matrix. Linear Algebra and its Applications, 321(1–3): 113–121, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/31/article.
pdf.

Gross:2000:NDP

[Gro00b] Jürgen Groß. Nonnegativedefinite and positive-definite solutions to the matrix equation AXA^{∗} = B — revisited. Linear Algebra and its Applications, 321(1–3): 123–129, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/32/article.
pdf.

Gross:2001:LPO

[Gro01] Jürgen Groß. Löwner partial ordering and space preordering of Hermitian nonnegative definite matrices. Linear Algebra and its Applications, 326(1–3):215–223, March 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/148/24/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/148/24/38/article.
pdf.

Gross:2006:RSP

[Gro06] Jürgen Groß. Remarks on the sharp partial order and the ordering of squares of matrices. Linear Algebra and its Applications, 417(1):87– 93, August 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Group:2008:ZFS

[Gro08] AIM Minimum Rank Special Graphs Work Group. Zero forcing sets and the minimum rank of graphs. Linear Algebra and its Applications, 428 (7):1628–1648, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ghareghani:2008:GCS

[GRTR08] N. Ghareghani, F. Ramezani, and B. Tayfeh-Rezaie. Graphs cospectral with starlike trees. Linear Algebra and its Applications, 429(11–12):2691– 2701, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gallivan:2006:SRE

[GRV06] K. Gallivan, X. Rao, and P. Van Dooren. Singular Riccati equations stabilizing large-scale systems. Linear Algebra and its Applications, 415(2–3):359–372, June 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guralnick:2000:CPT

[GS00] Robert M. Guralnick and B. A. Sethuraman. Commuting pairs and triples of matrices and related varieties. Linear Algebra and its Applications, 310(1–3):139–148, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/127/25/37/article.
pdf.

Sun:2002:CNA

[gS02] Ji guang Sun. Condition numbers of algebraic Riccati equations in the Frobenius norm. Linear Algebra and its Applications, 350(1– 3):237–261, July 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Godunov:2003:SNA

[GS03a] S. K. Godunov and M. Sadkane. Some new algorithms for the spectral dichotomy methods. Linear Algebra and its Applications, 358(1– 3):173–194, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/207/25/
35/abstract.html.

Sun:2003:PAMb

[gS03b] Ji guang Sun. Perturbation analysis of the matrix equation X + Q + A^{H}(X̂ − C)^{(} − 1)A. Linear Algebra and its Applications, 372(1):

33–51, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gaiduk:2004:GCF

[GS04a] Tat’yana N. Gaiduk and Vladimir V. Sergeichuk. Generic canonical form of pairs of matrices with zeros. Linear Algebra and its Applications, 380(1):241– 251, March 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grishin:2004:FSD

[GS04b] Denis Grishin and Thomas Strohmer. Fast scattered data approximation with Neumann and other boundary conditions. Linear Algebra and its Applications, 391(1):99–123, November 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gao:2005:ETC

[GS05] Yubin Gao and Yanling Shao. Exponents of two-colored digraphs with two cycles. Linear Algebra and its Applications, 407(1):263–276, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2006:SRT

[GS06] Ji-Ming Guo and Jia-Yu Shao. On the spectral radius of trees with fixed diameter. Linear Algebra and its Applications, 413(1):131– 147, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Goncalves:2008:IGN

[GS08] Maria Inez Cardoso Gonçalves and Ahmed Ramzi Sourour. Isometries of a generalized numerical radius. Linear Algebra and its Applications, 429(7):1478–1488, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gao:2009:GEP

[GS09a] Yubin Gao and Yanling Shao. Generalized exponents of primitive two-colored digraphs. Linear Algebra and its Applications, 430(5–6): 1550–1565, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gao:2009:PTC

[GS09b] Yubin Gao and Yanling Shao. On k-primitivity of two classes of digraphs. Linear Algebra and its Applications, 430(11–12):2922– 2928, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guersenzvaig:2009:CFP

[GS09c] Natalio H. Guersenzvaig and Fernando Szechtman. A closed formula for the product in simple integral extensions. Linear Algebra and its Applications, 430(8–9):2464– 2466, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gustafson:2009:SMM

[GS09d] Karl Gustafson and George P. H. Styan. Superstochastic matrices and magic Markov chains. Linear Algebra and its Applications, 430(10):2705– 2715, May 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gutknecht:2009:BGB

[GS09e] Martin H. Gutknecht and Thomas Schmelzer. The block grade of a block Krylov space. Linear Algebra and its Applications, 430(1):174– 185, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gowda:2003:SIB

[GSR03] M. Seetharama Gowda, Yoon Song, and G. Ravindran. On some interconnections between strict monotonicity, globally uniquely solvable, and P properties in semidefinite linear complementarity problems. Linear Algebra and its Applications, 370(1):355– 368, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gowda:2004:SPL

[GST04] M. Seetharama Gowda, Roman Sznajder, and J. Tao. Some P -properties for linear transformations on Euclidean Jordan algebras. Linear Algebra and its Applications, 393(1):203–232, December 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Genin:2002:RCS

[GŞV02] Y. Genin, R. Ştefan, and P. Van Dooren. Real and complex stability radii of polynomial matrices. Linear Algebra and its Applications, 351–352:381–410, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sun:2003:PAMa

[gSX03] Ji gunag Sun and Shu-Fang Xu. Perturbation analysis of the maximal solution of the matrix equation X + A^{∗}X^{1}A = P. II. Linear Algebra and its Applications, 362 (C):211–228, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/211/25/
                                        42/abstract.html.

Gasso:2002:FRM

[GT02a] M. Gassó and Juan R. Torregrosa. A P LU-factorization of rectangular matrices by the Neville elimination. Linear Algebra and its Applications, 357(1–3):163–171, December 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/203/25/
                                        36/abstract.html.

Guo:2002:CSL

[GT02b] Ji-Ming Guo and Shang-Wang Tan. A conjecture on the second largest eigenvalue of a tree with perfect matchings. Linear Algebra and its Applications, 347(1–3):9–15, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/194/27/
                                        28/abstract.html.

Guo:2004:NSL

[GT04] Ji-Ming Guo and Shang-Wang Tan. A note on the second largest eigenvalue of a tree with perfect matchings. Linear Algebra and its Applications, 380(1):125– 134, March 15, 2004. CO-DEN LAAPAW. ISSN 0024-

Griffin:2006:PMPa

[GT06a] Kent Griffin and Michael J. Tsatsomeros. Principal minors, Part I: A method for computing all the principal minors of a matrix. Linear Algebra and its Applications, 419(1):107–124, November 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Griffin:2006:PMPb

[GT06b] Kent Griffin and Michael J. Tsatsomeros. Principal minors, Part II: The principal minor assignment problem. Linear Algebra and its Applications, 419(1):125– 171, November 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gross:2006:IPT

[GT06c] Jürgen Groß and Yongge Tian. Invariance properties of a triple matrix product involving generalized inverses. Linear Algebra and its Applications, 417(1):94– 107, August 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gasso:2008:BFQ

[GT08] Maria T. Gassó and Juan R. Torregrosa. Bidiagonal factorizations and quasioscillatory rectangular matrices. Linear Algebra and its Applications, 429(8–9):1886– 1893, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grassmann:2009:SCT

[GT09] Winfried K. Grassmann and Javad Tavakoli. Spectrum of certain tridiagonal matrices when their dimension goes to infinity. Linear Algebra and its Applications, 431(8):1208–1217, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gowda:2009:SIT

[GTM09] M. Seetharama Gowda, Jiyuan Tao, and Melania Moldovan. Some inertia theorems in Euclidean Jordan algebras. Linear Algebra and its Applications, 430(8–9): 1992–2011, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gross:2001:QFC

[GTT01] Jürgen Groß, Götz Trenkler, and Sven-Oliver Troschke. Quaternions: further contributions to a matrix oriented approach. Linear Algebra and its Applications, 326(1– 3):205–213, March 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/148/24/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/148/24/37/article.
pdf.

Gu:2001:GIM

[Gu01] Chuanqing Gu. Generalized inverse matrix Padé approximation on the basis of scalar products. Linear Algebra and its Applications, 322(1– 3):141–167, January 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/141/24/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/141/24/32/article.
pdf.

Guerry:2008:ESV

[Gue08] Marie-Anne Guerry. On the evolution of stock vectors in a deterministic integervalued Markov system. Linear Algebra and its Applications, 429(8–9):1944–1953, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gugercin:2008:ISK

[Gug08] Serkan Gugercin. An iterative SVD–Krylov based method for model reduction of large-scale dynamical systems. Linear Algebra and its Applications, 428(8–9):1964– 1986, April 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gunnlaugsdottir:2003:MSC

[Gun03] Elísabet Gunnlaugsdóttir. Monoidal structure of the category of u_{q}^{+}-modules. Linear Algebra and its Applications, 365(C):183–199, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
47/abstract.html.

Guo:2002:NMN

[Guo02] Chun-Hua Guo. A note on the minimal nonnegative solution of a nonsymmetric algebraic Riccati equation. Linear Algebra and its Applications, 357(1–3): 299–302, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
47/abstract.html.

Guo:2003:LPL

[Guo03a] Ji-Ming Guo. The limit points of Laplacian spectra of graphs. Linear Algebra and its Applications, 362(C): 121–128, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

Guo:2003:LSR

[Guo03b] Ji-Ming Guo. On the Laplacian spectral radius of a tree. Linear Algebra and its Applications, 368(C):379– 385, July 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
50/abstract.html.

Guo:2004:NSQ

[Guo04] Chun-Hua Guo. Numerical solution of a quadratic eigenvalue problem. Linear Algebra and its Applications, 385(1):391–406, July 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2005:NUB

[Guo05a] Ji-Ming Guo. A new upper bound for the Laplacian spectral radius of graphs. Linear Algebra and its Applications, 400(1):61–66, May 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2005:SLL

[Guo05b] Ji-Ming Guo. On the second largest Laplacian eigenvalue of trees. Linear Algebra and its Applications, 404(1): 251–261, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2005:BGW

[Guo05c] Shu-Guang Guo. On bicyclic graphs whose second largest eigenvalue does not exceed 1. Linear Algebra and its Applications, 407(1):201–210, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2005:SRU

[Guo05d] Shu-Guang Guo. The spectral radius of unicyclic and bicyclic graphs with n vertices and k pendant vertices. Linear Algebra and its Applications, 408(1):78– 85, October 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2006:ELS

[Guo06a] Ji-Ming Guo. The effect on the Laplacian spectral radius of a graph by adding or grafting edges. Linear Algebra and its Applications, 413(1): 59–71, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2006:NET

[Guo06b] Ji-Ming Guo. A note on the kth eigenvalue of trees. Linear Algebra and its Applications, 413(1):148–154, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2006:LSR

[Guo06c] Ji-Ming Guo. On the Laplacian spectral radius of trees with fixed diameter. Linear Algebra and its Applications, 419(2–3):618–629, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2007:NCN

[Guo07a] Chun-Hua Guo. A new class of nonsymmetric algebraic Riccati equations. Linear Algebra and its Applications, 426(2–3):636–649, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2007:SRB

[Guo07b] Shu-Guang Guo. On the spectral radius of bicyclic graphs with n vertices and diameter d. Linear Algebra and its Applications, 422(1): 119–132, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2008:LPL

[Guo08] Ji-Ming Guo. On limit points of Laplacian spectral radii of graphs. Linear Algebra and its Applications, 429(7):

1705–1718, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gustafson:2000:EOT

[Gus00] Karl Gustafson. An extended operator trigonometry. Linear Algebra and its Applications, 319(1–3): 117–135, November 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/138/24/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/138/24/31/article.
pdf.

Gustafson:2002:OTS

[Gus02] Karl Gustafson. Operator trigonometry of statistics and econometrics. Linear Algebra and its Applications, 354 (1–3):141–158, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
38/abstract.html.

Gustafson:2009:OTH

[Gus09]  Karl Gustafson.  Opera-

tor trigonometry of Hotelling correlation, Frobenius condition, Penrose twistor. Linear Algebra and its Applications, 430(10):2762–2770, May 1, 2009. CODEN LAAPAW. ISSN 0024-

Guterman:2001:LPM

[Gut01] Alexander Guterman. Linear preservers for matrix inequalities and partial orderings. Linear Algebra and its Applications, 331 (1–3):75–87, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/31/article.
pdf.

Gutman:2008:GWE

[Gut08] Ivan Gutman. On graphs whose energy exceeds the number of vertices. Linear Algebra and its Applications, 429(11–12):2670–2677, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gracia:2006:SCI

[GV06a] Juan-Miguel Gracia and Francisco E. Velasco. Stability of controlled invariant subspaces. Linear Algebra and its Applications, 418(2–3):416–434, October 15, 2006. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Grubisic:2006:RAP

[GV06b] Luka Grubišić and Krešimir Veselić. On Ritz approximations for positive definite operators I (theory). Linear Algebra and its Applications, 417(2–3):397–422, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gheorghe:2009:QMC

[GV09a] Dana Gheorghe and Florian-Horia Vasilescu. Quotient morphisms, compositions, and Fredholm index. Linear Algebra and its Applications, 431(11):2049–2061, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gracia:2009:NSS

[GV09b] Juan-Miguel Gracia and Francisco E. Velasco. Nearest southeast submatrix that makes multiple a prescribed eigenvalue. Part 1. Linear Algebra and its Applications, 430(4):1196–1215, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gemignani:2008:FQF

[GVD08] Luca Gemignani, Marc Van Barel, and Steven Delvaux. Fast QR factorization of Cauchy-like matrices. Linear Algebra and its Applications, 428(2–3):697–711, January 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Groenen:2004:ICA

[GvdV04] Patrick J. F. Groenen and Michel van de Velden. Inverse correspondence analysis. Linear Algebra and its Applications, 388(1):221–238, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gallivan:2003:MRT

[GVV03] K. Gallivan, A. Vandendorpe, and P. Van Dooren. Model reduction via truncation: an interpolation point of view. Linear Algebra and its Applications, 375(1):115– 134, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ghosh:2002:MDS

[GW02] B. K. Ghosh and X. A. Wang. On minimal degree simultaneous pole assignment problems. Linear Algebra and its Applications, 351–352:411– 433, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gau:2003:CNR

[GW03a] Hwa-Long Gau and Pei Yuan Wu. Condition for the numerical range to contain an elliptic disc. Linear Algebra and its Applications, 364 (C):213–222, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
35/abstract.html.

Gau:2003:FBP

[GW03b] Hwa-Long Gau and Pei Yuan Wu. Finite Blaschke products of contractions. Linear Algebra and its Applications, 368 (C):359–370, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
48/abstract.html.

Gau:2004:CMR

[GW04a] Hwa-Long Gau and Pei Yuan Wu. Companion matrices: reducibility, numerical ranges and similarity to contractions. Linear Algebra and its Applications, 383(1):127– 142, May 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gau:2004:NRC

[GW04b] Hwa-Long Gau and Pei Yuan Wu. Numerical range circumscribed by two polygons. Linear Algebra and its Applications, 382(1):155– 170, May 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gau:2004:NRN

[GW04c] Hwa-Long Gau and Pei Yuan Wu. Numerical range of a normal compression II. Linear Algebra and its Applications, 390(1):121–136, October 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gau:2007:NRC

[GW07] Hwa-Long Gau and Pei Yuan Wu. Numerical ranges of companion matrices. Linear Algebra and its Applications, 421(2–3):202–218, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gau:2008:NRN

[GW08] Hwa-Long Gau and Pei Yuan Wu. Numerical ranges of nilpotent operators. Linear Algebra and its Applications, 429(4):716–726, August 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gau:2009:IFN

[GW09a] Hwa-Long Gau and Pei Yuan Wu. Inner functions of numerical contractions. Linear Algebra and its Applications, 430(8–9):2182–2191, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2009:SDT

[GW09b] Jun Guo and Kaishun Wang. Suborbits of m-dimensional totally isotropic subspaces under finite singular classical groups. Linear Algebra and its Applications, 430(8– 9):2063–2069, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2009:LSR

[GW09c] Shu-Guang Guo and Yan-Feng Wang. The Laplacian spectral radius of tricyclic graphs with n vertices and k pendant vertices. Linear Algebra and its Applications, 431(1–2):139– 147, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2009:ASBa

[GWL09a] Jun Guo, Kaishun Wang, and Fenggao Li. Association schemes based on maximal isotropic subspaces in singular classical spaces. Linear Algebra and its Applications, 430(2–3):747–755, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-

Guo:2009:ASBb

[GWL09b] Jun Guo, Kaishun Wang, and Fenggao Li. Association schemes based on maximal totally isotropic subspaces in singular pseudosymplectic spaces. Linear Algebra and its Applications, 431(10):1898–1909, October 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2006:LSG

[GWW06] Wenbin Guo, Musheng Wei, and Mingjie Wang. On least squares g-inverses and minimum norm g-inverses of a bordered matrix. Linear Algebra and its Applications, 416(2–3):627–642, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gao:2009:LGO

[GX09a] You Gao and Juan Xu. Lattices generated by orbits of subspaces under finite singular pseudo-symplectic groups I. Linear Algebra and its Applications, 431(9):1455–1476, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gu:2009:SSH

[GX09b] Chuanqing Gu and Huiyan Xue. A shift-splitting hierarchical identification method for solving Lyapunov matrix equations. Linear Algebra and its Applications, 430(5– 6):1517–1530, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guo:2009:NMN

[GYY09] Ji-Ming Guo, Weigen Yan, and Yeong-Nan Yeh. On the nullity and the matching number of unicyclic graphs. Linear Algebra and its Applications, 431(8):1293–1301, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guglielmi:2001:APF

[GZ01] N. Guglielmi and M. Zennaro. On the asymptotic properties of a family of matrices. Linear Algebra and its Applications, 322(1–3): 169–192, January 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/141/24/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/141/24/33/article.
                                        pdf.

Guglielmi:2003:LPF

[GZ03] N. Guglielmi and M. Zennaro. On the limit products of a family of matrices. Linear Algebra and its Applications, 362(C):11–27, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/211/25/
27/abstract.html.

Gutman:2006:LEG

[GZ06] Ivan Gutman and Bo Zhou. Laplacian energy of a graph. Linear Algebra and its Applications, 414(1):29–37, April 1, 2006. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Giambruno:2007:MJP

[GZ07] Antonio Giambruno and Mikhail Zaicev. Multialternating Jordan polynomials and codimension growth of matrix algebras. Linear Algebra and its Applications, 422(2–3):372–379, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Guglielmi:2008:AFE

[GZ08] Nicola Guglielmi and Marino Zennaro. An algorithm for finding extremal polytope norms of matrix families. Linear Algebra and its Applications, 428(10):2265– 2282, May 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Gegenhasi:2005:PDT

[GZHT05] Gegenhasi, Jun-Xiao Zhao, Xing-Biao Hu, and Hon-Wah Tam. Pfaffianization of the discrete three-dimensional three wave interaction equation. Linear Algebra and its Applications, 407(1):277–295, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Golub:2000:LSS

[GZZ00] Gene H. Golub, Zhenyue Zhang, and Hongyuan Zha. Large sparse symmetric eigenvalue problems with homogeneous linear constraints: the Lanczos process with inner-outer iterations. Linear Algebra and its Applications, 309(1–3):289–306, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/39/article.
pdf.

Ha:2002:PPS

[Ha02]  Kil-Chan Ha.  Positive

projections onto spin factors. Linear Algebra and its Applications, 348(1–3): 105–113, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
34/abstract.html.

Ha:2003:CAP

[Ha03] Kil-Chan Ha. A class of atomic positive linear maps in matrix algebras. Linear Algebra and its Applications, 359(1–3):277–290, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
46/abstract.html.

Hadjidimos:2004:ECP

[Had04]  A. Hadjidimos.  An ex-

tended compact profile iterative method criterion for sparse H-matrices. Linear Algebra and its Applications, 389(1):329–345, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadjidimos:2005:EER

[Had05] A. Hadjidimos. On the equivalence of extrapolation and Richardson’s iteration and its applications. Linear Algebra and its Applications, 402(1): 165–192, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadeler:2008:QPS

[Had08] K. P. Hadeler. Quiescent phases and stability. Linear Algebra and its Applications, 428(7):1620–1627, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Haemers:2008:SRG

[Hae08] Willem H. Haemers. Strongly regular graphs with maximal energy. Linear Algebra and its Applications, 429(11–12):2719–2723, December 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hagos:2002:SRG

[Hag02] Elias M. Hagos. Some results on graph spectra. Linear Algebra and its Applications, 356(1–3):103–111, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/202/25/
                                        34/abstract.html.

Hagler:2007:OOC

[Hag07] Brian A. Hagler. Ordered OLPS and CF Nth numerators. Linear Algebra and its Applications, 422(1):100– 118, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Han:2000:GIS

[Han00] Guo-Niu Han. Généralisation de l’identité de Scott sur les permanents. (French) [Generalization of Scott’s identity for permanents]. Linear Algebra and its Applications, 311 (1–3):25–34, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/128/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/128/25/27/article.
pdf.

Han:2002:SPC

[Han02a] Bin Han. Symmetry property and construction of wavelets with a general dilation matrix. Linear Algebra and its Applications, 353(1–3):207– 225, September 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hansen:2002:CTF

[Han02b] Frank Hansen. Convex trace functions of several variables. Linear Algebra and its Applications, 341(1–3): 309–315, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
51/abstract.html.

Hanna:2003:CLA

[Han03] L. A.-M. Hanna. On the classification of the Lie algebras L_{r}^{s}. Linear Algebra and its Applications, 370(1):251– 256, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Han:2005:SKR

[Han05] Yang Han. Subrepresentations of Kronecker representations. Linear Algebra and its Applications, 402(1):150– 164, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hanusa:2007:PCM

[Han07] Christopher R. H. Hanusa. Pseudo-centrosymmetric matrices, with applications to counting perfect matchings. Linear Algebra and its Applications, 427(2–3):206–217, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hansen:2009:SOM

[Han09] Frank Hansen. Some operator monotone functions. Linear Algebra and its Applications, 430(2–3):795–799, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hao:2004:CTF

[Hao04] Pengwei Hao. Customizable triangular factorizations of matrices. Linear Algebra and its Applications, 382(1): 135–154, May 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hartfiel:2000:DEG

[Har00a] D. J. Hartfiel. Difference equations in a general setting. Linear Algebra and its Applications, 305(1–3): 173–185, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/120/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/120/25/37/article.
pdf.

Harville:2000:SDT

[Har00b] David A. Harville. Simple derivations for two Jacobians of basic importance in multivariate statistics. Linear Algebra and its Applications, 321(1–3):145–152, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/34/article.
pdf.

Harrison:2004:RPP

[Har04]  Peter G. Harrison.  Re-

versed processes, product forms and a non-product form. Linear Algebra and its Applications, 386(1):359– 381, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hartwig:2005:TMA

[Har05] Brian Hartwig. Three mutually adjacent Leonard pairs. Linear Algebra and its Applications, 408(1):19–39, October 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hartfiel:2006:EPS

[Har06]  D. J. Hartfiel.  Embed-

ded population systems in Markov set-chains. Linear Algebra and its Applications, 418(2–3):489–497, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Harada:2007:MVK

[Har07a]  Tetsuo Harada.  Multi-

plicative versions of Klyachko’s theorem in finite factors. Linear Algebra and its Applications, 425(1):102– 108, August 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hartwig:2007:TAF

[Har07b] Brian Hartwig. The tetrahedron algebra and its finitedimensional irreducible modules. Linear Algebra and its Applications, 422(1):219– 235, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hasanov:2005:PDS

[Has05] Vejdi I. Hasanov. Positive definite solutions of the matrix equations X ± A ∗ X^{−q}A = Q. Linear Algebra and its Applications, 404(1): 166–182, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hermida-Alonso:2003:DFE

[HAT03] Jose A. Hermida-Alonso and M. T. Trobajo. The dynamic feedback equivalence over principal ideal domains. Linear Algebra and its Applications, 368(C):197–208, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
38/abstract.html.

Hazod:2005:NCC

[Haz05] Wilfried Hazod. Normalisers and centralisers of compact matrix groups. An elementary approach. Linear Algebra and its Applications, 410(1):96–111, November 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hayashi:2000:ACM

[HB00] Kentaro Hayashi and Peter M. Bentler. The asymptotic covariance matrix of maximum-likelihood estimates in factor analysis: the case of nearly singular matrix of estimates of unique variances. Linear Algebra and its Applications, 321(1–3): 153–173, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/35/article.
pdf.

Herman:2008:CVC

[HB08] Allen Herman and A. Rahnamai Barghi. The character values of commutative quasi-thin schemes. Linear Algebra and its Applications, 429(11–12):2663–2669, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hasfura-Buenaga:2005:AOP

[HBHS05] Julio-Roberto Hasfura-Buenaga, Allen Holder, and Jeffrey Stuart. The asymptotic optimal partition and extensions of the Nonsubstitution Theorem. Linear Algebra and its Applications, 394(1):145– 167, January 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2006:JIA

[HBLZ06] Li-Ping Huang, Tao Ban, De-Qiong Li, and Kang Zhao. Jordan isomorphisms and additive rank preserving maps on symmetric matrices over PID. Linear Algebra and its Applications, 419(2–3):311– 325, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2003:AMS

[HC03a] Jinchuan Hou and Jianlian Cui. Additive maps on standard operator algebras preserving invertibilities or zero divisors. Linear Algebra and its Applications, 359(1–3): 219–233, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
42/abstract.html.

Hou:2003:RPL

[HC03b] Jinchuan Hou and Jianlian Cui. Rank-1 preserving linear maps on nest algebras. Linear Algebra and its Applications, 369(C):

263–277, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/218/25/
                                        41/abstract.html.

Hu:2004:UTM

[HC04] Yong-Jian Hu and Gong-Ning Chen. A unified treatment for the matrix Stieltjes moment problem. Linear Algebra and its Applications, 380(1):227–239, March 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2006:GBT

[HC06] Li-Ping Huang and Yong-Yu Cai. Geometry of block triangular matrices over a division ring. Linear Algebra and its Applications, 416(2– 3):643–676, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Herman:2008:FAS

[HC08] Gabor T. Herman and Wei Chen. A fast algorithm for solving a linear feasibility problem with application to intensity-modulated radiation therapy. Linear Algebra and its Applications, 428(5– 6):1207–1217, March 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hardouin:2009:ISI

[HCLC09] Laurent Hardouin, Bertrand Cottenceau, Mehdi Lhommeau, and Euriell Le Corronc. Interval systems over idempotent semiring. Linear Algebra and its Applications, 431(5–7):855–862, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Harper:2003:EIP

[HD03] L. H. Harper and D. Dreier. The edge-isoperimetric problem on the 600-vertex regular solid. Linear Algebra and its Applications, 368 (C):209–228, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
39/abstract.html.

Hossain:2009:PSI

[HDA09] S. Hossain, Kjell A. Doksum, and S. Ejaz Ahmed. Positive shrinkage, improved pretest and absolute penalty estimators in partially linear models. Linear Algebra and its Applications, 430(10): 2749–2761, May 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hu:2007:BNP

[HdABC07] Yong-Jian Hu, K. d. A. Boubakar, and Gong-Ning Chen. On boundary Nevanlinna–Pick interpolation for Carathéodory matrix functions. Linear Algebra and its Applications, 423(2–3):209–229, June 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ho:2000:GHB

[HDB00] Ming-Tzu Ho, Aniruddha Datta, and S. P. Bhattacharyya. Generalizations of the Hermite-Biehler theorem: the complex case. Linear Algebra and its Applications, 320(1–3):23–36, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/26/article.
pdf.

Heinig:2001:CHM

[Hei01a] Georg Heinig. Chebyshev-Hankel matrices and the splitting approach for centrosymmetric Toeplitz-plus-Hankel matrices. Linear Algebra and its Applications, 327(1–3):181–196, April 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/39/abstract.
html;   http://www.elsevier.

nl/gej%2Dng/10/30/19/152/
25/39/article.pdf.

Heinig:2001:EMS

[Hei01b] Georg Heinig. Not every matrix is similar to a Toeplitz matrix. Linear Algebra and its Applications, 332–334(1): 519–531, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
52/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/52/article.
pdf.

Heinig:2002:KST

[Hei02a] Georg Heinig. Kernel structure of Toeplitz-plus-Hankel matrices. Linear Algebra and its Applications, 340(1– 3):1–13, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
27/abstract.html.

Heinig:2002:RTM

[Hei02b] Georg Heinig. On the reconstruction of Toeplitz matrix inverses from columns. Linear Algebra and its Applications, 350(1–3):199–212, July 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hendrickson:2007:LSA

[Hen07] Bruce Hendrickson. Latent semantic analysis and Fiedler retrieval. Linear Algebra and its Applications, 421(2– 3):345–355, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Herzog:2003:GMO

[Her03] Gerd Herzog. Generalized M-matrices and ordered Banach algebras. Linear Algebra and its Applications, 363 (C):125–131, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/212/25/
                                        36/abstract.html.

Hasanov:2006:PDS

[HES06] Vejdi I. Hasanov and Salah M. El-Sayed. On the positive definite solutions of nonlinear matrix equation X + A^{∗}X^{−δ}A = Q. Linear Algebra and its Applications, 412(2–3):154–160, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hansen:2004:RMT

[HFR04] Olaf Hansen, Silva Fischer, and Ronny Ramlau. Regularization of Mellin-type inverse problems with an application to oil engineering.

Linear Algebra and its Applications, 391(1):125–147, November 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

He:2009:ALE

[HG09] Dao-Jiang He and Da-Wei Guo. Admissibility of linear estimators with respect to inequality constraints under matrix loss function. Linear Algebra and its Applications, 430(1):241–250, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hegland:2007:CTS

[HGC07] Markus Hegland, Jochen Garcke, and Vivien Challis. The combination technique and some generalisations. Linear Algebra and its Applications, 420(2–3):249– 275, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2002:UGM

[HGW02] Yaoping Hou, Ivan Gutman, and Ching-Wah Woo. Unicyclic graphs with maximal energy. Linear Algebra and its Applications, 356(1–3): 27–36, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/202/25/
                                        28/abstract.html.

Haefner:2000:PGS

[HH00] Jeremy Haefner and Trae Holcomb. The Picard group of a structural matrix algebra. Linear Algebra and its Applications, 304(1–3): 69–101, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/118/25/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/118/25/28/article.
                                        pdf.

Hladnik:2005:OMT

[HH05] Milan Hladnik and John Holbrook. Optimal measures for Toeplitz matrices. Linear Algebra and its Applications, 402(1):319–334, June 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hladnik:2009:EPO

[HH09] Milan Hladnik and John Holbrook. Extreme points and optimal measures. Linear Algebra and its Applications, 430(8–9):2553–2561, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hackney:2009:LIV

[HHL^{+}09] Philip Hackney, Benjamin Harris, Margaret Lay, Lon H. Mitchell, Sivaram K. Narayan, and Amanda Pascoe. Linearly independent vertices and minimum semidefinite rank. Linear Algebra and its Applications, 431(8):1105– 1115, September 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Higham:2001:P

[HHLP01] Nicholas Higham, Roger Horn, Thomas J. Laffey, and Ferran Puerta. Preface. Linear Algebra and its Applications, 332–334(1):1, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/25/article.
pdf.

Haber:2009:CFI

[HHM09] Eldad Haber, Stefan Heldmann, and Jan Modersitzki. A computational framework for image-based constrained registration. Linear Algebra and its Applications, 431(3– 4):459–470, July 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2006:AHP

[HHT06] Min-Jei Huang, Chao-Ya Huang, and Tzong-Mo Tsai.

Applications of Hilbert’s projective metric to a class of positive nonlinear operators. Linear Algebra and its Applications, 413(1):202– 211, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadwin:2000:NRM

[HHW00] D. W. Hadwin, K. J. Harrison, and J. A. Ward. Numerical ranges and matrix completions. Linear Algebra and its Applications, 315(1– 3):145–154, August 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/132/25/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/132/25/33/article.
                                        pdf.

Hadwin:2004:CRN

[HHY04] Don Hadwin, Jinchuan Hou, and Hassan Yousefi. Completely rank-nonincreasing linear maps on spaces of operators. Linear Algebra and its Applications, 383(1):213– 232, May 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2009:NMP

[HHZ09] Jinchuan Hou, Kan He, and Xiuling Zhang. Nonlinear maps preserving numerical radius of indefinite skew products of operators. Linear Algebra and its Applications, 430(8–9):2240–2253, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hasanov:2006:TPE

[HI06] V. I. Hasanov and I. G. Ivanov. On two perturbation estimates of the extreme solutions to the equations X ± A^{∗}X^{−1}A = Q. Linear Algebra and its Applications, 413 (1):81–92, February 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hiai:2009:MEF

[Hia09] Fumio Hiai. Monotonicity for entrywise functions of matrices. Linear Algebra and its Applications, 431(8):1125– 1146, September 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Higham:2000:QFC

[Hig00] Nicholas J. Higham. QR factorization with complete pivoting and accurate computation of the SVD. Linear Algebra and its Applications, 309(1–3):153–174, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/

                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/126/25/34/article.
                                        pdf.

Higham:2006:GPS

[Hig06] Desmond J. Higham. Greedy pathlengths and small world graphs. Linear Algebra and its Applications, 416(2–3): 745–758, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Higham:2008:CSE

[Hig08] Nicholas J. Higham. Cayley, Sylvester, and Early Matrix Theory. Linear Algebra and its Applications, 428(1): 39–43, January 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hille:2003:QCP

[Hil03] Lutz Hille. Quivers, cones and polytopes. Linear Algebra and its Applications, 365 (C):215–237, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/214/25/
                                        29/abstract.html.

Hildebrand:2007:ESS

[Hil07a]  Roland Hildebrand.  Ex-

actness of sums of squares relaxations involving 3 × 3 matrices and Lorentz cones.

Linear Algebra and its Applications, 426(2–3):815–840, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hillar:2007:ABM

[Hil07b] Christopher J. Hillar. Advances on the Bessis–Moussa– Villani trace conjecture. Linear Algebra and its Applications, 426(1):130–142, October 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hildebrand:2008:SDL

[Hil08a] Roland Hildebrand. Semidefinite descriptions of lowdimensional separable matrix cones. Linear Algebra and its Applications, 429(4):901– 932, August 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hill:2008:SIS

[Hil08b] Jordan Dale Hill. Standard identities for skew-symmetric matrices. Linear Algebra and its Applications, 429(4):792– 803, August 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hirzallah:2005:ISP

[Hir05] Omar Hirzallah. Inequalities for sums and products of operators. Linear Algebra and its Applications, 407(1):32– 42, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hirzallah:2009:CIH

[Hir09] Omar Hirzallah. Commutator inequalities for Hilbert space operators. Linear Algebra and its Applications, 431(9):1571–1578, October 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hasanov:2004:IPE

[HIU04] V. I. Hasanov, I. G. Ivanov, and F. Uhlig. Improved perturbation estimates for the matrix equations X ± A^{∗}X^{−1}A = Q. Linear Algebra and its Applications, 379 (1):113–135, March 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2008:AMP

[HJ08] Jinchuan Hou and Meiyan Jiao. Additive maps preserving Jordan zero-products on nest algebras. Linear Algebra and its Applications, 429(1): 190–208, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Han:2009:PFR

[HJM09] Deguang Han, Wu Jing, and Ram N. Mohapatra. Perturbation of frames and Riesz bases in Hilbert C^{∗}-modules. Linear Algebra and its Applications, 431(5–7):746–759, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hogben:2005:CCP

[HJR05] Leslie Hogben, Charles R. Johnson, and Robert Reams. The copositive completion problem. Linear Algebra and its Applications, 408(1):207– 211, October 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hirzallah:2000:MYI

[HK00] Omar Hirzallah and Fuad Kittaneh. Matrix Young inequalities for the Hilbert– Schmidt norm. Linear Algebra and its Applications, 308 (1–3):77–84, March 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/125/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/125/25/28/article.
pdf.

Haemers:2002:HTG

[HK02a] Willem H. Haemers and Elisabeth Kuijken. The Hermitian two-graph and its code. Linear Algebra and its Applications, 356(1–3):

79–93, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/202/25/
32/abstract.html.

Hirzallah:2002:NIW

[HK02b] Omar Hirzallah and Fuad Kittaneh. Norm inequalities for weighted power means of operators. Linear Algebra and its Applications, 341 (1–3):181–193, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
41/abstract.html.

Hershkowitz:2003:PPM

[HK03] Daniel Hershkowitz and Nathan Keller. Positivity of principal minors, sign symmetry and stability. Linear Algebra and its Applications, 364(C):105–124, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
38/abstract.html.

Humphries:2005:TII

[HK05] S. Humphries and C. Krattenthaler. Trace identities from identities for determinants. Linear Algebra and its Applications, 411(1):328– 342, December 1, 2005. CO-DEN LAAPAW. ISSN 0024-

Hirzallah:2007:ISD

[HK07] Omar Hirzallah and Fuad Kittaneh. Inequalities for sums and direct sums of Hilbert space operators. Linear Algebra and its Applications, 424(1):71–82, July 1, 2007. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hwang:2005:NZT

[HKKL05] Suk-Geun Hwang, Ik-Pyo Kim, Si-Ju Kim, and Sang-Gu Lee. The number of zeros of a tight sign-central matrix. Linear Algebra and its Applications, 407(1):296–310, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hwang:2003:TSC

[HKKZ03] Suk-Geun Hwang, Ik-Pyo Kim, Si-Ju Kim, and Xiao-Dong Zhang. Tight signcentral matrices. Linear Algebra and its Applications, 371(1):225–240, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hwang:2003:SIC

[HKSS03] Suk-Geun Hwang, Arnold R. Kräuter, Bryan L. Shader, and Jia-Yu Shao. Special issue on the Combinatorial Matrix Theory Conference. Linear Algebra and its Applications, 373(1):1–3, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Herzog:2000:OSP

[HL00] Gerd Herzog and Roland Lemmert. On ordered spaces of polynomials. Linear Algebra and its Applications, 320 (1–3):199–203, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/39/article.
pdf.

Hou:2001:SNN

[HL01] Yaoping Hou and Jiongsheng Li. Square nearly nonpositive sign pattern matrices. Linear Algebra and its Applications, 327(1–3): 41–51, April 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/28/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/152/
25/28/article.pdf.

Hou:2002:BLE

[HL02] Yaoping Hou and Jiongsheng Li. Bounds on the largest eigenvalues of trees with a given size of matching. Linear Algebra and its Applications, 342(1–3): 203–217, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/173/27/
40/abstract.html.

Han:2003:RPS

[HL03] Yang Han and Mulan Liu. Regular points in system spaces. Linear Algebra and its Applications, 365(C):201– 213, May 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
44/abstract.html.

Han:2004:IPI

[HL04a] Song-Chol Han and Hong-Xing Li. Indices and periods of incline matrices. Linear Algebra and its Applications, 387(1):143–165, August 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Han:2004:IIM

[HL04b] Song-Chol Han and Hong-Xing Li. Invertible incline matrices and Cramer’s rule over inclines. Linear Algebra and its Applications, 389(1): 121–138, September 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Han:2004:SIH

[HL04c] Song-Chol Han and Hong-Xing Li. The semigroup of incline Hall matrices. Linear Algebra and its Applications, 390(1):183–196, October 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hu:2004:WLM

[HL04d] Guang-Da Hu and Mingzhu Liu. The weighted logarithmic matrix norm and bounds of the matrix exponential. Linear Algebra and its Applications, 390(1):145– 154, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hwang:2004:PDS

[HL04e] Suk-Geun Hwang and Eun-Young Lee. Permanents of doubly stochastic matrices with diagonal submatrices of fixed order. Linear Algebra and its Applications, 376(1): 97–108, January 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hentzel:2005:RRN

[HL05a] Irvin Roy Hentzel and Alicia Labra. On representations on right nilalgebras of right nilindex four. Linear Algebra and its Applications, 404(1): 389–400, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hyman:2005:RNH

[HL05b] James M. Hyman and Jia Li. The reproductive number for an HIV model with differential infectivity and staged progression. Linear Algebra and its Applications, 398 (1):101–116, March 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Han:2007:MOI

[HL07a] Song-Chol Han and Hong-Xing Li. Maximum orderindex of matrices over commutative inclines: an answer to an open problem. Linear Algebra and its Applications, 420(1):228–234, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hentzel:2007:GJA

[HL07b] Irvin Roy Hentzel and Alicia Labra. Generalized Jordan algebras. Linear Algebra and its Applications, 422(1): 326–330, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2007:SRM

[HL07c] Li-Ping Huang and Zhuo-Jun Liu. Similarity reduction of matrix over a quaternion division ring. Linear Algebra and its Applications, 427(2–3):317–332, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See corrigendum [HL08b].

Huang:2008:ECT

[HL08a] Fengying Huang and Bolian Liu. Exponents of a class of two-colored digraphs with two cycles. Linear Algebra and its Applications, 429(2– 3):658–672, July 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2008:ESR

[HL08b] Li-Ping Huang and Zhuo-Jun Liu. Erratum to “Similarity reduction of matrix over a quaternion division ring” [Linear Algebra Appl. 428 (2008) 765–780]. Linear Algebra and its Applications, 429 (4):948, August 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See [HL07c].

Huang:2008:WSR

[HL08c] Li-Ping Huang and Zhuo-Jun Liu. WITHDRAWN: Similarity reduction of matrix over a quaternion division ring. Linear Algebra and its Applications, 428(4):765– 780, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2009:SDA

[HL09] Tsung-Ming Huang and Wen-Wei Lin. Structured doubling algorithms for weakly stabilizing hermitian solutions of algebraic Riccati equations. Linear Algebra and its Applications, 430(5– 6):1452–1478, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Han:2004:SEI

[HLG04] Song-Chol Han, Hong-Xing Li, and Yun-Dong Gu. Standard eigenvectors of incline matrices. Linear Algebra and its Applications, 389(1): 235–248, September 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hall:2004:BSP

[HLR04] Frank J. Hall, Zhongshan Li, and Bhaskara Rao. From Boolean to sign pattern matrices. Linear Algebra and its Applications, 393(1):233– 251, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hautphenne:2008:NIE

[HLR08] Sophie Hautphenne, Guy Latouche, and Marie-Ange Remiche. Newton’s iteration for the extinction probability of a Markovian binary tree. Linear Algebra and its Applications, 428(11–12):2791– 2804, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2000:EBS

[HLS00] Tingzhu Huang, Wen Li, and Weiwei Sun. Estimates for bounds of some numerical characters of matrices. Linear Algebra and its Applications, 319(1–3): 137–145, November 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/138/24/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/138/24/32/article.
pdf.

Hadwin:2004:ATM

[HLT04] D. Hadwin, D. R. Larson, and D. Timotin. Approximation theory and matrix completions. Linear Algebra and its Applications, 377(1):165– 179, January 15, 2004. CO-

Hall:2001:SSP

[HLW01] Frank J. Hall, Zhongshan Li, and Di Wang. Symmetric sign pattern matrices that require unique inertia. Linear Algebra and its Applications, 338(1–3): 153–169, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/39/article.
pdf.

Han:2005:NIM

[HLW05] Song-Chol Han, Hong-Xing Li, and Jia-Yin Wang. On nilpotent incline matrices. Linear Algebra and its Applications, 406(1):201–217, September 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hager:2006:OGM

[HLW06] William W. Hager, Yong Liu, and Tan F. Wong. Optimization of generalized mean square error in signal processing and communication. Linear Algebra and its Applications, 416(2–3):815– 834, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2008:SGG

[HLW08] Yaoping Hou, Tiangang Lei, and Chingwah Woo. On the sandpile group of the graph K_{3} × C_{n}. Linear Algebra and its Applications, 428(8– 9):1886–1898, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hoffnung:2006:KTS

[HLY06] Leonard Hoffnung, Ren-Cang Li, and Qiang Ye. Krylov type subspace methods for matrix polynomials. Linear Algebra and its Applications, 415(1):52–81, May 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Haemers:2008:SCL

[HLZ08] Willem H. Haemers, Xiaogang Liu, and Yuanping Zhang. Spectral characterizations of lollipop graphs. Linear Algebra and its Applications, 428(11–12):2415– 2423, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hu:2009:BGE

[HLZ09] Zhiquan Hu, Shuchao Li, and Chunfeng Zhu. Bicyclic graphs with exactly two main eigenvalues. Linear Algebra and its Applications, 431(10):

1848–1857, October 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hansen:2002:CDA

[HM02] Pierre Hansen and Hadrien Mélot. Computers and discovery in algebraic graph theory. Linear Algebra and its Applications, 356(1–3): 211–230, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/202/25/
41/abstract.html.

Huang:2004:PAG

[HM04] Qianglian Huang and Jipu Ma. Perturbation analysis of generalized inverses of linear operators in Banach spaces. Linear Algebra and its Applications, 389(1):355– 364, September 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2006:SCG

[HM06] Qianglian Huang and Jipu Ma. On the semi-continuity of generalized inverses in Banach algebras. Linear Algebra and its Applications, 419(1): 172–179, November 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hinrichsen:2001:MPR

[HMH01] D. Hinrichsen, W. Manthey, and U. Helmke. Minimal partial realization by descriptor systems. Linear Algebra and its Applications, 326 (1–3):45–84, March 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/148/24/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/148/24/27/article.
                                        pdf.

Horn:2009:PD

[HMP09] Roger A. Horn, Dennis I. Merino, and Agnes T. Paras. The ψ_{S} polar decomposition. Linear Algebra and its Applications, 431(8):1249–1256, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hechler:2009:CGF

[HMR09] Jochen Hechler, Bernhard Mößner, and Ulrich Reif. C^{1}-continuity of the generalized four-point scheme. Linear Algebra and its Applications, 430(11–12):3019– 3029, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hauke:2001:IEP

[HMS01] Jan Hauke, Augustyn Markiewicz, and Tomasz Szulc. Interand extrapolatory properties of matrix partial orderings. Linear Algebra and its Applications, 332–334(1): 437–445, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
46/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/46/article.
pdf.

Huang:2008:SSM

[HMT^{+}08] Ting-Zhu Huang, Gu-Fang Mou, Gui-Xian Tian, Zhongshan Li, and Di Wang. Subdirect sums of P (P_{0})-matrices and totally nonnegative matrices. Linear Algebra and its Applications, 429(7):1730– 1743, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadwin:2004:GVM

[HN04] Don Hadwin and Eric Nordgren. A general view of multipliers and composition operators. Linear Algebra and its Applications, 383(1):187– 211, May 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hrivnak:2009:TCL

[HN09] Jiří Hrivnák and Petr Novotný. Twisted cocycles of lie algebras and corresponding invariant functions. Linear Algebra and its Applications, 430(4):1384–1403, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hrivnak:2006:GCP

[HNPT06] J. Hrivnák, P. Novotný, J. Patera, and J. Tolar. Graded contractions of the Pauli graded sl(3, C). Linear Algebra and its Applications, 418(2–3):498–550, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadjidimos:2003:MMI

[HNT03] A. Hadjidimos, D. Noutsos, and M. Tzoumas. More on modifications and improvements of classical iterative schemes for M-matrices. Linear Algebra and its Applications, 364(C):253–279, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/213/25/
                                        41/abstract.html.

Hanke:2000:QNA

[HNV00] Martin Hanke, James G. Nagy, and Curtis Vogel. Quasi-Newton approach to nonnegative image restorations. Linear Algebra and its Applications, 316(1–3): 223–236, September 1, 2000.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/38/article.
pdf.

Han:2003:RCP

[HNX03] Lixing Han, Michael Neumann, and Jianhong Xu. On the roots of certain polynomials arising from the analysis of the Nelder–Mead simplex method. Linear Algebra and its Applications, 363 (C):109–124, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
35/abstract.html.

Holbrook:2001:ACO

[HO01] John Holbrook and Matjaž Omladič. Approximating commuting operators. Linear Algebra and its Applications, 327(1–3):131–149, April 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/37/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/152/
25/37/article.pdf.

Hanzon:2002:SSC

[HO02] Bernard Hanzon and Raimund J. Ober. State space calculations for discrete probability densities. Linear Algebra and its Applications, 350(1– 3):67–87, July 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ho:2003:RES

[Ho03] Mark C. Ho. A rough estimate for the spectral radius of the sampling operator. Linear Algebra and its Applications, 375(1):51– 61, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hochstenbach:2009:JDT

[Hoc09] M. E. Hochstenbach. A Jacobi–Davidson type method for the generalized singular value problem. Linear Algebra and its Applications, 431 (3–4):471–487, July 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hoffman:2000:GVT

[Hof00] Alan J. Hoffman. Gersgorin variations I: on a theme of Pupkov and Solov’ev. Linear Algebra and its Applications, 304(1–3):173–177, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/118/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/118/25/35/article.
pdf.

Hoffman:2008:NLF

[Hof08]  Alan Hoffman.  A note

from LAA’s first editor-in-chief. Linear Algebra and its Applications, 428(1):4, January 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hogben:2000:IMM

[Hog00] Leslie Hogben. Inverse Mmatrix completions of patterns omitting some diagonal positions. Linear Algebra and its Applications, 313 (1–3):173–192, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/130/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/130/25/36/article.
pdf.

Hogben:2001:GTM

[Hog01] Leslie Hogben. Graph theoretic methods for matrix completion problems. Linear Algebra and its Applications, 328(1–3):161–202, May 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej%2Dng/10/
                                        30/19/153/25/32/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej%2Dng/10/30/19/153/
                                        25/32/article.pdf.

Hogben:2002:SMS

[Hog02] Leslie Hogben. The symmetric M-matrix and symmetric inverse M-matrix completion problems. Linear Algebra and its Applications, 353(1–3): 159–168, September 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hogben:2003:MCP

[Hog03] Leslie Hogben. Matrix completion problems for pairs of related classes of matrices. Linear Algebra and its Applications, 373(1):13– 29, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hogben:2007:CCP

[Hog07] Leslie Hogben. The copositive completion problem: Unspecified diagonal entries. Linear Algebra and its Applications, 420(1):160–162, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hogben:2008:ORM

[Hog08]  Leslie Hogben.  Orthog-

onal representations, minimum rank, and graph complements. Linear Algebra and its Applications, 428(11–12): 2560–2568, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Holtz:2000:ADM

[Hol00] Olga Holtz. Applications of the duality method to generalizations of the Jordan canonical form. Linear Algebra and its Applications, 310(1–3):11–17, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/127/25/28/article.
pdf.

Holladay:2002:DME

[Hol02] Kenneth W. Holladay. The distance matrix eigensystem of an equally spaced row of points. Linear Algebra and its Applications, 347 (1–3):17–58, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
29/abstract.html.

Holtz:2003:HBR

[Hol03] Olga Holtz. Hermite–Biehler, Routh–Hurwitz, and total positivity. Linear Algebra and its Applications, 372(1): 105–110, October 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Holtz:2005:IEP

[Hol05] Olga Holtz. The inverse eigenvalue problem for symmetric anti-bidiagonal matrices. Linear Algebra and its Applications, 408(1):268– 274, October 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Holladay:2008:CPD

[Hol08] Kenneth W. Holladay. Characteristic polynomials of distance matrices of one dimensional sets. Linear Algebra and its Applications, 428(11– 12):2614–2638, June 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hong:2002:FLM

[Hon02] Shaofang Hong. On the factorization of LCM matrices on gcd-closed sets. Linear Algebra and its Applications, 345(1–3):225–233, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/175/27/
                                        37/abstract.html.

Hong:2006:NMA

[Hon06] Shaofang Hong. Nonsingularity of matrices associated with classes of arithmetical functions on LCMclosed sets. Linear Algebra and its Applications, 416(1): 124–134, July 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hong:2008:DPP

[Hon08] Shaofang Hong. Divisibility properties of power GCD matrices and power LCM matrices. Linear Algebra and its Applications, 428(4): 1001–1008, February 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hoover:2007:SCS

[Hoo07] Melissa Meehan Hoover. Symplectic commutator subgroups. Linear Algebra and its Applications, 423(2–3): 305–323, June 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hanzon:2006:BRD

[HOP06] Bernard Hanzon, Martine Olivi, and Ralf L. M. Peeters. Balanced realizations of discrete-time stable all-pass systems and the tangential Schur algorithm. Linear Algebra and its Applications, 418(2–3):793–820, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Horn:2002:NMD

[Hor02] Roger A. Horn. Normal matrices with a dominant eigenvalue and an eigenvector with no zero entries. Linear Algebra and its Applications, 357 (1–3):35–44, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/203/25/
                                        28/abstract.html.

Horvath:2004:PMV

[Hor04] Zoltán Horváth. On the positivity of matrix-vector products. Linear Algebra and its Applications, 393(1):253– 258, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Horton:2007:IZF

[Hor07] Matthew D. Horton. Ihara zeta functions of digraphs. Linear Algebra and its Applications, 425(1):130–142, August 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2003:EDT

[Hou03] Xiang-Dong Hou. Elementary divisors of tensor products and p-ranks of binomial matrices. Linear Algebra and its Applications, 374(1):255– 274, November 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hebisch:2001:UBQ

[HOZ01] W. Hebisch, R. Olkiewicz, and B. Zegarliński. On upper bound for the quantum entropy. Linear Algebra and its Applications, 329(1–3): 89–96, May 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/154/25/30/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/154/
25/30/article.pdf.

Hartwig:2000:CAM

[HP00] Robert E. Hartwig and K. M. Prasad. Creation and annihilation in matrix theory. Linear Algebra and its Applications, 305(1–3): 47–65, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/120/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/120/25/29/article.
pdf.

Hwang:2001:DSM

[HP01a] Suk-Geun Hwang and Sung-Soo Pyo. Doubly stochastic matrices whose powers eventually stop. Linear Algebra and its Applications, 330 (1–3):25–30, June 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/155/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/155/25/27/article.
pdf.

Hwang:2001:MMV

[HP01b] Suk-Geun Hwang and Sung-Soo Pyo. Matrix majorization via vector majorization. Linear Algebra and its Applications, 332–334 (1):15–21, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/27/article.
pdf.

Haefner:2002:ATO

[HP02] Jeremy Haefner and Christopher J. Pappacena. Automorphisms of tiled orders. Linear Algebra and its Applications, 347(1–3):275–282, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
43/abstract.html.

Hochstenbach:2003:BEC

[HP03a] Michiel E. Hochstenbach and Bor Plestenjak. Backward error, condition numbers, and pseudospectra for the multiparameter eigenvalue problem. Linear Algebra and its Applications, 375(1):63– 81, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Horn:2003:TAT

[HP03b] Roger A. Horn and Gregory G. Piepmeyer. Two applications of the theory of primary matrix functions. Linear Algebra and its Applications, 361(C):99–106, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
33/abstract.html.

Holmes:2004:OFE

[HP04a] Roderick B. Holmes and Vern I. Paulsen. Optimal frames for erasures. Linear Algebra and its Applications, 377(1):31–51, January 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hood:2004:SFP

[HP04b] Jeffrey Hood and David Perkinson. Some facets of the polytope of even permutation matrices. Linear Algebra and its Applications, 381(1): 237–244, April 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hwang:2004:IEP

[HP04c] Suk-Geun Hwang and Sung-Soo Pyo. The inverse eigenvalue problem for symmetric doubly stochastic matrices. Linear Algebra and its Applications, 379(1):77– 83, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See comments [Fan10, XLG^{+}13].

Hinrichsen:2007:GBP

[HP07] Diederich Hinrichsen and Anthony J. Pritchard. Gershgorin– Brualdi perturbations and Riccati equations. Linear Algebra and its Applications, 425(2–3):374–403, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hiai:2009:RMP

[HP09] Fumio Hiai and Dénes Petz. Riemannian metrics on positive definite matrices related to means. Linear Algebra and its Applications, 430(11–12): 3105–3130, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Havlicek:2000:LGI

[HPP00] Miloslav Havlíček, Jiří Patera, and Edita Pelantová. On Lie gradings III. Gradings of the real forms of classical Lie algebras. Linear Algebra and its Applications, 314(1–3):1–47, July 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/131/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/131/25/25/article.
pdf.

Hartwig:2005:DTM

[HPP05] R. E. Hartwig, P. Patrício, and R. Puystjens. Diagonalizing triangular matrices via orthogonal Pierce decompositions. Linear Algebra and its Applications, 401(1):381– 391, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Peng:2007:BSM

[hPyHZ07] Zhuo hua Peng, Xi yan Hu, and Lei Zhang. The bisymmetric solutions of the matrix equation A_{1}X_{1}B_{1} + A_{2}X_{2}B_{2} + ⋯ + A_{l}X_{l}B_{l} = C and its optimal approximation. Linear Algebra and its Applications, 426(2–3):583– 595, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2008:AMD

[HQ08] Jinchuan Hou and Xiaofei Qi. Additive maps derivable at some points on J - subspace lattice algebras. Linear Algebra and its Applications, 429(8–9):1851–1863, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2009:GJD

[HQ09] Jinchuan Hou and Xiaofei Qi. Generalized Jordan derivation on nest algebras. Linear Algebra and its Applications, 430(5–6):1479–1485, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Heinig:2001:CSC

[HR01] Georg Heinig and Karla Rost. Centro-symmetric and centro-skewsymmetric Toeplitz matrices and Bezoutians. Linear Algebra and its Applications, 343–344 (1):195–209, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/174/27/
                                        35/abstract.html.

Heinig:2003:CCS

[HR03a] Georg Heinig and Karla Rost. Centrosymmetric and centro-skewsymmetric Toeplitz-plus-Hankel matrices and Bezoutians. Linear Algebra and its Applications, 366(C):257–281, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
40/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Hille:2003:VTR

[HR03b] Lutz Hille and Gerhard Röhrle. Variation on a theme of Richardson. Linear Algebra and its Applications, 365 (C):239–246, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
32/abstract.html.

Heinig:2004:SAH

[HR04] Georg Heinig and Karla Rost. Split algorithms for hermitian Toeplitz matrices with arbitrary rank profile. Linear Algebra and its Applications, 392(1):235–253, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hasani:2007:LPR

[HR07] A. M. Hasani and M. Radjabalipour. On linear preservers of (right) matrix majorization. Linear Algebra and its Applications, 423(2– 3):255–261, June 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2001:LEQ

[HS01] Liping Huang and Wasin So. On left eigenvalues of a quaternionic matrix. Linear Algebra and its Applications, 323(1–3):105–116, January 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Holtz:2002:OPG

[HS02] Olga Holtz and Hans Schneider. Open problems on GKKmatrices. Linear Algebra and its Applications, 345(1– 3):263–267, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/175/27/
                                        41/abstract.html.

Hochstenbach:2003:TSA

[HS03] Michiel E. Hochstenbach and Gerard L. G. Sleijpen. Twosided and alternating Jacobi– Davidson. Linear Algebra and its Applications, 358(1– 3):145–172, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/

gej-ng/10/30/19/207/25/
34/abstract.html.

Horn:2004:CSM

[HS04] Roger A. Horn and Vladimir V. Sergeichuk. Congruences of a square matrix and its transpose. Linear Algebra and its Applications, 389(1):347–353, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Horn:2006:CFC

[HS06a] Roger A. Horn and Vladimir V. Sergeichuk. Canonical forms for complex matrix congruence and ^{⋆} congruence. Linear Algebra and its Applications, 416(2–3):1010–1032, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Horn:2006:RAM

[HS06b] Roger A. Horn and Vladimir V. Sergeichuk. A regularization algorithm for matrices of bilinear and sesquilinear forms. Linear Algebra and its Applications, 412(2–3):380– 395, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hnetynkova:2007:LTC

[HS07a] Iveta Hnětynková and Zdeněk Strakoš. Lanczos tridiagonalization and core problems. Linear Algebra and its Applications, 421(2–3):243– 251, March 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special Issue in honor of Miroslav Fiedler.

Holbrook:2007:MZA

[HS07b] John Holbrook and Jean-Pierre Schoch. Moving zeros among matrices. Linear Algebra and its Applications, 424 (1):83–95, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Horn:2008:CMB

[HS08] Roger A. Horn and Vladimir V. Sergeichuk. Canonical matrices of bilinear and sesquilinear forms. Linear Algebra and its Applications, 428(1):193– 223, January 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huo:2009:MPA

[HS09] Xiaoming Huo and Andrew K. Smith. Matrix perturbation analysis of local tangent space alignment. Linear Algebra and its Applications, 430(2–3):732–746, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2005:GM

[HSL05] Ting-Zhu Huang, Shu-Qian Shen, and Hou-Biao Li. On generalized H-matrices. Linear Algebra and its Applications, 396(1):81–90, February 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hutchinson:2008:SMM

[HST08] Ryan Hutchinson, Roxana Smarandache, and Jochen Trumpf. On superregular matrices and MDP convolutional codes. Linear Algebra and its Applications, 428(11– 12):2585–2596, June 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hochstenbach:2008:EIR

[HSZ08] Michiel E. Hochstenbach, David A. Singer, and Paul F. Zachlin. Eigenvalue inclusion regions from inverses of shifted matrices. Linear Algebra and its Applications, 429(10):2481–2496, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hajdu:2001:ADT

[HT01] Lajos Hajdu and Robert Tijdeman. An algorithm for discrete tomography. Linear Algebra and its Applications, 339(1–3):147–169, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/169/27/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/169/27/36/article.
                                        pdf.

Higham:2002:MPP

[HT02] Nicholas J. Higham and Françoise Tisseur. More on pseudospectra for polynomial eigenvalue problems and applications in control theory. Linear Algebra and its Applications, 351–352:435– 453, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Higham:2003:BEM

[HT03] Nicholas J. Higham and Françoise Tisseur. Bounds for eigenvalues of matrix polynomials. Linear Algebra and its Applications, 358(1– 3):5–22, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/207/25/
                                        27/abstract.html.

Hou:2005:NHT

[HT05a] Yaoping Hou and Feng Tian. A note on Hoffman-type identities of graphs. Linear Algebra and its Applications, 402 (1):143–149, June 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2005:QIR

[HT05b] Huajun Huang and Tin-Yau Tam. On the QR iterations of real matrices. Linear Algebra and its Applications, 408 (1):161–176, October 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hansen:2007:DAM

[HT07a] Frank Hansen and Jun Tomiyama. Differential analysis of matrix convex functions. Linear Algebra and its Applications, 420(1):102– 116, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2007:ABQ

[HT07b] Huajun Huang and Tin-Yau Tam. An asymptotic behavior of QR decomposition. Linear Algebra and its Applications, 424(1):96– 107, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadjidimos:2008:PEC

[HT08a] A. Hadjidimos and M. Tzoumas. The principle of extrapolation and the Cayley Transform. Linear Algebra and its Applications, 429(10):2465– 2480, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Heydari:2008:CPS

[HT08b] Abbas Heydari and Bijan Taeri. On the characteristic polynomial of a special class of graphs and spectra of balanced trees. Linear Algebra and its Applications, 429(7): 1744–1757, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadjidimos:2009:NEM

[HT09a] A. Hadjidimos and M. Tzoumas. Nonstationary Extrapolated Modulus Algorithms for the solution of the Linear Complementarity Problem. Linear Algebra and its Applications, 431(1–2):197–210, July 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hadjidimos:2009:OCE

[HT09b] A. Hadjidimos and M. Tzoumas. On the optimal complex extrapolation of the complex Cayley transform. Linear Algebra and its Applications, 430(2–3):619–632, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Haynes:2007:PPI

[HTK07] Ronald D. Haynes, Manfred R. Trummer, and Shannon C. Kennedy. Persistently positive inverses of perturbed M-matrices. Linear Algebra and its Applications, 422(2– 3):742–754, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Higham:2002:DDH

[HTV02] Nicholas J. Higham, Françoise Tisseur, and Paul M. Van Dooren. Detecting a definite Hermitian pair and a hyperbolic or elliptic quadratic eigenvalue problem, and associated nearness problems. Linear Algebra and its Applications, 351–352:455–474, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2000:ESS

[Hua00a]  Liping Huang.  The ex-

plicit solutions and solvability of linear matrix equations. Linear Algebra and its Applications, 311(1–3): 195–199, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/128/25/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/128/25/39/article.
pdf.

Huang:2000:TQA

[Hua00b] Liping Huang. On two questions about quaternion matrices. Linear Algebra and its Applications, 318(1–3): 79–86, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/137/24/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/137/24/29/article.
                                        pdf.

Huang:2008:SIS

[Hua08a] Rong Huang. Sign idempotent sign patterns similar to nonnegative sign patterns. Linear Algebra and its Applications, 428(11–12):2524– 2535, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2008:SIH

[Hua08b] Rong Huang. Some inequalities for the Hadamard product and the Fan product of matrices. Linear Algebra and its Applications, 428 (7):1551–1559, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2009:SPS

[Hua09]  Rong Huang.  Sign k-

potent sign patterns and ray k-potent ray patterns that allow k-potence. Linear Algebra and its Applications, 430(8–9):2149–2156, April 15, 2009. CODEN LAAPAW. ISSN 0024-

Huhtanen:2005:ANI

[Huh05] Marko Huhtanen. Aspects of nonnormality for iterative methods. Linear Algebra and its Applications, 394(1):119– 144, January 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huhtanen:2006:RLK

[Huh06] Marko Huhtanen. Real linear Kronecker product operations. Linear Algebra and its Applications, 418(1):347– 361, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huhtanen:2007:HRY

[Huh07] Marko Huhtanen. How real is your matrix? Linear Algebra and its Applications, 424(1): 304–319, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hunter:2005:SDM

[Hun05] Jeffrey J. Hunter. Stationary distributions and mean first passage times of perturbed Markov chains. Linear Algebra and its Applications, 410(1):217–243, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hunter:2006:MTA

[Hun06] Jeffrey J. Hunter. Mixing times with applications to perturbed Markov chains. Linear Algebra and its Applications, 417(1):108–123, August 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hunter:2008:VFP

[Hun08] Jeffrey J. Hunter. Variances of first passage times in a Markov chain with applications to mixing times. Linear Algebra and its Applications, 429(5–6):1135–1162, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hunter:2009:CMT

[Hun09] Jeffrey J. Hunter. Coupling and mixing times in a Markov chain. Linear Algebra and its Applications, 430(10): 2607–2621, May 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hendrickx:2001:FDS

[HV01] Jef Hendrickx and Marc Van Barel. Fast direct solution methods for symmetric banded Toeplitz systems, based on the sine transform. Linear Algebra and its Applications, 343–344(1): 211–232, March 1, 2001.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
36/abstract.html.

Hachez:2003:EHQ

[HV03] Y. Hachez and P. Van Dooren. Elliptic and hyperbolic quadratic eigenvalue problems and associated distance problems. Linear Algebra and its Applications, 371(1):31–44, September 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ho:2008:NNM

[HV08] Ngoc-Diep Ho and Paul Van Dooren. Non-negative matrix factorization with fixed row and column sums. Linear Algebra and its Applications, 429(5–6):1020–1025, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hogben:2007:FMC

[HvdH07] Leslie Hogben and Hein van der Holst. Forbidden minors for the class of graphs G with ξ(G) ≤ 2. Linear Algebra and its Applications, 423 (1):42–52, May 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huhtanen:2005:RLE

[HvP05] Marko Huhtanen and Jan von Pfaler. The real linear eigenvalue problem in C^{n}. Linear Algebra and its Applications, 394(1):169– 199, January 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hardin:2007:OLP

[HvS07] Hanna M. Härdin and Jan H. van Schuppen. Observers for linear positive systems. Linear Algebra and its Applications, 425(2–3):571–607, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hodgess:2000:TDS

[HW00] Erin M. Hodgess and William W. S. Wei. Temporal disaggregation of stationary bivariate time series. Linear Algebra and its Applications, 321(1–3):175–196, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/140/24/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/140/24/36/article.
                                        pdf.

Hachez:2005:ASS

[HW05a] Yvan Hachez and Hugo J. Woerdeman. Approximating sums of squares with a single square. Linear Algebra and its Applications, 399(1): 187–201, April 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2005:GSH

[HW05b] Li-Ping Huang and Zhe-Xian Wan. Geometry of skew-Hermitian matrices. Linear Algebra and its Applications, 396(1):127–157, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hua:2007:UGG

[HW07a] Hongbo Hua and Maolin Wang. Unicyclic graphs with given number of pendent vertices and minimal energy. Linear Algebra and its Applications, 426(2–3):478–489, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2007:IBE

[HW07b] Ting-Zhu Huang and Lin Wang. Improving bounds for eigenvalues of complex matrices using traces. Linear Algebra and its Applications, 426(2–3):841–854, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2006:SGS

[HWC06] Yaoping Hou, Chingwah Woo, and Pingge Chen. On the sandpile group of the square cycle C_{n}^{2}. Linear Algebra and its Applications, 418(2–3):457–467, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hemasinha:2003:PBL

[HWKS03] Rohan Hemasinha, James R. Weaver, Stephen J. Kirkland, and Jeffrey L. Stuart. Properties of the Brualdi–Li tournament matrix. Linear Algebra and its Applications, 361(C):63–73, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
31/abstract.html.

Hartwig:2001:SAR

[HWW01] Robert E. Hartwig, Guorong Wang, and Yimin Wei. Some additive results on Drazin inverse. Linear Algebra and its Applications, 322(1–3): 207–217, January 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/141/24/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/141/24/35/article.
pdf.

Hu:2008:NBG

[HXL08] Shengbiao Hu, Tan Xuezhong, and Bolian Liu. On the nullity of bicyclic graphs. Linear Algebra and its Applications, 429(7):1387–1391, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hiai:2002:IIU

[HZ02] Fumio Hiai and Xingzhi Zhan. Inequalities involving unitarily invariant norms and operator monotone functions. Linear Algebra and its Applications, 341(1–3): 151–169, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
39/abstract.html.

Hilscher:2003:SDS

[HZ03] Roman Hilscher and Vera Zeidan. Symplectic difference systems: variable stepsize discretization and discrete quadratic functionals. Linear Algebra and its Applications, 367(C):67–104, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
32/abstract.html.

Hiai:2004:SVS

[HZ04a] Fumio Hiai and Xingzhi Zhan. Submultiplicativity vs subadditivity for unitarily invariant norms. Linear Algebra and its Applications, 377(1):155–164, January 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2004:RIL

[HZ04b] Jinchuan Hou and Xiuling Zhang. Ring isomorphisms and linear or additive maps preserving zero products on nest algebras. Linear Algebra and its Applications, 387 (1):343–360, August 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hou:2005:ZPP

[HZ05a] Jinchuan Hou and Liankuo Zhao. Zero-product preserving additive maps on symmetric operator spaces and self-adjoint operator spaces. Linear Algebra and its Applications, 399(1):235–244, April 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2005:DPG

[HZ05b] Jianguo Huang and Jieyong Zhou. A direct proof and a generalization for a Kantorovich type inequality. Linear Algebra and its Applications, 397(1):185– 192, March 1, 2005. CO-

Hilscher:2006:CID

[HZ06] Roman Hilscher and Vera Zeidan. Coupled intervals for discrete symplectic systems. Linear Algebra and its Applications, 419(2–3):750–764, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hu:2002:HSO

[HZM02] Bo Hu, Guisheng Zhai, and Anthony N. Michel. Hybrid static output feedback stabilization of second-order linear time-invariant systems. Linear Algebra and its Applications, 351–352:475–485, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hu:2005:MMI

[HZY05] Yongjian Hu, Xiuping Zhang, and Zhenghong Yang. Mixed mean inequalities for several positive definite matrices. Linear Algebra and its Applications, 395(1):247– 263, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dokovic:2002:CSR

[– DI02] Dragomir Ž. – Doković and Khakim D. Ikramov. On the congruence of square real matrices. Linear Algebra and its Applications, 353(1–3): 149–158, September 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Iyigun:2009:CAD

[IBI09] Cem Iyigun and Adi Ben-Israel. Contour approximation of data: A duality theory. Linear Algebra and its Applications, 430(10):2771– 2780, May 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Iglesias:2001:LIS

[Igl01] Pablo A. Iglesias. Logarithmic integrals and system dynamics: an analogue of Bode’s sensitivity integral for continuous-time, timevarying systems. Linear Algebra and its Applications, 343–344(1):451–471, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
47/abstract.html.

Ivanov:2001:ME

[IHM01] I. G. Ivanov, V. I. Hasanov, and B. V. Minchev. On matrix equations X ± A^{∗}X^{2}A = I. Linear Algebra and its Applications, 326(1–3):27– 44, March 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/148/24/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/148/24/26/article.
pdf.

Ilmonen:2008:EMJ

[IHM08] Pauliina Ilmonen, Pentti Haukkanen, and Jorma K. Merikoski. On eigenvalues of meet and join matrices associated with incidence functions. Linear Algebra and its Applications, 429 (4):859–874, August 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See comment [AKYD16].

Ilic:2009:LCT

[II09] Aleksandar Ilić and Milovan Ilić. Laplacian coefficients of trees with given number of leaves or vertices of degree two. Linear Algebra and its Applications, 431(11):2195– 2202, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ito:2006:ROC

[IK06] Kazufumi Ito and Karl Kunisch. Reduced order control based on approximate inertial manifolds. Linear Algebra and its Applications, 415 (2–3):531–541, June 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ito:2009:CMG

[IK09] Masatoshi Ito and Eizaburo Kamei. A complement to monotonicity of generalized Furuta-type operator functions. Linear Algebra and its Applications, 430(1):544– 546, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ikramov:2001:HSR

[Ikr01] Khakim D. Ikramov. Hamiltonian square roots of skew-Hamiltonian matrices revisited. Linear Algebra and its Applications, 325(1–3): 101–107, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/147/24/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/147/24/31/article.
                                        pdf.

Ikramov:2007:NME

[Ikr07a] Khakim D. Ikramov. A note on the matrix equation XA− AX = X^{p}. Linear Algebra and its Applications, 424 (2–3):390–392, July 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ikramov:2007:CM

[Ikr07b] Khakim D. Ikramov. On condiagonalizable matrices. Linear Algebra and its Applications, 424(2–3):456–465, July 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Inglot:2006:TDD

[IL06] Tadeusz Inglot and Teresa Ledwina. Towards data driven selection of a penalty function for data driven Neyman tests. Linear Algebra and its Applications, 417 (1):124–133, August 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ilisevic:2008:GBP

[Ili08] Dijana Ilišević. Generalized bicircular projections via the operator equation αX^{2}AY^{2}+ βXAY + A = 0. Linear Algebra and its Applications, 429(8–9):2025–2029, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ilic:2009:EUC

[Ili09a] Aleksandar Ilić. The energy of unitary Cayley graphs. Linear Algebra and its Applications, 431(10):1881–1889, October 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ilic:2009:OTL

[Ili09b] Aleksandar Ilić. On the ordering of trees by the Laplacian coefficients. Linear Algebra and its Applications, 431(11):2203–2212, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ipsen:2009:PIC

[IMS^{+}09] Ilse Ipsen, Julio Moro, Peter Semrl, Jia-Yu Shao, Pei Yuan Wu, and Richard Brualdi. Preface to the 14th ILAS Conference Proceedings Shanghai 2007. Linear Algebra and its Applications, 430(5–6):1441, March 1, 2009. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Indulal:2009:SBD

[Ind09] G. Indulal. Sharp bounds on the distance spectral radius and the distance energy of graphs. Linear Algebra and its Applications, 430(1):106– 113, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ito:2003:GNR

[INOY03] Masatoshi Ito, Hiroshi Nakazato, Kazuyoshi Okubo, and Takeaki Yamazaki. On generalized numerical range of the Aluthge transformation. Linear Algebra and its Applications, 370(1):147–161, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ionescu:2001:FBA

[IO01] Vlad Ionescu and Cristian Oarǎ. The four-block Adamjan-Arov-Krein problem for discrete-time systems. Linear Algebra and its Applications, 328(1–3): 95–119, May 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/153/25/28/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/153/
25/28/article.pdf.

Ibarrola:2007:CGP

[IPP07] P. Ibarrola and A. Pérez-Palomares. A characterization of a Gaussian process in terms of sufficient estimators. Linear Algebra and its Applications, 426(2–3):280–289, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ibarrola:2009:DLM

[IPP09] P. Ibarrola and A. Pérez-Palomares. A decomposition of a linear model. Linear Algebra and its Applications, 430(4):1017–1024, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ipsen:2000:ARP

[Ips00] Ilse C. F. Ipsen. Absolute and relative perturbation bounds for invariant subspaces of matrices. Linear Algebra and its Applications, 309(1– 3):45–56, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/126/25/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/126/25/28/article.
                                        pdf.

Ipsen:2003:NUA

[Ips03] Ilse C. F. Ipsen. A note on unifying absolute and relative perturbation bounds. Linear Algebra and its Applications, 358(1–3):239–253, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/207/25/
                                        38/abstract.html.

Im:2008:GAC

[IS08] Bokhee Im and Jonathan D. H. Smith. Generic adjoints in comtrans algebras of bilinear spaces. Linear Algebra and its Applications, 428(4): 953–961, February 1, 2008. CODEN LAAPAW. ISSN

Im:2009:TPC

[IS09] Bokhee Im and Jonathan D. H. Smith. Trilinear products and comtrans algebra representations. Linear Algebra and its Applications, 430 (1):17–26, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ishibashi:2004:SOG

[Ish04] Hiroyuki Ishibashi. Structure of the orthogonal group O_{n}(V) over L-rings. Linear Algebra and its Applications, 390(1):357–368, October 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ishibashi:2006:LPA

[Ish06] Hiroyuki Ishibashi. Length problems for automorphisms of modules over local rings. Linear Algebra and its Applications, 418(1):269–276, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ito:2007:IPL

[IT07a] Tatsuro Ito and Paul Terwilliger. q-Inverting pairs of linear transformations and the q-tetrahedron algebra. Linear Algebra and its Applications, 426(2–3):516–532, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ito:2007:TPK

[IT07b] Tatsuro Ito and Paul Terwilliger. Tridiagonal pairs of Krawtchouk type. Linear Algebra and its Applications, 427(2–3):218–233, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ibarrola:2004:BFP

[IV04] Pilar Ibarrola and Ricardo Vélez. On Behrens–Fisher problem for continuous time Gaussian processes. Linear Algebra and its Applications, 389(1):63–76, September 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ivanov:2005:PAS

[Iva05] I. G. Ivanov. Perturbation analysis for solutions of X ± A ∗ X^{−n}A = Q. Linear Algebra and its Applications, 395(1):313–331, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ip:2007:IWD

[IWL07] Wai Cheung Ip, Heung Wong, and Jin Shan Liu. Inverse Wishart distributions based on singular elliptically contoured distribution. Linear Algebra and its Applications, 420(2–3):424–432, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dokovic:2007:UAZ

[– DJ07] Dragomir Ž. – Doković and Charles R. Johnson. Unitarily achievable zero patterns and traces of words in A and A. Linear Algebra and its Applications, 421(1):63– 68, February 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jain:2006:BMA

[Jai06] Sapna Jain. Bursts in mmetric array codes. Linear Algebra and its Applications, 418(1):130–141, October 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jain:2007:CTB

[Jai07] Sapna Jain. Campopianotype bounds in non-Hamming array coding. Linear Algebra and its Applications, 420 (1):135–159, January 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jamison:2007:BPS

[Jam07] James E. Jamison. Bicircular projections on some Banach spaces. Linear Algebra and its Applications, 420(1): 29–33, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jansson:2001:QRB

[Jan01] Christian Jansson. Quasiconvex relaxations based on interval arithmetic. Linear Algebra and its Applications, 324(1–3):27–53, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/144/24/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/144/24/27/article.
pdf.

Jiang:2000:STL

[JB00] Eric P. Jiang and Michael W. Berry. Solving total leastsquares problems in information retrieval. Linear Algebra and its Applications, 316 (1–3):137–156, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/33/article.
pdf.

Jungers:2008:FPR

[JB08] Raphaël M. Jungers and Vincent D. Blondel. On the finiteness property for rational matrices. Linear Algebra and its Applications, 428(10): 2283–2295, May 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2002:PME

[JD02] Charles R. Johnson and António Leal Duarte. On the possible multiplicities of the eigenvalues of a Hermitian matrix whose graph is a tree. Linear Algebra and its Applications, 348 (1–3):7–21, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
28/abstract.html.

Johnson:2003:IEP

[JDS03a] Charles R. Johnson, António Leal Duarte, and Carlos M. Saiago. Inverse eigenvalue problems and lists of multiplicities of eigenvalues for matrices whose graph is a tree: the case of generalized stars and double generalized stars. Linear Algebra and its Applications, 373(1):311–330, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2003:RPM

[JDS^{+}03b] Charles R. Johnson, António Leal Duarte, Carlos M. Saiago, Brian D. Sutton, and Andrew J. Witt. On the relative position of multiple eigenvalues in the spectrum of an Hermitian matrix with a given graph. Linear Algebra and its Applications, 363 (C):147–159, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/212/25/
                                        38/abstract.html.

Johnson:2008:SMM

[JDS08] Charles R. Johnson, António Leal Duarte, and Carlos M. Saiago. The structure of matrices with a maximum multiplicity eigenvalue. Linear Algebra and its Applications, 429(4):875–886, August 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jeffries:2005:BTS

[Jef05] Clark Jeffries. Bipartite and tripartite systems and matrices from genetic control research. Linear Algebra and its Applications, 409(1):70– 78, November 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jelonek:2003:AAI

[Jel03] Zbigniew Jelonek. Affine automorphisms that are isometries. Linear Algebra and its Applications, 364(C):125– 134, May 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
28/abstract.html.

Johnson:2001:GSL

[JF01] Charles R. Johnson and Susana Furtado. A generalization of Sylvester’s law of inertia. Linear Algebra and its Applications, 338(1–3): 287–290, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
48/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/48/article.
pdf.

Jimenez-Gestal:2008:TDF

[JGPI08] Clara Jiménez-Gestal and José M. Pérez-Izquierdo. Ternary derivations of finitedimensional real division algebras. Linear Algebra and its Applications, 428(8–9): 2192–2219, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jarlebring:2009:PTP

[JH09] Elias Jarlebring and Michiel E. Hochstenbach. Polynomial two-parameter eigenvalue problems and matrix pencil methods for stability of delay-differential equations. Linear Algebra and its Applications, 431(3–4):369– 380, July 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jiang:2005:GMD

[JHL05] Yi Jiang, William W. Hager, and Jian Li. The geometric mean decomposition. Linear Algebra and its Applications, 396(1):373–384, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ji:2003:ASP

[Ji03a]  Guoxing Ji.  Asymptotic

similarity-preserving linear maps on B(H). Linear Algebra and its Applications, 368 (C):371–378, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/217/25/
                                        49/abstract.html.

Ji:2003:SPL

[Ji03b]  Guoxing Ji.  Similarity-

preserving linear maps on B(H). Linear Algebra and its Applications, 360(C):249– 257, February 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
42/abstract.html.

Ji:2005:EEG

[Ji05] Jun Ji. Explicit expressions of the generalized inverses and condensed Cramer rules. Linear Algebra and its Applications, 404(1):183– 192, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ji:2009:AJMb

[Ji09] Peisheng Ji. Additivity of Jordan maps on Jordan algebras. Linear Algebra and its Applications, 431(1–2):179– 188, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jiang:2000:MIS

[Jia00a] Jiming Jiang. A matrix inequality and its statistical application. Linear Algebra and its Applications, 307(1– 3):131–144, March 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/124/25/34/article.
                                        pdf.

Jiang:2000:PSO

[Jia00b] Qingtang Jiang. Parametrizations of symmetric orthogonal multifilter banks with different filter lengths. Linear Algebra and its Applications, 311(1–3):79–96, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/128/25/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/128/25/31/article.
                                        pdf.

Jiang:2004:DMB

[Jia04] Jiming Jiang. Dispersion matrix in balanced mixed ANOVA models. Linear Algebra and its Applications, 382(1):211–219, May 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jiang:2005:PES

[Jia05] Er-Xiong Jiang. Perturbation in eigenvalues of a symmetric tridiagonal matrix. Linear Algebra and its Applications, 399(1):91–107, April 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jiang:2009:PTH

[Jia09] Chung-Chou Jiang. On products of two Hermitian operators. Linear Algebra and its Applications, 430(1):553– 557, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jiang:2003:NES

[JK03] Chung-Chou Jiang and Kung-Hwang Kuo. On normal extensions of submatrices. Linear Algebra and its Applications, 370(1):301–314, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jeon:2006:OS

[JK06a] Ho Jeon and Hyoun Kim. On operators satisfying T^{∗}|T^{2}|T ≥ T^{∗}|T |^{2}T. Linear Algebra and its Applications, 418(2–3): 854–862, October 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jin:2006:PCT

[JK06b] Yi Jin and Bahman Kalantari. A procedure of Chvátal for testing feasibility in linear programming and matrix scaling. Linear Algebra and its Applications, 416 (2–3):795–798, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jung:2009:SNM

[JKL09] Changdo Jung, Hyun-Min Kim, and Yongdo Lim. On the solution of the nonlinear matrix equation X^{n} = f(X). Linear Algebra and its Applications, 430(8–9):2042– 2052, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jung:2003:ETC

[JKLP03] Il Bong Jung, Eungil Ko, Chunji Li, and Sang Soo Park. Embry truncated complex moment problem. Linear Algebra and its Applications, 375(1):95–114, December 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2004:EMP

[JKO04] Charles R. Johnson, Brenda Kroschel, and Matjaž Omladič. Eigenvalue multiplicities in principal submatrices. Linear Algebra and its Applications, 390(1):111– 120, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2002:CCH

[JKRT02] N. W. Johnson, R. Kellerhals, J. G. Ratcliffe, and S. T. Tschantz. Commensurability classes of hyperbolic Coxeter groups. Linear Algebra and its Applications, 345(1– 3):119–147, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/175/27/
33/abstract.html.

Jin:2003:STB

[JL03] Xiao-Qing Jin and Siu-Long Lei. Sine transform based preconditioners for solving constant-coefficient firstorder PDEs. Linear Algebra and its Applications, 366 (C):283–294, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
41/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Ji:2009:AJMa

[JL09] Peisheng Ji and Zhongyan Liu. Additivity of Jordan maps on standard Jordan operator algebras. Linear Algebra and its Applications, 430 (1):335–343, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jing:2003:AMP

[JLL03] Wu Jing, Pengtong Li, and Shijie Lu. Additive mappings that preserve rank one nilpotent operators. Linear Algebra and its Applications, 367 (C):213–224, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/216/25/
                                        40/abstract.html.

Johnson:2005:PMS

[JLMC05] Charles R. Johnson, Stefan Leichenauer, Peter McNamara, and Roberto Costas. Principal minor sums of (A + tB)^{m}. Linear Algebra and its Applications, 411(1):386– 389, December 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jacob:2002:CAP

[JLMP02] Birgit Jacob, Juliette Leblond, Jean-Paul Marmorat, and Jonathan R. Partington. A constrained approximation problem arising in parameter identification. Linear Algebra and its Applications, 351–352:487–500, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2008:STS

[JLN08] Charles R. Johnson, Terry D. Lenker, and Sivaram K. Narayan. Schur-type stability properties and complex diagonal scaling. Linear Algebra and its Applications, 429(10):2497–2520, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2001:PLS

[JLY01] Charles R. Johnson, Suzanne A. Lewis, and Donald Y. Yau. Possible line sums for a qualitative matrix. Linear Algebra and its Applications, 327(1– 3):53–60, April 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/29/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/152/
25/29/article.pdf.

Jung:2009:NCG

[JLY09] Changdo Jung, Hosoo Lee, and Takeaki Yamazaki. On a new construction of geometric mean of n-operators. Linear Algebra and its Applications, 431(9):1477–1488, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jokar:2009:SSU

[JM09] Sadegh Jokar and Volker Mehrmann. Sparse solutions to underdetermined Kronecker product systems. Linear Algebra and its Applications, 431(12):2437–2447, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jiang:2008:UMD

[JMN08] Yunjiang Jiang, Lon H. Mitchell, and Sivaram K.

Narayan. Unitary matrix digraphs and minimum semidefinite rank. Linear Algebra and its Applications, 428(7):1685–1695, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2009:MSS

[JMP09] C. R. Johnson, C. Marijuán, and M. Pisonero. Matrices and spectra satisfying the Newton inequalities. Linear Algebra and its Applications, 430(11–12):3030– 3046, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jbilou:2004:SSC

[JMT04] K. Jbilou, A. Messaoudi, and K. Tabaâ. Some Schur complement identities and applications to matrix extrapolation methods. Linear Algebra and its Applications, 392(1): 195–210, November 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2005:DSE

[JMT05] Charles R. Johnson, Robert D. Masson, and Michael W. Trosset. On the diagonal scaling of Euclidean distance matrices to doubly stochastic matrices. Linear Algebra and its Applications, 397(1): 253–264, March 1, 2005. CO-

Jain:2006:LBC

[JN06] Sapna Jain and Ki-Bong Nam. Lower bounds for codes correcting moderatedensity bursts of fixed length with Lee weight consideration. Linear Algebra and its Applications, 418(1):122– 129, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jain:2005:SPC

[JNL05] Sapna Jain, Ki-Bong Nam, and Ki-Suk Lee. On some perfect codes with respect to Lee metric. Linear Algebra and its Applications, 405 (1):104–120, August 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2002:UMS

[JO02] Charles R. Johnson and Kazuyoshi Okubo. Uniqueness of matrix square roots under a numerical range condition. Linear Algebra and its Applications, 341(1–3): 195–199, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
42/abstract.html.

Johnson:2004:STP

[JO04] C. R. Johnson and D. D. Olesky. Sums of totally positive matrices. Linear Algebra and its Applications, 392(1): 1–9, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2005:RSI

[JO05] C. R. Johnson and D. D. Olesky. Rectangular submatrices of inverse M-matrices and the decomposition of a positive matrix as a sum. Linear Algebra and its Applications, 409(1):87– 99, November 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jocic:2009:INB

[Joc09a] Danko R. Jocić. Interpolation norms between row and column spaces and the norm problem for elementary operators. Linear Algebra and its Applications, 430(11–12): 2961–2974, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jocic:2009:MEO

[Joc09b] Danko R. Jocić. Multipliers of elementary operators and comparison of row and column space Schatten p norms. Linear Algebra and its Applications, 431(11):2062– 2070, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2005:CES

[Joh05] Warren P. Johnson. Confluent q-extensions of some classical determinants. Linear Algebra and its Applications, 411(1):281–294, December 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jondrup:2009:FNC

[Jøn09]  Søren Jøndrup.  Factor-

ization in noncommutative curves. Linear Algebra and its Applications, 431(5–7): 985–990, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2001:UMS

[JOR01] Charles R. Johnson, Kazuyoshi Okubo, and Robert Reams. Uniqueness of matrix square roots and an application. Linear Algebra and its Applications, 323(1–3):51–60, January 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jorgensen:2005:RAM

[Jør05] L. K. Jørgensen. Rank of adjacency matrices of directed (strongly) regular graphs.

Linear Algebra and its Applications, 407(1):233–241, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jaume:2004:RCS

[JP04] Daniel A. Jaume and Rubén Puente. Representability of convex sets by analytical linear inequality systems. Linear Algebra and its Applications, 380(1):135– 150, March 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jungers:2008:EAD

[JPB08] Raphaël M. Jungers, Vladimir Protasov, and Vincent D. Blondel. Efficient algorithms for deciding the type of growth of products of integer matrices. Linear Algebra and its Applications, 428(10): 2296–2311, May 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2000:SRS

[JR00] Charles R. Johnson and Robert Reams. Semidefiniteness without real symmetry. Linear Algebra and its Applications, 306(1–3): 203–209, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

gej-ng/10/30/19/121/25/
41/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/121/25/41/article.
pdf.

Johnson:2005:STC

[JR05] Charles R. Johnson and Robert Reams. Spectral theory of copositive matrices. Linear Algebra and its Applications, 395(1):275– 281, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jbilou:2006:PML

[JR06] K. Jbilou and A. J. Riquet. Projection methods for large Lyapunov matrix equations. Linear Algebra and its Applications, 415(2– 3):344–358, June 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jagels:2009:EKS

[JR09] Carl Jagels and Lothar Reichel. The extended Krylov subspace method and orthogonal Laurent polynomials. Linear Algebra and its Applications, 431(3–4):441– 458, July 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2000:PDC

[JS00a] Charles R. Johnson and Ronald L. Smith. The positive definite completion problem relative to a subspace. Linear Algebra and its Applications, 307(1–3):1– 14, March 1, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/25/article.
pdf.

Johnson:2000:PAG

[JS00b] Charles R. Johnson and David P. Stanford. Patterns that allow given row and column sums. Linear Algebra and its Applications, 311 (1–3):97–105, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/128/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/128/25/32/article.
pdf.

Johnson:2001:APM

[JS01a] Charles R. Johnson and Ronald L. Smith. Almost principal minors of inverse M-matrices. Linear Algebra and its Applications, 337 (1–3):253–265, November 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

gej-ng/10/30/19/167/27/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/167/27/38/article.
pdf.

Johnson:2001:LIP

[JS01b] Charles R. Johnson and Ronald L. Smith. Linear interpolation problems for matrix classes and a transformational characterization of M-matrices. Linear Algebra and its Applications, 330 (1–3):43–48, June 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/155/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/155/25/29/article.
pdf.

Jensen:2004:MEC

[JS04] D. R. Jensen and S. S. Srinivasan. Matrix equivalence classes with applications. Linear Algebra and its Applications, 388(1):249– 260, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jayaraman:2007:WMV

[JS07a] Sachindranath Jayaraman and K. C. Sivakumar. Weak monotonicity of A versus 1inverses nonnegative on the range space of A. Linear Algebra and its Applications, 427(2–3):171–175, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Johnson:2007:CPP

[JS07b] Charles R. Johnson and Ronald L. Smith. Corrigendum to: “Positive, path product, and inverse M-matrices” [Linear Algebra Appl. 421 (2007) 328– 337]. Linear Algebra and its Applications, 423(2–3): 519, June 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See [JS07c].

Johnson:2007:PPP

[JS07c] Charles R. Johnson and Ronald L. Smith. Positive, path product, and inverse M-matrices. Linear Algebra and its Applications, 421(2– 3):328–337, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See corrigendum [JS07b].

Jenek:2003:CSM

[JSU03] Sławomir Jenek, Tomasz Szulc, and Frank Uhlig. On the class of D_{k}-symmetrizable matrices. Linear Algebra and its Applications, 363 (C):133–145, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/

gej-ng/10/30/19/212/25/
37/abstract.html.

Johnson:2005:OGS

[JSWT05] Charles R. Johnson, Tomasz Szulc, and Dominika Wojtera-Tyrakowska. Optimal Gersgorinstyle estimation of extremal singular values. Linear Algebra and its Applications, 402 (1):46–60, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jain:2004:NMA

[JT04a] S. K. Jain and John Tynan. Nonnegative matrices A with AA^{#} ≥ 0. Linear Algebra and its Applications, 379(1): 381–394, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jordan:2004:TPC

[JT04b] Cristina Jordán and Juan R. Torregrosa. The totally positive completion problem. Linear Algebra and its Applications, 393(1):259–274, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jordan:2009:CDT

[JTeG09] Cristina Jordán, Juan-R. Torregrosa, and Ramadán el Ghamry. The completable digraphs for the totally nonnegative completion problem. Linear Algebra and its Applications, 430(5–6):1675– 1690, March 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jordan:2001:GCC

[JTU01] C. Jordán, J. R. Torregrosa, and A. M. Urbano. Graphs and controllability completion problems. Linear Algebra and its Applications, 332–334(1):355–370, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
42/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/42/article.
pdf.

Jordan:2003:CPM

[JTU03] C. Jordán, J. R. Torregrosa, and A. M. Urbano. Completions of partial P -matrices with acyclic or non-acyclic associated graph. Linear Algebra and its Applications, 368(C):25–51, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
27/abstract.html.

Jiang:2003:SME

[JW03] Tongsong Jiang and Musheng Wei. On solutions of the matrix equations X −AXB = C and X − AX̄B = C. Linear Algebra and its Applications, 367(C):225–233, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
41/abstract.html.

Ji:2005:LTD

[JW05] Peisheng Ji and Lin Wang. Lie triple derivations of TUHF algebras. Linear Algebra and its Applications, 403(1):399–408, July 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ji:2006:SPL

[JW06] Guoxing Ji and Baowei Wu. Similarity preserving linear maps on upper triangular matrix algebras. Linear Algebra and its Applications, 414 (1):278–287, April 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Jin:2008:SSE

[JW08] Xiao-Qing Jin and Yi-Min Wei. A survey and some extensions of T. Chan’s preconditioner. Linear Algebra and its Applications, 428(2–3): 403–412, January 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ji:2007:ECC

[JY07a] Jun Ji and Bo Yang. Eigenvalue comparisons for a class of boundary value problems of second order difference equations. Linear Algebra and its Applications, 420(1): 218–227, January 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ji:2007:ECS

[JY07b] Jun Ji and Bo Yang. Eigenvalue comparisons for second order difference equations with Neumann boundary conditions. Linear Algebra and its Applications, 425 (1):171–183, August 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2009:ADC

[jYxL09] Shang jun Yang and Xiao xin Li. Algorithms for determining the copositivity of a given symmetric matrix. Linear Algebra and its Applications, 430(2–3):609–618, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2001:GVD

[jYzWbZ01] Shang jun Yang, Hua zhang Wu, and Quan bing Zhang. Generalization of Vandermonde determinants. Linear Algebra and its Applications, 336(1–3):201–204, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
40/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/166/27/40/article.
pdf.

Knyazev:2006:PRQ

[KA06] Andrew V. Knyazev and Merico E. Argentati. On proximity of Rayleigh quotients for different vectors and Ritz values generated by different trial subspaces. Linear Algebra and its Applications, 415(1):82– 95, May 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kahoui:2004:SLN

[Kah04] M’hammed El Kahoui. Subresultants and locally nilpotent derivations. Linear Algebra and its Applications, 380 (1):253–261, March 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kahan:2006:TSS

[Kah06]  W. Kahan.  Is there a

small skew Cayley transform with zero diagonal? Linear Algebra and its Applications, 417(2–3):335–341, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kakimura:2008:SSL

[Kak08] Naonori Kakimura. Signsolvable linear complementarity problems. Linear Algebra and its Applications, 429 (2–3):606–616, July 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kalantari:2003:SPM

[Kal03a] Bahman Kalantari. Semidefinite programming and matrix scaling over the semidefinite cone. Linear Algebra and its Applications, 375(1):221– 243, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kalauch:2003:MPO

[Kal03b] Anke Kalauch. On maximum principles for M-operators. Linear Algebra and its Applications, 371(1):209–224, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kalra:2006:CEC

[Kal06]  Deepti Kalra.  Complex

equiangular cyclic frames and erasures. Linear Algebra and its Applications, 419(2–3): 373–399, December 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kaplan:2000:TCM

[Kap00] Wilfred Kaplan. A test for copositive matrices. Linear Algebra and its Applications, 313(1–3):203–206, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/130/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/130/25/38/article.
pdf.

Kaplan:2001:CP

[Kap01] Wilfred Kaplan. A copositivity probe. Linear Algebra and its Applications, 337(1–3): 237–251, November 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/167/27/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/167/27/37/article.
pdf.

Konstantinov:2003:PBC

[KAP^{+}03] Mihail Konstantinov, Vera Angelova, Petko Petkov, Dawei Gu, and Vassilios Tsachouridis. Perturbation bounds for coupled matrix Riccati equations. Linear Algebra and its Applications, 359(1–3):197–218, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/208/25/
                                        41/abstract.html.

Karampetakis:2004:SSD

[Kar04] N. P. Karampetakis. On the solution space of discrete time AR-representations over a finite time horizon. Linear Algebra and its Applications, 382(1):83–116, May 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Karaa:2005:LRC

[Kar05a] Samir Karaa. A limit result concerning the QR factorization of banded Toeplitz matrices. Linear Algebra and its Applications, 405(1):41– 44, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Karzanov:2005:CCT

[Kar05b] Alexander V. Karzanov. Concave cocirculations in a triangular grid. Linear Algebra and its Applications, 400 (1):67–89, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Karzanov:2007:MOV

[Kar07] Alexander V. Karzanov. On 0, ±1 matrices, odd vectors, and bisubmodular polyhedra. Linear Algebra and its Applications, 422(1):17– 21, April 1, 2007. CO-

Karaev:2009:NRS

[Kar09] M. T. Karaev. On the numerical radius of some C_{00}- contractions. Linear Algebra and its Applications, 431(5– 7):818–822, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kattchee:2004:EKD

[Kat04] K. Kattchee. Elasticities of Krull domains with finite divisor class group. Linear Algebra and its Applications, 384(1):171–185, June 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Keel:2002:RCP

[KB02a] L. H. Keel and S. P. Bhattacharyya. Root counting, phase unwrapping, stability and stabilization of discrete time systems. Linear Algebra and its Applications, 351–352:501–518, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kokol-Bukovsek:2002:MMS

[KB02b] Damjana Kokol-Bukovšek. More on matrix semigroup homomorphisms. Linear Algebra and its Applications, 346(1–3):73–95, May 1, 2002.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/189/27/
31/abstract.html.

Kovacec:2007:CCD

[KBdP07] Alexander Kovacec, Natália Bebiano, and João da Providência. On the corners of certain determinantal ranges. Linear Algebra and its Applications, 426(1):96–108, October 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kurowicka:2003:PPD

[KC03] Dorota Kurowicka and Roger Cooke. A parameterization of positive definite matrices in terms of partial correlation vines. Linear Algebra and its Applications, 372(1):225– 251, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kurowicka:2006:CPP

[KC06] D. Kurowicka and R. M. Cooke. Completion problem with partial correlation vines. Linear Algebra and its Applications, 418(1):188– 200, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Koev:2007:AEC

[KD07] Plamen Koev and Froilán Dopico. Accurate eigenvalues of certain sign regular matrices. Linear Algebra and its Applications, 424(2– 3):435–447, July 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Koliha:2007:MPI

[KDC07] J. J. Koliha, Dragan Djordjević, and Dragana Cvetković. Moore–Penrose inverse in rings with involution. Linear Algebra and its Applications, 426(2–3):371–381, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Keller:2008:ME

[Kel08] Joseph B. Keller. Multiple eigenvalues. Linear Algebra and its Applications, 429(8–9):2209–2220, October 16, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kestenband:2000:CIC

[Kes00] Barbu C. Kestenband. The correlations with identity companion automorphism, of finite Desarguesian planes. Linear Algebra and its Applications, 304(1–3):1–31, January 1, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/118/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/118/25/25/article.
pdf.

Keska:2001:VHT

[Kes01] Sergiusz Keska. A variant of the Hausdorff theorem for multi-index matrices II. Linear Algebra and its Applications, 327(1–3): 17–26, April 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/26/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/152/
25/26/article.pdf.

Kestenband:2007:CFD

[Kes07] Barbu C. Kestenband. The correlations of finite Desarguesian planes of square order defined by diagonal matrices. Linear Algebra and its Applications, 423(2–3):366– 385, June 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kozeratska:2004:PCA

[KFGK04] L. Kozeratska, J. F. Forbes, R. G. Goebel, and J. V. Kresta. Perturbed cones for analysis of uncertain multicriteria optimization problems. Linear Algebra and its Applications, 378(1):203– 229, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kovacec:2009:HPC

[KG09] Alexander Kovačec and Maria Celeste Gouveia. The Hankel pencil conjecture. Linear Algebra and its Applications, 431(9):1509–1525, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2001:AEP

[KH01a] Sang Dong Kim and Yong HunLee. Analysis on eigenvalues for preconditioning cubic spline collocation method of elliptic equations. Linear Algebra and its Applications, 327(1–3):1–15, April 15, 2001. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/25/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/152/
25/25/article.pdf.

Korkee:2001:BDM

[KH01b] Ismo Korkee and Pentti Haukkanen. Bounds for determinants of meet matrices associated with incidence functions. Linear Algebra and its Applications, 329(1– 3):77–88, May 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej%2Dng/10/
                                        30/19/154/25/29/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej%2Dng/10/30/19/154/
                                        25/29/article.pdf.

Korkee:2003:MJM

[KH03] Ismo Korkee and Pentti Haukkanen. On meet and join matrices associated with incidence functions. Linear Algebra and its Applications, 372(1):127–153, October 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Korkee:2008:DMJ

[KH08] Ismo Korkee and Pentti Haukkanen. On the divisibility of meet and join matrices. Linear Algebra and its Applications, 429(8–9):1929– 1943, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kennedy:2009:IPP

[KH09] Shannon C. Kennedy and Ronald D. Haynes. Inverse positivity of perturbed tridiagonal M-matrices. Linear Algebra and its Applications, 430(8–9):2312–2323, April 15, 2009. CODEN LAAPAW. ISSN 0024-

Khare:2009:VSU

[Kha09]  Apoorva Khare.  Vector

spaces as unions of proper subspaces. Linear Algebra and its Applications, 431(9): 1681–1686, October 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kamvar:2004:AMC

[KHG04] Sepandar Kamvar, Taher Haveliwala, and Gene Golub. Adaptive methods for the computation of PageRank. Linear Algebra and its Applications, 386(1):51–65, July 15, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2005:AHS

[Kim05] Sang Og Kim. Automorphisms of Hilbert space effect algebras. Linear Algebra and its Applications, 402(1): 193–198, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kintzel:2005:PPF

[Kin05] Ulric Kintzel. Procrustes problems in finite dimensional indefinite scalar product spaces. Linear Algebra and its Applications, 402(1): 1–28, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirkland:2000:SPS

[Kir00] Steve Kirkland. On the sequence of powers of a stochastic matrix with large exponent. Linear Algebra and its Applications, 310(1– 3):109–122, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/127/25/34/article.
pdf.

Kirkland:2003:UBP

[Kir03a]  S. Kirkland.  An upper

bound on the Perron value of an almost regular tournament matrix. Linear Algebra and its Applications, 361(C):7–22, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
28/abstract.html.

Kirkland:2003:NLP

[Kir03b] Steve Kirkland. A note on limit points for algebraic connectivity. Linear Algebra and its Applications, 373(1):5–11, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirkland:2004:DBC

[Kir04] S. Kirkland. Digraph-based conditioning for Markov chains. Linear Algebra and its Applications, 385(1):81– 93, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirkland:2005:SSD

[Kir05] Steve Kirkland. Sensitivity of the stable distribution vector for a size-classified population model. Linear Algebra and its Applications, 398(1): 3–23, March 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirkland:2006:CES

[Kir06] S. Kirkland. Conditioning of the entries in the stationary vector of a Google-type matrix. Linear Algebra and its Applications, 418(2–3):665– 681, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirkland:2007:CLI

[Kir07]  Steve Kirkland.  Con-

structably Laplacian integral graphs. Linear Algebra and its Applications, 423 (1):3–21, May 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kittaneh:2004:NIS

[Kit04] Fuad Kittaneh. Norm inequalities for sums and differences of positive operators. Linear Algebra and its Applications, 383(1):85– 91, May 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kittaneh:2009:SVI

[Kit09] Fuad Kittaneh. Singular value inequalities for commutators of Hilbert space operators. Linear Algebra and its Applications, 430(8–9):2362– 2367, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kwak:2001:CPG

[KK01] Jin Ho Kwak and Young Soo Kwon. Characteristic polynomials of graph bundles having voltages in a dihedral group. Linear Algebra and its Applications, 336(1–3): 99–118, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/166/27/34/article.
pdf.

Koca:2006:MSC

[KKABAF06] Mehmet Koca, Ramazan Koç, Muataz Al-Barwani, and Shadia Al-Farsi. Maximal subgroups of the Coxeter group W (H_{4}) and quaternions. Linear Algebra and its Applications, 412(2–3):441– 452, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Koca:2007:CGO

[KKK07] Mehmet Koca, Ramazan Koç, and Nazife Ö. Koca. The Chevalley group G_{2}(2) of order 12 096 and the octonionic root system of E_{7}. Linear Algebra and its Applications, 422(2–3):808– 823, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2008:GCP

[KKL08] Dongseok Kim, Hye Kyung Kim, and Jaeun Lee. Generalized characteristic polynomials of graph bundles. Linear Algebra and its Applications, 429(4):688–697, August 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kalogeropoulos:2007:RDI

[KKP07] Grigoris I. Kalogeropoulos, Athanasios D. Karageorgos, and Athanasios A. Pantelous. On the relation of the Drazin inverse and matrix pencil theory methods for the study of generalized linear systems.

Linear Algebra and its Applications, 427(2–3):197–205, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kalogeropoulos:2008:SCM

[KKP08] Grigoris I. Kalogeropoulos, Athanasios D. Karageorgos, and Athanasios A. Pantelous. A stabilization criterion for matrix pencils under bilinear transformation. Linear Algebra and its Applications, 428(11–12):2852– 2862, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2007:CIH

[KL07a] Sejong Kim and Yongdo Lim. A converse inequality of higher order weighted arithmetic and geometric means of positive definite operators. Linear Algebra and its Applications, 426(2–3):490– 496, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Koo:2007:PFS

[KL07b] Yoo Young Koo and Jae Kun Lim. Perturbation of frame sequences and its applications to shift-invariant spaces. Linear Algebra and its Applications, 420(2–3): 295–309, January 15, 2007. CODEN LAAPAW. ISSN

Klein:2000:GWF

[Kle00]  André Klein.  A gener-

alization of Whittle’s formula for the information matrix of vector-mixed time series. Linear Algebra and its Applications, 321(1–3): 197–208, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/37/article.
pdf.

Klein:2004:ISB

[Kle04] Thomas Klein. Invariant symmetric block matrices for the design of mixture experiments. Linear Algebra and its Applications, 388(1):261– 278, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Klemes:2007:AIC

[Kle07] Ivo Klemeš. Alexandrov’s inequality and conjectures on some Toeplitz matrices. Linear Algebra and its Applications, 422(1):164– 185, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kleinsteuber:2009:SJA

[Kle09] Martin Kleinsteuber. A sort-Jacobi algorithm for semisimple Lie algebras. Linear Algebra and its Applications, 430 (1):155–173, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kubicki:2001:TPM

[KLM01] Grzegorz Kubicki, Jenő Lehel, and Michał Morayne. Totally positive matrices and totally positive hypergraphs. Linear Algebra and its Applications, 331(1– 3):193–202, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/156/25/
                                        40/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/156/25/40/article.
                                        pdf.

Koukouvinos:2003:AFF

[KLMS03] C. Koukouvinos, E. Lappas, M. Mitrouli, and Jennifer Seberry. An algorithm to find formulae and values of minors for Hadamard matrices: II. Linear Algebra and its Applications, 371(1):111–124, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Klyachko:2000:RWS

[Kly00] Alexander A. Klyachko. Random walks on symmetric spaces and inequalities for matrix spectra. Linear Algebra and its Applications, 319(1–3):37–59, November 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/138/24/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/138/24/27/article.
pdf.

Kwon:2009:GMC

[KLY09] K. H. Kwon, Lance L. Littlejohn, and G. J. Yoon. Ghost matrices and a characterization of symmetric Sobolev bilinear forms. Linear Algebra and its Applications, 431(1– 2):104–119, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Karcanias:2004:STB

[KM04] Nicos Karcanias and Marilena Mitrouli. System theoretic based characterisation and computation of the least common multiple of a set of polynomials. Linear Algebra and its Applications, 381(1): 1–23, April 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kravvaritis:2007:EMA

[KM07] Christos Kravvaritis and Marilena Mitrouli. Evaluation of minors associated to weighing matrices. Linear Algebra and its Applications, 426(2–3):774–809, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Klin:2004:DSR

[KMMZ04] Mikhail Klin, Akihiro Munemasa, Mikhail Muzychuk, and Paul-Hermann Zieschang. Directed strongly regular graphs obtained from coherent algebras. Linear Algebra and its Applications, 377 (1):83–109, January 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirkland:2002:GEA

[KMNS02] Stephen J. Kirkland, Jason J. Molitierno, Michael Neumann, and Bryan L. Shader. On graphs with equal algebraic and vertex connectivity. Linear Algebra and its Applications, 341(1– 3):45–56, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/171/27/
                                        31/abstract.html.

Kirkland:2005:P

[KMOT05] S. Kirkland, J. J. McDonald, D. D. Olesky, and M. J. Tsatsomeros. Preface. Linear Algebra and its Applications, 409(1):1– 12, November 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Konstantinov:2000:PSL

[KMP00] M. Konstantinov, V. Mehrmann, and P. Petkov. On properties of Sylvester and Lyapunov operators. Linear Algebra and its Applications, 312 (1–3):35–71, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/129/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/129/25/27/article.
pdf.

Koukouvinos:2000:GGE

[KMS00] C. Koukouvinos, M. Mitrouli, and Jennifer Seberry. Growth in Gaussian elimination for weighing matrices, W ( n, n 1). Linear Algebra and its Applications, 306(1–3): 189–202, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
40/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/121/25/40/article.
                                        pdf.

Koukouvinos:2001:AFF

[KMS01] C. Koukouvinos, M. Mitrouli, and Jennifer Seberry. An algorithm to find formulae and values of minors for Hadamard matrices. Linear Algebra and its Applications, 330(1–3):129–147, June 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/155/25/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/155/25/34/article.
                                        pdf.

Klein:2005:RPF

[KMS05a] André Klein, Guy Mélard, and Peter Spreij. On the resultant property of the Fisher information matrix of a vector ARMA process. Linear Algebra and its Applications, 403(1):291–313, July 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Koukouvinos:2005:VMS

[KMS05b] C. Koukouvinos, M. Mitrouli, and Jennifer Seberry. Values of minors of some infinite families of matrices constructed from supplementary difference sets and their application to the growth problem. Linear Algebra and its Applications, 406(1):218–234, September 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kravvaritis:2005:GPS

[KMS05c] C. Kravvaritis, M. Mitrouli, and Jennifer Seberry. On the growth problem for skew and symmetric conference matrices. Linear Algebra and its Applications, 403(1):183– 206, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Klein:2000:CEF

[KMZ00] André Klein, Guy Mélard, and Toufik Zahaf. Construction of the exact Fisher information matrix of Gaussian time series models by means of matrix differential rules. Linear Algebra and its Applications, 321(1–3): 209–232, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/38/article.
pdf.

Kirkland:2000:RMC

[KN00a] Stephen J. Kirkland and Michael Neumann. Regular Markov chains for which the transition matrix has large exponent. Linear Algebra and its Applications, 316(1– 3):45–65, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/28/article.
pdf.

Klein:2000:DDE

[KN00b] A. Klein and H. Neudecker. A direct derivation of the exact Fisher information matrix of Gaussian vector state space models. Linear Algebra and its Applications, 321(1–3): 233–238, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/39/article.
pdf.

Kuba:2001:RDS

[KN01] Attila Kuba and Maurice Nivat. Reconstruction of discrete sets with absorption. Linear Algebra and its Applications, 339(1–3): 171–194, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/169/27/
37/abstract.html;   http:

Knuppel:2003:FAI

[KN03a] Frieder Knüppel and Klaus Nielsen. k-fold anti-invariant subspaces of a linear mapping. Linear Algebra and its Applications, 375(1):13– 19, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Knyazev:2003:GTP

[KN03b] Andrew V. Knyazev and Klaus Neymeyr. A geometric theory for preconditioned inverse iteration III: A short and sharp convergence estimate for generalized eigenvalue problems. Linear Algebra and its Applications, 358(1–3):95–114, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/207/25/
31/abstract.html.

King:2004:NTN

[KN04] Christopher King and Michael Nathanson. New trace norm inequalities for 2 × 2 blocks of diagonal matrices. Linear Algebra and its Applications, 389(1):77–93, September 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

King:2006:ECQ

[KN06a] Christopher King and Michael Nathanson. On the existence of a common quadratic Lyapunov function for a rank one difference. Linear Algebra and its Applications, 419(2–3):400–416, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Knuppel:2006:ECN

[KN06b] Frieder Knüppel and Klaus Nielsen. The extended covering number of SL_{n} is n + 1. Linear Algebra and its Applications, 418(2–3):634–656, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirkland:2009:CED

[KN09] Stephen J. Kirkland and Michael Neumann. The case of equality in the Dobrushin– Deutsch–Zenger bound. Linear Algebra and its Applications, 431(12):2373–2394, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

King:2005:MPE

[KNR05] Christopher King, Michael Nathanson, and Mary Beth Ruskai. Multiplicativity properties of entrywise positive maps. Linear Algebra and its Applications, 404(1): 367–379, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Knutson:2000:SAG

[Knu00] Allen Knutson. The symplectic and algebraic geometry of Horn’s problem. Linear Algebra and its Applications, 319(1–3):61–81, November 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/138/24/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/138/24/28/article.
pdf.

Kirkland:2007:TMW

[KNX07] S. J. Kirkland, Michael Neumann, and Jianhong Xu. Transition matrices for wellconditioned Markov chains. Linear Algebra and its Applications, 424(1):118–131, July 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kuzma:2006:AMS

[KO06] B. Kuzma and M. Orel. Additive mappings on symmetric matrices. Linear Algebra and its Applications, 418(1):277– 291, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kwak:2008:ATE

[KO08] Jin Ho Kwak and Ju-Mok Oh. Arc-transitive elementary abelian covers of the octahedron graph. Linear Algebra and its Applications, 429(8–9):2180–2198, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kobilinsky:2000:RIN

[Kob00] A. Kobilinsky. Reparametrisation of interest in nonuniform factorial designs. Linear Algebra and its Applications, 321(1–3):239–280, December 15, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/140/24/
                                        40/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/140/24/40/article.
                                        pdf.

Kobal:2005:BCF

[Kob05] Damjan Kobal. Belitskii’s canonical form for 5 × 5 upper triangular matrices under upper triangular similarity. Linear Algebra and its Applications, 403(1):178– 182, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kolotilina:2002:OCB

[Kol02] L. Yu. Kolotilina. Optimally conditioned block matrices. Linear Algebra and its Applications, 340(1–3): 55–67, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
30/abstract.html.

Kolotilina:2003:GOB

[Kol03a] L. Yu. Kolotilina. Generalizations of the Ostrowski-Brauer theorem. Linear Algebra and its Applications, 364(C):65–80, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
30/abstract.html.

Kolotilina:2003:NSC

[Kol03b] L. Yu. Kolotilina. Nonsingularity/singularity criteria for nonstrictly block diagonally dominant matrices. Linear Algebra and its Applications, 359(1–3): 133–159, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
37/abstract.html.

Kolotilina:2005:BRR

[Kol05] Lily Kolotilina. Book review:

Richard S. Varga, Geršgorin and his circles, Springer series in computational mathematics vol. 36, Springer-Verlag, Berlin, Heidelberg, New York (2004) ISBN-10 3-540-21100-4, x + 226 pp. Linear Algebra and its Applications, 402(1): 397–402, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kolotilina:2008:STS

[Kol08]  L. Yu. Kolotilina.  On

the sharpness of two-sided bounds for the Perron Root and of the related eigenvalue inclusion sets. Linear Algebra and its Applications, 429(10):2521–2539, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kolotilina:2009:BIN

[Kol09] L. Yu. Kolotilina. Bounds for the infinity norm of the inverse for certain M- and Hmatrices. Linear Algebra and its Applications, 430(2–3): 692–702, January 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Komaroff:2008:ENT

[Kom08] N. Komaroff. Enhancements to the von Neumann trace inequality. Linear Algebra and its Applications, 428(4):738– 741, February 1, 2008. CO-

Konstantinova:2008:SPC

[Kon08] Elena Konstantinova. Some problems on Cayley graphs. Linear Algebra and its Applications, 429(11–12):2754– 2769, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kosowski:2000:REP

[Kos00] Przemysław Kosowski. The relative error in the Pruess method for Sturm–Liouville problems. Linear Algebra and its Applications, 309(1– 3):325–337, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
41/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/41/article.
pdf.

Kosir:2003:CHT

[Koš03] Tomaž Košir. The Cayley– Hamilton theorem and inverse problems for multiparameter systems. Linear Algebra and its Applications, 367(C):155–163, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/216/25/
                                        36/abstract.html.

Kosem:2006:IB

[Kos06] Tomaž Kosem. Inequalities between ||f(A + B)|| and ||f(A) + f(B)||. Linear Algebra and its Applications, 418 (1):153–160, October 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kovacs:2002:ACD

[Kov02] István Kovács. On automorphisms of circulant digraphs on p^{m} vertices, p an odd prime. Linear Algebra and its Applications, 356(1–3): 231–252, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/202/25/
                                        42/abstract.html.

Kim:2007:IAS

[KOvdD07] In-Jae Kim, D. D. Olesky, and P. van den Driessche. Inertially arbitrary sign patterns with no nilpotent realization. Linear Algebra and its Applications, 421(2– 3):264–283, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kozyakin:2009:AAS

[Koz09] Victor Kozyakin. On accuracy of approximation of the spectral radius by the Gelfand formula. Linear Algebra and its Applications, 431(11):2134–2141, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kalantari:2001:DLB

[KP01] Bahman Kalantari and Thomas H. Pate. A determinantal lower bound. Linear Algebra and its Applications, 326(1–3): 151–159, March 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/148/24/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/148/24/33/article.
pdf.

Kosir:2002:SIS

[KP02] Tomaž Košir and Bor Plestenjak. On stability of invariant subspaces of commuting matrices. Linear Algebra and its Applications, 342(1–3): 133–147, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/173/27/
36/abstract.html.

Kala:2006:TLO

[KP06] Radosław Kala and Paweł Pordzik. Two local operators and the BLUE. Linear Algebra and its Applications, 417 (1):134–139, August 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kiukas:2008:NIE

[KP08] J. Kiukas and J.-P. Pellonpää. A note on infinite extreme correlation matrices. Linear Algebra and its Applications, 428(11–12):2501– 2508, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirmani:2009:MPC

[KP09] S. N. U. A. Kirmani and Alan M. Polansky. Multivariate process capability via Löwner ordering. Linear Algebra and its Applications, 430(10):2681–2689, May 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kalogeropoulos:2004:CJC

[KPK04] G. Kalogeropoulos, P. Psarrakos, and N. Karcanias. On the computation of the Jordan canonical form of regular matrix polynomials. Linear Algebra and its Applications, 385(1):117–130, July 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kuijper:2007:PDP

[KPP07] Margreta Kuijper, Raquel Pinto, and Jan Willem Polderman. The predictable degree property and row reducedness for systems over a finite ring. Linear Algebra and its Applications, 425(2–3):776–796, September 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kirkland:2001:LSH

[KPT01] S. Kirkland, P. J. Psarrakos, and M. J. Tsatsomeros. On the location of the spectrum of hypertournament matrices. Linear Algebra and its Applications, 323(1–3):37–49, January 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Koiran:2004:RTV

[KPV04] Pascal Koiran, Natacha Portier, and Gilles Villard. A rank theorem for Vandermonde matrices. Linear Algebra and its Applications, 378 (1):99–107, February 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Koh:2003:HM

[KR03] Youngmee Koh and Sangwook Ree. On k-hypertournament matrices. Linear Algebra and its Applications, 373(1):183– 195, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2004:IIM

[KR04] K. H. Kim and F. W. Roush. Inclines and incline matrices: a survey. Linear Algebra and its Applications, 379(1):457– 473, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kermani:2005:TCS

[KR05] H. Momenaee Kermani and M. Radjabalipour. On triangularizability of the commutant of a single matrix. Linear Algebra and its Applications, 406(1):159–164, September 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Koliha:2006:NRL

[KR06] J. J. Koliha and V. Rakočević. The nullity and rank of linear combinations of idempotent matrices. Linear Algebra and its Applications, 418(1): 11–14, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kaneko:2007:TIP

[KR07] Osamu Kaneko and Paolo Rapisarda. On the Takagi interpolation problem. Linear Algebra and its Applications, 425(2–3):453–470, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kopparty:2008:MRP

[KR08] Swastik Kopparty and K. P. S. Bhaskara Rao. The minimum rank problem: A counterexample. Linear Algebra and its Applications, 428 (7):1761–1765, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kratz:2001:BMD

[Kra01] Werner Kratz. Banded matrices and difference equations. Linear Algebra and its Applications, 337(1–3): 1–20, November 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/167/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/167/27/27/article.
pdf.

Krause:2003:ACD

[Kra03a] Amrey Krause. Algebraic computations in derived categories. Linear Algebra and its Applications, 365 (C):247–266, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
34/abstract.html.

Krause:2003:SPA

[Kra03b] Henning Krause. A short proof for Auslander’s defect formula. Linear Algebra and its Applications, 365 (C):267–270, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/214/25/
                                        42/abstract.html.

Kramar:2004:IGS

[Kra04]  Marjeta Kramar.  Irre-

ducible groups with submultiplicative spectrum. Linear Algebra and its Applications, 378(1):273–282, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Krattenthaler:2005:ADC

[Kra05] C. Krattenthaler. Advanced determinant calculus: A complement. Linear Algebra and its Applications, 411(1): 68–166, December 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kressner:2006:PQA

[Kre06] Daniel Kressner. The periodic QR algorithm is a disguised QR algorithm. Linear Algebra and its Applications, 417(2–3):423–433, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kribs:2005:QCP

[Kri05] David W. Kribs. A quantum computing primer for operator theorists. Linear Algebra and its Applications, 400(1): 147–167, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kronewitter:2001:UNG

[Kro01] F. Dell Kronewitter. Using noncommutative Gröbner bases in solving partially prescribed matrix inverse completion problems. Linear Algebra and its Applications, 338(1–3):171–199, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
40/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/40/article.
pdf.

Kruglyak:2003:DSM

[KRS03] Stanislav Kruglyak, Vyacheslav Rabanovich, and Yuriĭ Samoĭlenko. Decomposition of a scalar matrix into a sum of orthogonal projections. Linear Algebra and its Applications, 370(1):217– 225, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Koliha:2004:DSP

[KRS04] J. J. Koliha, V. Rakočević, and I. Straškraba. The difference and sum of projectors. Linear Algebra and its Applications, 388(1):279–288, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Katre:2009:SLS

[KRS09] S. A. Katre, R. A. Rao, and D. N. Sheth. Solving linear systems via Pfaffians. Linear Algebra and its Applications, 430(4):968– 975, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Krupnik:2004:LJC

[Kru04] Naum Krupnik. On the lengths of Jordan chains for eigenvalues on the boundary of numerical range of quadratic operator polynomial. Linear Algebra and its Applications, 385(1):131– 147, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2000:LSS

[KS00a] Si-Ju Kim and Bryan L. Shader. Linear systems with signed solutions. Linear Algebra and its Applications, 313(1–3):21–40, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/130/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/130/25/26/article.
pdf.

Krafft:2000:CM

[KS00b] Olaf Krafft and Martin Schaefer. Centrogonal Matrices. Linear Algebra and its Applications, 306(1–3): 145–154, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/121/25/36/article.
pdf.

Klein:2001:SEV

[KS01] Andre Klein and Peter Spreij. On Stein’s equation, Vandermonde matrices and Fisher’s information matrix of time series processes. Part I: The autoregressive moving average process. Linear Algebra and its Applications, 329(1– 3):9–47, May 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/154/25/26/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/154/
25/26/article.pdf.

Kim:2002:MWS

[KS02] Si-Ju Kim and Bryan L. Shader. On matrices which have signed null-spaces. Linear Algebra and its Applications, 353(1–3):245–255, September 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kukush:2004:NMI

[KS04] Alexander Kukush and Hans Schneeweiss. A note on a matrix inequality for generalized means. Linear Algebra and its Applications, 388(1):289– 294, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Klein:2005:SSE

[KS05] André Klein and Peter Spreij. On the solution of Stein’s equation and Fisher’s information matrix of an AR-MAX process. Linear Algebra and its Applications, 396(1): 1–34, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

King:2006:SCN

[KS06a] Christopher King and Robert Shorten. Singularity conditions for the non-existence of a common quadratic Lyapunov function for pairs of third order linear time invariant dynamic systems. Linear Algebra and its Applications, 413(1):24–35, February 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Klein:2006:BSS

[KS06b] André Klein and Peter Spreij. The Bezoutian, state space realizations and Fisher’s information matrix of an ARMA process. Linear Algebra and its Applications, 416(1):160–174, July 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Klein:2006:EEF

[KS06c] André Klein and Peter Spreij. An explicit expression for the Fisher information matrix of a multiple time series process. Linear Algebra and its Applications, 417(1):140– 149, August 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Krupnik:2006:SMG

[KS06d] Naum Krupnik and Ilya M. Spitkovsky. Sets of matrices with given joint numerical range. Linear Algebra and its Applications, 419(2–3):569– 585, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kwak:2006:ZFL

[KS06e] Jin Ho Kwak and Iwao Sato. Zeta functions of line, middle, total graphs of a graph and their coverings. Linear Algebra and its Applications, 418 (1):234–256, October 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kurata:2007:GMO

[KS07a] Hiroshi Kurata and Tadashi Sakuma. A group majorization ordering for Euclidean distance matrices. Linear Algebra and its Applications, 420(2–3):586–595, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kurmayya:2007:NMP

[KS07b] T. Kurmayya and K. C. Sivakumar. Nonnegative Moore–Penrose inverses of Gram operators. Linear Algebra and its Applications, 422(2–3):471–476, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2008:FPV

[KS08] In-Jae Kim and Bryan L. Shader. On Fiedler- and Parter-vertices of acyclic matrices. Linear Algebra and its Applications, 428(11–12): 2601–2613, June 1, 2008. CO-

Klein:2009:MDC

[KS09] André Klein and Peter Spreij. Matrix differential calculus applied to multiple stationary time series and an extended Whittle formula for information matrices. Linear Algebra and its Applications, 430(2–3):674–691, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2005:NPM

[KSH05] Byeong Moon Kim, Byung Chul Song, and Woonjae Hwang. Nonnegative primitive matrices with exponent 2. Linear Algebra and its Applications, 407(1):162–168, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2007:PGG

[KSH07] Byeong Moon Kim, Byung Chul Song, and Woonjae Hwang. Primitive graphs with given exponents and minimum number of edges. Linear Algebra and its Applications, 420(2–3):648–662, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2008:WTT

[KSH08] Byeong Moon Kim, Byung Chul Song, and Woonjae Hwang. Wielandt type theorem for Cartesian product of digraphs. Linear Algebra and its Applications, 429(4):841– 848, August 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Karner:2003:SDR

[KSU03] Herbert Karner, Josef Schneid, and Christoph W. Ueberhuber. Spectral decomposition of real circulant matrices. Linear Algebra and its Applications, 367(C):301– 311, July 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/216/25/
                                        47/abstract.html.

Kratz:2008:RFC

[KT08] Werner Kratz and Markus Tentler. Recursion formulae for the characteristic polynomial of symmetric banded matrices. Linear Algebra and its Applications, 428(11–12): 2482–2500, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kojima:2006:CSD

[KTKR06] Chiaki Kojima, Kiyotsugu Takaba, Osamu Kaneko, and Paolo Rapisarda. A characterization of solutions of the discrete-time algebraic Riccati equation based on quadratic difference forms. Linear Algebra and its Applications, 416(2–3):1060– 1082, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kharaghani:2008:EM

[KTR08] H. Kharaghani and B. Tayfeh-Rezaie. On the energy of (0, 1)-matrices. Linear Algebra and its Applications, 429(8–9):2046–2051, October 16, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kulpa:2001:DRI

[Kul01] Zenon Kulpa. Diagrammatic representation for interval arithmetic. Linear Algebra and its Applications, 324(1– 3):55–80, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/144/24/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/144/24/28/article.
pdf.

Kung:2006:MBB

[Kun06] Joseph P. S. Kung. Minimal blocks of binary evenweight vectors. Linear Algebra and its Applications, 416 (2–3):288–297, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kuroda:2004:QGS

[Kur04] Mitsuru Kuroda. Quadratic Gauss sums on matrices. Linear Algebra and its Applications, 384(1):187–198, June 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kuzma:2002:AIP

[Kuz02a] B. Kuzma. Additive idempotence preservers. Linear Algebra and its Applications, 355(1–3):103–117, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/201/27/
                                        34/abstract.html.

Kuzma:2002:AMD

[Kuz02b]  Bojan Kuzma.  Additive

mappings decreasing rank one. Linear Algebra and its Applications, 348(1–3): 175–187, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/195/27/
                                        39/abstract.html.

Katkova:2006:SCT

[KV06] Olga M. Katkova and Anna M. Vishnyakova. On sufficient conditions for the total positivity and for the multiple positivity of matrices. Linear Algebra and its Applications, 416(2–3):1083–1097, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Karampetakis:2009:FMI

[KV09] N. P. Karampetakis and S. Vologiannidis. On the fundamental matrix of the inverse of a polynomial matrix and applications to ARMA representations. Linear Algebra and its Applications, 431(11):2261–2276, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kaliuzhnyi-Verbovetskyi:2009:SRF

[KVV09] Dmitry S. Kaliuzhnyi-Verbovetskyi and Victor Vinnikov. Singularities of rational functions and minimal factorizations: The noncommutative and the commutative setting. Linear Algebra and its Applications, 430(4):869– 889, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kim:2000:BIF

[KY00] Sang Wook Kim and Jae Heon Yun. Block ILU factorization preconditioners for a blocktridiagonal H-matrix. Linear Algebra and its Applications, 317(1–3):103–125, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/136/25/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/136/25/29/article.
                                        pdf.

Khalifeh:2008:MMC

[KYAA08] M. H. Khalifeh, H. Yousefi-Azari, and A. R. Ashrafi. A matrix method for computing szeged and vertex PI indices of join and composition of graphs. Linear Algebra and its Applications, 429(11–12):2702–2709, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Kye:2003:FSU

[Kye03] Seung-Hyeok Kye. Facial structures for unital positive linear maps in the twodimensional matrix algebra. Linear Algebra and its Applications, 362(C):57–73, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/211/25/
                                        29/abstract.html.

Kuzmin:2003:RIM

[KZ03] S. G. Kuz’min and A. N. Zubkov. Rings of invariants of 2 × 2 matrices in positive characteristic. Linear Algebra and its Applications, 365 (C):271–278, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
41/abstract.html.

Li:2005:CSS

[LA05] Xiezhang Li and Edward Arroyo. A condition for the superiority of the (2, 2)-step methods over the related Chebyshev method. Linear Algebra and its Applications, 403(1):143–158, July 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ladkani:2008:PCT

[Lad08] Sefi Ladkani. On the periodicity of Coxeter transformations and the non-negativity of their Euler forms. Linear Algebra and its Applications, 428(4):742–753, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lancaster:2005:IVS

[Lan05] Peter Lancaster. Isospectral vibrating systems. Part 1. The spectral method. Linear Algebra and its Applications, 409(1):51–69, November 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lancaster:2008:MUS

[Lan08] Peter Lancaster. Modelupdating for self-adjoint quadratic eigenvalue problems. Linear Algebra and its Applications, 428(11–12): 2778–2790, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laszlo:2003:GBR

[Lás03] Ákos László. A generalization of the bounded real lemma. Linear Algebra and its Applications, 372(1):79– 103, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lasarow:2006:RIG

[Las06]  Andreas Lasarow.  On

rank invariance of generalized Schwarz–Pick–Potapov block matrices of matrixvalued Carathéodory functions. Linear Algebra and its Applications, 413(1):36– 58, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laszlo:2007:PBW

[Lás07] L. László. Proof of Böttcher and Wenzel’s conjecture on commutator norms for 3-by-3 matrices. Linear Algebra and its Applications, 422(2– 3):659–663, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lavri:2003:IIC

[Lav03a] Boris Lavri. The isometries and the G-invariance of certain seminorms. Linear Algebra and its Applications, 374(1):31–40, November 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lavric:2003:MOM

[Lav03b] Boris Lavrič. Monotonicity and ^{∗} orthant-monotonicity of certain maximum norms. Linear Algebra and its Applications, 367(C):29–36, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
28/abstract.html.

Lavric:2005:ICM

[Lav05] Boris Lavrič. The isometries of certain maximum norms. Linear Algebra and its Applications, 405(1):249– 263, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2000:TLE

[LB00] Bolian Liu and Zhou Bo. On the third largest eigenvalue of a graph. Linear Algebra and its Applications, 317(1–3): 193–200, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/136/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/136/25/33/article.
pdf.

Lin:2001:GSC

[LB01] Zhiping Lin and N. K. Bose. A generalization of Serre’s conjecture and some related issues. Linear Algebra and its Applications, 338(1–3): 125–138, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/36/article.
pdf.

Liang:2004:UBP

[LB04] Heng Liang and Fengshan Bai. An upper bound for the permanent of (0, 1)-matrices. Linear Algebra and its Applications, 377(1):291–295, January 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2009:TIH

[LC09] Qingbing Liu and Guoliang Chen. On two inequalities for the Hadamard product and the Fan product of matrices. Linear Algebra and its Applications, 431(5– 7):974–984, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2006:AOR

[LCF06] Herng-Jer Lee, Chia-Chi Chu, and Wu-Shiung Feng. An adaptive-order rational Arnoldi method for modelorder reductions of linear time-invariant systems. Linear Algebra and its Applications, 415(2–3):235–261, June 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2003:DWT

[LCN03] Fu-Rong Lin, Wai-Ki Ching, and Michael K. Ng. Discrete wavelet transforms for Toeplitz matrices. Linear Algebra and its Applications, 370(1):269–285, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2004:PRL

[LCN04] Fu-Rong Lin, Wai-Ki Ching, and Michael K. Ng. Preconditioning regularized least squares problems arising from high-resolution image reconstruction from lowresolution frames. Linear Algebra and its Applications, 391(1):149–168, November 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lopez:2002:SFS

[LCP02] Marco A. López, María J. Cánovas, and Juan Parra. Stability of the feasible set for linear inequality systems: A carrier index set approach. Linear Algebra and its Applications, 357(1–3): 83–105, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/203/25/
                                        31/abstract.html.

Loly:2009:MSS

[LCTS09] Peter Loly, Ian Cameron, Walter Trump, and Daniel Schindel. Magic square spectra. Linear Algebra and its Applications, 430(10):2659– 2680, May 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2009:NLB

[LCW09] Yao-Tang Li, Fu-Bin Chen, and De-Feng Wang. New lower bounds on eigenvalue of the Hadamard product of an M-matrix and its inverse. Linear Algebra and its Applications, 430(4):1423–1431, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2002:IIO

[LD02] Baibing Li and Bart De Moor. Identification of influential observations on total least squares estimates. Linear Algebra and its Applications, 348(1–3):23–39, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
29/abstract.html.

Lenzing:2009:CRF

[LdlP09] Helmut Lenzing and José A. de la Peña. A chebysheff recursion formula for Coxeter polynomials. Linear Algebra and its Applications, 430(4): 947–956, February 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2007:LDP

[LDS07] Dayan Liu, Xiankun Du, and Xiaosong Sun. The linear dependence problem for power linear maps. Linear Algebra and its Applications, 426(2–3):706–715, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2008:QLK

[LDS08] Dayan Liu, Xiankun Du, and Xiaosong Sun. Quadratic linear Keller maps of nilpotency index three. Linear Algebra and its Applications, 429(1): 12–17, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2000:LNA

[Lee00] Gwang-Yeon Lee. k-Lucas numbers and associated bipartite graphs. Linear Algebra and its Applications, 320 (1–3):51–61, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/139/24/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/139/24/28/article.
                                        pdf.

Lee:2009:MRC

[Lee09a] Eun-Young Lee. A matrix reverse Cauchy–Schwarz inequality. Linear Algebra and its Applications, 430(2–3): 805–810, January 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2009:FPD

[Lee09b]  Tsiu-Kwen Lee.  Finite-

ness properties of differential polynomials. Linear Algebra and its Applications, 430(8– 9):2030–2041, April 15, 2009.

Legisa:2004:APO

[Leg04]  Peter Legiša.  Automor-

phisms of M_{n}, partially ordered by rank subtractivity ordering. Linear Algebra and its Applications, 389(1): 147–158, September 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2002:CSR

[LEL02] Wen Li, Ludwig Elsner, and Linzhang Lu. Comparisons of spectral radii and the theorem of Stein– Rosenberg. Linear Algebra and its Applications, 348(1– 3):283–287, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
49/abstract.html.

Lemos:2004:NSC

[Lem04]  Manoel Lemos.  Non-

separating cocircuits in binary matroids. Linear Algebra and its Applications, 382(1):171–178, May 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lesnjak:2000:SEL

[Leš00] Gorazd Lešnjak. Semigroups of EP linear transformations. Linear Algebra and its Applications, 304(1–3): 109–118, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/118/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/118/25/30/article.
pdf.

Lesnjak:2001:NCC

[Leš01] Gorazd Lešnjak. A note on constrained convergence. Linear Algebra and its Applications, 325(1–3):75–80, March 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/147/24/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/147/24/29/article.
pdf.

Lesnjak:2002:APN

[Leš02]  Gorazd Lešnjak.  Addi-

tive preservers of numerical range. Linear Algebra and its Applications, 345(1– 3):235–253, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/175/27/
38/abstract.html.

Lesnjak:2004:NAV

[Leš04] Gorazd Lešnjak. A note on an algebraic version of Cochran’s theorem. Linear Algebra and its Applications, 388(1):295–299, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2003:PDS

[LG03] Xin-Guo Liu and Hua Gao. On the positive definite solutions of the matrix equations X^{s} ± A^{T} X^{−t}A = I_{n}. Linear Algebra and its Applications, 368(C):83–97, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
30/abstract.html.

Luo:2004:RTF

[LG04] Jingxiang Luo and Winfried K. Grassmann. Ratetilting for fast simulation of level/phase processes. Linear Algebra and its Applications, 386(1):261–284, July 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2006:OTT

[LG06] Wenshui Lin and Xiaofeng Guo. Ordering trees by their largest eigenvalues. Linear Algebra and its Applications, 418(2–3):450–456, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lungo:2001:SID

[LGH01] Alberto Del Lungo, Paolo Gronchi, and Gabor T. Herman. Special issue on Discrete Tomography. Linear Algebra and its Applications, 339(1–3):1–2, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/169/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/169/27/27/article.
pdf.

Lin:2001:PAE

[LgS01] Wen-Wei Lin and Ji guang Sun. Perturbation analysis for the eigenproblem of periodic matrix pairs. Linear Algebra and its Applications, 337(1–3):157–187, November 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/167/27/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/167/27/35/article.
pdf.

Lenzing:2003:AFQ

[LH03] Helmut Lenzing and Liane Hasenberg. Additive functions on quivers. Linear Algebra and its Applications, 365 (C):279–289, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
46/abstract.html.

Liu:2004:SPS

[LH04] Jianzhou Liu and Yunqing Huang. Some properties on Schur complements of H-matrices and diagonally dominant matrices. Linear Algebra and its Applications, 389(1):365–380, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2005:SUL

[LH05] Dong Liu and Naihong Hu. Steinberg unitary Leibniz algebras. Linear Algebra and its Applications, 405(1):279– 303, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2006:NIB

[LHF06] Xiao-Qin Liu, Ting-Zhu Huang, and Ying-Ding Fu. A note on improvement on bounds for nonmaximal eigenvalues of symmetric positive matrices. Linear Algebra and its Applications, 419(2–3):612–617, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Huang:2008:DPS

[lHH08] Wen ling Huang and Hans Havlicek. Diameter preserving surjections in the geometry of matrices. Linear Algebra and its Applications, 429(1):376–386, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2008:ISS

[LHL08] Hou-Biao Li, Ting-Zhu Huang, and Hong Li. Inclusion sets for singular values. Linear Algebra and its Applications, 428(8–9):2220–2235, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2002:IPR

[LHS02] Zhongshan Li, Frank J. Hall, and Jeffrey L. Stuart. Irreducible powerful ray pattern matrices. Linear Algebra and its Applications, 342(1– 3):47–58, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/173/27/
                                        30/abstract.html.

Li:2005:RPR

[LHS05] Zhongshan Li, Frank J. Hall, and Jeffrey L. Stuart. Reducible powerful ray pattern matrices. Linear Algebra and its Applications, 399(1):125– 140, April 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2007:LBM

[LHSL07] Hou-Biao Li, Ting-Zhu Huang, Shu-Qian Shen, and Hong Li. Lower bounds for the minimum eigenvalue of Hadamard product of an M-matrix and its inverse. Linear Algebra and its Applications, 420(1): 235–247, January 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2004:SCG

[LHZ04] Jianzhou Liu, Yunqing Huang, and Fuzhen Zhang. The Schur complements of generalized doubly diagonally dominant matrices. Linear Algebra and its Applications, 378(1):231–244, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2000:SPC

[Li00a] Chi-Kwong Li. A simple proof of the Craig-Sakamoto theorem. Linear Algebra and its Applications, 321(1–3): 281–283, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/140/24/
                                        41/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/140/24/41/article.
                                        pdf.

Li:2000:RPT

[Li00b] Ren-Cang Li. Relative perturbation theory: IV. sin 2θ theorems. Linear Algebra and its Applications, 311 (1–3):45–60, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/128/25/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/128/25/29/article.
                                        pdf.

Li:2001:CMO

[Li01] Xiezhang Li. Is a Chebyshev method optimal for an elliptic region also optimal for a nearly elliptic region? Linear Algebra and its Applications, 338(1–3): 37–51, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/168/27/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/168/27/30/article.
                                        pdf.

Li:2003:REU

[Li03] Ruiming Li. Relationship of eigenvalues for USAOR iterative method applied to a class of p-cyclic matrices. Linear Algebra and its Applications, 362(C):101–108, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/211/25/
32/abstract.html.

Li:2005:DUG

[Li05] Jianxiang Li. The determinant of a unicyclic graph’s neighborhood matrix. Linear Algebra and its Applications, 394(1):201–216, January 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2006:SPI

[Li06a] Chi-Kwong Li. A short proof of interlacing inequalities on normalized Laplacians. Linear Algebra and its Applications, 414(2–3):425– 427, April 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2006:LOA

[Li06b] Jing-Rebecca Li. Low order approximation of the spherical nonreflecting boundary kernel for the wave equation. Linear Algebra and its Applications, 415(2–3):455– 468, June 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2008:VMC

[Li08a] Ren-Cang Li. Vandermonde matrices with Chebyshev nodes. Linear Algebra and its Applications, 428(8–9):1803– 1832, April 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2008:NGP

[Li08b] Shuchao Li. On the nullity of graphs with pendent vertices. Linear Algebra and its Applications, 429(7): 1619–1628, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2008:PBU

[Li08c] Wen Li. On the perturbation bound in unitarily invariant norms for subunitary polar factors. Linear Algebra and its Applications, 429 (2–3):649–657, July 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2008:MPI

[Li08d] Yuan Li. The Moore–Penrose inverses of products and differences of projections in a C^{∗}-algebra. Linear Algebra and its Applications, 428(4): 1169–1177, February 1, 2008. CODEN LAAPAW. ISSN

Liess:2009:RLC

[Lie09] Otto Liess. Remarks on the Lax conjecture for hyperbolic polynomials. Linear Algebra and its Applications, 430(8– 9):2123–2132, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2005:AMB

[Lim05a] M. H. Lim. Additive mappings between Hermitian matrix spaces preserving rank not exceeding one. Linear Algebra and its Applications, 408(1):259–267, October 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2005:RON

[Lim05b] Ming-Huat Lim. Rank-one nonincreasing additive mappings on second symmetric product spaces. Linear Algebra and its Applications, 402 (1):263–271, June 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2005:IMP

[Lim05c] Yongdo Lim. The inverse mean problem of geometric mean and contraharmonic means. Linear Algebra and its Applications, 408(1):221– 229, October 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2006:NAM

[Lim06] Ming-Huat Lim. A note on additive mappings decreasing rank one. Linear Algebra and its Applications, 414(2– 3):428–434, April 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2007:HPM

[Lim07a] Yongdo Lim. Hilbert’s projective metric on Lorenz cones and Birkhoff formula for Lorentzian compressions. Linear Algebra and its Applications, 423(2–3):246–254, June 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2007:SMW

[Lim07b] Yongdo Lim. On the symmetric matrix word equation XBX^{2}B^{3}X^{2}BX = A. Linear Algebra and its Applications, 427(1):77–86, November 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2008:APN

[Lim08a] Ming-Huat Lim. Additive preservers of non-zero decomposable tensors. Linear Algebra and its Applications, 428 (1):239–253, January 1, 2008.

Lim:2008:ALM

[Lim08b] Yongdo Lim. On Ando– Li–Mathias geometric mean equations. Linear Algebra and its Applications, 428(8– 9):1767–1777, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2009:LRO

[Lim09a]  Ming-Huat Lim.  Linear

rank one preservers between spaces of matrices with zero trace. Linear Algebra and its Applications, 430(11–12): 2982–2996, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2009:NEB

[Lim09b] Yongdo Lim. Nonlinear equations based on jointly homogeneous mappings. Linear Algebra and its Applications, 430(1):279–285, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2009:SNM

[Lim09c] Yongdo Lim. Solving the nonlinear matrix equation X = Q + ∑_{i=1}^{m} M_{i}^{∗}X^{δ_{i}}M_{i}^{∗} via a contraction principle. Linear Algebra and its Applications, 430(4):1380–1383, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Linden:2002:CRZ

[Lin02a] Hansjörg Linden. Containment regions for zeros of polynomials from numerical ranges of companion matrices. Linear Algebra and its Applications, 350(1–3):125– 145, July 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lindqvist:2002:CPF

[Lin02b] Bo Henry Lindqvist. On comparison of the Perron– Frobenius eigenvalues of two ML-matrices. Linear Algebra and its Applications, 353(1–3):257–266, September 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2003:CPM

[Lin03] Ying-Fen Lin. Commutativitypreserving maps on Lie ideals of prime algebras. Linear Algebra and its Applications, 371(1):361–368, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2007:OOC

[Lin07] C.-S. Lin. On operator order and chaotic operator order for two operators. Linear Algebra and its Applications, 425(1):1–6, August 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2008:EFI

[Lin08a] Fu-Rong Lin. An explicit formula for the inverse of band triangular Toeplitz matrix. Linear Algebra and its Applications, 428(2–3):520–534, January 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2008:LSR

[Lin08b] Liang Lin. The Laplacian spectral radius of graphs on surfaces. Linear Algebra and its Applications, 428(4):973– 977, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liping:2001:CQM

[Lip01] Huang Liping. Consimilarity of quaternion matrices and complex matrices. Linear Algebra and its Applications, 331(1–3):21–30, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/27/article.
pdf.

Lippert:2005:FTE

[Lip05] Ross A. Lippert. Fixing two eigenvalues by a minimal perturbation. Linear Algebra and its Applications, 406(1): 177–200, September 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lipyanski:2008:AES

[Lip08]  R. Lipyanski.  Automor-

phisms of the endomorphism semigroups of free linear algebras of homogeneous varieties. Linear Algebra and its Applications, 429(1):156– 180, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2002:LIM

[Liu02a] Shuangzhe Liu. Local influence in multivariate elliptical linear regression models. Linear Algebra and its Applications, 354(1–3): 159–174, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/200/27/
                                        39/abstract.html.

Liu:2002:SII

[Liu02b] Shuangzhe Liu. Several inequalities involving Khatri– Rao products of positive semidefinite matrices. Linear Algebra and its Applications, 354(1–3):175–186, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
40/abstract.html.

Liu:2003:GEB

[Liu03a] Bolian Liu. Generalized exponents of boolean matrices. Linear Algebra and its Applications, 373(1):169–182, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2003:FLS

[Liu03b] Weijun Liu. Finite linear spaces admitting a projective group PSU(3, q) with q even. Linear Algebra and its Applications, 374(1):291– 305, November 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Livshits:2004:DPS

[Liv04] Leo Livshits. On decomposability of periodic semigroups of non-negative matrices. Linear Algebra and its Applications, 383(1):163– 174, May 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2001:CLP

[LJ01] BoLian Liu and Wen Jiang. On a conjecture of Lewin’s problem. Linear Algebra and its Applications, 323(1–3): 201–206, January 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2004:JEM

[LJ04] Pengtong Li and Wu Jing. Jordan elementary maps on rings. Linear Algebra and its Applications, 382(1):237– 245, May 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2000:CPG

[LK00] Jaeun Lee and Hye Kyung Kim. Characteristic polynomials of graphs having a semifree action. Linear Algebra and its Applications, 307 (1–3):35–46, March 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/27/article.
pdf.

Lee:2003:LAF

[LK03] Gwang-Yeon Lee and Jin-Soo Kim. The linear algebra of the k-Fibonacci matrix. Linear Algebra and its Applications, 373(1):75– 87, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-

Leventides:2009:ZAM

[LK09] John Leventides and Nicos Karcanias. Zero assignment of matrix pencils by additive structured transformations. Linear Algebra and its Applications, 431(8):1380–1396, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lambrou:2002:STF

[LL02] M. S. Lambrou and W. E. Longstaff. Small transitive families of subspaces in finite dimensions. Linear Algebra and its Applications, 357 (1–3):229–245, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
41/abstract.html.

Lu:2003:JIS

[LL03] Fangyan Lu and Pengtong Li. Jordan isomorphisms of J -subspace lattice algebras. Linear Algebra and its Applications, 371(1):255–264, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ling:2004:JMN

[LL04] Zhengchu Ling and Fangyan Lu. Jordan maps of nest algebras. Linear Algebra and its Applications, 387(1):361– 368, August 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2005:NEP

[LL05a] Chi-Kwong Li and Ren-Cang Li. A note on eigenvalues of perturbed Hermitian matrices. Linear Algebra and its Applications, 395(1):183– 190, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2005:AJE

[LL05b] Pengtong Li and Fangyan Lu. Additivity of Jordan elementary maps on nest algebras. Linear Algebra and its Applications, 400(1):327– 338, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lawson:2006:SSM

[LL06a] Jimmie Lawson and Yongdo Lim. Solving symmetric matrix word equations via symmetric space machinery. Linear Algebra and its Applications, 414(2–3):560– 569, April 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2006:FPP

[LL06b] Weijun Liu and Jinglei Li. Finite projective planes admitting a projective linear group PSL(2, q). Linear Algebra and its Applications, 413(1): 121–130, February 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lawson:2008:CTP

[LL08a] Jimmie Lawson and Yongdo Lim. Γ-Commuting triples of positive definite matrices and midpoint equations on unitary matrices. Linear Algebra and its Applications, 428(8– 9):2030–2039, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2008:JEO

[LL08b] Pengtong Li and Fangyan Lu. Jordan elementary operators on J -subspace lattice algebras. Linear Algebra and its Applications, 428(11–12): 2675–2687, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2008:SCP

[LL08c] Jer-Shyong Lin and Cheng-Kai Liu. Strong commutativity preserving maps on Lie ideals. Linear Algebra and its Applications, 428(7): 1601–1609, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2008:SRQ

[LL08d] Huiqing Liu and Mei Lu. On the spectral radius of quasitree graphs. Linear Algebra and its Applications, 428(11– 12):2708–2714, June 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2009:TGG

[LL09a] Shuchao Li and Xuechao Li. On tricyclic graphs of a given diameter with minimal energy. Linear Algebra and its Applications, 430(1):370– 385, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liao:2009:GLD

[LL09b] Ping-Bao Liao and Cheng-Kai Liu. On generalized Lie derivations of Lie ideals of prime algebras. Linear Algebra and its Applications, 430(4):1236–1242, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2008:SPS

[LLHK08] Jianzhou Liu, Jicheng Li, Zhuohong Huang, and Xu Kong. Some properties of Schur complements and diagonal-Schur complements of diagonally dominant matrices.

Linear Algebra and its Applications, 428(4):1009–1030, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2003:IPL

[LLM03] Pengtong Li, Fangyan Lu, and Jipu Ma. Invertibility preserving linear maps on J -subspace lattice algebras. Linear Algebra and its Applications, 372(1):155– 166, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2004:LSR

[LLT04a] Huiqing Liu, Mei Lu, and Feng Tian. On the Laplacian spectral radius of a graph. Linear Algebra and its Applications, 376(1):135– 141, January 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2004:SRG

[LLT04b] Huiqing Liu, Mei Lu, and Feng Tian. On the spectral radius of graphs with cut edges. Linear Algebra and its Applications, 389(1):139–145, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lu:2005:BLS

[LLT05] Mei Lu, Huiqing Liu, and Feng Tian. Bounds of Laplacian spectrum of graphs based on the domination number. Linear Algebra and its Applications, 402(1):390– 396, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lu:2006:NUB

[LLT06] Mei Lu, Huiqing Liu, and Feng Tian. A new upper bound for the spectral radius of graphs with girth at least 5. Linear Algebra and its Applications, 414(2– 3):512–516, April 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2007:SRU

[LLT07] Huiqing Liu, Mei Lu, and Feng Tian. On the spectral radius of unicyclic graphs with fixed diameter. Linear Algebra and its Applications, 420(2–3):449–457, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2009:MTC

[LLY09] Muhuo Liu, Bolian Liu, and Zhifu You. The majorization theorem of connected graphs. Linear Algebra and its Applications, 431(5–7):

553–557, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Livshits:2001:TCP

[LM01] Leo Livshits and Gordon MacDonald. n-transitivity and the complementation property. Linear Algebra and its Applications, 329(1–3): 157–169, May 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/154/25/33/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/154/
25/33/article.pdf.

Li:2002:IIT

[LM02] Chi-Kwong Li and Roy Mathias. Interlacing inequalities for totally nonnegative matrices. Linear Algebra and its Applications, 341(1– 3):35–44, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
30/abstract.html.

Li:2003:RIP

[LM03] Pengtong Li and Jipu Ma. Ring isomorphisms and pentagon subspace lattices. Linear Algebra and its Applications, 367(C):59–66, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/216/25/
                                        31/abstract.html.

Li:2004:USS

[LM04a] Chi-Kwong Li and Tom Milligan. Uniqueness of the solutions of some completion problems. Linear Algebra and its Applications, 392(1):91– 102, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Loewy:2004:SNI

[LM04b] R. Loewy and J. J. McDonald. The symmetric nonnegative inverse eigenvalue problem for 5 × 5 matrices. Linear Algebra and its Applications, 393(1):275–298, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Le:2006:ITM

[LM06a] Hien Thu Le and Judith Joanne McDonald. Inverses of M-type matrices created with irreducible eventually nonnegative matrices. Linear Algebra and its Applications, 419(2–3):668–674, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Letac:2006:ERD

[LM06b] Gérard Letac and Hélène Massam. Extremal rays and duals for cones of positive definite matrices with prescribed zeros. Linear Algebra and its Applications, 418(2–3):737–750, October 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Loehr:2006:BMA

[LM06c] Nicholas A. Loehr and Anthony Mendes. Bijective matrix algebra. Linear Algebra and its Applications, 416 (2–3):917–944, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2009:AHG

[LM09a] Xueliang Li and Hongping Ma. All hypoenergetic graphs with maximum degree at most 3. Linear Algebra and its Applications, 431(11):2127–2133, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Loewy:2009:NPT

[LM09b] Raphael Loewy and Volker Mehrmann. A note on Potter’s theorem for quasicommutative matrices. Linear Algebra and its Applications, 430(7):1812–1825, April 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Luzon:2009:RMR

[LM09c] Ana Luzón and Manuel A. Morón. Riordan matrices in the reciprocation of quadratic polynomials. Linear Algebra and its Applications, 430(8– 9):2254–2270, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Langer:2001:NCB

[LMMT01] H. Langer, A. Markus, V. Matsaev, and C. Tretter. A new concept for block operator matrices: the quadratic numerical range. Linear Algebra and its Applications, 330(1–3):89–112, June 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/155/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/155/25/32/article.
pdf.

Louka:2009:IEL

[LMT09] M. A. Louka, N. M. Missirlis, and F. I. Tzaferis. The impact of eigenvalue locality on the convergence behavior of the PSD method for two-cyclic matrices. Linear Algebra and its Applications, 430(8–9):1929–1944, April 15, 2009. CODEN LAAPAW. ISSN 0024-

Lemmerling:2003:EIS

[LMV03] Philippe Lemmerling, Nicola Mastronardi, and Sabine Van Huffel. Efficient implementation of a structured total least squares based speech compression method. Linear Algebra and its Applications, 366(C):295–315, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
42/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Lancaster:2003:WCS

[LMZ03] P. Lancaster, A. S. Markus, and F. Zhou. A wider class of stable gyroscopic systems. Linear Algebra and its Applications, 370(1):257– 267, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lenker:2003:SSM

[LN03] Terry Lenker and Sivaram Narayan. Spectrally stable matrices. Linear Algebra and its Applications, 362 (C):177–189, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/211/25/
                                        39/abstract.html.

Lu:2004:LPR

[LN04] Linzhang Lu and Michael K. Ng. Localization of Perron roots. Linear Algebra and its Applications, 392(1):103– 117, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lins:2006:ILM

[LN06] Brian Lins and Roger Nussbaum. Iterated linear maps on a cone and Denjoy–Wolff theorems. Linear Algebra and its Applications, 416(2– 3):615–626, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2008:NAA

[LN08] Jongwoo Lee and Ki-Bong Nam. Non-associative algebras containing the matrix ring. Linear Algebra and its Applications, 429(1): 72–78, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Logofet:2005:STL

[Log05] Dmitrii O. Logofet. Strongerthan-Lyapunov notions of matrix stability, or how “flowers” help solve problems in mathematical ecology. Linear Algebra and its Applications, 398(1):75– 100, March 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lomadze:2002:RMT

[Lom02] Vakhtang Lomadze. Rosenbrock models and their homotopy equivalence. Linear Algebra and its Applications, 351–352:519–532, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lomadze:2007:SIV

[Lom07a] Vakhtang Lomadze. State and internal variables for linear systems. Linear Algebra and its Applications, 425(2–3):534–547, September 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lomadze:2007:WLD

[Lom07b] Vakhtang Lomadze. When are linear differentiationinvariant spaces differential? Linear Algebra and its Applications, 424(2–3):540–554, July 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lomadze:2009:HDD

[Lom09] Vakhtang Lomadze. How to define the dual of a higherdimensional linear system.

Linear Algebra and its Applications, 431(11):2084–2101, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Longstaff:2004:BTI

[Lon04] W. E. Longstaff. Burnside’s theorem: irreducible pairs of transformations. Linear Algebra and its Applications, 382(1):247–269, May 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lopatin:2007:BPL

[Lop07] A. A. Lopatin. On block partial linearizations of the pfaffian. Linear Algebra and its Applications, 426(1):109– 129, October 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Longstaff:2000:SEM

[LP00] W. E. Longstaff and Oreste Panaia. Single elements of matrix incidence algebras. Linear Algebra and its Applications, 318(1–3): 117–126, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/137/24/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/137/24/32/article.
pdf.

Li:2001:LOP

[LP01a] Chi-Kwong Li and Edward Poon. Linear operators preserving directional majorization. Linear Algebra and its Applications, 325(1–3): 141–146, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/147/24/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/147/24/33/article.
pdf.

Li:2001:SIE

[LP01b] Chi-Kwong Li and Yiu-Tung Poon. Spectral inequalities and equalities involving products of matrices. Linear Algebra and its Applications, 323(1–3):131–143, January 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2001:CIB

[LP01c] Jiong-Sheng Li and Yong-Liang Pan. de Caen’s inequality and bounds on the largest Laplacian eigenvalue of a graph. Linear Algebra and its Applications, 328(1– 3):153–160, May 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/153/25/31/abstract.
html;   http://www.elsevier.

                                        nl/gej%2Dng/10/30/19/153/
                                        25/31/article.pdf.

Li:2007:SPM

[LP07] Chi-Kwong Li and Edward Poon. Schur product of matrices and numerical radius (range) preserving maps. Linear Algebra and its Applications, 424(1): 8–24, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2009:SUS

[LP09] Chi-Kwong Li and Yiu-Tung Poon. The sum of unitary similarity orbits containing only special operators. Linear Algebra and its Applications, 431(12):2336– 2345, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lur:2003:SNF

[LPG03] Yung-Yih Lur, Chin-Tzong Pang, and Sy-Ming Guu. On simultaneously nilpotent fuzzy matrices. Linear Algebra and its Applications, 367(C):37–45, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/216/25/
                                        29/abstract.html.

Lerer:2004:EMN

[LPR04] L. Lerer, M. A. Petersen, and A. C. M. Ran. Existence of minimal nonsquare Jsymmetric factorizations for self-adjoint rational matrix functions. Linear Algebra and its Applications, 379(1):159– 178, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2005:LMT

[LPS05] Chi-Kwong Li, Yiu-Tung Poon, and Nung-Sing Sze. Linear maps transforming the higher numerical ranges. Linear Algebra and its Applications, 400(1):291–311, May 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liang:2008:CG

[LPW08] Hao Liang, Yong-Liang Pan, and Jian Wang. The critical group of K_{m} × P_{n}. Linear Algebra and its Applications, 428(11–12):2723– 2729, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Luk:2000:FEA

[LQ00] Franklin T. Luk and Sanzheng Qiao. A fast eigenvalue algorithm for Hankel matrices. Linear Algebra and its Applications, 316(1–3):

171–182, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/35/article.
pdf.

Liu:2000:FPR

[LR00] Guoyang Liu and Lewis C. Robertson. Free product Z_{3} ∗ Z_{3} of rotations with rational entries. Linear Algebra and its Applications, 307(1– 3):183–192, March 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/38/article.
pdf.

Langlois:2002:CCE

[LR02] Yves Langlois and Thomas Ransford. Counterexample to a conjecture of Elsner on the spectral variation of matrices. Linear Algebra and its Applications, 349(1– 3):193–195, July 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2003:STR

[LR03] Guoyang Liu and Lewis C. Robertson. Structure theorem for the rotation group over Q. Linear Algebra and its Applications, 362(C): 201–209, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/211/25/
41/abstract.html.

Lomonosov:2004:SPB

[LR04] Victor Lomonosov and Peter Rosenthal. The simplest proof of Burnside’s theorem on matrix algebras. Linear Algebra and its Applications, 383(1):45–47, May 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Labra:2005:RTA

[LR05a] Alicia Labra and Cristián Reyes. Representations on train algebras of rank 3. Linear Algebra and its Applications, 400(1):91– 97, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lancaster:2005:CFS

[LR05b] P. Lancaster and L. Rodman. Canonical forms for symmetric/skew-symmetric real matrix pairs under strict equivalence and congruence. Linear Algebra and its Applications, 406(1):1–76, September 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2005:ICR

[LR05c] Chi-Kwong Li and Leiba Rodman. Inverse closed raynonsingular cones of matrices. Linear Algebra and its Applications, 400(1):203– 230, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2009:PSR

[LR09] Chi-Kwong Li and Leiba Rodman. Preservers of spectral radius, numerical radius, or spectral norm of the sum on nonnegative matrices. Linear Algebra and its Applications, 430(7):1739– 1761, April 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lienemann:2006:MMM

[LRK06] Jan Lienemann, Evgenii B. Rudnyi, and Jan G. Korvink. MST MEMS model order reduction: Requirements and benchmarks. Linear Algebra and its Applications, 415 (2–3):469–498, June 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lindner:2004:FSB

[LRR04] Marko Lindner, Vladimir S. Rabinovich, and Steffen Roch. Finite sections of band operators with slowly oscillating coefficients. Linear Algebra and its Applications, 390(1):19–26, October 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2002:LTB

[LRŠ02] Chi-Kwong Li, Leiba Rodman, and Peter Šemrl. Linear transformations between matrix spaces that map one rank specific set into another. Linear Algebra and its Applications, 357(1–3): 197–208, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
38/abstract.html.

Lagomasino:2008:MMR

[LRW08] G. López Lagomasino, L. Reichel, and L. Wunderlich. Matrices, moments, and rational quadrature. Linear Algebra and its Applications, 429(10):2540–2554, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lemos:2007:PLC

[LRWW07] Manoel Lemos, Talmage James Reid, Bryan Williams, and Haidong Wu. Pairs of largest circuits in 3-connected matroids. Linear Algebra and its Applications, 427(2–3):313– 316, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lewkowicz:2005:CIS

[LRY05] Izchak Lewkowicz, Leiba Rodman, and Elad J. Yarkoni. Convex invertible sets and matrix sign function. Linear Algebra and its Applications, 396(1):329–352, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2000:NRP

[LS00a] Ren-Cang Li and G. W. Stewart. A new relative perturbation theorem for singular subspaces. Linear Algebra and its Applications, 313 (1–3):41–51, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/130/25/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/130/25/27/article.
                                        pdf.

Li:2000:MGS

[LS00b] Wen Li and Weiwei Sun. Modified Gauss–Seidel type methods and Jacobi type methods for Z-matrices. Linear Algebra and its Applications, 317(1–3):227–240, September 15, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/136/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/136/25/38/article.
pdf.

Li:2001:CRP

[LS01] Wen Li and Weiwei Sun. Comparison results for parallel multisplitting methods with applications to AOR methods. Linear Algebra and its Applications, 331(1– 3):131–144, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/35/article.
pdf.

Lewis:2002:QES

[LS02a] Adrian S. Lewis and Hristo S. Sendov. Quadratic expansions of spectral functions. Linear Algebra and its Applications, 340(1–3): 97–121, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
34/abstract.html.

Li:2002:OM

[LS02b] Chi-Kwong Li and Hans Schneider. Orthogonality of matrices. Linear Algebra and its Applications, 347(1– 3):115–122, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/194/27/
                                        34/abstract.html.

Li:2002:CCC

[LS02c] Yuan-Chuan Li and Mau-Hsiang Shih. Contractibility of compact contractions in Hilbert space. Linear Algebra and its Applications, 341 (1–3):369–378, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/171/27/
                                        57/abstract.html.

Lu:2002:ALE

[LS02d] Chang-Yu Lu and Ning-Zhong Shi. Admissible linear estimators in linear models with respect to inequality constraints. Linear Algebra and its Applications, 354 (1–3):187–194, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/200/27/
                                        41/abstract.html.

Langville:2004:P

[LS04a] Amy N. Langville and William J. Stewart. Preface. Linear Algebra and its Applications, 386(1):1– 2, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2004:LMT

[LS04b] Chi-Kwong Li and Nung-Sing Sze. Linear maps transforming H-unitary matrices. Linear Algebra and its Applications, 377(1):111– 124, January 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laffey:2006:NRS

[LŠ06a] Thomas J. Laffey and Helena Šmigoc. Nonnegative realization of spectra having negative real parts. Linear Algebra and its Applications, 416(1):148–159, July 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lesnjak:2006:IJS

[LS06b] Gorazd Lešnjak and Nung-Sing Sze. On injective Jordan semi-triple maps of matrix algebras. Linear Algebra and its Applications, 414(1): 383–388, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2006:DPG

[LS06c] Chi-Kwong Li and Sebastian J. Schreiber. On dispersal and population growth for multistate matrix models. Linear Algebra and its Applications, 418(2–3):900–912, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laffey:2007:CSL

[LS07a] Thomas J. Laffey and Helena Smigoc. Common solution to the Lyapunov equation for 2 × 2 complex matrices. Linear Algebra and its Applications, 420(2–3):609– 624, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laffey:2007:CNS

[LŠ07b] Thomas J. Laffey and Helena Šmigoc. Construction of nonnegative symmetric matrices with given spectrum. Linear Algebra and its Applications, 421(1):97– 109, February 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laffey:2007:TCE

[LS07c] Thomas J. Laffey and Helena Smigoc. Tensor conditions for the existence of a common solution to the Lyapunov equation. Linear Algebra and its Applications, 420(2–3):672– 685, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laffey:2008:SPS

[LŠ08a] Thomas J. Laffey and Helena Šmigoc. Spectra of principal submatrices of nonnegative matrices. Linear Algebra and its Applications, 428 (1):230–238, January 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2008:NFP

[LS08b] Yuan-Chuan Li and Mau-Hsiang Shih. The normed finiteness property of compact contraction operators. Linear Algebra and its Applications, 428(10):2319–2323, May 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laffey:2009:CLS

[LŠ09a] Thomas J. Laffey and Helena Šmigoc. Common Lyapunov solutions for two matrices whose difference has rank one. Linear Algebra and its Applications, 431(1– 2):228–240, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Landweber:2009:DHA

[LS09b] Peter S. Landweber and Eugene R. Speer. On D. Hägele’s approach to the Bessis–Moussa–Villani conjecture. Linear Algebra and its Applications, 431 (8):1317–1324, September 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lv:2009:EEE

[LS09c] Haiyan Lv and Yuming Shi. Error estimate of eigenvalues of perturbed second-order discrete Sturm– Liouville problems. Linear Algebra and its Applications, 430(8–9):2389–2415, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2009:LSR

[LSC09a] Jianxi Li, Wai Chee Shiu, and Wai Hong Chan. The Laplacian spectral radius of some graphs. Linear Algebra and its Applications, 431(1– 2):99–103, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2009:SRL

[LSC09b] Jianxi Li, Wai Chee Shiu, and Wai Hong Chan. Some results on the Laplacian eigenvalues of unicyclic graphs. Linear Algebra and its Applications, 430(8–9):2080– 2093, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2006:ILR

[LSD06] Yuan Li, Xiu-Hong Sun, and Hong-Ke Du. The intersection of left (right) spectra of 2 × 2 upper triangular operator matrices. Linear Algebra and its Applications, 418 (1):112–121, October 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2009:DMT

[LSF09] Yue Liu, Jia-Yu Shao, and Min Fang. The decomposability of the matrix with two determinantal regional components. Linear Algebra and its Applications, 431(11):2187–2194, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2008:BDR

[LSH08] Yue Liu, Jia-Yu Shao, and Chang-Xiang He. On the boundaries of the determinantal regions of ray pattern matrices. Linear Algebra and its Applications, 428(11–12): 2699–2707, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lovasz:2000:COR

[LSS00] L. Lovász, M. Saks, and A. Schrijver. A correction: orthogonal representations and connectivity of graphs. Linear Algebra and its Applications, 313(1– 3):101–105, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/130/25/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/130/25/29/article.
                                        pdf.

Li:2007:MPN

[LŠS07] Chi-Kwong Li, Peter Šemrl, and Nung-Sing Sze. Maps preserving the nilpotency of products of operators. Linear Algebra and its Applications, 424(1):222–239, July 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Leite:2008:ATI

[LST08] Ricardo S. Leite, Nicolau C. Saldanha, and Carlos Tomei. An atlas for tridiagonal isospectral manifolds. Linear Algebra and its Applications, 429(1):387–402, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Leite:2003:PPI

[LT03a] R. S. Leite and C. Tomei. Parameterization by polytopes of intersections of orbits by conjugation. Linear Algebra and its Applications, 361 (C):223–243, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
41/abstract.html.

Loewy:2003:CRC

[LT03b] Raphael Loewy and Bit-Shun Tam. CP rank of completely positive matrices of order 5. Linear Algebra and its Applications, 363 (C):161–176, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
39/abstract.html.

Ledoux:2004:CAM

[LT04a] James Ledoux and Laurent Truffet. Comparison and aggregation of max-plus linear systems. Linear Algebra and its Applications, 378(1):245– 272, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2004:SLP

[LT04b] Shwu-Huey Lin and Bit-Shun Tam. Strong linear preservers of symmetric doubly stochastic or doubly substochastic matrices. Linear Algebra and its Applications, 379(1):179– 200, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2005:OPK

[LT05] Chi-Kwong Li and Tin-Yau Tam. Operator properties of T and K(T). Linear Algebra and its Applications, 401(1): 173–191, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lebtahi:2007:NGP

[LT07a] Leila Lebtahi and Néstor Thome. A note on kgeneralized projections. Linear Algebra and its Applications, 420(2–3):572–575, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2007:PMP

[LT07b] Ming-Huat Lim and Joshua Juat-Huan Tan. Preservers of matrix pairs with bounded distance. Linear Algebra and its Applications, 422(2– 3):517–525, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2007:JIU

[LT07c] Cheng-Kai Liu and Wan-Yu Tsai. Jordan isomorphisms of upper triangular matrix rings. Linear Algebra and its Applications, 426(1):143– 148, October 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Luk:2008:ILA

[LT08] Franklin T. Luk and Daniel M. Tracy. An improved LLL algorithm. Linear Algebra and its Applications, 428(2–3): 441–452, January 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lim:2009:PPB

[LT09] Ming-Huat Lim and Joshua Juat-Huan Tan. Preservers of pairs of bivectors with bounded distance. Linear Algebra and its Applications, 430(1):564–573, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2002:LMP

[LTT02] Chi-Kwong Li, Bit-Shun Tam, and Nam-Kiu Tsing. Linear maps preserving permutation and stochastic matrices. Linear Algebra and its Applications, 341(1–3): 5–22, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
28/abstract.html.

Liu:2009:RHS

[LTT09] Yonghui Liu, Yongge Tian, and Yoshio Takane. Ranks of Hermitian and skew-Hermitian solutions to the matrix equation AXA^{∗} = B. Linear Algebra and its Applications, 431(12):2359– 2372, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2002:NRN

[LTW02] Chi-Kwong Li, Bit-Shun Tam, and Pei Yuan Wu. The numerical range of a nonnegative matrix. Linear Algebra and its Applications, 350(1– 3):1–23, July 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lu:2002:AJM

[Lu02a] Fangyan Lu. Additivity of Jordan maps on standard operator algebras. Linear Algebra and its Applications, 357 (1–3):123–131, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/203/25/
                                        33/abstract.html.

Lu:2002:MMO

[Lu02b] Fangyan Lu. Multiplicative mappings of operator algebras. Linear Algebra and its Applications, 347(1– 3):283–291, May 15, 2002.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
44/abstract.html.

Lu:2002:PCP

[Lu02c] Linzhang Lu. Perron complement and Perron root. Linear Algebra and its Applications, 341(1–3):239–248, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
45/abstract.html.

Lu:2003:JTM

[Lu03] Fangyan Lu. Jordan triple maps. Linear Algebra and its Applications, 375(1):311– 317, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lu:2009:CDJ

[Lu09] Fangyan Lu. Characterizations of derivations and Jordan derivations on Banach algebras. Linear Algebra and its Applications, 430(8– 9):2233–2239, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lur:2005:ASN

[Lur05]  Yung-Yih Lur.  On the

asymptotic stability of nonnegative matrices in max algebra. Linear Algebra and its Applications, 407(1):149–161, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lur:2006:MVG

[Lur06a] Yung-Yih Lur. A max version of the generalized spectral radius theorem. Linear Algebra and its Applications, 418 (1):336–346, October 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lur:2006:NGR

[Lur06b]  Yung-Yih Lur.  A note

on a gap result for norms of semigroups of matrices. Linear Algebra and its Applications, 419(2–3):368–372, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Luz:2007:CDL

[Luz07] Carlos J. Luz. A characterization of Delsarte’s linear programming bound as a ratio bound. Linear Algebra and its Applications, 423(1): 99–108, May 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lemmens:2003:ILN

[LV03] Bas Lemmens and Onno Van Gaans. Iteration of linear p-norm nonexpansive maps.

Linear Algebra and its Applications, 371(1):265–276, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Leibfritz:2006:ROO

[LV06] F. Leibfritz and S. Volkwein. Reduced order output feedback control design for PDE systems using proper orthogonal decomposition and nonlinear semidefinite programming. Linear Algebra and its Applications, 415(2–3):542– 575, June 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lien:2000:DGK

[LW00] Magnhild Lien and William Watkins. Dual graphs and knot invariants. Linear Algebra and its Applications, 306 (1–3):123–130, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/121/25/34/article.
pdf.

Li:2001:IPG

[LW01] Xiezhang Li and Yimin Wei. An improvement on the perturbation of the group inverse and oblique projection. Linear Algebra and its Applications, 338(1–3): 53–66, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/168/27/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/168/27/31/article.
                                        pdf.

Lockhart:2003:CAC

[LW03] Jody M. Lockhart and William P. Wardlaw. Chain addition cycles. Linear Algebra and its Applications, 362(C):1–10, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/211/25/
                                        26/abstract.html.

Li:2007:NRD

[LW07] Xiezhang Li and Yimin Wei. A note on the representations for the Drazin inverse of 2×2 block matrices. Linear Algebra and its Applications, 423 (2–3):332–338, June 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2009:ELT

[LWD09] Jie Lin, Yan Wang, and Shaoqiang Deng. T^{∗}-extension of Lie triple systems. Linear Algebra and its Applications, 431(11):2071–2083, November 1, 2009. CO-

Liu:2007:BGI

[LwW07] Ping Liu and Yu wen Wang. The best generalized inverse of the linear operator in normed linear space. Linear Algebra and its Applications, 420(1):9–19, January 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2008:CSS

[LWWZ08] Lijing Lin, Yimin Wei, Ching-Wah Woo, and Jieyong Zhou. On the convergence of splittings for semidefinite linear systems. Linear Algebra and its Applications, 429(10):2555–2566, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2005:PCR

[LWZ05] Chi-Kwong Li, Hugo Woerdeman, and Fuzhen Zhang. Preface and Conference Report. Linear Algebra and its Applications, 399(1):1– 2, April 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2009:CQC

[LWZ09] Lijing Lin, Yimin Wei, and Naimin Zhang. Convergence and quotient convergence of iterative methods for solving singular linear equations with index one. Linear Algebra and its Applications, 430(5– 6):1665–1674, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lenders:2008:FSI

[LX08] Patrick Lenders and Jingling Xue. Factorization of singular integer matrices. Linear Algebra and its Applications, 428(4):1046–1055, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lin:2004:AGB

[LXW04] Zhiping Lin, Li Xu, and Qinghe Wu. Applications of Gröbner bases to signal and image processing: a survey. Linear Algebra and its Applications, 391(1):169–202, November 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2005:COD

[LY05] Chun-Hsien Li and Suh-Yuh Yang. On a conjecture in D-optimal designs with n ≡ 0mod4. Linear Algebra and its Applications, 400(1):279– 290, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2006:BBP

[LY06] Bolian Liu and Lihua You. Bounds on the base of primitive nearly reducible sign pattern matrices. Linear Algebra and its Applications, 418(2–3):863–881, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2009:NUT

[LY09a] Hanyu Li and Hu Yang. A note on an unusual type of generalized polar decomposition. Linear Algebra and its Applications, 431(5–7):518– 526, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lur:2009:CGS

[LY09b] Yung-Yih Lur and Wen-Wei Yang. Continuity of the generalized spectral radius in max algebra. Linear Algebra and its Applications, 430(8– 9):2301–2311, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2007:EMS

[LYY07] Bolian Liu, Lihua You, and Gexin Yu. On extremal matrices of second largest exponent by Boolean rank. Linear Algebra and its Applications, 422(1):186–197, April 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2000:DNR

[LZ00] Chi-Kwong Li and Alexandru Zaharia. Decomposable numerical ranges on orthonormal tensors. Linear Algebra and its Applications, 308(1– 3):139–152, March 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/125/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/125/25/32/article.
pdf.

Li:2007:BGM

[LZ07] Xueliang Li and Jianbin Zhang. On bicyclic graphs with maximal energy. Linear Algebra and its Applications, 427(1):87–98, November 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laszkiewicz:2008:AMR

[LZ08a] Beata Laszkiewicz and Krystyna Zieţak. Approximation by matrices with restricted spectra. Linear Algebra and its Applications, 428(4):1031– 1040, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2008:APS

[LZ08b] Tsiu-Kwen Lee and Yiqiang Zhou. Algebraic prime subalgebras in simple Artinian algebras. Linear Algebra and its Applications, 428(4):881– 889, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lee:2009:RIG

[LZ09a] Tsiu-Kwen Lee and Yiqiang Zhou. Right ideals generated by an idempotent of finite rank. Linear Algebra and its Applications, 431(11):2118– 2126, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2009:PAM

[LZ09b] Jing Li and Yuhai Zhang. Perturbation analysis of the matrix equation X − A^{∗}X^{−p}A = Q. Linear Algebra and its Applications, 431(9):1489–1501, October 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Li:2009:RCG

[LZ09c] Ren-Cang Li and Wei Zhang. The rate of convergence of GMRES on a tridiagonal Toeplitz linear system. II. Linear Algebra and its Applications, 431(12):2425–2436, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liu:2009:LEU

[LZS09] Ruifang Liu, Mingqing Zhai, and Jinlong Shu. The least eigenvalue of unicyclic graphs with n vertices and k pendant vertices. Linear Algebra and its Applications, 431(5– 7):657–665, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Lu:2007:LBL

[LzZT07] Mei Lu, Lian zhu Zhang, and Feng Tian. Lower bounds of the Laplacian spectrum of graphs based on diameter. Linear Algebra and its Applications, 420(2–3):400–406, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mahony:2003:CTR

[MA03] R. Mahony and P.-A. Absil. The continuous-time Rayleigh quotient flow on the sphere. Linear Algebra and its Applications, 368 (C):343–357, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
47/abstract.html.

Martinez-Avendano:2006:EOC

[MA06] Rubén A. Martínez-Avendaño. Eigenmatrices and operators commuting with finiterank operators. Linear Algebra and its Applications, 419(2–3):739–749, December 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mayo:2007:FSG

[MA07] A. J. Mayo and A. C. Antoulas. A framework for the solution of the generalized realization problem. Linear Algebra and its Applications, 425(2–3):634–662, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ma:2009:BLB

[Ma09] Hongping Ma. Bounds on the local bases of primitive, non-powerful, minimally strong signed digraphs. Linear Algebra and its Applications, 430(2–3):718–731, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Maesumi:2008:ONC

[Mae08] Mohsen Maesumi. Optimal norms and the computation of joint spectral radius of matrices. Linear Algebra and its Applications, 428(10):

2324–2338, May 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Maher:2007:SNI

[Mah07] P. J. Maher. Some norm inequalities concerning generalized inverses, 2. Linear Algebra and its Applications, 420(2–3):517–525, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mailybaev:2001:TVD

[Mai01] Alexei A. Mailybaev. Transformation to versal deformations of matrices. Linear Algebra and its Applications, 337(1–3):87–108, November 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/167/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/167/27/31/article.
pdf.

Malamud:2001:CJI

[Mal01a] S. M. Malamud. A converse to the Jensen inequality, its matrix extensions and inequalities for minors and eigenvalues. Linear Algebra and its Applications, 322(1– 3):19–41, January 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/141/24/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/141/24/25/article.
pdf.

Mallik:2001:ITM

[Mal01b] Ranjan K. Mallik. The inverse of a tridiagonal matrix. Linear Algebra and its Applications, 325(1–3): 109–139, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/147/24/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/147/24/32/article.
pdf.

Malejki:2009:ALE

[Mal09] Maria Malejki. Asymptotics of large eigenvalues for some discrete unbounded Jacobi matrices. Linear Algebra and its Applications, 431(10): 1952–1970, October 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Maponi:2007:SLS

[Map07] P. Maponi. The solution of linear systems by using the Sherman–Morrison formula. Linear Algebra and its Applications, 420(2–3):276–294, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Markiewicz:2002:EEC

[Mar02] Augustyn Markiewicz. Estimation and experiments comparison with respect to the matrix risk. Linear Algebra and its Applications, 354 (1–3):213–222, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/200/27/
                                        43/abstract.html.

Maritz:2005:STM

[Mar05] Pieter Maritz. Sir Thomas Muir, 1844–1934. Linear Algebra and its Applications, 411(1):3–67, December 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Massey:2009:ORS

[Mas09] Pedro G. Massey. Optimal reconstruction systems for erasures and for the q-potential. Linear Algebra and its Applications, 431(8):1302–1316, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Matache:2001:NRC

[Mat01] Valentin Matache. Numerical ranges of composition operators. Linear Algebra and its Applications, 331 (1–3):61–74, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/30/article.
pdf.

Mateja:2002:CSI

[Mat02] J. Mateja. Convergence of scaled iterates by the Jacobi method. Linear Algebra and its Applications, 349(1– 3):17–53, July 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Matsuura:2003:CST

[Mat03]  Masaya Matsuura.  On

the Craig-Sakamoto theorem and Olkin’s determinantal result. Linear Algebra and its Applications, 364 (C):321–323, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
43/abstract.html.

Mattina:2004:GIC

[Mat04] Daniela La Mattina. On the graded identities and cocharacters of the algebra of 3 × 3 matrices. Linear Algebra and its Applications, 384(1): 55–75, June 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mathai:2005:PMV

[Mat05a] A. M. Mathai. A pathway to matrix-variate gamma and normal densities. Linear Algebra and its Applications, 396(1):317–328, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Matolcsi:2005:GEN

[Mat05b] Máté Matolcsi. A geometric estimate on the norm of product of functionals. Linear Algebra and its Applications, 405(1):304–310, August 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Matsumoto:2005:PPP

[Mat05c] Sho Matsumoto. α-Pfaffian, Pfaffian point process and shifted Schur measure. Linear Algebra and its Applications, 403(1):369–398, July 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mathai:2007:RVU

[Mat07] A. M. Mathai. Random volumes under a general matrixvariate model. Linear Algebra and its Applications, 425 (1):162–170, August 1, 2007. CODEN LAAPAW. ISSN

Malik:2007:SOP

[MB07] Saroj Malik and P. Bhimasankaram. Shorted operators of partitioned matrices and application. Linear Algebra and its Applications, 425 (1):150–161, August 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mirman:2001:NRP

[MBLV01] Boris Mirman, Vladimir Borovikov, Lev Ladyzhensky, and Robert Vinograd. Numerical ranges, Poncelet curves, invariant measures. Linear Algebra and its Applications, 329(1–3):61–75, May 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/154/25/28/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/154/
25/28/article.pdf.

McCullough:2001:FOV

[McC01] Scott McCullough. Factorization of operator-valued polynomials in several noncommuting variables. Linear Algebra and its Applications, 326(1–3):193–203, March 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/148/24/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/148/24/36/article.
                                        pdf.

McCluskey:2005:SCL

[McC05] C. Connell McCluskey. A strategy for constructing Lyapunov functions for nonautonomous linear differential equations. Linear Algebra and its Applications, 409(1): 100–110, November 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McDonald:2003:PSN

[McD03] Judith J. McDonald. The peripheral spectrum of a nonnegative matrix. Linear Algebra and its Applications, 363 (C):217–235, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/212/25/
                                        42/abstract.html.

McEachin:2002:CBS

[McE02] R. McEachin. A connection between Schur multiplication and Fourier interpolation. Linear Algebra and its Applications, 341(1–3): 379–389, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/

gej-ng/10/30/19/171/27/
58/abstract.html.

Martin:2006:EED

[MCEC06] R. J. Martin, N. Chauhan, J. A. Eccleston, and B. S. P. Chan. Efficient experimental designs when most treatments are unreplicated. Linear Algebra and its Applications, 417(1):163–182, August 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McGillivray:2006:PAS

[McG06] Barbara McGillivray. A probabilistic algorithm for the secant defect of Grassmann varieties. Linear Algebra and its Applications, 418(2–3):708– 718, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McMillen:2009:EDB

[McM09] Tyler McMillen. On the eigenvalues of double band matrices. Linear Algebra and its Applications, 431(10): 1890–1897, October 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marin:2002:LTC

[MD02] Jean-Michel Marin and Thierry Dhorne. Linear Toeplitz covariance structure models with optimal estimators of variance components. Linear Algebra and its Applications, 354(1–3):195–212, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
42/abstract.html.

Mosic:2009:MPI

[MD09] Dijana Mosić and Dragan S. Djordjević. Moore–Penroseinvertible normal and Hermitian elements in rings. Linear Algebra and its Applications, 431(5–7):732–745, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mosic:2009:EER

[MDK09] Dijana Mosić, Dragan S. Djordjević, and J. J. Koliha. EP elements in rings. Linear Algebra and its Applications, 431(5–7):527–535, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mehl:2000:ATA

[Meh00]  Christian Mehl.  Anti-

triangular and anti- m - Hessenberg forms for Hermitian matrices and pencils. Linear Algebra and its Applications, 317(1–3):143– 176, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/136/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/136/25/31/article.
pdf.

Mehrmann:2008:RB

[Meh08] Volker Mehrmann. Ralph Byers 1955–2007. Linear Algebra and its Applications, 428(11–12):2410– 2414, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Meijer:2005:MAH

[Mei05] Erik Meijer. Matrix algebra for higher order moments. Linear Algebra and its Applications, 410(1):112– 134, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Meini:2006:NME

[Mei06] Beatrice Meini. Nonlinear matrix equations and structured linear algebra. Linear Algebra and its Applications, 413(2–3):440–457, March 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Melman:2000:SCM

[Mel00] A. Melman. Symmetric centrosymmetric matrix-vector multiplication. Linear Algebra and its Applications, 320 (1–3):193–198, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/38/article.
pdf.

Melman:2001:TSE

[Mel01] A. Melman. A two-step even-odd split Levinson algorithm for Toeplitz systems. Linear Algebra and its Applications, 338(1–3): 219–237, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
42/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/42/article.
pdf.

Melman:2009:SIS

[Mel09] A. Melman. Spectral inclusion sets for structured matrices. Linear Algebra and its Applications, 431(5–7):633– 656, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Merris:2005:IIG

[Mer05] Russell Merris. Immanantal invariants of graphs. Linear Algebra and its Applications, 401(1):67–75, May 15, 2005.

Matsumoto:2009:GRB

[MF09] Akemi Matsumoto and Masatoshi Fujii. Generalizations of reverse Bebiano–Lemos–Providência inequality. Linear Algebra and its Applications, 430(5– 6):1544–1549, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Munoz-Fernandez:2008:LSM

[MFPPSS08] G. A. Muñoz-Fernández, N. Palmberg, D. Puglisi, and J. B. Seoane-Sepúlveda. Lineability in subsets of measure and function spaces. Linear Algebra and its Applications, 428(11–12):2805– 2812, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Miranian:2003:SRR

[MG03] L. Miranian and M. Gu. Strong rank revealing LU factorizations. Linear Algebra and its Applications, 367(C):1–16, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
26/abstract.html.

Moldovan:2009:SDD

[MG09] Melania M. Moldovan and M. Seetharama Gowda. Strict diagonal dominance and a Geršgorin type theorem in Euclidean Jordan algebras. Linear Algebra and its Applications, 431(1–2): 148–161, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Michiels:2006:PSR

[MGWN06] Wim Michiels, Kirk Green, Thomas Wagenknecht, and Silviu-Iulian Niculescu. Pseudospectra and stability radii for analytic matrix functions with application to timedelay systems. Linear Algebra and its Applications, 418 (1):315–335, October 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Matejas:2006:QCE

[MH06] J. Matejaš and V. Hari. Quadratic convergence estimate of scaled iterates by J-symmetric Jacobi method. Linear Algebra and its Applications, 417(2–3):434–465, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Manthey:2007:BCF

[MH07] Wilfried Manthey and Uwe Helmke. Bruhat canonical form for linear systems.

Linear Algebra and its Applications, 425(2–3):261–282, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Meng:2005:SSO

[MHZ05] Chunjun Meng, Xiyan Hu, and Lei Zhang. The skewsymmetric orthogonal solutions of the matrix equation AX = B. Linear Algebra and its Applications, 402(1): 303–318, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Meng:2005:COP

[MHZU05] Chunjun Meng, Xiyan Hu, Lei Zhang, and Frank Uhlig. Completion of an orthogonal projector and a related inverse mapping problem. Linear Algebra and its Applications, 403(1):229– 247, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Miklavic:2009:TAD

[Mik09] Štefko Miklavič. The Terwilliger algebra of a distanceregular graph of negative type. Linear Algebra and its Applications, 430(1):251– 270, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mills:2003:ISC

[Mil03] Mark A. Mills. The intersection of the similarity and conjunctivity equivalence classes. Linear Algebra and its Applications, 362 (C):129–136, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/211/25/
35/abstract.html.

Minoux:2001:EMM

[Min01] M. Minoux. Extension of MacMahon’s Master Theorem to pre-semi-rings. Linear Algebra and its Applications, 338(1–3):19–26, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/28/article.
pdf.

Miranda:2000:SVD

[Mir00] Hector F. Miranda. Singular values, diagonal elements, and extreme matrices. Linear Algebra and its Applications, 305(1–3): 151–159, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/120/25/
35/abstract.html;   http:

Mirman:2003:UMC

[Mir03] Boris Mirman. UB-matrices and conditions for Poncelet polygon to be closed. Linear Algebra and its Applications, 360(C):123–150, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
33/abstract.html.

Merikoski:2003:CLB

[MK03] Jorma Kaarlo Merikoski and Ravinder Kumar. Characterizations and lower bounds for the spread of a normal matrix. Linear Algebra and its Applications, 364(C):13–31, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
25/abstract.html.

Merikoski:2005:UBS

[MK05a] Jorma K. Merikoski and Ravinder Kumar. Upper bounds for singular values. Linear Algebra and its Applications, 401(1):371– 379, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Merikoski:2005:LBN

[MK05b] Jorma Kaarlo Merikoski and Ravinder Kumar. Lower bounds for the numerical radius. Linear Algebra and its Applications, 410(1):135–142, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Miao:2002:IAD

[ML02] Zhengke Miao and Jiongsheng Li. Inertially arbitrary (2r − 1)-diagonal sign patterns. Linear Algebra and its Applications, 357(1–3): 133–141, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/203/25/
                                        34/abstract.html.

Mastronardi:2004:IRS

[MLK^{+}04] N. Mastronardi, P. Lemmerling, A. Kalsi, D. P. O’Leary, and S. Van Huffel. Implementation of the regularized structured total least squares algorithms for blind image deblurring. Linear Algebra and its Applications, 391(1): 203–221, November 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mata-Lorenzo:2000:CP

[MLR00] Luis E. Mata-Lorenzo and Lázaro Recht. Convexity properties of Tr[(a^{∗}a)^{n}].

Linear Algebra and its Applications, 315(1–3):25–38, August 15, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/132/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/132/25/26/article.
pdf.

Marek:2003:CTS

[MM03] Ivo Marek and Petr Mayer. Convergence theory of some classes of iterative aggregation/disaggregation methods for computing stationary probability vectors of stochastic matrices. Linear Algebra and its Applications, 363(C):177–200, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
40/abstract.html.

Malik:2005:PQR

[MM05] Madhur Malik and S. R. Mohan. On Q and R_{0} properties of a quadratic representation in linear complementarity problems over the secondorder cone. Linear Algebra and its Applications, 397(1): 85–97, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mangasarian:2006:AVE

[MM06a] O. L. Mangasarian and R. R. Meyer. Absolute value equations. Linear Algebra and its Applications, 419(2–3):359– 367, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McAsey:2006:GRE

[MM06b] Michael McAsey and Libin Mou. Generalized Riccati equations arising in stochastic games. Linear Algebra and its Applications, 416(2– 3):710–723, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marco:2007:FAA

[MM07] Ana Marco and José-Javier Martínez. A fast and accurate algorithm for solving Bernstein–Vandermonde linear systems. Linear Algebra and its Applications, 422(2– 3):616–628, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McDonald:2008:LCC

[MM08] Judith J. McDonald and DeAnne M. Morris. Level characteristics corresponding to peripheral eigenvalues of a nonnegative matrix. Linear Algebra and its Applications, 429(7):1719– 1729, October 1, 2008. CO-

Mackey:2004:RAH

[MMT04] D. Steven Mackey, Niloufer Mackey, and Françoise Tisseur. G-reflectors: analogues of Householder transformations in scalar product spaces. Linear Algebra and its Applications, 385(1):187– 213, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mehl:2004:DSM

[MMX04] Christian Mehl, Volker Mehrmann, and Hongguo Xu. On doubly structured matrices and pencils that arise in linear response theory. Linear Algebra and its Applications, 380(1): 3–51, March 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Molitierno:2003:TPM

[MN03] Jason J. Molitierno and Michael Neumann. On trees with perfect matchings. Linear Algebra and its Applications, 362(C):75–85, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/211/25/
30/abstract.html.

Morris:2007:GHM

[MN07] Walter D. Morris Jr. and Makoto Namiki. Good hidden P -matrix sandwiches. Linear Algebra and its Applications, 426(2–3):325–341, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mense:2008:AML

[MN08] C. Mense and R. Nabben. On algebraic multi-level methods for non-symmetric systems — comparison results. Linear Algebra and its Applications, 429(10):2567–2588, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mohan:2001:MPS

[MND01a] S. R. Mohan, S. K. Neogy, and A. K. Das. More on positive subdefinite matrices and the linear complementarity problem. Linear Algebra and its Applications, 338(1–3): 275–285, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/168/27/
                                        47/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/168/27/47/article.
                                        pdf.

Mohan:2001:CFC

[MND01b] S. R. Mohan, S. K. Neogy, and A. K. Das. On the classes of fully copositive and fully semimonotone matrices. Linear Algebra and its Applications, 323(1– 3):87–97, January 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

MacGillivray:2008:PDS

[MNOvdD08] G. MacGillivray, S. Nasserasr, D. D. Olesky, and P. van den Driessche. Primitive digraphs with smallest large exponent. Linear Algebra and its Applications, 428(7):1740– 1752, April 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mehrmann:2008:GPF

[MNV08] Volker Mehrmann, Reinhard Nabben, and Elena Virnik. Generalisation of the Perron– Frobenius theory to matrix pencils. Linear Algebra and its Applications, 428(1):20– 38, January 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mohan:2005:BRG

[Moh05] S. R. Mohan. Book review: Guorong Wang, Yimin Wei and Sanzheng Qiao, Review of Generalized Inverses: Theory and Computations, Graduate Series in Mathematics vol. 5, Science Press, Beijing (2004) ISBN-10 7-03-012437-5, ISBN-13 978-7-03-012437-1. Linear Algebra and its Applications, 406(1):310– 312, September 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mohar:2007:LCA

[Moh07] Bojan Mohar. On the Laplacian coefficients of acyclic graphs. Linear Algebra and its Applications, 422(2–3): 736–741, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Molnar:2002:CMM

[Mol02] Lajos Molnár. Conditionally multiplicative maps on the set of all bounded observables preserving compatibility. Linear Algebra and its Applications, 349(1–3):197– 201, July 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Molnar:2003:PHS

[Mol03] Lajos Molnár. Preservers on Hilbert space effects. Linear Algebra and its Applications, 370(1):287–300, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Molteni:2004:ATT

[Mol04] G. Molteni. About two trigonometric matrices. Linear Algebra and its Applications, 382(1):39–59, May 1, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Molitierno:2005:SRS

[Mol05a] Jason J. Molitierno. The spectral radius of submatrices of Laplacian matrices for trees and its comparison to the Fiedler vector. Linear Algebra and its Applications, 406(1):253–271, September 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Molnarova:2005:GMP

[Mol05b] Monika Molnárová. Generalized matrix period in maxplus algebra. Linear Algebra and its Applications, 404(1): 345–366, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Molitierno:2006:ACG

[Mol06a] Jason J. Molitierno. On the algebraic connectivity of graphs as a function of genus. Linear Algebra and its Applications, 419(2–3):519–531, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Molnar:2006:MJT

[Mol06b] Lajos Molnár. Multiplicative Jordan triple isomorphisms on the self-adjoint elements of von Neumann algebras. Linear Algebra and its Applications, 419(2–3):586–600, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Molinari:2008:DBT

[Mol08a] Luca Guido Molinari. Determinants of block tridiagonal matrices. Linear Algebra and its Applications, 429(8–9): 2221–2226, October 16, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Molitierno:2008:SRS

[Mol08b] Jason J. Molitierno. The spectral radius of submatrices of Laplacian matrices for graphs with cut vertices. Linear Algebra and its Applications, 428(8–9):1987– 1999, April 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Molnar:2009:MPH

[Mol09] Lajos Molnár. Maps preserving the harmonic mean or the parallel sum of positive operators. Linear Algebra and its Applications, 430(11–12): 3058–3065, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-

Monov:2008:FSP

[Mon08] Vladimir V. Monov. A family of symmetric polynomials of the eigenvalues of a matrix. Linear Algebra and its Applications, 429(8–9):2199– 2208, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Moraca:2008:BNM

[Mor08] Nenad Morača. Bounds for norms of the matrix inverse and the smallest singular value. Linear Algebra and its Applications, 429(10):2589– 2601, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Moision:2007:CLJ

[MOS07] Bruce Moision, Alon Orlitsky, and Paul H. Siegel. On codes with local joint constraints. Linear Algebra and its Applications, 422(2– 3):442–454, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Moslehian:2009:OEH

[Mos09] M. S. Moslehian. Operator extensions of Hua’s inequality. Linear Algebra and its Applications, 430(4): 1131–1139, February 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McCarthy:2002:BTC

[MOSZ02] Nicholas McCarthy, David Ogilvie, Ilya Spitkovsky, and Nahum Zobin. Birkhoff’s theorem and convex hulls of Coxeter groups. Linear Algebra and its Applications, 347 (1–3):219–231, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
40/abstract.html.

Mourad:2006:NBS

[Mou06] Bassam Mourad. A note on the boundary of the set where the decreasingly ordered spectra of symmetric doubly stochastic matrices lie. Linear Algebra and its Applications, 416(2–3):546– 558, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Maroulas:2000:FMP

[MP00] J. Maroulas and P. Psarrakos. On factorization of matrix polynomials. Linear Algebra and its Applications, 304(1–3):131–139, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/118/25/
32/abstract.html;   http:

Mathai:2005:SCM

[MP05] A. M. Mathai and Serge B. Provost. Some complex matrix-variate statistical distributions on rectangular matrices. Linear Algebra and its Applications, 410(1):198– 216, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marek:2006:NLG

[MP06] Ivo Marek and Ivana Pultarová. A note on local and global convergence analysis of iterative aggregation– disaggregation methods. Linear Algebra and its Applications, 413(2–3):327–341, March 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Matkovic:2006:VJI

[MPP06] A. Matković, J. Pečarić, and I. Perić. A variant of Jensen’s inequality of Mercer’s type for operators with applications. Linear Algebra and its Applications, 418(2–3):551– 564, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Merino:2010:PDM

[MPP10] Dennis I. Merino, Agnes T. Paras, and Diane Christine P. Pelejo. On the ϕ_{J} polar decomposition of matrices. Linear Algebra and its Applications, 432(5):1165–1175, February 15, 2010. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See related work [DoMP09].

Micic:2000:CII

[MPS00] Jadranka Mićić, Josip Pečarić, and Yuki Seo. Complementary inequalities to inequalities of Jensen and Ando based on the Mond-Pečarić method. Linear Algebra and its Applications, 318(1–3): 87–107, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/137/24/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/137/24/30/article.
pdf.

Micic:2003:FOP

[MPS03] Jadranka Mićić, Josip Pečarić, and Yuki Seo. Function order of positive operators based on the Mond-Pečarić method. Linear Algebra and its Applications, 360(C): 15–34, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

Malinowska:2006:ELS

[MPS06] I. Malinowska, P. Pawlas, and D. Szynal. Estimation of location and scale parameters for the Burr XII distribution using generalized order statistics. Linear Algebra and its Applications, 417(1):150– 162, August 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marijuan:2007:MSC

[MPS07] Carlos Marijuán, Miriam Pisonero, and Ricardo L. Soto. A map of sufficient conditions for the real nonnegative inverse eigenvalue problem. Linear Algebra and its Applications, 426(2–3):690– 705, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Maroulas:2002:PFT

[MPT02] J. Maroulas, P. J. Psarrakos, and M. J. Tsatsomeros. Perron–Frobenius type results on the numerical range. Linear Algebra and its Applications, 348 (1–3):49–62, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/

                                        gej-ng/10/30/19/195/27/
                                        31/abstract.html.

Magajna:2004:OIS

[MPT04] Bojan Magajna, Marko Petkovšek, and Aleksej Turnšek. An operator inequality and self-adjointness. Linear Algebra and its Applications, 377(1):181–194, January 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Martins:2003:WHU

[MQ03] Rosa Amélia Martins and João Filipe Queiró. 2-widths of the Hölder unit balls. Linear Algebra and its Applications, 361(C):245–255, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/210/25/
                                        42/abstract.html.

Michael:2006:OSN

[MQ06] T. S. Michael and Thomas Quint. Optimal strategies for node selection games on oriented graphs: Skew matrices and symmetric games. Linear Algebra and its Applications, 412(2–3):77–92, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mehl:2001:CNM

[MR01] Christian Mehl and Leiba Rodman. Classes of normal matrices in indefinite inner products. Linear Algebra and its Applications, 336(1– 3):71–98, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/166/27/33/article.
pdf.

Mehl:2002:SMR

[MR02a] Christian Mehl and Leiba Rodman. Symmetric matrices with respect to sesquilinear forms. Linear Algebra and its Applications, 349(1– 3):55–75, July 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Moricz:2002:ERS

[MR02b] Ferenc Móricz and Billy E. Rhoades. An equivalent reformulation of summability by weighted mean methods, revisited. Linear Algebra and its Applications, 349(1– 3):187–192, July 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mullhaupt:2002:LGM

[MR02c] Andrew P. Mullhaupt and Kurt S. Riedel. Low grade matrices and matrix fraction representations. Linear Algebra and its Applications, 342 (1–3):187–201, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/173/27/
                                        39/abstract.html.

Mastnak:2009:SFM

[MR09] Mitja Mastnak and Heydar Radjavi. Structure of finite, minimal nonabelian groups and triangularization. Linear Algebra and its Applications, 430(7):1838– 1848, April 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mehl:2007:EMS

[MRR07] Christian Mehl, André C. M. Ran, and Leiba Rodman. Extension to maximal semidefinite invariant subspaces for hyponormal matrices in indefinite inner products. Linear Algebra and its Applications, 421(1):110–116, February 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marek:2000:CTC

[MS00a] Ivo Marek and Daniel B. Szyld. Comparison theorems for the convergence factor of iterative methods for singular matrices. Linear Algebra and its Applications, 316(1– 3):67–87, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/29/article.
pdf.

Matos:2000:CPF

[MS00b] Isabel T. Matos and Fernando C. Silva. A completion problem over the field of real numbers. Linear Algebra and its Applications, 320(1– 3):63–77, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/29/article.
pdf.

Marcellan:2001:MAT

[MS01a] Francisco Marcellán and Franciszek Hugon Szafraniec. A matrix algorithm towards solving the moment problem of Sobolev type. Linear Algebra and its Applications, 331 (1–3):155–164, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/37/article.
pdf.

Mizuno:2001:ZFD

[MS01b] Hirobumi Mizuno and Iwao Sato. Zeta functions of digraphs. Linear Algebra and its Applications, 336(1–3): 181–190, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/166/27/
                                        38/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/166/27/38/article.
                                        pdf.

Mizuno:2002:WZF

[MS02a] Hirobumi Mizuno and Iwao Sato. Weighted zeta functions of digraphs. Linear Algebra and its Applications, 355(1–3):35–48, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/201/27/
                                        30/abstract.html.

Murthy:2002:PDS

[MS02b] N. L. Bhanu Murthy and Murali K. Srinivasan. The polytope of degree sequences of hypergraphs. Linear Algebra and its Applications, 350 (1–3):147–170, July 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Malyshev:2003:CNL

[MS03a] A. Malyshev and M. Sadkane. Condition numbers for Lanczos bidiagonalization with complete reorthogonalization. Linear Algebra and its Applications, 371(1):315– 331, September 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Martins:2003:EJP

[MS03b] Enide Andrade Martins and Fernando C. Silva. On the eigenvalues of Jordan products. Linear Algebra and its Applications, 359(1–3): 249–262, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
44/abstract.html.

Martins:2003:IPJ

[MS03c] Enide Andrade Martins and Fernando C. Silva. On the invariant polynomials of Jordan products. Linear Algebra and its Applications, 360(C): 173–189, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
35/abstract.html.

Malyshev:2004:CPM

[MS04a] A. N. Malyshev and M. Sadkane. Componentwise pseudospectrum of a matrix. Linear Algebra and its Applications, 378(1):283–288, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marek:2004:ASM

[MS04b] Ivo Marek and Daniel B. Szyld. Algebraic Schwarz methods for the numerical solution of Markov chains. Linear Algebra and its Applications, 386(1):67–81, July 15, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Massey:2004:GSC

[MS04c] Pedro Massey and Demetrio Stojanoff. Generalized Schur complements and P - complementable operators. Linear Algebra and its Applications, 393(1):299–318, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mencherini:2004:USM

[MS04d] Lorenzo Mencherini and Sergio Spagnolo. Uniformly symmetrizable 3 × 3 matrices. Linear Algebra and its Applications, 382(1):25– 38, May 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Meshulam:2004:LLD

[MŠ04e] Roy Meshulam and Peter Šemrl. Locally linearly dependent operators and reflexivity of operator spaces. Linear Algebra and its Applications, 383(1):143–150, May 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mirman:2005:CCP

[MS05] Boris Mirman and Pradeep Shukla. A characterization of complex plane Poncelet curves. Linear Algebra and its Applications, 408(1):86– 119, October 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mason:2006:SDS

[MS06] Oliver Mason and Robert Shorten. On the simultaneous diagonal stability of a pair of positive linear systems. Linear Algebra and its Applications, 413(1):13– 23, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mizuno:2007:NBZ

[MS07] Hirobumi Mizuno and Iwao Sato. A new Bartholdi zeta function of a digraph. Linear Algebra and its Applications, 423(2–3):498–511, June 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McDonald:2008:SAR

[MS08] Judith J. McDonald and Jeffrey Stuart. Spectrally arbitrary ray patterns. Linear Algebra and its Applications, 429(4):727–734, August 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Meerbach:2005:EBR

[MSF05] Eike Meerbach, Christof Schütte, and Alexander Fischer. Eigenvalue bounds on restrictions of reversible nearly uncoupled Markov chains. Linear Algebra and its Applications, 398(1):141– 160, March 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Martins:2006:BFG

[MSR06] N. Martins, R. Severino, and J. Sousa Ramos. Bowen– Franks groups of reducible bimodal subshifts of finite type. Linear Algebra and its Applications, 414(1):125– 137, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mason:2007:KYP

[MSS07] Oliver Mason, Robert Shorten, and Selim Solmaz. On the Kalman–Yacubovich–Popov lemma and common Lyapunov solutions for matrices with regular inertia. Linear Algebra and its Applications, 420(1):183–197, January 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Moghaddamfar:2008:DMR

[MSS08] A. R. Moghaddamfar, S. Navid Salehy, and S. Nima Salehy. The determinants of matrices with recursive entries. Linear Algebra and its Applications, 428(11–12):2468– 2481, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Misiurewicz:2006:ILM

[MST06] Michał Misiurewicz, John G. Stevens, and Diana M. Thomas. Iterations of linear maps over finite fields. Linear Algebra and its Applications, 413(1):218–234, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mehrmann:2009:NBM

[MSW09] V. Mehrmann, C. Schröder, and D. S. Watkins. A new block method for computing the Hamiltonian Schur form. Linear Algebra and its Applications, 431(3–4):350– 368, July 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Martin:2002:FSL

[MT02] Heath M. Martin and Alexander Tovbis. On formal solutions of linear matrix differential-difference equations. Linear Algebra and its Applications, 345(1–3): 29–42, April 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/175/27/
                                        29/abstract.html.

Mazorchuk:2004:RDF

[MT04] Volodymyr Mazorchuk and Lyudmila Turowska. Radical ∗-doubles of finitedimensional algebras. Linear Algebra and its Applications, 390(1):293–309, October 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Masri:2005:NER

[MT05] Ibrahim Masri and Andrew Tonge. Norm estimates for random multilinear Hankel forms. Linear Algebra and its Applications, 402(1):255– 262, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Molnar:2006:TSS

[MT06] Lajos Molnár and Werner Timmermann. Transformations on the sets of states and density operators. Linear Algebra and its Applications, 418(1):75–84, October 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

MacGillivray:2005:SAS

[MTvdD05] G. MacGillivray, R. M. Tifenbach, and P. van den Driessche. Spectrally arbitrary star sign patterns. Linear Algebra and its Applications, 400(1): 99–119, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mueller:2009:NFL

[Mue09] Markus Mueller. Normal form for linear systems with respect to its vector relative degree. Linear Algebra and its Applications, 430(4): 1292–1312, February 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Muller:2003:NAV

[Mül03] Jürgen Müller. A note on applications of the ‘Vector Enumerator’ algorithm. Linear Algebra and its Applications, 365(C):291–300, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
30/abstract.html.

Muldoon:2005:GHF

[Mul05] Martin E. Muldoon. Generalized hyperbolic functions, circulant matrices and functional equations. Linear Algebra and its Applications, 406(1):272–284, September 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Merikoski:2004:BPL

[MV04] Jorma Kaarlo Merikoski and Ari Virtanen. The best possible lower bound for the Perron root using traces. Linear Algebra and its Applications, 388(1):301–313, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Markovsky:2007:LVR

[MV07] Ivan Markovsky and Sabine Van Huffel. Left vs right representations for solving weighted low-rank approximation problems. Linear Algebra and its Applications, 422(2–3):540–552, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mastronardi:2008:SBA

[MVV08] N. Mastronardi, M. Van Barel, and R. Vandebril. A Schur-based algorithm for computing bounds to the smallest eigenvalue of a symmetric positive definite Toeplitz matrix. Linear Algebra and its Applications, 428(2–3):479–491, January 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ming:2001:SRT

[MW01a] Guo Ji Ming and Tan Shang Wang. On the spectral radius of trees. Linear Algebra and its Applications, 329(1– 3):1–8, May 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/154/25/25/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/154/
25/25/article.pdf.

Ming:2001:RBM

[MW01b] Guo Ji Ming and Tan Shang Wang. A relation between the matching number and Laplacian spectrum of a graph. Linear Algebra and its Applications, 325 (1–3):71–74, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/147/24/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/147/24/28/article.
pdf.

Mayer:2003:FPI

[MW03a] Günter Mayer and Ingo Warnke. On the fixed points of the interval function [f]([x]) = [A][x] + [b]. Linear Algebra and its Applications, 363(C):201–216, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/212/25/
                                        41/abstract.html.

McKay:2003:AEM

[MW03b] Brendan D. McKay and Xiaoji Wang. Asymptotic enumeration of 0–1 matrices with equal row sums and equal column sums. Linear Algebra and its Applications, 373(1): 273–287, November 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Masaro:2008:RODa

[MW08a] Joe Masaro and Chi Song Wong. Robustness of Aoptimal designs. Linear Algebra and its Applications, 429(7):1392–1408, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Masaro:2008:RODb

[MW08b] Joe Masaro and Chi Song Wong. Robustness of optimal designs for correlated random variables. Linear Algebra and its Applications, 429(7): 1639–1646, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

McGuire:2009:CHM

[MW09] Gary McGuire and Harold N. Ward. Cocyclic Hadamard matrices from forms over finite Frobenius rings. Linear Algebra and its Applications, 430(7):1730–1738, April 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Myskova:2005:ISM

[Myš05] Helena Myšková. Interval systems of max-separable linear equations. Linear Algebra and its Applications, 403(1): 263–272, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Myskova:2006:CSI

[Myš06] Helena Myšková. Control solvability of interval systems of max-separable linear equations. Linear Algebra and its Applications, 416 (2–3):215–223, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Myung:2009:TMN

[Myu09] Sung Myung. Transfer maps and nonexistence of joint determinant. Linear Algebra and its Applications, 431(9): 1633–1651, October 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Miao:2000:LES

[MZ00] Zhengke Miao and Kemin Zhang. The local exponent sets of primitive digraphs. Linear Algebra and its Applications, 307(1– 3):15–33, March 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/26/article.
pdf.

Medina:2004:FSP

[MZ04] Gonzalo Medina and Alexander Zavadskij. The four subspace problem: An elementary solution. Linear Algebra and its Applications, 392(1): 11–23, November 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

MarquesdeSa:2005:NIF

[MZ05] E. Marques de Sá and Yu Lin Zhang. On the number of invariant factors of matrix products. Linear Algebra and its Applications, 401(1):393– 399, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-

Morgan:2006:HRA

[MZ06] Ronald B. Morgan and Min Zeng. A harmonic restarted Arnoldi algorithm for calculating eigenvalues and determining multiplicity. Linear Algebra and its Applications, 415(1):96–113, May 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Moslehian:2008:OEI

[MZ08] Mohammad Sal Moslehian and Fuzhen Zhang. An operator equality involving a continuous field of operators and its norm inequalities. Linear Algebra and its Applications, 429(8–9):2159–2167, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Mondie:2000:SFL

[MZK00] S. Mondié, P. Zagalak, and V. Kučera. State feedback in linear control theory. Linear Algebra and its Applications, 317(1–3):177– 192, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/136/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/136/25/32/article.
                                        pdf.

Nabben:2007:RBS

[Nab07] Reinhard Nabben. On relationships between several classes of Z-matrices, Mmatrices and nonnegative matrices. Linear Algebra and its Applications, 421(2–3): 417–439, March 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nahtman:2006:MPI

[Nah06] Tatjana Nahtman. Marginal permutation invariant covariance matrices with applications to linear models. Linear Algebra and its Applications, 417(1):183– 210, August 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nakazi:2000:TDA

[Nak00]  Takahiko Nakazi.  Two-

dimensional Q-algebras. Linear Algebra and its Applications, 315(1–3):197–205, August 15, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/132/25/
                                        37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/132/25/37/article.
                                        pdf.

Nakic:2003:RPT

[Nak03] Ivica Nakić. Relative perturbation theory for a class of diagonalizable Hermitian matrix pairs. Linear Algebra and its Applications, 358(1– 3):195–217, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/207/25/
36/abstract.html.

Narayanan:2005:SFC

[Nar05] H. Narayanan. On set functions that can be extended to convex functionals. Linear Algebra and its Applications, 402(1):74–100, June 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ng:2003:HPB

[NB03] Michael K. Ng and Zhong-Zhi Bai. A hybrid preconditioner of banded matrix approximation and alternating direction implicit iteration for symmetric Sinc–Galerkin linear systems. Linear Algebra and its Applications, 366 (C):317–335, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
43/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Nakazato:2005:OMS

[NBdP05] Hiroshi Nakazato, Natália Bebiano, and João da Providência. J-orthostochastic matrices of size 3 × 3 and numerical ranges of Krein space operators. Linear Algebra and its Applications, 407(1):211–232, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nakazato:2008:NRH

[NBdP08a] Hiroshi Nakazato, Natália Bebiano, and João da Providência. The J-numerical range of a J-Hermitian matrix and related inequalities. Linear Algebra and its Applications, 428(11–12):2995– 3014, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nakazato:2008:PDE

[NBdP08b] Hiroshi Nakazato, Natália Bebiano, and João da Providência. Product of diagonal entries of the unitary orbit of a 3-by-3 normal matrix. Linear Algebra and its Applications, 429 (4):698–715, August 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ng:2000:CTP

[NCCY00] Michael K. Ng, Raymond H. Chan, Tony F. Chan, and Andy M. Yip. Cosine transform preconditioners for high resolution image reconstruction. Linear Algebra and its Applications, 316(1–3): 89–104, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/30/article.
pdf.

Noutsos:2005:ABC

[NCV05] D. Noutsos, S. Serra Capizzano, and P. Vassalos. Asymptotic behavior of the condition number of two-level Toeplitz matrix sequences. Linear Algebra and its Applications, 395(1):121–140, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Noutsos:2008:CFM

[NCV08] D. Noutsos, S. Serra Capizzano, and P. Vassalos. The conditioning of FD matrix sequences coming from semielliptic differential equations. Linear Algebra and its Applications, 428(2–3):600–624, January 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Neogy:2005:PPT

[ND05] S. K. Neogy and A. K. Das. Principal pivot transforms of some classes of matrices. Linear Algebra and its Applications, 400(1):243– 252, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nagarajan:2002:PTT

[NDS02] K. R. Nagarajan, M. Paul Devasahayam, and T. Soundararajan. Products of three triangular matrices over commutative rings. Linear Algebra and its Applications, 348(1–3):1–6, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/195/27/
                                        27/abstract.html.

Nedoma:2001:SVS

[Ned01a] Josef Nedoma. On solving vague systems of linear equations with pattern-shaped columns. Linear Algebra and its Applications, 324(1–3): 107–118, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/144/24/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/144/24/30/article.
                                        pdf.

Nedoma:2001:PRV

[Ned01b] Josef Nedoma. Positively regular vague matrices. Linear Algebra and its Applications, 326(1–3):85–100, March 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/148/24/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/148/24/28/article.
pdf.

Neumann:2000:IPC

[Neu00] Michael Neumann. Inverses of Perron complements of inverse M-matrices. Linear Algebra and its Applications, 313(1–3):163–171, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/130/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/130/25/35/article.
pdf.

Neumaier:2001:GLS

[Neu01]  Arnold Neumaier.  Gen-

eralized Lyapunov-Schmidt reduction for parametrized equations at near singular points. Linear Algebra and its Applications, 324(1–3): 119–131, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/144/24/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/144/24/31/article.
                                        pdf.

Neymeyr:2001:GTPa

[Ney01a] Klaus Neymeyr. A geometric theory for preconditioned inverse iteration. I: Extrema of the Rayleigh quotient. Linear Algebra and its Applications, 322(1– 3):61–85, January 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/141/24/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/141/24/28/article.
                                        pdf.

Neymeyr:2001:GTPb

[Ney01b] Klaus Neymeyr. A geometric theory for preconditioned inverse iteration. II: Convergence estimates. Linear Algebra and its Applications, 322(1–3):87–104, January 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/141/24/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/141/24/29/article.
                                        pdf.

Neymeyr:2006:GTP

[Ney06] Klaus Neymeyr. A geometric theory for preconditioned inverse iteration IV: On the fastest convergence cases. Linear Algebra and its Applications, 415(1):114– 139, May 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Neymeyr:2009:PEI

[Ney09] Klaus Neymeyr. On preconditioned eigensolvers and Invert–Lanczos processes. Linear Algebra and its Applications, 430(4):1039–1056, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Niezgoda:2000:LPG

[Nie00] Marek Niezgoda. On lattice property of group induced cone orderings. Linear Algebra and its Applications, 307 (1–3):77–87, March 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/30/article.
pdf.

Nievergelt:2001:HHF

[Nie01a]  Yves Nievergelt.  Hyper-

spheres and hyperplanes fitted seamlessly by algebraic constrained total leastsquares. Linear Algebra and its Applications, 331 (1–3):43–59, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/156/25/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/156/25/29/article.
                                        pdf.

Niezgoda:2001:LMP

[Nie01b]  Marek Niezgoda.  Linear

maps preserving group majorization. Linear Algebra and its Applications, 330(1– 3):113–127, June 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/155/25/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/155/25/33/article.
                                        pdf.

Nievergelt:2004:FCS

[Nie04a]  Yves Nievergelt.  Fitting

conics of specific types to data. Linear Algebra and its Applications, 378(1):1–30, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Niezgoda:2004:CEC

[Nie04b]  Marek Niezgoda.  Cases

of equality for certain inequalities involving a Gdoubly stochastic map. Linear Algebra and its Applications, 378(1):167–182, February 1, 2004. CO-

Niezgoda:2007:ERH

[Nie07a] Marek Niezgoda. An extension of a result of Hwang and Pyo. Linear Algebra and its Applications, 420(1):163– 174, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Niezgoda:2007:LST

[Nie07b] Marek Niezgoda. Laguerre– samuelson type inequalities. Linear Algebra and its Applications, 422(2–3):574–581, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Niezgoda:2007:LPN

[Nie07c] Marek Niezgoda. On linear preservers of normal maps. Linear Algebra and its Applications, 426(1):149– 158, October 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Niezgoda:2008:SRN

[Nie08] Marek Niezgoda. Some remarks on normal maps with applications to eigenvalues and singular values of matrices. Linear Algebra and its Applications, 428(8–9):2116– 2129, April 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Niezgoda:2009:CAO

[Nie09] Marek Niezgoda. Commutators and accretive operators. Linear Algebra and its Applications, 431(8):1192–1198, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nikiforov:2006:EDD

[Nik06a] Vladimir Nikiforov. Eigenvalues and degree deviation in graphs. Linear Algebra and its Applications, 414(1): 347–360, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nikiforov:2006:EED

[Nik06b] Vladimir Nikiforov. Eigenvalues and extremal degrees of graphs. Linear Algebra and its Applications, 419(2–3): 735–738, December 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nikiforov:2006:WSR

[Nik06c] Vladimir Nikiforov. Walks and the spectral radius of graphs. Linear Algebra and its Applications, 418(1):257– 268, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nikiforov:2007:BGEa

[Nik07a] Vladimir Nikiforov. Bounds on graph eigenvalues I. Linear Algebra and its Applications, 420(2–3):667–671, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nikiforov:2007:BGEb

[Nik07b] Vladimir Nikiforov. Bounds on graph eigenvalues II. Linear Algebra and its Applications, 427(2–3):183–189, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nikiforov:2007:CNS

[Nik07c] Vladimir Nikiforov. Chromatic number and spectral radius. Linear Algebra and its Applications, 426(2–3):810– 814, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nikiforov:2007:EFS

[Nik07d] Vladimir Nikiforov. Eigenvalues and forbidden subgraphs I. Linear Algebra and its Applications, 422(1):284– 290, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nikiforov:2008:EFG

[Nik08a] V. Nikiforov. The energy of C_{4}-free graphs of bounded degree. Linear Algebra and its Applications, 428(11–12): 2569–2573, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nikiforov:2008:SCO

[Nik08b] Vladimir Nikiforov. A spectral condition for odd cycles in graphs. Linear Algebra and its Applications, 428 (7):1492–1498, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nikiforov:2009:MSR

[Nik09] Vladimir Nikiforov. The maximum spectral radius of C_{4}-free graphs of given order and size. Linear Algebra and its Applications, 430(11–12): 2898–2905, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nishizawa:2002:ECD

[Nis02] Michitomo Nishizawa. Evaluation of a certain qdeterminant. Linear Algebra and its Applications, 342 (1–3):107–115, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/173/27/
34/abstract.html.

Naqvi:2004:ENM

[NM04] Sarah Carnochan Naqvi and Judith J. McDonald. Eventually nonnegative matrices are similar to seminonnegative matrices. Linear Algebra and its Applications, 381(1): 245–258, April 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nagatou:2001:EME

[NN01] K. Nagatou and M. T. Nakao. An enclosure method of eigenvalues for the elliptic operator linearized at an exact solution of nonlinear problems. Linear Algebra and its Applications, 324(1–3): 81–106, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/144/24/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/144/24/29/article.
pdf.

Nomura:2005:RSD

[Nom05a] Kazumasa Nomura. A refinement of the split decomposition of a tridiagonal pair. Linear Algebra and its Applications, 403(1): 1–23, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2005:TPA

[Nom05b] Kazumasa Nomura. Tridiagonal pairs and the Askey– Wilson relations. Linear Algebra and its Applications, 397(1):99–106, March 1, 2005. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2005:TPH

[Nom05c] Kazumasa Nomura. Tridiagonal pairs of height one. Linear Algebra and its Applications, 403(1):118–142, July 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Norenberg:2003:ECH

[Nör03] R. Nörenberg. From elementary calculations to Hall polynomials. Linear Algebra and its Applications, 365 (C):301–309, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/214/25/
                                        43/abstract.html.

Noutsos:2006:PFP

[Nou06]  Dimitrios Noutsos.  On

Perron–Frobenius property of matrices having some negative entries. Linear Algebra and its Applications, 412(2–3):132–153, January 15, 2006. CODEN

Noutsos:2008:SRT

[Nou08] Dimitrios Noutsos. On Stein– Rosenberg type theorems for nonnegative and Perron– Frobenius splittings. Linear Algebra and its Applications, 429(8–9):1983–1996, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Novak:2009:SCS

[Nov09] Nika Novak. Sums of commuting square-zero transformations. Linear Algebra and its Applications, 431(10): 1768–1773, October 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nakazato:2001:SNR

[NP01] Hiroshi Nakazato and Panayiotis Psarrakos. On the shape of numerical range of matrix polynomials. Linear Algebra and its Applications, 338 (1–3):105–123, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/35/article.
pdf.

Nyman:2007:TSV

[NP07] Adam Nyman and Christopher J. Pappacena. Twosided vector spaces. Linear Algebra and its Applications, 420(2–3):339–360, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ng:2000:NAC

[NPP00] Michael K. Ng, Robert J. Plemmons, and Felipe Pimentel. A new approach to constrained total least squares image restoration. Linear Algebra and its Applications, 316(1–3):237–258, September 1, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/39/article.
pdf.

Ng:2002:ITM

[NRW02] Michael K. Ng, Karla Rost, and You-Wei Wen. On inversion of Toeplitz matrices. Linear Algebra and its Applications, 348(1–3): 145–151, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
37/abstract.html.

Nieuwenhuis:2002:SMN

[NS02] Herman J. Nieuwenhuis and Lambert Schoonbeek. Stability of matrices with negative diagonal submatrices. Linear Algebra and its Applications, 353(1–3):183–196, September 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nieslony:2003:AFF

[NS03] A. Nieslony and G. Steidl. Approximate factorizations of Fourier matrices with nonequispaced knots. Linear Algebra and its Applications, 366(C):337–351, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
44/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Nath:2007:NSA

[NS07a] Milan Nath and Bhaba Kumar Sarma. On the nullspaces of acyclic and unicyclic singular graphs. Linear Algebra and its Applications, 427(1):42–54, November 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nitica:2007:SMP

[NS07b] V. Nitica and I. Singer. The structure of max-plus hyperplanes. Linear Algebra and its Applications, 426(2–3): 382–414, October 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Neumann:2008:NPP

[NS08a] Michael Neumann and Nung-Sing Sze. A note on the perturbation of positive matrices by normal and unitary matrices. Linear Algebra and its Applications, 428(1):224– 229, January 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nitica:2008:CMMa

[NS08b] V. Nitica and I. Singer. Contributions to max–min convex geometry. I: Segments. Linear Algebra and its Applications, 428(7):1439–1459, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nitica:2008:CMMb

[NS08c] V. Nitica and I. Singer. Contributions to max–min convex geometry. II: Semispaces and convex sets. Linear Algebra and its Applications, 428(8–9):2085–2115, April 15, 2008. CODEN LAAPAW. ISSN 0024-

Niezgoda:2001:NPR

[NT01] M. Niezgoda and Tin-Yau Tam. On the norm property of G(c)-radii and Eaton triples. Linear Algebra and its Applications, 336(1–3): 119–130, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/166/27/35/article.
pdf.

Neudecker:2002:TFM

[NT02] Heinz Neudecker and Götz Trenkler. Third and fourth moment matrices of vecX in multivariate analysis. Linear Algebra and its Applications, 354(1–3):223–229, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
44/abstract.html.

Nomura:2006:DAL

[NT06a] Kazumasa Nomura and Paul Terwilliger. The determinant of AA^{⋆} − A^{⋆}A for a Leonard pair A, A^{⋆}. Linear Algebra and its Applications, 416 (2–3):880–889, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2006:MUA

[NT06b] Kazumasa Nomura and Paul Terwilliger. Matrix units associated with the split basis of a Leonard pair. Linear Algebra and its Applications, 418(2–3):775–787, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2006:STF

[NT06c] Kazumasa Nomura and Paul Terwilliger. Some trace formulae involving the split sequences of a Leonard pair. Linear Algebra and its Applications, 413(1):189–201, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2007:BLP

[NT07a] Kazumasa Nomura and Paul Terwilliger. Balanced Leonard pairs. Linear Algebra and its Applications, 420 (1):51–69, January 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2007:LTT

[NT07b] Kazumasa Nomura and Paul Terwilliger. Linear transformations that are tridiagonal with respect to both eigenbases of a Leonard pair.

Linear Algebra and its Applications, 420(1):198–207, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2007:SDT

[NT07c] Kazumasa Nomura and Paul Terwilliger. The split decomposition of a tridiagonal pair. Linear Algebra and its Applications, 424(2– 3):339–345, July 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2008:STPa

[NT08a] Kazumasa Nomura and Paul Terwilliger. Sharp tridiagonal pairs. Linear Algebra and its Applications, 429(1): 79–99, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2008:STPb

[NT08b] Kazumasa Nomura and Paul Terwilliger. The structure of a tridiagonal pair. Linear Algebra and its Applications, 429(7):1647–1662, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2008:SEL

[NT08c] Kazumasa Nomura and Paul Terwilliger. The switching element for a Leonard pair.

Linear Algebra and its Applications, 428(4):1083–1108, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2008:TCT

[NT08d] Kazumasa Nomura and Paul Terwilliger. Towards a classification of the tridiagonal pairs. Linear Algebra and its Applications, 429(2–3):503– 518, July 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2009:TMB

[NT09a] Kazumasa Nomura and Paul Terwilliger. Transition maps between the 24 bases for a Leonard pair. Linear Algebra and its Applications, 431(5– 7):571–593, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nomura:2009:TPC

[NT09b] Kazumasa Nomura and Paul Terwilliger. Tridiagonal pairs and the μ-conjecture. Linear Algebra and its Applications, 430(1):455–482, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Noutsos:2009:NCR

[NT09c] Dimitrios Noutsos and Michael J. Tsatsomeros. On the numerical characterization of the reachability cone for an essentially nonnegative matrix. Linear Algebra and its Applications, 430(4):1350–1363, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nakic:2003:WKF

[NV03] Ivica Nakić and Krešimir Veselić. Wielandt and Ky-Fan theorem for matrix pairs. Linear Algebra and its Applications, 369(C):77– 93, August 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
29/abstract.html.

Neumann:2004:IBC

[NX04] M. Neumann and J. Xu. Improved bounds for a condition number for Markov chains. Linear Algebra and its Applications, 386(1):225– 241, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nan:2000:CCF

[NY00] Jizhu Nan and Hong You. The conjugacy classes of fixed point free elements in GL_{n}(F) and SL_{n}(F). Linear Algebra and its Applications, 305(1–3):161–171, January 15, 2000. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/120/25/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/120/25/36/article.
                                        pdf.

Nakazi:2002:TDC

[NY02] Takahiko Nakazi and Takanori Yamamoto. Two dimensional commutative Banach algebras and von Neumann inequality. Linear Algebra and its Applications, 341(1–3): 273–290, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/171/27/
                                        48/abstract.html.

Nakagami:2007:CDR

[NY07a] Yoshiomi Nakagami and Haruo Yanai. On Cullis’ determinant for rectangular matrices. Linear Algebra and its Applications, 422(2–3): 422–441, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Nakazi:2007:FDS

[NY07b] Takahiko Nakazi and Takanori Yamamoto. Finite dimensional semisimple Q-algebras. Linear Algebra and its Applications, 420(2–3):407–423, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-

Ng:2006:AEL

[NZ06] Che Tat Ng and Weinian Zhang. An algebraic equation for linear operators. Linear Algebra and its Applications, 412(2–3):303–325, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ozturk:2000:ICM

[ÖA00] Fikri Öztürk and Fikri Akdeniz. Ill-conditioning and multicollinearity. Linear Algebra and its Applications, 321(1–3):295–305, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
43/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/43/article.
pdf.

Oara:2009:ZCG

[OA09] Cristian Oară and Raluca Andrei. Zero cancellation for general rational matrix functions. Linear Algebra and its Applications, 431(11):2000– 2023, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Omran:2001:RPD

[OB01] Mohammad Othman Omran and Wayne Barrett. The real positive definite completion problem for a 4cycle. Linear Algebra and its Applications, 336(1–3): 131–166, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/166/27/36/article.
pdf.

Oblak:2008:JFM

[Obl08] Polona Oblak. Jordan forms for mutually annihilating nilpotent pairs. Linear Algebra and its Applications, 428 (7):1476–1491, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Okay:2007:ADL

[OCK07] Irfan Okay, Stephen L. Campbell, and Peter Kunkel. The additional dynamics of least squares completions for linear differential algebraic equations. Linear Algebra and its Applications, 425(2–3):471–485, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Okay:2009:CID

[OCK09] Irfan Okay, Stephen L. Campbell, and Peter Kunkel. Completions of implicitly defined linear time varying vector fields. Linear Algebra and its Applications, 431(9): 1422–1438, October 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oliveira:2002:CPL

[OdAJ02] Carla Silva Oliveira, Nair Maria Maia de Abreu, and Samuel Jurkiewicz. The characteristic polynomial of the Laplacian of graphs in (a, b)- linear classes. Linear Algebra and its Applications, 356 (1–3):113–121, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/202/25/
35/abstract.html.

OKeeffe:2002:GBS

[OF02] Henry O’Keeffe and Patrick Fitzpatrick. Gröbner basis solutions of constrained interpolation problems. Linear Algebra and its Applications, 351–352:533–551, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ohno:2008:QOS

[Ohn08]  Hiromichi Ohno.  Quasi-

orthogonal subalgebras of matrix algebras. Linear Algebra and its Applications, 429(8–9):2146–2158, October 16, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oikhberg:2008:PPN

[Oik08] Timur Oikhberg. Products of projections in von Neumann algebras. Linear Algebra and its Applications, 429(4):759– 775, August 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oishi:2001:FEM

[Ois01] Shin’ichi Oishi. Fast enclosure of matrix eigenvalues and singular values via rounding mode controlled computation. Linear Algebra and its Applications, 324 (1–3):133–146, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/144/24/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/144/24/32/article.
pdf.

Dokovic:2006:SIG

[– Dok06] Dragomir Ž. – Doković. Structure of isometry group of bilinear spaces. Linear Algebra and its Applications, 416 (2–3):414–436, July 15, 2006. CODEN LAAPAW. ISSN

Okninski:2001:SZE

[Okn01] Jan Okniński. Semigroups of zero entropy. Linear Algebra and its Applications, 323(1–3):207–211, January 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oktac:2004:SDL

[Okt04] Asuman Oktaç. Student discussions on a linear algebra problem in a distanceeducation course. Linear Algebra and its Applications, 379(1):439–455, March 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Okubo:2003:WUI

[Oku03] K. Okubo. On weakly unitarily invariant norm and the Aluthge transformation. Linear Algebra and its Applications, 371(1):369–375, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Okubo:2005:STR

[Oku05] Susumu Okubo. Symmetric triality relations and structurable algebras. Linear Algebra and its Applications, 396(1):189–222, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Okubo:2006:WUI

[Oku06]  Kazuyoshi Okubo.  On

weakly unitarily invariant norm and the λ-Aluthge transformation for invertible operator. Linear Algebra and its Applications, 419(1):48– 52, November 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

OLeary:2003:SSP

[O’L03] Dianne P. O’Leary. Scaling symmetric positive definite matrices to prescribed row sums. Linear Algebra and its Applications, 370(1):185– 191, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

OLeary:2008:GHG

[O’L08] Dianne P. O’Leary. Gene Howard Golub, 1932–2007. Linear Algebra and its Applications, 428(11–12):2405– 2409, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Olkin:2002:ADR

[Olk02] Ingram Olkin. The 70th anniversary of the distribution of random matrices: A survey. Linear Algebra and its Applications, 354(1–3): 231–243, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
45/abstract.html.

Omidi:2008:LSC

[Omi08a] G. R. Omidi. On a Laplacian spectral characterization of graphs of index less than 2. Linear Algebra and its Applications, 429(11–12):2724– 2731, December 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Omidi:2008:SCG

[Omi08b] G. R. Omidi. The spectral characterization of graphs of index less than 2 with no path as a component. Linear Algebra and its Applications, 428 (7):1696–1705, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Omidi:2009:SLS

[Omi09] G. R. Omidi. On a signless Laplacian spectral characterization of T -shape trees. Linear Algebra and its Applications, 431(9):1607–1615, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Omladic:2004:VCT

[Oml04] Matjaž Omladič. A variety of commuting triples. Linear Algebra and its Applications, 383(1):233–245, May 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Omladic:2000:MRC

[OO00] Matjaž Omladič and Vesna Omladič. More on restricted canonical correlations. Linear Algebra and its Applications, 321(1–3): 285–293, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/140/24/
                                        42/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/140/24/42/article.
                                        pdf.

Olshevsky:2005:KTB

[OO05] Alex Olshevsky and Vadim Olshevsky. Kharitonov’s Theorem and Bezoutians. Linear Algebra and its Applications, 399(1):285–297, April 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Olshevsky:2006:TPM

[OOT06] Vadim Olshevsky, Ivan Oseledets, and Eugene Tyrtyshnikov. Tensor properties of multilevel Toeplitz and related matrices. Linear Algebra and its Applications, 412 (1):1–21, January 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oruc:2000:EFV

[OP00] Halil Oruç and George M. Phillips. Explicit factorization of the Vandermonde matrix. Linear Algebra and its Applications, 315(1–3): 113–123, August 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/132/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/132/25/30/article.
pdf.

Oliveira:2005:SPM

[OP05] Ricardo C. L. F. Oliveira and Pedro L. D. Peres. Stability of polytopes of matrices via affine parameterdependent Lyapunov functions: Asymptotically exact LMI conditions. Linear Algebra and its Applications, 405 (1):209–228, August 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ohno:2007:QOS

[OPS07] Hiromichi Ohno, Dénes Petz, and András Szántó. Quasiorthogonal subalgebras of 4× 4 matrices. Linear Algebra and its Applications, 425 (1):109–118, August 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Olesky:2002:MGG

[ORvdD02] D. D. Olesky, Aidan Roy, and P. van den Driessche. Maximal graphs and graphs with maximal spectral radius. Linear Algebra and its Applications, 346 (1–3):109–130, May 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/189/27/
                                        33/abstract.html.

Oliveira:2000:ESP

[OS00] Suely Oliveira and David E. Stewart. Exponential splittings of products of matrices and accurately computing singular values of long products. Linear Algebra and its Applications, 309(1– 3):175–190, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/126/25/
                                        35/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/126/25/35/article.
                                        pdf.

Okubo:2001:CCM

[OS01] Kazuyoshi Okubo and Ilya Spitkovsky. On the characterization of 2×2-contraction matrices. Linear Algebra and its Applications, 325(1– 3):177–189, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/147/24/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/147/24/36/article.
pdf.

ORyan:2003:CHD

[OS03] Manuel O’Ryan and Daniel B. Shapiro. Centers of higher degree forms. Linear Algebra and its Applications, 371(1): 301–314, September 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oliveira:2005:NCP

[OS05] A. R. L. Oliveira and D. C. Sorensen. A new class of preconditioners for large-scale linear systems from interior point methods for linear programming. Linear Algebra and its Applications, 394(1): 1–24, January 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oara:2009:MIC

[OS09] Cristian Oară and Şerban Sabău. Minimal indices cancellation and rank revealing factorizations for rational matrix functions. Linear Algebra and its Applications, 431(10):1785–1814, October 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oseledets:2008:IOL

[Ose08a] Ivan Oseledets. The integral operator with logarithmic kernel has only one positive eigenvalue. Linear Algebra and its Applications, 428 (7):1560–1564, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oseledets:2008:OKL

[Ose08b] Ivan Oseledets. Optimal Karatsuba-like formulae for certain bilinear forms in GF(2). Linear Algebra and its Applications, 429(8–9): 2052–2066, October 16, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ostrovskii:2003:MVP

[Ost03] M. I. Ostrovskii. Minimalvolume projections of cubes and totally unimodular matrices. Linear Algebra and its Applications, 364(C):91– 103, May 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/213/25/
                                        27/abstract.html.

Ostrowski:2006:PES

[Ost06] Tadeusz Ostrowski. Population equilibrium with support in evolutionary matrix games. Linear Algebra and its Applications, 417(1):211– 219, August 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Olesky:2002:ETN

[OSvdD02] D. D. Olesky, Bryan Shader, and P. van den Driessche. Exponents of tuples of nonnegative matrices. Linear Algebra and its Applications, 356 (1–3):123–134, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/202/25/
36/abstract.html.

Osikiewicz:2001:IAH

[OT01] Beth Osikiewicz and Andrew Tonge. An interpolation approach to Hardy– Littlewood inequalities for norms of operators on sequence spaces. Linear Algebra and its Applications, 331(1–3):1–9, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/25/article.
pdf.

Oseledets:2006:UAC

[OT06] Ivan Oseledets and Eugene Tyrtyshnikov. A unifying approach to the construction of circulant preconditioners. Linear Algebra and its Applications, 418(2–3):435–449, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Omidi:2007:STD

[OT07] G. R. Omidi and K. Tajbakhsh. Starlike trees are determined by their Laplacian spectrum. Linear Algebra and its Applications, 422(2–3):654– 658, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Osaka:2009:DPS

[OT09] Hiroyuki Osaka and Jun Tomiyama. Double piling structure of matrix monotone functions and of matrix convex functions. Linear Algebra and its Applications, 431(10): 1825–1832, October 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Oliveira:2003:SCT

[OTC03] Vilma A. Oliveira, Marcelo C. M. Teixeira, and Lúcia Cossi. Stabilizing a class of time delay systems using the Hermite-Biehler theorem. Linear Algebra and its Applications, 369(C):203– 216, August 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
37/abstract.html.

Otsuka:2001:GCC

[Ots01] Naohisa Otsuka. Generalized controlled and conditioned invariances for linear ω-periodic discrete-time systems. Linear Algebra and its Applications, 327(1–3):207– 223, April 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/41/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/152/
25/41/article.pdf.

Oseledets:2009:MIC

[OTZ09] Ivan Oseledets, Eugene Tyrtyshnikov, and Nickolai Zamarashkin. Matrix inversion cases with size-independent tensor rank estimates. Linear Algebra and its Applications, 431(5–7):558–570, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ouchi:2002:PMU

[O’u02]  Moto O’uchi.  Pseudo-

multiplicative unitaries associated with inclusions of finite-dimensional C^{∗}- algebras. Linear Algebra and its Applications, 341(1–3): 201–218, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/171/27/
                                        43/abstract.html.

OMeara:2006:ASD

[OV06a] K. C. O’Meara and C. Vinsonhaler. On approximately simultaneously diagonalizable matrices. Linear Algebra and its Applications, 412(1): 39–74, January 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Onn:2006:PMM

[OV06b] Shmuel Onn and Ernesto Vallejo. Permutohedra and minimal matrices. Linear Algebra and its Applications, 412(2–3):471–489, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Olesky:2000:GSD

[OvdDV00] D. D. Olesky, P. van den Driessche, and J. H. Verner. Graphs with the same determinant as a complete graph. Linear Algebra and its Applications, 312(1–3): 191–195, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/129/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/129/25/37/article.
pdf.

Ovsienko:2003:SNM

[Ovs03] Serge Ovsienko. Strongly nilpotent matrices and Gelfand-Zetlin modules. Linear Algebra and its Applications, 365 (C):349–367, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
52/abstract.html.

Ovtchinnikov:2006:CREa

[Ovt06a] E. Ovtchinnikov. Cluster robust error estimates for the Rayleigh–Ritz approximation I: Estimates for invariant subspaces. Linear Algebra and its Applications, 415(1): 167–187, May 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ovtchinnikov:2006:CREb

[Ovt06b] E. Ovtchinnikov. Cluster robust error estimates for the Rayleigh–Ritz approximation II: Estimates for eigenvalues. Linear Algebra and its Applications, 415(1):188– 209, May 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ovtchinnikov:2006:CRP

[Ovt06c] E. Ovtchinnikov. Cluster robustness of preconditioned gradient subspace iteration eigensolvers. Linear Algebra and its Applications, 415(1): 140–166, May 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ou:2007:DLA

[OWY07] Shikun Ou, Dengyin Wang, and Ruiping Yao. Derivations of the Lie algebra of strictly upper triangular matrices over a commutative ring. Linear Algebra and its Applications, 424(2–3):378– 383, July 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Okoh:2003:CAK

[OZ03a] F. Okoh and F. A. Zorzitto. Curves arising from Kronecker modules. Linear Algebra and its Applications, 365 (C):311–348, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/214/25/
                                        45/abstract.html.

Okoh:2003:CARa

[OZ03b] Frank Okoh and Frank Zorzitto. Commutative algebras of rational function matrices as endomorphisms of Kronecker modules I. Linear Algebra and its Applications, 374(1):41–62, November 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Okoh:2003:CARb

[OZ03c] Frank Okoh and Frank Zorzitto. Commutative algebras of rational function matrices as endomorphisms of Kronecker modules II. Linear Algebra and its Applications, 374(1):63–85, November 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pinar:2004:RSL

[PA04] Mustafa Ç. Pınar and Orhan Arıkan. On robust solutions to linear least squares problems affected by data uncertainty and implementation errors with application to stochastic signal modeling. Linear Algebra and its Applications, 391(1):223–243, November 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Packer:2002:SCM

[Pac02] Asa Packer. Symmetric conference matrices and locally largest regular crosspolytopes in cubes. Linear Algebra and its Applications, 357(1–3):1–13, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/203/25/
                                        26/abstract.html.

Pan:2000:ECR

[Pan00] C.-T. Pan. On the existence and computation of rank-revealing LU factorizations. Linear Algebra and its Applications, 316(1–3): 199–222, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/134/24/
                                        37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/134/24/37/article.
                                        pdf.

Pan:2002:SUB

[Pan02] Yong-Liang Pan. Sharp upper bounds for the Laplacian graph eigenvalues. Linear Algebra and its Applications, 355(1–3):287–295, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/201/27/
                                        44/abstract.html.

Pappacena:2000:MPG

[Pap00] Christopher J. Pappacena. Matrix pencils and a generalized Clifford algebra. Linear Algebra and its Applications, 313(1–3):1–20, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/130/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/130/25/25/article.
pdf.

Parlett:2002:MDD

[Par02a] Beresford N. Parlett. The (matrix) discriminant as a determinant. Linear Algebra and its Applications, 355(1– 3):85–101, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/201/27/
33/abstract.html.

Parthasarathy:2002:BMG

[Par02b] K. R. Parthasarathy. The Bush matrix over a Galois field and error correcting quantum codes. Linear Algebra and its Applications, 341(1–3):23–34, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
29/abstract.html.

Parlett:2009:RCG

[Par09a] Beresford N. Parlett. A result complementary to Gersgorin’s Circle Theorem. Linear Algebra and its Applications, 431(1–2):20–27, July 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Parlett:2009:RFM

[Par09b] Beresford N. Parlett. Review of Functions of Matrices: Theory and Computation, Nicholas J. Higham. SIAM, Philadelphia, PA (2008), xx + 425 pp. Linear Algebra and its Applications, 430(11–12): 3139–3140, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pascasio:2001:ICT

[Pas01] Arlene A. Pascasio. An inequality on the cosines of a tight distance-regular graph. Linear Algebra and its Applications, 325(1–3): 147–159, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/147/24/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/147/24/34/article.
                                        pdf.

Patricio:2001:MPI

[Pat01] Pedro Patrício. The Moore– Penrose inverse of von Neumann regular matrices over a ring. Linear Algebra and its Applications, 332–334(1):

469–483, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
49/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/49/article.
pdf.

Patricio:2003:MPI

[Pat03] Pedro Patrício. The Moore– Penrose inverse of a factorization. Linear Algebra and its Applications, 370(1):227– 235, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pate:2004:ERF

[Pat04] Thomas H. Pate. Extreme rays and faces for the cone of class functions non-negative on the set of Gram matrices. Linear Algebra and its Applications, 390(1):27– 45, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pate:2008:PC

[Pat08a] Thomas H. Pate. On permanental compounds. Linear Algebra and its Applications, 429(5–6):1093–1101, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pate:2008:TIP

[Pat08b] Thomas H. Pate. Twisted inner products and contraction inequalities on spaces of contravariant and covariant tensors. Linear Algebra and its Applications, 429(7): 1489–1503, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Parlett:2006:AOM

[PB06] B. N. Parlett and E. Barszcz. Another orthogonal matrix. Linear Algebra and its Applications, 417(2–3):342–346, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Petrovic:2009:BGW

[PBA09] Miroslav Petrović, Bojana Borovićanin, and Tatjana Aleksić. Bicyclic graphs for which the least eigenvalue is minimum. Linear Algebra and its Applications, 430(4): 1328–1335, February 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Bai:2009:LAI

[pBhSzZ09] Rui pu Bai, Cai hong Shen, and Yao zhong Zhang. 3-Lie algebras with an ideal N. Linear Algebra and its Applications, 431(5–7):673– 700, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-

Bai:2008:SLD

[pBlWyXwA08] Rui pu Bai, Xiao ling Wang, Wen ying Xiao, and Hong wei An. Structure of low dimensional n-Lie algebras over a field of characteristic 2. Linear Algebra and its Applications, 428(8–9):1912– 1920, April 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Petrovic:2006:IHG

[PBR06] Miroslav Petrović, Bojana Borovićanin, and Zoran Radosavljević. The integral 3-harmonic graphs. Linear Algebra and its Applications, 416(2–3):298–312, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Petrovic:2004:GMT

[PBT04] Miroslav Petrović, Bojana Borovićanin, and Aleksandar Torgašev. On graphs with at most three Laplacian eigenvalues greater than or equal to two. Linear Algebra and its Applications, 380(1):173– 184, March 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pison-Casares:2004:SLS

[PCVT04] Pilar Pisón-Casares and Alberto Vigneron-Tenorio. Nsolutions to linear systems over Z. Linear Algebra and its Applications, 384(1):135– 154, June 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Parlett:2000:RRR

[PD00] Beresford N. Parlett and Inderjit S. Dhillon. Relatively robust representations of symmetric tridiagonals. Linear Algebra and its Applications, 309(1–3): 121–151, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/33/article.
pdf.

Peisheng:2007:JMT

[Pei07] Ji Peisheng. Jordan maps on triangular algebras. Linear Algebra and its Applications, 426(1):190–198, October 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pena:2001:DCT

[Peñ01] J. M. Peña. Determinantal criteria for total positivity. Linear Algebra and its Applications, 332–334(1): 131–137, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/33/article.
pdf.

Pena:2003:NSR

[Peñ03a] J. M. Peña. On nonsingular sign regular matrices. Linear Algebra and its Applications, 359(1–3): 91–100, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
33/abstract.html.

Pena:2003:CDI

[Peñ03b] Javier Peña. A characterization of the distance to infeasibility under block-structured perturbations. Linear Algebra and its Applications, 370(1):193–216, September 1, 2003. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pena:2004:CST

[Peñ04] J. M. Peña. Characterizations and stable tests for the Routh–Hurwitz conditions and for total positivity. Linear Algebra and its Applications, 393(1):319–332, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Penzl:2006:AMR

[Pen06]  Thilo Penzl.  Algorithms

for model reduction of large dynamical systems. Linear Algebra and its Applications, 415(2–3):322–343, June 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pena:2008:HOL

[Peñ08]  J. M. Peña.  Hierarchi-

cal open Leontief models. Linear Algebra and its Applications, 428(11–12):2549– 2559, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pena:2009:PSM

[Peñ09] J. M. Peña. Positive symmetric matrices with exactly one positive eigenvalue. Linear Algebra and its Applications, 430(5–6):1566–1573, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Peperko:2008:MVG

[Pep08] Aljoša Peperko. On the max version of the generalized spectral radius theorem. Linear Algebra and its Applications, 428(10):2312– 2318, May 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pauca:2000:SMR

[PEPS00] V. P. Pauca, B. L. Ellerbroek, R. J. Plemmons, and X. Sun. Structured matrix representations of two-parameter Hankel transforms in adaptive optics. Linear Algebra and its Applications, 316(1– 3):29–43, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/27/article.
pdf.

Peretz:2003:MSA

[Per03] Dvora Peretz. Mathematics, simple or abstract? Cognitive abstraction processes in learning mathematics. Linear Algebra and its Applications, 361(C):23–39, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
29/abstract.html.

Pereira:2007:WLM

[Per07] Rajesh Pereira. Weak logmajorization, Mahler measure and polynomial inequalities. Linear Algebra and its Applications, 421(1):117– 121, February 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pickmann:2007:EIE

[PES07] Hubert Pickmann, Juan Egaña, and Ricardo L. Soto. Extremal inverse eigenvalue problem for bordered diagonal matrices. Linear Algebra and its Applications, 427(2–3):256–271, December 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Petek:2002:SCP

[Pet02] Tatjana Petek. Spectrum and commutativity preserving mappings on triangular matrices. Linear Algebra and its Applications, 357(1–3): 107–122, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/203/25/
                                        32/abstract.html.

Petek:2005:NUS

[Pet05]  Tatjana Petek.  A note

on unitary similarity preserving linear mappings on B(H). Linear Algebra and its Applications, 394(1):217– 224, January 1, 2005. CO-DEN LAAPAW. ISSN 0024-

Petkovic:2008:NHO

[Pet08] Miodrag S. Petković. On new higher order families of simultaneous methods for finding polynomial zeros. Linear Algebra and its Applications, 429(10):2602–2622, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Petrovic:2009:SRM

[Pet09] Zoran Z. Petrović. Spaces of real matrices of fixed small rank. Linear Algebra and its Applications, 431(8):1199– 1207, September 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pfander:2008:IRB

[Pfa08] Götz E. Pfander. On the invertibility of “rectangular” bi-infinite matrices and applications in time–frequency analysis. Linear Algebra and its Applications, 429(1):331– 345, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pan:2005:SRV

[PFL05] Yong-Liang Pan, Yi-Zheng Fan, and Jiong-Sheng Li. Spectral rational variation in two places for adjacency matrix is impossible. Linear Algebra and its Applications, 404(1):275–282, July 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pang:2001:NSC

[PG01] Chin-Tzong Pang and Sy-Ming Guu. A note on the sequence of consecutive powers of a nonnegative matrix in max algebra. Linear Algebra and its Applications, 330 (1–3):209–213, June 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/155/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/155/25/37/article.
pdf.

Puystjens:2004:DIM

[PG04] R. Puystjens and M. C. Gouveia. Drazin invertibility for matrices over an arbitrary ring. Linear Algebra and its Applications, 385(1):105– 116, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Paszke:2006:GCC

[PGRO06] Wojciech Paszke, Krzysztof Gałkowski, Eric Rogers, and David H. Owens. H_{∞} and guaranteed cost control of discrete linear repetitive processes. Linear Algebra and its Applications, 412(2–3): 93–131, January 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ptak:2000:RJN

[PH00] Vlastimil Pták and Olga Holtz. A remark on the Jordan normal form of matrices. Linear Algebra and its Applications, 310(1–3):5–7, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/127/25/26/article.
pdf.

Poplin:2004:DIC

[PH04] Phillip L. Poplin and Robert E. Hartwig. Determinantal identities over commutative semirings. Linear Algebra and its Applications, 387(1): 99–132, August 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Provost:2009:ICM

[PH09] Serge B. Provost and Hyung Tae Ha. On the inversion of certain moment matrices. Linear Algebra and its Applications, 430(10):2650–2658, May 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Phillips:2004:SPI

[Phi04] D. Paul Phillips. Some partial inverse eigenvalue problems: recovering diagonal entries of symmetric matrices. Linear Algebra and its Applications, 380(1):263– 270, March 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Peeters:2007:CLS

[PHO07] Ralf L. M. Peeters, Bernard Hanzon, and Martine Olivi. Canonical lossless state-space systems: Staircase forms and the Schur algorithm. Linear Algebra and its Applications, 425(2–3):404–433, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Peng:2006:TIE

[PHZ06] Juan Peng, Xi-Yan Hu, and Lei Zhang. Two inverse eigenvalue problems for a special kind of matrices. Linear Algebra and its Applications, 416(2–3):336– 347, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Picci:2007:MTI

[Pic07]  Giorgio Picci.  A mod-

ule theoretic interpretation of multiplicity and rank of a stationary random process. Linear Algebra and its Applications, 425(2–3):443–452, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pihlak:2004:MI

[Pih04] Margus Pihlak. Matrix integral. Linear Algebra and its Applications, 388(1):315– 325, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pittenger:2003:UPB

[Pit03] Arthur O. Pittenger. Unextendible product bases and the construction of inseparable states. Linear Algebra and its Applications, 359(1–3): 235–248, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
43/abstract.html.

Pituk:2009:NIA

[Pit09] Mihály Pituk. Nonnegative iterations with asymptotically constant coefficients. Linear Algebra and its Applications, 431(10):1815–1824, October 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pang:2007:AIS

[PJ07] Yongfeng Pang and Guoxing Ji. Algebraic isomorphisms and strongly double triangle subspace lattices. Linear Algebra and its Applications, 422(1):265– 273, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Parrilo:2008:AJS

[PJ08] Pablo A. Parrilo and Ali Jadbabaie. Approximation of the joint spectral radius using sum of squares. Linear Algebra and its Applications, 428 (10):2385–2402, May 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pires:2008:SLC

[PK08] M. Pires and V. G. Kravchenko. On a “stability” in the linear complementarity problem. Linear Algebra and its Applications, 429(7):1810– 1822, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Patra:2008:EAC

[PL08] K. L. Patra and A. K. Lal. The effect on the algebraic connectivity of a tree by grafting or collapsing of edges. Linear Algebra and its Applications, 428(4):855– 864, February 1, 2008. CO-

Plesnik:2007:FOP

[Ple07] Ján Plesník. Finding the orthogonal projection of a point onto an affine subspace. Linear Algebra and its Applications, 422(2–3):455– 470, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pang:2002:NPM

[PLG02] Chin-Tzong Pang, Yung-Yih Lur, and Sy-Ming Guu. A new proof of Mayer’s theorem. Linear Algebra and its Applications, 350(1–3):273– 278, July 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Liao:2006:MNP

[pLLfY06] An ping Liao, Yuan Lei, and Shi fang Yuan. The matrix nearness problem for symmetric matrices associated with the matrix equation [A^{T} XA, B^{T} XB] = [C, D]. Linear Algebra and its Applications, 418(2–3):939–954, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Plum:2001:CAE

[Plu01] Michael Plum. Computerassisted enclosure methods for elliptic differential equations. Linear Algebra and its Applications, 324(1–3): 147–187, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/144/24/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/144/24/33/article.
pdf.

Parlett:2000:IDA

[PM00] Beresford N. Parlett and Osni A. Marques. An implementation of the dqds algorithm (positive case). Linear Algebra and its Applications, 309(1–3):217–259, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/37/article.
pdf.

Pecaric:2005:SFR

[PM05] Josip Pečarić and Jadranka Mićić. Some functions reversing the order of positive operators. Linear Algebra and its Applications, 396(1):175– 187, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Peterson:2008:LR

[PM08] Michael Peterson and Chris Monico. F_{2} Lanczos revisited. Linear Algebra and its Applications, 428(4):1135– 1150, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Peria:2005:WMM

[PMS05] Francisco D. Martínez Pería, Pedro G. Massey, and Luis E. Silvestre. Weak matrix majorization. Linear Algebra and its Applications, 403(1): 343–368, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Poon:2008:SMM

[Poo08]  Edward Poon.  Schur-

multiplicative maps preserving unitarily invariant norms. Linear Algebra and its Applications, 428(4):865–870, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Popa:2001:EAO

[Pop01] Constantin Popa. Extension of an approximate orthogonalization algorithm to arbitrary rectangular matrices. Linear Algebra and its Applications, 331(1–3):181– 192, July 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/39/article.
pdf.

Popescu:2002:MIS

[Pop02] Gelu Popescu. Meromorphic interpolation in several variables. Linear Algebra and its Applications, 357(1–3): 173–196, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
37/abstract.html.

Popa:2008:CKE

[Pop08]  Constantin Popa.  Con-

strained Kaczmarz extended algorithm for image reconstruction. Linear Algebra and its Applications, 429(8–9): 2247–2267, October 16, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Potts:2003:FAD

[Pot03] Daniel Potts. Fast algorithms for discrete polynomial transforms on arbitrary grids. Linear Algebra and its Applications, 366 (C):353–370, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/

45/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Pournaki:2001:OBS

[Pou01] M. R. Pournaki. On the orthogonal basis of the symmetry classes of tensors associated with certain characters. Linear Algebra and its Applications, 336(1–3): 255–260, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
44/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/166/27/44/article.
pdf.

Patricio:2001:ANR

[PP01a] Pedro Patrício and Roland Puystjens. About the von Neumann regularity of triangular block matrices. Linear Algebra and its Applications, 332–334(1):485–502, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
50/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/50/article.
pdf.

Peluso:2001:SIT

[PP01b] R. Peluso and T. Politi. Some improvements for two-sided bounds on the inverse of diagonally dominant tridiagonal matrices. Linear Algebra and its Applications, 330 (1–3):1–14, June 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/155/25/
                                        25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/155/25/25/article.
                                        pdf.

Puerta:2002:GSC

[PP02] F. Puerta and X. Puerta. On the geometry of the set of controllability subspaces of a pair (A, B). Linear Algebra and its Applications, 351–352:585–599, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Patricio:2004:DMP

[PP04a] Pedro Patrício and Roland Puystjens. Drazin–Moore– Penrose invertibility in rings. Linear Algebra and its Applications, 389(1):159–173, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Patricio:2004:GIT

[PP04b] Pedro Patrício and Roland Puystjens. Generalized invertibility in two semigroups of a ring. Linear Algebra and its Applications, 377(1):125– 139, January 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Papathanasiou:2008:DMP

[PP08] Nikolaos Papathanasiou and Panayiotis Psarrakos. The distance from a matrix polynomial to matrix polynomials with a prescribed multiple eigenvalue. Linear Algebra and its Applications, 429(7): 1453–1477, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Popovych:2009:EDC

[PP09] Dmytro R. Popovych and Roman O. Popovych. Equivalence of diagonal contractions to generalized IWcontractions with integer exponents. Linear Algebra and its Applications, 431(5–7): 1096–1104, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Puerta:2002:GRK

[PPF02] X. Puerta, F. Puerta, and J. Ferrer. Global reduction to the Kronecker canonical form of a C^{r}-family of timeinvariant linear systems. Linear Algebra and its Applications, 346(1–3):27–45, May 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/189/27/
                                        29/abstract.html.

Pauca:2006:NMF

[PPP06] V. Paul Pauca, J. Piper, and Robert J. Plemmons. Nonnegative matrix factorization for spectral data analysis. Linear Algebra and its Applications, 416(1):29– 47, July 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Puerta:2004:VDO

[PPT04] F. Puerta, X. Puerta, and S. Tarragona. Versal deformations in orbit spaces. Linear Algebra and its Applications, 379(1):329–343, March 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pan:2005:CCI

[pPyHZ05] Xiao ping Pan, Xi yan Hu, and Lei Zhang. A class of constrained inverse eigenproblem and associated approximation problem for skew symmetric and centrosymmetric matrices. Linear Algebra and its Applications, 408(1):66–77, October 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Puerta:2002:CDM

[PPZ02] F. Puerta, X. Puerta, and I. Zaballa. A cellular decomposition of the manifold of observable conditioned invariant subspaces. Linear Algebra and its Applications, 351–352:601–621, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Puerta:2005:PCS

[PPZ05] F. Puerta, X. Puerta, and I. Zaballa. On the parametrization of the controllability subspaces of a controllable pair. Linear Algebra and its Applications, 401(1):429–444, May 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Prazmowska:2009:MAG

[PPZ09] M. Prazmowska, K. Prazmowski, and M. Zynel. Metric affine geometry on the universe of lines. Linear Algebra and its Applications, 430(11–12):3066– 3079, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Papendieck:2000:MEP

[PR00] Britta Papendieck and Peter Recht. On maximal entries in the principal eigenvector of graphs. Linear Algebra and its Applications, 310 (1–3):129–138, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/127/25/36/article.
pdf.

Pecaric:2001:BWN

[PR01] Josip Pečarić and Ivan Perićand Rajko Roki. On bounds for weighted norms for matrices and integral operators. Linear Algebra and its Applications, 326(1–3): 121–135, March 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/148/24/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/148/24/31/article.
pdf.

Petersen:2002:MNS

[PR02a] M. A. Petersen and A. C. M. Ran. Minimal nonsquare spectral factors. Linear Algebra and its Applications, 351–352:553–565, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Petersen:2002:NSF

[PR02b] M. A. Petersen and A. C. M. Ran. Nonsquare spectral factors via factorizations of a unitary function. Linear Algebra and its Applications, 351–352:567–583, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pittenger:2002:CSP

[PR02c] Arthur O. Pittenger and Morton H. Rubin. Convexity and the separability problem of quantum mechanical density matrices. Linear Algebra and its Applications, 346(1–3):47–71, May 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/189/27/
30/abstract.html.

Peric:2004:GIC

[PR04a] Ivan Perić and Rajna Rajić. Grüss inequality for completely bounded maps. Linear Algebra and its Applications, 390(1):287–292, October 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pittenger:2004:MUB

[PR04b] Arthur O. Pittenger and Morton H. Rubin. Mutually unbiased bases, generalized spin matrices and separability. Linear Algebra and its Applications, 390(1):255– 278, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pollatsek:2004:PIC

[PR04c] Harriet Pollatsek and Mary Beth Ruskai. Permutationally invariant codes for quantum error correction. Linear Algebra and its Applications, 392(1): 255–288, November 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pecaric:2007:DWE

[PR07] Josip Pečarić and Rajna Rajić. The Dunkl–Williams equality in pre-Hilbert C^{∗}- modules. Linear Algebra and its Applications, 425(1):16– 25, August 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Prajitura:2002:WFD

[Prǎ02] Gabriel T. Prǎjiturǎ. When is HyperlatT = Hyperlatf(T) in finite dimension? Linear Algebra and its Applications, 357(1–3):209–215, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
39/abstract.html.

Previtali:2005:UAQ

[Pre05] Andrea Previtali. Unitriangular actions on quadratic forms and character degrees. Linear Algebra and its Applications, 408(1):120–150, October 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Prokip:2005:SMC

[Pro05] Volodymyr Prokip. On similarity of matrices over commutative rings. Linear Algebra and its Applications, 399 (1):225–233, April 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Protasov:2008:ENL

[Pro08] V. Yu. Protasov. Extremal L_{p}-norms of linear operators and self-similar functions. Linear Algebra and its Applications, 428(10):2339– 2356, May 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Phatak:2002:AVU

[PRP02] A. Phatak, P. M. Reilly, and A. Penlidis. The asymptotic variance of the univariate PLS estimator. Linear Algebra and its Applications, 354 (1–3):245–253, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
46/abstract.html.

Passi:2002:IJD

[PRS02] Inder Bir S. Passi, Klaus W. Roggenkamp, and Marcos Soriano. Integral Jordan decomposition of matrices. Linear Algebra and its Applications, 355(1–3):241–261, November 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/201/27/
42/abstract.html.

Pereira:2005:ATS

[PRV05] Ricardo Pereira, Paula Rocha, and Paolo Vettori. Algebraic tools for the study of quaternionic behavioral systems. Linear Algebra and its Applications, 400(1):121– 140, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pan:2001:SMN

[PRW01] Victor Y. Pan, Youssef Rami, and Xinmao Wang. Structured matrices and Newton’s iteration: unified approach. Linear Algebra and its Applications, 343–344(1): 233–265, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/

Pryde:2009:CSE

[Pry09a] A. J. Pryde. Comparison of subdominant eigenvalues of some linear search schemes. Linear Algebra and its Applications, 431(9):1439–1442, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pryporova:2009:QCM

[Pry09b] Olga Pryporova. Qualitative convergence of matrices. Linear Algebra and its Applications, 431(1–2): 28–38, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Przybylska:2002:MDE

[Prz02] Maria Przybylska. On matrix differential equations and abstract FG algorithm. Linear Algebra and its Applications, 346(1–3):155–175, May 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/189/27/
35/abstract.html.

Petek:2003:SCP

[PS03] Tatjana Petek and Humberto Sarria. Spectrum and commutativity preserving mappings on H_{2}. Linear Algebra and its Applications, 364 (C):317–319, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
42/abstract.html.

Petkovic:2007:IAC

[PS07] Marko D. Petković and Predrag S. Stanimirović. Interpolation algorithm for computing Drazin inverse of polynomial matrices. Linear Algebra and its Applications, 422(2–3):526–539, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Parlett:2008:MPR

[PS08] Beresford Parlett and Gilbert Strang. Matrices with prescribed Ritz values. Linear Algebra and its Applications, 428(7):1725–1739, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Psarrakos:2000:NRL

[Psa00] Panayiotis J. Psarrakos. Numerical range of linear pencils. Linear Algebra and its Applications, 317(1–3):127– 141, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/136/25/

30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/136/25/30/article.
pdf.

Powell:2005:ESL

[PSvdW05] James A. Powell, Ivan Slapničar, and Wopke van der Werf. Epidemic spread of a lesion-forming plant pathogen—analysis of a mechanistic model with infinite age structure. Linear Algebra and its Applications, 398(1):117–140, March 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Puntanen:2000:P

[PSW00] Simo Puntanen, George P. H. Styan, and Hans Joachim Werner. Preface. Linear Algebra and its Applications, 321(1–3):1–2, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/24/article.
pdf.

Puntanen:2002:P

[PSW02] Simo Puntanen, George P. H. Styan, and Hans Joachim Werner. Preface. Linear Algebra and its Applications, 354(1–3):1–2, October 15, 2002. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/200/27/
                                        28/abstract.html.

Puntanen:2004:P

[PSW04] Simo Puntanen, George P. H. Styan, and Hans Joachim Werner. Preface. Linear Algebra and its Applications, 388(1):1–2, September 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Puntanen:2005:P

[PSW05] Simo Puntanen, George P. H. Styan, and Hans Joachim Werner. Preface. Linear Algebra and its Applications, 410(1):1–2, November 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Psarrakos:2003:PEN

[PT03] Panayiotis J. Psarrakos and Michael J. Tsatsomeros. The Perron eigenspace of nonnegative almost skew-symmetric matrices and Levinger’s transformation. Linear Algebra and its Applications, 360 (C):43–57, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/209/25/
                                        29/abstract.html.

Psarrakos:2004:PPF

[PT04] Panayiotis J. Psarrakos and Michael J. Tsatsomeros. A primer of Perron–Frobenius theory for matrix polynomials. Linear Algebra and its Applications, 393(1):333–351, December 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Plonka:2005:FNS

[PT05] Gerlind Plonka and Manfred Tasche. Fast and numerically stable algorithms for discrete cosine transforms. Linear Algebra and its Applications, 394(1):309– 345, January 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pascasio:2006:PCS

[PT06a] Arlene A. Pascasio and Paul Terwilliger. The pseudo-cosine sequences of a distance-regular graph. Linear Algebra and its Applications, 419(2–3):532–555, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Psarrakos:2006:BLF

[PT06b] Panayiotis J. Psarrakos and Michael J. Tsatsomeros. Bounds for Levinger’s function of nonnegative almost skew-symmetric matrices.

Linear Algebra and its Applications, 416(2–3):759–772, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Puerta:2001:CRI

[Pue01] Ferran Puerta. Conference report: 8th ILAS Conference in Barcelona. Linear Algebra and its Applications, 332–334 (1):607–609, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
57/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/57/article.
pdf.

Puerta:2005:VDG

[Pue05] X. Puerta. Versal deformations in generalized flag manifolds. Linear Algebra and its Applications, 401(1):445– 451, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Puente:2007:CCB

[Pue07] Rubén Puente. Cyclic convex bodies and optimization moment problems. Linear Algebra and its Applications, 426(2–3):596–609, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pugliese:2004:NMP

[Pug04] Paolo Pugliese. On the normal matrix of the polynomial LS problem over the Chebyshev points. Linear Algebra and its Applications, 378(1): 61–69, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pultarova:2007:LCA

[Pul07] Ivana Pultarová. Local convergence analysis of iterative aggregation–disaggregation methods with polynomial correction. Linear Algebra and its Applications, 421(1): 122–137, February 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pastravanu:2006:GMD

[PV06a] Octavian Pastravanu and Mihail Voicu. Generalized matrix diagonal stability and linear dynamical systems. Linear Algebra and its Applications, 419(2–3):299–310, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pereira:2006:ISS

[PV06b] Rajesh Pereira and Mohammad Ali Vali. Inequalities for the spectra of symmetric doubly stochastic matrices. Linear Algebra and its Applications, 419(2–3):643–647, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Plestenjak:2008:CAP

[PVV08] Bor Plestenjak, Marc Van Barel, and Ellen Van Camp. A Cholesky LR algorithm for the positive definite symmetric diagonal-plus-semiseparable eigenproblem. Linear Algebra and its Applications, 428(2–3):586–599, January 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Peng:2002:SFH

[PW02] Irine Peng and Shayne Waldron. Signed frames and Hadamard products of Gram matrices. Linear Algebra and its Applications, 347(1– 3):131–157, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
36/abstract.html.

Petriu:2004:AMV

[PW04] Dorina C. Petriu and C. Murray Woodside. Approximate mean value analysis based on Markov chain aggregation by composition. Linear Algebra and its Applications, 386(1): 335–358, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-

Pan:2008:DSI

[PW08a] Victor Y. Pan and Xinmao Wang. Degeneration of structured integer matrices modulo an integer. Linear Algebra and its Applications, 429(8–9):2113–2130, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Plischke:2008:DRJ

[PW08b] Elmar Plischke and Fabian Wirth. Duality results for the joint spectral radius and transient behavior. Linear Algebra and its Applications, 428(10):2368–2384, May 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Pronzato:2005:KTI

[PWZ05] Luc Pronzato, Henry P. Wynn, and Anatoly Zhigljavsky. Kantorovich-type inequalities for operators via D-optimal design theory. Linear Algebra and its Applications, 410(1):160–169, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Peng:2003:NSC

[PX03] Jigen Peng and Zong-Ben Xu. A note on Söderlind’s conjecture. Linear Algebra and its Applications, 367 (C):47–57, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
30/abstract.html.

Pan:2009:APE

[PY09] Victor Y. Pan and Xiaodong Yan. Additive preconditioning, eigenspaces, and the inverse iteration. Linear Algebra and its Applications, 430 (1):186–203, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379508003492.

Pan:2000:SAC

[PZ00] Victor Y. Pan and Ailong Zheng. Superfast algorithms for Cauchy-like matrix computations and extensions. Linear Algebra and its Applications, 310 (1–3):83–108, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/127/25/33/article.
pdf.

Qiu:2004:UDA

[QC04] Li Qiu and Tongwen Chen.

Unitary dilation approach to contractive matrix completion. Linear Algebra and its Applications, 379(1):345– 352, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Queiro:2009:ICN

[QD09] João Filipe Queiró and António Leal Duarte. Imbedding conditions for normal matrices. Linear Algebra and its Applications, 430(7): 1806–1811, April 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Qi:2009:ALL

[QH09] Xiaofei Qi and Jinchuan Hou. Additive Lie (ξ-Lie) derivations and generalized Lie (ξ-Lie) derivations on nest algebras. Linear Algebra and its Applications, 431(5–7):843– 854, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Qi:2009:PSM

[Qi09] Houduo Qi. Positive semidefinite matrix completions on chordal graphs and constraint nondegeneracy in semidefinite programming. Linear Algebra and its Applications, 430(4):1151–1164, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Queyranne:2006:BLI

[QT06] Maurice Queyranne and Fabio Tardella. Bimonotone linear inequalities and sublattices of R^{n}. Linear Algebra and its Applications, 413(1): 100–120, February 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Quint:2007:NRS

[Qui07]  Thomas Quint.  A new

rating system for duplicate bridge. Linear Algebra and its Applications, 422(1):236– 249, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Qu:2008:NHR

[QX08] Biao Qu and Naihua Xiu. A new halfspace-relaxation projection method for the split feasibility problem. Linear Algebra and its Applications, 428(5–6):1218–1229, March 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Qiu:2007:SHP

[QZ07] Li Qiu and Xingzhi Zhan. On the span of Hadamard products of vectors. Linear Algebra and its Applications, 422 (1):304–307, April 1, 2007. CODEN LAAPAW. ISSN

Rabta:2008:SSP

[RA08a] Boualem Rabta and Djamil Aïssani. Strong stability and perturbation bounds for discrete Markov chains. Linear Algebra and its Applications, 428(8–9):1921–1927, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Revesz:2008:EAC

[RA08b] Sz. Gy. Révész and A. San Antolín. Equivalence of Aapproximate continuity for self-adjoint expansive linear maps. Linear Algebra and its Applications, 429(7):1504– 1521, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rabanovich:2004:EML

[Rab04] Vyacheslav Rabanovich. Every matrix is a linear combination of three idempotents. Linear Algebra and its Applications, 390(1):137– 143, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rada:2003:CPC

[Rad03a] Juan Rada. Characteristic polynomial of catacondensed systems. Linear Algebra and its Applications, 367 (C):243–253, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
43/abstract.html.

Radjabalipour:2003:AMN

[Rad03b] M. Radjabalipour. Additive mappings on von Neumann algebras preserving absolute values. Linear Algebra and its Applications, 368 (C):229–241, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
40/abstract.html.

Rada:2009:MIE

[Rad09] Juan Rada. The McClelland inequality for the energy of digraphs. Linear Algebra and its Applications, 430(2–3): 800–804, January 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ralha:2003:OSR

[Ral03] Rui Ralha. One-sided reduction to bidiagonal form. Linear Algebra and its Applications, 358(1–3):219–238, January 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/207/25/
37/abstract.html.

Rao:2000:SPS

[Rao00] C. Radhakrishna Rao. Statistical proofs of some matrix inequalities. Linear Algebra and its Applications, 321 (1–3):307–320, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
44/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/44/article.
pdf.

Rao:2009:PIM

[Rao09] K. P. S. Bhaskara Rao. Products of idempotent matrices over integral domains. Linear Algebra and its Applications, 430(10):2690–2695, May 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ramezani:2009:NSC

[RBTR09] F. Ramezani, N. Broojerdian, and B. Tayfeh-Rezaie. A note on the spectral characterization of θ-graphs. Linear Algebra and its Applications, 431(5–7):626–632, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ren:2003:ISO

[RDxC03] Fang-Guo Ren, Hong-Ke Du, and Huai xin Cao. The intersection of the spectra of operator completions. Linear Algebra and its Applications, 371(1):103–109, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rebaza:2007:SSF

[Reb07] Jorge Rebaza. Smooth Schur factorizations in the continuation of separatrices. Linear Algebra and its Applications, 421(1):138–156, February 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Renaud:2001:MFG

[Ren01] P. F. Renaud. A matrix formulation of Grüss inequality. Linear Algebra and its Applications, 335(1–3): 95–100, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/165/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/165/27/31/article.
pdf.

Renes:2007:ETF

[Ren07] Joseph M. Renes. Equiangular tight frames from Paley tournaments. Linear Algebra and its Applications, 426(2–3):497–501, October 15, 2007. CODEN LAAPAW. ISSN 0024-

Reurings:2006:CMN

[Reu06] Martine C. B. Reurings. Contractive maps on normed linear spaces and their applications to nonlinear matrix equations. Linear Algebra and its Applications, 418(1): 292–311, October 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Remm:2003:ASA

[RG03] Elisabeth Remm and Michel Goze. Affine structures on abelian Lie groups. Linear Algebra and its Applications, 360(C):215–230, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
39/abstract.html.

Rosales:2001:PIF

[RGG01] J. C. Rosales and J. I. García-García. Primary ideals of finitely generated commutative cancellative monoids. Linear Algebra and its Applications, 336(1–3):219–230, October 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
42/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/166/27/42/article.
                                        pdf.

Rieck:2000:MOP

[Rie00]  M. Q. Rieck.  Maximal

orthogonality and pseudoorthogonality with applications to generalized inverses. Linear Algebra and its Applications, 315(1–3): 155–173, August 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/132/25/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/132/25/34/article.
                                        pdf.

Rihm:2001:AIM

[Rih01]  Robert Rihm.  Accelera-

tion of iteration methods for interval fixed point problems. Linear Algebra and its Applications, 324(1–3): 189–207, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/144/24/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/144/24/34/article.
                                        pdf.

Rowlinson:2002:SCS

[RJ02] P. Rowlinson and P. S. Jackson. Star complements and switching in graphs. Linear Algebra and its Applications, 356(1–3):135–143, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/202/25/
37/abstract.html.

Raissouli:2000:CMC

[RK00] Mustapha Raissouli and Ali Kacha. Convergence of matrix continued fractions. Linear Algebra and its Applications, 320(1–3):115–129, November 15, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/32/article.
pdf.

Raissouli:2003:CFE

[RL03] Mustapha Raïssouli and Fatima Leazizi. Continued fraction expansion of the geometric matrix mean and applications. Linear Algebra and its Applications, 359(1– 3):37–57, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
29/abstract.html.

Ren:2001:SSN

[RM01] Bin Ren and Dao Ji Meng. Some 2-step nilpotent Lie algebras I. Linear Algebra and its Applications, 338(1– 3):77–98, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/168/27/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/168/27/33/article.
                                        pdf.

Rojo:2006:TBA

[RM06] Oscar Rojo and Luis Medina. Tight bounds on the algebraic connectivity of Bethe trees. Linear Algebra and its Applications, 418(2–3):840– 853, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2009:SGB

[RM09a] Oscar Rojo and Luis Medina. Spectra of generalized Bethe trees attached to a path. Linear Algebra and its Applications, 430(1):483– 503, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rommes:2009:ESL

[RM09b] Joost Rommes and Nelson Martins. Exploiting structure in large-scale electrical circuit and power system problems. Linear Algebra and its Applications, 431 (3–4):318–333, July 15, 2009.

Rogers:2002:KPB

[RO02] E. Rogers and D. H. Owens. Kronecker product based stability tests and performance bounds for a class of 2D continuous-discrete linear systems. Linear Algebra and its Applications, 353(1–3): 33–52, September 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Robinson:2002:OIM

[Rob02] Donald W. Robinson. Outer inverses of matrices. Linear Algebra and its Applications, 348(1–3):247–258, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
45/abstract.html.

Robinson:2005:CA

[Rob05] Donald W. Robinson. The classical adjoint. Linear Algebra and its Applications, 411(1):254–276, December 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Robinson:2006:AAU

[Rob06] Donald W. Robinson. An alternative approach to unitoidness. Linear Algebra and its Applications, 413(1):72– 80, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Roche:2009:JIM

[Roc09] Jennifer L. Roche. On Jordan ideals and matrix rings. Linear Algebra and its Applications, 431(9):1675–1680, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rodriguez:2005:SAR

[Rod05] J. A. Rodríguez. A spectral approach to the Randić index. Linear Algebra and its Applications, 400(1):339– 344, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rodman:2006:SVU

[Rod06]  Leiba Rodman.  Similar-

ity vs unitary similarity and perturbation analysis of sign characteristics: Complex and real indefinite inner products. Linear Algebra and its Applications, 416(2–3):945– 1009, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rodman:2007:CFM

[Rod07] Leiba Rodman. Canonical forms for mixed symmetric– skewsymmetric quaternion matrix pencils. Linear Algebra and its Applications, 424(1):184–221, July 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rodman:2008:PHS

[Rod08]  Leiba Rodman.  Pairs

of hermitian and skewhermitian quaternionic matrices: Canonical forms and their applications. Linear Algebra and its Applications, 429(5–6):981–1019, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rodman:2009:RCM

[Rod09] Leiba Rodman. Review of: J-contractive Matrix Valued Functions and Related Topics, by Damir Z. Arov and Harry Dym. Cambridge University Press (2008), xi + 575 pages, ISBN-13 978-0-521-88300-9, hardback. Linear Algebra and its Applications, 431(9):1702–1705, October 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2002:SLM

[Roj02]  Oscar Rojo.  The spec-

trum of the Laplacian matrix of a balanced binary tree. Linear Algebra and its Applications, 349(1–3):203– 219, July 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2005:IBL

[Roj05]  Oscar Rojo.  Improved

bounds for the largest eigenvalue of trees. Linear Algebra and its Applications, 404(1): 297–304, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2006:SCR

[Roj06a] Oscar Rojo. On the spectra of certain rooted trees. Linear Algebra and its Applications, 414(1):218–243, April 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2006:SST

[Roj06b] Oscar Rojo. The spectra of some trees and bounds for the largest eigenvalue of any tree. Linear Algebra and its Applications, 414(1):199– 217, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2007:NUB

[Roj07a] Oscar Rojo. A nontrivial upper bound on the largest Laplacian eigenvalue of weighted graphs. Linear Algebra and its Applications, 420(2–3):625–633, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2007:LET

[Roj07b] Oscar Rojo. On the sth Laplacian eigenvalue of trees of order st + 1. Linear Algebra and its Applications, 425 (1):143–149, August 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2007:SGO

[Roj07c] Oscar Rojo. The spectra of a graph obtained from copies of a generalized Bethe tree. Linear Algebra and its Applications, 420(2–3):490–507, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2008:NUB

[Roj08a] Oscar Rojo. New upper bounds on the spectral radius of unicyclic graphs. Linear Algebra and its Applications, 428(4):754–764, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2008:SWG

[Roj08b]  Oscar Rojo.  Spectra of

weighted generalized Bethe trees joined at the root. Linear Algebra and its Applications, 428(11–12):2961– 2979, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2009:CTU

[Roj09a] Oscar Rojo. Computing tight upper bounds on the algebraic connectivity of certain graphs. Linear Algebra and its Applications, 430(1):532– 543, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2009:SCG

[Roj09b]  Oscar Rojo.  Spectra of

copies of a generalized Bethe tree attached to any graph. Linear Algebra and its Applications, 431(5–7):863–882, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rosales:2005:FGN

[Ros05] J. C. Rosales. Fundamental gaps of numerical semigroups generated by two elements. Linear Algebra and its Applications, 405(1):200– 208, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rost:2006:GHN

[Ros06] Karla Rost. Georg Heinig November 24, 1947–May 10, 2005: A personal memoir and appreciation by Karla Rost. Linear Algebra and its Applications, 413(1):1–12, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rosales:2009:ASN

[Ros09] J. C. Rosales. Atoms of the set of numerical semigroups with fixed Frobenius number. Linear Algebra and its Applications, 430(1):41– 51, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rowlinson:2007:CCS

[Row07] P. Rowlinson. Co-cliques and star complements in extremal strongly regular graphs. Linear Algebra and its Applications, 421(1):157– 162, February 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rodman:2000:LPM

[RP00] Leiba Rodman and PeterŠemrl. Linear preservers of minimal rank. Linear Algebra and its Applications, 310(1–3):73–82, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/127/25/32/article.
                                        pdf.

Rojo:2003:NIE

[RP03] Oscar Rojo and Matilde Peña. A note on the integer eigenvalues of the Laplacian matrix of a balanced binary tree. Linear Algebra and its Applications, 362 (C):293–300, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/211/25/
                                        49/abstract.html.

Rajagopal:2005:GEN

[RP05] Ravikiran Rajagopal and Lee C. Potter. A generic effective Nullstellensatz. Linear Algebra and its Applications, 397(1):1–15, March 1, 2005. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ran:2002:NME

[RR02] André C. M. Ran and Martine C. B. Reurings. On the nonlinear matrix equation X + A^{∗}F (X)A = Q: solutions and perturbation theory. Linear Algebra and its Applications, 346(1–3): 15–26, May 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/

gej-ng/10/30/19/189/27/
28/abstract.html.

Ran:2004:NME

[RR04a] André C. M. Ran and Martine C. B. Reurings. A nonlinear matrix equation connected to interpolation theory. Linear Algebra and its Applications, 379(1):289– 302, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2004:DSU

[RR04b] Oscar Rojo and Héctor Rojo. A decreasing sequence of upper bounds on the largest Laplacian eigenvalue of a graph. Linear Algebra and its Applications, 381(1):97– 116, April 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2004:SRS

[RR04c] Oscar Rojo and Héctor Rojo. Some results on symmetric circulant matrices and on symmetric centrosymmetric matrices. Linear Algebra and its Applications, 392(1): 211–233, November 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Renardy:2005:DSA

[RR05] M. Renardy and D. L. Russell. A discrete systems approach to cardinal spline Hermite interpolation. Linear Algebra and its Applications, 406(1):77–98, September 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2007:EFE

[RR07a] Oscar Rojo and María Robbiano. An explicit formula for eigenvalues of Bethe trees and upper bounds on the largest eigenvalue of any tree. Linear Algebra and its Applications, 427(1):138– 150, November 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2007:SSW

[RR07b] Oscar Rojo and María Robbiano. On the spectra of some weighted rooted trees and applications. Linear Algebra and its Applications, 420(2–3):310–328, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ravi:2002:OFI

[RRH02] M. S. Ravi, Joachim Rosenthal, and Uwe Helmke. Output feedback invariants. Linear Algebra and its Applications, 351–352:623–637, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rohn:2001:P

[RRY01] J. Rohn, S. M. Rump, and T. Yamamoto. Preface. Linear Algebra and its Applications, 324(1–3):1–2, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/144/24/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/144/24/24/article.
pdf.

Radjabalipour:2004:BTM

[RRY04] M. Radjabalipour, P. Rosenthal, and B. R. Yahaghi. Burnside’s theorem for matrix rings over division rings. Linear Algebra and its Applications, 383(1):29–44, May 15, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Robbin:2000:EVM

[RS00] Joel W. Robbin and Dietmar A. Salamon. The exponential Vandermonde matrix. Linear Algebra and its Applications, 317(1–3):225– 226, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/136/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/136/25/37/article.
pdf.

Rovnyak:2001:SIC

[RS01] J. Rovnyak and L. A. Sakhnovich. Some indefinite cases of spectral problems for canonical systems of difference equations. Linear Algebra and its Applications, 343–344(1):267–289, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/174/27/
                                        38/abstract.html.

Robbe:2002:PEB

[RS02] M. Robbé and M. Sadkane. A priori error bounds on invariant subspace approximations by block Krylov subspaces. Linear Algebra and its Applications, 350(1–3): 89–103, July 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Resta:2003:NDP

[RS03a] Giovanni Resta and Giovanni Sburlati. On the number of different permanents of some sparse (0, 1)-circulant matrices. Linear Algebra and its Applications, 375(1):197–209, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2003:ECN

[RS03b] Oscar Rojo and Ricardo L. Soto. Existence and construction of nonnegative matrices with complex spectrum. Linear Algebra and its Applications, 368(C):53–69, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
28/abstract.html.

Rodman:2005:MFP

[RS05a] L. Rodman and I. M. Spitkovsky. 3 × 3 matrices with a flat portion on the boundary of the numerical range. Linear Algebra and its Applications, 397(1):193– 207, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2005:SAM

[RS05b] Oscar Rojo and Ricardo Soto. The spectra of the adjacency matrix and Laplacian matrix for some balanced trees. Linear Algebra and its Applications, 403(1): 97–117, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rota:2005:PCH

[RS05c] Gian-Carlo Rota and Joel Alvin Stein. A problem of Cayley from 1857 and how he could have solved it. Linear Algebra and its Applications, 411(1): 167–253, December 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Resta:2006:CPD

[RS06a] Giovanni Resta and Giovanni Sburlati. Computing permanents via determinants for some classes of sparse matrices. Linear Algebra and its Applications, 418(1):33– 43, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Robbe:2006:EIB

[RS06b] Mickaël Robbé and Miloud Sadkane. Exact and inexact breakdowns in the block GM-RES method. Linear Algebra and its Applications, 419(1): 265–285, November 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2009:GPS

[RS09] Oscar Rojo and Ricardo L. Soto. Guo perturbation for symmetric nonnegative circulant matrices. Linear Algebra and its Applications, 431(5– 7):594–607, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2000:ANU

[RSR00] Oscar Rojo, Ricardo Soto, and Héctor Rojo. An always nontrivial upper bound for Laplacian graph eigenvalues. Linear Algebra and its Applications, 312(1–3): 155–159, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/129/25/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/129/25/34/article.
pdf.

Rajeshwari:2000:NFC

[RSS00] K. N. Rajeshwari, K. Santaram, and Pramod K. Sharma. A note on feedback cyclicity of C[Y]. Linear Algebra and its Applications, 320(1–3):1–13, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/24/article.
pdf.

Radjabalipour:2001:AMO

[RST01] M. Radjabalipour, K. Seddighi, and Y. Taghavi. Additive mappings on operator algebras preserving absolute values. Linear Algebra and its Applications, 327(1–3):197– 206, April 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/40/abstract.
html;   http://www.elsevier.

                                        nl/gej%2Dng/10/30/19/152/
                                        25/40/article.pdf.

Ruskai:2002:ACP

[RSW02] Mary Beth Ruskai, Stanislaw Szarek, and Elisabeth Werner. An analysis of completely-positive tracepreserving maps on M_{2}. Linear Algebra and its Applications, 347(1–3):159–187, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/194/27/
                                        37/abstract.html.

Ren:2005:ABS

[RSW05] Guojing Ren, Yuming Shi, and Yi Wang. Asymptotic behavior of solutions of perturbed linear difference systems. Linear Algebra and its Applications, 395(1):283– 302, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rognes:2001:SCM

[RT01a] Øystein Rognes and Xue-Cheng Tai. Subspace correction methods for nonsymmetric parabolic problems. Linear Algebra and its Applications, 332–334(1): 205–234, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/157/25/

38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/38/article.
pdf.

Romeo:2001:SVP

[RT01b] Mario Romeo and Paolo Tilli. On some variational problems in the theory of unitarily invariant norms and Hadamard products. Linear Algebra and its Applications, 326(1–3):161–172, March 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/148/24/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/148/24/34/article.
pdf.

Rada:2003:PCM

[RT03] Juan Rada and Antonio Tineo. Polygonal chains with minimal energy. Linear Algebra and its Applications, 372 (1):333–344, October 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rejto:2006:RKN

[RT06] Lídia Rejtő and Gábor Tusnády. Reconstruction of Kauffman networks applying trees. Linear Algebra and its Applications, 417(1):220– 244, August 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rukhin:2001:PCM

[Ruk01] Andrew L. Rukhin. Pattern correlation matrices and their properties. Linear Algebra and its Applications, 327(1–3):105–114, April 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/152/25/34/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/152/
25/34/article.pdf.

Rump:2001:CEB

[Rum01a] Siegfried M. Rump. Computational error bounds for multiple or nearly multiple eigenvalues. Linear Algebra and its Applications, 324(1–3): 209–226, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/144/24/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/144/24/35/article.
pdf.

Rump:2001:SVM

[Rum01b] Siegfried M. Rump. Selfvalidating methods. Linear Algebra and its Applications, 324(1–3):3–13, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/144/24/

25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/144/24/25/article.
pdf.

Rump:2003:M

[Rum03a] Siegfried M. Rump. On P -matrices. Linear Algebra and its Applications, 363 (C):237–250, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
43/abstract.html.

Rump:2003:PFT

[Rum03b] Siegfried M. Rump. Perron– Frobenius theory for complex matrices. Linear Algebra and its Applications, 363 (C):251–273, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
44/abstract.html.

Rump:2006:EPS

[Rum06] Siegfried M. Rump. Eigenvalues, pseudospectrum and structured perturbations. Linear Algebra and its Applications, 413(2–3):567–593, March 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Raats:2005:MRC

[RvdGM05] V. M. Raats, B. B. van der Genugten, and J. J. A.

Moors. Multivariate regression with consecutively added dependent variables. Linear Algebra and its Applications, 410(1):170–197, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Reyes:2000:NAT

[RVZ00] Enrique Reyes, Rafael H. Villarreal, and Leticia Zárate. A note on affine toric varieties. Linear Algebra and its Applications, 318(1–3): 173–179, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/137/24/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/137/24/35/article.
pdf.

Radkova:2009:CCI

[Rvz09] D. Radkova and A. J. van zanten. Constacyclic codes as invariant subspaces. Linear Algebra and its Applications, 430(2–3):855–864, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Roorda:2001:OAR

[RW01] Berend Roorda and Siep Weiland. Optimal angle reduction — a behavioral approach to linear system approximation. Linear Algebra and its Applications, 337(1–3): 189–235, November 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/167/27/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/167/27/36/article.
pdf.

Rakocevic:2002:WDI

[RW02] Vladimir Rakočević and Yimin Wei. A weighted Drazin inverse and applications. Linear Algebra and its Applications, 350(1–3): 25–39, July 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rewienski:2006:MOR

[RW06] Michał Rewieński and Jacob White. Model order reduction for nonlinear dynamical systems based on trajectory piecewise-linear approximations. Linear Algebra and its Applications, 415 (2–3):426–454, June 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2002:LCA

[rWQ02] Guo rong Wang and Lin Qiu. Leverrier-Chebyshev algorithm for the singular pencils. Linear Algebra and its Applications, 345 (1–3):1–8, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/175/27/
                                        27/abstract.html.

Radjabalipour:2008:OSI

[RY08] Mehdi Radjabalipour and Bamdad R. Yahaghi. On one-sided ideals of rings of continuous linear operators. Linear Algebra and its Applications, 429(5–6):1184–1190, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Roca:2001:IFA

[RZ01] A. Roca and I. Zaballa. Invariant factor assignment under state feedback and output injection. Linear Algebra and its Applications, 332–334 (1):401–436, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/157/25/
                                        45/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/157/25/45/article.
                                        pdf.

Roca:2004:FIO

[RZ04] A. Roca and I. Zaballa. Feedback invariants via output injection. Linear Algebra and its Applications, 385(1):407– 442, July 1, 2004. CO-

Roca:2005:PAI

[RZ05] A. Roca and I. Zaballa. A polynomial approach to the invariant factors assignment problem via state feedback and output injection. Linear Algebra and its Applications, 401(1):135– 158, May 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dokovic:2008:QMU

[– DS08] Dragomir Ž. – Doković and Benjamin H. Smith. Quaternionic matrices: Unitary similarity, simultaneous triangularization and some trace identities. Linear Algebra and its Applications, 428(4):890– 910, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sorensen:2002:SEA

[SA02] D. C. Sorensen and A. C. Antoulas. The Sylvester equation and approximate balanced reduction. Linear Algebra and its Applications, 351–352:671–700, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Salam:2008:OSH

[SAAF08] A. Salam, E. Al-Aidarous, and A. El Farouk. Optimal symplectic Householder transformations for SR decomposition. Linear Algebra and its Applications, 429(5–6):1334–1353, September 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stevanovic:2008:SRG

[SAH08] Dragan Stevanović, Mustapha Aouchiche, and Pierre Hansen. On the spectral radius of graphs with a given domination number. Linear Algebra and its Applications, 428(8– 9):1854–1864, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sakhnovich:2001:SLI

[Sak01] Alexander Sakhnovich. Szegő limits for infinite Toeplitz matrices determined by the Taylor series of two rational functions. Linear Algebra and its Applications, 343–344 (1):291–302, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/174/27/
                                        39/abstract.html.

Sakhnovich:2004:ETP

[Sak04]  L. A. Sakhnovich.  Ex-

tremal trigonometric and power polynomials in several variables. Linear Algebra and its Applications, 389(1): 283–293, September 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Salam:2000:VHD

[Sal00] A. Salam. On vector Hankel determinants. Linear Algebra and its Applications, 313 (1–3):127–139, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/130/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/130/25/32/article.
pdf.

Salgado:2009:ASA

[Sal09] G. Salgado. Algebraic structures in F^{N} associated to linear transformations. Linear Algebra and its Applications, 431(5–7):991–997, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sato:2005:PCM

[Sat05]  Masahisa Sato.  Periodic

Coxeter matrices and their associated quadratic forms. Linear Algebra and its Applications, 406(1):99–108, September 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sato:2006:FLG

[Sat06] Iwao Sato. L-functions for line graphs of semiregular bipartite graphs. Linear Algebra and its Applications, 416 (2–3):373–388, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sato:2008:SWC

[Sat08] Iwao Sato. The stochastic weighted complexity of a group covering of a digraph. Linear Algebra and its Applications, 429(8–9):1905– 1914, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Savchenko:2003:PRP

[Sav03] S. V. Savchenko. On the Perron roots of principal submatrices of co-order one of irreducible nonnegative matrices. Linear Algebra and its Applications, 361 (C):257–277, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/210/25/
                                        43/abstract.html.

Savchenko:2004:PSC

[Sav04] S. V. Savchenko. Principal submatrices of co-order one with the biggest Perron root. Linear Algebra and its Applications, 389(1):1–22, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Savchenko:2005:CJF

[Sav05] S. V. Savchenko. On the change of the Jordan form under the transition from the adjacency matrix of a vertextransitive digraph to its principal submatrix of co-order one. Linear Algebra and its Applications, 394(1):225– 235, January 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Savas:2006:DRV

[Sav06a] Berkant Savas. Dimensionality reduction and volume minimization—generalization of the determinant minimization criterion for reduced rank regression problems. Linear Algebra and its Applications, 418(1):201– 214, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Savchenko:2006:NMT

[Sav06b] S. V. Savchenko. Normal matrices and their principal submatrices of co-order one. Linear Algebra and its Applications, 419(2–3):556–568, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Soto:2005:CSR

[SBM05] Ricardo Soto, Alberto Borobia, and Julio Moro. On the comparison of some realizability criteria for the real nonnegative inverse eigenvalue problem. Linear Algebra and its Applications, 396(1):223–241, February 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stuart:2002:IPP

[SBS02] Jeffrey L. Stuart, LeRoy Beasley, and Bryan Shader. Irreducible, pattern k-potent ray pattern matrices. Linear Algebra and its Applications, 346(1–3):261–271, May 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/189/27/
                                        39/abstract.html.

Sburlati:2005:VPC

[Sbu05] Giovanni Sburlati. On the values of permanents of (0, 1) circulant matrices with three ones per row. Linear Algebra and its Applications, 408 (1):284–297, October 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sburlati:2008:PPC

[Sbu08] Giovanni Sburlati. On the parity of permanents of circulant matrices. Linear Algebra and its Applications, 428(8– 9):1949–1955, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shi:2001:STH

[SC01] Yuming Shi and Shaozhu Chen. Spectral theory of higher-order discrete vector Sturm–Liouville problems. Linear Algebra and its Applications, 323(1–3):7– 36, January 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Serra-Capizzano:2006:GCG

[SC06]  Stefano  Serra-Capizzano.

The GLT class as a generalized Fourier analysis and applications. Linear Algebra and its Applications, 419(1): 180–233, November 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Serra-Capizzano:2007:SAM

[SC07]  Stefano  Serra-Capizzano.

The spectral approximation of multiplication operators via asymptotic (structured) linear algebra. Linear Algebra and its Applications, 424(1): 154–176, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Song:2008:NPP

[SC08] Xiaofei Song and Chongguang Cao. A note on k-potence preserving maps. Linear Algebra and its Applications, 429(7):1579–1586, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schmidt:2000:CJD

[Sch00a] Dieter Schmidt. Construction of the Jordan decomposition by means of Newton’s method. Linear Algebra and its Applications, 314 (1–3):75–89, July 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/131/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/131/25/27/article.
pdf.

Schneider:2000:IGA

[Sch00b] Hans Schneider. Introduction to a group of articles in tribute to Vlastimil Pták. Linear Algebra and its Applications, 310(1–3):1–4, May 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/127/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/127/25/25/article.
pdf.

Scherer:2002:SFD

[Sch02a] Carsten W. Scherer. Structured finite-dimensional controller design by convex optimization. Linear Algebra and its Applications, 351–352: 639–669, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schneider:2002:HWD

[Sch02b] Hans Schneider. Helmut Wielandt 19 December 1910– 14 February 2001: A personal memoir. Linear Algebra and its Applications, 353(1–3):1– 3, September 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schneider:2002:WPE

[Sch02c] Hans Schneider. Wielandt’s proof of the exponent inequality for primitive nonnegative matrices. Linear Algebra and its Applications, 353(1–3):5–10, September 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schmoeger:2003:OE

[Sch03] Christoph Schmoeger. On the operator equation e^{A} = e^{B}. Linear Algebra and its Applications, 359(1–3): 169–179, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/208/25/
                                        39/abstract.html.

Schafer:2004:NSM

[Sch04] Uwe Schäfer. A new subclass of P -matrices. Linear Algebra and its Applications, 393(1): 353–364, December 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schaffrin:2006:NCT

[Sch06] Burkhard Schaffrin. A note on Constrained Total Least-Squares estimation. Linear Algebra and its Applications, 417(1):245–258, August 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schmoeger:2008:CSC

[Sch08] Christoph Schmoeger. Characterizations of some classes of relatively regular operators. Linear Algebra and its Applications, 429(1):302– 310, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schilders:2009:SIL

[Sch09a] Wil H. A. Schilders. Solution of indefinite linear systems using an LQ decomposition for the linear constraints. Linear Algebra and its Applications, 431(3–4):

381–395, July 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schmoeger:2009:GPB

[Sch09b] Christoph Schmoeger. Generalized projections in Banach algebras. Linear Algebra and its Applications, 430(2–3):601–608, January 15, 2009. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schwartz:2009:ECP

[Sch09c] Moshe Schwartz. Efficiently computing the permanent and Hafnian of some banded Toeplitz matrices. Linear Algebra and its Applications, 430(4):1364–1374, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sciriha:2002:PRT

[Sci02] Irene Sciriha. Polynomial reconstruction and terminal vertices. Linear Algebra and its Applications, 356(1–3): 145–156, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/202/25/
38/abstract.html.

Sciriha:2009:GCE

[Sci09] Irene Sciriha. Graphs with a common eigenvalue deck. Linear Algebra and its Applications, 430(1):78–85, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Serra-Capizzano:2009:TED

[SCS09] Stefano Serra-Capizzano and Debora Sesana. Tools for the eigenvalue distribution in a non-Hermitian setting. Linear Algebra and its Applications, 430(1):423–437, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stanimirovic:2000:FRD

[SD00] Predrag S. Stanimirović and Dragan S. Djordjević. Fullrank and determinantal representation of the Drazin inverse. Linear Algebra and its Applications, 311(1–3): 131–151, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/128/25/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/128/25/34/article.
pdf.

Stevanovic:2007:WRI

[SdAdFDV07] Dragan Stevanović, Nair M. M. de Abreu, Maria A. A. de Freitas, and Renata Del-Vecchio. Walks and regular integral graphs. Linear Algebra and its Applications, 423(1):119–135, May 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Silva:2003:LNS

[SdL03] Manuela S. Silva and Teresa P. de Lima. Looking for nonnegative solutions of a Leontief dynamic model. Linear Algebra and its Applications, 364 (C):281–316, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
40/abstract.html.

Sun:2006:RHP

[SDL06] Xiaosong Sun, Xiankun Du, and Dayan Liu. On the range of a Hadamard power of a positive semidefinite matrix. Linear Algebra and its Applications, 416(2–3):868– 871, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sandovici:2009:IFP

[SdS09] Adrian Sandovici and Henk de Snoo. An index formula for the product of linear relations. Linear Algebra and its Applications, 431(11):2160– 2171, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sandovici:2005:SLR

[SdSW05] Adrian Sandovici, Henk de Snoo, and Henrik Winkler. The structure of linear relations in Euclidean spaces. Linear Algebra and its Applications, 397(1):141– 169, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sandovici:2007:ADN

[SdSW07] Adrian Sandovici, Henk de Snoo, and Henrik Winkler. Ascent, descent, nullity, defect, and related notions for linear relations in linear spaces. Linear Algebra and its Applications, 423 (2–3):456–497, June 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seddik:2001:NRE

[Sed01] Ameur Seddik. The numerical range of elementary operators II. Linear Algebra and its Applications, 338(1–3): 239–244, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/168/27/
                                        43/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/168/27/43/article.
                                        pdf.

Seddik:2004:SON

[Sed04] Ameur Seddik. On some operator norm inequalities. Linear Algebra and its Applications, 389(1):183–187, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seddik:2007:ROO

[Sed07] Ameur Seddik. Rank one operators and norm of elementary operators. Linear Algebra and its Applications, 424(1):177–183, July 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seddighin:2009:ATS

[Sed09] Morteza Seddighin. Antieigenvalue techniques in statistics. Linear Algebra and its Applications, 430(10):2566– 2580, May 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Semrl:2003:HFT

[Šem03] Peter Šemrl. Hua’s fundamental theorems of the geometry of matrices and related results. Linear Algebra and its Applications, 361 (C):161–179, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/10/30/19/210/25/
                                        38/abstract.html.

Semancikova:2006:OMM

[Sem06a] Blanka Semančíková. Orbits in max–min algebra. Linear Algebra and its Applications, 414(1):38–63, April 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Semrl:2006:MMS

[Šem06b] Peter Šemrl. Maps on matrix spaces. Linear Algebra and its Applications, 413(2– 3):364–393, March 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Semancikova:2007:OCC

[Sem07] Blanka Semančíková. Orbits and critical components of matrices in max–min algebra. Linear Algebra and its Applications, 426(2–3):415–447, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Semrl:2008:CPM

[Šem08] Peter Šemrl. Commutativity preserving maps. Linear Algebra and its Applications, 429(5–6):1051–1070, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sengupta:2004:KFP

[Sen04] Debasis Sengupta. On the Kalman filter with possibly degenerate and correlated errors. Linear Algebra and its Applications, 388(1):327–340, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sendov:2007:HOD

[Sen07] Hristo S. Sendov. The higherorder derivatives of spectral functions. Linear Algebra and its Applications, 424(1): 240–281, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seo:2000:SRC

[Seo00] Yuki Seo. Specht’s ratio and complementary inequalities to Golden-Thompson type inequalities on the Hadamard product. Linear Algebra and its Applications, 320(1–3): 15–22, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/25/article.
pdf.

Sergeichuk:2000:CML

[Ser00] Vladimir V. Sergeichuk. Canonical matrices for linear matrix problems. Linear Algebra and its Applications, 317(1–3):53–102, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/136/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/136/25/28/article.
pdf.

Sergeichuk:2004:CCM

[Ser04] Vladimir V. Sergeichuk. Computation of canonical matrices for chains and cycles of linear mappings. Linear Algebra and its Applications, 376(1):235–263, January 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sergeev:2007:MPD

[Ser07] Sergeĭ Sergeev. Max-plus definite matrix closures and their eigenspaces. Linear Algebra and its Applications, 421(2–3):182–201, March 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sergeichuk:2008:CMI

[Ser08] Vladimir V. Sergeichuk. Canonical matrices of isometric operators on indefinite inner product spaces. Linear Algebra and its Applications, 428(1):154–192, January 1, 2008. CO-

Sergeev:2009:MAP

[Ser09] Sergeĭ Sergeev. Max algebraic powers of irreducible matrices in the periodic regime: An application of cyclic classes. Linear Algebra and its Applications, 431(8):1325–1339, September 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Severo:2008:MCB

[Sev08] Norman C. Severo. Matrices with convolutions of binomial functions and Krawtchouk matrices. Linear Algebra and its Applications, 429(1): 50–56, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shao:2003:SPA

[SG03] Yanling Shao and Yubin Gao. Sign patterns that allow diagonalizability. Linear Algebra and its Applications, 359 (1–3):113–119, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
35/abstract.html.

Sznajder:2006:PCT

[SG06] Roman Sznajder and M. Seetharama Gowda. The Q-property of composite transformations and the P -property of Steintype transformations on selfdual and symmetric cones. Linear Algebra and its Applications, 416(2–3):437–451, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shao:2008:ECS

[SG08] Yanling Shao and Yubin Gao. Exponents of 2-coloring of symmetric digraphs. Linear Algebra and its Applications, 428(7):1538–1550, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stone:2009:FVG

[SG09] Eric A. Stone and Alexander R. Griffing. On the Fiedler vectors of graphs that arise from trees by Schur complementation of the Laplacian. Linear Algebra and its Applications, 431(10): 1869–1880, October 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shao:2008:OTC

[SGS08] Jia-Yu Shao, Ji-Ming Guo, and Hai-Ying Shan. The ordering of trees and connected graphs by algebraic connectivity. Linear Algebra and its Applications, 428 (7):1421–1438, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shao:2000:ACM

[SH00a] Jia-Yu Shao and Zhi-Xiang Hu. Algebraic constructions of the minimal forbidden digraphs of strong sign nonsingular matrices. Linear Algebra and its Applications, 317(1–3):1–12, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/136/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/136/25/25/article.
pdf.

Shao:2000:NMG

[SH00b] Jia-Yu Shao and Suk-Geun Hwang. Nearly L-matrices and generalized row sign balanced matrices. Linear Algebra and its Applications, 317 (1–3):41–52, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/136/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/136/25/27/article.
pdf.

Shao:2002:MDS

[SH02] Jia-Yu Shao and Jin-Ling He. Matrices with doubly signed generalized inverses. Linear Algebra and its Applications, 355(1–3): 71–84, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/201/27/
                                        32/abstract.html.

She:2005:NSP

[SH05] Yiyuan She and Pengwei Hao. On the necessity and sufficiency of PLUS factorizations. Linear Algebra and its Applications, 400(1):193– 202, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shen:2008:SGT

[SH08] Jin Shen and Yaoping Hou. On the sandpile group of 3 × n twisted bracelets. Linear Algebra and its Applications, 429(8–9):1894–1904, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Szanto:2009:FKI

[SH09] Csaba Szántó and Alexandru Horváth. Formulas for Kronecker invariants using a representation theoretical approach. Linear Algebra and its Applications, 430(2–3): 664–673, January 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shao:2000:DSS

[Sha00] Jia-Yu Shao. On the digraphs of sign solvable linear systems. Linear Algebra and its Applications, 313 (1–3):115–126, July 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/130/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/130/25/31/article.
pdf.

Shader:2003:SSS

[Sha03] Bryan L. Shader. Set-systems with signed solutions. Linear Algebra and its Applications, 361(C):121–132, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
35/abstract.html.

Sharifi:2009:DPI

[Sha09] Kamran Sharifi. Descriptions of partial isometries on Hilbert C^{∗}-modules. Linear Algebra and its Applications, 431(5–7):883–887, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shen:2000:CSM

[She00] Jianhong Shen. Compactification of a set of matrices with convergent infinite products. Linear Algebra and its Applications, 311(1– 3):177–186, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/128/25/
37/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/128/25/37/article.
pdf.

Shen:2001:SVG

[She01] Jianhong Shen. On the singular values of Gaussian random matrices. Linear Algebra and its Applications, 326 (1–3):1–14, March 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/148/24/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/148/24/24/article.
pdf.

Sherstnev:2003:TDR

[She03] Anatoly N. Sherstnev. Twodimensional reductions of the cone of positive diagonal operators in 2. Linear Algebra and its Applications, 370(1): 375–378, September 1, 2003.

Shih:2001:KCC

[Shi01] Mau-Hsiang Shih. König chain for compact matrix sets. Linear Algebra and its Applications, 330(1–3): 205–208, June 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/155/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/155/25/36/article.
pdf.

Shi:2004:RMR

[Shi04] Yuming Shi. On the rank of the matrix radius of the limiting set for a singular linear Hamiltonian system. Linear Algebra and its Applications, 376(1):109– 123, January 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shi:2006:WTT

[Shi06] Yuming Shi. Weyl–Titchmarsh theory for a class of discrete linear Hamiltonian systems. Linear Algebra and its Applications, 416(2–3):452– 519, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shi:2007:BLS

[Shi07a] Lingsheng Shi. Bounds on the (laplacian) spectral radius of graphs. Linear Algebra and its Applications, 422(2–3): 755–770, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shi:2007:SRG

[Shi07b] Lingsheng Shi. The spectral radii of a graph and its line graph. Linear Algebra and its Applications, 422(1): 58–66, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shi:2009:SRI

[Shi09] Lingsheng Shi. The spectral radius of irregular graphs. Linear Algebra and its Applications, 431(1–2):189–196, July 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shparlinski:2006:ESC

[Shp06] Igor Shparlinski. On the energy of some circulant graphs. Linear Algebra and its Applications, 414(1):378– 382, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Song:2003:EPM

[SHRC03] Seok-Zun Song, Suk-Geun Hwang, Seog-Hoon Rim, and Gi-Sang Cheon. Extremes of permanents of (0,1)-matrices. Linear Algebra and its Applications, 373(1):197–210, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shao:2003:NNE

[SHS03a] Jia-Yu Shao, Jin-Ling He, and Hai-Ying Shan. Number of nonzero entries of S^{2}NS matrices and matrices with signed generalized inverses. Linear Algebra and its Applications, 373(1):223– 239, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Skoug:2003:LHP

[SHS03b] Muriel J. Skoug, Richard D. Hill, and Joseph R. Siler. On a lattice of hermitianpreserving cones. Linear Algebra and its Applications, 367(C):291–300, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
46/abstract.html.

Shu:2002:SUB

[SHWR02] Jin-Long Shu, Yuan Hong, and Kai Wen-Ren. A sharp upper bound on the largest eigenvalue of the Laplacian matrix of a graph. Linear Algebra and its Applications, 347(1–3):123–129, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/194/27/
                                        35/abstract.html.

Shen:2005:GSG

[SHZ05] Xiaoling Shen, Yaoping Hou, and Yuanping Zhang. Graph Z_{n} and some graphs related to Z_{n} are determined by their spectrum. Linear Algebra and its Applications, 404(1): 58–68, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stevanovic:2009:LCU

[SI09] Dragan Stevanović and Aleksandar Ilić. On the Laplacian coefficients of unicyclic graphs. Linear Algebra and its Applications, 430(8–9): 2290–2300, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sidi:2001:DGT

[Sid01]  Avram Sidi.  DGMRES:

A GMRES-type algorithm for Drazin-inverse solution of singular non-symmetric linear systems. Linear Algebra and its Applications, 335 (1–3):189–204, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/

gej-ng/10/30/19/165/27/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/165/27/39/article.
pdf.

Silva:2000:NIF

[Sil00] Fernando C. Silva. On the number of invariant factors of partially prescribed matrices and control theory. Linear Algebra and its Applications, 311(1–3):1–12, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/128/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/128/25/25/article.
pdf.

Sivic:2008:VCT

[Šiv08] Klemen Šivic. On varieties of commuting triples. Linear Algebra and its Applications, 428(8–9):2006–2029, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sivasubramanian:2009:ADM

[Siv09] Sivaramakrishnan Sivasubramanian. q-Analogs of distance matrices of 3hypertrees. Linear Algebra and its Applications, 431(8):1234–1248, September 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sridhar:2000:PPM

[SK00] R. Sridhar and S. N. Kabadi. On pseudomatroid property of matrices. Linear Algebra and its Applications, 304(1– 3):33–43, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/118/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/118/25/26/article.
pdf.

Scherrer:2008:AIP

[SK08a] Alexander Scherrer and Karl-Heinz Küfer. Accelerated IMRT plan optimization using the adaptive clustering method. Linear Algebra and its Applications, 428(5– 6):1250–1271, March 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Suss:2008:BCS

[SK08b] Philipp Süss and Karl-Heinz Küfer. Balancing control and simplicity: A variable aggregation method in intensity modulated radiation therapy planning. Linear Algebra and its Applications, 428(5– 6):1388–1405, March 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Song:2008:RMT

[SKBS08] Seok-Zun Song, Kyung-Tae Kang, LeRoy B. Beasley, and Nung-Sing Sze. Regular matrices and their strong preservers over semirings. Linear Algebra and its Applications, 429(1):209–223, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Salimbahrami:2006:ORL

[SL06] Behnam Salimbahrami and Boris Lohmann. Order reduction of large scale secondorder systems using Krylov subspace methods. Linear Algebra and its Applications, 415(2–3):385–405, June 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Slapnicar:2003:HAS

[Sla03] Ivan Slapničar. Highly accurate symmetric eigenvalue decomposition and hyperbolic SVD. Linear Algebra and its Applications, 358(1–3): 387–424, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/207/25/
45/abstract.html.

Santos:2001:DFC

[SLP01] J. Santos, L. C. Longoria, and J. C. Palacios. Detection of flaws in construction columns using 3D reconstruction and manipulation. Linear Algebra and its Applications, 339(1–3): 195–204, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/169/27/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/169/27/38/article.
pdf.

Shao:2006:IPD

[SLR06] Jia-Yu Shao, Yue Liu, and Ling-Zhi Ren. The inverse problems of the determinantal regions of ray pattern and complex sign pattern matrices. Linear Algebra and its Applications, 416(2–3):835– 843, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shaked-Monderer:2009:NUB

[SM09a] Naomi Shaked-Monderer. A note on upper bounds on the cp-rank. Linear Algebra and its Applications, 431(12):2407–2413, December 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sweet:2009:MDS

[SM09b] L. G. Sweet and J. A. Mac-Dougall. The maximum dimension of a subspace of nilpotent matrices of index 2. Linear Algebra and its Applications, 431(8):1116–1124, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

belcastro:2002:MFP

[smbH02] sarah-marie belcastro and Thomas C. Hull. Modelling the folding of paper into three dimensions using affine transformations. Linear Algebra and its Applications, 348(1– 3):273–282, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
48/abstract.html.

Smith:2001:CM

[Smi01a] Ronald L. Smith. On characterizing Z-matrices. Linear Algebra and its Applications, 338(1–3):99–104, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/34/article.
pdf.

Smith:2001:FSM

[Smi01b] S. P. Smith. The factorability of symmetric matrices and some implications for statistical linear models. Linear Algebra and its Applications, 335(1–3): 63–80, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/165/27/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/165/27/29/article.
                                        pdf.

Smithies:2001:FRH

[Smi01c] Laura Smithies. Finite rank harmonic operator-valued functions. Linear Algebra and its Applications, 336(1– 3):15–20, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/166/27/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/166/27/28/article.
                                        pdf.

Smigoc:2004:IEP

[Šmi04] Helena Šmigoc. The inverse eigenvalue problem for nonnegative matrices. Linear Algebra and its Applications, 393(1):365–374, December 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Smith:2008:PDC

[Smi08a] Ronald L. Smith. The positive definite completion problem revisited. Linear Algebra and its Applications, 429(7): 1442–1452, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Smithies:2008:NPD

[Smi08b] Laura Smithies. A note on polar decomposition based geršgorin-type sets. Linear Algebra and its Applications, 429(10):2623–2627, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shorten:2009:APB

[SMK09] Robert Shorten, Oliver Mason, and Christopher King. An alternative proof of the Barker, Berman, Plemmons (BBP) result on diagonal stability and extensions. Linear Algebra and its Applications, 430(1):34–40, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Smoktunowicz:2008:BMS

[Smo08] Alicja Smoktunowicz. Block matrices and symmetric perturbations. Linear Algebra and its Applications, 429(10):2628–2635, November 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sumi:2009:RTS

[SMS09] Toshio Sumi, Mitsuhiro Miyazaki, and Toshio Sakata. Rank of 3-tensors with 2 slices and Kronecker canonical forms. Linear Algebra and its Applications, 431(10): 1858–1868, October 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shorten:2009:TRC

[SN09] Robert Shorten and Kumpati S. Narendra. On a theorem of Redheffer concerning diagonal stability. Linear Algebra and its Applications, 431(12):2317–2329, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

So:2004:HRE

[So04] Wasin So. The high road to an exponential formula. Linear Algebra and its Applications, 379(1):69–75, March 1, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sobczyk:2008:GMA

[Sob08] Garret Sobczyk. Geometric matrix algebra. Linear Algebra and its Applications, 429(5–6):1163–1173, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Solovev:2006:PIM

[Sol06] Sergey I. Solov’ëv. Preconditioned iterative methods for a class of nonlinear eigenvalue problems. Linear Algebra and its Applications, 415(1): 210–229, May 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Song:2000:IHP

[Son00] Yongzhong Song. On an inequality for the Hadamard product of an M-matrix and its inverse. Linear Algebra and its Applications, 305(1– 3):99–105, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/120/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/120/25/32/article.
pdf.

Song:2002:NVS

[Son02] Yongzhong Song. A note on the variation of the spectrum of an arbitrary matrix. Linear Algebra and its Applications, 342(1–3): 41–46, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/173/27/
29/abstract.html.

Soto:2003:ECN

[Sot03] Ricardo L. Soto. Existence and construction of nonnegative matrices with prescribed spectrum. Linear Algebra and its Applications, 369(C): 169–184, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/218/25/
                                        34/abstract.html.

Soto:2006:RCS

[Sot06] Ricardo L. Soto. Realizability criterion for the symmetric nonnegative inverse eigenvalue problem. Linear Algebra and its Applications, 416 (2–3):783–794, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Soules:2005:PBN

[Sou05] George W. Soules. Permanental bounds for nonnegative matrices via decomposition. Linear Algebra and its Applications, 394(1):73– 89, January 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Simoncini:2004:BSK

[SP04] V. Simoncini and M. Pennacchio. The behavior of symmetric Krylov subspace methods for solving Mx = (M − γI)v. Linear Algebra and its Applications, 380(1): 53–71, March 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shomron:2009:LAM

[SP09] Noam Shomron and Beresford N. Parlett. Linear algebra meets Lie algebra: The Kostant–Wallach theory. Linear Algebra and its Applications, 431(10):1745–1767, October 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Spiegel:2000:CSM

[Spi00] Eugene Spiegel. Completing symplectic matrices. Linear Algebra and its Applications, 312(1–3):93–100, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/129/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/129/25/29/article.
pdf.

Spiegel:2005:IIA

[Spi05] E. Spiegel. Involutions in incidence algebras. Linear Algebra and its Applications, 405(1):155–162, August 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Soto:2006:ABT

[SR06] Ricardo L. Soto and Oscar Rojo. Applications of a Brauer theorem in the nonnegative inverse eigenvalue problem. Linear Algebra and its Applications, 416(2– 3):844–856, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Singer:2000:REP

[SS00] Sanja Singer and Saša Singer. Rounding-error and perturbation bounds for the indefinite QR factorization. Linear Algebra and its Applications, 309(1–3):103–119, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/32/article.
pdf.

Shao:2001:MSG

[SS01] Jia-Yu Shao and Hai-Ying Shan. Matrices with signed generalized inverses. Linear Algebra and its Applications, 322(1–3):105–127, January 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/141/24/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/

Shao:2002:SPM

[SS02] Jia-Yu Shao and Hai-Ying Shan. The solution of a problem on matrices having signed generalized inverses. Linear Algebra and its Applications, 345(1–3): 43–70, April 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/175/27/
30/abstract.html.

Schwetlick:2003:ICS

[SS03a] Hubert Schwetlick and Uwe Schnabel. Iterative computation of the smallest singular value and the corresponding singular vectors of a matrix. Linear Algebra and its Applications, 371(1):1–30, September 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shader:2003:ENM

[SS03b] Bryan L. Shader and Saib Suwilo. Exponents of nonnegative matrix pairs. Linear Algebra and its Applications, 363(C):275–293, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
45/abstract.html.

Singer:2003:REP

[SS03c] Sanja Singer and Saša Singer. Rounding error and perturbation bounds for the symplectic QR factorization. Linear Algebra and its Applications, 358(1–3): 255–279, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/207/25/
                                        39/abstract.html.

Schmale:2004:CMP

[SS04a] Wiland Schmale and Pramod K. Sharma. Cyclizable matrix pairs over C[x] and a conjecture on Töplitz pencils. Linear Algebra and its Applications, 389(1):33–42, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shan:2004:MTS

[SS04b] Hai-Ying Shan and Jia-Yu Shao. Matrices with totally signed powers. Linear Algebra and its Applications, 376 (1):215–224, January 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shao:2005:DRC

[SS05a] Jia-Yu Shao and Hai-Ying Shan. The determinantal regions of complex sign pattern matrices and ray pattern matrices. Linear Algebra and its Applications, 395(1):211– 228, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stevanovic:2005:RHC

[SS05b] Dragan Stevanović and Ivan Stanković. Remarks on hyperenergetic circulant graphs. Linear Algebra and its Applications, 400(1):345–348, May 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Salinelli:2006:CMY

[SS06a] Ernesto Salinelli and Carlo Sgarra. Correlation matrices of yields and total positivity. Linear Algebra and its Applications, 418(2–3):682–692, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Salvi:2006:VSC

[SS06b] Rodolfo Salvi and Norma Zagaglia Salvi. On very sparse circulant (0, 1) matrices. Linear Algebra and its Applications, 418(2–3):565–575, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Saranya:2006:GII

[SS06c] C. R. Saranya and K. C. Sivakumar. Generalized inverses of an invertible infinite matrix over a finite field. Linear Algebra and its Applications, 418(2–3):468–479, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seidman:2006:SRP

[SS06d] Thomas I. Seidman and Hans Schneider. The spectral radius in partially ordered algebras. Linear Algebra and its Applications, 417(2–3):347– 369, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sun:2006:ESO

[SS06e] Huaqing Sun and Yuming Shi. Eigenvalues of secondorder difference equations with coupled boundary conditions. Linear Algebra and its Applications, 414(1):361– 372, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Saez-Schwedt:2007:FCR

[SS07a] A. Sáez-Schwedt. Feedback cyclization for rings with finite stable range. Linear Algebra and its Applications, 427(2–3):234–241, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Salinelli:2007:SSC

[SS07b] Ernesto Salinelli and Carlo Sgarra. Shift, slope and curvature for a class of yields correlation matrices. Linear Algebra and its Applications, 426(2–3):650–666, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shi:2007:UBS

[SS07c] Lingsheng Shi and Zhipeng Song. Upper bounds on the spectral radius of bookfree and/or K_{2,l}-free graphs. Linear Algebra and its Applications, 420(2–3):526–529, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Silva:2007:LSEa

[SS07d] Fernando C. Silva and Rita Simões. On the Lyapunov and Stein equations. Linear Algebra and its Applications, 420(2–3):329–338, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Silva:2007:LSEb

[SS07e] Fernando C. Silva and Rita Simões. On the Lyapunov and Stein equations, II. Linear Algebra and its Applications, 426(2–3):305–311, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stegeman:2007:KUC

[SS07f] Alwin Stegeman and Nicholas D. Sidiropoulos. On Kruskal’s uniqueness condition for the Candecomp/Parafac decomposition. Linear Algebra and its Applications, 420(2–3): 540–552, January 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Simic:2008:PRG

[SS08a] Slobodan K. Simić and Zoran Stanić. On the polynomial reconstruction of graphs whose vertex-deleted subgraphs have spectra bounded from below by −2. Linear Algebra and its Applications, 428(8–9):1865–1873, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Singer:2008:OBR

[SS08b] Sanja Singer and Saša Singer. Orthosymmetric block reflectors. Linear Algebra and its Applications, 429(5–6):1354– 1385, September 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sander:2009:TDE

[SS09a] T. Sander and J. W. Sander. Tree decomposition by eigenvectors. Linear Algebra and its Applications, 430(1):133– 144, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seidel:2009:BAS

[SS09b] Markus Seidel and Bernd Silbermann. Banach algebras of structured matrix sequences. Linear Algebra and its Applications, 430(4):1243–1281, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shang:2009:FPM

[SS09c] Ying Shang and Michael K. Sain. Fixed poles in the model matching problem for systems over semirings. Linear Algebra and its Applications, 430(8–9):2368–2388, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shivakumar:2009:RIM

[SS09d] P. N. Shivakumar and K. C. Sivakumar. A review of infinite matrices and their applications. Linear Algebra and its Applications, 430(4):976– 998, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seidman:2005:CTU

[SSA05] Thomas I. Seidman, Hans Schneider, and Marina Arav. Comparison theorems using general cones for norms of iteration matrices. Linear Algebra and its Applications, 399(1):169–186, April 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sergeev:2009:VSS

[SSB09] Sergeĭ Sergeev, Hans Schneider, and Peter Butkovič. On visualization scaling, subeigenvectors and Kleene stars in max algebra. Linear Algebra and its Applications, 431(12):2395–2406, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shao:2004:IST

[SSG04] Yanling Shao, Liang Sun, and Yubin Gao. Inertia sets of two classes of symmetric sign patterns. Linear Algebra and its Applications, 384 (1):85–95, June 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Schmid:2000:DMC

[SSMQ^{+}00] Michael Schmid, Rainer Steinwandt, Jörn Müller-Quade, Martin Rötteler, and Thomas Beth. Decomposing a matrix into circulant and diagonal factors. Linear Algebra and its Applications, 306 (1–3):131–143, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/121/25/35/article.
pdf.

Saez-Schwedt:2006:CGR

[SSSG06] A. Sáez-Schwedt and T. Sánchez-Giralda. A characterization of GCU rings with the FC_{2} property. Linear Algebra and its Applications, 413(2–3): 297–302, March 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Saez-Schwedt:2008:CAB

[SSSG08] A. Sáez-Schwedt and T. Sánchez-Giralda. Coefficient assignability and a block decomposition for systems over rings. Linear Algebra and its Applications, 429(5–6):1277–1287, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seberry:2005:CTG

[SSW05] Jennifer Seberry, Sarah A. Spence, and Tadeusz A. Wysocki. A construction technique for generalized complex orthogonal designs and applications to wireless communications. Linear Algebra and its Applications, 405(1):163–176, August 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Song:2008:QGV

[SSW08] Guang’ai Song, Yucai Su, and Yuezhu Wu. Quantization of generalized Virasoro-like algebras. Linear Algebra and its Applications, 428(11–12): 2888–2899, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shao:2005:RSL

[SSY05] Jia-Yu Shao, Hai-Ying Shan, and Li-Hua You. Ray solvable linear systems and ray S^{2}NS matrices. Linear Algebra and its Applications, 395(1):229–246, January 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Slapnicar:2000:RPT

[ST00] Ivan Slapničar and Ninoslav Truhar. Relative perturbation theory for hyperbolic eigenvalue problem. Linear Algebra and its Applications, 309(1–3):57–72, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/

gej-ng/10/30/19/126/25/
29/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/29/article.
pdf.

Seeger:2003:EIC

[ST03a] Alberto Seeger and Mounir Torki. On eigenvalues induced by a cone constraint. Linear Algebra and its Applications, 372(1):181–206, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Slapnicar:2003:RPT

[ST03b] Ivan Slapničar and Ninoslav Truhar. Relative perturbation theory for hyperbolic singular value problem. Linear Algebra and its Applications, 358(1–3): 367–386, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/207/25/
44/abstract.html.

Spielman:2007:SPW

[ST07] Daniel A. Spielman and Shang-Hua Teng. Spectral partitioning works: Planar graphs and finite element meshes. Linear Algebra and its Applications, 421(2–3): 284–305, March 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seo:2008:CAH

[ST08] Yuki Seo and Masaru Tominaga. A complement of the Ando–Hiai inequality. Linear Algebra and its Applications, 429(7):1546–1554, October 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stanic:2007:GWS

[Sta07] Zoran Stanić. On graphs whose second largest eigenvalue equals 1 — the star complement technique. Linear Algebra and its Applications, 420(2–3):700–710, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stanic:2009:DCF

[Sta09a] Zoran Stanić. On determination of caterpillars with four terminal vertices by their Laplacian spectrum. Linear Algebra and its Applications, 431(11):2035–2048, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stanic:2009:NSG

[Sta09b] Zoran Stanić. On nested split graphs whose second largest eigenvalue is less than 1. Linear Algebra and its Applications, 430(8–9):2200– 2211, April 15, 2009. CO-

Sustik:2007:EET

[STDH07] Mátyás A. Sustik, Joel A. Tropp, Inderjit S. Dhillon, and Robert W. Heath Jr. On the existence of equiangular tight frames. Linear Algebra and its Applications, 426(2–3):619–635, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stepniak:2000:MRC

[Stȩ00a] Czesław Stȩpniak. On a matrix result in comparison of linear experiments. Linear Algebra and its Applications, 321(1–3):321–325, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
45/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/45/article.
pdf.

Stevanovic:2000:WCC

[Ste00b] Dragan Stevanović. When can the components of NEPS of connected bipartite graphs be almost cospectral? Linear Algebra and its Applications, 311(1–3):35–44, May 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/128/25/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/128/25/28/article.
                                        pdf.

Stefanovski:2001:CFL

[Ste01a]  Jovan Stefanovski.  Con-

densed forms of linear control system under output feedback. Linear Algebra and its Applications, 328(1–3): 1–55, May 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej%2Dng/10/
                                        30/19/153/25/25/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej%2Dng/10/30/19/153/
                                        25/25/article.pdf.

Stewart:2001:GST

[Ste01b] G. W. Stewart. A generalization of Saad’s theorem on Rayleigh–Ritz approximations. Linear Algebra and its Applications, 327(1–3):115– 119, April 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic). URL http://www.
                                        elsevier.nl/gej%2Dng/10/
                                        30/19/152/25/35/abstract.
                                        html;   http://www.elsevier.
                                        nl/gej%2Dng/10/30/19/152/
                                        25/35/article.pdf.

Stevanovic:2002:CNC

[Ste02a]  Dragan Stevanović.  On

the components of NEPS of connected bipartite graphs.

Linear Algebra and its Applications, 356(1–3):67–78, November 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/202/25/
31/abstract.html.

Stewart:2002:BEB

[Ste02b] G. W. Stewart. Backward error bounds for approximate Krylov subspaces. Linear Algebra and its Applications, 340(1–3):81–86, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
32/abstract.html.

Stevanovic:2003:BLE

[Ste03a] Dragan Stevanović. Bounding the largest eigenvalue of trees in terms of the largest vertex degree. Linear Algebra and its Applications, 360 (C):35–42, February 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/209/25/
28/abstract.html.

Stevanovic:2003:RGM

[Ste03b] Dragan Stevanović. Remarks on graphs with majority of eigenvalues at most 1. Linear Algebra and its Applications, 367(C):337–340, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/216/25/
                                        49/abstract.html.

Stewart:2004:ELP

[Ste04] G. W. Stewart. An Elsnerlike perturbation theorem for generalized eigenvalues. Linear Algebra and its Applications, 390(1):1–5, October 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Steiner:2005:VTF

[Ste05a] Marcel Steiner. Variations on a theme of Frobenius about almost commuting unitary matrices. Linear Algebra and its Applications, 400 (1):1–13, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stepniak:2005:MER

[Ste05b] Czesław Stepniak. Matrix equations with restraints and their statistical applications. Linear Algebra and its Applications, 410(1):279–287, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stefanovski:2006:SFN

[Ste06a] Jovan Stefanovski. Spectral factorization of nonsymmetric polynomial matrices. Linear Algebra and its Applications, 412(2–3):412– 440, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stewart:2006:NGH

[Ste06b] G. W. Stewart. A note on generalized and hypergeneralized projectors. Linear Algebra and its Applications, 412(2–3):408–411, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stewart:2006:PSP

[Ste06c] Michael Stewart. Perturbation of the SVD in the presence of small singular values. Linear Algebra and its Applications, 419(1):53– 77, November 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stender:2007:GIP

[Ste07a] Torben Stender. A generalization of imaginary parts of eigenvalues for matrices: Chain rotation numbers. Linear Algebra and its Applications, 426(1):53– 70, October 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stevanovic:2007:RPA

[Ste07b] Dragan Stevanović. Research problems from the Aveiro Workshop on Graph Spectra. Linear Algebra and its Applications, 423(1):172– 181, May 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stewart:2007:GIA

[Ste07c] Michael Stewart. A generalized isometric Arnoldi algorithm. Linear Algebra and its Applications, 423(2–3):183– 208, June 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stenger:2008:FSZ

[Ste08] Frank Stenger. Fourier series for zeta function via Sinc. Linear Algebra and its Applications, 429(10):2636– 2639, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stegeman:2009:UCC

[Ste09] Alwin Stegeman. On uniqueness conditions for Candecomp/Parafac and Indscal with full column rank in one mode. Linear Algebra and its Applications, 431(1– 2):211–227, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Strohmer:2001:FSS

[Str01]  Thomas Strohmer.  Four

short stories about Toeplitz matrix calculations. Linear Algebra and its Applications, 343–344(1):321–344, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
41/abstract.html.

Strikwerda:2002:PAA

[Str02] John C. Strikwerda. A probabilistic analysis of asynchronous iteration. Linear Algebra and its Applications, 349(1–3):125–154, July 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Strasek:2003:UPJ

[Str03a] Rok Strašek. Uniform primeness of the Jordan algebra of hermitian quaternion matrices. Linear Algebra and its Applications, 367 (C):235–242, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
42/abstract.html.

Strauss:2003:BOT

[Str03b] Vladimir Strauss. On the bicommutant for one type of Jsymmetric nilpotent algebras in Krein spaces. Linear Algebra and its Applications, 372 (1):167–180, October 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Strasek:2006:OQH

[Str06] Rok Strašek. On Euclidean algebra of Hermitian operators on a quaternionic Hilbert space. Linear Algebra and its Applications, 416(2–3):580– 587, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Strohmer:2008:NET

[Str08] Thomas Strohmer. A note on equiangular tight frames. Linear Algebra and its Applications, 429(1):326–330, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sturm:2000:SOS

[Stu00] Jos F. Sturm. Similarity and other spectral relations for symmetric cones. Linear Algebra and its Applications, 312(1–3):135–154, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/129/25/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/129/25/33/article.
pdf.

Stuart:2003:RPP

[Stu03] Jeffrey Stuart. Reducible pattern k-potent ray pattern matrices. Linear Algebra and its Applications, 362 (C):87–99, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/211/25/
31/abstract.html.

Stykel:2002:NSP

[Sty02a] Tatjana Stykel. Numerical solution and perturbation theory for generalized Lyapunov equations. Linear Algebra and its Applications, 349(1–3):155–185, July 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stykel:2002:SIT

[Sty02b] Tatjana Stykel. Stability and inertia theorems for generalized Lyapunov equations. Linear Algebra and its Applications, 355(1–3): 297–314, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/201/27/
45/abstract.html.

Stykel:2006:BTM

[Sty06] Tatjana Stykel. Balanced truncation model reduction for semidiscretized Stokes equation. Linear Algebra and its Applications, 415(2–3): 262–289, June 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shaska:2009:HAG

[SU09] T. Shaska and V. Ustimenko. On the homogeneous algebraic graphs of large girth and their applications. Linear Algebra and its Applications, 430(7): 1826–1837, April 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Suciu:2009:QIS

[Suc09]  Laurian Suciu.  Quasi-

isometries in semi-Hilbertian spaces. Linear Algebra and its Applications, 430(8–9): 2474–2487, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sudo:2002:SGA

[Sud02] Takahiro Sudo. Structure of group C^{∗}-algebras of the generalized disconnected Mautner groups. Linear Algebra and its Applications, 341 (1–3):317–326, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
52/abstract.html.

Sudo:2004:SRA

[Sud04] Takahiro Sudo. Stable rank of C^{∗}-algebras of type I. Linear Algebra and its Applications, 383(1):65–76, May 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sun:2002:GSU

[Sun02] Yeong-Jeu Sun. Global stabilizability of uncertain systems with time-varying delays via dynamic observerbased output feedback. Linear Algebra and its Applications, 353(1–3):91–105, September 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sun:2005:SSP

[Sun05] Ji-Guang Sun. On the sensitivity of the spectral projection. Linear Algebra and its Applications, 395(1):83– 94, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sun:2008:SDR

[Sun08] Yidong Sun. The Star of David rule. Linear Algebra and its Applications, 429(8–9):1954–1961, October 16, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Singh:2001:MMF

[SV01] Mandeep Singh and H. L. Vasudeva. Monotone matrix functions of two variables. Linear Algebra and its Applications, 328(1–3): 131–152, May 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.
elsevier.nl/gej%2Dng/10/
30/19/153/25/30/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/153/
25/30/article.pdf.

Steerneman:2005:PM

[Sv05] Ton Steerneman and Frederieke van Perlo-ten Kleij. Properties of the matrix A − XY ∗. Linear Algebra and its Applications, 410(1):70– 86, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

smithies:2006:SVD

[sV06a] Laura smithies and Richard S. Varga. Singular value decomposition Geršgorin sets. Linear Algebra and its Applications, 417(2–3):370–380, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Soto:2006:PCM

[SV06b] M. J. Soto and J. L. Vicente. Polyhedral cones and monomial blowing-ups.

Linear Algebra and its Applications, 412(2–3):362–372, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Santesso:2007:ZPP

[SV07] Paolo Santesso and Maria Elena Valcher. On the zero pattern properties and asymptotic behavior of continuoustime positive system trajectories. Linear Algebra and its Applications, 425(2–3):283– 302, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Satnoianu:2005:SRM

[SvdD05] R. A. Satnoianu and P. van den Driessche. Some remarks on matrix stability with application to Turing instability. Linear Algebra and its Applications, 398(1):69– 74, March 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sleijpen:2003:UHR

[SvdE03] Gerard L. G. Sleijpen and Jasper van den Eshof. On the use of harmonic Ritz pairs in approximating internal eigenpairs. Linear Algebra and its Applications, 358(1–3): 115–137, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/207/25/
                                        32/abstract.html.

Svalbe:2001:RTI

[SvdS01] Imants Svalbe and Dimitri van der Spek. Reconstruction of tomographic images using analog projections and the digital Radon transform. Linear Algebra and its Applications, 339(1–3): 125–145, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/169/27/
                                        35/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/169/27/35/article.
                                        pdf.

Sonneveld:2009:NMF

[SvKHO09] P. Sonneveld, J. J. I. M. van Kan, X. Huang, and C. W. Oosterlee. Nonnegative matrix factorization of a correlation matrix. Linear Algebra and its Applications, 431 (3–4):334–349, July 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Srivastava:2002:RMU

[SvR02] Muni S. Srivastava and Dietrich von Rosen. Regression models with unknown singular covariance matrix. Linear Algebra and its Applications, 354(1–3): 255–273, October 15, 2002.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
47/abstract.html.

Smith:2005:CMF

[SvW05] Kirby C. Smith and Leon van Wyk. A concrete matrix field description of some Galois fields. Linear Algebra and its Applications, 403(1): 159–164, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shen:2003:NAP

[SW03] Jian Shen and Cynthia J. Wyels. On the number of arcs in primitive digraphs with large exponents. Linear Algebra and its Applications, 364 (C):243–251, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
33/abstract.html.

Shu:2004:SUB

[SW04] Jinlong Shu and Yarong Wu. Sharp upper bounds on the spectral radius of graphs. Linear Algebra and its Applications, 377(1):241–248, January 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shi:2009:STS

[SW09] Guoliang Shi and Hongyou Wu. Spectral theory of Sturm–Liouville difference operators. Linear Algebra and its Applications, 430 (2–3):830–846, January 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Seberry:2003:MDS

[SXKM03] Jennifer Seberry, Tianbing Xia, Christos Koukouvinos, and Marilena Mitrouli. The maximal determinant and subdeterminants of ±1 matrices. Linear Algebra and its Applications, 373(1):297– 310, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Stacho:2004:BPS

[SZ04] L. L. Stachó and B. Zalar. Bicircular projections on some matrix and operator spaces. Linear Algebra and its Applications, 384(1):9– 20, June 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Spivey:2008:SPP

[SZ08] Michael Z. Spivey and Andrew M. Zimmer. Symmetric polynomials, Pascal matrices, and Stirling matrices. Linear Algebra and its Applications, 428(4):1127–1134, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Szabo:2007:CMM

[Sza07] V. E. S. Szabó. A class of matrix monotone functions. Linear Algebra and its Applications, 420(1):79– 85, January 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Sheng:2007:MSS

[SZZ07] Yu Qiu Sheng, Bao Dong Zheng, and Xian Zhang. Maps on spaces of symmetric matrices preserving idempotence. Linear Algebra and its Applications, 420(2–3):576– 585, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tabachnikova:2001:ISL

[Tab01] O. M. Tabachnikova. Infinite systems of linear equations and mathematical logic. Linear Algebra and its Applications, 343–344(1):473– 475, March 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
48/abstract.html.

Tadej:2006:PEC

[Tad06] Wojciech Tadej. Permutation equivalence classes of Kronecker products of unitary Fourier matrices. Linear Algebra and its Applications, 418(2–3):719–736, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tagami:2004:SAS

[Tag04] Makoto Tagami. Symmetric association schemes attached to finite upper half planes over rings. Linear Algebra and its Applications, 376(1): 225–234, January 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Takahashi:2000:STS

[Tak00] K. Takahashi. On sums of three square-zero matrices. Linear Algebra and its Applications, 306(1–3): 45–57, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
30/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/121/25/30/article.
pdf.

Takane:2004:MSR

[Tak04] Yoshio Takane. Matrices with special reference to applications in psychometrics.

Linear Algebra and its Applications, 388(1):341–361, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tam:2004:PGE

[Tam04] Bit-Shun Tam. The Perron generalized eigenspace and the spectral cone of a cone-preserving map. Linear Algebra and its Applications, 393(1):375–429, December 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tan:2001:PMD

[Tan01] Yijia Tan. On the powers of matrices over a distributive lattice. Linear Algebra and its Applications, 336(1– 3):1–14, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/166/27/27/article.
pdf.

Tan:2003:EMD

[Tan03a] Yijia Tan. On the eigenproblem of matrices over distributive lattices. Linear Algebra and its Applications, 374(1): 87–106, November 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tang:2003:LOP

[Tan03b] Xiao Min Tang. Linear operators preserving adjoint matrix between matrix spaces. Linear Algebra and its Applications, 372(1):287–293, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tan:2005:TMD

[Tan05a] Yi-Jia Tan. On the transitive matrices over distributive lattices. Linear Algebra and its Applications, 400(1): 169–191, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tang:2005:ARP

[Tan05b] Xiao-Min Tang. Additive rank-1 preservers between hermitian matrix spaces and applications. Linear Algebra and its Applications, 395(1): 333–342, January 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tan:2007:IMA

[Tan07] Yijia Tan. On invertible matrices over antirings. Linear Algebra and its Applications, 423(2–3):428–444, June 1, 2007. CODEN LAAPAW. ISSN 0024-

Tan:2008:NMA

[Tan08] Yijia Tan. On nilpotent matrices over antirings. Linear Algebra and its Applications, 429(5–6):1243–1253, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tanaka:2009:BFR

[Tan09a] Hajime Tanaka. A bilinear form relating two Leonard systems. Linear Algebra and its Applications, 431(10): 1726–1739, October 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tanaka:2009:NSH

[Tan09b] Hajime Tanaka. A note on the span of Hadamard products of vectors. Linear Algebra and its Applications, 430(2–3):865–867, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tao:2006:MRS

[Tao06] Yunxing Tao. More results on singular value inequalities of matrices. Linear Algebra and its Applications, 416 (2–3):724–729, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tarazaga:2005:FCE

[Tar05] Pablo Tarazaga. Faces of the cone of Euclidean distance matrices: Characterizations, structure and induced geometry. Linear Algebra and its Applications, 408 (1):1–13, October 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tendeiro:2009:STT

[TBK09] Jorge N. Tendeiro, Jos M. F. Ten Berge, and Henk A. L. Kiers. Simplicity transformations for three-way arrays with symmetric slices, and applications to Tucker-3 models with sparse core arrays. Linear Algebra and its Applications, 430(4):924– 940, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

tenBerge:2006:STS

[tBS06] Jos M. F. ten Berge and Alwin Stegeman. Symmetry transformations for square sliced three-way arrays, with applications to their typical rank. Linear Algebra and its Applications, 418(1):215– 224, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

tenBerge:2004:TRI

[tBSR04] Jos M. F. ten Berge, Nikolaos D. Sidiropoulos, and Roberto Rocci. Typical rank and indscal dimensionality for symmetric three-way arrays of order I × 2 × 2 or I × 3 × 3. Linear Algebra and its Applications, 388(1):363– 377, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trefethen:2005:SP

[TE05] Lloyd N. Trefethen and Mark Embree. Spectra and pseudospectra: the behavior of nonnormal matrices and operators. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-11946-5. xviii + 606 pp. The behavior of nonnormal matrices and operators.

Terhal:2001:FIP

[Ter01a] Barbara M. Terhal. A family of indecomposable positive linear maps based on entangled quantum states. Linear Algebra and its Applications, 323(1–3):61–73, January 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Terwilliger:2001:TLT

[Ter01b] Paul Terwilliger. Two linear transformations each tridiagonal with respect to an eigenbasis of the other. Linear Algebra and its Applications, 330(1–3):149–203, June 15, 2001. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/155/25/
                                        35/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/155/25/35/article.
                                        pdf.

Terwilliger:2002:SAD

[Ter02] Paul Terwilliger. The subconstituent algebra of a distanceregular graph; thin modules with endpoint one. Linear Algebra and its Applications, 356(1–3):157–187, November 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/202/25/
                                        39/abstract.html.

Teranishi:2003:ESS

[Ter03] Yasuo Teranishi. Equitable switching and spectra of graphs. Linear Algebra and its Applications, 359(1–3): 121–131, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/208/25/
                                        36/abstract.html.

Terwilliger:2004:LPR

[Ter04] Paul Terwilliger. Leonard pairs and the q-Racah polynomials. Linear Algebra and its Applications, 387(1):235– 276, August 1, 2004. CO-DEN LAAPAW. ISSN 0024-

Tam:2008:ULM

[TFZ08] Bit-Shun Tam, Yi-Zheng Fan, and Jun Zhou. Unoriented Laplacian maximizing graphs are degree maximal. Linear Algebra and its Applications, 429(4):735– 758, August 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Takahashi:2002:CFP

[TG02] Satoru Takahashi and Bijoy K. Ghosh. Canonical forms and parameter identification problems in perspective systems. Linear Algebra and its Applications, 351–352:701–717, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Thornburg:2000:OIM

[TH00] Mark L. Thornburg and Richard D. Hill. Order intervals of matrices. Linear Algebra and its Applications, 305(1–3):33–46, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/120/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/120/25/28/article.
pdf.

Tam:2001:IN

[TH01] Tin-Yau Tam and William C. Hill. On G-invariant norms. Linear Algebra and its Applications, 331(1–3):101– 112, July 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/156/25/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/156/25/33/article.
                                        pdf.

Telek:2004:AIM

[THH04] M. Telek, A. Horváth, and G. Horváth. Analysis of inhomogeneous Markov reward models. Linear Algebra and its Applications, 386(1):383– 405, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Thielemann:2004:BSR

[Thi04] Henning Thielemann. Bounds for smoothness of refinable functions. Linear Algebra and its Applications, 391(1):245– 260, November 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tian:2009:NSP

[THZ09] Gui-Xian Tian, Ting-Zhu Huang, and Bo Zhou. A note on sum of powers of the Laplacian eigenvalues of bipartite graphs. Linear Algebra and its Applications, 430(8–9):2503–2510, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Takamatsu:2008:IRD

[TI08] Mizuyo Takamatsu and Satoru Iwata. Index reduction for differential–algebraic equations by substitution method. Linear Algebra and its Applications, 429(8–9): 2268–2277, October 16, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tian:2000:CTB

[Tia00]  Yongge Tian.  Complet-

ing triangular block matrices with maximal and minimal ranks. Linear Algebra and its Applications, 321(1–3): 327–345, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
46/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/46/article.
pdf.

Tian:2002:ULB

[Tia02] Yongge Tian. Upper and lower bounds for ranks of matrix expressions using generalized inverses. Linear Algebra and its Applications, 355 (1–3):187–214, November 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/201/27/
                                        39/abstract.html.

Ticozzi:2005:OCA

[Tic05] Francesco Ticozzi. Optimal commuting approximation of Hermitian operators. Linear Algebra and its Applications, 400(1):319–325, May 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tikhonov:2006:NDM

[Tik06] Oleg E. Tikhonov. A note on definition of matrix convex functions. Linear Algebra and its Applications, 416 (2–3):773–775, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tilli:2003:UBC

[Til03] Paolo Tilli. Universal bounds on the convergence rate of extreme Toeplitz eigenvalues. Linear Algebra and its Applications, 366(C):403– 416, June 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/215/25/
                                        47/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Tisseur:2000:BEC

[Tis00] Françoise Tisseur. Backward error and condition of polynomial eigenvalue problems. Linear Algebra and its Applications, 309(1–3): 339–361, April 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/126/25/
42/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/126/25/42/article.
pdf.

Tifenbach:2009:DID

[TK09] R. M. Tifenbach and S. J. Kirkland. Directed intervals and the dual of a graph. Linear Algebra and its Applications, 431(5–7):792– 807, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tkachev:2005:UFM

[Tka05] V. G. Tkachev. Ullemar’s formula for the moment map, II. Linear Algebra and its Applications, 404(1):380– 388, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Thorndyke:2008:RRM

[TKX08] B. Thorndyke, A. Koong, and L. Xing. Reducing respiratory motion artifacts in radionuclide imaging through retrospective stacking: A simulation study. Linear Algebra and its Applications, 428(5–6):1325–1344, March 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trendafilov:2002:MPP

[TL02] Nickolay T. Trendafilov and Ross A. Lippert. The multimode Procrustes problem. Linear Algebra and its Applications, 349(1–3):245– 264, July 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Truhar:2003:TGI

[TL03] Ninoslav Truhar and Ren-Cang Li. A sin2Θ theorem for graded indefinite Hermitian matrices. Linear Algebra and its Applications, 359 (1–3):263–276, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/208/25/
                                        45/abstract.html.

Tanaka:2000:SFS

[TM00] R. Tanaka and K. Murota. Symmetric failures in symmetric control systems. Linear Algebra and its Applications, 318(1–3):145–172, October 15, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/137/24/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/137/24/34/article.
pdf.

Torre-Mayo:2007:NIE

[TMARAE^{+}07] J. Torre-Mayo, M. R. Abril-Raymundo, E. Alarcia-Estévez, C. Marijuán, and M. Pisonero. The nonnegative inverse eigenvalue problem from the coefficients of the characteristic polynomial. EBL digraphs. Linear Algebra and its Applications, 426(2–3):729–773, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tonge:2000:ECM

[Ton00]  Andrew Tonge.  Equiva-

lence constants for matrix norms: a problem of Goldberg. Linear Algebra and its Applications, 306(1–3): 1–13, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
25/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/121/25/25/article.
                                        pdf.

Toni:2004:LBZ

[Ton04] A. Toni. Lower bounds on zero-one matrices. Linear Algebra and its Applications, 376(1):275–282, January 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tovbis:2000:NFN

[Tov00] Alexander Tovbis. Normal forms of “near similarity” transformations and linear matrix equations. Linear Algebra and its Applications, 317(1–3):13–40, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/136/25/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/136/25/26/article.
                                        pdf.

Toyonaga:2009:NEM

[Toy09] Kenji Toyonaga. Numerical enclosure for multiple eigenvalues of an Hermitian matrix whose graph is a tree. Linear Algebra and its Applications, 431(11):1989–1999, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trenkler:2005:MVS

[TP05] Götz Trenkler and Simo Puntanen. A multivariate version of Samuelson’s inequality. Linear Algebra and its Applications, 410(1):143–149, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tarfulea:2009:IFP

[TP09a] Andrei Tarfulea and Robert Perlis. An Ihara formula for partially directed graphs. Linear Algebra and its Applications, 431(1–2):73–85, July 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tian:2009:EEU

[TP09b] Yongge Tian and Simo Puntanen. On the equivalence of estimations under a general linear model and its transformed models. Linear Algebra and its Applications, 430 (10):2622–2641, May 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Taher:2001:LRR

[TR01] Rajae Ben Taher and Mustapha Rachidi. Linear recurrence relations in the algebra of matrices and applications. Linear Algebra and its Applications, 330(1–3):15–24, June 15, 2001. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/155/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/155/25/26/article.
pdf.

Taher:2002:SEF

[TR02] R. Ben Taher and M. Rachidi. Some explicit formulas for the polynomial decomposition of the matrix exponential and applications. Linear Algebra and its Applications, 350(1–3):171–184, July 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Taher:2003:MPE

[TR03] Rajae Ben Taher and Mustapha Rachidi. On the matrix powers and exponential by the r-generalized Fibonacci sequences methods: the companion matrix case. Linear Algebra and its Applications, 370(1):341–353, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2002:SDG

[Tre02] William F. Trench. Spectral distribution of generalized Kac–Murdock–Szegő matrices. Linear Algebra and its Applications, 347(1– 3):251–273, May 15, 2002.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
42/abstract.html.

Trench:2003:AEDb

[Tre03a] William F. Trench. Absolute equal distribution of families of finite sets. Linear Algebra and its Applications, 367 (C):131–146, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
34/abstract.html.

Trench:2003:AEDa

[Tre03b] William F. Trench. Absolute equal distribution of the spectra of Hermitian matrices. Linear Algebra and its Applications, 366 (C):417–431, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
48/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Trench:2003:NAZ

[Tre03c] William F. Trench. A note on asymptotic zero distribution of orthogonal polynomials. Linear Algebra and its Applications, 375(1):275– 281, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2004:CPM

[Tre04a] William F. Trench. Characterization and properties of matrices with generalized symmetry or skew symmetry. Linear Algebra and its Applications, 377(1):207– 218, January 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2004:ENA

[Tre04b] William F. Trench. An elementary note on asymptotic properties of Toeplitz and multilevel Toeplitz matrices. Linear Algebra and its Applications, 382(1):231– 235, May 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2004:HHS

[Tre04c] William F. Trench. Hermitian, Hermitian R-symmetric, and Hermitian R-skew symmetric Procrustes problems. Linear Algebra and its Applications, 387(1):83–98, August 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2004:IEA

[Tre04d]  William F. Trench.  In-

verse eigenproblems and associated approximation problems for matrices with generalized symmetry or skew symmetry. Linear Algebra and its Applications, 380(1): 199–211, March 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2004:MPS

[Tre04e] William F. Trench. Minimization problems for (R, S)- symmetric and (R, S)-skew symmetric matrices. Linear Algebra and its Applications, 389(1):23–31, September 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trent:2004:CTB

[Tre04f] Tavan T. Trent. An H^{2}- corona theorem on the bidisk for infinitely many functions. Linear Algebra and its Applications, 379(1):213– 227, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2005:MMI

[Tre05] William F. Trench. Multilevel matrices with involutory symmetries and skew symmetries. Linear Algebra and its Applications, 403(1):

53–74, July 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2006:ASP

[Tre06] William F. Trench. Asymptotic spectral properties of totally symmetric multilevel Toeplitz matrices. Linear Algebra and its Applications, 416(2–3):696–709, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2008:CPM

[Tre08] William F. Trench. Characterization and properties of matrices with k-involutory symmetries. Linear Algebra and its Applications, 429(8–9):2278–2290, October 16, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2009:NLS

[Tre09a] William F. Trench. On nonautonomous linear systems of differential and difference equations with Rsymmetric coefficient matrices. Linear Algebra and its Applications, 431(11):2109– 2117, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2009:PMB

[Tre09b] William F. Trench. Properties of multilevel block \underline{α}- circulants. Linear Algebra and its Applications, 431(10): 1833–1847, October 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trench:2009:PUB

[Tre09c] William F. Trench. Properties of unilevel block circulants. Linear Algebra and its Applications, 430(8–9):2012– 2025, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Trenn:2009:NFP

[Tre09d]  Stephan Trenn.  A nor-

mal form for pure differential algebraic systems. Linear Algebra and its Applications, 430(4):1070–1084, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tarazaga:2001:PFT

[TRH01] Pablo Tarazaga, Marcos Raydan, and Ana Hurman. Perron–Frobenius theorem for matrices with some negative entries. Linear Algebra and its Applications, 328(1– 3):57–68, May 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic). URL http://www.

elsevier.nl/gej%2Dng/10/
30/19/153/25/26/abstract.
html;   http://www.elsevier.
nl/gej%2Dng/10/30/19/153/
25/26/article.pdf.

Trishin:2002:REF

[Tri02] Ivan M. Trishin. On reduction of elements of the full matrix superalgebra to a block-diagonal form by conjugation. Linear Algebra and its Applications, 357(1–3): 59–82, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
30/abstract.html.

Trumpf:2007:OLT

[Tru07] Jochen Trumpf. Observers for linear time-varying systems. Linear Algebra and its Applications, 425(2–3):303– 312, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tian:2001:REI

[TS01] Yongge Tian and George P. H. Styan. Rank equalities for idempotent and involutary matrices. Linear Algebra and its Applications, 335 (1–3):101–117, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/165/27/

32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/165/27/32/article.
pdf.

Tam:2003:LEC

[TS03] Bit-Shun Tam and Hans Schneider. Linear equations over cones and Collatz– Wielandt numbers. Linear Algebra and its Applications, 363(C):295–332, April 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/212/25/
46/abstract.html.

Truhar:2006:RRB

[TS06] Ninoslav Truhar and Ivan Slapničar. Relative residual bounds for indefinite Hermitian matrices. Linear Algebra and its Applications, 417(2–3):466–477, September 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tian:2009:SME

[TS09] Yongge Tian and George P. H. Styan. On some matrix equalities for generalized inverses with applications. Linear Algebra and its Applications, 430(10):2716– 2733, May 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tsatsomeros:2000:PPT

[Tsa00] Michael J. Tsatsomeros. Principal pivot transforms: properties and applications. Linear Algebra and its Applications, 307(1–3):151–165, March 1, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/36/article.
pdf.

Tsatsomeros:2001:CEC

[Tsa01] Michael Tsatsomeros. A criterion for the existence of common invariant subspaces of matrices. Linear Algebra and its Applications, 322(1– 3):51–59, January 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/141/24/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/141/24/27/article.
pdf.

Tang:2007:SM

[TSOA07] A. Kevin Tang, Alp Simsek, Asuman Ozdaglar, and Daron Acemoglu. On the stability of P -matrices. Linear Algebra and its Applications, 426(1):22–32, October 1, 2007. CODEN

Tian:2009:MGI

[TT09] Yongge Tian and Yoshio Takane. More on generalized inverses of partitioned matrices with Banachiewicz–Schur forms. Linear Algebra and its Applications, 430(5–6):1641– 1655, March 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Takagi:2002:PII

[TTM02] Hiroyuki Takagi, Sin-Ei Takahasi, and Takeshi Miura. Polynomial identities that imply commutativity for rings. Linear Algebra and its Applications, 341(1–3): 299–307, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
50/abstract.html.

Tanahashi:2002:IPS

[TU02] Kôtarô Tanahashi and Atsushi Uchiyama. Isolated point of spectrum of pquasihyponormal operators. Linear Algebra and its Applications, 341(1–3):345–350, January 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/

                                        gej-ng/10/30/19/171/27/
                                        54/abstract.html.

Tanahashi:2004:IPS

[TUC04] Kôtarô Tanahashi, Atsushi Uchiyama, and Muneo Chō. Isolated points of spectrum of (p, k)-quasihyponormal operators. Linear Algebra and its Applications, 382(1):221– 229, May 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tuma:2006:RMP

[Tům06] Miroslav Tůma. Review of Matrix Preconditioning Techniques and Applications by Ke Chen. Cambridge University Press, 2005, xxiv + 568 pages, ISBN-13 978-0-521-83828-3, ISBN-10 0-521-83828-2. Linear Algebra and its Applications, 418(2–3): 955–957, October 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tupan:2008:SFF

[Tup08a]  Alexandru Tupan.  On

sesquilinear forms over fields with involution. Linear Algebra and its Applications, 428 (1):254–258, January 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tupan:2008:SDO

[Tup08b] Alexandru Tupan. On the similarity of dual operators. Linear Algebra and its Applications, 429(1):290– 292, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tupan:2009:SSD

[Tup09] Alexandru Tupan. Similarity of some dual operators in characteristic two. Linear Algebra and its Applications, 431(5–7):888–894, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Turnsek:2000:EOO

[Tur00] Aleksej Turnšek. Elementary operators and orthogonality. Linear Algebra and its Applications, 317(1–3): 207–216, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/136/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/136/25/35/article.
pdf.

Turnsek:2005:OPJ

[Tur05] Aleksej Turnšek. On operators preserving James’ orthogonality. Linear Algebra and its Applications, 407(1): 189–195, September 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tervo:2008:OCM

[TVB08] J. Tervo, M. Vauhkonen, and E. Boman. Optimal control model for radiation therapy inverse planning applying the Boltzmann transport equation. Linear Algebra and its Applications, 428(5–6):1230– 1249, March 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Torney:2000:SLC

[TW00] David C. Torney and Jun Wang. Systems of linear congruences with individual moduli. Linear Algebra and its Applications, 318(1–3): 195–208, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/137/24/
                                        37/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/137/24/37/article.
                                        pdf.

Tan:2003:DPP

[TW03] Victor Tan and Fei Wang. On determinant preserver problems. Linear Algebra and its Applications, 369(C): 311–317, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/218/25/
                                        44/abstract.html.

Tuncel:2005:SEU

[TW05] Levent Tunçel and Henry Wolkowicz. Strengthened existence and uniqueness conditions for search directions in semidefinite programming. Linear Algebra and its Applications, 400(1): 31–60, May 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tang:2009:DLA

[TX09] Xiao-Min Tang and Jin-Li Xu. The derivation Lie algebra of the higher rank Virasoro-like algebra and its automorphism groups. Linear Algebra and its Applications, 430(8–9):2170–2181, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Takane:2005:WGT

[TY05] Yoshio Takane and Haruo Yanai. On the Wedderburn– Guttman theorem. Linear Algebra and its Applications, 410(1):267–278, November 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tang:2006:SLP

[TY06] Xiao-Min Tang and Ya-Qin Yang. Strong linear preservers of rank reverse permutability on triangular matrices. Linear Algebra and its Applications, 414(1):84– 96, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Takane:2007:ACE

[TY07] Yoshio Takane and Haruo Yanai. Alternative characterizations of the extended Wedderburn–Guttman theorem. Linear Algebra and its Applications, 422(2–3):701– 711, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Takane:2008:RO

[TY08] Yoshio Takane and Haruo Yanai. On ridge operators. Linear Algebra and its Applications, 428(8–9):1778– 1790, April 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Takane:2009:NSC

[TY09] Yoshio Takane and Haruo Yanai. On the necessary and sufficient condition for the extended Wedderburn– Guttman theorem. Linear Algebra and its Applications, 430(11–12):2890– 2895, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tyrtyshnikov:2004:KPA

[Tyr04] Eugene Tyrtyshnikov. Kroneckerproduct approximations for some function-related matrices. Linear Algebra and its Applications, 379(1):423– 437, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tyrtyshnikov:2001:TER

[TZ01] E. E. Tyrtyshnikov and N. L. Zamarashkin. Toeplitz eigenvalues for Radon measures. Linear Algebra and its Applications, 343–344(1):345– 354, March 1, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
42/abstract.html.

Tyrtyshnikov:2003:GET

[TZ03] E. E. Tyrtyshnikov and N. L. Zamarashkin. A general equidistribution theorem for the roots of orthogonal polynomials. Linear Algebra and its Applications, 366 (C):433–439, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
49/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

Tadej:2008:DUM

[TŻ08] Wojciech Tadej and Karol Życzkowski. Defect of a unitary matrix. Linear Algebra and its Applications, 429 (2–3):447–481, July 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Uchiyama:2002:MMO

[Uch02] Mitsuru Uchiyama. Mixed matrix (operator) inequalities. Linear Algebra and its Applications, 341(1–3): 249–257, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/171/27/
                                        46/abstract.html.

Uddin:2009:ODW

[Udd09] Nizam Uddin. E-optimal designs when errors follow a conditional autonormal process. Linear Algebra and its Applications, 430(10):2642– 2649, May 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Uhlig:2000:CMT

[Uhl00] Frank Uhlig. Challenges in Matrix Theory 2000. Linear Algebra and its Applications, 304(1–3):179, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/118/25/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/118/25/36/article.
pdf.

Uhlig:2001:CWG

[Uhl01] Frank Uhlig. Constructive ways for generating (generalized) real orthogonal matrices as products of (generalized) symmetries. Linear Algebra and its Applications, 332–334(1):459–467, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
48/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/48/article.
pdf.

Uhlig:2003:NUB

[Uhl03] Frank Uhlig. A new unified, balanced, and conceptual approach to teaching linear algebra. Linear Algebra and its Applications, 361 (C):147–159, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
37/abstract.html.

Uhlig:2004:RIC

[Uhl04] Frank Uhlig. Report on the 10th ILAS Conference “Challenges in Matrix Theory” at Auburn University in June 2002. Linear Algebra and its Applications, 379(1):503– 535, March 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Unknown:2001:I

[Unk01] Unknown. Index. Linear Algebra and its Applications, 323(1–3):213, January 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Urbano:2004:NCP

[Urb04] Ana María Urbano. Nilpotent completions of partial upper triangular matrices. Linear Algebra and its Applications, 390(1):209– 254, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Valcher:2000:NLS

[Val00] Maria Elena Valcher. Nonnegative linear systems in the behavioral approach: the autonomous case. Linear Algebra and its Applications, 319 (1–3):147–162, November 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/138/24/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/

Valcher:2002:SSF

[Val02a] Maria Elena Valcher. On some special features which are peculiar to discrete time behaviors with trajectories on Z_{+}. Linear Algebra and its Applications, 351–352:719– 737, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vallejo:2002:CMM

[Val02b] Ernesto Vallejo. The classification of minimal matrices of size 2 × q. Linear Algebra and its Applications, 340(1– 3):169–181, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
39/abstract.html.

Valles:2008:CNT

[Val08] A. Fernández Valles. Common non-trivial invariant closed cones for commuting contractions. Linear Algebra and its Applications, 428(11– 12):2955–2960, June 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vallentin:2009:SSP

[Val09] Frank Vallentin. Symmetry in semidefinite programs.

Linear Algebra and its Applications, 430(1):360–369, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

VanIseghem:2004:SCF

[Van04] Jeannette Van Iseghem. Stieltjes continued fraction and QD algorithm: scalar, vector, and matrix cases. Linear Algebra and its Applications, 384(1):21–42, June 1, 2004. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

VanMieghem:2007:NTL

[Van07] Piet Van Mieghem. A new type of lower bound for the largest eigenvalue of a symmetric matrix. Linear Algebra and its Applications, 427(1):119–129, November 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Varah:2003:PDH

[Var03] J. M. Varah. Positive definite Hankel matrices of minimal condition. Linear Algebra and its Applications, 368 (C):303–314, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
44/abstract.html.

Varga:2004:GHC

[Var04] Richard S. Varga. Geršgorin and his circles, volume 36 of Springer series in computational mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-21100-4 (hardcover). ISSN 0179-3632. x + 226 pp. LCCN QA193 .V37 2004.

URL http://www.loc.gov/
catdir/enhancements/fy0815/
2004104814-d.html.

Vargas:2006:MLF

[Var06]  J. A. Vargas.  Matrices

of linear forms and curves. Linear Algebra and its Applications, 418(2–3):363–379, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Verwer:2009:USI

[VB09a] J. G. Verwer and M. A. Botchev. Unconditionally stable integration of Maxwell’s equations. Linear Algebra and its Applications, 431(3–4):300–317, July 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vonBelow:2009:ITU

[vB09b] Joachim von Below. An index theory for uniformly locally finite graphs. Linear Algebra and its Applications, 431(1– 2):1–19, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vaish:2004:WIM

[VC04] Akhil K. Vaish and N. Rao Chaganty. Wishartness and independence of matrix quadratic forms for Kronecker product covariance structures. Linear Algebra and its Applications, 388(1): 379–388, September 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanDam:2007:GGD

[vD07] E. R. van Dam. Graphs with given diameter maximizing the spectral radius. Linear Algebra and its Applications, 426(2–3):454–457, October 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vandenDriessche:2009:RCA

[vdD09] P. van den Driessche. Review of: A Combinatorial Approach to Matrix Theory and its Applications, by Richard A. Brualdi and Dragos Cvetković, CRC Press (2008). Linear Algebra and its Applications, 430(8–9): 2501–2502, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanDam:2003:WGD

[vDH03a] Edwin R. van Dam and Willem H. Haemers. Which graphs are determined by their spectrum? Linear Algebra and its Applications, 373(1):241–272, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanderHolst:2003:GWP

[vdH03b] Hein van der Holst. Graphs whose positive semi-definite matrices have nullity at most two. Linear Algebra and its Applications, 375(1):1–11, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanderHolst:2008:MCG

[vdH08a] Hein van der Holst. The maximum corank of graphs with a 2-separation. Linear Algebra and its Applications, 428 (7):1587–1600, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanderHolst:2008:TCG

[vdH08b] Hein van der Holst. Threeconnected graphs whose maximum nullity is at most three. Linear Algebra and its Applications, 429(2–3):625– 632, July 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanDam:2007:CGG

[vDHK07] E. R. van Dam, W. H. Haemers, and J. H. Koolen. Cospectral graphs and the generalized adjacency matrix. Linear Algebra and its Applications, 423(1):33– 41, May 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanDam:2007:MSR

[vDK07] E. R. van Dam and R. E. Kooij. The minimal spectral radius of graphs with a given diameter. Linear Algebra and its Applications, 423 (2–3):408–419, June 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanderMee:2006:FSP

[vdMRS06] Cornelis van der Mee, Giuseppe Rodriguez, and Sebastiano Seatzu. Fast superoptimal preconditioning of multiindex Toeplitz matrices. Linear Algebra and its Applications, 418(2–3):576–590, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanderMee:2001:SFB

[vdMSR01] Cornelis V. M. van der Mee, Sebastiano Seatzu, and Giuseppe Rodriguez. Spectral factorization of biinfinite multi-index block Toeplitz matrices. Linear Algebra and its Applications, 343–344(1):355–380, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/174/27/
43/abstract.html.

vanderMee:2003:SIM

[vdMSR03] Cornelis V. M. van der Mee, Sebastiano Seatzu, and Giuseppe Rodriguez. Semi-infinite multi-index perturbed block Toeplitz systems. Linear Algebra and its Applications, 366(C):459– 482, June 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
51/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

vanDoorn:2008:SQD

[vDP08] Erik A. van Doorn and Philip K. Pollett. Survival in a quasi-death process. Linear Algebra and its Applications, 429(4):776– 791, August 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanDam:2005:CDT

[vDS05] E. R. van Dam and E. Spence. Combinatorial designs with two singular values II. Partial geometric designs. Linear Algebra and its Applications, 396(1):303–316, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vandeVelden:2000:EPR

[vdVN00] Michel van de Velden and Heinz Neudecker. On an eigenvalue property relevant in correspondence analysis and related methods. Linear Algebra and its Applications, 321(1–3):347–364, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
47/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/47/article.
pdf.

vanderWoude:2001:SLM

[vdW01] Jacob W. van der Woude. A simplex-like method to compute the eigenvalue of an irreducible (max,+)-system. Linear Algebra and its Applications, 330(1–3):67–87, June 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/155/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/155/25/31/article.
pdf.

vanderWoude:2002:GCF

[vdW02] J. W. van der Woude. The generic canonical form of a regular structured matrix pencil. Linear Algebra and its Applications, 353(1–3): 267–288, September 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

vanderWoude:2003:ECS

[vdW03] J. W. van der Woude. On the existence of a common solution X to the matrix equations A_{i}XB_{j} = C_{ij}, (i, j) ⊂ Γ. Linear Algebra and its Applications, 375(1):135–145, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vladimirov:2000:SPD

[VEB00] Alexander Vladimirov, Ludwig Elsner, and Wolf-Jürgen Beyn. Stability and paracontractivity of discrete linear inclusions. Linear Algebra and its Applications, 312 (1–3):125–134, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/129/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/129/25/32/article.
pdf.

Vermeer:2008:OSR

[Ver08] J. Vermeer. Orthogonal similarity of a real matrix and its transpose. Linear Algebra and its Applications, 428 (1):382–392, January 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Veselic:2000:PTE

[Ves00] K. Veselić. Perturbation theory for the eigenvalues of factorised symmetric matrices. Linear Algebra and its Applications, 309(1–3): 85–102, April 15, 2000. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/126/25/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/126/25/31/article.
                                        pdf.

Veselic:2003:BES

[Ves03] Krešimir Veselić. Bounds for exponentially stable semigroups. Linear Algebra and its Applications, 358(1–3): 309–333, January 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/207/25/
                                        42/abstract.html.

Grudzinski:2009:SMP

[vGKN09] Olaf v. Grudzinski, Frieder Knüppel, and Klaus Nielsen.

Supernormal mappings: Part 1. Orthogonally indecomposable modules and applications. Linear Algebra and its Applications, 431(1–2): 39–55, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vardi:2001:SSR

[VHK01] Eilat Vardi, Gabor T. Herman, and T. Yung Kong. Speeding up stochastic reconstructions of binary images from limited projection directions. Linear Algebra and its Applications, 339(1– 3):75–89, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/169/27/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/169/27/32/article.
pdf.

VanBarel:2003:SMS

[VHK03] Marc Van Barel, Georg Heinig, and Peter Kravanja. A superfast method for solving Toeplitz linear least squares problems. Linear Algebra and its Applications, 366(C):441–457, June 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/215/25/
50/abstract.html. Special

issue on Structured Matrices: Analysis, Algorithms and Applications.

VanHuffel:2004:P

[VHM04] Sabine Van Huffel, Per Christian Hansen, and Peter Maass. Preface. Linear Algebra and its Applications, 391 (1):1, November 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vidunas:2007:NLP

[Vid07] Raimundas Vidūnas. Normalized Leonard pairs and Askey–Wilson relations. Linear Algebra and its Applications, 422(1):39–57, April 1, 2007. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vidar:2008:TPS

[Vid08] Melvin A. Vidar. Tridiagonal pairs of shape (1, 2, 1). Linear Algebra and its Applications, 429(1):403–428, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Villarreal:2008:RCM

[Vil08] Rafael H. Villarreal. Rees cones and monomial rings of matroids. Linear Algebra and its Applications, 428(11–12): 2933–2940, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vinod:2000:FMI

[Vin00]  H. D. Vinod.  Founda-

tions of multivariate inference using modern computers. Linear Algebra and its Applications, 321(1–3): 365–385, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/140/24/
48/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/48/article.
pdf.

Virnik:2008:SAP

[Vir08] Elena Virnik. Stability analysis of positive descriptor systems. Linear Algebra and its Applications, 429(10):2640– 2659, November 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Visick:2000:QVO

[Vis00] George Visick. A quantitative version of the observation that the Hadamard product is a principal submatrix of the Kronecker product. Linear Algebra and its Applications, 304(1–3): 45–68, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/118/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/118/25/27/article.
                                        pdf.

Vieira:2006:MVQ

[VJ06] A. C. Vieira and S. M. Alves Jorge. On minimal varieties of quadratic growth. Linear Algebra and its Applications, 418(2–3):925–938, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vera-Lopez:2003:CCU

[VLA03] Antonio Vera-López and J. M. Arregi. Conjugacy classes in unitriangular matrices. Linear Algebra and its Applications, 370(1):85– 124, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vera-Lopez:2004:CUM

[VLA04] Antonio Vera-López and J. M. Arregi. Computing in unitriangular matrices over finite fields. Linear Algebra and its Applications, 387 (1):193–219, August 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vera-Lopez:2008:ENC

[VLAOVL08] A. Vera-López, J. M. Arregi, Leyre Ormaetxea, and F. J. Vera-López. The exact number of conjugacy classes of the Sylow p-subgroups of GL(n, q) modulo (q − 1)^{13}. Linear Algebra and its Applications, 429(2–3):617– 624, July 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vincent:2004:ESF

[VM04] Jean-Marc Vincent and Corine Marchand. On the exact simulation of functionals of stationary Markov chains. Linear Algebra and its Applications, 386(1):285– 310, July 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Volaufova:2004:SEP

[Vol04] Júlia Volaufová. Some estimation problems in multistage linear models. Linear Algebra and its Applications, 388(1):389–397, September 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Voss:2007:NJJ

[Vos07]  Heinrich Voss.  A new

justification of the Jacobi– Davidson method for large eigenproblems. Linear Algebra and its Applications, 424 (2–3):448–455, July 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

VanBarel:2001:BHM

[VPV01] Marc Van Barel, Vlastimil Pták, and Zdeněk Vavřín. Bézout and Hankel matrices associated with row reduced matrix polynomials, Barnetttype formulas. Linear Algebra and its Applications, 332–334(1):583–606, August 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/157/25/
56/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/157/25/56/article.
pdf.

Viswanath:2006:MCL

[VR06] G. Viswanath and B. Sundar Rajan. Matrix characterization of linear codes with arbitrary Hamming weight hierarchy. Linear Algebra and its Applications, 412(2–3):396– 407, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Verde-Star:2004:GGP

[VS04a] Luis Verde-Star. Groups of generalized Pascal matrices. Linear Algebra and its Applications, 382(1):179– 194, May 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Verde-Star:2004:LMA

[VS04b]  Luis Verde-Star.  A La-

grange matrices approach to confluent Cauchy matrices. Linear Algebra and its Applications, 389(1):249–268, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Verde-Star:2005:FM

[VS05] Luis Verde-Star. Functions of matrices. Linear Algebra and its Applications, 406(1):285– 300, September 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vandebril:2006:NSC

[VV06] Raf Vandebril and Marc Van Barel. Necessary and sufficient conditions for orthogonal similarity transformations to obtain the Arnoldi(Lanczos)–Ritz values. Linear Algebra and its Applications, 414(2–3):435– 444, April 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vaserstein:2000:TDR

[VW00] L. Vaserstein and E. Wheland. Two-dimensional representations of the free group in two generators over an arbitrary field. Linear Algebra and its Applications, 307(1– 3):145–150, March 1, 2000.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/35/article.
pdf.

Vanluyten:2008:SNM

[VWD08] Bart Vanluyten, Jan C. Willems, and Bart De Moor. Structured nonnegative matrix factorization with applications to hidden Markov realization and clustering. Linear Algebra and its Applications, 429(7):1409–1424, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vong:2008:CAS

[VWJ08] Seak-Weng Vong, Wei Wang, and Xiao-Qing Jin. Convergence analysis of superoptimal PCG algorithm for Toeplitz systems with a Fisher–Hartwig singularity. Linear Algebra and its Applications, 428(2–3):535–549, January 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Vettori:2002:MTA

[VZ02] Paolo Vettori and Sandro Zampieri. Module theoretic approach to controllability of convolutional systems. Linear Algebra and its Applications, 351–352:739– 759, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2009:PEM

[WA09] Dong Q. Wang and S. E. Ahmed. Point estimation for multi-spectral distributed random matrices. Linear Algebra and its Applications, 430(10):2581–2591, May 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wada:2007:SRC

[Wad07] Shuhei Wada. On some refinement of the Cauchy– Schwarz inequality. Linear Algebra and its Applications, 420(2–3):433–440, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Walker:2005:LDQ

[Wal05a] Ronald A. Walker. Linear dependence of quotients of analytic functions of several variables with the least subcollection of generalized Wronskians. Linear Algebra and its Applications, 408(1):151– 160, October 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Walker:2005:IBN

[Wal05b] Stephen G. Walker. Improving bounds for nonmaximal eigenvalues of positive matrices. Linear Algebra and its Applications, 397(1):133– 139, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Waldron:2009:CEF

[Wal09] Shayne Waldron. On the construction of equiangular frames from graphs. Linear Algebra and its Applications, 431(11):2228–2242, November 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2001:CTQ

[Wan01] Tai-Lin Wang. Convergence of the tridiagonal qr algorithm. Linear Algebra and its Applications, 322(1– 3):1–17, January 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/141/24/
                                        24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/141/24/24/article.
                                        pdf.

Wang:2002:ROL

[Wan02] Guorong Wang. The reverse order law for the Drazin inverses of multiple matrix products. Linear Algebra and its Applications, 348(1– 3):265–272, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
47/abstract.html.

Wang:2003:LTP

[Wan03a] Yi Wang. Linear transformations preserving logconcavity. Linear Algebra and its Applications, 359 (1–3):161–167, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
38/abstract.html.

Wanless:2003:LBM

[Wan03b] Ian M. Wanless. A lower bound on the maximum permanent in Λ_{n}^{k}. Linear Algebra and its Applications, 373(1): 153–167, November 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2004:SME

[Wan04] Qing-Wen Wang. A system of matrix equations and a linear matrix equation over arbitrary regular rings with identity. Linear Algebra and its Applications, 384(1):43– 54, June 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wanat:2006:SDS

[Wan06] Magdalena Wanat. The αscalar diagonal stability of block matrices. Linear Algebra and its Applications, 414 (1):304–309, April 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2007:STS

[Wan07a] Li Wang. Semiconvergence of two-stage iterative methods for singular linear systems. Linear Algebra and its Applications, 422(2–3):824– 838, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2007:SPS

[Wan07b] Xiaochang Wang. On the solvability of polynomial systems arising in control. Linear Algebra and its Applications, 425(2–3):313–322, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2009:SPP

[Wan09a]  Heping Wang.  Shape-

preserving properties of ω, q-Bernstein polynomials. Linear Algebra and its Applications, 430(4):957–967, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2009:CMG

[Wan09b] Hsin-Ju Wang. Co-maximal graph of non-commutative rings. Linear Algebra and its Applications, 430(2–3):633– 641, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2009:UBS

[Wan09c] Ping Wang. An upper bound for ||A^{−1}||_{∞} of strictly diagonally dominant M-matrices. Linear Algebra and its Applications, 431(5–7):511–517, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2009:CCF

[Wan09d] Wei Wang. A counterexample to a conjecture of Friedland. Linear Algebra and its Applications, 430(8–9):2026– 2029, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2009:OHT

[Wan09e] Wen-Huan Wang. Ordering of Hückel trees according to minimal energies. Linear Algebra and its Applications, 430(2–3):703–717, January 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Watts:2000:TWC

[Wat00] Valerie L. Watts. A tree whose complement is not eigensharp. Linear Algebra and its Applications, 306 (1–3):183–188, February 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/121/25/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/121/25/39/article.
pdf.

Waterhouse:2001:ALC

[Wat01a] William C. Waterhouse. Asymmetric linear correlations of projective planes over fields. Linear Algebra and its Applications, 323(1–3): 99–104, January 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Watts:2001:BRK

[Wat01b] Valerie L. Watts. Boolean rank of Kronecker products. Linear Algebra and its Applications, 336(1–3): 261–264, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
45/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/166/27/45/article.
pdf.

Waterhouse:2004:PJD

[Wat04a] William C. Waterhouse. Polynomials for the Jordan decomposition in characteristic p. Linear Algebra and its Applications, 392(1):39– 44, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Watkins:2004:HSL

[Wat04b] David S. Watkins. On Hamiltonian and symplectic Lanczos processes. Linear Algebra and its Applications, 385 (1):23–45, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Waterhouse:2005:SAH

[Wat05] William C. Waterhouse. The structure of alternating-Hamiltonian matrices. Linear Algebra and its Applications, 396(1):385–390, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Waterhouse:2006:RCC

[Wat06] William C. Waterhouse. Real congruences of complex subspaces of 2 × 2 symmetric complex matrices. Linear Algebra and its Applications, 414(2–3):502–505, April 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2001:SCC

[WB01] Chuan-Long Wang and Zhong-Zhi Bai. Sufficient conditions for the convergent splittings of non-Hermitian positive definite matrices. Linear Algebra and its Applications, 330(1–3):215–218, June 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/155/25/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/155/25/38/article.
pdf.

Wang:2008:CCS

[WB08] Li Wang and Zhong-Zhi Bai. Convergence conditions for splitting iteration methods for non-Hermitian linear systems. Linear Algebra and its Applications, 428(2–3):453– 468, January 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Chu:2004:PEP

[wCL04] Eric King wah Chu and Wen-Wei Lin. Perturbation of eigenvalues for periodic matrix pairs via the Bauer–Fike theorems. Linear Algebra and its Applications, 378(1):183– 202, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wei:2005:GIS

[WD05] Yimin Wei and Huaian Diao. On group inverse of singular Toeplitz matrices. Linear Algebra and its Applications, 399(1):109–123, April 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2006:HPN

[WD06] Yaokun Wu and Aiping Deng. Hoffman polynomials of nonnegative irreducible matrices and strongly connected digraphs. Linear Algebra and its Applications, 414 (1):138–171, April 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2009:MSA

[Wd09] Shun Wang and Eric Sturler de. Multilevel sparse approximate inverse preconditioners for adaptive mesh refinement. Linear Algebra and its Applications, 431(3– 4):409–426, July 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wenchang:2009:FDD

[Wen09] Chu Wenchang. Finite differences and determinant identities. Linear Algebra and its Applications, 430(1):215– 228, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Werpachowski:2008:ARP

[Wer08] Roman Werpachowski. On the approximation of real powers of sparse, infinite, bounded and Hermitian matrices. Linear Algebra and its Applications, 428(1):316– 323, January 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Werpachowski:2009:SGD

[Wer09] Roman Werpachowski. On the solutions of generalized discrete Poisson equation. Linear Algebra and its Applications, 430(8–9):1877– 1885, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2004:LBP

[WF04] Mingsheng Wang and Dengguo Feng. On Lin–Bose problem. Linear Algebra and its Applications, 390(1):279– 285, October 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2008:LMP

[WFJ08] Meili Wang, Li Fang, and Guoxing Ji. Linear maps preserving idempotency of products or triple Jordan products of operators. Linear Algebra and its Applications, 429(1):181–189, July 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wei:2002:IBM

[WG02a] Musheng Wei and Wenbin Guo. On g-inverses of a bordered matrix: revisited. Linear Algebra and its Applications, 347(1–3):189–204, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
38/abstract.html.

Wei:2002:ROL

[WG02b] Musheng Wei and Wenbin Guo. Reverse order laws for least squares g-inverses and minimum norm g-inverses of products of two matrices. Linear Algebra and its Applications, 342(1–3): 117–132, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/173/27/
35/abstract.html.

Wang:2009:SST

[WGL09] Kaishun Wang, Jun Guo, and Fenggao Li. Suborbits of subspaces of type (m, k) under finite singular general linear groups. Linear Algebra and its Applications, 431(8):1360–1366, September 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2009:NRI

[WGT09] Pei Yuan Wu, Hwa-Long Gau, and Ming Cheng Tsai. Numerical radius inequality for C_{0} contractions. Linear Algebra and its Applications, 430(5–6):1509–1516, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2008:SNG

[WHAB08] Jianfeng Wang, Qiongxiang Huang, Xinhui An, and Francesco Belardo. Some notes on graphs whose spectral radius is close to \frac{3}{2} \sqrt{2}. Linear Algebra and its Applications, 429(7):1606–1618, October 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2009:NSC

[WHBM09a] Jianfeng Wang, Qiongxiang Huang, Francesco Belardo, and Enzo M. Li Marzi. A note on the spectral characterization of dumbbell graphs. Linear Algebra and its Applications, 431(10):1707–1714, October 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2009:GWS

[WHBM09b] Jianfeng Wang, Qiongxiang Huang, Francesco Belardo, and Enzo M. Li Marzi. On graphs whose signless Laplacian index does not exceed 4.5. Linear Algebra and its Applications, 431(1–2):162– 178, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

White:2000:DDS

[Whi00] R. E. White. Domain decomposition splittings. Linear Algebra and its Applications, 316(1–3):105–112, September 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/134/24/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/134/24/31/article.
pdf.

Wiesner:2001:BMP

[Wie01] Emilie B. Wiesner. Backward minimal points for bounded linear operators on finitedimensional vector spaces. Linear Algebra and its Applications, 338(1–3):251–259, November 15, 2001. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
45/abstract.html;   http:
//www.elsevier.nl/gej-ng/

                                        10/30/19/168/27/45/article.
                                        pdf.

Wilkening:2007:ACJ

[Wil07] Jon Wilkening. An algorithm for computing Jordan chains and inverting analytic matrix functions. Linear Algebra and its Applications, 427(1): 6–25, November 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wild:2009:IAU

[Wil09] Marcel Wild. Incidence algebras that are uniquely determined by their zero– nonzero matrix pattern. Linear Algebra and its Applications, 430(4):1007–1016, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wimmer:2001:SBC

[Wim01] Harald K. Wimmer. Similarity of block companion and block Toeplitz matrices. Linear Algebra and its Applications, 343–344(1): 381–387, March 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/174/27/
                                        44/abstract.html.

Winkler:2000:RMS

[Win00] Joab R. Winkler. A resultant matrix for scaled Bernstein polynomials. Linear Algebra and its Applications, 319(1–3):179–191, November 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/138/24/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/138/24/35/article.
pdf.

Winkler:2003:CMR

[Win03] Joab R. Winkler. A companion matrix resultant for Bernstein polynomials. Linear Algebra and its Applications, 362(C):153–175, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/211/25/
38/abstract.html.

Wirth:2002:GSR

[Wir02] Fabian Wirth. The generalized spectral radius and extremal norms. Linear Algebra and its Applications, 342(1– 3):17–40, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/173/27/
28/abstract.html.

Wirth:2005:GSR

[Wir05] Fabian Wirth. The generalized spectral radius is strictly increasing. Linear Algebra and its Applications, 395(1): 141–153, January 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wirth:2009:BDM

[Wir09] Jens Wirth. Block-diagonalisation of matrices and operators. Linear Algebra and its Applications, 431(5–7):895–902, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2002:CSM

[WJL02] Yao-Kun Wu, Rui-Zhong Jia, and Qiao Li. g-Circulant solutions to the (0, 1) matrix equation A^{m} = J_{n}. Linear Algebra and its Applications, 345(1–3):195–224, April 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/175/27/
                                        36/abstract.html.

Wang:2009:OTP

[WK09] Wen-Huan Wang and Li-Ying Kang. Ordering of the trees with a perfect matching by minimal energies. Linear Algebra and its Applications, 431(5–7):946– 961, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2003:AS

[WL03] Yaokun Wu and Qiao Li. An approach to solving A^{k} = J − I. Linear Algebra and its Applications, 373(1):121– 142, November 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wei:2007:CSL

[WL07] Qiong Wei and Pengtong Li. Centralizers of J -subspace lattice algebras. Linear Algebra and its Applications, 426 (1):228–237, October 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2008:LAV

[WL08] Kaishun Wang and Zengti Li. Lattices associated with vector spaces over a finite field. Linear Algebra and its Applications, 429(2–3):439– 446, July 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2009:LTD

[WL09a] Heng-Tai Wang and Qing-Guo Li. Lie triple derivation of the Lie algebra of strictly upper triangular matrix over a commutative ring. Linear Algebra and its Applications, 430(1):66– 77, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2009:RLN

[WL09b] Qing-Wen Wang and Cheng-Kun Li. Ranks and the leastnorm of the general solution to a system of quaternion matrix equations. Linear Algebra and its Applications, 430(5–6):1626–1640, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wei:2006:RPB

[WLBZ06] Yimin Wei, Xiezhang Li, Fanbin Bu, and Fuzhen Zhang. Relative perturbation bounds for the eigenvalues of diagonalizable and singular matrices — Application of perturbation theory for simple invariant subspaces. Linear Algebra and its Applications, 419(2–3): 765–771, December 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2007:SCM

[WLL07] Jinn-Wen Wu, Kuang-Yow Lian, and Peter Lancaster. A stabilization criterion for matrices. Linear Algebra and its Applications, 422(1):22– 28, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wen:2007:CTI

[WlW07] Jiajin Wen and Wan lan Wang. Chebyshev type inequalities involving permanents and their applications. Linear Algebra and its Applications, 422(1):295– 303, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2002:SRC

[WM02] Li Yun Wang and Dao Ji Meng. Some results on complete Lie superalgebras. Linear Algebra and its Applications, 355(1–3):1–14, November 1, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/201/27/
28/abstract.html.

Wang:2003:SCL

[WM03] Li Yun Wang and Dao Ji Meng. Some complete Lie superalgebras. Linear Algebra and its Applications, 369 (C):339–349, August 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/218/25/
46/abstract.html.

Wang:2005:NMG

[WM05] Chuan-Long Wang and Guo-Yan Meng. Nonstationary multisplittings with general weighting matrices for mildly nonlinear systems. Linear Algebra and its Applications, 407(1):1–10, September 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Willson:2005:ERD

[WO05] A. N. Willson, Jr. and H. J. Orchard. An efficient resultant for determining reciprocal zeros in polynomials. Linear Algebra and its Applications, 411(1):309–327, December 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Woerdeman:2002:CTP

[Woe02] Hugo J. Woerdeman. The Carathéodory-Toeplitz problem with partial data. Linear Algebra and its Applications, 342(1–3):149–161, February 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/173/27/
                                        37/abstract.html.

Woerdeman:2004:SPN

[Woe04] Hugo J. Woerdeman. The separability problem and normal completions. Linear Algebra and its Applications, 376(1):85–95, January 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wojciechowski:2003:DMP

[Woj03] Piotr J. Wojciechowski. Directed maximal partial orders of matrices. Linear Algebra and its Applications, 375 (1):45–49, December 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wood:2002:KPE

[Woo02] Jeffrey Wood. Key problems in the extension of modulebehaviour duality. Linear Algebra and its Applications, 351–352:761–798, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wright:2006:RMH

[Wri06] Stephen E. Wright. Real matrix homotopies based on Householder reflections. Linear Algebra and its Applications, 412(2–3):538–545, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2002:CBS

[WSL02] Qing-Wen Wang, Jian-Hua Sun, and Shang-Zhi Li. Consistency for bi(skew)symmetric solutions to systems of generalized Sylvester equations over a finite central algebra. Linear Algebra and its Applications, 353(1–3):169–182, September 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2009:IES

[WSL09] Bao-Feng Wu, Jia-Yu Shao, and Yue Liu. Interlacing eigenvalues on some operations of graphs. Linear Algebra and its Applications, 430(4):1140–1150, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wittig:2006:MOR

[WSW06] Tilmann Wittig, Rolf Schuhmann, and Thomas Weiland. Model order reduction for large systems in computational electromagnetics. Linear Algebra and its Applications, 415(2–3):499– 530, June 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Tan:2009:SRW

[wThY09] Shang wang Tan and Yan hong Yao. On the spectral radius of weighted trees with fixed diameter and weight set. Linear Algebra and its Applications, 431(1–2): 86–98, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2002:NRA

[Wu02] Pei Yuan Wu. Numerical range of Aluthge transform of operator. Linear Algebra and its Applications, 357(1–3): 295–298, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/203/25/
46/abstract.html.

Wu:2005:BEE

[Wu05a] Chai Wah Wu. On bounds of extremal eigenvalues of irreducible and m-reducible matrices. Linear Algebra and its Applications, 402(1):29– 45, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2005:RRR

[Wu05b] Chai Wah Wu. On Rayleigh– Ritz ratios of a generalized Laplacian matrix of directed graphs. Linear Algebra and its Applications, 402(1):207– 227, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2005:MNP

[Wu05c]  Huazhang Wu.  On the

multiple Nevanlinna–Pick interpolation problems in some matrix-valued function classes. Linear Algebra and its Applications, 399(1):71– 90, April 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2006:DTR

[Wu06a] Gang Wu. A dynamic thick restarted semi-refined ABLE algorithm for computing a few selected eigentriplets of large nonsymmetric matrices. Linear Algebra and its Applications, 416(2–3):313– 335, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2006:OBS

[Wu06b]  Guoxing Wu.  Optimal

bounds for the spectral variation of two regular matrix pairs. Linear Algebra and its Applications, 418(2–3):891– 899, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2006:EPP

[Wu06c] Yaokun Wu. Even poset and a parity result for binary linear code. Linear Algebra and its Applications, 418(2–3): 591–594, October 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2008:SPC

[Wu08a] Chai Wah Wu. On some properties of contracting matrices. Linear Algebra and its Applications, 428(11–12): 2509–2523, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2008:UBS

[Wu08b] Guoxing Wu. On upper bounds for the spectral variation of two regular matrix pairs. Linear Algebra and its Applications, 429(2– 3):497–502, July 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2008:ALE

[Wu08c] Jian-Hong Wu. Admissibility of linear estimators in multivariate linear models with respect to inequality constraints. Linear Algebra and its Applications, 428(8– 9):2040–2048, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2008:ISD

[Wu08d] Yaokun Wu. Invariant subspace, determinant and characteristic polynomials. Linear Algebra and its Applications, 428(7):1753–1760, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2009:JAD

[Wu09] Jing Wu. On Jordan allderivable points of B(H).

Linear Algebra and its Applications, 430(4):941–946, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Walker:2008:LBL

[WV08] S. G. Walker and P. Van Mieghem. On lower bounds for the largest eigenvalue of a symmetric matrix. Linear Algebra and its Applications, 429(2–3):519–526, July 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2009:SRQ

[WvdWC09] Qing-Wen Wang, J. W. van der Woude, and Hai-Xia Chang. A system of real quaternion matrix equations with applications. Linear Algebra and its Applications, 431(12):2291–2303, December 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2002:INN

[wW02] Ming wei Wang. An inequality for non-negative matrices II. Linear Algebra and its Applications, 348(1– 3):259–264, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/

gej-ng/10/30/19/195/27/
46/abstract.html.

Wang:2007:MRB

[WW07] Weiping Wang and Tianming Wang. Matrices related to the Bell polynomials. Linear Algebra and its Applications, 422(1):139– 154, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wallace:2008:PFI

[WW08a] Dorothy Wallace and Walker White. Pseudoprime factorizations of integer matrices. Linear Algebra and its Applications, 429(1):142– 155, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Winkler:2008:SCS

[WW08b] Henrik Winkler and Harald Woracek. On semibounded canonical systems. Linear Algebra and its Applications, 429(5–6):1082–1092, September 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2009:MTR

[WWJ09] Minghui Wang, Musheng Wei, and Zhigang Jia. Mixedtype reverse order law of ( AB ) ^{(13)}. Linear Algebra and its Applications, 430(5–6):1691– 1699, March 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2004:GIT

[WWQ04] Guorong Wang, Yimin Wei, and Sanzheng Qiao. Generalized inverses: theory and computations, volume 5 of Graduate series in mathematics. Science Press, Beijing, China, 2004. ISBN 7-03-012437-5 (hardcover). xi + 294 pp. LCCN QA188 .W362 2004.

Wang:2006:EAT

[WX06a] Wei Wang and Cheng-Xian Xu. An excluding algorithm for testing whether a family of graphs are determined by their generalized spectra. Linear Algebra and its Applications, 418(1):62– 74, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2006:NST

[WX06b] Wei Wang and Cheng-Xian Xu. Note the T -shape tree is determined by its Laplacian spectrum. Linear Algebra and its Applications, 419 (1):78–81, November 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2006:SCS

[WX06c] Wei Wang and Cheng-Xian Xu. On the spectral characterization of T -shape trees.

Linear Algebra and its Applications, 414(2–3):492–501, April 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2007:NGS

[WX07a] Wei Wang and Cheng-Xian Xu. Note: On the generalized spectral characterization of graphs having an isolated vertex. Linear Algebra and its Applications, 425(1):210– 215, August 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2007:SRS

[WX07b] Wei Wang and Cheng-Xian Xu. Some results on the spectral reconstruction problem. Linear Algebra and its Applications, 427(1):151– 159, November 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2005:SRT

[WXH05] Baofeng Wu, Enli Xiao, and Yuan Hong. The spectral radius of trees on k pendant vertices. Linear Algebra and its Applications, 395(1):343– 349, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2004:DJA

[WY04] Xing Tao Wang and Hong You. Decomposition of Jordan automorphisms of strictly triangular matrix algebra over local rings. Linear Algebra and its Applications, 392(1):183–193, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2005:DGI

[WY05] Susan Y. J. Wu and Arthur L. B. Yang. Division and the Giambelli identity. Linear Algebra and its Applications, 406(1):301–309, September 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2006:DPS

[WY06a] Dengyin Wang and Qiu Yu. Derivations of the parabolic subalgebras of the general linear Lie algebra over a commutative ring. Linear Algebra and its Applications, 418(2–3):763–774, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2006:DLA

[WY06b] Xing Tao Wang and Hong You. Decomposition of Lie automorphisms of upper triangular matrix algebra over commutative rings. Linear Algebra and its Applications, 419(2–3):466–474, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Willems:2007:BDR

[WY07] Jan C. Willems and Yutaka Yamamoto. Behaviors defined by rational functions. Linear Algebra and its Applications, 425(2–3):226–241, September 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2008:MUT

[WY08] Zhongying Wang and Hong You. Maps on uppertriangular matrix algebras preserving k-potences. Linear Algebra and its Applications, 429(8–9):1915–1928, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wyman:2002:PZS

[Wym02] Bostwick F. Wyman. Poles, zeros, and sheaf cohomology. Linear Algebra and its Applications, 351–352:799–807, August 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2007:ALL

[WYZ07] Dengyin Wang, Qiu Yu, and Yanxia Zhao. Automorphisms of a linear Lie algebra over a commutative ring. Linear Algebra and its Applications, 423(2–3):324– 331, June 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2007:PAO

[WZ07] Yuwen Wang and Hao Zhang. Perturbation analysis for oblique projection generalized inverses of closed linear operators in Banach spaces. Linear Algebra and its Applications, 426(1):1– 11, October 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wu:2008:EES

[WZCL08] Junliang Wu, Limin Zou, Xiangping Chen, and Shengjie Li. The estimation of eigenvalues of sum, difference, and tensor product of matrices over quaternion division algebra. Linear Algebra and its Applications, 428(11–12): 3023–3033, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Wang:2000:HPI

[WZZ00] Bo-Ying Wang, Xiuping Zhang, and Fuzhen Zhang.

On the Hadamard product of inverse M-matrices. Linear Algebra and its Applications, 305(1–3):23–31, January 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/120/25/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/120/25/27/article.
pdf.

Xiao:2001:MAP

[XCX01] MinTang Xiao, GuangCao Chong, and Zhang Xian. Modular automorphisms preserving idempotence and Jordan isomorphisms of triangular matrices over commutative rings. Linear Algebra and its Applications, 338(1–3): 145–152, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/38/article.
pdf.

Xie:2003:SME

[XD03] Huiqing Xie and Hua Dai. On the sensitivity of multiple eigenvalues of nonsymmetric matrix pencils. Linear Algebra and its Applications, 374(1):143–158, November 15, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2008:MAA

[XDS08] Genjiu Xu, Theo S. H. Driessen, and Hao Sun. Matrix analysis for associated consistency in cooperative game theory. Linear Algebra and its Applications, 428 (7):1571–1586, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2009:MAD

[XDS09] Genjiu Xu, Theo Driessen, and Hao Sun. Matrix approach to dual similar associated consistency for the Shapley value. Linear Algebra and its Applications, 430(11–12):2896– 2897, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2008:PFM

[XH08] Qingxiang Xu and Xiaoxia Hu. Particular formulae for the Moore–Penrose inverses of the partitioned bounded linear operators. Linear Algebra and its Applications, 428(11–12):2941– 2946, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xie:2006:SCI

[XHpS06] Dongxiu Xie, Xiyan Hu, and Yan ping Sheng. The solvability conditions for the inverse eigenproblems of symmetric and generalized centro-symmetric matrices and their approximations. Linear Algebra and its Applications, 418(1):142– 152, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2009:MSR

[XHSZ09] Mimi Xu, Yuan Hong, Jinlong Shu, and Mingqing Zhai. The minimum spectral radius of graphs with a given independence number. Linear Algebra and its Applications, 431(5–7):937– 945, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xiang:2003:IHP

[Xia03] Shuhuang Xiang. On an inequality for the Hadamard product of an M-matrix or an H-matrix and its inverse. Linear Algebra and its Applications, 367(C):17–27, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
27/abstract.html.

Xiang:2006:NMR

[Xia06] Hua Xiang. A note on the minimax representation for the subspace distance and singular values. Linear Algebra and its Applications, 414(2–3):470–473, April 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xie:2008:REE

[XJ08] Bao-Hua Xie and Yue-Ping Jiang. P U(2, 1) with regular elliptic elements. Linear Algebra and its Applications, 428(4):732–737, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xie:2004:NLA

[XL04] Jinhai Xie and Fangyan Lu. A note on 2-local automorphisms of digraph algebras. Linear Algebra and its Applications, 378(1):93– 98, February 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xuezhong:2005:NUG

[XL05] Tan Xuezhong and Bolian Liu. On the nullity of unicyclic graphs. Linear Algebra and its Applications, 408 (1):212–220, October 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2013:CSN

[XLG^{+}13] Wei-Ru Xu, Ying-Jie Lei, Xian-Ming Gu, Yong Lu, and Yan-Ru Niu. Comment on “A note on the inverse eigenvalue problem for symmetric doubly stochastic matrices”. Linear Algebra and its Applications, 439(8): 2256–2262, October 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379513003881.

See [HP04c, Fan10].

Xin:2002:SSS

[XM02] Xin Xin and Tsutomu Mita. A simple state-space design of an interactor for a nonsquare system via system matrix pencil approach. Linear Algebra and its Applications, 351–352:809–823, August 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2008:FSS

[XQ08] Wei Xu and Sanzheng Qiao. A fast symmetric SVD algorithm for square Hankel matrices. Linear Algebra and its Applications, 428(2–3):550– 563, January 15, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2008:NST

[XQW08] Wei Xu, Sanzheng Qiao, and Yimin Wei. A note on the scaled total least squares problem. Linear Algebra and its Applications, 428(2–3): 469–478, January 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2004:CM

[XS04] Guang-Hui Xu and Jia-Yu Shao. Conditional Smatrices. Linear Algebra and its Applications, 376(1):187– 206, January 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2008:PSD

[XS08] Qingxiang Xu and Lijuan Sheng. Positive semi-definite matrices of adjointable operators on Hilbert C^{∗}-modules. Linear Algebra and its Applications, 428(4):992–1000, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2008:SSO

[XSG08] Qingxiang Xu, Lijuan Sheng, and Yangyang Gu. The solutions to some operator equations. Linear Algebra and its Applications, 429(8–9):1997– 2024, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2001:PAM

[Xu01] Shu-Fang Xu. Perturbation analysis of the maximal solution of the matrix equation X + A^{∗}X^{−1}A = P. Linear Algebra and its Applications, 336(1–3): 61–70, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/166/27/32/article.
pdf.

Xu:2003:SLM

[Xu03] Hongguo Xu. An SVD-like matrix decomposition and its applications. Linear Algebra and its Applications, 368(C):1–24, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
26/abstract.html.

Xu:2004:CPM

[Xu04] Changqing Xu. Completely positive matrices. Linear Algebra and its Applications, 379(1):319–327, March 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2006:EPD

[Xu06] Hongguo Xu. On equivalence of pencils from discretetime and continuous-time control. Linear Algebra and its Applications, 414(1):97– 124, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2007:TBD

[Xu07] Hongguo Xu. Transformations between discrete-time and continuous-time algebraic Riccati equations. Linear Algebra and its Applications, 425(1):77–101, August 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2008:CHP

[Xu08] Qingxiang Xu. Common hermitian and positive solutions to the adjointable operator equations AX = C, XB = D. Linear Algebra and its Applications, 429(1): 1–11, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2009:MPI

[Xu09]  Qingxiang Xu.  Moore–

Penrose inverses of partitioned adjointable operators on Hilbert C^{∗}-modules. Linear Algebra and its Applications, 430(11–12):2929– 2942, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-

Xue:2001:BPT

[Xue01] Jungong Xue. Blockwise perturbation theory for nearly uncoupled Markov chains and its application. Linear Algebra and its Applications, 326(1–3):173–191, March 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/148/24/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/148/24/35/article.
pdf.

Xiang:2006:PAG

[XWD06] Hua Xiang, Yimin Wei, and Huaian Diao. Perturbation analysis of generalized saddle point systems. Linear Algebra and its Applications, 419 (1):8–23, November 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xi:2003:CAC

[XX03] Changchang Xi and Dajing Xiang. Cellular algebras and Cartan matrices. Linear Algebra and its Applications, 365(C):369–388, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
35/abstract.html.

Xu:2009:SBS

[XX09] Guang-Hui Xu and Chang-Qing Xu. Sharp bounds for the spectral radius of digraphs. Linear Algebra and its Applications, 430(5– 6):1607–1612, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xu:2009:RHM

[XXZ09] Changqing Xu, Zhaodi Xu, and Fuzhen Zhang. Revisiting hua-marcus-bellman-ando inequalities on contractive matrices. Linear Algebra and its Applications, 430(5– 6):1499–1508, March 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Xiang:2006:CAB

[XZ06] Shuhuang Xiang and Shenglei Zhang. A convergence analysis of block accelerated over-relaxation iterative methods for weak block H-matrices to partition π. Linear Algebra and its Applications, 418(1):20– 32, October 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhu:2009:SIE

[xZxGpL09] Sheng xin Zhu, Tong xiang Gu, and Xing ping Liu. Solving inverse eigenvalue problems via Householder and rank-one matrices. Linear Algebra and its Applications, 430(1):318– 334, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yagle:2001:CCM

[Yag01] Andrew E. Yagle. A convergent composite mapping Fourier domain iterative algorithm for 3-D discrete tomography. Linear Algebra and its Applications, 339 (1–3):91–109, December 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/169/27/
33/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/169/27/33/article.
pdf.

Yahaghi:2004:ISM

[Yah04] Bamdad R. Yahaghi. On irreducible semigroups of matrices with traces in a subfield. Linear Algebra and its Applications, 383(1):17– 28, May 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yahaghi:2006:AAM

[Yah06] Bamdad R. Yahaghi. On F -algebras of algebraic matrices over a subfield F of the center of a division ring. Linear Algebra and its Applications, 418(2–3):599–613, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yahaghi:2008:STC

[Yah08] Bamdad R. Yahaghi. On simultaneous triangularization of collections of compact operators. Linear Algebra and its Applications, 428(4): 1151–1168, February 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yamamoto:2001:SME

[Yam01] Nobito Yamamoto. A simple method for error bounds of eigenvalues of symmetric matrices. Linear Algebra and its Applications, 324(1–3): 227–234, February 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/144/24/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/144/24/36/article.
                                        pdf.

Yamazaki:2002:NRA

[Yam02] Takeaki Yamazaki. On numerical range of the Aluthge transformation. Linear Algebra and its Applications, 341 (1–3):111–117, January 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/171/27/
35/abstract.html.

Yamazaki:2006:EKI

[Yam06] Takeaki Yamazaki. An extension of Kantorovich inequality to n-operators via the geometric mean by Ando– Li–Mathias. Linear Algebra and its Applications, 416(2– 3):688–695, July 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2001:OPI

[Yan01a] Changsen Yang. An order preserving inequality via Furuta inequality, II. Linear Algebra and its Applications, 331(1–3):89–100, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/
32/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/32/article.
pdf.

Yang:2001:PBM

[Yan01b] Z. H. Yang. Polynomial Bezoutian matrix with respect to a general basis. Linear Algebra and its Applications, 331(1–3):165–179, July 1, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/156/25/

38/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/156/25/38/article.
pdf.

Yang:2002:CCS

[Yan02a] Weiguo Yang. Convergence in the Cesàro sense and strong law of large numbers for nonhomogeneous Markov chains. Linear Algebra and its Applications, 354(1–3): 275–288, October 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/200/27/
48/abstract.html.

Yang:2002:MI

[Yan02b] Xiaojing Yang. A matrix inequality. Linear Algebra and its Applications, 348 (1–3):41–47, June 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/195/27/
30/abstract.html.

Yang:2003:NCH

[Yan03] Xiaojing Yang. Necessary conditions of Hurwitz polynomials. Linear Algebra and its Applications, 359(1–3): 21–27, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
27/abstract.html.

Yang:2006:CCI

[Yan06] Qingjie Yang. Conjugacy classes in integral symplectic groups. Linear Algebra and its Applications, 418(2–3): 614–624, October 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See erratum [Yan07].

Yang:2007:ECC

[Yan07] Qingjie Yang. Erratum to: “Conjugacy classes in integral symplectic groups” [Linear Algebra Appl. 418 (2006) 614–624]. Linear Algebra and its Applications, 420(2– 3):726, January 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). See [Yan06].

Yang:2009:SLL

[Yan09a] Weiguo Yang. Strong law of large numbers for countable nonhomogeneous Markov chains. Linear Algebra and its Applications, 430(11–12): 3008–3018, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2009:GLF

[Yan09b]  Yongzhi Yang.  General-

ized Leibniz functional matrices and factorizations of some well-known matrices. Linear Algebra and its Applications, 430(1):511–531, January 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2000:GCCa

[YC00a] Zheng-Hong Yang and Gong-Ning Chen. Generalizedconfluent Cauchy and Cauchy– Vandermonde matrices. Linear Algebra and its Applications, 308(1–3):31–64, March 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/125/25/
26/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/125/25/26/article.
pdf.

Yong:2000:DCE

[YC00b] Xuerong Yong and Shiyi Cao. Disproof of a conjecture on the existence of the pathrecursive period for a connected graph. Linear Algebra and its Applications, 317 (1–3):201–205, September 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/136/25/
34/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/136/25/34/article.
pdf.

You:2001:GIS

[YC01] Hong You and Jianlong Chen. Generalized inverses of a sum of morphisms. Linear Algebra and its Applications, 338(1–3):261–273, November 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/168/27/
46/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/168/27/46/article.
pdf.

Ye:2008:OES

[Ye08] Qiang Ye. Optimal expansion of subspaces for eigenvector approximations. Linear Algebra and its Applications, 428(4):911–918, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yekutieli:2004:SB

[Yek04] Amnon Yekutieli. On the structure of behaviors. Linear Algebra and its Applications, 392(1):159–181, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yeredor:2004:MDE

[Yer04] Arie Yeredor. Multiple delays estimation for chirp signals using structured total least squares. Linear Algebra and its Applications, 391(1):261– 286, November 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yeung:2004:PAC

[Yeu04] Man-Chung Yeung. Probabilistic analysis of complex Gaussian elimination without pivoting. Linear Algebra and its Applications, 384(1):109– 134, June 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yanagi:2005:TIT

[YFK05] Kenjiro Yanagi, Shigeru Furuichi, and Ken Kuriyama. On trace inequalities and their applications to noncommutative communication theory. Linear Algebra and its Applications, 395(1):351– 359, January 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Ye:2009:LEG

[YFL09] Miao-Lin Ye, Yi-Zheng Fan, and Dong Liang. The least eigenvalue of graphs with given connectivity. Linear Algebra and its Applications, 430(4):1375–1379, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yopp:2000:CCD

[YH00a] David A. Yopp and Richard D. Hill. On completely copositive and decomposable linear transformations. Linear Algebra and its Applications, 312(1–3):1–12, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/129/25/
25/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/129/25/25/article.
pdf.

Yopp:2000:CCP

[YH00b] David A. Yopp and Richard D. Hill. On the cone of completely positive linear transformations. Linear Algebra and its Applications, 304(1– 3):119–129, January 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/118/25/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/118/25/31/article.
pdf.

Yang:2004:FBM

[YH04a] Yongzhi Yang and Heidi Holtti. The factorization of block matrices with generalized geometric progression rows. Linear Algebra and its Applications, 387(1):51– 67, August 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2004:CPV

[YH04b] Zheng-Hong Yang and Yong-Jian Hu. Confluent polynomial Vandermonde-like matrices: displacement structures, inversion formulas and fast algorithm. Linear Algebra and its Applications, 382 (1):61–82, May 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yin:2004:CRA

[Yin04] Qingxiang Yin. Construction of real antisymmetric and bi-antisymmetric matrices with prescribed spectrum data. Linear Algebra and its Applications, 389(1):95–106, September 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yin:2005:CRS

[Yin05] Qingxiang Yin. Construction of real symmetric and per-antisymmetric matrices with prescribed spectrum data. Linear Algebra and its Applications, 408(1):275– 283, October 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yanagi:2005:GSI

[YKF05] Kenjiro Yanagi, Ken Kuriyama, and Shigeru Furuichi. Generalized Shannon inequalities based on Tsallis relative operator entropy. Linear Algebra and its Applications, 394 (1):109–118, January 1, 2005. CODEN LAAPAW. ISSN

Yu:2006:MET

[YL06] Aimei Yu and Xuezheng Lv. Minimum energy on trees with k pendent vertices. Linear Algebra and its Applications, 418(2–3):625– 633, October 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yu:2008:LSR

[YL08] Aimei Yu and Mei Lu. Laplacian spectral radius of trees with given maximum degree. Linear Algebra and its Applications, 429(8–9):1962–1969, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yin:2009:SPE

[YLF09] Xiaoyan Yin, Sanyang Liu, and Liang Fang. Solutions and perturbation estimates for the matrix equation X^{s} + A^{∗}X^{−t}A = Q. Linear Algebra and its Applications, 431(9):1409–1421, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

You:2005:IND

[YLS05] Lihua You, Bolian Liu, and Jian Shen. r-indecomposable and r-nearly decomposable matrices. Linear Algebra and its Applications, 407(1): 105–116, September 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yu:2004:SRG

[YLT04] Aimei Yu, Mei Lu, and Feng Tian. On the spectral radius of graphs. Linear Algebra and its Applications, 387(1): 41–49, August 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yu:2005:CGW

[YLT05a] Aimei Yu, Mei Lu, and Feng Tian. Characterization on graphs which achieve a Das’ upper bound for Laplacian spectral radius. Linear Algebra and its Applications, 400(1):271–277, May 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yu:2005:OTT

[YLT05b] Aimei Yu, Mei Lu, and Feng Tian. Ordering trees by their Laplacian spectral radii. Linear Algebra and its Applications, 405(1):45– 59, August 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2009:FII

[YLT09] Zhongpeng Yang, Shuangzhe Liu, and Götz Trenkler.

Further inequalities involving the Khatri–Rao product. Linear Algebra and its Applications, 430(10):2696– 2704, May 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2009:MSI

[YLX09] Haijian Yang, Qingguo Li, and Hongru Xu. A multiplicative Schwarz iteration scheme for solving the linear complementarity problem with an H-matrix. Linear Algebra and its Applications, 430(4):1085–1098, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

You:2006:ROP

[YLZ06] Hong You, Shaowu Liu, and Guodong Zhang. Rank one preserving R-linear maps on spaces of self-adjoint operators on complex Hilbert spaces. Linear Algebra and its Applications, 416(2–3): 568–579, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2009:DSA

[YLZ09] Zhenghong Yang, Cheng Li, and Zhengzhi Zhang. Displacement structure approach to q-adic Chebyshev– Vandermonde-like matrices. Linear Algebra and its Applications, 430(11–12):2906– 2921, June 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2007:GPF

[YM07] Yongzhi Yang and Catherine Micek. Generalized Pascal functional matrix and its applications. Linear Algebra and its Applications, 423 (2–3):230–245, June 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yakhlef:2001:PNM

[YMP01] Hossain O. Yakhlef, Francisco Marcellán, and Miguel A. Piñar. Perturbations in the Nevai matrix class of orthogonal matrix polynomials. Linear Algebra and its Applications, 336(1–3): 231–254, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/166/27/
                                        43/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/166/27/43/article.
                                        pdf.

Yong:2000:PCF

[Yon00] Xuerong Yong. Proof of a conjecture of Fiedler and Markham. Linear Algebra and its Applications, 320(1– 3):167–171, November 15, 2000. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
35/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/35/article.
pdf.

Yoshino:2007:MHI

[Yos07] Takashi Yoshino. A modified Heinz’s inequality. Linear Algebra and its Applications, 420(2–3):686–699, January 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yousefi:2007:AR

[You07] Hassan Yousefi. Asymptotic reflexivity. Linear Algebra and its Applications, 422(2– 3):604–615, April 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Peng:2003:RAR

[yPyH03] Zhen yun Peng and Xi yan Hu. The reflexive and antireflexive solutions of the matrix equation AX = B. Linear Algebra and its Applications, 375(1):147–155, December 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Peng:2003:CJM

[yPyHZ03] Zhen yun Peng, Xi yan Hu, and Lei Zhang. On the construction of a Jacobi matrix from its mixed-type eigenpairs. Linear Algebra and its Applications, 362(C): 191–200, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/211/25/
40/abstract.html.

Yuan:2008:MSR

[YSL08] Xi-Ying Yuan, Jia-Yu Shao, and Yue Liu. The minimal spectral radius of graphs of order n with diameter n − 4. Linear Algebra and its Applications, 428(11–12):2840– 2851, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

You:2007:BBI

[YSS07] Lihua You, Jiayu Shao, and Haiying Shan. Bounds on the bases of irreducible generalized sign pattern matrices. Linear Algebra and its Applications, 427(2–3):285–300, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shao:2006:SLE

[ySZyY06] Jia yu Shao, Li Zhang, and Xi ying Yuan. On the second Laplacian eigenvalues of trees of odd order. Linear Algebra and its Applications, 419(2–3):475–485, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

You:2004:APR

[YT04] Hong You and Xiao Min Tang. Additive preservers of rank-additivity on the spaces of symmetric and alternate matrices. Linear Algebra and its Applications, 380(1):185– 198, March 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yanai:2006:NDR

[YTI06] Haruo Yanai, Yoshio Takane, and Hidetoki Ishii. Nonnegative determinant of a rectangular matrix: Its definition and applications to multivariate analysis. Linear Algebra and its Applications, 417 (1):259–274, August 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yuan:2000:ORT

[Yua00] Jin-Yun Yuan. The Ostrowski-Reich theorem for SOR iterations: extensions to the rank deficient case. Linear Algebra and its Applications, 315(1– 3):189–196, August 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/132/25/
36/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/132/25/36/article.
                                        pdf.

You:2007:PPM

[YW07] Hong You and Zhongying Wang. k-Potence preserving maps without the linearity and surjectivity assumptions. Linear Algebra and its Applications, 426(1):238– 254, October 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2000:GCCb

[YWS00] Guang-Hong Yang, Jian Liang Wang, and Yeng Chai Soh. Guaranteed cost control for discrete-time linear systems under controller gain perturbations. Linear Algebra and its Applications, 312(1– 3):161–180, June 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/129/25/
                                        35/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/129/25/35/article.
                                        pdf.

Yan:2003:TDR

[YY03] Zhibin Yan and Hong You. Two-dimensional representations of the free group with two generators over a finite field. Linear Algebra and its Applications, 359(1– 3):29–36, January 15, 2003.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/208/25/
28/abstract.html.

Yan:2005:ZCR

[YYZ05] Guo-Guang Yan, Arthur L. B. Yang, and Joan J. Zhou. The zrank conjecture and restricted Cauchy matrices. Linear Algebra and its Applications, 411(1):371– 385, December 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2007:SFS

[YZ07] Xiande Yang and Yiqiang Zhou. Some families of strongly clean rings. Linear Algebra and its Applications, 425(1):119–129, August 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2009:MPI

[YZ09] Li Yang and Ling Zhang. Maps on B(H) preserving involution. Linear Algebra and its Applications, 431(5– 7):666–672, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2006:SCB

[yZtLC06] Cheng yi Zhang, Yao tang Li, and Feng Chen. On Schur complement of block diagonally dominant matrices. Linear Algebra and its Applications, 414(2–3):533– 546, April 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2007:EDS

[yZXtL07] Cheng yi Zhang, Chengxian Xu, and Yao tang Li. The eigenvalue distribution on Schur complements of Hmatrices. Linear Algebra and its Applications, 422(1):250– 264, April 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Dokovic:2004:TNF

[– DZ04] Dragomir Ž. – Doković and Kaiming Zhao. Tridiagonal normal forms for orthogonal similarity classes of symmetric matrices. Linear Algebra and its Applications, 384(1): 77–84, June 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zaimi:2004:IFA

[Zaï04] Toufik Zaïmi. On the integer form of the additive Hilbert’s Theorem 90. Linear Algebra and its Applications, 390 (1):175–181, October 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zalar:2003:SNM

[Zal03] Bojana Zalar. On semigroups of normal matrices. Linear Algebra and its Applications, 368(C):191–195, July 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/217/25/
37/abstract.html.

Zarikian:2006:APA

[Zar06] Vrej Zarikian. Alternatingprojection algorithms for operator-theoretic calculations. Linear Algebra and its Applications, 419(2–3):710– 734, December 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zavadskij:2003:TEP

[Zav03] Alexander Zavadskij. Tame equipped posets. Linear Algebra and its Applications, 365(C):389–465, May 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/214/25/
37/abstract.html.

Zavadskij:2007:KPR

[Zav07] Alexander G. Zavadskij. On the Kronecker Problem and related problems of Linear Algebra. Linear Algebra and its Applications, 425(1):26– 62, August 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zavadskij:2008:MPC

[Zav08] A. G. Zavadskij. A matrix problem over a central quadratic skew field extension. Linear Algebra and its Applications, 428(1):393– 399, January 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zheng:2009:SCP

[ZBY09] Bing Zheng, Zhong-Zhi Bai, and Xi Yang. On semiconvergence of parameterized Uzawa methods for singular saddle point problems. Linear Algebra and its Applications, 431(5–7):808– 817, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhou:2005:ESM

[ZD05] Bin Zhou and Guang-Ren Duan. An explicit solution to the matrix equation AX − XF = BY. Linear Algebra and its Applications, 402(1): 345–366, June 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zemke:2006:HEE

[Zem06] Jens-Peter M. Zemke. Hessenberg eigenvalue–eigenmatrix relations. Linear Algebra and its Applications, 414(2–3):

589–606, April 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zemke:2007:APK

[Zem07] Jens-Peter M. Zemke. Abstract perturbed Krylov methods. Linear Algebra and its Applications, 424(2–3): 405–434, July 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zeng:2001:SAS

[Zen01] Lisheng Zeng. Some applications of spectral theory of nonnegative matrices to input-output models. Linear Algebra and its Applications, 336(1–3): 205–218, October 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/166/27/
41/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/166/27/41/article.
pdf.

Zeng:2008:GST

[Zen08] Guangxing Zeng. A generalization of Stickelberger’s theorem. Linear Algebra and its Applications, 428(11–12): 2880–2887, June 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2007:NPS

[ZF07] Guang Zhang and Wenying Feng. On the number of positive solutions of a nonlinear algebraic system. Linear Algebra and its Applications, 422(2–3):404– 421, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2006:GBN

[ZFLW06] Jian-Hua Zhang, Shan Feng, Hong-Xia Li, and Rui-Hua Wu. Generalized biderivations of nest algebras. Linear Algebra and its Applications, 418(1):225–233, October 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2006:LCN

[ZFW06] Jian-Hua Zhang, Shan Feng, and Rui-Hua Wu. Local 2cocycles of nest subalgebras of factor von Neumann algebras. Linear Algebra and its Applications, 416(2–3):908– 916, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhou:2009:FCS

[ZG09] Yihui Zhou and M. Seetharama Gowda. On the finiteness of the cone spectrum of certain linear transformations on Euclidean Jordan algebras. Linear Algebra and its Applications, 431(5–7):772– 782, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2006:LPP

[ZgC06] Xian Zhang and Chong guang Cao. Linear kpower/k-potent preservers between matrix spaces. Linear Algebra and its Applications, 412(2–3):373–379, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2009:CAL

[ZGL09] Shangli Zhang, Wenhao Gui, and Gang Liu. Characterization of admissible linear estimators in the general growth curve model with respect to an incomplete ellipsoidal restriction. Linear Algebra and its Applications, 431(1– 2):120–131, July 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhao:2002:BME

[ZH02] Shuqin Zhao and Yuan Hong. On the bounds of maximal entries in the principal eigenvector of symmetric nonnegative matrix. Linear Algebra and its Applications, 340(1– 3):245–252, January 1, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.com/
                                        gej-ng/10/30/19/172/27/
                                        44/abstract.html.

Zhang:2000:SCM

[Zha00] Fuzhen Zhang. Schur complements and matrix inequalities in the Löwner ordering. Linear Algebra and its Applications, 321(1–3): 399–410, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/140/24/
                                        50/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/140/24/50/article.
                                        pdf.

Zhang:2001:CDI

[Zha01] Liping Zhang. A characterization of the Drazin inverse. Linear Algebra and its Applications, 335(1–3): 183–188, September 15, 2001. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/165/27/
                                        38/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/30/19/165/27/38/article.
                                        pdf.

Zhang:2003:LOP

[Zha03a] Xian Zhang. Linear operators that preserve pairs of matrices which satisfy extreme rank properties — a supplementary version. Linear Algebra and its Applications, 375(1):283–290, December 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2003:HPD

[Zha03b] Yuhai Zhang. On Hermitian positive definite solutions of matrix equation X + A^{∗}X^{−2}A = I. Linear Algebra and its Applications, 372 (1):295–304, October 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhan:2004:SMI

[Zha04a] Xingzhi Zhan. On some matrix inequalities. Linear Algebra and its Applications, 376 (1):299–303, January 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2004:IPM

[Zha04b] Xian Zhang. Idempotencepreserving maps without the linearity and surjectivity assumptions. Linear Algebra and its Applications, 387(1): 167–182, August 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2004:LAP

[Zha04c] Xian Zhang. Linear/additive preservers of ranks 2 and 4 on alternate matrix spaces over fields. Linear Algebra and its Applications, 392(1):25– 38, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2004:TCG

[Zha04d] Xiao-Dong Zhang. On the two conjectures of Graffiti. Linear Algebra and its Applications, 385(1):369– 379, July 1, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2004:SER

[Zha04e]  Xiao-Dong Zhang.  The

smallest eigenvalue for reversible Markov chains. Linear Algebra and its Applications, 383(1):175–186, May 15, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2004:TSU

[Zha04f] Xiao-Dong Zhang. Two sharp upper bounds for the Laplacian eigenvalues. Linear Algebra and its Applications, 376 (1):207–213, January 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2005:APR

[Zha05a] Xian Zhang. Additive preservers of rank on alternate matrix spaces over fields and applications. Linear Algebra and its Applications, 397(1):

325–343, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2005:LAP

[Zha05b] Xian Zhang. Linear/additive preservers of rank 2 on spaces of alternate matrices over fields. Linear Algebra and its Applications, 396(1):91– 102, February 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2005:EEN

[Zha05c] Xiao-Dong Zhang. Eigenvectors and eigenvalues of nonregular graphs. Linear Algebra and its Applications, 409 (1):79–86, November 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2005:NDS

[Zha05d] Xiao-Dong Zhang. A note on doubly stochastic graph matrices. Linear Algebra and its Applications, 407(1):196–200, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhan:2006:CPI

[Zha06] Xingzhi Zhan. Completion of a partial integral matrix to a unimodular matrix. Linear Algebra and its Applications, 414(1):373– 377, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhan:2007:ENP

[Zha07a] Xingzhi Zhan. Extremal numbers of positive entries of imprimitive nonnegative matrices. Linear Algebra and its Applications, 424(1):132– 138, July 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2007:GTT

[Zha07b] Fuzhen Zhang. Geršgorin type theorems for quaternionic matrices. Linear Algebra and its Applications, 424(1):139–153, July 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2007:OTA

[Zha07c] Xiao-Dong Zhang. Ordering trees with algebraic connectivity and diameter. Linear Algebra and its Applications, 427(2–3):301–312, December 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2009:ACD

[Zha09] Xiao-Dong Zhang. Algebraic connectivity and doubly stochastic tree matrices. Linear Algebra and its Applications, 430(5–6):1656– 1664, March 1, 2009. CO-

Zheng:2002:DMC

[Zhe02] Baodong Zheng. Decomposition of matrices into commutators of involutions. Linear Algebra and its Applications, 347(1–3):1–7, May 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
27/abstract.html.

Zhou:2000:LAA

[Zho00] Jianhua Zhou. Lie algebras with almost dimensionally nilpotent inner derivations. Linear Algebra and its Applications, 320(1–3): 97–114, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
31/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/31/article.
pdf.

Zhou:2003:RCO

[Zho03] Fei Zhou. A rank criterion for the order of a pole of a matrix function. Linear Algebra and its Applications, 362 (C):287–292, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/211/25/
                                        48/abstract.html.

Zhou:2008:SPL

[Zho08] Bo Zhou. On sum of powers of the Laplacian eigenvalues of graphs. Linear Algebra and its Applications, 429(8–9): 2239–2246, October 16, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhu:2007:ADP

[Zhu07] Jun Zhu. All-derivable points of operator algebras. Linear Algebra and its Applications, 427(1):1–5, November 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhu:2008:QTF

[Zhu08] Xinyun Zhu. The N = 1 quivers of types A_{n} and D_{n} have finite representation type. Linear Algebra and its Applications, 428(4):919– 929, February 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhou:2003:SCI

[ZHZ03] Fu-Zhao Zhou, Xi-Yan Hu, and Lei Zhang. The solvability conditions for the inverse eigenvalue problems of centro-symmetric matrices. Linear Algebra and its Applications, 364(C):147– 160, May 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/213/25/
32/abstract.html.

Zhao:2008:LSS

[ZHZ08] Lijun Zhao, Xiyan Hu, and Lei Zhang. Least squares solutions to AX = B for bisymmetric matrices under a central principal submatrix constraint and the optimal approximation. Linear Algebra and its Applications, 428(4): 871–880, February 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zippin:2003:CES

[Zip03] M. Zippin. A characterization of the Euclidean space. Linear Algebra and its Applications, 361(C):3–6, March 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/210/25/
27/abstract.html.

Zitan:2009:TAR

[Zit09] Fouad Zitan. Train algebras of rank 3 with finiteness conditions. Linear Algebra and its Applications, 431(5–7): 1081–1087, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zivkovic:2006:CSM

[Živ06] Miodrag Živković. Classification of small (0, 1) matrices. Linear Algebra and its Applications, 414(1):310– 346, April 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zizler:2004:SRM

[Ziz04] Peter Zizler. The spectral radius of a multivariate sampling operator. Linear Algebra and its Applications, 385(1):463–474, July 1, 2004. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2004:LDN

[ZJC04] Jian-Hua Zhang, Guo-Xing Ji, and Huai-Xin Cao. Local derivations of nest subalgebras of von Neumann algebras. Linear Algebra and its Applications, 392(1):61– 69, November 15, 2004. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2002:AGB

[ZJL02] Shenggui Zhang, Zhaolin Jiang, and Sanyang Liu. An application of the Gröbner basis in computation for the minimal polynomials and inverses of block circulant matrices. Linear Algebra and its Applications, 347(1– 3):101–114, May 15, 2002.

CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/194/27/
33/abstract.html.

Zhang:2002:LSM

[ZL02] Xiao-Dong Zhang and Jiong-Sheng Li. The Laplacian spectrum of a mixed graph. Linear Algebra and its Applications, 353(1–3):11–20, September 15, 2002. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2003:LEM

[ZL03] Xiao-Dong Zhang and Rong Luo. The Laplacian eigenvalues of mixed graphs. Linear Algebra and its Applications, 362(C):109–119, March 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/211/25/
33/abstract.html.

Zhang:2009:OTL

[ZLC09] Xiao-Dong Zhang, Xia-Ping Lv, and Ya-Hong Chen. Ordering trees by the Laplacian coefficients. Linear Algebra and its Applications, 431(12):2414–2424, December 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhai:2009:SRB

[ZLS09] Mingqing Zhai, Ruifang Liu, and Jinlong Shu. On the spectral radius of bipartite graphs with given diameter. Linear Algebra and its Applications, 430(4):1165–1170, February 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zaslavsky:2003:CJC

[ZM03] Boris G. Zaslavsky and Judith J. McDonald. A characterization of Jordan canonical forms which are similar to eventually nonnegative matrices with the properties of nonnegative matrices. Linear Algebra and its Applications, 372(1):253– 285, October 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2008:DVB

[ZM08] Yin Zhang and Michael Merritt. Dose-volume-based IMRT fluence optimization: A fast least-squares approach with differentiability. Linear Algebra and its Applications, 428(5–6):1365–1387, March 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zobin:2000:SRQ

[Zob00] Nahum Zobin. Some remarks on quasi-equivalence of bases in Fréchet spaces. Linear Algebra and its Applications, 307(1–3):47–67, March 1, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/124/25/
28/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/124/25/28/article.
pdf.

Zavorin:2003:CSG

[ZOE03] Ilya Zavorin, Dianne P. O’Leary, and Howard Elman. Complete stagnation of GMRES. Linear Algebra and its Applications, 367 (C):165–183, July 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/216/25/
37/abstract.html.

Zhang:2006:LDC

[ZPY06] Jian-Hua Zhang, Fang-Fang Pan, and Ai-Li Yang. Local derivations on certain CSL algebras. Linear Algebra and its Applications, 413(1):93– 99, February 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2007:AMB

[ZQ07] Yanxia Zhang and Li Qiu. On the angular metrics between linear subspaces. Linear Algebra and its Applications, 421(1):163–170, February 1, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zheng:2005:TPP

[ZS05] Baodong Zheng and Yuqiu Sheng. On transformation partly preserving a nonsingular bilinear function. Linear Algebra and its Applications, 403(1):273–284, July 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhai:2009:MLS

[ZSL09] Mingqing Zhai, Jinlong Shu, and Zhonghua Lu. Maximizing the Laplacian spectral radii of graphs with given diameter. Linear Algebra and its Applications, 430(8– 9):1897–1905, April 15, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhu:2009:GCP

[ZTK09] Jiandong Zhu, Yu-Ping Tian, and Jing Kuang. On the general consensus protocol of multi-agent systems with double-integrator dynamics. Linear Algebra and its Applications, 431(5–7):701–715, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zusmanovich:2005:LDC

[Zus05] Pasha Zusmanovich. Lowdimensional cohomology of current Lie algebras and analogs of the Riemann tensor for loop manifolds. Linear Algebra and its Applications, 407(1):71–104, September 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhao:2000:APG

[ZW00] Xiqiang Zhao and Tianming Wang. The algebraic properties of the generalized Pascal functional matrices associated with the exponential families. Linear Algebra and its Applications, 318(1– 3):45–52, October 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/137/24/
27/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/137/24/27/article.
pdf.

Zhang:2003:OLR

[ZW03] Zhenyue Zhang and Lixin Wu. Optimal low-rank approximation to a correlation matrix. Linear Algebra and its Applications, 364 (C):161–187, May 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/213/25/
                                        31/abstract.html.

Zhang:2006:LTD

[ZWC06] Jian-Hua Zhang, Bao-Wei Wu, and Huai-Xin Cao. Lie triple derivations of nest algebras. Linear Algebra and its Applications, 416(2–3):559– 567, July 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zaris:2006:IZL

[ZWPR06] P. Zaris, J. Wood, H. Pillai, and E. Rogers. On invariant zeros of linear systems of PDEs. Linear Algebra and its Applications, 417(1):275– 297, August 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhai:2009:MSR

[ZWS09] Mingqing Zhai, Yarong Wu, and Jinlong Shu. Maximizing the spectral radius of bicyclic graphs with fixed girth. Linear Algebra and its Applications, 431(5–7):716– 723, August 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zou:2000:LET

[ZWY00] Yong Zou, Zheng Wang, and Xuerong Yong. On the largest kth eigenvalues of trees with n ≡ 0 (mod k). Linear Algebra and its Applications, 320(1–3): 173–182, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/30/19/139/24/
36/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/139/24/36/article.
pdf.

Zhu:2005:CGJ

[ZX05a] Jun Zhu and Changping Xiong. Characterization of generalized Jordan ∗-left derivations on real nest algebras. Linear Algebra and its Applications, 404(1):325– 344, July 15, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhu:2005:GDM

[ZX05b] Jun Zhu and Changping Xiong. Generalized derivable mappings at zero point on some reflexive operator algebras. Linear Algebra and its Applications, 397(1):367– 379, March 1, 2005. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhu:2007:DMU

[ZX07] Jun Zhu and Changping Xiong. Derivable mappings at unit operator on nest algebras. Linear Algebra and its Applications, 422(2–3):721– 735, April 15, 2007. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhu:2008:ADP

[ZXZ08] Jun Zhu, Changping Xiong, and Renyuan Zhang. Allderivable points in the algebra of all upper triangular matrices. Linear Algebra and its Applications, 429(4):804– 818, August 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhu:2009:ADP

[ZXZ09] Jun Zhu, Changping Xiong, and Lin Zhang. All-derivable points in matrix algebras. Linear Algebra and its Applications, 430(8–9):2070– 2079, April 15, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zheng:2002:PCT

[ZY02] Baodong Zheng and Hong You. Products of commutators of transvections over local rings. Linear Algebra and its Applications, 357(1– 3):45–57, December 15, 2002. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/30/19/203/25/
                                        29/abstract.html.

Zhang:2006:JDT

[ZY06] Jian-Hua Zhang and Wei-Yan Yu. Jordan derivations of triangular algebras. Linear Algebra and its Applications, 419(1):251–255, November 1, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yan:2008:UVC

[zY08] Zi zong Yan. A unified version of Cauchy–Schwarz and Wielandt inequalities. Linear Algebra and its Applications, 428(8–9):2079–2084, April 15, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2003:RCH

[ZyC03] Xian Zhang and Mei yu Cheng. The rank-constrained Hermitian nonnegative-definite and positive-definite solutions to the matrix equation AXA^{∗} = B. Linear Algebra and its Applications, 370(1):163–174, September 1, 2003. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zygmunt:2002:MCP

[Zyg02] Marcin J. Zygmunt. Matrix Chebyshev polynomials and continued fractions. Linear Algebra and its Applications, 340(1–3):155–168, January 1, 2002. CODEN LAAPAW.

ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.com/
gej-ng/10/30/19/172/27/
38/abstract.html.

Zhang:2005:LAN

[ZYP05] Jian-Hua Zhang, Ai-Li Yang, and Fang-Fang Pan. Local automorphisms of nest subalgebras of factor von Neumann algebras. Linear Algebra and its Applications, 402 (1):335–344, June 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2006:LMP

[ZYP06] Jian-Hua Zhang, An-Li Yang, and Fang-Fang Pan. Linear maps preserving zero products on nest subalgebras of von Neumann algebras. Linear Algebra and its Applications, 412(2–3):348–361, January 15, 2006. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2000:GMW

[ZZ00] Bao-Xue Zhang and Xian-Hai Zhu. Gauss–Markov and weighted least-squares estimation under a general growth curve model. Linear Algebra and its Applications, 321(1–3):387–398, December 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL

http://www.elsevier.nl/

gej-ng/10/30/19/140/24/
49/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/30/19/140/24/49/article.
pdf.

Zhang:2005:CQG

[ZZ05] Xiaoxia Zhang and Ervin Yunwei Zhao. The compact quantum group U_{q}(2) I. Linear Algebra and its Applications, 408(1):244–258, October 1, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2008:NMP

[ZZ08a] Jian-Hua Zhang and Fang-Juan Zhang. Nonlinear maps preserving Lie products on factor von Neumann algebras. Linear Algebra and its Applications, 429(1):18– 30, July 1, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2008:LSR

[ZZ08b] Xiaoling Zhang and Heping Zhang. The Laplacian spectral radius of some bipartite graphs. Linear Algebra and its Applications, 428 (7):1610–1619, April 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2009:SGD

[ZZ09] Xiaoling Zhang and Heping Zhang. Some graphs determined by their spectra. Linear Algebra and its Applications, 431(9):1443–1454, October 1, 2009. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zhang:2008:DSO

[ZZJ08] Shifang Zhang, Huaijie Zhong, and Qiaofen Jiang. Drazin spectrum of operator matrices on the Banach space. Linear Algebra and its Applications, 429(8–9):2067–2075, October 16, 2008. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Zha:2006:QCQ

[ZZY06] Hongyuan Zha, Zhenyue Zhang, and Wenlong Ying. A quadratically convergent QRlike method without shifts for the Hermitian eigenvalue problem. Linear Algebra and its Applications, 417(2–3): 478–495, September 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).